A Complete Bibliography of Publications in the British Journal for the History of Science: 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

21 October 2024 Version 1.20

Title word cross-reference

$10.00 [Cro96c]. $100.00 [Lar96a, Cam90]. $107.00 [Ste95]. $109.00 [Bro96b]. $112.00 [Nor94a]. $115.00 [Bow94b, Bra92a]. $117.00 [McC98]. $12.95 [Edg91a, Jor91, Spa91]. $120.00 [Sch91a, Mor95c, Sch91a]. $13.95 [Jon93]. $135.00 [Bro99c]. $139.00 [Non93a]. $14.95 [Boh95, Dan92, Law96a, Mor92a, Smi99a, Cuo95, Ost94, Spa95]. $140 [Aly90]. $140.00 [Hen98b]. $149.00 [Hac93c]. $15.00 [Di 95a, Hun92]. $15.95 [Daw98a, Roh99a]. $150.00 [Hod96]. $16.00 [Gol91b, Smi90a]. $16.50 [Gre92c]. $16.95 [Bis98, Kus95c, Mol96]. $17.00 [Hen92b]. $17.95 [Hem95, Kus97, Mor98, Sho99, Smi99b].

$170.00 [Hel96b]. $18.25 [Bow92c]. $18.95 [Blo93a, Mic98, Ost98]. $189.00 [Cas94]. $19.00 [Bro93a]. $19.50 [Mur96, War96a]. $19.95 [Mur92a, Nig99b, Nut93, Ste91a, Ver94, Boh93, Gol94]. $195 [Buc92]. $195.00 [Smi92b]. $20.00 [Hil91c]. $20.75 [Bow92c]. $21.95 [Sha97, Hac90e]. $22.50 [Nig99a, Edg90b, Por91a]. $22.95 [Lyn99, Poo99, Pyl93a, Yeo96b]. $23.95 [Gro93]. $24.00 [Chu91]. $24.95 [Boh95, Har96c, Ben93, Kus95e, Law91a, Wis96]. $25.00 [Aga95a, Bev90]. $250.00 [War96b]. $26.95 [Hen98a]. $27.50 [Ras96a, Aga96a, Bro94b, McK97b]. $27.95 [Sho99, Dea93, Wit95a]. $28.75 [Can90b]. $280.00 [HD97]. $29.00 [Hil93]. $29.95 [Bug96, Slo97, Spa91, Cro95, Fox92, GG91c, Har98b, Hir95, Ili95b, Law96c, Olb91a, Pal96a, Roh99b, Uim95, War91b]. $30.00 [Har92c]. $31.50 [Kra93b]. $32.00 [Bro93a]. $32.50 [Edg91a, Ver95]. $32.95 [Goo97]. $33.00 [Spa98a]. $34.50 [Ari90, Har92a, Kra92, Pum90]. $34.95 [Bru96, Aga94a, Cri90, Hor99b, Pau90]. $35 [Nea98]. $35.00 [Mor92a, BV91, Gre96, Le 91, O’B96a, Rav99, Smi92d, Smi93c]. $36.00 [Mol98b]. $36.25 [Sut97]. $36.50 [Ost99]. $37.00 [Hac95]. $37.95 [Bow93c]. $38 [Nic96a]. $38.00 [Hug92c, Kni98b]. $38.45 [Har93a, Kra90b]. $39.00 [Aga95a]. $39.50 [Edg91a, Jor91, Smi99a, Bar93b, Bra95, Bro91a, Coo96, Gou98b, Kra90a, Mag96a, Woo98]. $39.95 [Bis98, Dea92, Fie94, Luc96b, Smi90a, Yav92]. $40 [Woo95, Por90a]. $40.00 [Mol94a, Pal97, Tei93]. $42.50 [Hil91a, Hoc92a, Hor96b, Por90b]. $42.95 [Hem95, Mal93b]. $43.00 [War95b]. $43.95 [Fie94, Luc93]. $44.00 [Nye93b]. $44.50 [Dan92, Ger91, Ink90, Kra91a, Pal91b]. $44.95 [Law96a, Bow93a, Het95, Hoc93a]. $45.00 [Bro99c, Gre96, Hen92b, Ger95, Hoc93b, Kni90c, Pum92]. $45.95 [Cro96b]. $46.00 [Hug92a]. $47.50 [Bro91b, Kri91]. $47.95 [Man93b]. $48.00 [Kni95]. $49 [Jac90]. $49.50 [Bro99c, Hal92d, Bra98, Coo90, Fre93, Hal91a, Hen99, Hil91b, Kre96, Mar93a, Maz96, Ras96b, Smi91a, Smi96, Wat91, Wil91b]. $49.95 [Aga96a, Bro99c, Hun92, Smi99b, Bro93b, Bro98, Hes95b, Kus98, Mal95a, McE95, Sha93a, She99]. $50.00 [Law92e]. $51.50 [Mol95]. $51.75 [Mur96, Law93f]. $52.00 [Mor98]. $54.00 [Hal92d, Div98]. $54.50 [She92]. $54.90 [Kus96b]. $54.95 [Mic98, Nig99b, Ost98, Dol94, Hen95a, Law93c, Pyl93b, Sha97, Tim98, Wei94]. $55.00 [Lar96a, Hac92b]. $57.00 [Nut93, Old97]. $57.50 [Law93b]. $59.50 [Blo93a, Ste91a, Bar92, Bos92, Bro90e, But92a, Cha90c, Kra91b, Por91c]. $59.95 [Aga97c, Car95, Dol98b, Har98h, Hor97a, Jor96, Mar96b, Ost93c, Ost93d, Pic98, Pri93, Tur93c, Yos99]. $599 [Hug96]. $60.00 [Gro93, Bai91]. $64.95 [Har96c, Poo99, Aga97b, Can96, Har96d, Hig94, Ili98, Jac96]. $65 [Mur92a]. $65.00 [Lar96a, Bro90a, Jam93a, Lar96a]. $66.75 [Nic93a]. $67.00 [Hen93a]. $69.00 [Jam92c, Whi90a]. $69.50 [Smi93b]. $69.95 [Gol93, Hen96, Mal94, Mon93, Wil97]. $7.50 [Gre92c]. $70.00 [Nig99a, Har95a, Joh94]. $72.00 [Lyn99]. $732.00 [Lar96a]. $74.95 [Hor96a, Hor97b, Mil95]. $75.00 [Bug96, Goo97, Hen98a, Lar96a, Ras96a, Far99]. $77.00 [Hug92c]. $78.00 [Ams92]. $79.95 [Nig99c]. $8.95 [Edg91a]. $84.00 [Lar96a]. $85.75 [Osl97]. $86.00 [Mil94]. $86.25 [Can90b]. $89.00 [Smi94d, Elm97]. $89.95 [Bru96, Slo97]. $9.25 [Ari90]. $9.95 [Bro93d, Non91b]. $92.00 [Hun98]. $93.00 [Lar96a]. $95.00 [Mul91, Fra90, Hod93b]. $96.00 [Rag95]. $96.25 [War91a]. $99.00 [Hod93a, Div90]. $99.50 [Nic93c].

. [Goo91b].

0 [Cra92c, Di 93b, Di 93d, Gre91, Kni92a, Law91b, Law92a, Law93d, Mor91c, Mor93, Nye92b, Olb91b, Osb90a, Smi92a, Tur91]. 0-00-217956-3 [Bru91]. 0-00-686173-3 [McE93]. 0-00-686176-8 [Gre95]. 0-00-686177-6 [Wil95a]. 0-00-686178-4 [Blo98]. 0-00-686184-9 [Nic94b]. 0-04-440522-7 [Bra91]. 0-04-445919-X [Mur92b]. 0-040551138-1 [Tha91]. 0-0818-4379-0 [Nye93b]. 0-11-290060-7 [But94]. 0-11-515302-0 [Col93]. 0-12-515175-6 [Aga97b]. 0-12-587445-6 [Cri90]. 0-13-217555-X [Hac92a]. 0-14-043528-X [MK98]. 0-19-215361-7 [Ben90a]. 0-19-261198-4 [Law93g]. 0-19-261651-X [Por91b]. 0-19-261747-8 [Pic92b]. 0-19-286162-X [Twe93c]. 0-19-289296-7 [Stu98]. 0-19-505203X [Kni90a]. 0-19-506040-7 [Pyl93a]. 0-19-506380-5 [Jon93]. 0-19-506447-X [Law92a]. 0-19-563562-0 [Har97]. 0-19-721795-8 [Elm91]. 0-19-812983-1 [Ley90]. 0-19-818369-0 [Law98b]. 0-19-822991-7 [Lar90]. 0-19-824093-7 [Mol98a]. 0-19-824240-9 [Yeo92]. 0-19-824250-6 [Cla92c]. 0-19-824485-1 [Whi90c]. 0-19-824812-1 [Ait90]. 0-19-824876-8 [Hen93b]. 0-19-824900-4 [McN92]. 0-19-824979-9 [Ree90]. 0-19-824998-5 [Str90]. 0-19-852394-7 [Rav99]. 0-19-853918-5 [Wil91c]. 0-19-853945-2 [Hal94a]. 0-19-853985-1 [Mal94]. 0-19-854053-1 [Sec98]. 0-19-854684-X [Har93c]. 0-19-857725-7 [Smi92c]. 0-19-858180-7 [Mor91c]. 0-224-04202-5 [La 97]. 0-226-03713-4 [Law93b]. 0-226-04559-5 [Ili95b]. 0-226-07884-1 [Can90b]. 0-226-07886-8 [Can90b]. 0-226-07887-6 [Goo97]. 0-226-07888-4 [Goo97]. 0-226-10073-1 [Yeo96b]. 0-226-11279-9 [Hen98a]. 0-226-11280-2 [Hen98a]. 0-226-12146-1 [Mor92a]. 0-226-12148-8 [Mor92a]. 0-226-13476-8 [Yeo96a]. 0-226-14346-5 [Hod92a]. 0-226-27903-0 [Ari90]. 0-226-28217-1 [Gro93]. 0-226-28219-8 [Gro93]. 0-226-28545-6 [Aga95a]. 0-226-28546-4 [Aga95a]. 0-226-29205-2 [Dol90a]. 0-226-31633-5 [Cam90]. 0-226-32658-6 [Mor98]. 0-226-32659-4 [Mor98]. 0-226-35551-0 [Jor94]. 0-226-35551-9 [Jor94]. 0-226-48111-5 [Nig99a]. 0-226-48112-3 [Nig99a]. 0-226-48230-8 [Nut93]. 0-226-48231-6 [Nut93]. 0-226-49794-1 [Bow92c]. 0-226-49797-6 [Bow92c]. 0-226-49798-4 [Bow92c]. 0-226-49799-2 [Bow92c]. 0-226-52086-2 [Ost94]. 0-226-53236-4 [Nic93a]. 0-226-53237-2 [Nic93a]. 0-226-54234-3 [Pum95a]. 0-226-54235-1 [Pum95a]. 0-226-61086-1 [Ras96a]. 0-226-61088-8 [Ras96a]. 0-226-66952-1 [Har98b]. 0-226-71202-8 [Boh93]. 0-226-73104-9 [Law93f]. 0-226-76129-0 [O’B92]. 0-226-76130-4 [O’B92]. 0-226-84922-9 [Mur96]. 0-226-84991-0 [Mur96]. 0-226-87054-5 [Kra92]. 0-226-90215-3 [Di 93d]. 0-226-91065-2 [Bug96]. 0-226-91066-0 [Bug96]. 0-229-13630-2 [Rou97]. 0-229-13634-5 [Rou97]. 0-229-13820-8 [Mau95]. 0-229-13824-0 [Mau95]. 0-233-98681-2 [Bru93]. 0-253-20571-9 [Spa91]. 0-253-34266-X [Hef96]. 0-253-34287-2 [Di 93c]. 0-253-35405-6 [Le 91]. 0-253-35549-4 [Spa91]. 0-262-01121-2 [Bow92a]. 0-262-03137-X [Olb91b]. 0-262-03143-4 [Twe90]. 0-262-05040-4 [Gra91]. 0-262-05051-X [Aga98]. 0-262-07138-X [Kra93b]. 0-262-08086-9 [Smi93a]. 0-262-08187-3 [CK91]. 0-262-08198-9 [Nye92b]. 0-262-08200-4 [Twe92c]. 0-262-09028-7 [GG91b]. 0-262-11172-1 [Blo93c]. 0-262-12146-8 [Twe91c]. 0-262-13168-X [Twe91b]. 0-262-13232-X [Buc90]. 0-262-13259-1 [Hug92b]. 0-262-13278-8 [Twe93b]. 0-262-14046-2 [Twe92b]. 0-262-14048-9 [Luc92]. 0-262-15031-X [Hud92]. 0-262-15038-7 [Twe91a]. 0-262-16123-0 [Twe92a]. 0-262-19304-3 [Law93g]. 0-262-19342-6 [Joh96]. 0-262-19347-7 [Edg95]. 0-262-23145-X [For91]. 0-262-23167-0 [Aga95b]. 0-262-24037-8 [Hir95]. 0-262-53082-1 [Twe90]. 0-262-69167-1 [Edg95]. 0-268-00690-3 [Bro93d]. 0-268-00923-6 [Non91b]. 0-268-01029-3 [Sha97]. 0-268-01032-3 [Sha97]. 0-271-00629-3 [Por91a]. 0-292-71339-5 [Hil93]. 0-297-81514-8 [Rou97]. 0-299-13600-0 [Jon94]. 0-299-13604-3 [Jon94]. 0-299-14350-3 [Sho96a]. 0-299-14354-6 [Sho96a]. 0-300-04131-4 [Smi92d]. 0-300-04343-0 [Pal91a]. 0-300-04521-2 [Pal91a]. 0-300-05359-2 [Sha96]. 0-300-05588-9 [Hak95b]. 0-300-05705-9 [Jam93a]. 0-300-06382-2 [Har98d]. 0-313-23719-0 [Blu90]. 0-330-32013-0 [Jam94a]. 0-333-44159-1 [Bow91b]. 0-333-44419-1 [Law92c]. 0-333-46213-0 [Law90b]. 0-333-48331-6 [Goo90a]. 0-333-51059-3 [Sho93c]. 0-333-54291-6 [Mor92f]. 0-333-55077-3 [Jam92b]. 0-335-09317-5 [Stu91]. 0-335-09318-3 [Stu91]. 0-340-55861-X [Ost93b]. 0-374-14660-8 [Aga96a]. 0-404-61474-4 [Por90b]. 0-415-00037-8 [Edg90a]. 0-415-00625-2 [Mor90a]. 0-415-01988-5 [Hal92a]. 0-415-02063-8 [Bur98]. 0-415-03122-2 [Mor90a]. 0-415-03423-X [Law90c]. 0-415-03557-0 [Edg90a]. 0-415-03712-3 [Far94]. 0-415-03732-8 [McC95]. 0-415-04424-3 [Jon92]. 0-415-04771-4 [Hen95b]. 0-415-06042-7 [Hor98]. 0-415-11029-7 [O’B96b]. 0-415-11545-0 [Get96]. 0-415-18613-7 [Smi99b]. 0-415-18641-5 [Smi99b]. 0-436-09478-9 [Can90a]. 0-444-98749-5 [Hug92c]. 0-444-98753-3 [Hug92c]. 0-465-02989-2 [Aga98]. 0-465-02997-3 [McK97b]. 0-485-11390-2 [Rob93]. 0-485-11394-5 [Kus94]. 0-485-11433-X [Smi94a]. 0-485-24011-4 [Hay98a]. 0-485-24211 [Hay98a]. 0-486-26173-5 [Fie91]. 0-520-05859-3 [Por90b]. 0-520-06248-5 [Mor90d]. 0-520-06846-7 [Pum92]. 0-520-07022-4 [BV91]. 0-520-08210-9 [Kni95]. 0-521-23503-0 [Bro90e]. 0-521-23646-0 [Aly90]. 0-521-24254-1 [Wil91b]. 0-521-24775-6 [Sch91a]. 0-521-25625-9 [Buc92]. 0-521-26621-1 [Mur92a]. 0-521-26634-3 [Bro91a]. 0-521-29681-1 [HS90]. 0-521-30104-1 [Ink90]. 0-521-30142-4 [Pum90]. 0-521-30278-1 [Wil90a]. 0-521-30517-9 [Pic93a]. 0-521-30675-2 [Blo94b]. 0-521-30694-9 [Hal91a]. 0-521-30984-0 [Kra91b]. 0-521-31105-5 [Ore90]. 0-521-31977-3 [Mur92a]. 0-521-32109-3 [She92]. 0-521-32210-3 [Ore90]. 0-521-32361-4 [Sch91a]. 0-521-32467X [Fra90]. 0-521-32831-4 [Smi92b]. 0-521-32937-X [Bar93b]. 0-521-33115-3 [Hod91]. 0-521-33297-4 [Wei94]. 0-521-33330-X [Osb90a]. 0-521-33499-3 [Bra92a]. 0-521-33537-X [Smi91a]. 0-521-33818-2 [Bow91a]. 0-521-34023-3 [Olb91a]. 0-521-34045-4 [Pal91b]. 0-521-34108-6 [Har92b]. 0-521-34194-9 [Hil91b]. 0-521-34468-8 [Sha97]. 0-521-34804-8 [Pum94]. 0-521-34915-X [Har92b]. 0-521-35042-5 [Por91c]. 0-521-35077-8 [Coo90]. 0-521-35158-8 [Fie93]. 0-521-35168-5 [Wil97]. 0-521-35287-8 [Hen90a]. 0-521-35363-7 [Het95]. 0-521-35365-3 [Mar93a]. 0-521-35366-1 [Kri91]. 0-521-35434-X [Kra90a]. 0-521-35556-7 [Bro90a]. 0-521-35601-6 [Gol94]. 0-521-35602-4 [Ste91a]. 0-521-35685-7 [Ste91a]. 0-521-36113-3 [Lev94]. 0-521-36144-3 [Mon93]. 0-521-36231-8 [War96b]. 0-521-36261-X [Cha90c]. 0-521-36439-6 [Kra91a]. 0-521-36550-3 [Blo93a]. 0-521-36619-4 [Jor91]. 0-521-36698-4 [Jor91]. 0-521-36751-4 [Blo93a]. 0-521-37165-1 [Jor96]. 0-521-37213-5 [Smi93b]. 0-521-37327-1 [Bos92]. 0-521-37398-0 [Ger91]. 0-521-37419-7 [Bar92]. 0-521-37435-9 [Mal93a]. 0-521-38014-6 [Dan92]. 0-521-38045-6 [Wat91]. 0-521-38084-7 [Man93b]. 0-521-38089-8 [Hen91b]. 0-521-38135-5 [Bru96]. 0-521-38153-3 [Mar96b]. 0-521-38240-8 [Ver94]. 0-521-38249-1 [Sha93a]. 0-521-38373-0 [Sim92]. 0-521-38470-2 [Kra94]. 0-521-38884-8 [Dan92]. 0-521-39032-X [Mil95]. 0-521-39150-4 [But92a]. 0-521-39215-2 [Har92a]. 0-521-39323-X [Jac93]. 0-521-39414-7 [Law93c]. 0-521-39510-0 [Mar94b]. 0-521-39699-9 [Mar94b]. 0-521-40048-1 [Fre93]. 0-521-40049-X [Pri93]. 0-521-40099-6 [Aga93]. 0-521-40322-7 [Hor97a]. 0-521-40385-5 [Har96d]. 0-521-40386-3 [Hoc93a]. 0-521-40418-5 [Tur93c]. 0-521-40471-1 [Can96]. 0-521-40503-3 [Ost93c]. 0-521-40507-6 [Mal94]. 0-521-41095-9 [Ost93d]. 0-521-41109-2 [Bow93a]. 0-521-41373-7 [Mor95c]. 0-521-41398-2 [Car95]. 0-521-41408-3 [Aga94b]. 0-521-41522-5 [Law96a]. 0-521-41539-X [Wil94a]. 0-521-41573-X [Aga97c]. 0-521-41617-5 [Aga94a]. 0-521-41700-7 [Gol93]. 0-521-41775-9 [Bro93b]. 0-521-41933-6 [Hig94]. 0-521-42152-7 [Kus97]. 0-521-42534-4 [Law96a]. 0-521-43172-7 [Jac96]. 0-521-43182-4 [Dol94]. 0-521-43268-5 [Pic98]. 0-521-43323-1 [Law98c]. 0-521-43344-4 [Hen96]. 0-521-43423-8 [Hod96]. 0-521-43427-0 [Mal95a]. 0-521-43496-3 [Hen95a]. 0-521-44489-6 [Aga96a]. 0-521-45394-1 [Slo97]. 0-521-45518-9 [Hor96a, Hor97b]. 0-521-46108-1 [Hem95]. 0-521-46121-9 [Hug98]. 0-521-46649-0 [Hem95]. 0-521-46785-3 [Kus95c]. 0-521-47258-X [Mic98]. 0-521-47309-8 [Smi99b]. 0-521-47347-0 [Met96b]. 0-521-47540-6 [Kus96b]. 0-521-47564-3 [Bru96]. 0-521-47849-9 [Mic98]. 0-521-47884-7 [Kla95]. 0-521-47889-8 [Smi99b]. 0-521-48251-8 [Smi98]. 0-521-48303-4 [Dol98b, Dol99]. 0-521-48930-5 [Smi98]. 0-521-49741-8 [Har98h]. 0-521-55149-8 [Hay98b]. 0-521-55321-1 [Tim98]. 0-521-55325-3 [Nig99b]. 0-521-55699-6 [Nig99b]. 0-521-55894-8 [Slo97]. 0-521-56004-7 [Ful98]. 0-521-56100-0 [Ost98]. 0-521-56102-7 [Yos99]. 0-521-56130-2 [Bro99c]. 0-521-56224-4 [Nig99c]. 0-521-56413-1 [Poo99]. 0-521-56474-3 [Poo99]. 0-521-56796-3 [Ost98]. 0-521-58864-2 [Gau99]. 0-521-59196-1 [Ili98]. 0-565-01048-4 [Bro91c]. 0-565-01106-5 [Di 93f]. 0-565-01109-X [Di 93f]. 0-565-011448 [Di 93f].

0-567-08600-3 [She99]. 0-571-16182-0 [Odd91]. 0-631-15785-9 [Kni90b]. 0-631-16107-4 [Smi91c]. 0-631-16404-9 [Kus98]. 0-631-16492-8 [Swa92]. 0-631-16528-2 [Smi93d]. 0-631-16649-1 [Law92d]. 0-631-16816-8 [Jam94b]. 0-631-16818-4 [Bro92b]. 0-631-16849-4 [Hoc92d]. 0-631-16851-6 [Hoc92d]. 0-631-16977-6 [Lee90]. 0-631-16978-4 [Lee90]. 0-631-17295-5 [Hut92]. 0-631-17887-2 [Cro95]. 0-631-17906-2 [Ost93a]. 0-631-18535-6 [Cor93]. 0-631-18536-0 [Cor93]. 0-631-19384-7 [Har96c]. 0-631-19385-5 [Har96c]. 0-674-16296-X [Hod90a]. 0-674-17553-0 [Pal98]. 0-674-19274-0 [Whi90d]. 0-674-34171-6 [Man97]. 0-674-49660-4 [Kat90]. 0-674-60483-0 [Di 93b]. 0-674-62835-7 [Hoc90a]. 0-674-65760-8 [Ver90]. 0-674-74580-9 [Pau90]. 0-674-80290-X [Bev90]. 0-679-40104-0 [Bro94b]. 0-691-00099-9 [War96a]. 0-691-01136-2 [Nea98]. 0-691-01189-3 [Hen99]. 0-691-02142-2 [Roh99b]. 0-691-02997-0 [Gou98b]. 0-691-03204-1 [Fie94]. 0-691-03290-4 [Gre96]. 0-691-03367-6 [Bra98]. 0-691-03397-8 [Kre96]. 0-691-03418-4 [Coo96]. 0-691-03439-7 [Bra95]. 0-691-03442-7 [Pal96a]. 0-691-03717-5 [Law96c]. 0-691-03724-8 [Kus95e]. 0-691-03745-0 [Maz96]. 0-691-03759-0 [Ras96b]. 0-691-03776-0 [Wis96]. 0-691-03794-9 [Mag96a]. 0-691-04454-6 [Woo98]. 0-691-04804-5 [Uim95]. 0-691-05691-9 [Gre96]. 0-691-08465-3 [Har90b]. 0-691-08512-9 [Mor93]. 0-691-08533-1 [War91b]. 0-691-08543-9 [GG91c]. 0-691-08758-X [Smi96]. 0-691-08788-1 [O’B96a]. 0-691-10250-3 [War95b]. 0-7103-0428-5 [Hoc92c]. 0-7148-8038-8 [Hac90d]. 0-7181-3430-3 [Bro93c]. 0-7181-3641-1 [Goo96]. 0-7185-1357-6 [Ben94b]. 0-7190-1903-6 [Bar90a].

0-7190-2227-4 [Sta90b]. 0-7190-2886-8 [Pal90]. 0-7190-3025-0 [Har90a]. 0-7190-3302-0 [Nye92a]. 0-7201-2114-0 [Mar92a]. 0-7294-0374-2 [Cas93]. 0-7306-2494-3 [Bro95d]. 0-7450-0682-X [Kus95b]. 0-7484-0642-5 [Fyf99]. 0-7484-0696-4 [Bro99c]. 0-7484-0720-0 [Bro99c]. 0-7486-0359-X [Blo93b]. 0-7486-0521-5 [Sec96]. 0-7503-0042-6 [Law92b]. 0-7503-0147-3 [HD97]. 0-7503-0453-7 [Bro99c]. 0-7509-1334-7 [Lyn99]. 0-7509-1343-6 [Lyn99]. 0-7618-0708-X [Ost99]. 0-7735-0655-1 [For90a]. 0-7735-0689-6 [Bra90]. 0-7735-0791-4 [Rud92a]. 0-7735-0793-0 [Rud92a]. 0-7885-0273-5 [Sho99]. 0-7885-0274-3 [Sho99]. 0-7914-1821-9 [Mol96]. 0-7914-2011-6 [Cuo95]. 0-7923-0095-5 [Mul91]. 0-7923-0145-5 [Div90]. 0-7923-0420-9 [Non91a]. 0-7923-045109 [Tur91]. 0-7923-0625-2 [Rus91]. 0-7923-0719-4 [Jam91]. 0-7923-0825-5 [Hac93c]. 0-7923-0898-0 [Mar96c]. 0-7923-0958-8 [Ams92]. 0-7923-1019-5 [Hod93a]. 0-7923-1054-3 [Cas94]. 0-7923-1308-9 [Jam92c]. 0-7923-1380-1 [Nor94a]. 0-7923-1520-0 [Mil94]. 0-7923-1577-4 [Non93a]. 0-7923-1658-4 [Smi94d]. 0-7923-1777-7 [Bow94b]. 0-7923-1926-5 [Sch95b]. 0-7923-2336-X [Ste95]. 0-7923-2399-8 [Hel96b]. 0-7923-3114-1 [Bro96b]. 0-7923-3617-8 [Hen98b]. 0-8014-2062-8 [Kra90b]. 0-8014-2248-5 [Yav92]. 0-8014-2340-6 [Bro91b]. 0-8014-2358-9 [Can91a]. 0-8014-2573-5 [Fie94]. 0-8014-2641-3 [Har93a]. 0-8014-2723-1 [Luc93]. 0-8014-2947-1 [Fis98b]. 0-8018-3101-6 [Mar93c]. 0-8018-3102-4 [Mar96a]. 0-8018-3678-6 [Hod90b]. 0-8018-4002-3 [Joh93]. 0-8018-4098-8 [Hug92a]. 0-8018-4126-7 [Bow92b]. 0-8018-4151-8 [Sho93b]. 0-8018-4292-1 [Bai93]. 0-8018-4421-5 [Har95b].

0-8018-4452-5 [Wri94a]. 0-8018-4744-3 [Smi95b]. 0-8018-4745-1 [Smi95b]. 0-8018-4790-7 [Spa98b]. 0-8018-4801-6 [Rou97]. 0-8018-4967-5 [Hor96b]. 0-8018-5141-6 [Har98g]. 0-8018-5152-1 [Aga98]. 0-8018-5173-4 [Hor99a]. 0-8018-5281-1 [Har98c]. 0-8018-5355-9 [Har99b]. 0-8018-5502-0 [Roh99a]. 0-8018-5503-9 [Roh99a]. 0-8018-5523-3 [Daw98a]. 0-8039-9100-2 [Har93b]. 0-8047-1673-0 [Law91b]. 0-8047-1713-3 [Edg91a]. 0-8047-1714-1 [Edg91a]. 0-8047-1721-4 [Edg91a]. 0-8047-1722-2 [Edg91a]. 0-8047-1796-6 [Hoc92a]. 0-8047-1879-2 [Hoc93b]. 0-8047-1950-0 [Smi93c]. 0-8047-2985-9 [Smi99a]. 0-8047-3276-0 [Smi99a]. 0-8057-8613-9 [Boh95]. 0-8057-8638-4 [Boh95]. 0-8078-2220-5 [Bis98]. 0-8078-4531-0 [Bis98]. 0-8108-2165-6 [Byn90]. 0-8108-2570-8 [Law93e]. 0-8122-1297-5 [Cam92]. 0-8122-3307-7 [Por97]. 0-8122-3308-5 [Por97]. 0-8122-8185-3 [Sha92]. 0-8122-8205-1 [Gre91]. 0-8122-8207-8 [Cam92]. 0-8130-0847-6 [Jac90]. 0-8132-0738-X [Mal93b]. 0-8133-1575-1 [Law98a]. 0-8135-1529-7 [Hen92b]. 0-8135-1530-0 [Hen92b]. 0-8135-1700-1 [Dea93]. 0-8135-1823-7 [Nic96a]. 0-8138-0883-9 [Smi90a]. 0-8153-0493-5 [Mol98b]. 0-8153-1801-3 [Lar96a]. 0-8153-1801-4 [Lar96a]. 0-8153-1802-2 [Lar96a]. 0-8153-1803-0 [Lar96a]. 0-8153-1804-9 [Lar96a]. 0-8153-1805-7 [Lar96a]. 0-8153-1806-5 [Lar96a]. 0-8153-1807-3 [Lar96a]. 0-8153-1808-1 [Lar96a]. 0-8153-1809-X [Lar96a]. 0-8153-1811-1 [Lar96a]. 0-8173-0390-1 [Hac90e]. 0-8173-0536-X [Bow93c]. 0-8176-3148-8 [Whi90a]. 0-8176-3148-8/Europe [Whi90a]. 0-8176-3479-7 [Kra93a]. 0-8211-8252-3 [Bro92c]. 0-8223-1253-0 [Bow94a]. 0-8223-1479-7 [Kus96a]. 0-8223-1494-0 [Kus96a]. 0-8240-0039-0 [Law92e]. 0-8240-5171-8 [Hen93a]. 0-8240-6639-1 [Hod93b]. 0-8240-7446-7 [Hal92c]. 0-8248-1177-1 [Bro90c]. 0-8412-3522-8 [Hor99b]. 0-8476-7637-4 [Hil91a]. 0-8476-7751-6 [Ger95]. 0-85115-321-6 [Joh94]. 0-85115-395-X [McC98]. 0-85115-506-5 [Pum91b]. 0-85115-523-5 [Pum91a]. 0-85115-594-4 [Elm97]. 0-85186-294-2 [Mor92c]. 0-85198-689-7 [War91a]. 0-85274-001-8 [Cra92c]. 0-85274-029-4 [Hac92c]. 0-85274-141-3 [Shr92]. 0-85296-895-7 [Mor99]. 0-85323-027-7 [Hal91b]. 0-85323-187-7 [Bra92b]. 0-85389-350-0 [Moo92]. 0-85389-351-9 [Moo92]. 0-85481-072-2 [Smi92a]. 0-85481-078-1 [Pri91]. 0-85964-226-7 [Hen94b]. 0-85989-340-5 [Wea91]. 0-86027-020-3 [Wil90a]. 0-86078-280-8 [Bry92a]. 0-86078-281-6 [Ung92]. 0-86078-394-4 [Gou95a]. 0-86078-464-9 [Kni98a]. 0-86078-465-7 [Luc98]. 0-86341-212-2 [Hac93a]. 0-86341-218-1 [Hor92]. 0-86341-227-0 [Hac92d]. 0-86341-234-3 [Jam92a]. 0-86341-237-8 [Mar93b]. 0-86341-255-6 [Jam93b]. 0-86341-281-5 [Hac94]. 0-87023-881-7 [Hor95]. 0-87023-882-5 [Hor95]. 0-87169-171-X [Por90a]. 0-87169-184-1 [Bai91]. 0-87169-191-4 [Tei93]. 0-87169-214-7 [Kra98]. 0-87169-774-2 [Hun92]. 0-87169-801-3 [Hil91c]. 0-87169-833-1 [Di 95a]. 0-87413-379-3 [Sho93c]. 0-87474-752-X [Edg90b]. 0-88135-066-4 [Hal92d]. 0-88135-071-0 [Dea92]. 0-88135-089-3 [Law91a]. 0-88135-092-3 [Kni90c]. 0-88135-098-2 [Pyl93b]. 0-88135-152-0 [McE95]. 0-88135-154-7 [Spa95]. 0-88135-157-1 [Hes95b]. 0-88135-158-X [Wit95a]. 0-88135-164-4 [Luc96b]. 0-88135-166-0 [Bro98]. 0-88135-189-X [Cro96b]. 0-88318-622-5 [Hoc92b]. 0-88318-719-1 [Hoc92b]. 0-89263-314-X [Ver95]. 0-89766-545-7 [Bur93]. 0-89766-546-5 [Bur93]. 0-901805-51-3 [Ben94b]. 0-903364-06-9 [Gou91]. 0-9041-8023-9 [BD90]. 0-905958-91-8 [Law91c]. 0-905958-98-5 [Law91c]. 0-906450-08-X [Nic94a]. 0-906795-43-5 [Jon91]. 0-906795-81-8 [Sch91b]. 0-907582-67-2 [Mac90b]. 0-907628-79-6 [Gou92]. 0-912289-83-X [Hen90b]. 0-917853-29-6 [Smi90a]. 0-918102-16-2 [Gol91b]. 0-918102-21-9 [Cro96c]. 0-931292-23-9 [Gre92c]. 0-931292-24-7 [Gre92c]. 0-933951-66-3 [Kni98b]. 0-934223-11-4 [Kni92a]. 0-934223-18-1 [Rus94]. 0-934223-34-3 [Har96a]. 0-942299-72-8 [Sut97]. 0-9502557-8-5 [Bry91]. 0-9504478-1-1 [Moo92]. 0-9504478-2-X [Moo92]. 0-9507258-2-X [Twe93a]. 0-9513889-0-8 [Jam90]. 0-9532213-0-X [Har99a]. 0-9635309-8-4 [Mur95b]. 0-9635924-0-8 [Mol94a]. 00 [Bro95a]. 000 [Cle92, Cre91b, Di 93a, Fin92, Rap92]. 0151-4105 [Cro96d]. 0193-5380 [Bro93a]. 0252-7960 [Woo95]. 0296-8916 [Hal90d]. 045109 [Tur91]. 0925-6806 [Fie92]. 0960-8788 [Hen94a]. 0963-6625 [Kni92b].

1 [Aga94c, Gre91, Law93d, Pyl91]. 1-174 [Har93b]. 1-415-01306-2 [Smi91b]. 1-55553-058-3 [Hoc91]. 1-55655-217-3 [Hen94b]. 1-85072-096-7 [Tur93a]. 1-85196-085-6 [Ans96]. 1-85285-067-1 [Osl98]. 1-85296-354-9 [Hen90b]. 1-85315-118-1 [Cra92a]. 1-85315-119-X [Cra92a]. 1-85315-158-0 [Law93d]. 1-85315-249-8 [Pal97]. 1-85343-181-8 [Soy96]. 1-85463-012-1 [Por95a]. 1-85728-292-2 [Law96b]. 1-85728-622-7 [Far99]. 1-85928-287-3 [Met99]. 1-85928-338-1 [Har98a]. 1-871768-00-4 [Lan90]. 1-871768-01-2 [Fau96b]. 1-878822-07-1 [Sim94]. 1-899702-00-8 [Old97]. 1-90 [Ano82c, Ano82b]. 10.00 [Nic94a]. 10.95 [Hud92]. 100 [BD90, Bur94]. 1000 [Mac83]. 114ème [vH92]. 116e [Ano93q, Dav94]. 116th [Ano93q]. 118 [Kus95d]. 118.00 [Jam94d]. 12.95 [Can90a, Law91c, Sec96, Shr92]. 12.99 [O’B96b, Smi99b, Wil95a]. 125.00 [Ili97, Jam92c]. 128 [Ver91]. 13.95 [Kus95b]. 130.00 [Ams92]. 14-22 [BG^{+}92, Out94]. 14.20 [Cam92]. 14.95 [Ben90a, Edg90a, Hen90b, Mor90a, Mor92c, Pum94, Sim92]. 14.99 [Odd91, Ost93b]. 142.00 [Hun98]. 1450 [Lin92]. 148 [Di 93e]. 15-19 [Bry91]. 15.00 [Hal91b]. 15.95 [Bra90, Hac92a]. 15.99 [Get96]. 150.00 [Hen95b, Mil94, Rus91, Smi94d]. 15th [Ano82b, NW82a]. 16.00 [Tur93a]. 16.95 [Swa92]. 1610 [Wey92]. 1610/11 [Hel96a, Kra91c]. 165.00 [Bro96b]. 1660s [Mur92b, U^{+}91]. 1670s [Ili95c]. 1676 [LM23]. 168.00 [Rag95]. 1683 [LHH91]. 1690s [Hun92, Str87]. 16th [BW90, Sch91b]. 17.50 [Hod90b]. 17.95 [Aga95b, Twe91b]. 175.00 [Hod93a, Ste95]. 178.00 [Maz98b]. 178FF [Cro94]. 17th [Hac92b, Kle94, Grm90, Pic87, Poi04]. 18.00 [Bai93]. 18.99 [Har99a]. 180.00 [Bow94b]. 1830s [Top92]. 1837 [OS92b]. 185.00 [Mul91]. 1862 [Buc92]. 1872 [vLvHH^{+}88]. 1877 [EJR94, EJR95]. 18th [Hac92b, Kle94, Poi04, Woo95, Bec91b, BaJANJ84, FHR90, Roc88, Sch01, SGS00]. 18th-Century [Woo95]. 19.00 [Jam92a, Mar93b]. 19.50 [Pic92b]. 19.95 [Bar90a, Bro92b, For90a, Hod90a, Law90c, Ost93a]. 190F [Hil98a]. 1911 [War95a]. 1913-65 [Ano82c]. 1914 [Dau98]. 1920s [Bow98, May97b]. 1930s [Bow98, Kuz87]. 1939 [HP85, PvM85]. 1940 [Pes84]. 1945 [PCA91]. 1948 [Ano97d]. 195-286 [Bro93a]. 195.00 [Hen98b, Nor94a]. 1950s [Kra91b]. 195FF [Out94]. 1960s [Fis98a]. 1979 [Ano82a]. 198.00 [Kra93a, Mar94a]. 1985 [A^{+}87]. 1986 [Sil90]. 1987 [CM90]. 1988 [BG^{+}92]. 1990 [Van91]. 1990s [Mur92b, U^{+}91]. 1991 [Ano93q]. 1992 [Ped92b]. 1995 [Ano96o]. 19th [Ano82b, CM90, Hac90e, Kni89, KI92, Mar94a, NW82a, Dau98, Jun89, Man88, PP95, Sch89]. 19th-century [Hac90e, Man88]. 1st [PCA91].

2 [HS91, Pyl91, Pyl92, Smi92a, vH92]. 2-07-072992-3 [Hal95]. 2-08-211208-X [Gol95b]. 2-13-044774-0 [Bow93b]. 2-13-046972-8 [Cro96d]. 2-225-84057-1 [Bow95]. 2-228-88277-1 [Pyl91]. 2-251-34504-3 [HS91]. 2-251-34505-1 [Non93b]. 2-252-62032-X [Non94]. 2-7084-0367-2 [Kel95]. 2-7116-1034-9 [Pyl92]. 2-7116-1038-1 [Neu92]. 2-7116-9427-5 [Chu91]. 2-7116-9755-X [Out94]. 2-7351-0685-3 [Hil98a]. 2-7355-0171-X [GG90]. 2-7355-0208-2 [vH92]. 2-7355-0272-4 [Dav94]. 2-85367-196-8 [GG96]. 2-85704-384-8 [Cro94]. 2-86563-212-1 [Osb90b]. 2-86639-079-2 [Fin93c]. 2-89007-646-6 [Chu91]. 2-908212-13-7 [Bow93d]. 2-908212-14-5 [Bow93d]. 20 [Bru91]. 20.00 [Bro93c, Bru93, Goo96, Por95a]. 200 [Har92c]. 200.00 [Sho93a]. 200FF [Bow93d]. 20th [Aga97b, CM90, Hes95b, HW81, Lin93, Neb95]. 20th-century [Hes95b, Lin93]. 22.00 [Fau96b, Gou91]. 22.50 [Blo93c, Cla92c, Di 93c, For91, Gra91, Hen90a, Law98a, Ley90, Str90, Twe92c]. 22.95 [Kla95]. 23.50 [Hoc90a]. 23.95 [Ver90]. 24.00 [Kni90a]. 24.55 [Mor90d]. 24.70 [Har90b]. 24.75 [Hug92b]. 24.95 [Joh96, Twe91a]. 240.00 [Non93a]. 24000 [Pan91a]. 25 [Sha96, Wil90a]. 25-27 [Ver95]. 25.00 [Har98g, Kni90b, La 97, Law98b, Smi91c, Wea91]. 250FF [Bow93d]. 250th [AL87]. 26.95 [Luc92]. 260.00 [Hac93c]. 27.50 [Ait90, Whi90c, Wri94a]. 27.95 [Dol90a, Hod92a]. 28.95 [Mur95b]. 280.00 [Hel96b]. 29 [Ano93q]. 29.00 [Jam93b, Nye93b]. 29.50 [Sim94]. 29.95 [Bow94a, Nye92a, Pal90, Pal98]. 291 [Hal90a]. 298.00 [HD97]. 2nd [Hen90b].

3 [Di 92, Di 93d, HS91]. 3-05-000075-9 [HD97]. 3-05-000766-4 [HD97]. 3-05-001920-4 [HD97]. 3-05-002602-2 [HD97]. 3-11-011507-7 [Pos90]. 3-406-01660-X [Met93]. 3-406-01661-8 [Met93]. 3-406-35092-5 [Sch93c]. 3-437-11271-6 [Ver91]. 3-437-11305-4 [Di 93e]. 3-465-02590-3 [Kus95d]. 3-486-26273-4 [Bro94a]. 3-486-56337-8 [Fuc99]. 3-515-05020-5 [Lor90]. 3-515-05214-3 [Mol92]. 3-515-05369-7 [Gre92c]. 3-515-05373-5 [Alt91]. 3-515-05385-9 [Hin91]. 3-515-05584-3 [Gre93b]. 3-515-05865-6 [New92]. 3-515-05874-5 [Hel96a]. 3-518-57887-1 [Wei90]. 3-593-34763-6 [Har94b]. 3-7186-5200-5 [Nye93b]. 3-7186-5792-9 [Div98]. 3-7643-1974-7 [Koe91]. 3-7643-2316-7 [Jam94d]. 3-7643-2552-6 [Mar94a]. 3-7643-2728-6 [Mol94b]. 3-7643-2919-X [Met96a]. 3-7643-3148-8 [Whi90a]. 3-7643-5003-2 [Bro95a]. 3-7643-5311-2 [Maz98b]. 3-7692-1225-8 [Di 92]. 3-7995-7639-8 [Gre93a]. 3-9 [vH92]. 3-928068-03-2 [Eck92]. 3-928068-04-0 [Eck92]. 3-928186-09-4 [Hug96]. 3-928186-15-9 [Hes95a]. 3-928577-13-1 [Tur93b]. 30 [Hac94, Hak95b, Mac90b]. 30.00 [Blo93b, Fin93a, Hut92, Jam94b, Lar90, Mol98a, Por91b, Wil91c, Yeo92]. 30.50 [Joh93]. 300.00 [Cas94, Dav94, Hal94c]. 31 [Ano96o]. 31.50 [Bow92a, CK91, Smi93a]. 32 [Hac92d]. 32.50 [Hef96, Hod91]. 32/Wellcome [Spa98a]. 32467X [Fra90]. 33.00 [Bow91b, Har99b]. 33.65 [Hoc91]. 33.75 [Twe92a]. 34.50 [Bow92b]. 34.95 [Bro95d]. 35 [Hen91b, Pic93a]. 35.00 [Bry91, But94, Har90a, Har98d, Jac93, Law92c, Law92d, Met96b, Smi93d, Smi94a, Sta90b]. 35.50 [Fis98b]. 35.95 [Twe92b].

350th [vLP82]. 36.50 [Blu90]. 360 [Pug84]. 36000 [Per93]. 365 [Pos90]. 37.00 [Har95b, Hor99a]. 37.50 [Goo90a, Gou92, Hay98b, Ree90, Soy96]. 38.00 [Har94b, Rob93]. 39.50 [Hac92c, Pum91a]. 39.95 [Man97]. 3rd [Gou91]. 3UH [Fau96b].

4 [Kni92a, Maz98a]. 4-87378-299-6 [Mol94b]. 4.45 [Fie91]. 4.95 [Jam90]. 40 [Aga93, Alb92, Fie93, Ful98, Tha91]. 40.00 [Aga94b, Hes95a, Jam92b, Jon92, Law96b, Mar92a, McC95, Osl98, Smi92c]. 40.50 [Buc90, Twe91c]. 41.50 [Aga98, Har98c]. 41.95 [Rus94]. 42.00 [Jon91]. 42.50 [Luc98]. 42.75 [GG91b]. 44 [Hac93a]. 44.95 [Twe93b]. 45 [Hac90d]. 45.00 [Dol99, Har98a, Har96a, Hoc92c, Hug98, Law92b, McK96, Mor99, Pum91b]. 450FB [Rus92]. 47.25 [Byn90]. 47.50 [Kni98a, McN92, Ung92]. 48.00 [Dol90b]. 480 [Bow93b]. 49.50 [Bry92a, Sec98]. 49.95 [Ans96, Kra94].

5.00 [Bra92b]. 50 [Eck92]. 50.00

[Kus94, Law98c, Spa98a]. 50000 [Fin93b]. 505203X [Kni90a]. 51.00 [Tur93b]. 54.00 [Mar93c, Mar96a]. 55.00 [Hal94a, Hen93b, Lev94, Wil94a]. 565 [Hen92a]. 58.00 [Gre93b]. 59.00 [Hor92]. 59.65 [Law93e]. 59.95 [Fyf99].

6.00 [Law90a]. 6.99 [Jam94a]. 60.00 [Elm91, Har93c, Jam91, Mur92b, Non91a, Sch95b]. 600 [Lin92]. 65.00 [Hal92a, Smi91b]. 68 [Hin91].

7 [Mor91c]. 70 [Alt91, Hal90d]. 70.00 [Wri96b]. 700-1300 [Hil91a]. 74-7670-165-9 [Goo91b]. 75.00 [Blo94b, Far94, Sch91b]. 7th [Bry91].

8 [Cra92c, Di 92, Fie92, Pri91]. 8.00

[Ben94b]. 8.75 [Yeo96a]. 8.99 [Bra91, McE93, Nic94b]. 8/9 [Hal90d]. 8/Europe [Whi90a]. 80.00 [Lan90]. 800 [C^{+}95]. 81-7036-245-8 [Har93b]. 81-85150-11-7 [Har92d]. 81-85150-19-2 [Har92c]. 84 [Maz98a, Oli92]. 84-00-06745-2 [War90a]. 84-00-07056-9 [Oli92]. 84-206-9412-6 [Mal96]. 84-7283-286-4 [Maz98a]. 84-7670-164-0 [Goo91b]. 84-7670-191-8 [Goo91b]. 84-7670-204-3 [Goo91b]. 84-7670-205-1 [Goo91b]. 84-7670-216-7 [Goo91b]. 84-7670-217-5 [Goo91b]. 84-7822-154-9 [Mal98]. 84.00 [Koe91]. 85.00 [Hor98]. 85000 [Bal91b]. 87-7304-236-6 [Hal94c]. 87-7492-910-0 [Hac95]. 87-7709-010-1 [Jam94c]. 88 [BG^{+}92, Sch93c]. 88-09-20793-9 [Lar96b]. 88-15-02088-8 [Cre91b]. 88-15-03647-4 [Fin93b]. 88-204-6373-3 [Pan91a]. 88-204-6663-5 [Per93]. 88-204-6733-X [Rap92]. 88-211-9566-X [Fin93a]. 88-222-3761-7 [Bal91b]. 88-222-4196-7 [Mor96a]. 88-222-4411-7 [Gou98a]. 88-339-0563-3 [Di 93a]. 88-7761-046-8 [Ben93]. 88-86166-24-9 [Di 98]. 88.00 [Gre93a].

9 [Mor93, Nye92b, Oli92, Pyl92]. 9.95 [Col93]. 9.99 [Gre95, MK98, Mor92f]. 90 [Fie92]. 90-04-09378-8 [Fie92]. 90-04-09641-8 [Rag95]. 90-04-09740-6 [Mol95]. 90-04-10118-7 [Hun98]. 90-04-10371-6 [Osl97]. 90-272-4553-3 [Sho93a]. 90-277-2647-7 [Dol90b]. 90-351-0786-1 [Hac93b]. 90-365-0379-5 [Alb92]. 90-5183-121-8 [Hac92b]. 90-5183-818-2 [Spa98a]. 90-5235-071-X [Ili97]. 90-73368-05-7 [McK96]. 90.00 [Mal93a, Mol95]. 905 [Nic93b]. 91-554-3157-7 [Whi97]. 92-9092-125-0 [Wri96b]. 95.00 [Hal92c, Law93g]. 951-710-001-6 [Bro96a]. 98.00 [Hel96a, Met96a]. 99.00 [Mar96c]. 9DX [Jam90]. 9th [PCA91].

= [AHY90, GP92, KA81, SLT88, VB80].

Âge [Osb90b, BBC88].

A$ [Bro95d]. A$25.00 [Kni92b]. A-M. [Blo82]. A. [Bec99, Gou92, Hac92a, Hut92, Kra97, Mar94a, Mur96, Osl98, Ost93a, Smi92a, Tra92, PHvMW93]. A.D [Hil91a, Nut93]. A.D. [C^{+}95, Lin92, Mac83]. Aaserud [Kri91]. Abbey [BaJANJ84, FM00, Gay98]. Abbott [Jam92d]. Abbri [Cle92, Cro96b]. Abhandlungen [Bro94a, Kus95d, Van91]. Abraham [Maz96, DR95, Hen92a]. abridgement [NR86]. Absolute [Bar89a, Fra90, Gra91, Ear89]. Absolutism [Ili95b, Bia93]. abstract [Wus84]. abyss [Ros84b]. Academia [Ben93]. Academic [Aga97b, Bow94b, Bro96b, Cas94, Cri90, Div90, Div98, Dol90b, Hel96b, Hen98b, Hod93a, Hun92, Jam92c, Mar96c, Mil94, Nic93c, Non93a, Nor94a, Nye93b, Sch95b, Smi94d, Ste95, Swa88, Tim98, Eng83, Mar92b]. Académie [McC98, Str87, Stu95]. Academy [Baa90, Bro90a, Bro96a, Bur93, Cro92, Fie92, Frä89, Hal94c, Hun92, Kni90c, Mor95c, Ped92b, Hey83, Kal95a, Vuc84]. according [Fie92, Hen91a]. Account [Ano96o, Hul88, Lee96, Sta84a]. Accounting [Nig99b, Pow96]. Achievement [Luc98, Gou95b]. Acid [Rus91, Ben90b]. Ackerman [O’B96a]. Acres [vH92]. across [Di 93c, Kni96, Pan91b]. Acta [Whi97]. Actes [Dav94, Out94, Ano93q, BG^{+}92, Pic87]. actinometry [Vos97]. Action [Alb92, Ste91c]. Active [Hac93a, Cal90]. activists [Kuz87]. actors [Kai94]. acts [BG^{+}92, Pic87]. actuels [B^{+}95, Cro96d]. AD [Bru96]. Ada [Ste85]. Adam [Cra92c, Hac92c, Law92b, Shr92, Hak95b, Spe82]. Adams [Cra92a, ATKL08]. adaptations [Sul82a]. additions [But94, Sar87].

Address [Can97, Kni96, Bro99a]. Addresses [Har95a]. Adelard [Mol94b, BEoCF92, ECL92]. Administration [Cro95, Don93]. Adosa [Fin93c]. Adriaan [Hen86b]. Adrian [Bro93c, Goo96, Hod92a, Law96b]. Adrien [KW87]. adult [Sut99]. Advanced [Ost99, Wil90a]. Advancement [Mor84b, Hol86]. advent [Pye85]. Adventurer [Ost93a, Hal92b]. Advisors [Edg91a, Yor89]. Advocate [Pum90, Joy87]. aeris [SSH85]. aerodynamics [HV82]. Aeronautical [Edg90b, Pen89]. aesthetics [Ano04, WM85]. aether [Gol83]. Affair [Lit91]. Affiliation [Lar96a, Kal95b]. Affinität [Bro95a, Kle94]. Affinities [Sho93a]. affinity [Adl87, Kle94]. Africa [G^{+}09]. Aftalion [Cam92]. after [AS96, Can83, Har98h, Nic96b, Sir87, Wil84]. afterword [Ost94]. against [JK84]. Agassiz [Di 93d, Win91]. Age [Can90a, Can96, Hac92c, Hod90a, Mau95, Por97, Pum90, Pum92, Rou97, Swa92, Twe90, Bec95, Ber85, BBC88, Bur83, Cau89, Cer89, Cor88a, Hak95a, Ili95a, Joy87, Kni86, Kni88, Kni89, Mor84a, Tre97, Tur96, Wai90]. Agency [Jam91, Wri96b, Goo90c]. Agenda [Bow93c, Rai91]. Ages [Ano82c, Mol95, Gol91a, NW82b, Sor83]. agnosticism [Lig87]. Agricola [HS91, McE95, Ber93a]. Agricultural [Pal90, Coo89a, Pri83]. agriculturist [McC84]. Ahmad [Hil93]. AI [McK97b, Cre93]. AIDS [Fox92]. aim [PB83]. aims [Max84, Pum91c]. Air [Edg90b, Pen89, SSH85]. air-pump [SSH85]. Airy [Rot93b]. Aiton [Dun92b, Dun92a]. Akademgorodok [Woo98, Jos97]. Akademie [HD97]. Al [Hil93, LN95, Mal98, Mül93, Smi92a, AS89, vdW85]. Al-Andalusi [Hil93]. al-Haytham [Smi92a, AS89]. al-Khwarizmi [vdW85]. Alabama [Bow93c, Hac90e]. Alan [Mal93a, Mal94, Mur92b, Str90, Ver90, Aga93, Blo94b, Jam90, Pal97, Pic98, Stu91]. Albany [Cuo95, Mol96]. albero [La 94]. Albert [Ari90, HS91, Oli92, Ver95, War95a, War96a, BH87, Hel96a, Pai82, SCS^{+}87, Woo95]. Albertini [Law91a]. Albertus [Zam92]. Albrecht [Hes95a]. Alchemic [Gre93a]. Alchemical [Gre92c, Hor99b, Pri91, Mor91a, Per89, RCRC98, Wey92]. Alchemie [Wey92]. Alchemist [Man97, New94]. Alchemistische [Gre93a, Wey92]. Alchemists [Gre96, Pat94]. Alchemy [Gre96, Hun90, Man93b, Wey92, Dob91, GGC90, Kri82, Smi94c]. alchimia [Bal91b, GGC90]. Aldershot [Bry92a, Gou95a, Har98a, Kni98a, Luc98, Met99, Ung92]. Aldrich [Twe93b]. Aldrovandi [Fin91]. Aleksandr [FS88]. Alexander [Jor94, Kra94, Aga94a, Hei88, Roh99a, TFC93]. Alexandria [Aly90, Jon86, VoC89]. Alfons [Har98h, Mal98]. Alfred [Bro94b, Bro84, Law92a, Smi92c, WS91a]. Algar [Hor92]. Algarotti [Fin93a, Ara91, Fin91]. Algebrâ [Viè83]. Algebra [Mar94a, Maz98b, Cor96, Fis94, KI92, Sin91, Viè83, vdW85]. algèbre [Neu92, Sin91]. Alianza [Mal96]. alias [Kno90]. Alice [Hun92, Hun92]. Alison [She92]. Allan [Tre97, Kus96a]. Allen [Pri93]. Allied [Mor90d, Sto88]. Allison [Bro96b]. Alloys [Buc90]. Almagest [PT84]. Alok [Hil93]. alpacas [GL88]. also [Bow92c]. Alter [Mac90b]. Alteration [Far99, Poo98]. Alternative [Law90b, Coo88]. Alternativen [Di 92, Jun89]. alternatives [Jun89]. Altertumswissenschaft [Hin91]. Altruism [Bar93b, Cro91a]. Alvarez [AT87]. Alvaro [Goo91b, Fre90b]. Alvey [OO89, Twe91a]. am [AH93, Bro94a, Hes95a, Sch92b, Wei90]. Ambiguity [Ste97a]. Ambiguous [Mor93, Per92]. Amedeo [Mor84c]. America [Aga97c, Aga98, Blu90, Hen94b, Hor95, Hor96b, Lar96a, Luc92, Mal93b, Mor92d, Nic93b, Ost99, Uim95, BEFH10, Doe96, Edw96, FL88, HV82, Kuz87, Lun94, Mis95, NBR^{+}95, Nye90, RR81, RT95, Seg94, Swa88, T^{+}85, War86]. American [Bai91, Bow92b, Bow93c, Bro90c, Bug96, Di 95a, Gre91, Gre92c, Hil91c, Hoc92b, Hor99b, Hun92, Jon93, Jon94, Kra98, Lar96a, Man97, Mau95, Nic93a, Nye93b, Por90a, Sho99, Sil90, Smi90a, Tei93, Bru87a, Deg91, Gos87, HL86, Kal95b, Lie87, Liv87, Mai91, MR89, Mit92, MM10, Mor89, Moy83, New94, Ost84, Pit93, Rai91, Rei85, SB92, Who82, Wri94b]. Ames [Hel96b, Smi90a]. Amherst [Hor95]. Among [Mur96, Rud85, SW95, Van93]. Ampère [Blo82]. Amphion [Fin93c]. AMS [Por90b]. Amsterdam [Div98, Hac92b, Hac93b, Hug92c, Sho93a, Spa98a]. analogy [Tau93]. analyseos [Viè83]. analysing [Hes99]. Analysis [Law92b, Maz96, DR95, Ham90, Hea86, Hea90]. Analyst [Mil94, BJB92]. analytic [Viè83]. Analytical [Smi99b, Bec95, Sha98]. Analytics [Non93a, Wal92]. Anamika [Har92c, Har92d]. Anatomical [Har98a, Law93g, Sch91b, BW90, Cun97, RT92]. Anatomy [Di 93f, Arm83, Liv99, OS92b]. Ancaster [Ost99]. ancestors [Des82a, Des82b, Des84]. Ancient [Bai91, Bar94, Cam90, Cla99a, Cla92a, Cla95a, Edg90b, Fre94, Get96, Kra98, O’B96b, Whi90a, CK90, HW87, Kre82, Llo83, Pen89]. Ancients [Har98a, Cun97]. Andalusi [Hil93]. Anderson [But92a, But94, Gou95a]. André [Bru93]. Andreas [Fuc99]. Andrew [Bru96, Elm91, Lee90, Met99, Mur92b, Spa98a, BD90, Har98a, Pal90, Ste91a]. Andrews [Mar93c]. Anfänge [Gol91a, Gre93b]. Angeles [BV91, Kni95, Por90b, Twe92b, Twe93b]. Angeli [Pan91a, Per93, Rap92]. angles [Cha90d]. Anglican [Cor88b]. Angular [Hac92a, Cha94]. aniline [Tra92]. Animal [Cro91b, Mor92a, Mor93, Ras96a, Sch91b, BW90, Nyh95, Per92, Rus86b]. Animals [Alb92, Mor82, Ste91c]. Anita [Tur93a]. Ann [Mic98, Har98g]. Anna [Mar96b, War96a, Pal91a]. Annabella [Bra91]. Annals [Bur93]. Anne [Hen93b, War96a, CMNH92, Har90b, Har96c, Roh99b, Sha96]. annexes [CLR87]. Annibale [Sha97]. Annie [Pyl92]. Anning [Tor95]. anniversary [AL87, vLP82]. Annotated [Byn90, Hud92, Woo95, Coo89b, DeV82, Ogi86, Por83]. annotations [Coo90]. Announcements [Ano90m]. Annual [Bow94a, Mar92a, Kni92b, Wei92]. annum [Cro96d, Hen94a]. Ant [Bar93b, Cro91a]. Antarctic [Lev94, Fog92]. Anthology [Lar96a, Smi92c, NBR^{+}95, WS91b, WS91a]. Anthony [Day99, Law90b, Por91a, Rus94]. anthropological [ME81, Ver85]. Anthropologie [Ver91, Atr86]. Anthropology [Bow93a, Wat91, Atr86, Cru90, Kuk91, Kuk93, Sto84]. anti [Bow83, Hac84]. anti-Darwinian [Bow83]. anti-submarine [Hac84]. Antibiotics [Swa92, Wai90]. Anticipations [KK87]. Antievolution [Lar96a, BN95, RT95]. Antievolutionism [Num95a]. antiques [SB92]. Antiquity [Bai93, Bow93c, Fie91, Get96, Hil91c, Hud92, Whi90c, Ali86, Cro90b, Gre92a, O’M89, Ogi86, Rai91, Sor83, Whi90b]. Antoine [Cro94, Cro95, Don93, Poi93]. Anton [Di 95a, vBDGO93]. Antonello [Di 93a, Mül93]. Antoni [Maz98a, vLP82]. Antonio [Mal96, Goo91b, Rap92, Ree90]. ants [Cla97]. Anwendungen [Koe91, Sch89]. Anziehungskraft [Adl87]. Ape [McK96]. Apeman [McK96]. Apologia [JK84]. Apotheker [Di 92]. Appendix [vLvHH^{+}88]. Appleyard [Jam94a]. applications [Sch89, WW86]. Applied [Mil95]. appointment [Ric97]. apprenticeships [Lan90, WW88a]. Approach [Dol90a, Gie88]. Approaches [Bur87, Law96a, Ner87]. Approaching [Law96c, Mic95]. Appropriated [Non93a, Wal92]. Appropriating [Fri89, Kra90b]. approximation [Kol92]. après [Bow93d, Gay92]. April [Ano93q, Mar96a, EJR94, EJR95]. Arabic [Bur98, Wea91, Dun90, RM96]. Arato [Fin93a]. arch [Old86]. Archaeological [Bow91a, Tri89]. Archaeology [Jor91, MMVZ90, Mur92a, She92, Gut89, PW89]. Archeology [Mur95b, CT93]. Archetypes [Des82a, Des82b, Des84]. Archibald [Jam94c]. Archiv [Gre92c]. Archive [Ano91p, Ano92a, Ano98a]. Archives [Bro96b, Sch95b, BKH90, PLP92, RR90, SW93, Top97, TW87]. Arctic [Hig94, Lev93]. argent [Poi04]. Argument [Sha92, Dea91, PB82]. arguments [Des89b]. Ariew [Mal93b]. Aristarchus [Van85]. aristocracy [Bec95]. Aristotelean [Fur88]. Aristotelianism [Bru87b]. Aristotle [Bos92, Mol98a, Non93a, Wea91, Dun90, Eva87, Fre95, Mor82, Sch83, Wal92, War90b]. arithmetical [Edw87]. arithmetice [BJ91, Mol92]. Armagh [Moo92]. armed [Mac83]. Armin [Alt91]. armonia [Di 93a]. Arne [Luc96b]. Arouet [Cas93]. Arrhenius [Bro98, Cra96]. arrived [Fis94]. Art [Bro91c, Dea92, Fie94, Har98d, Law93g, Osl98, Poi04, Rav99, Sch91b, Aly90, Bea91, BB84, BW90, Cro96a, Edg91b, Fie97, Gil83, HA87, Noa99, PS91, Sta84a, Sta91, Tom84, Ung92, Viè83, VoC89]. arte [Lan86]. Artful [Joh96, Sta94]. Arthur [Aga98, Lar96a, Olb91b, Sut97, BN95, Cro95]. Arti [Di 98]. Articles [Har95a]. Artificial [Hac91, McK97b, Twe91b, Cre93, THBH92b]. artis [Mol92, BJ91]. artisan [Ili95c]. artisans [Sec94]. Arts [Hil91c, Por97, GJ86, Whi90b]. Arturo [Wri96b]. AS50.00 [Kni92b]. Asclepius [Cop92]. Asdepius [Mon93]. Ash [Har98h, Kus96b]. Ashcraft [Ost93d]. Asia [Mar10]. Asian [Ald85]. Aspects [Fie94, Hes95b, Buc85, Lin93, Ste97a].

Aspray [Aga98, Bow92a]. assistance [Hen92b, Hod93b, Hug96]. Associated [Har96a, Kni92a, Rus94, Sho93c]. Association [Kni92b, Mor84b, Pur99, Sho99, Smi90a, Soy96, Bro90a, Edg90a, Hac94, Hor92, Jam92a, Jam93b, Law90b, Moo92, Sch91b]. astrologia [Bal91b, GGC90]. Astrology [Bra91, O’B96a, O’B96b, Bar94, GGC90, Kit89, Smo94, Zam92]. Astronomer [Cha90c, HD97, FFLW97, How89]. astronomiae [Nor94a, Zam92]. Astronomical [DTS^{+}87, Kri88, STT87, Tay87, TST87, Wil90a, BR96, Ger89, Hea86, Hea90, Mol90, Mur85, Mus98, PR90]. astronomie [Pic87]. astronomisches [Alt91, Ger89]. Astronomy [Aga94b, Aga97c, Bra95, Hac90d, Hac92a, Kra98, Mol98b, Moo92, Wil91b, Wil95a, Wil97, Ash94, Ben87, Cha94, Cla95a, Cla95b, DeV82, Doe96, EBT96, HK84, Het88, Hir83, Kel91, Nor94b, Pic87, Rob92, Smi82, Ste87, Ste94, Sul82b, TW95]. astrophotography [Pan97]. astrophysicist [Cha83, Ost84]. Astrophysics [Wil91b, Wil97, DeV82, Ost84, TW95]. asylum [Tom84]. asylum-keeping [Tom84]. Atherton [Can91a]. Athlone [Hay98a, Kus94, Rob93, Smi94a]. Atlanta [Sho99]. Atlantic [Wit95a, Eri94]. Atlantis [Woo98, Jos97]. atmospheric [Jan98]. atom [HA84, Wil87b]. Atomic [Kel83, Smy89, Edg91a]. atomism [Roc84]. Atomist [Pum90, Joy87]. Atoms [Fre93, Osl91]. atop [Dra81]. Atti [Cle92]. Attitudes [Bru91, Di 98, Jam92c, DH90, Maz95a, Ros91]. attraction [Far95]. attractive [Adl87]. Attributed [Pri91, Per89]. Auditors [Cri94, Cri96]. August [Pic99]. Augustan [Pum95b]. Augustus [Ric97]. Aurora [Chr93]. auspices [Smy89]. Aussagen [Gro89, Hin91]. Austin [Hil93]. Austoker [Por91b]. Australasian [Kni92b].

Australia [Bro95d, Fin93d, Moy86]. Australian [Bro90a, Bro91c, Nor89]. Austrian [Blo93b]. Author [Jam90, Ost94, Her91]. Authority [Mor90d, Lar91, Out84, Red87, Sto88]. authors [Ano91r, Ano92o, Ano93p, Ano94o, Ano96p]. Autobiographical [Bro90c, MR89]. Autobiography [Smi98, GH84, GH96, Tod83]. Autographs [Bra92a]. autour [BD95, B^{+}95, Cro96d, Mor96a]. available [Bow92c]. avait [Lan87]. Avebury [Mur92b, U^{+}91]. aviation [Gil83]. Avicenna [Sir87]. Avogadro [Mor84c]. Avoiding [Har96a, MGBE95]. Avril [vH92, Ano93q]. avril-4 [Ano93q]. Awareness [Ben90a, Whi88]. Ayers [Gau99].

B [Bru93, Cam90, Cra92a, Fra90, Gou98a, Hac93a, Har98g, Kat90, Lar96a, Law93g, Mar93c, Mar96a, MW95, Nic96a, Olb91a, Ost98, Rud92a, Smi92b, Smi92d, Wil97]. B. [Nye93b, Smi93b]. B.C [C^{+}95, Nut93]. B.C. [Lin92, RL97]. Baatz [Bur93]. Babbage [Twe91c, Twe92b, Twe93b, Ash94, BCKR89, Lin90]. Bachelier [Dim93]. Back [Aga97a, Ano90a, Ano90c, Ano90e, Ano91a, Ano91c, Ano91e, Ano91g, Ano92b, Ano92d, Ano92f, Ano92h, Ano93a, Ano93c, Ano93e, Ano93g, Ano94a, Ano94c, Ano94e, Ano94g, Ano95a, Ano95c, Ano95e, Ano95g, Ano96a, Ano96c, Ano96e, Ano96g, Bro99b, Cra92c, Fau96a, Kus95a, McK97a, Mor92e, Mur95a, Smi90b, Sta90a, Ste97b, Kai94, Roc90]. Background [Non93a, Can93b, McC81, McI85, Wal92]. Bacon [Ree90, Sha93a, BL83, Mar92c, PR88, PR90, Vic87]. Baconian [Sho96a, Smi94b]. Baden [Cor88b]. Baer [Di 95a, BEFH10, vBDGO93]. Bailie [Hud92]. Baily [Ash94]. Baker [Hor99a]. Bakken [Law93e, OS92a]. Bakker [Hac93b].

Bala [Har93b]. Balaguer [Maz98a, Mal95b]. Balan [Osb90b]. Balance [Col93, BD95, HE92, Mor96a]. Balbinot [Smi93b]. ballooning [Gil83]. Baltimore [Aga98, Bai93, Bow92b, Daw98a, Har98c, Har99b, Har95b, Har98g, Hod90b, Hor96b, Hor99a, Hug92a, Joh93, Mar93c, Mar96a, Mol94a, Nye93b, Roh99a, Rou97, Sho93b, Smi95b, Spa98b, Wri94a, KT87]. Band [HD97]. Bang [Aga94a, SN93]. Bank [Aga94c, Nic96a, Hag92]. Baptista [Pag82]. Barbara [Joh96, Law93g]. Barber [Cas93]. Barbour [Fra90]. Barcelona [Maz98a]. Barker [Law93b, Mal93b]. Barker-Benfield [Law93b]. Barnard [Aga96a, She95]. Barnes [Bow98]. Baron [Smi92b]. Baroque [Wit95a, Eri94]. Barrow [Hal91a, Fei90]. Barry [Rud92a]. Barsanti [La 94]. Bartholin [Jam94c]. Barton [O’B96b]. Bartram [BB82]. Basalla [HS90]. based [Gun84]. Basel [Bro95a, Jam94d, Koe91, Kra93a, Mar94a, Maz98b, Met96a, Mol94b, Whi90a]. Basic [Aga98, McK97b]. Basil [Bro92b, Hoc92d, Kni90b, Law92d, Lee90, Smi91c, Swa92]. Basingstoke [Law92c]. basis [Sch92b]. Batchelor [Hug98]. Bauer [Bro91c, Hor96a, Hor97b, Nor89]. Baumeister [Bro94a, Sch92b]. Baustatik [Koe91, Sch89]. Bayertz [Wei90]. BC [Bru96, Sho99]. Bd [HP85, LM23, Met93, Van91]. Be [Hey95, Hun98, F^{+}88]. Beagle [Bro90e, DK88, DD94]. Beatty [Hod91]. Beauchamp [Mor99]. Beaujouan [Ung92]. Beaumarchais [vPvP90]. Beaune [Out94]. Because [Bea91]. Bechler [Cas94]. Bechmann [Kel95]. Beck [Met93, War96a]. Beck’schen [Sch93c]. become [War90c]. Becquemont [Bow93d]. Beddoes [Sta84b]. Bedeutung [Fri91, New92]. Been [Kus95b, May99]. Beer [Law98b]. Beeson [Gou91]. Before [Fei90, FC96, Gos87, Hal91a, Lar96a, Met99, Bar84, Bro84, Ger89, NC81, Num95a, Sti86, Web88]. Beginning [Buc90]. Beginnings [Nut93, Bro97b, Cra84, Gol91a, Lin92]. Behalf [Mar93b, Hac92d, Hac93a]. behavior [Gra87]. behavioural [Cou96]. Beiheft [Gre92c]. Beiträge [Eck92, Hel96a, Tur93b]. Beitragen [Gre93b, Gol91a]. Belfast [Moo92]. Belief [Ost99, Dur85, Hoo97]. Bell [Lar96a, Rei85]. Bellarmin [Chu91]. Bellarmin-Vrin [Chu91]. Beller [War95b]. Belles [HS91, Non94, Non93b]. Below [Bry88]. Beltrán [Mal96]. Bénédict [Woo95]. Benfey [Bro92c, Cam92]. Benfield [Law93b]. Bengal [Har93b, Bal91a]. Benjamin [Law92d, Sho93a, War91b, Bro89, Cro84, Jam92d]. Bennett [Gou95a, Hac90d, Moo92]. Benninga [Rus91]. Benoit [Out94]. Bensaude [Cro96b, Gol95b]. Bensaude-Vincent [Cro96b, Gol95b]. Bentham [dCC09]. Benzolfest [Sch93a]. Bérard [Men90]. Beretta [Lar96b, McE95]. Berg [Mac90b]. Bergia [Smi93b]. Bergren [Mol98b]. Berichte [Bro94a]. Berkel [Hac93b]. Berkeley [BV91, Can91a, Cro96c, Gol91b, Hun92, Kni95, Mil94, Mor90d, Por90b, Pum92, Ath90]. Berlin [Bro95a, Eck92, HD97, Jam94d, Koe91, Kra93a, Mar94a, Met96a, Mol94b, Tur93b, HRP90]. Berliner [Eck92, Tur93b]. Bermannus [HS91, AHY90]. Bernadette [Cro96b, Gol95b]. Bernard [BD90, Cro96c, Hen92b, Osb90b, GGSGB94, Nig99a]. Bernd [Sho93a]. Bernhard [New92]. Berrios [Hay98a]. Bert [Hac93b, McK96]. Berthollet [SG80]. Bertoloni [Hal94a]. Bertotti [Smi93b]. Berzelius [Bro93b, MF92]. besoin [Lan87]. Bessemer [Bar84]. Bethe [Edg91a]. Bethesda [Hen94b]. Bethlehem [Kni92a, Rus94]. Betty [Man93b]. Between [And84, Hey83, Mor90a, Sho93a, Smi92b, Ano90k, BM82, BaJANJ84, Ber95, Clu88, De 96, Fie92, Gre86, Hen91a, Lan91, Lar96b, Pro93, Ros86, Sch89, Sha83, STW90, Tau93, BBC88]. Beverley [Sch95b]. Beyond [Aga94b, Ald85, Cer89, Kni92a, Kuz87, Rob92, SG90, Twe90, FS84, Nea98, Tay96]. Beziehung [Koe91, Sch89]. Bhativedanta [Mur95b]. bi [Out94]. bi-centenaire [Out94]. Biagioli [Ili95b]. Bianchini [Alt91, Ger89, Ger89]. Bible [Ili98, Har98f]. biblical [Man93a]. Bibliographic [Woo95]. Bibliographical [Elm91, Sch91b, BW90]. Bibliographies [Ano82c, Jon91, Sch91b, NW82b, NW82a, Ano82b]. Bibliography [Ano82c, Ano82b, Bow92c, Byn90, Fau96b, Hud92, Jon91, Coo90, DeV82, Fru82, Le 90, Mul91, NW82b, NW82a, Ogi86, Por83, Rot93a, Sar87]. Biblioteca [Mor96a]. bicentenaire [BG^{+}92]. bicentenary [BG^{+}92, GS88, WBP97]. Bicentennial [Wil90a, Way87]. Biddell [Rot93b]. Big [Aga94a, GH92, Hoc93b, Aga94c, CW93, SN93]. Biggemann [HD97]. Bildung [Dau98, Fuc99]. Bill [Lar91, Nye92a]. Binghamton [Tha91]. bio [Fru82]. bio-bibliography [Fru82]. Biobibliography [WW86, Coo89b]. Biochemical [Tei93, Fru82, Fru90]. Biographical [DM96, Hor98, Hud92, Büh81, O’C88, Ogi86]. Biographies [Cro95]. Biography [Edg90b, Fie93, Har90a, Hen91b, Hoc92a, Kra92, La 97, Law98c, Ait85, Car89, FS88, FIF86, Kra90c, Lev85, Mor84c, Mul91, Pen89, Sha93b, Shi83, SY96, Spe82, TC90, Wal91, Wal90, Gro95a]. Biologen [Har94b, Dei92]. Biological [Bro91b, Mil89, Grm90, NLH^{+}84]. biologie [Bow95, CM93, MC^{+}93]. biologique [Grm90, Pyl91]. Biologist [Di 93e, Ver95, Wei91]. Biologists [Dei92]. Biology [Blo93a, Bow92b, Mil95, NLH^{+}84, Nyh95, Ras96a, Len82, Mai91, MC^{+}93, Pro86, Reh83, Rus88, Sap90]. biomedical [Ver85]. Biotechnology [Hor96a, Hor97b, Ken86, Rus91, Ver94, Ben90b, Bud93, THBB92b, Yox83]. Birch [Gun84]. Birefringent [Jam94c, Bar91]. Birkhauser [Jam94d, Bro95a, Koe91, Kra93a, Mar94a, Maz98b, Met96a, Mol94b, Whi90a]. Birnie [Sho93c, CHG91]. Birt [Jan98]. birth [AL87, JK84]. birthday [vLP82]. bis [Gol91a, Gre93b]. Bishop [Bow98]. Bitbol [Pyl92]. Bitbol-Hespériès [Pyl92]. Bjerknes [Kra90b, Fri89]. BJH [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h]. BJHS [Ano93r]. Blacker [Vig82]. Blackett [Nye99]. Blackwell [Bro92b, Cor93, Cro95, Har96c, Hoc92d, Hut92, Jam94b, Kni90b, Kus98, Law92d, Lee90, Ost93a, Smi91c, Smi93d, Swa92]. Blaise [Loe87]. Blake [Nye93b]. Blake-Coleman [Nye93b]. Blanchard [GG96]. Blaxter [Hor97a]. Blay [Hal95]. Blessings [Hor95, Seg94]. Blitz [Smi94d]. Bloch [Osb90b]. Blondel [Bry91, Mor96a]. blood [WKB88]. Bloomington [Di 93c, Hef96, Le 91, Spa91]. Blut [Wei90, WKB88]. Boag [Hug92c]. Boas [Ost93c]. Bodies [Mal94, Sin91, Sha93c, Dra81]. Bodson [Rus92]. Body [Law93g, Por97, Sta91, Arm83, RT92]. Boethius [Lor90, Van91]. Boethius-Texte [Van91]. Bohr [Hen92a, HP85, Kri91, PHvMW93, PvM85, Aas90, Hen84b, HP85, Kai94, MR82f, MR82g, PHvMW93, PvM85]. Bokhara [Ald85]. Bollati [Di 93a].

Bologna [Cre91b, Fin93b]. Bomb [Twe93c, Pou93, Smy89]. Bones [Twe93a, Bry92b]. Bonnet [And82]. Boockmann [Met93]. Book [Aga93, Aga94a, Aga94b, Aga95a, Aga95b, Aga96a, Aga97b, Aga97c, Aga98, Ait90, Alb92, Alt91, Aly90, Ams92, Ano92j, Ano95i, Ans96, Ari90, Bai91, Bai93, Bal91b, Bar92, Bar90a, Bar93b, Ben90a, Ben93, Ben94b, BV91, Bev90, Bis98, Blo93a, Blo98, Blo93b, Blo93c, Blo94b, Blu90, Boh93, Boh95, Bos92, Bow92a, Bow94a, Bow91a, Bow91b, Bow92c, Bow92b, Bow93d, Bow93a, Bow93b, Bow93c, Bow94b, Bow95, Bra90, Bra91, Bra92a, Bra92b, Bra95, Bra98, BD90, Bro90a, Bro93b, Bro93a, Bro94a, Bro95a, Bro98, Bro96a, Bro91a, Bro92b, Bro93c, Bro93d, Bro94b, Bro90c, Bro92c, Bro96b, Bro99c, Bro90e, Bro91b, Bro91c, Bro95d, Bru91, Bru93, Bru96, Bry91, Bry92a, Buc90, Buc92, Bug96, Bur93, Bur98, But92a, But94, Byn90, CK91, Cam90, Cam92, Can90a, Can90b, Can91a, Can96]. Book [Car95, Cas93, Cas94, Cha90c, Chu91, Cla92c, Cle92, Col93, Coo90, Coo96, Cor93, Cra92a, Cra92c, Cre91b, Cri90, Cro94, Cro95, Cro96b, Cro96c, Cro96d, Cuo95, Dan92, Dav94, Daw98a, Dea92, Dea93, Di 92, Di 93a, Di 93b, Di 93c, Di 93f, Di 93d, Di 93e, Di 95a, Di 98, Div90, Div98, Dol94, Dol98b, Dol99, Dol90a, Dol90b, Eck92, Edg90a, Edg90b, Edg91a, Edg95, Elm91, Elm97, Far94, Far99, Fau96b, Fie91, Fie92, Fie93, Fie94, Fin92, Fin93a, Fin93b, Fin93c, Fis98b, For90a, For91, Fox92, Fra90, Fre93, Fuc99, Ful98, Fyf99, Gau99, Ger91, Ger95, Get96, Gol91b, Gol93, Gol94, Gol95b, Goo90a, Goo96, Goo97, Goo91b, Gou91, Gou92, Gou95a, Gou98a, Gou98b, GG90, GG91b, GG91c, GG96, Gra91, Gre91, Gre95, Gre92c, Gre93a, Gre93b]. Book [Gre96, Gro93, Hac90d, Hac90e, Hac92a, Hac92b, Hac92c, Hac92d, Hac93a, Hac93b, Hac93c, Hac94, Hac95, Hak95b, Hal90d, Hal90a, Hal91a, Hal91b, Hal92c, Hal92a, Hal92d, Hal94a, Hal94c, Hal95, Har92a, Har98a, Har98b, Har98c, Har99b, Har99a, Har90a, Har93a, Har95a, Har96a, Har98d, Har90b, Har92b, Har92c, Har92d, Har93b, Har93c, Har95b, Har96c, Har96d, Har97, Har98g, Har94b, Har98h, Hay98a, Hay98b, Hef96, Hel96a, Hel96b, Hem95, Hen90a, Hen90b, Hen91b, Hen92a, Hen92b, Hen93a, Hen93b, Hen94a, Hen94b, Hen95a, Hen95b, Hen96, Hen98a, Hen98b, Hen99, Hes95a, Hes95b, Het95, Hig94, Hil91b, Hil91a, Hil91c, Hil93, Hil98a, Hin91, Hir95, Hoc90a, Hoc91, Hoc92b, Hoc92a, Hoc92c, Hoc92d, Hoc93a, Hoc93b, Hod90a, Hod90b, Hod91, Hod92a, Hod93a, Hod93b, Hod96, HS90, HS91, Hor92, Hor95]. Book [Hor96a, Hor96b, Hor97a, Hor97b, Hor98, Hor99a, Hor99b, Hud92, Hug92c, Hug92a, Hug92b, Hug96, Hug98, Hun92, HD97, Hun98, Hut92, Ili95b, Ili97, Ili98, Ink90, Jac96, Jac90, Jac93, Jam90, Jam91, Jam92a, Jam92b, Jam92c, Jam93a, Jam93b, Jam94a, Jam94b, Jam94c, Jam94d, Joh93, Joh96, Joh94, Jon86, Jon91, Jon92, Jon93, Jon94, Jor91, Jor94, Jor96, Kat90, Kel95, Kla95, Kni90a, Kni90b, Kni90c, Kni92a, Kni92b, Kni95, Kni98a, Kni98b, Koe91, Kra90a, Kra90b, Kra91a, Kra91b, Kra92, Kra93a, Kra93b, Kra94, Kra98, Kre96, Kri91, Kus94, Kus95b, Kus95c, Kus95d, Kus95e, Kus96a, Kus96b, Kus97, Kus98, La 94, La 97, Lan90, Lar90, Lar96b, Lar96a, Law90a, Law90b, Law90c, Law91b, Law91a, Law91c, Law92a, Law92b, Law92c, Law92d, Law92e, Law93g, Law93b, Law93c, Law93d, Law93e]. Book [Law93f, Law96b, Law96a, Law96c, Law98b, Law98a, Law98c, Le 91, Lee90, Lev94, Ley90, Lor90, Luc92, Luc93, Luc96b, Luc98, Lyn99, Mac90b, Mag96a, Mal93b, Mal93a, Mal94, Mal95a, Mal96, Mal98, Man93b, Man97, Mar92a, Mar93a, Mar93b, Mar93c, Mar94a, Mar96b, Mar96a, MK98, Mar94b, Mar96c, Mau95, Maz96, Maz98a, Maz98b, McC98, McC95, McE93, McE95, McK96, McK97b, McN92, Met93, Met96a, Met96b, Met99, Mic98, Mil95, Mil94, Mol92, Mol94a, Mol94b, Mol95, Mol96, Mol98a, Mol98b, Mon93, Moo92, Mor90a, Mor92a, Mor90d, Mor91c, Mor92c, Mor92d, Mor92f, Mor93, Mor95c, Mor96a, Mor98, Mor99, Mül93, Mul91, Mur92a, Mur92b, Mur95b, Mur96, Nea98, Neu92, New92, Nic93a, Nic93c, Nic93b, Nic94a, Nic94b, Nic96a, Nig99a, Nig99b, Nig99c, Non91b, Non91a]. Book [Non93a, Non94, Non93b, Nor94a, Nut93, Nye92a, Nye92b, Nye93b, O’B92, O’B96a, O’B96b, Odd91, Olb91a, Olb91b, Old97, Oli92, Ore90, Osb90a, Osb90b, Osl97, Osl98, Ost93b, Ost93c, Ost93d, Ost93a, Ost94, Ost98, Ost99, Out94, Pal90, Pal91a, Pal91b, Pal96a, Pal97, Pal98, Pan91a, Pau90, Per93, Pic92b, Pic93a, Pic98, Poo99, Por90b, Por90a, Por91a, Por91b, Por91c, Por95a, Por97, Pos90, Pri91, Pri93, Pum90, Pum91b, Pum91a, Pum92, Pum94, Pum95a, Pyl91, Pyl92, Pyl93a, Pyl93b, Rag95, Rap92, Ras96a, Ras96b, Rav99, Ree90, Rob93, Roh99b, Roh99a, Rou97, Rud92a, Rus91, Rus92, Rus94, Sch91a, Sch93c, Sch95b, Sch91b, Sec96, Sec98, Sha97, Sha92, Sha93a, Sha96, She99, She92, Sho93a, Sho93b, Sho93c, Sho96a, Sho99, Shr92, Sil90, Sim92, Sim94, Slo97]. Book [Smi90a, Smi91b, Smi91c, Smi91a, Smi92c, Smi92b, Smi92d, Smi92a, Smi93b, Smi93a, Smi93c, Smi93d, Smi94d, Smi94a, Smi95b, Smi96, Smi98, Smi99a, Smi99b, Soy96, Spa91, Spa94, Spa95, Spa98a, Spa98b, Sta90b, Ste91a, Ste95, Str90, Stu91, Stu98, Sut97, Swa92, Tei93, Tha91, Tim98, Tur93a, Tur93b, Tur91, Tur93c, Twe90, Twe91a, Twe91b, Twe91c, Twe92a, Twe92b, Twe92c, Twe93a, Twe93b, Twe93c, Uim95, Ung92, Van91, Ver90, Ver94, Ver95, Ver91, War91a, War90a, War91b, War95a, War95b, War96a, War96b, Wat91, Wea91, Wei90, Wei94, Whi97, Whi90a, Whi90c, Whi90d, WM85, Wil90a, Wil91c, Wil91b, Wil94a, Wil95a, Wil97, Wis96, Wit95a, Woo95, Woo98, Wri94a, Wri96b, Yav92, Yeo92, Yeo96a, Yeo96b, Yos99, vH92, Cla99a, Cla92a, Cla95a, OW10, Pat94, Ped92a]. Books [Aga98, AS89, Ano90g, Ano90h, Ano90i, Ano91i, Ano91j, Ano91k, Ano91l, Ano91m, Ano92k, Ano92l, Ano92m, Ano93i, Ano93j, Ano93k, Ano93l, Ano93m, Ano94i, Ano94j, Ano94k, Ano94l, Ano94m, Ano95j, Ano95k, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano97a, Ano97b, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano99a, Ano99b, Ano99c, Ano99d, Bry91, Har99a, Hen93a, Jam94a, Law93e, MK98, McK97b, OS92a, Smi92a, Soy96, Sut97, Got92, Mas98, CL87]. Boole [Mac85]. Boon [Ben94b]. Borders [Hir95, YK94]. Borelli [Mel98]. Boringhieri [Di 93a]. Borodin [FS88]. Bossi [Ste95]. Boston [Bow94b, Bro95a, Cas94, Hac93c, Hoc91, Jam94d, Koe91, Kra93a, Mar94a, Mar96c, Met96a, Mol94b, Nor94a, Rus91, Ste95, Tha91, Whi90a]. Botanic [Har93c]. Botanical [Sch91b, GEMM83, All86]. Botanik [Di 92, Jun89]. botanists [All86]. Botany [Dol98b, Dol99, Har98g, Hun92, Jun89, Hen86b, MR96, Sht96]. bound [Pai86]. boundaries [Hes99, Mal05]. Boussingault [McC84]. Bouvier [Woo95]. Bowdoin [Mur96]. Bowers [Jam92a]. Bowler [Boh95, Bro92b, Hod90b, Nic94b, Smi91c]. Boyd [Twe91d]. Boydell [Elm97, Joh94, McC98, Pum91b, Pum91a]. Boyle [Ans96, Hen94b, Ost98, Ost99, Ale85, BKH90, BHW94, Cla99b, Cle93, Dea95, Fig99, Hoo97, Hun90, Hun99b, Kah99, SSH85, Sha93b]. Boyling [Can99]. Bradwardine [Lor90, BM89]. Brahe [Fie93, Wil91b, Ros86, TC90]. Brain [Har90b, Law91b, Soy96, Har87, NC81, Smi92e, Sta89b]. Bramwell [Pal91a]. breeding [Rus86b]. Breger [HD97]. Brenner [Cle92]. Bret [Cro96c]. Bretagna [Pan91a]. Brewer [Far94]. Brian [Pic98, Hil91b, Jam92a, Mol95, Mon93, Twe91a]. Briarfield [Fau96b]. Bridgewater [Top92]. Bridgman [Hoc92a, Wal90]. Bridson [Sch91b]. Brief [May99]. Briefe [HD97, LM23, LS76]. Briefen [vLvHH^{+}88]. Briefwechsel [Alt91, Ger89, HP85, HD97, LHH91, PHvMW93, PvM85, LHB04, LS76, vL76]. bright [Moy86]. Brill [Fie92, Hun98, Mol95, Osl97, Rag95]. Bringing [HV82, Kai94, Pal96b]. Bristol [Bro99c, Cra92c, Hac92c, Hen94a, HD97, Law92b, Shr92]. Britain [Blu90, Bru91, Dol94, Elm97, Gol93, Jon92, Law92b, Law93b, Law93c, Mac90b, McC95, Nye92a, Por90b, Twe91a, Alt87, Arm83, BB92, Ber85, BR91, Bry88, Can83, Coo84, DH90, For98, Fox88, FL88, Gol92b, Goo90b, Ham90, H^{+}82, Hun95b, Luc90, May99, Maz92, McG84, OO89, Rus93, Ste92, TBDA10, Wel93, Wil94b, Wyn82, Yeo93]. Britannica [FC82]. British [All86, Bow93a, Bro91c, Bug96, Byn90, CK91, Fau96b, Hac93a, Har98d, Hoc92d, Lee90, Nic94a, Pal90, Sta90b, Ano97c, Ano97d, Ano98g, Ano99e, Bro97d, Cal90, Can94, Can97, Coo89a, Coo89b, Cre91a, CC94, FB82, Fis94, Gun84, Hen89, Hen94a, Hul99, Kuk91, Kuk93, Lie88, LLDM94, Mac88a, Mor84b, O’C88, Pat83b, Poc88, Reh83, Sto84, Top97, THBR92, WW75, WW86, WW93, Wil91a, Wil81, Wri94b, FC82]. Broadcasting [Hac94, Woo92]. Brock [McE93, Nig99c]. Brodine [Di 93c]. Broman [Tim98]. Brontë [Hay98b, Shu96, Shu04]. Brooke [Kni98a, Ost99, She99]. brothers [Gil83]. Brotons [Mal98]. Brown [Har99b, Kra91b, Lar96a, Nic93b, War91b, BN95]. Browne [La 97]. Brownian [Dim93, Mai90]. Bruce [Gre92c, Hoc93b, Bow91a, Bra95, Gre92c, Har93a, Hug96]. Bruna [GG91b]. Brundage [Por91a]. Brunner [Kus98].

Bruno [Kra91a, Kus95b, Ver90]. Brunswick [Dea93, Hen92b, Nic96a]. Brush [Har96a, Smi90a]. Bryan [Jam94a]. Bryant [Nye92b]. Bryden [But94, Twe93a]. Bryder [Por91b]. BSHS [Nic94a, May99]. Buchwald [Can90b, Goo97, Jam94c]. Bucke [Sho86]. Bud [Ben94b, Ver94]. Buffon [BG^{+}92, Out94, BG^{+}92]. builder [Sch92b]. building [Gil83, Wri87]. buildings [FG96]. burdens [Hol86]. Bureau [Mor92a, Cro91b]. Bürgerliche [Fuc99, Dau98]. Burkhardt [Hod90a, Hod96]. Burnby [Lan90]. Busard [Mol92, Mol94b]. Business [CK89, Gre96, Hod90a, Wil91c, Ash94, Lie88, Rei85, Smi94c, Yox83]. Butler [Ben94b]. Butters [Gou98a]. Bynum [Hen95b]. Byron [Lar96a, NN95].

C [Bro90c, Can94, Daw98a, Di 93f, Fie94, Fox92, Gou91, Gou92, GG93a, GG93b, Hac92b, Hal92a, Har96a, Hir95, Kra92, Lar96a, Law92c, MR89, Mor90e, NN95, Nut93, Nye92b, Nye93b, Osl98, Pum94, Roh99a, Sch93c, Sch95b, Slo97, Smi99b, Tur91, Twe91c, Wil91b, YK94, WD82, Car86, Ric91]. C. [Bry91, Gou91, Hac92b, Jor96, Out94, Rob96a, Sha98, Sil90, Ver95]. ca [Got92, Hen93a, Pum92]. CAB [War91a]. cabala [Bal91b, GGC90]. cable [Hun96]. Caesar [Sho99, RL97]. Cahan [Mor98]. calcul [Bea91, Ung92]. Calculating [Aga97b, Neb95, Twe93b, Ash94, MKW92]. calculation [Bea91, Ste99]. Calculations [Poo99, DR97]. Calendar [Far99, Hod96, BSK^{+}85, BSK^{+}94, Poo98]. Calendars [Cla95b, Kra98, Cla95a]. Calendrical [DR97, Poo99]. California [BV91, Cro96c, Gol91b, Hun92, Kni95, Mor90d, Por90b, Pum92]. Called [Daw98a, Hel96b, BEoCF92, ECL92, Mol94b, PTS^{+}96, vGA94]. Callick [Hac93a]. Calling [Yav92, Ole91]. Calvert [Cha93].

Cambrian [Sec86]. Cambridge [Aga93, Aga94a, Aga94b, Aga95b, Aga96a, Aga97c, Aga98, Aly90, Bar92, Bar93b, Bev90, Blo93a, Blo93c, Blo94b, Bos92, Bow92a, Bow91a, Bow93a, Bra92a, Bro90a, Bro93b, Bro91a, Bro99c, Bro90e, Bru96, Buc90, Buc92, But92a, CK91, Can96, Car95, Cha90c, Coo90, Dan92, Di 93b, Dol94, Dol98b, Dol99, Edg95, Fie93, For91, Fra90, Fre93, Ful98, Gau99, Ger91, Gol93, Gol94, GG91b, Gra91, Hal91a, Har92a, Har92b, Har96d, Har98h, Hay98b, Hem95, Hen90a, Hen90b, Hen91b, Hen94b, Hen95a, Hen96, Het95, Hig94, Hil91b, Hir95, Hoc90a, Hoc93a, Hod90a, Hod91, Hod96, HS90, Hor96a, Hor97a, Hor97b, Hud92, Hug92c, Hug92b, Hug98, Ili98, Ink90, Jac96, Jac93, Joh96, Jor91, Jor96, Kat90, Kla95, Kra90a, Kra91a, Kra91b, Kra93b, Kra94, Kri91, Kus95c, Kus96b, Kus97, Law93g]. Cambridge [Law93c, Law96a, Law98c, Lev94, Luc92, Mal93a, Mal94, Mal95a, Man93b, Man97, Mar93a, Mar96b, Mar94b, Met96b, Mic98, Mil95, Mon93, Mor95c, Mur92a, Nig99b, Nig99c, Nye92b, Olb91a, Olb91b, Ore90, Osb90a, Ost93c, Ost93d, Ost98, Pal91b, Pal98, Pau90, Pic93a, Pic98, Poo99, Por91c, Pri93, Pum90, Pum94, Rou97, Sch91a, Sha97, Sha93a, She92, Sim92, Slo97, Smi91a, Smi92b, Smi93b, Smi93a, Smi98, Smi99b, Ste91a, Tim98, Tur93c, Twe90, Twe91a, Twe91b, Twe91c, Twe92a, Twe92b, Twe92c, Twe93b, Ver94, War95b, War96b, Wat91, Wei94, Whi90d, Wil90a, Wil91b, Wil94a, Wil97, Yos99, Bec95, Ben97, CW87, Fyf97, GA98, Gou97, Hen84a, KBRS90, Ron83, SSK88, Sel10, YED97]. camera [HA87]. campaign [Gay98]. Campanella [Per93, Ern91]. Campbell [Aga98, Wil91c]. Campbell-Kelly [Aga98, Wil91c]. Campus [Har94b]. Can [Chu91]. Canada [Bra90, Hal92d]. Canadian [Hig94, Lev93]. Cancer [Law92c, Pal98, Fuj96, Ric91]. Caneva [Smi96]. Canguilhem [Olb91b, Sut97, CD94]. Canham [Hor99b]. Cannell [Smi94a]. Cannizzaro [Roc84]. Canon [Sir87]. Canton [Bro98, Dea92, Hal90a, Hal92d, Hes95b, Kni90c, Law91a, Luc96b, McE95, Pyl93b, Wit95a]. Cantor [Hal92a, Jam92b, Mor92f, She99]. Cape [La 97]. capitalist [Cou96]. Captain [BD90, DJS97, Wri87]. Carber [Har96a]. Cardano [Hen99, Sir97]. Carding [Ben94b]. Cardwell [Gre95, Har90a, Pic99]. Care [Aga95b, Gre92c, Hor99a, Bak96, Mor91b, Web88, Wie93]. Career [Pyl93b, Can94, She91]. careers [Dei92]. Carl [Jon93]. Carolina [Bis98]. Carozzi [Woo95]. Carrive [Osb90b]. cart [Wes95]. cart-wheels [Wes95]. Cartesian [Ait90, Cla92c, Cla89, Gro91, Hey83]. Cartography [Cam90, HW87, HW03]. Case [Hem95, Met96b, Des89b, Dim93, EC87, Hun99b, Kus95g, Mai90, Mei91, Pet94, Sch84, Sha93b, Wes95, Zan86]. caso [Cre91b, Des89b]. Casper [Can96]. Cassiano [Fin93c, FDLJ89]. Castagnola [Bal91b]. Castle [Wey92]. Cata [Fin93b]. Cata-logazione [Fin93b]. Catalana [Maz98a]. Catalog [Woo95]. catalogazione [Olm92]. cataloging [Olm92]. Catalogue [Elm91, Hen94b, Lar96b, AL87, Ber95, CL87, FC82, RWD90]. Categories [Hil93]. Cathedrals [For90a, SP88]. Catherine [Kus95b, Can90a, Coo96, Hil98a]. Cathode [Bro99c, Dah97]. Catholic [Di 98, Mal93b, Maz95a]. Caton [Jac90]. Catoptrica [Mol94b, EiT92, ET92, Tak92]. Cattenius [Hel96a, Kra91c, Kra91c]. Caufield [Can90a]. Causality [Osl97, vR95]. Cause [Dra81]. causis [OH85]. Cavendish [Kni98b, Kim95]. Cecilia [GH84, GH96, Smi98]. celebrating [AL87]. celeste [Non93b, GP92]. celestial [Cha90d, GP92, GB95b, Rot93b]. Cell [Di 93e, Wei91]. cellulaire [Chu91, Duc87]. cellular [Duc87]. celluloid [Fri83]. census [CHU87]. centenaire [Out94]. Centenary [Bow94b, Bro99c, Sar92, Spr97]. Center [Har98d, Hes95b, Lin93]. centers [Kri88]. central [Lan87]. centrale [Lan87]. centrality [Gra84]. Centre [Bry91, Ost99, Ano82a]. Centres [Wil90a, Ste99]. centring [CW93]. Centuries [Hac92b, Hun98, Kle94, Wil97, Cre91a, NW82a, Ano82b, Hey95, Poi04]. Century [Aga97b, BV91, Bro93a, Bro95d, Di 93b, Di 98, Elm97, For90a, Gau99, Gol91b, Gre92c, Ham90, Har98c, Har90b, Hen90b, Hig94, Hil91c, Hor99b, Hud92, Hug96, Hun92, Jam92b, Jam92c, Joh93, Jun89, Kni89, Lan90, Lar96a, Law92b, Law93b, Len82, Mar94a, Pal91a, Por90b, Por97, Rag95, Rud92a, Sch91b, Sho96a, Sim94, Woo95, And92, Ara91, BW90, CC94, Fis94, FM00, Ful91, G^{+}09, Gou97, GJ86, Hac90e, Har92e, Her88, Hes95b, Hui92, Jac95a, Kel98, KI92, Lev93, Luc96a, May97a, McC85, Mil99, Mor95a, NBR^{+}95, Ogi86, Pic90, Rud88, Sch95a, Sec94, Sha83, Wri96a, Ali86, And84, Ano04, Arm83, AW87, BB92, Bec91b, BaJANJ84, Bra89, Bry88, Buc85, Buc89, Bur99, BP02, Can90b, Can91b, Cha82, CJ87, Cle93, CM90, Coo84, Dau98, Dem83]. Century [Dol98a, Eng83, Fin93d, Fox87, Fox88, FHR90, Fre92, GA98, Gol87, Gra84, Gre82, Gre92b, Grm90, Hal84, Har82a, Har87, HW81, Hun95b, Kni86, Kni88, Kro91, Lev90, Lin96, Lin93, Mal05, Man88, Mas98, Maz95a, Mor91b, Moy83, Mur85, Neb95, PP95, Pic87, Pri83, RCRC98, Reh83, RT95, Roc88, Roc84, Ros91, Sch88, Sch89, Sch01, SP88, Sho86, SGS00, Sla82, Smi94b, Swa88, TBDA10, Tur83, Tur87, Wal81, WW88a, WR92, Whi90b, WD82]. Certain [MT83]. Certainty [Law91b, Mag96a, Mat95, Sha83, Sta89b]. Ceruzzi [Twe90]. Cesare [Ili97]. CH [Met93, Hac95]. Châtelet [ZH06]. Chadwick [Por91a, Bru88a]. Chadwyck [Hen94b]. Chadwyck-Healey [Hen94b].

Chain [Bos92, War90b]. chair [Ric97, AH93]. Chaldecott [Day99]. Challenge [Pri93, Deb91, Lie88]. Chalmers [Stu91]. Chalubinski [Ams92]. chamber [Cha97]. Chambers [Hen90b, Yeo96b, M^{+}89, Mil90]. Chambéry [Ano93q]. chambre [Cha84]. Chance [Aga93, Dan92, Ger95, Hod91, Hod92b, EM94, G^{+}89, Hac90a]. Chandos [War91b]. Chandra [Kra92, Wal91]. Chandrasekhar [Kra92, Wal91]. Change [Bos92, Bro96a, But92a, Div98, Elm97, Har98h, Kal95a, Osl97, Wri94a, AS96, BC90, BB84, Fox95, HL86, Hun95b, LeG88, Mar10, McC93, Moy83, RA91, Rus83b, Sch93b, War90b, Win93, vR95]. Changed [Hod91, G^{+}89]. Changes [PP87]. Changing [Ful98, Hor99b, McK96, Nye93b, Dol96, Gre82, RCRC98]. Channell [Pyl93a]. Chant [Edg90a]. chantiers [B^{+}95, Cro96d]. chaos [PS84a, PS84b]. Chapel [Bis98]. chapitres [BN82]. Chapman [Hac92a]. Chaps [Bar90b]. Chaptal [BV90]. Chapter [Rus91, Ben90b, Kel91]. character [TB97]. Chard [Edg90b, Har98d, Pen89]. Charismatic [Kus95e, Nol94]. Charlbury [Har99a]. Charles [Bro92b, Bro90e, Hod93b, Hod96, La 97, Smi92c, Twe91c, Twe92b, Twe93b, And82, BCKR89, BGH^{+}87, Bow90, Bow92c, Bro95c, BSK^{+}85, BS10, BSK^{+}94, Bur00, Can97, DK88, DDG90, De 96, GL88, Her91, Hil91a, Lin90, MK98, Spa91, Stu98, Vig82]. Charlotte [Hay98b, Shu96, Shu04]. Charts [BD90, DJS97]. check [WW75]. check-list [WW75]. Checklist [Hen93a, Got92, Law90a]. ’Chemical [FTG91, BV90, Bro93a, Cam92, Cro96c, Div90, Gre92c, Hor99b, Jam93b, Kni95, Kni98a, Mor91b, Nig99c, Par99, Pri93, Roc84, Tei93, Aft91, B^{+}95, Bro97a, Bro95b, Cle93, Coh96, Cre91a, Deb91, Don88, Fru90, GGSGB94, Huf82, KN88, Mor91a, NPYGDL86, NhTWL87, Nye93a, SG80, THBR92, TW87]. Chemie [Adl87, Bro94a, Bro95a, Fri91, Kle94, New92, Sch84, Sch92b]. Chemist [Col93, Mor91c, FS88, McC84, Rus86a, Sta84b, Tod83, Wil90c, HE92]. Chemistry [Bro93b, But92a, Cro96b, Gol91b, Hor99b, Kni95, Law93c, Mal94, McE93, McE95, Mor92c, Mor92d, NPYGDL86, Rob93, T^{+}85, THBB92a, Adl87, And84, Ber93a, Bro92a, Büt83, CHU87, Don91, Fri91, Gol92b, Hol85, Kle94, Kni92d, MF92, Nye93a, RCRC98, RA91, R^{+}85, Sch84, Sch92b, Sha93c, TT86, THBM92, TW87, KN88, NhTWL87, WH91]. Chernin [Kra94]. Chernyak [Law92a]. Chertok [Smi93c]. Chester [Mol94b, BEoCF92, ECL92]. chez [BH90, Pyl92]. Chicago [Aga95a, Ari90, Boh93, Bow92c, Bug96, Cam90, Can90b, Di 93d, Dol90a, Goo97, Gro93, Har98b, Hen90b, Hen98a, Hod92a, Ili95b, Jor94, Kra92, Law93b, Law93f, Mor92a, Mor98, Mur96, Nic93a, Nic93b, Nig99a, Nut93, O’B92, Ost94, Pum95a, Ras96a, Yeo96a, Yeo96b]. Childbirth [Law96b, Wil94c]. Chimica [Cle92]. chimique [BG80, B^{+}95, Cro96d, SG80, SG80]. China [But92a, Hen95a, Ano87c, Huf93, Huf95, KN88, NLH^{+}84, NR86, NhTWL87, yQ85, RA91, Wri96a]. Chloe [Har98d]. Chouet [Hey83]. CHP [Rus93]. CHP/DH [Rus93]. Chrisogono [BR96]. Christensen [Hac95]. Christiaan [Coh91]. Christian [O’B96a, Ost99, Hoo97, Smo94]. Christiane [Di 95a]. Christianson [Aga96a, Fie93]. Christie [Hac90d, Hal92a]. Christine [Ink90, Mor96a]. Christoph [Eck92]. Christopher [Blu90, Por90b, Ben82, Law92b, Por90b]. Chronicled [Fyf99, Dav99]. Chronicles [Can90a, Cau89]. chronologies [Wil87a]. Church [Moo92, Sha97, FC94, Fan96]. Chymiatria [Gre92c, Mor91b]. chymicall [CL87]. Ciència [Maz98a, Lan87, Oli92, OE^{+}90, LN95, Mal98]. Cientificas [War90a, Oli92]. Cimento [Mel98]. Circle [Gre92c, Hac90d, Hac92a, Ben87, Cha90d, Cha94, Gay96, Mor91a]. circuit [Jor90]. circumstance [Pet89]. Cirey [dG01]. Cité [Bry91, CM90, Cha84]. Cittadino [Pal91b]. City [Bro93d, Non91b, Woo98, Cha84, Jos97, CM90]. civ [Non93b]. civil [Dau98]. civilisation [Ano87c, KN88, NLH^{+}84, NR86, NhTWL87]. Civilization [Goo90a, Mei91]. Civilizations [Ano82c, Ano82b, NW82b, NW82a]. Civilizing [Har95b, Pye93]. Clagett [Bai91, Kra98]. Clara [Har98b]. Clarendon [Ait90, Cla92c, Hal94a, Hen93b, Lar90, Law93g, Ley90, Mal94, McN92, Mol98a, Str90, Whi90c, Wil91c, Yeo92]. clarifying [DD94]. Clark [Lar96a, Mur92b, She99, CMN95]. Clarke [Ait90]. Clásicos [Goo91b]. classes [Bec95]. Classic [Nic93b]. Classical [Das88, Hil91a, Mol98b, Sho99, BBC88, Har82b, Hil84, Sta90c]. Classics [Sul82b, Gje84]. classificazioni [La 94]. classify [Gol87]. Classique [Osb90b, BBC88]. classroom [Buc98]. Claude [GG96, Fin93c, SG80]. Claude-Louis [SG80]. Claw [Kla95]. Clayton [Hor92]. cleanliness [WKB88]. clergyman [Eng83]. Clerk [Buc92, Gol83, Har98e, Hen86a, Yos99]. Clermont [Fin93c]. Clermont-Ferrand [Fin93c]. Clifford [Hor92]. climate [Liv99, Mar10]. Clinical [AJ^{+}89, Hay98a, Law91a, Law91c, BP95, Büt83, PB99, Rob89]. Clio [Bra91, Hac93b, Spa98a, Chr93, Dij90, Kit89]. Clock [Hen99, Sir97]. Clockmaking [Bee89, Gou91]. Clocks [Kra98, Cla95a, Cla95b]. clos [Bla93, Hal95]. Close [LSS90, Shr92]. Closed [Aga98, Bla93, Edw96]. closure [EC87]. Cloth [Twe90, Whi90d]. cloud [Cha97].

Club [Bar90b]. Clulee [Mor90a]. Co [Pos90]. Coalfield [Twe91d]. Coast [Hac90e, Man88, HS99]. Coastal [BD90, DJS97]. Cockcroft [HA84]. coelestis [FC82]. coffee [Ste99]. coffee-houses [Ste99]. Cognitive [Dol90a, Por97, Tur91, Gie88, MH88]. Cohen [Hen92b, Hen98a, Sch93c]. Cold [Aga98, Hui92, Sim94, Edw96]. Colección [Goo91b]. Coleman [Nye93b]. Coleridge [Ley90, Wyl89]. Colette [Kni98b]. Colin [Ost93b, Bro95d, Edg90a]. collaboration [Bur98]. colleagues [AT87]. Collected [Bry92a, GF^{+}03, Hen94b, War96a, War96b, BH87, DD95, SCS^{+}87, War95a, KvDC^{+}37, LM23, LS76, vPSBD^{+}04]. Collection [Hac92c, Har95a, Jon86, McC81, O’C88]. Collections [But94, Hen94b, Ano91p, Ano92a, Ano98a, Cle83, HAB92, TW87]. College [Law90b, Rob96a, Sme97, Smi90a, Daw98b]. collegio [Wal84]. Colliding [Le 91, Ste90b]. Collie [Sho93c]. Collins [Bru91, Gol94, Twe91b]. Colloque [BG^{+}92, Out94, A^{+}87, Pic87]. Colloques [Rus92, vH92]. Cologne [Fie92, Mol95, Osl97]. Colonial [For90a, Har96d, LLDM94, Gro95b, Hen86b, Kum96, Moy86, SP88]. Colonialism [Hef96, Osb94]. Colonisation [Har92d, San91]. color [Bla83, Sep88]. Colored [Mal94, Sha93c]. colour [Cam90, Dol98a]. colours [Her83]. comburentibus [BL83]. comes [Mor84a]. Comet [Sho99, RL97]. comic [Ros94b]. Comité [Dav94, GG90]. commemorating [vLP82]. Commemoration [Di 93f, GT92]. Commentaries [Fyf99, Nic93b, Dav99]. Commentary [Lor90, Mil94, Mol94b, Smi92a, AT87, BM89, Can99, Di 93f, EiT92, ET92, Hac93b, HS91, OH85, Tak92]. commentated [ZH92]. commentées [Fin93c]. commerce [Ste99]. Commercial [Nig99b, Cat88, Pow96]. Commerical [Jac90]. Commission [Bry91]. Committee [Por91b, AB89]. Communicating [Sho93c, Jen91]. communication [Bec91b, Kel95]. communications [Pri83]. Communities [Aga97c, Doe96]. Community [Hun92, Nic93a, Sha96, Gol95a, Gos87, Huf82, Hul99, Mit92]. Companion [BP93b, Hal92a, Hen95b, O^{+}90, Smi99b, Sel10, Tra92, YED97]. Company [Hun92, CW87, FB82]. Comparative [Di 93f, Di 93d, Gli87, OS92b, Wil87d, Win91]. Compiled [Ams92, Bow92c]. Complete [Blo93c, Cas93, FMF85, Kit92, VB80]. completes [VB80]. completion [MR00]. complex [Ken86]. Components [Hac93a, Cal90]. Composite [McN92, FS91]. Composition [Gro89]. Computer [Aga98, CK91, CKA96, Twe90, Cer89, Hen89, Mor84a, NOF96]. Computers [Aga98, Edw96]. Computing [Bow92a, Sch93b, Twe91a, Twe92b, Twe93b, Asp90, OO89]. Comte [Pan91a]. Concept [Bra92b, A^{+}87, Bow95, CM93, McI85, MC^{+}93, Row90a, Wus84]. Conception [Cuo95, Had94]. Conceptions [Can96, Hak95a]. Concepts [Mor95a, CJ87, THBS92]. Conceptual [Bec91a, Cas94, Har82a, Hul88, Moy83]. conceptualisation [Bla83]. conceptualization [Bla83]. Concise [Hen90b, M^{+}89, Mil90]. conclusion [Ari90]. Conduct [Sha96, Gol95a]. Conductors [Nye93b, BC92]. Confer [Bra91, Kit89]. Conference [Ano87a, Hac95, Ver95, A^{+}87, BG^{+}92, CM90, Pic87]. confidence [Tom84]. Conflicts [Har94a]. Congrès [Ano93q, Dav94, vH92]. Congress [Ano93q]. conjunction [Lee90]. connaissances [Rus92]. Connection [Kle94]. connus [vPvP90]. Conquest [Hod90a, Que88]. Conrad [Bru96]. Consciousness [Nea98, Por90b, Fox88, Tay96]. Consejo [Oli92, War90a]. consent [Coo84].

Consequences [Kus94, Ric92a, Wol87].

Conservation

[Smi96, Sta90b, Can93a, Mac88a]. Conservative [Pal90, Mei91, Coo89a]. conservatives [Bec95]. consisting [Lar96a]. Console [Gol87]. Constructing [Dan90, FG96, Kus95c, Pic84]. Construction [Bev90, Boh93, Hug92b, Kra90b, Aga94c, Fri89, Mac90a, Ric92b, Rus89, WT82]. Constructive [Cha88]. consultancy [Twe91d]. Consultations [Law91a, AJ^{+}89]. consumer [Mor96b]. Consumo [Goo91b]. Consumption [BP93a, Far94]. contemporaries [Zam92]. Contemporary [Bec82, BV90, She92, Ner87, PW89]. Content [Non93a, Wal92]. contentious [ZH92]. contested [Liv93]. Context [Blo93b, Bro93a, Cro96b, Har92c, Mar94b, Nig99a, Nut93, Nye93b, War95b, Kai94, Kum91, Lin92, MW95, PT92, TR92, Ver85]. contexts [Hes99]. Continent [Gue81]. Continents [Le 91, Ore90, LeG88, Ste90b]. continued [Vic91]. Continuity [BB84, Mal93b, BA91, Kre82]. continuum [Sor83]. contours [Fin91]. contra [JK84, Sep88]. Contrasts [Fru90, Tei93]. contributed [Cre91a, CC94]. Contribution [Di 93e, BC90, Cou96, Dim93, Hen86b, Wei91, Wus84, Yat83]. Contributions [Rus92, Bro91a, Fie93, Gol91a]. Control [Mor95c, Ben86, Cro92, Yox83]. controverses [BD95, Mor96a]. Controversies [Sec98, BD95, EC87, Hal83b, Sho96b]. Controversy [Kre96, Mor93, Pum95a, Sec86, De 96, Fri91, Kra97, Kri82, Luc96a, Mos93, Per92, Pro93, Rud85, Sul82a, Tur94]. Convegno [Cle92]. Conway [Hen93b, CMNH92]. Cook [BD90, DJS97]. Cooke [Tur93a, McC92]. Cooper [Pal90, CL87]. cooperative [Swa88]. Coopersmith [Luc93]. Cooter [Byn90, Law90b]. Copenhagen [Fie92, Hal94c, Jam94c]. Copenhaver [Mon93]. Copernican [Fie91, Pum95a, Cro90b, GB95a, Mos93]. Copernicana [Fie92]. Copernicanism [Sha97, FC94, Fan96, Gra84]. copernicano [GB95a, Mal96]. Copernicus [Alt91, Fie92, Ger89, Ger89, Hen91a, Ros84a]. copies [Twe93a]. Copper [BC92, Nye93b]. Corbey [McK96]. Corbusier [Loa98]. cord [NC81]. Cornell [Bro91b, Can91a, Fie94, Fis98b, Har93a, Kra90b, Luc93, Yav92]. Corner [Gay98]. Cornwall [Kni98b]. corpis [Per89]. Corps [Neu92, Sin91, Fis98a, GG92]. Corpus [Hal90d, Mon93, Pri91, Cop92]. corpuscles [Ale85]. corrected [Sha97]. Correction [Ano90j, Ano91n, Ano93n, Ano96n, Ano99f]. Correia [Har98b]. Correspondance [BaJANJ84, LF97]. Correspondence [Ano04, Ano05, BaJANJ84, Cro96c, Di 95a, Hen93b, Hod96, HD97, Rus99, Sho93c, Smi92b, War96a, vBDGO93, BSK^{+}85, BS10, BSK^{+}94, Bur00, CHG91, CMNH92, FFLW97, GGSGB94, HRP90, HP85, LF97, LS76, Mor84b, Nap96, PHvMW93, STW90, vL76]. Corresponding [Sec94]. Corry [Maz98b]. cosas [Goo91b, ML89]. cose [Ara91, Fin93a]. Cosenza [Cle92]. Cosmic [Fur89, Osb90a, Van85]. cosmographicum [KA81]. cosmologists [Fur87]. Cosmology [BBBM90, Het95, Smi93b, Wil95a, Fie88, Nor94b, Pau93]. Cosmos [Hen96, Roh99a, vH97, Gra94]. Costabel [Gab90]. Cotes [Gow83]. Cotterell [Hil91b]. Coudert [Bro96b]. couleur [Bla83]. Coulomb [Mor96a, BD95]. Council [Por91b, Bro97d, Ano91s, Ano92p, Ano93s, Ano96q, Ano98g, Ano99e, AB89, Ben94a, SW93]. Count [Gou95a, ABR93, Ste99, Wey92]. Counter [Por97]. Counter-rational [Por97]. counting [Ger89]. Countries [Hac92b, Ano90k]. country [De 96]. country-house [De 96]. Countrymen [Mor90e]. County [LN95]. Courier [vPvP90]. cours [Lan87]. courses [Lan87]. Court [Gre92c, Luc96a, Daw98b, Mor91a]. Courtier [Ili95b, Bia93]. Cover [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h]. Covetous [Sch95b, Sou93]. Coyne [Ben93, Sha97]. Crabs [Spa98b, PR95]. Crabtree [Hak95b]. Cradle [Kel83]. Crafting [Fuj96, Pal98]. Craftsmen [Mal95a, FJ93]. Craib [Cra92a, Ada89]. Craig [Twe91c]. Crawford [Bro98, Hoc93a, Hod90a]. Crawford-Burkhardt [Hod90a]. Crayfishes [Spa98b, PR95]. Creation [Goo97, Kal95b, Lar96a, Maz96, Num95b, Yeo96b, Buc94, CS94, DR95, Hen84b, Pro93, Sor83, Blo82]. Creation-Evolution [Lar96a, Num95b]. Creationism [Bro94b, Lar96a, NBR^{+}95, RT95]. Creationist [Lar96a, NN95]. Creationists [Bro94b, Num92]. creative [Loa98]. creativity [Hol85]. credibility [Hun96]. Cremo [Mur95b]. crests [Jan98]. Crevier [McK97b]. criminal [Bur99]. crise [Gre86]. Crisis [Hoc92a, Osl97, Gre86, Wal90, vR95]. Crispini [Cle92]. Critical [Ano82c, Ano82b, Hac92a, Mil94, Mol94b, PW89, Sha97, She92, BL83, CF86, Cas93, Cha90d, Cha94, EiT92, ET92, Gle82, KRS83, Liu92, NW82b, OH85, Pum91c, Rot93a, SG80, Tak92, WW88b]. Criticism [Law93g, Jun89, Man93a, Sta91, WM85]. Critics [Jon92, Maz92]. Critique [Hun98, Smi93c, CS92, Hey95, Mei91, SG80, Spa94, SP98]. Crocker [Mul91]. Crombie [Gou92, Osl98]. Cromwell [Noa99]. Cronin [Bar93b]. Crookes [Gay96]. Crosland [Mor95c]. Crowcroft [Mor92a]. Crowds [Tur93c, vG92]. Crowe [Fie91]. Crump [Wat91]. Crusaders [Who82]. Crystal [Jam94c, Bar91]. crystallography [Sch89]. Crystals [ES90, Hoc92b]. CS [Bry92a]. CT [Kni98b, Sha96]. CT13 [Jam90]. CTHS [vH92]. Cult [Kus95e, Sha98, Smi99b, Nol94]. Cultivating [Har98g, Sht96]. cultivation [Pat83a]. Cultural [Blo93b, Bro93b, Di 93c, Hoc92a, Kat90, Law98b, Luc98, TBDA10, Bee96, Coo84, Gou95b, MF92, Pan91b, Rud88, TR92, Wal90]. Culture [Bur93, Dav94, Gre96, Ili95b, Joh93, Kus96b, Law93b, Law93c, Law98a, Mau95, Mol96, Mor98, Pum92, Roh99b, Rou97, Spa98a, Ash95, Baa90, BB92, Bia93, Bur87, CK90, Dau98, Gol92b, Gou97, Har96b, Høy94, Ili95c, KC95, Kro91, Kum96, Liv87, Mei91, Mil99, Mor96b, OS06, Ols90, Roc88, Smi94c, Sut95, You85, vHC95]. Cultures [JSS96, Slo97, For98, ME81]. Cumulative [Ano82c, Ano82b, NW82b, NW82a]. Cunningham [Har98a, Met99, Spa98a, Ste91a]. CUP [Hen90b]. Cure [Pic92b, Swa92, Wai90, Wea90]. Curie [Gir86]. Curiosities [Spa98b, PR95]. Curiosity [BS91, Jam92a, SL87]. Current [Luc96b, Mar93a, Sie91]. Curriculum [Pum91c]. Curtis [Bra98, Wil97]. curves [Eng84]. Cuvier [Lau87b, Out84]. cvii [Ans96]. cxv [Hac93c]. Cybernetics [Twe92c, WM85, Hei91]. cycles [Cou96]. cyclooctatetraene [THBM92]. Cynthia [Bev90].

D [Aga94a, Bra92a, But94, Fox92, Hug92c, Jam93b, Kra94, Lan90, Law93g, Mal93a, Mil95, Mol98b, Out94, Pyl91, Rou97, Sch91b]. D. [Smi94a, Twe93a, WHS92]. D.C [Hor99b, Sil90]. D.D [Gun84]. Dahl [Bro99c]. d’Ailly [O’B96a, Smo94]. daily [Gun84]. dal [FDLJ89, Fin93c]. D’alembert [Yeo96a]. dalesman [Spe82]. Dalitz [War96b]. Dalla [Mül93]. Dalton [Roc84]. Dame [Bro93d, Non91b, Sha97]. Dan [Hac95]. Daniel [Blu90, Bow93d, Gau99, Gro93, Mar93a, McK97b]. Danielle [Fin93c]. Danish [Hal94c, Jam94c, Ped92b]. dans [A^{+}87, dG01]. Danske [Hal94c]. Danziger [Kus95c]. Dario [Rap92]. dark [Ros84b]. Darkness [Whi90d]. Darnton [Mas98]. Darwin [Bar93b, Boh93, Bow93d, Bro92b, Bro93c, Bro90e, Di 93c, Hod93a, Hod93b, Hod96, La 97, Smi92d, BGH^{+}87, Bec92, Bow90, Bro95c, BSK^{+}85, BS10, BSK^{+}94, Bur00, CW87, Cro91a, DK88, DDG90, De 96, DM91, Ful91, Gay92, Her91, Lau87b, Pan91b, RR91, Ric92b, Rit90, Rob96b, Sec91, Wil91d, You85]. Darwinian [Hod90b, Pal91b, Bow83, Bow88, Cit90]. Darwinism [Boh95, Di 93e, Jon93, Bec92, Bow83, Bow93e, Deg91, Dur85, Jun89, MW95, Tor92, Wei91]. Darwinisme [Bow93d, Bow93b, Bec92, Tor92]. darwinismo [Mül93]. Darwinismus [Di 92, Jun89]. Darwinist [Cro84]. Daston [Hod91]. Data [Mol94a, EMMS93a, EMMS93b]. Dauben [Maz96]. Daughters [Har98g, Sht96]. Daum [Fuc99]. Davenport [Bru91, Wil94a]. Davenport-Hines [Bru91, Wil94a]. David [BD90, Cam90, Hen90b, Jam93b, Mor92f, Mor98, Mur92b, Cor93, Dol98b, Dol99, Jam91, Jam94b, Kat90, Kus97, Luc92, Mau95, McC98, Mor92c, Nut93, Osb90a, Ost93b, Pum94, Pyl93a, Rob93, Smi92b, Smi94d, SP98, War91a]. Davies [Old97]. Davis [Bro99c, Fyf99, Lar96a, Ost98]. Davy [Jam94b, Kni92c]. Dawkins [Twe91d]. Dawson [Smi99b]. Day [Har99a, Hor98, Sch91b, BW90]. DC [Mal93b]. De-centring [CW93]. Dead [Rud92a, Lev90]. Dean [Hen90a]. Dear [Sha92]. Dearth [Hor97a, BR95]. Death [Bru91, Ric87, BG^{+}92, DH90]. Debate [Dol94, Cor88b, Cro86, Dic82, Kai94, Liv99, Mas98, Wel93, Yeo93, Smi82]. Debates [Lar96a, Num95b, B^{+}95]. Débats [Cro96d, B^{+}95]. Debunkers [Mau95]. Debus [Pri93]. débuts [Lan87, Pic87]. decades [Bow83]. December [Ano92p, Ano96o, EJR94, EJR95, Mar96a]. deception [Mel98]. decisions [Pug84, Wyn82]. Decline [Jon93, Mol95, Deg91, Law93a, THBS92]. découverte [Six87]. Dee [Elm91, Har99a, Mor90a, Clu88, DF98, RWD90]. Deep [Law93f, McC82, Rud92b]. Deepak [Har92c, Har97]. defectives [Jac95a]. defence [JK84]. Defenders [Mau95]. defense [Gra84]. Deficiency [Lar91]. Defied [Pum92, Ols90]. Defining [Dol94, Yeo93]. Definition [McE95, Ber93a]. Degler [Jon93]. Deichmann [Har94b]. Deified [Pum92, Ols90]. Deist [Sul82a]. del [Cle92, Fin93b, Fre90b, GB95a, Goo91b, Lan86, Mal96, Olm92, Rey90]. Delahide [Mol94a]. Delaporte [Sut97]. Delaroche [Men90]. Delft [vLP82]. Delhi [Har92c, Har92d, Har93b, Har97]. della [Cle92, Di 93a, Fin93b, La 94, Mül93, Olm92]. Demainbray [Mor90e]. democrat [Sta84b]. Democritus [Dic82]. demon [Gol83]. Demonstration [Law98a, Sut95]. demonstrations [Wes95]. demonstrators [Pum95b]. Denis [Ost94]. Denmark [Hac95]. Department [McK96, Rob96a, Hul99]. Derek [Cha90c]. derivatives [Tra92]. Dershowitz [Poo99]. Desargues [DFG87]. Descarte [Gro93]. Descartes [Osl97, Pyl92, Pyl93b, BH90, Cla82, Gar92, She91, Sor87, Wil82, vR95]. Description [Roh99a, vH97]. Design [Aga93, Hem95, Pet94, For98, Hod92b, Pet89, Sch89]. Desmond [Bro93c, Goo96, Har93c, Hod92a, Ait90, Jac93]. d’espèce [A^{+}87]. Desperately [Ver93]. desperation [Ric97]. Dessi [Cre91b]. destitute [Ric87]. Determinism [Edg95, SM94]. Detlef [HD97]. Deutsch [Bru93]. deutsche [Dau98, Fuc99]. Deutschen [CM90]. Deutscher [Di 92]. Deutsches [Bro94a]. Deutschland [CM90, Di 92, Jun89, Sch93c, Wei90, WKB88]. deux [Gre86]. Developing [Bug96, Wri94b]. Development [For90a, Fyf99, Goo90a, Hac92a, Hac93a, Jac90, Mur92a, Twe93b, BEFH10, BB84, BBC88, BM84, BR91, Cal90, Cat88, Cha88, Cha90d, Cha94, CM90, Dav99, Har82a, HRP90, Hul88, MMVZ90, MKW92, MR82d, MR82e, NC81, Ost84, SO85, Sha93b, SP88, Smy89, Zan86]. developments [R^{+}85]. développement [BBC88]. Développment [Osb90b]. deviance [Jac95a]. devices [Ole84]. Devil [Goo96, Des94]. Devonian [Rud85]. DeWolf [Edg91a]. Dfl [Alb92, Ams92, Bow94b, Bro96b, Cas94, Hac92b, Hac93c, Hel96b, Hen98b, Hod93a, Hug92c, Hun98, Ili97, Jam92c, McK96, Mil94, Mul91, Non91a, Non93a, Nor94a, Rus91, Sho93a, Smi94d, Spa98a, Ste95, Wri96b]. DH [Rus93]. d’Histoire [Bry91, Dav94, Woo95, Ano93q, Cro96d, Fin93c, Rus92]. d’Historia [Maz98a]. diagram [How96]. Dialog [GB95a]. Diálogo [GB95a, Mal96]. Dialogue [Ano05, Sha97, AHY90, Dra81, Hen84b, HS91, PS84a, PS84b]. Dialogus [SSH85]. Diana [Olb91a]. Diane [Hod90a]. Diaries [Har99a, DF98]. ’Diary [Gun84, Bro90e, Hor92, DK88, PCA91]. Dictionary [Gro95a, Hen90b, Hor98, Hud92, DM96, M^{+}89, Mil90, O’C88, Ogi86, PB^{+}89, Gun84]. did [Pum95b, War90c]. Didactic [Gre92c, Mor91b]. Diderot [Yeo96a, Dul88, LSR95, OS06]. didn’t [Ste99]. Diego [Aga97b, Mur95b, Nic93c]. Diesel [Tho87]. Dietz [Pum95a]. Difference [Twe91c, Lin90]. Different [Kra93b, Gol92a]. differential [Gra85]. differentiation [Eng84, Mar92b]. Difficult [Mal93a, WHS92]. Difficultés [CMDM82]. difficulties [MR82f, MR82g, CMDM82]. diffusion [Noa99, Sch01]. Dijksterhuis [Hac93b]. Dijon [BG^{+}92, Out94]. Dilemma [Edg95, Twe93c, Pou93, SM94]. dilemmas [Hei86]. Dimensions [Kni92b, Van85]. d’influences [Ano90k, vH92]. Dioramas [Whi97, Won93]. Dirac [Hen91b, War96b, DD95, Kra90c, KW87]. Direct [Smi92a, AS89]. Directory [But94, HAB92]. Dirk [Alb92]. Disciple [Goo96, Des94]. disciplinary [Aga97c]. Discipline [Yav92, Goo91a, Mar92b, Ole91]. Disciplines [Kni95, Nye93a]. Discontinuities [Por97]. Discours [LSR95]. Discourse [Aga98, Yeo96a, vLMB88, Edw96, LSR95]. Discoverer [Aga94a, SN93]. Discovering [AT87]. Discovery [Bro99c, Fra90, Har93c, Mur96, Non93a, Pic92b, Bar89a, Bar89b, Bli82, DF97, Des92, MF86, MR82a, OP87, Pic90, PB82, PB83, Sch84, Sec91, Six87, Van93, Wal92, Wea90]. discussion [ST94]. Discussione [Cre91b]. discussioni [Des89b]. Disease [Law96c, Olb91a, Pal97, DPP88, EE95, Mic95]. Disenchanted [Sch88]. dish [Aga94c]. Disorders [Hay98a, BP95, PB99]. Displacement [Mar93a, Sie91]. display [For98]. disputata [Law93a, Mol95]. dispute [Sec86, ZH92]. Disputes [Tur93b, EC87, ZH92]. dissection [Ric87]. dissent [AL87]. dissolution [GB95b]. distance [Fie92]. distances [Hen91a]. distinguished [Cha83]. distributed [Sha97]. disturbed [Bro97b]. Divide [Mor90d, Sto88]. Divided [Hac90d, Ben87]. Dividing [Cha90d, Cha94, Hac92a]. divine [Gol91a, Jon88]. divinity [Dur85]. Division [Wri96b, HF87]. dix [Sch01]. dix-huitiéme [Sch01]. DKK [Hac95, Hal94c, Jam94c]. DM [Alt91, Bro95a, Di 93e, Eck92, Gre92c, Gre93a, Gre93b, Har94b, Hel96a, Hes95a, Hin91, HD97, Kus95d, Maz98b, Met93, Met96a, Mol92, New92, Pos90, Sch93c, Tur93b, Ver91]. DM148.00 [Fuc99]. DM48 [Lor90, Wei90]. DNA [Kri82]. Dobbs [Man93b]. Doctor [Twy88]. doctors [Pes84]. documentary [Ano87a, RR81]. Documentation [Sch91b, OP87]. Documents [Lar96a, NBR^{+}95, vPvP90]. Doel [Aga97c]. Does [Bro99a, Edg95, SM94]. dogma [Car88]. Dohrn [Di 95a, vBDGO93]. Dolby [Ful98]. Domenico [Hal94a]. Domesticating [Pur99]. domestication [Goo91a]. Dominic [Whi90c]. Dominis [BR96]. Don [Cra92a, War96a, Eng83, Ada89]. Donald [Pic99, Gre95, Har90a]. Donne [Slo85]. Donoghue [Ost94]. Donovan [Cro95]. Dordrecht [Ams92, Bow94b, Bro96b, Cas94, Div90, Dol90b, Hac93c, Hel96b, Hen98b, Hod93a, Jam91, Jam92c, Mar96c, Mil94, Mul91, Non91a, Non93a, Nor94a, Rus91, Sch95b, Smi94d, Ste95, Tur91]. Dorinda [Law96a]. Döring [HD97]. Dorn [Sho93b]. Dorothy [Smi95b]. Dörries [Mor96a]. Doskey [Por90a]. Double [Har90b, Har87]. doubts [Ili95c]. Douglas [Coo90, Mil94]. Dover [Fie91]. Dr [CHG91, Jam90, Sho93c]. Dr. [GG93a, GG93b]. Drake [Non91a, LS90]. Drakenstein [Hen86b]. drapery [Cha84]. drapière [Cha84]. Drawings [Bro91c, Nor89]. Dream [Blo93c, Bra90, Kit92]. Dresden [Kra91b]. Drifting [LeG88, Le 91, Ore90, Ste90b]. Dritte [HD97, LS76]. Drive [Edg95, SM94]. Drosophila [Pal96b]. Droz [Spa94]. Dualität [Koe91, Sch89]. duality [Sch89, Ste97a]. Dublin [Old97, Wil90a]. Ducal [Gou98a, But96]. Duchenne [Pal97, EE95]. Duchesneau [Chu91]. Due [SB92]. Duffin [Fox92, Hen92b]. Dugald [TB97]. Duhem [Mar96c, Jak84]. Duke [Bow94a, Kus96a]. duly [Cla97]. Dumont [Ver91]. d’un [SG80]. d’une [BV93, Gol95b]. Dunsink [Wil90a, Way87]. Dunstan [Wea91]. Durham [Bow94a, Kus96a]. during [Cro96c, For90a, GGSGB94, Hun92, SP88, Str87]. Duris [Spa94]. Dutch [Bur99, Dij90, Hen86b]. Dutton [Olb91a]. duty [Man93a]. dye [THBH92b]. Dyestuffs [Rus94, THBH92a, Tra93]. Dyke [Ost99]. Dynamical [Fra90, Bar89a]. Dynamics [Fra90, Kni95, Luc98, Bar89b, CD81, Gou95b, Nye93a]. Dystrophy [Pal97, EE95].

E. [Bow94a, Can94, Cra92a, Fyf99, Hac93a, Hac93b, Kra94, Whi90d]. e.V [Lil85]. Eagle [Bev90]. Early [BH87, Bro90c, Bro91b, CK91, Can90b, Coo96, Cuo95, Dol94, Far99, Gou92, Gre92c, Hen95a, Hun98, Joh93, Lar96a, Law93f, Luc98, Mal93b, Mal95a, Mar96a, O’B92, Por90b, Pri93, Pum91b, Pum92, Twe92a, BA91, Buc89, Can94, Can97, CMN95, Coh96, Col87, Cre91a, Cro90a, Deb91, Dol98a, Don91, EJR94, EJR95, Eis83, Eis93, FJ93, Fin91, Fox88, Gou95b, Had94, Hen89, Hey95, Huf93, Huf95, Hun96, Hun82, Hun89, Jac95a, Jac83, Jam92d, Kal95b, Kro91, MR89, Mil89, Mor91b, Mor84b, Ost84, Pan97, Phi87, Poc88, Poo98, PJP91, Reh83, Rud92b, Rus86a, Rus86b, Sch95a, Sec91, Sir90, Sor83, SCS^{+}87, Ste99, THBH92b, VoC89, Wil95b, Yeo93]. Earman [Gra91]. Earth [Fie92, Hug92b, Sho93a, Car88, Gle82, Nye99, Por83, Ros84b, Car88, Gra84, Mac90a, RR91]. Earth1 [Hen91a]. earthquake [Kel98]. East [MM10]. Easy [Mal94, Sha93c]. Échanges [Ano90k, vH92]. Eckert [Hoc92b]. Eclipse [Joh96, Bow83, Sta94]. Eco [Ost94]. École [Lan87]. Ecological [MS87, Nic93b]. Ecologists [Mor92a, Cro91b, Söd86]. ecologization [Söd86]. Ecology [Bra89, Nic93a, Nic93b, Nic96a, Pal91a, Pal91b, Cit90, Hag92, McI85, Mit92, Nic96b]. Econometric [Ger91, Mor90b]. Economic [BEFH10, Ger95, GG91b, Kla95, Luc98, Sch91a, Ben86, Cla92b, EM94, Gou95b, II90]. Economics [GG91c, Kla95, Por91c, MS87, Mir89, Sch90a]. economies [Cou96]. Economy [Bow94a, Sch91a, Mar92b, Wei92]. Ecosystem [Nic96a, Hag92]. ed [Ams92, Ans96, Bal91b, Bow94a, Bow91b, Bow93b, Bow94b, Bow95, Bra90, Bra91, Bro90a, Bro93a, Bro99c, Buc90, Buc92, Bur98, Cra92c, Di 95a, Di 98, Div90, Div98, Dol90b, Edg90a, Fre93, Fyf99, Ger95, Goo90a, Hac95, Hal90d, Hal90a, Hal91a, Hal92d, Har99a, Har95a, Har92c, Hen90b, Hen92b, Hen93b, Hen94b, Hes95a, Hes95b, Hoc91, Hor96a, Hor97b, HD97, Kla95, Kni90b, Kni90c, Kni92a, Kus96a, Kus97, Lar96a, Law90b, Law91a, Law92d, Luc98, Mul91, Nic94a, Nig99a, Nig99b, Non94, Por90b, Por90a, Rag95, Ras96b, Rou97, Sec96, Sec98, Sha92, Sil90, Smi90a, Smi91b, Smi92b, Smi95b, Spa95, Spa98b, Tha91, Tur91, Van91, War96b, Wea91]. Eddington [Jor96, Cha83, Kil94]. Edens [Har96d, Gro95b]. Edgar [Tre97]. Edgerton [Fie94]. Edgeworth [Ger95, EM94]. Edicions [Mal98]. edifice [Aga94c]. Edinburgh [Blo93b, Hen90b, Sec96, She99]. Edison [Mar93c, Mar96a, EJ^{+}07, ER93, EJR95, FIF86]. Edited [Daw98a, Hen93b, Met93, Yeo96b, Di 93f, Fie92, Hac93b, Hen92b, Hen94a, Hil93, Kni92b, Mal96, MK98, Mic98, Mil94, Mor98, Mur92a, Ost98, Rap92, Sch91a, Smi92c, Smi98, Sut97, Twe93b]. éditées [Fin93c]. Éditer [Dul88]. editing [Ano87a]. Edition [Aly90, Hen93b, Hil98a, Mil94, Mol94b, Ost94, Tur93b, ZH92, BL83, Blo98, Bow92c, CF86, Cas93, Daw98a, Di 93c, EiT92, ET92, Hen90b, Jam94a, Kus95c, Kus97, Lan90, Law93d, Mül93, OH85, SG80, Sch91a, Sha97, Smi98, Stu98, Tak92, Twe93a, Twe93c, Ver94, VoC89, WW88b, SG80]. Editions [Bow93d, Pyl91, vH92, Fin93c]. Editor [Lar96a, Wil91b, Ams92, Gun84, HD97, Ano93r]. Editore [Bal91b, Fin92, Cle92]. Editorial [Ano90l, Ano90m, Ano91o, Ano91p, Ano91q, Ano93o, Ano94n, Mal96]. editrice [Fin93b]. edn [Gou91]. eds [Aga95a, BV91, Blo93b, BD90, Bro93b, Cam90, Cle92, Cro96b, Cro96c, Dea92, Dol98b, Dol99, Edg95, Elm91, Far94, Gau99, Gou95a, GG96, Hac92b, Hac93c, Hal92a, Har99b, Har96a, Har98d, Har92b, Har96c, Har98h, Hay98a, Hen95b, Hil98a, Hoc93b, Hor92, Hor98, Hug92c, HD97, Jam92a, Jam93b, Kra91b, Kra93a, Law98c, Lee90, Luc96b, Lyn99, Mal93b, Mal93a, Mal95a, Mar92a, Mar93c, Mar94a, Mar96b, Mar96a, Mar94b, McK96, McN92, Mor96a, Nic93b, Non91a, Oli92, Osb90b, Ost93b, Ost93d, Out94, Pic98, Por91b, Pum91a, Pum94, Rus92, Sch93c, She92, Sho93a, Slo97, Smi93b, Smi99b, Spa98a, Ste91a, Ste95, Str90, Ver95, Ver91, War95b, War96a, Wil91b, Wil97]. Education [Dav94, FG93, Joh96, Mar96b, Mol95, And92, Buc98, Dau98, How82, May97b, Sta94, Sta90c, Top92, Wei82]. Educator [Kra93b, Gol92a, Rob96a]. Edvard [Lin90, Twe91c]. Edward [Aga96a, Bro99c, Bru93, Law93e, Ost98, Poo99, Rus86a, RR90, Wal88, Blo93b, Fis91b, Har99a, Hen96, She95, WBP97]. Edwardian [Law90c, Tro89]. Edwards [Aga98]. Edwin [Aga94a, Aga96a, Por91a, Bru88a, Chr95, SN93]. eeuw [Bur99]. Effect [Bro98, Cra96]. Effects [Goo97, Mil95, Buc94, MJ89, Smi94e]. efficiency [Mar92b]. Egan [Col93]. Egg [Har98b, Cor97]. Egyptian [Bai91, Kra98, Cla99a, Cla92a, Cla95a]. Ehrlich [Bäu84]. Eighteenth [BV91, Bro93a, Gol91b, Har98c, Hun98, Joh93, Lan90, Law93b, Por90b, Por97, Rud92a, Wil97, Kel98, May97a, McC85, Mil99, Mor95a, Sch95a, And84, Ano04, BB92, BP02, Fox87, Fox88, Hey95, Kro91, Lev90, Lin96, Mal05, Mas98, TBDA10, WW88a]. Eighteenth-Century [Gol91b, Har98c, Law93b, Kel98, And84, BB92, BP02, Fox88, Lin96]. Eilhard [Bro94a, Sch92b]. Einstein [HP85, PHvMW93, PvM85, War95a, War95b, War96a, Wol87, BH87, FD85, HP85, Kai94, Lit91, Liu92, MR82f, MR82g, Pai82, PHvMW93, PvM85, Pye85, Ros82, ST94, SCS^{+}87, Wol87]. Eire [But94, HAB92]. Eisendrath [Smi99b]. Eisenstaedt [Kra93a]. Elaine [Coo90]. Eleanor [Cla92b]. Elective [Sho93a]. Electric [Goo97, Buc94, FIF86, Wis85]. Electrical [Hac92d, Hac93a, Mar93b, Mor99, Nye93b, BC90, BC92, Tun92, BaJANJ84, Pur99]. Electricity [Mor93, Mor99, Nye92a, Bea97, H^{+}82, Luc90, Noa99, Per92, SW93, SW93]. Electrification [Luc93, Coo92]. Electrifying [Luc92, Nye90]. électriques [BaJANJ84]. électrodynamique [Blo82]. Electromagnetic [Mar93a, Buc85, Hen86a, OP87, Sie91]. Electron [Bro99c, Spr97, Dah97, DF97]. Electrons [Hoc92b, ES90]. elementis [BJ91, Mol92]. Elements [Ano92n, Mol94b, Ano92n, Cas93, BEoCF92, ECL92]. Elena [Oli92]. Elgin [Sho93c, CHG91]. Elisabeth [Bro98, Hoc93a]. Elise [Roh99a]. Elizabeth [Har99b, Har96a, Kni92a, Uim95]. Elizabethan [Par87]. Elizabethans [Bud98]. Ellis [Hac92a, Law93d]. eloquence [NS02, NS02]. Elsevier [Hug92c]. elsewhere [Hal92d]. Elspeth [Hil91c]. Elton [Mor92a, Cro91b]. Embryo [Wea91, Dun90]. Emergence [Fox92, Gas99, Hir83, Kim95, THBH92a, Wei83]. emergency [Fis94]. Emergent [Bli92, Smi94d]. Emerging [Kno90, Sho93a]. Emerson [Aga96a, She95, Twe92a]. Emery [Pal97]. Emigré [Har98h, AS96]. Emil [Tur93b]. Emilie [ZH06]. Emily [Cla92c]. eminent [Kah99]. Emmanuel [Cro96c]. Emmy [vdW85]. Emphasis [Mol95, Law93a]. Empire [Buc91, Dol98b, Dol99, GG90, Gre96, Har92c, Har96c, Hod91, HJF89, Kum91, Pal91b, Smi91a, Sta90b, Vuc84, G^{+}89, Mac88a, MR96, Sta89a, Cit90, HJF89, Smi94c]. empirical [Jan98, Tau93]. Empiricism [Hod93a, Wil91d]. empirisme [Duc82]. employing [Hes99]. enchanted [Roh99b]. Encounter [Law98b, Bee96]. Encounters [Shr92, LSS90]. Encyclopaedia [Hen95b, Smi91b]. Encyclopedia [Bur98, RM96, Rei94, Yeo96a, BP93b, LSR95]. Encyclopédie [LSR95]. end [Gal87, PP95, Slo85]. Endeavour [BD90]. Ended [Ano96o]. Endless [Sim92, Her90]. ends [Red87, Sch95a]. enduring [Gje84, Kol92]. Energy [Bra92b, Buc91, Edg91a, Har82a, Luc96b, Sim92, Smi96, Can93a, Her90, MS87, Row90a, Rus93, Smy89]. Engagement [She99, BC98]. engend’ring [Rus86b]. Engineer [Kra93b, Gol92a, Vig82]. Engineering [Bug96, Div90, Hem95, Mar92b, Coh96, Hil84, Pet94, Wei82, Wri94b]. Engineers [Gre86, Hac92d, Hac93a, H^{+}82, Mar93b, Mor99, Bar93a, Hun96]. Engines [HL86, Twe91c, Lin90]. England [Bar90a, Cri90, Far99, Har92a, Har98g, Hod90a, Jam90, Law96b, Luc98, Ost93d, Por91a, Bru88a, Cle93, Con87, Gou95b, Hey81, How82, KK87, Opp85, Par87, Poo98, Por89, Pum95b, Ric88, Rus86b, Sha83, Sht96, Ste91b, Wil94c]. English [Ink90, Lor90, Mol94b, Mon93, Ost99, War96a, BL83, BM89, Cop92, Don91, Dra81, EiT92, ET92, Mac88b, Tak92, Vic87]. Engraving [Lan87]. Enigma [Lan82, Nor94a, Zam92]. Enlightenment [Ano07, Bro93b, Joh96, Law93g, Law93c, Law96a, Law98a, McE95, Ber93a, Gol92b, Sta94, Sut95, Das88, Han85, Hey83, Jac83, KK87, LL09, LP04, MF92, Out95, Sta91, ZH06]. Enough [Gra91, Ear89]. Enquiries [Har96a, BGH^{+}87, MGBE95]. Enquiry [Law92d, Ost98, Ben91, BDH96]. Enschede [Alb92]. enseignement [GG90, HJF89]. Entangled [Nic96a, Hag92]. Entdeckung [Sch84]. Enterprise [Gol91b, Liv93]. Enters [Gre92c, Twe90, Cer89, Mor91b]. Entertainment [Joh96, Sta94]. Enthusiasm [Hun98, Hey95]. entomologist [Cla92b]. Entre [BBC88, Osb90b, Ano90k, BaJANJ84, Gre86, vH92]. Entwicklung [CM90, Eck92, HRP90, Sch93c]. Environment [Nye92a, Luc90, MS87]. Environmental [Har96d, Nic94a, Nic94b, Bow92d, Sho93e]. Environmentalism [Har96d, Gro95b]. environments [FG96]. enzymes [Tei91]. epic [NPYGDL86]. Epicurean [Fre93, Osl91]. Epicurus [LL09]. epidemiology [Col87]. episode [Liu92]. episodes [Het88, Liv93]. Episteme [Smi94d]. epistemological [Des89b]. epistemologiche [Cre91b, Des89b]. epistémologie [GG96, PP95]. epistemology [PP95]. Epistolario [Rap92, VG98]. epistolieres [SGS00, SGS00]. equations [Gra85, MR82c, MR82e, MR87b]. equilibrio [Mül93]. Equilibrium [GG91b, II90]. Equivalence [Ber93b, Hal94a]. Era [Bro93b, Pum92, Fri81, MF92]. Erasmus [Ili97, Jam94c]. eretico [Red83]. Erhard [Koe91]. Eric [HD97, Bro91b, Dun92a, Dun92b]. Erich [Het95]. Eriksson [Wit95a]. Ernst [Di 95a, Per93, Twe93b]. Error [Hem95, Men90, Pet94]. ERS [Tur93b]. Erwin [MR87a]. Espaäna [Rey90, Lan86]. España [Goo91b, Lan89]. Españoles [Goo91b]. Essai [BG80, Neu92, Sin91]. Essay [For91, Sch91a, BEFH10, Edg91a, Law93d, MJ89, Red87, Sin91, Vos97, Wil90b, HD97]. Essays [Aly90, Bow91b, Bra92b, Goo90a, Gou95a, Har92c, Kni92a, Law98c, LS90, Mal93b, Mal93a, Mor98, Nic94a, Non91b, Non91a, Osb90a, PP87, Por91b, Rou97, She92, TT86, Wal81, Wes89, ABR93, AB89, BA91, Daw98a, Don88, Dur85, Fur89, GJ85, GR88, Hol86, JK84, Kum91, MH88, PW89, SG90, Sho93e, SY96, Sto83, Sto84, VoC89, WHS92, WBP97, vHC95]. Establishing [Hun89, Pum91b]. establishment [Ben97]. establishments [EMW82]. Estrange [ABR93, Bry92a, Gou95a]. Estudios [Goo91b, Rey90, Lan89]. età [Fin93b]. établis [Non93b]. Ethics [Fis98b, WM85, SG96]. ethnographic [Sto83]. Études [Fin93c, Hal90d]. Eucharist [Bow98]. Euclid [Mol94a, Mol94b, Mol98b, BEoCF92, EiT92, EMMS93a, EBT96, ECL92, ET92, EMMS93b, Tak92]. Eugene [Nea98, Pal91b]. eugenic [Raf88]. Eugenics [Jon92, Maz92, SP98, WKB88, Lar91, Maz92]. Eugenik [Wei90, WKB88]. Euler [Can96, Hak95a]. Europe [Ben94b, Cam90, Cuo95, Hil98a, Jam92a, Mal95a, Mar96b, Ost93b, Por97, Rus94, Ste95, Whi90a, Wri96b, BS91, C^{+}92, Eis83, Eis93, FJ93, FG93, GR91, GJ86, Had94, HW87, KRS94, Mar10, PB94, Tra93, vPvP90]. European [Cro96b, Fre93, Hac95, Har93c, Law93g, Law96a, Nut93, Por90a, Wea91, Wri96b, Ano90k, Des92, Dun90, Lin92, MD88, MM10, OS06, Osl91, RT92, Tyt82]. Européens [vH92, Ano90k]. Evaluative [McK96]. Evan [Bro93b]. Eve [Fie94, Har98b, Edg91b, Cor97]. Evelleen [Law92c]. even [Cla92b]. ever [Tor95]. Everitt [Har96a]. Everyday [Edg90a, Hod91, BC89, CB^{+}90, G^{+}89]. Everyone [Gol94, CP93]. evidence [Nye99]. evils [MJ89]. Evolution [Ben90a, Boh93, Bow98, Bro94b, Di 98, Hod92a, HS90, Lar96a, Smi94d, Bas88, Bli92, Bow83, CD81, Cha88, Des89a, Dur85, Kal95b, Lau87b, Maz95a, Num95b, Ray87, Ric92b, Rob96b, Whi88, Bro84, Lau87b]. Evolutionary [Bow94b, Jon94, Tau93, Yeo96b, CS94, Cor88a, Hul88, Pit93, Pro86, Sar92, Ver93]. evolutionism [Bec92]. Evolutionnisme [Bow93d, Bec92]. Evolving [Sal85]. Exact [Har95b, Mal93a, Pye93, WHS92]. exakten [Van91]. examining [WT82]. Exchange [Cuo95, Ger89, Had94, Ili95c]. Exchanges [Ano90k]. Exeter [Wea91]. Exhibiting [Bea97, Mor99]. Exhibition [Lar96b, AL87, Ber95]. Exotic [Hef96, Osb94]. Expand [Kra94, TFC93]. expanded [Di 93c]. expanding [Smi82]. Expansion [Har95b, Har96d, Gro95b, Pye93]. expectations [LLDM94]. Expenditure [Ano96o]. Experience [Har92d, Pum91b, Smi99a, Gra98, Hun89, San91, NS02]. Experiment [Can90b, Goo90a, Goo90c, Hut92, Jam91, LS90, Non91a, Ost93c, AH85, Buc89, Cha97, Dra81, Fra86, Hal90c, Hal91c, NS02, SW95]. Experimental [Bry92a, Law92e, Ost93c, Hal91c, Kre90, Men90, NC81, SSH85, Vos97, Pum95b]. Experiments [Bar91, Fri91, Hel96b, Jam94c, New92, De 96, Gal87, vGA94]. expert [Lar91]. Experts [Twe91b]. Explaining [Dol90a, Gie88, Rus93]. Exploration [Hig94, Smi91a, Bur94, Hol85, Lev93, Sta89a]. explorer [Ald85]. Exploring [Huf91, Hug92a]. explosive [Mei91]. Exposures [Can90a, Cau89]. expulsion [Dei92]. external [Ale85]. extra [Law90c, Tro89]. extra-ordinary [Law90c, Tro89]. Extraordinary [Mor91c, Wil90c]. extraterrestrial [Cro86, Dic82]. extreme [HF87]. Eye [Kre96, Ash94, Tur94].

F [Bro99c, Bry91, Cle92, Coo90, Cro96d, Edg91a, Gou91, Gou95a, Har95a, Har96a, Hen95b, Joh93, Mag96b, Mar96c, Pyl93a, Twe92b, Neu92]. F. [Cle92, Non94]. F.Fr [GG90]. F.R.S [Wal88]. F.R.S. [Gun84]. Faber [Odd91]. Fabri [Fin93c]. Fabric [Law93g, RT92, Sta85]. Fabrication [Stu91, Cha90a, Cha90b]. Faces [Man93b, Dob91, Ful91, Tyt82]. facsimile [Jam94c, Law93d]. Facsimilies [Bra92a]. Fact [Sho96a, Smi94b]. Faculteit [Alb92]. failed [Gay98, THBvdB92]. Failings [Jon92, Maz92]. Failure [CK91, Twe91c, Hen89, Lin90, SP98]. Fairfield [Bow93c, Rai91]. Faith [Ost99]. Falconer [Bro99c]. fall [THBvdB92]. Fallstudie [Eck92, Mei91, Sch84]. Fälschung [Wol87]. Families [Eng84]. family [Raf88, SL87]. Fantoli [Sha97]. Faraday [Jam92a, Jam92b, Jam93b, Mor92f, BS91, CGJ91, Can91b, FTG91, GJ85, Goo94, Jam92d]. Farben [Mor90d, Sto88]. farmer [Dem83]. Farrar [Aga96a]. fast [Adl87]. Fat [Out91]. Father [CMDM82]. Faulhaber [Met96a, Sch93b, Sch93b]. Faust [Rou97, Hay94]. Fear [Hoc90a, Wea88]. Febrer [Goo91b]. Fechners [Hei93, Kus95d, Hei93]. Fecondità [Di 93a]. Fee [Har99b]. Feeble [Bis98, Nol95]. Feeble-Minded [Bis98, Nol95]. Feeding [Aga95b, Wie93]. Feeling [Sho96a, Adl87, Smi94b]. Feingold [Hal91a]. Feldhay [Sha97]. Fellows [Hun82]. Femmes [SGS00]. Fenton [Har99a, Pal90]. Ferdinand [Bro91c, Nor89]. Ferdinando [Cro96b]. Fermat [SO85]. Ferran [Maz98a, Mal95b]. Ferrand [Fin93c]. Ferrier [Mol94a]. Festival [For98]. Festivals [For98]. fever [Col87]. FF [Bow93b, Dav94, Bow95, Fin93c, Gol95b, GG96, Hal95, Kel95, Non93b, Osb90b]. Fiasco [Sim94, Hui92]. Fiction [Shr92, LSS90]. Fictions [Smi99b, Sha98]. Field [Mal95a, Rav99, Hen86a]. Fields [Law98b, Bee96]. fieldwork [Sto83]. fifteen [H^{+}82]. fifteenth [GJ86]. fifteenth-and [GJ86]. fifty [All86, Hea86, Hea90]. Film [Law90a]. Fin [Har92b, Lar90, TP90, GG96, Har94c, PP95, Har89]. final [DD94]. Finn [Kri91]. Finnet [Bro95d]. Finnish [Bro96a]. Fire [Aga96a, Man97, CF86, New94, She95].

Firenze [La 94]. First [HD97, Mor92c, Sha97, Bow92c, Coh84, FFLW97, Grm90, H^{+}82, Olm92, Poi04, Wal88, WH91]. Fisch [McN92, Yeo92]. Fischer [Di 93e, Ver91]. Fisher [Bru93]. Fisherian [Har94a]. Fisheries [Mil95, Smi94e]. Fishes [PR95, Spa98b]. Fishing [Mil95, Smi94e]. fitness [Who82]. Fits [Mal94, Sha93c]. Flammarion [Gol95b]. Flamsteed [HD97, FC82, FFLW97]. Flash [Bro99c, Dah97]. Fleck [Ams92]. Fleeming [Mor94]. Fleetwood [Noa99]. Fliess [FMF85]. Flight [Twe90, Cer89, Har85]. Flora [Har93c, Har98g, Des92, Sht96]. Florence [Bal91b, Gou98a, Mor96a, But96]. Florey [Wil84]. Florida [Jac90]. Floris [Hen98a]. Flourens [NC81]. flow [Jor90]. Floyd [Wri87]. Fogg [Lev94]. Foley [Hel96b]. Folge [Bro94a]. Folgen [Wol87]. Folkerts [Mol94b]. Folkestone [Law93e]. folklore [Llo83]. fondamentaux [BN82]. Fondamenti [Cle92]. Fondements [Atr86]. font [Lil85]. Fontana [Blo98, Gre95, McE93, Nic94b, Wil95a, Bow92d, Bro92a, Car94, Nor94b, Smi97]. Food [Hor97a, BR95]. Forbes [HD97]. Forbidden [CT93, Mur95b]. Force [Bra98, De 95, Adl87, Har82a, Mac83]. Forced [AS96, Har98h]. forces [Blo94a, Pai86]. Foreword [Ili97, Mar96c, Ost99, Edg91a, Jor94, Ost94]. Form [Mol98a, Ras96a, BBC88, Bur87, Eme84, Fre95, Fur88, Nyh95]. Formation [Osb90b, BBC88, Huf82, Lie87]. Forme [Osb90b, BBC88]. Formed [NW82b, NW82a, Ano82c, Ano82b]. Formulae [Alb92, Ste91c]. formulation [MR82b]. Forschung [CM90, Sch93c, Dei92]. Forschungsförderung [Har94b]. fortuna [Pan91a]. fortune [dCC09]. fortunes [Tyt82]. Fossil [Spa91, Ste90a]. fossilist [Tor95]. Foucault [Alb92, Jor91, Gut89, Ste91c]. Foundation [Ano97d, Cas93, Hor99b, Kle94, MR82f, MR82g]. Foundations [Nic93b, Tur91, Lau87a]. founder [Old94]. Founders [Bow94b, O’C88, Sar92]. Founding [Smi99b, Sha98]. Four [Hoc93a, PTS^{+}96, Cra92b, Daw98a]. Fourcroy [BV90]. Fourteenth [Rag95, Fre92]. Fourteenth-Century [Rag95, Fre92]. Fox [Blu90, Div98, Mar96b, Por90b]. Fr.F [Hal90d]. fra [La 94]. Fragment [Jam90, Twy88]. France [Bow93b, Cro96d, Lau87b, McC95, Pes84, Pic93a, Pri93, Spa94, Ver90, And84, Ano82a, CM90, Cor88a, Cro96d, Deb91, KK87, Lat88, Lau87b, Out84, Pes84, Poi04, Pri83, Ram88, Sch01, TBDA10, Wei83, Wil94b, dCC09, dG01]. Frances [Joh94]. Francesco [Bal91b, Fin93a, Law91a, Ara91, GGC90]. Francis [Bro99c, Fyf99, HD97, Ree90, Sha93a, BM84, Bur94, Mar92c, PR88, PR90]. Francisco [Goo91b, Twe92b, Twe93b, Ara91, Coo90, Fre90b, GGC90]. Franco [Pan91a, Per93, Rap92, Fin93a]. Francois [Fin93c, Sut97, Bra98, Cas93, Chu91]. Francois-Marie [Cas93]. Frängsmyr [BV91, Bro93b, Hal92d, Kni90c, Spa95]. Frank [Lar96a, Mal95a, Mar92a, Mor92f, CMN95, Goo90a]. Frankfurt [Har94b, Kus95d, Wei90]. Frankfurt-on-Main [Kus95d]. Frankland [Rus86a, RR90]. Frankreich [Sch93c, CM90]. Frans [Sho93a]. Franz [Alt91, Blo93a, Gre92c, Gre93b, Hel96a, Hin91, Mol92, New92, Van91, Ver91, Sap90]. Frazer [Ack87]. Fred [Cam92, Hod93a]. Frederic [Aga95a, Gol91b]. Frederick [Law90c, Tro89, Di 93b, Hod96]. Frederik [Aga97b]. Free [Ost98, Soy96, BDH96, Hen94b]. Freedom [Fis98b, SG96]. French [Bla83, Get96, Har95b, Hef96, Mar96c, Met99, Mor95c, Pri93, AHY90, Ano87b, Ano90k, Ano92n, Ano93q, Atr86, A^{+}87, Bea91, BG^{+}92, Bec91b, Bec92, BaJANJ84, BV93, BH90, Bla93, BBC88, BD95, B^{+}95, CMDM82, Cha84, Cro92, Deb91, Duc87, FDLJ89, FM00, GP92, G^{+}09, Gay92, Gol87, Gre86, Grm90, HJF89, Lan87, Lau87b, LF97, MC^{+}93, NS02, Osb94, PP95, Pes84, Pic87, Poi04, Pye93, Roc88, SG80, Sch01, SGS00, Sin91, Six87, Tor92, Wei82, dCC09, dGB^{+}89, dG01, vPvP90]. Frenkel [Kra94]. Fresquet [Goo91b]. Freud [Blo93c, Hak95b, Kus98, Smi92d, Bru95, Cra93, FMF85, Gos87, Kit92, MF86, Rit90]. Freudenthal [Hal90d, Mol98a, Rag95]. Freudian [Kah99]. Friar [Met99, FC96]. Friederike [Met93]. Friedman [Kra90b]. Friedmann [Kra94, TFC93]. Friedrich [Eck92, Mei91, Met93]. Friends [Ans96, Hen93b, CMNH92]. Fritscher [New92]. Frog [Mor93, Per92]. Front [Ano90b, Ano90d, Ano90f, Ano91b, Ano91d, Ano91f, Ano91h, Ano92c, Ano92e, Ano92g, Ano92i, Ano93b, Ano93d, Ano93f, Ano93h, Ano94b, Ano94d, Ano94f, Ano94h, Ano95b, Ano95d, Ano95f, Ano95h, Ano96b, Ano96d, Ano96f, Ano96h]. Frontiers [Di 93c, Pan91b]. Fruton [Tei93]. Fry [Ash85]. Fryer [Wri96a]. Fuchs [Wil87b]. Fuchsine [THBvdB92]. Fujimura [Pal98]. Fukuolca [Mol94b]. Functionalism [Sto84]. Fund [Bec99]. Fundament [Bro94a, Sch92b]. Fundamental [Jor96, Kil94, MR82c, MR82e]. fundamentals [MR87b]. Funding [Hun92, Str87]. Funeral [Sho99, RL97]. Furley [Osb90a]. Fusion [Her90, Sim92, Sim94, Hui92]. Future [Hor95, Seg94, Bro99a, MJ89].

G [Ack87, Bow91a, But94, Elm91, Ful98, Gou95a, Hac90e, Hal92a, Har96a, Har98h, Hen94a, Hen99, Hug98, HD97, Kus96b, Lan90, Mor90d, Mor99, O’B92, Pri93, Sha98, Smi90a, Smi99b, Sto88]. G. [Bar92, BT96, Ben90a, Hac92c, Hug98, Law93b, Lev94, Str90, Wea91]. Gabriel [Smi92b]. Gad [Hal90d, Mol98a, Rag95]. Gainesville [Jac90]. Galaxy [Het95, Pau93]. Gale [Aga96a]. Galilean [Dra81]. Galiléenne [Non94, dGB^{+}89]. Galilei [Ari90, Mal96, Non93b]. Galileiani [Sha97]. Galileo [Dea93, Hug92a, Ili95b, Ili97, Non91b, Non91a, Non93a, Sha97, Ari90, Bia93, Dra81, FC94, Fan96, Huf91, LS90, MW94, Mal96, Non93b, Red83, Seg91, SB09, Wal81, Wal84, Wal92, Wes89]. Galison [Hoc93b]. Gallic [dGB^{+}89]. Gallimard [Hal95]. Galton [Bur94]. Galvani [Mor93, Per92]. Galvani-Volta [Mor93]. Game [Bow94a, Twe93c, Pou93, Wei92]. Games [Sho99, RL97]. Gandt [Bra98, Non94]. Gaposchkin [Smi98, GH84, GH96]. Garber [Gau99, Gro93, Kni92a]. Gardens [Har93c]. Garland [Hal92c, Har95a, Hen93a, Hod93b, Lar96a, Law92e, Mol98b]. Gary [Jor91, Smi93a]. gas [Wil81]. Gassendi [Fis98b, Pum90, Bru87b, Joy87, SG96]. Gatekeeper [Nig99c, Bro97a]. Gauld [Blo94b]. Gauss [Büh81]. Gavin [Sch91b]. Gavroglu [Hac93c]. Gayon [Bow93d, Out94]. GE [Rei85]. Gedichte [Met93, Hüb90]. Gehennical [Man97, New94]. Geison [Aga95a, Pal96a]. Gelehrten [Tur93b, ZH92]. Gender [Jor93, Law92d, Law98a, Uim95, Ben91, Lun94, Sut95]. Gene [Wei90, WKB88, Yox83]. General [Ben90a, Kra93a, Lar96a, Wil91b, Wil97, Whi88, Wis85, Gun84]. Generali [Rap92]. generation [Mor82]. generous [Tom84]. genes [WKB88]. Genese [Eck92, Mei91, Chu91, Duc87, BN82]. genesis [Goo90b, Mei91, Wus84, Duc87]. Genetic [Bug96, Kri82, Pal97, EE95, Wri94b]. Genetics [Bow94b, Dei96, Pal98, Fuj96, Maz92, Sar92]. Geneva [Spa94, Hey83]. Genevan [BaJANJ84]. Genève [Woo95]. genevois [BaJANJ84]. Genius [Hen90a, Kni90b, Man93b, Mur89, Dob91, Jak84, Sim88]. Genova [Fin93a]. gens [Roc88]. gentleman [Dem83]. gentleman-farmer [Dem83]. gentlemanly [Rud85]. Gentlemen [Mor84b, Sec94]. Geochemie [Fri91, New92]. geochemistry [Fri91]. Geoffrey [Hor99b, She99, Jam92b, Law98a, Mor92f]. Geographical [Cor93, Liv93]. Geographies [OW10]. Geography [Har98d, Har96c, Sho93b, CL96, Dor91, Nic96b]. Geological [Mor92b, Old97, CC94, Hal83b, Her91, Her95]. Geologist [Sec96, MS95, Rob96b, Spe82]. Geologists [Sar87]. Geology [Bow92c, Gre82, MK98, Sho93a, Twe91d, BGH^{+}87, Lau87a, LeG88, LR91, LS97, Mac95, Old94, Sar87, Sec86, Sec91]. Geometria [BM89]. Geometric [Whi90a]. Geometrica [Lor90]. geometrical [DFG87, Fie88]. géométrie [Non94, dGB^{+}89]. Geometry [Bra98, Cri90, Fie94, De 95, Edg91b, Ric88, Viè83, dGB^{+}89]. Geomorphology [Tha91]. Geophysical [Sil90]. Geophysics [Sil90]. Georg [Twe91c, HS91, Lin90]. George [Ben93, Daw98a, Lar96a, Man97, Rou97, Sha97, Sho93c, Smi92b, Smi94a, Can93b, HS90, Lev85, Lor90, Mac85, MH88, Mil94, Mol94a, New94, PN95, Rot93b, SL87, Sor99a]. Georges [Sut97, CD94, Olb91b, Out84, Sut97]. Georgian [Por95b]. Geoscience [Le 91, Ste90b]. Gerald [Aga95a, Pal96a]. Gerard [Gou95a, Cro94, ABR93, Bry92a]. Gerd [Hod91]. Gerda [Woo95]. Gerl [Alt91]. German [Daw98a, Di 93b, Gre92c, Har98h, Kus96b, Pal91b, Ras96a, Roh99b, Tim98, Adl87, AH93, Ash95, AS96, Cit90, CM90, Dau98, Dei92, Fri91, Ger89, Gol91a, Gre92b, Gro89, Har96b, Hay98a, Hei93, Hei86, Hei88, HRP90, HP85, Hüb90, Huf82, Jun89, KvDC^{+}37, Kle94, Kra91c, LM23, LHB04, LS76, LHH91, Len82, Lil85, Mai88, Mei91, Mor91a, Mun98, Nyh95, PHvMW93, PvM85, Sch93b, Sch89, Sch84, Sch92b, Ver85, WKB88, Wey92, Wol87, ZH92, vL76, vLvHH^{+}88, vPSBD^{+}04]. German-Speaking [Har98h, AS96]. Germana [Per93]. Germany [Di 93e, Har98c, Wei94, CM90, Dei96, Geu92, Har94c, Jun89, KK87, Lin96, Tho87, Wei91, WKB88]. Gersonides [Rag95, Fre92]. Gesammelte [HD97, Met93, KvDC^{+}37, vPSBD^{+}04]. Geschichte [AH93, Eck92, Hel96a, Hes95a, Hug96, Sch84, Tur93b, Van91, Wei90, WKB88]. Gestalt [Kus96b, Ash95]. Getting [Kni96]. Geuter [Wei94]. Giants [Ost94, Mer93]. Gibbon [WBP97]. Giere [Dol90a]. Gifford [BC98]. Gigerenzer [Hod91]. Gill [Hod93b]. Gillian [Law98b]. Gillmor [Sil90]. Ginneken [Tur93c]. Giorgio [GG91b]. Giotto [Fie94, Edg91b]. Giovanni [Di 98]. Girard [DFG87]. Girolamo [Hen99, Sir97]. Giroux [Aga96a]. Giuliana [Fin92]. Giuliano [Di 93c]. Giulio [La 94]. Giunti [Lar96b]. Giuseppe [Fin93b]. given [Aga98, Hen94b, Por90b, Sho93c, Bro94a, Bro96a, Bro92c, Bro91c, Bur93, Cas93, Goo91b, Gou95a, Hac93b, Har92d, Kat90, Kra98, Lar96b, Mal96, Mal98, Met99, Mol94b, Non94, Rou97, Sha92, Sho93b, Sho93c, Spa94, Twe90, War90a, Whi97]. Glasgow [BC98, Mar92b]. Glaxo [Wil94a, DHS92]. Gld [Mol95, Rag95]. Gleason [Hal92c]. Glimpsing [Hir83]. Global [Mar10, Men86]. Glory [Lin90, Twe91c]. Glossaries [Smi92a]. Glover [Hel96b]. God [Osl97, vR95]. Godlewska [Har96c]. Goethe [Adl87, AZW87, Sep88]. Goethes [Adl87]. Golden [Law91c, Swa92, Wai90]. Goldgar [Sha96]. Goldhammer [Gre93b, Olb91b, Sut97]. Goldstein [Hil98a, Kra93b]. Golem [Gol94, CP93]. Golinski [Law93c]. González [Goo91b].

Good [Bow91b, GR88]. Gooding [Jam91, Jam93b, Mor92f]. Goodman [Ost93b]. Goods [Far94, BP93a]. Gordon [Sho93c, CHG91, Old97]. Gordon/Niigata [Por95a]. gorillas [Out91]. Gottfried [HD97, Sho93a]. göttliche [Gol91a, Gre93b]. Gottlieb [Hen93a]. Goudaroulis [Hac93c]. Gouk [Luc98]. Goupil [Cro96c]. Govardhan [Mur95b]. governance [Mus98]. Government [CK91, Col93, HE92, Por91a, Bru88a, Goo88, Hen89, Smy89]. Gowing [Bow91b, GR88]. Goya [Por97]. grabado [Lan87]. Graf [Gre93a, Wey92]. Graham [Smi99a, Kus94, Mar92a, Sor99a]. Gran [Pan91a]. Grand [Ore90]. Grant [Hen96, Mar96c]. Graves [Law91c, Tay89]. Gray [Hil98a, BC90]. Great [Hac92c, Hal83b, KW87, yQ85, Rud85, McG84, Smi82]. greatest [Man93a, Tor95]. Greece [Llo83]. Greek [Bar92, Mon93, Osb90a, Cop92, Fur87, Fur89, Llo91, Ole84]. Green [Gro95b, Har96d, Smi94a, Can93b, Cla92b]. Greene [Bai93]. Greenhouse [Bro98, Cra96]. Greenwood [Blu90]. Gregg [Nic93a]. Gregorio [Hod93b]. Gregory [Daw98a, Di 93b]. Gresham [Mag96b]. Grew [Jon91, Le 90]. Grison [Cro96c]. Grmek [Fox92, Pyl91]. Groeben [Di 95a]. Grorgio [Pan91a]. Grosholz [Cla92c]. Gross [Hin91]. Grössing [Met93]. Grounds [Cav96, Kni98b]. Group [Twe92c, Gra85, Hei91, Sch89, Wus84]. Groups [Tei93, Fru90]. Grove [Har96d, War91a]. Growth [Hoc93b, Por91a, Bru88a, GH92, Mei91]. Gruber [Di 93f]. Gründerzeit [Eck92, Mei91]. Grundlegung [Bro95a, Kle94]. Gruppe [Koe91, Sch89]. Gruyter [Pos90]. Guagnini [Mar96b]. Guericke [Hel96b, vGA94]. guerra [Mül93]. guerres [Gre86]. Guicciardini [Bal91b, GGC90]. Guide [Ben94b, C^{+}92, Hen94b, Sch91b, Smi90a, BW90, Bru88b, TW87, Kel91]. Gunnar [Wit95a]. gunpowder [NPYGDL86]. Gunter [Ver91]. Gustav [Di 93e, Hei93, Kus95d, Ver91, Hei93]. Gutting [Jor91]. Guy [Mar96c, Ung92]. Guyton [Cro96c, GGSGB94].

H [Bro91a, Cas93, CHG91, FG96, GG93a, GG93b, Gre96, Har96d, Hig94, Jac96, Jon92, Law93d, Mar96c, McE93, Mor90a, Mor92c, Nic93b, Nig99c, Non91a, Ost99, Pal97, Pal98, Sch93c, Sho93c, Smi92c, Spa91, Tim98, Twe92a, War96b]. H. [Hen98a, Mol92, Ore90, Rus91]. Habitat [Whi97, Won93]. Hack [Tha91]. Hacking [Dan92]. Hadden [Cuo95]. Hagen [Nic96a]. Hak [BD90]. Hak-luyt [BD90]. Hakfoort [Can96]. Haldane [MW95]. half [Her88]. half-century [Her88]. Hall [Hut92, Mal94, Ost93c, Ost93a]. Halleux [HS91]. Halley [Van85]. Hambledon [Gou92, Osl98]. Hamilton [O’D83]. Hamlin [Law92b]. Hammond [Col93]. Hand [GG91b, II90]. Handbook [Mai88, Gje86]. Handbuch [Mai88, Pos90]. Hankins [Gou98b]. Hanle [Bro91a]. Hanns [Dol98b, Dol99]. Hans [Edg91a, Bro94a, Hod90a]. happiness [MJ89]. Harald [Edg90b]. Haramundanis [Smi98]. hardback [Bow92c, Jam94a]. hardcover [Aga95a, Aly90, Ari90, Bis98, Blo93a, Boh95, Bow92c, Bro99c, Bru96, Bug96, Bur93, Can90b, Cor93, Cra92a, Dan92, Di 93f, Edg90a, Edg91a, Edg95, Goo97, Goo91b, Gre92c, Gro93, Har92b, Har96c, Hay98a, Hem95, Hen92b, Hen98a, Hoc92b, Hoc92d, Hor95, Hug92c, Jon94, Jor91, Jor94, Kus96a, Law90b, Law96a, Lee90, Lyn99, Mar94b, Mau95, Met93, Mic98, Moo92, Mor92a, Mor98, Mur92a, Mur96, Nic93a, Nig99a, Nig99b, Nut93, O’B92, Ore90, Ost98, Pal91a, Poo99, Pum95a, Ras96a, Rou97, Rud92a, Sho96a, Sho99, Slo97, Smi95b, Smi98, Smi99a, Smi99b, Spa91, Ste91a, Stu91]. hardware [Mor84a]. Harley [Cam90]. Harman [Buc92, Mal93a, Yos99]. Harmke [Spa98a]. Harmondsworth [MK98]. Harmonic [Bra95, Ste94]. Harmonics [Ros94b]. Harold [Col93, Lar96a, CMN95, Sho93b]. Harriet [Twe93a]. Harrington [Har90b, Roh99b]. Harriot [Shi83]. Harris [Lar90]. Harrison [Ili98, Whi90d, Par87]. Harry [Hir95, Lar96a, Ost99, Gol94, Twe91b, YK94]. Hartmann [Gre92c, Mor91b]. Harvard [Bev90, Di 93b, Hoc90a, Hod90a, Kat90, Man97, Pal98, Pau90, Ver90, Whi90d]. Harvester [Kus95b]. Harwood [Div98, Nye93b]. Hatfield [Smi93a]. Haven [Hak95b, Har98d, Jam93a, Pal91a, Sha96, Smi92d]. Hawaii [Bro90c]. Haynes [Rou97]. Haytham [Smi92a, AS89]. Hazard [Ger95, EM94]. head [Rob96a]. Healey [Hen94b]. Healing [Hak95b, Har98c, Cra93, Lin96]. Health [Bar90a, Har98c, Lin96, Por89, Stu98, Web98, Lan89, Web88, Who82]. Heat [Har96a, Mol98a, Por91c, Fre95, Gil83, MGBE95, Mir89]. Heathen [Gay98]. Heavens [Bra95, Pum95a, Mos93, Ste94, FC82]. Hebraica [Rud92a, Lev90]. Hedin [Luc96b]. Hedley [Kni98a]. Heidelberger [Kus95d]. Heike [Hac93c, GG91a]. Heilbron [BV91]. Heims [Aga95b, Twe92c]. Heinrich [Goo97, Har95a, Aly90, Buc94]. Heinz [HD97]. Heinz-Jurgen [HD97]. Heirs [Mor90d, Sto88]. Heisenberg [HP85, PHvMW93, PvM85, HP85, PHvMW93, PvM85]. held [AL87, Ber95, Lar96b, Ver95]. Helden [Ari90, Ver95]. Hele [Jac93]. Helen [Har98d]. Helena [Bar93b]. Hélène [Hal90d]. Helge [Hen91b]. Hellenistic [B^{+}82, Mol98b, Ano82d, EBT96]. Helmholtz [Kre96, Mor98, Smi93a, Van91, Hat90, Tur94].

Helmholtz-Hering [Kre96]. Helminthology [War91a, Gro90]. Helmont [Cle93, Pag82]. Helmontian [Cle93]. Helmut [Hoc92b]. Helmuth [Met93, Hes95a]. Helsinki [Bro96a]. Hemel [Kus95b]. Hempstead [Kus95b]. henceforth [Pan97]. Henderson [Fie92]. Hendrik [Hen86b]. Hendry [CK91]. Henrika [Bow93a]. Henry [Bow93c, Daw98a, Edg91a, Har99b, Hut92, Jen91, Mul91, Sho93c, Cha93, CMNH92, FB97, Hal90c, Hem95, HC90, Jac83, Odd91, Rai91]. Hentschel [Eck92]. Herbert [Edg91a, HD97, Gay98]. Hercules [Kel83]. here [Gay98]. heredity [Rus86b, The94]. heretic [Sno99]. heretical [Sch93a]. Hering [Kre96, Tur94]. Heritage [Fie94, Hor99b, Ali86, Edg91b, Wal84]. Herman [Sim92]. Hermann [Van91, Mor98]. Hermetica [Mon93, Cop92]. Hermeticum [Mon93, Cop92]. Heron [SP98]. Herophilus [Aly90, VoC89]. Herries [Old97]. Herschel [Jam93a, Ash94, HK84, Jan98, Mus98, Sch92a, Vos97]. Hertwig [Di 93e, Wei91]. Hertz [Goo97, Har95a, Buc94, OP87]. Hertzsprung [HK84]. Hespériès [Pyl92]. Hess [HD97, Mau95]. Hessen [Gre92c, Mor91a]. Heuser [Tur93b]. Hevesy [Lev85]. Hevly [Hoc93b]. Hexenprozess [Hüb90, Met93]. Heyd [Hun98]. Hidden [Mur95b, CT93, Far95, GGC90]. Hideo [Hir95]. hierarchical [Sal85]. hierarchies [EMW82]. High [THBB92a, Wri94a, McC93]. Higher [Hil91a, Sta90c]. Hilger [Cra92c, Hac92c, Law92b, Shr92]. Hill [Bis98, Kni98b, Mur95b]. Hills [Car95]. Himself [Ans96]. Hines [Bru91, Wil94a]. Hints [Jam93b, FTG91]. Hiram [Jac90]. hispánica [Lan87]. Histoire [A^{+}87, Bow95, Bry91, CM93, FD87, MC^{+}93, Osb90b, Spa94, Atr86, BBC88, Bow93d, Gay92, Lau87b, Neu92, Rus92, Sin91]. Histoires [Hil98a]. Història [LN95, Mal98, Goo91b, ML89, FC82]. Historic [Fyf99, Pum92, Woo95, Dav99]. Historical [Aga95a, AB89, Bra92b, Bro96a, Byn90, CM90, Fau96b, Gou95a, Hal91b, Har93b, Hes95b, Hod90b, Kla95, Kus95c, Luc96b, McC81, Mil94, Por91b, Sha92, Tur91, ABR93, Bal91a, Bow88, Coo89b, Dan90, Dea91, DD94, Edg91a, KT87, Lin93, Liu92, MR82d, MR82e, ME81, NC81, Ols90, OE^{+}90, Pic93b, SO85, SY86, T^{+}85, ZH92, AH93]. históricas [Oli92, OE^{+}90]. historicized [Sto84]. Histories [Get96, Hem95, Fre94, Pet94]. Historiographical [Hen98a, Coh94, Hes99]. Historiography [Hac95, Jor93, Kra87, Vos97]. Historiques [Dav94, GG90]. Historischen [AH93, CM90, Hes95a, Tur93b]. historisches [ZH92]. History [Aga96b, Aga98, Alb92, AL87, Ano97c, Ano97d, Ano98g, Ano99e, Ano05, A^{+}87, Ben90a, Ben94b, Blo98, Blo94b, Boh95, Bow94a, Bow91a, Bow93a, Bow93c, Bra91, Bro98, Bro96b, Bro99c, Bro91b, Bro91c, Bro95d, Bro97d, Bru88c, Bry91, Bud97, Bur98, Büt83, CK89, Cam90, Cam92, Can97, Car95, Col93, Coo96, Cra92c, Cro96b, Cro96c, Dea92, Di 93f, Div98, DTS^{+}87, Edg95, EO90, For90a, Fox92, Gau99, Ger91, Get96, Gol91b, Gou91, GG91b, Gre95, Gre92c, Gre96, Hac90d, Hac92d, Hac93a, Hac94, Hac95, HSHSJ93, Hal90a, Hal92a, Hal92d, Hal94c, Har99b, Har96d, Hay98a, Hen94a, Hen95b, Hes95b, Hil98a, Hoc90a, Hor98, Hug96, Jac96, Jam90, Jam94d, Kip91, Kit89, Kni90a, Kni90b, Kni90c, Kni92a, Kni92b, Kni98a, Kra93a, Kus95c, Kus96b, Kus97, Lar96b, Law90a, Law90b, Law91a]. History [Law96a, Lee90, LS90, Lev94, Luc96b, Mal93b, Mal93a, Mar92a, May99, McC95, McE93, McE95, McK97b, Mic98, Mol98b, Mor92a, Mor99, Mur95b, Nic93c, Nic94a, Nic94b, Non91a, Nye92b, O’B96a, Odd91, Olb91b, Osl97, Ost98, Pal91a, Pal97, Pic92b, Por91b, Pum90, Pum91c, Pyl93b, Rob93, Rus91, STT87, Sch95b, She92, Sil90, Slo97, Sme97, Smi90a, Smi91b, Smo94, Soy96, Spa95, Spa98a, Spa98b, Stu98, Tay87, Tha91, Top97, TST87, Tur93a, Tur93c, Twe92b, Twe93a, Twe93b, Ver94, War91a, WKB88, Whi97, Wil90a, Wil91c, Wil91b, Wil94a, Wil95a, Wil97, Wit95a, Woo92, Yeo96b, Aft91, AH93, Ali86, All86, Ano82c, Ano82b, Ano93q, Ara91, Atr86, AB89, BA91, Bar93a, Ben87, Ben97, Ben90b, BP95, BBC88, Blo94a, Bow92d, Bra89, Bro92a, Bro95b, Bro99a, Bru88b, Bry92b]. history [Bry88, Bud93, Buf87, BP93b, CKA96, Can88, C^{+}92, Car94, Car88, Car86, CW87, Cha97, CS94, Cha82, Coh96, Coo88, CT93, Cre93, Cro91b, Dah97, DHS92, DM96, DeV82, Dol98a, EE95, FB97, FB82, Fin93d, Fis94, Fog92, Fox95, Fre94, Fru82, GS88, GA98, Gau92, Gay92, GEMM83, Gro90, HE92, HW03, HH86, HH92, HW81, Hen86b, HK84, HF87, Het88, Hil84, Hil94, Hom83, How96, How82, Hun85, Ing84, II90, JK84, JSS96, Joh84, Joy87, Kel91, KRS83, Kni92d, Kri82, Kuk91, Kuk93, Lau87b, Lea90, Lea94, Liu92, Liv93, MKW92, Maz95b, McC82, McC92, Men86, Mer89, Mil89, ML89, Mor90b, Mor90c, Moy83, Mur89, NW82b, NW82a, Nor89, Nor94b, O^{+}90, Old86, Old95, Ole84, Out91, Pal98, Pat94, Ped92b, Pet89, Phi87, Pic84]. history [Pic90, PR95, PW89, PT86, PB99, RM96, RR81, Roc90, Ron83, Ros84b, Rot93a, R^{+}85, Rus93, Sar87, SSK88, SG90, Sch84, Sec94, Sha83, SP88, SW89, Sho93e, Sin91, Sin97, Smi92e, SM94, Smi97, Ste91c, Sti86, TT86, Tri89, TW87, Twy88, Way87, Wea88, Wea90, Web98, Wei92, Wel93, WR92, WHS92, Whi88, Who82, Wil81, WD82, Wil94b, Won93, Wus84, aHH86, aHH88, aHH92, vdW85, LN95, MC^{+}93, Rai91, Wei92].

Hitler

[Dei92, Har94b, Roh99b, Dei92, Har96b]. HMS [DD94]. HMSO [But94, Col93]. Hobbes [Str90, Nap96, Pro93, SSH85, Sor86]. Hodder [Ost93b]. Hoddeson [Kra91b]. Hodge [Hal92a]. Hodgkin [Aga93]. Hofbauer [Sho93a]. Hofmann [Tra92]. Hohenheim [Daw98a, PTS^{+}96]. Hohenlohe [Gre93a, Wey92, Wey92]. Holbrook [But94]. Holism [Kus96b, Roh99b, Ash95, Har96b]. Hollister [Mar92a]. Hollister-Short [Mar92a]. Holmes [Gol91b, Wei94]. Holy [Gre96, Smi94c]. Home [Bro90a]. Homes [Aga95a]. homme [Hil98a]. Homoeopathy [Nic88]. honest [For85]. Honnecourt [Kel95, Bec91b]. Honolulu [Bro90c]. Honor [Kni92a, BEFH10, MH88, SG90]. Honour [Bow91b, LS90, Mal93a, Non91a, GR88, WHS92]. Hook [Old97, Her95]. Hooke [Pum91a, Ili95c, Ros84b]. Hoover [PLP92]. Hooykaas [Ost99]. Hope [Hen93b]. Hopkins [Aga98, Bai93, Bow92b, Daw98a, Har98c, Har99b, Har95b, Har98g, Hod90b, Hor96b, Hor99a, Hug92a, Joh93, Mar93c, Mar96a, Nye93b, Roh99a, Rou97, Sho93b, Smi95b, Spa98b, Wri94a]. Horace [Woo95, CW87]. Horace-Bénédict [Woo95]. horoscope [GGC90]. Hortus [Hen86b]. Horwood [Hac92a]. Hoskin [Wil91b]. Hourya [Neu92]. House [Hal92d, De 96]. houses [Ste99]. housewife [SW93]. Houston [Ver95]. Howard [War96a, Hor95, Wil84, Wri94a]. Howse [Cha90c]. Hoyningen [Jor94]. Hoyningen-Huene [Jor94]. Høyrup [Mol96]. Huainanzi [Lew95]. Hubble [Aga94a, Aga96a, Chr95, SN93]. Hubert [Mol94b]. Hübner [Met93]. Huene [Jor94]. Huertas [War90a]. Hufbauer [Hug92a]. Huff [Hen95a]. Hugh [Sec96, Sec98, MS95, Sho96b]. Hughes [Hoc92b]. Hugo [The94]. huitiéme [Sch01].

Huizenga [Sim94]. Hulin [GG90]. Hulin-Jing [GG90]. humaines [Bow95, CM93, MC^{+}93]. Human [Blo98, Jam91, Jon92, Jon93, Mur95b, Mur96, Smi95b, War91a, Wea91, CT93, Deg91, Dun90, Goo90c, Gra87, Gro90, Har94a, Kai94, MJ89, Maz92, Ros94a, Smi97, Sta85, Van93]. Humanism [Fie94, GJ86]. humanist [Slo85]. Humanists [Mal95a, FJ93]. Humanities [Mur92a, GJ86, MMVZ90]. Humboldt [Hei88, Roh99a, Hei88, Nic96b]. Humboldtian [Nic96b]. Humphry [Jam94b, Kni92c]. Hundred [Div90, All86, Hea86, Hea90, vLP82]. Hunt [Har93a, Sch91b]. Hunter [Ans96, Elm97, Hen94b, Lyn99, Mur92b, Nye92b, Ost98, Ost99, Pum91b, Pum91a, Spa91, Ste90a, BP02]. Hunterian [Di 93f, OS92b]. Hunting [Sta90b, Mac88a]. Hutton [Hen93b, Lyn99, Mul91, Tha91, Old94]. Huxley [Goo96, Sho93c, Ver95, CHG91, Des94, FG96, Jen91]. Huygens [Ano82a, Coh91]. Hydrographic [Hac90e, Man88]. Hygiene [Pau90, Liv99, Pro88]. Hyman [Mur92b, Tha91]. Hypatia [Ali86]. Hypnosis [Smi93c, CS92]. Hypnotism [Blo94b, Gau92]. hypothèse [Bow93d, Gay92]. Hypotheses [Ait90, Cla89]. hypothesis [AH85, Gay92]. Hysteria [Law96c, MF86, Mic95].

I. [Hen92b, Sto88]. Ian [GG91b, Hen90b, Hor98, Dan92, Ley90, Smi91b]. IBM [Twe92a, Pug84, PJP91]. Ibn [Hil93, Smi92a, AS89]. IBSN [Str90]. Icelandic [Jam94c, Bar91]. ICL [Wil91c, CK89]. iconography [HS99]. Idea [Fin92, Kni90b, Ree90, Cro86, Mur89, PR88]. Ideas [Aga95b, Ale85, Bro96b, Ger91, Kni92d, Kus94, Rob93, Sch95b, Ano05, Fre90b, Goo91b, Lau87b, Mor90b, Ric92a, Tau93, Wie93]. idées [Lau87b, Spa94]. identités [Hil98a]. identités/Mathematical [Hil98a]. Identity [Hil98a, Por90b, Rou97, Fox88, Tur96, War86]. Ideological [Boh93, Jan98, Ric92b]. Ideology [Can88, Olb91b, BM82, Llo83]. IEE [Hac92d, Hac93a, Hac94, Mor99]. Igor [Aga94a]. ihre [Gro89, Hin91, Wol87]. ihren [vLvHH^{+}88]. II [Bow92c, Cre91b, Fin93a, Gre93a, Mol94b, Pan91a, Roh99b, Smi92a, Wey92, Wol87, Bra92b, Gol91b, BEoCF92, Cla95a, ECL92, Goo88, Har96b, WW86, Wol87]. III [Blo93b, Bow92c, Fau96b, AS89, Lan87]. IIIe [Cle92]. Ilana [Ams92]. Ilie [Por97]. ill [For90a]. Illus [Rus91, Aga96a, Aga97b, Aga97c, Bro92b, Car95, Col93, Cra92a, Fau96b, Gol93, Gou92, Gou98a, Har98d, Har98g, Hel96b, Hen93a, Hoc93b, Hug92c, Hun92, Kra98, La 97, Lar96b, Lar96a, Law93g, Law93f, Luc96b, Mal98, Man93b, Man97, McE95, McK96, Mor92a, Mur92b, Old97, Pal97, Pal98, Pum92, Roh99b, Rud92a, Spa98b, Twe93b, Wri96b]. Illusions [Whi97, Won93]. illust [Kra92]. illustrated [HH86, HH92, Ron83, Sta84a, aHH86, aHH88, aHH92]. Illustration [Law93g, Sch91b, BW90, Dol98a]. Illustrations [Fyf99, CD81, Dav99, RT92]. im [CM90, Dau98, Di 92, Fuc99, Hel96a, Jun89, Kra91c, Sch93c]. Image [Ful98, Dol96, Jor90]. imagery [Bur94]. Images [Ano04, Blu90, Buc98, Hoc90a, Jac95a, Kla95, Rot93b, FL88, PLP92, Wea88]. Imagination [For91, Gou98b, Nic93c, Sho96a, HS95, Smi94b, Wil90b]. Imaginative [Har98d, CL96]. Imaging [Law93g, Sta91]. Immagini [La 94]. immobility [Gra84]. Immortal [Aga96a, She95]. Immunology [Jac96, Law92a, Maz95b, TC91]. Imperfect [Hor95, Seg94]. Imperial [Di 93e, Mus98, Ros86, Wei91].

Imperialism [Bal91a, Har93b, Har96d, Smi91a, Sta90b, Gro95b, Mac88a, Sta89a]. implications [Ver85]. Impolite [Gol95a, Sha96]. importance [Don91, Gil83]. Impulses [Smi95b, Ros94a]. Impurity [Law92b, Ham90]. ’Inattention [Mor90e]. Including [Hal94a, Ber93b, SSH85]. Income [Ano96o]. Incorporated [Har92a, Ste91b]. incorporating [Dra81]. Incubator [Hor99a, Bak96]. indeterminism [PB82]. Index [Ano91r, Ano92o, Ano93p, Ano94o, Ano96p, Byn90, Fau96b, Hug96, WW93, Coo89b, Gou91, WW75, WR92, Ano97c]. indexes [Whi90a]. India [Har93b, Des82c, Har92e, Kum96]. Indian [Har92c, Har92d, Har93c, Des92, Kum91, San91]. Indiana [Bro93d, Di 93c, Hef96, Le 91, Spa91]. Indianapolis [Di 93c, Hef96, Le 91]. Indians [ML89]. Indias [Goo91b, ML89]. indicators [T^{+}85]. Indies [Spa98b, PR95]. indigo [THBR92]. individual [Hen94a, Kni92b]. individuals [Bro93a]. indivisibilibus [WW88b]. Industrial [Don91, Goo90a, Hil91b, Hoc92b, Hul99, Ink90, Mar96b, Nye92b, Nye93b, CM90, CK90, ES90, FG93, HL86, Hun85, Ken86, Rei85, Tho87, Twe91d, Wis85, Mac88b]. Industrialisation [Har92a, Sch88]. industrialization [Ste91b]. Industrie [Bry91, CM90]. industrielle [CM90, Sch93c]. Industry [CK91, Cam92, CM90, Hac92d, McC95, Rus94, Aft91, Ber85, Hen89, Lie87, Mor90c, Poc88, Pug84, Swa88, THBR92, THBH92a, THBH92b, Tra93, Wes88, Wil81, Wil94b]. inédits [vPvP90]. inertia [yQ85]. Infancy [Kel83]. infini [Bla93, Hal95]. Infinity [Kre82, Pro93, Rav99, Bla93, Fie97]. Influence [Bro92b, Mar93b, Smi92d, Ano90k, Bow90, Rit90, Tun92]. Influential [Bar90b]. Information [Aga98, Hor96a, Hor97b, Ben86, CKA96, Ili95c, NOF96]. informed [Söd86]. infortunes [dCC09]. Ingénieurs [Gre86]. Ingold [Rob96a]. Ingrao [GG91b]. Inhalation [Law93d, SE91]. Inhibition [Smi92e, Soy96]. initiative [OO89, Twe91a]. innere [Hei93, Kus95d]. Innovating [CK91, Hen89]. Innovation [Hen98b, Mar93a, Sie91, Wes88, Ber85, Har94c, Lie88, McG95, THBH92b]. Inoculating [May97a]. Inquiry [Hen98a, Nig99b, Coh94, MJ89, PR90, Pow96, Wyn82]. Inquisition [Sha97]. insects [Cla97]. insider [De 92]. Inspiration [Ric97]. Institut [AH93, CM90, Hes95a, AH93]. Institute [AL87, Bro99c, CM90, Gre92c, Hoc92b, HD97, Kni92b, Moo92, Mur95b, Pri91, Sch91b, Smi92a, Spa98a, Twe92b, Twe93b, Wil90a]. Institution [Cas93, Edg90b, Hac92d, Hac93a, Mar93b, Mor99, Cra84, Hun82, PLP92]. Institutional [Har94c, Kni92b, Nut93, Hen94a, Lin92]. Institutionalization [Hal92d]. institutionalizing [GJ86]. Institutions [Bis98, Bro93a, Nol95]. Instruction [Twe93a, Bry92b]. Instrument [CW87, Lil85, McC92, Tur93a, Gas99, War90c]. Instruments [Bry91, Bry92a, But94, Gou95a, Gou98b, Hac90d, HS95, McC95, ABR93, Ben87, HAB92, McC81, McC82, Tur83, Wil94b]. insulin [Bli82]. Intellectual [Boh95, Coo96, Elm97, Hoc92a, Osl97, Hey81, Hun95b, Sha93b, TB97, Wal90, dG01]. intellectuelle [dG01]. Intelligence [McK97b, Cre93]. Intelligent [Twe91b]. Intensive [Hor99a, Bak96]. Inter [Aga97c, Nye92a, Luc90]. Inter-disciplinary [Aga97c]. Inter-War [Nye92a, Luc90]. interaction [THBM92]. Interdisciplinary [Blo93c, Doe96, Kit92]. interests [Sec94]. Interference [Jam94d, Kip91]. internal [Hei93]. International [Bro96b, Cam92, Hac94, Hoc92c, Kni92b, Sch95b, Tha91, War91a, Aft91, A^{+}87, BG^{+}92, CM90, Out94, Sar87, Woo92, Rei94]. Internationale [Bry91]. internationales [CM90]. Internationalism [Hoc93a, Cra92b, Hoc94]. Internet [Rou97, Tur96]. Interpretation [Bro93a, Jac96, Maz95b, SW93]. Interpretations [Law96c, Mic95]. interregnum [Gou96]. Introduction [Aga95b, Bow92c, Bra92a, Bro97b, Bro97c, Coo90, Di 95a, FJ98, Hac93b, Hen93b, Hun99a, Jam94c, Lor90, Mal94, Mol94a, Mol94b, Mon93, Mor95b, O’B92, Olb93, Sec93, Smi92a, Sor99b, Spa91, Yeo96b, Ari90, BL83, BM89, CLR87, Cop92, CK90, Di 93f, EiT92, ET92, Gou91, Gun84, Hal93, Hey83, HS91, Kni98b, Kra87, Lin84, Mor98, Sir90, Tak92, Twe92b]. Introductions [Mil94, Ams92]. introductory [Daw98a, Law93d, Old86]. inventario [Fin93b]. Inventing [Ink90, Mac88b]. Invention [Aga95b, Coo96, Jam93a, Met99, Nye93b, Rav99, Smi91c, Wie93, Bow89, Don94, Fie97, FC96, FIF86, Gil83, Sch92a, Wil95b]. Inventory [Hug96, WR92, Olm92]. Investigaciones [Oli92, War90a]. Investigation [Mal93a, Tra92, WHS92]. Investigative [Gol91b]. Investment [Cou96]. Invisible [Coo96, Gay96, GG91b, Hir83, II90, Wil95b]. Inward [Pai86]. Ionic [Bro98, Cra96]. Iowa [Smi90a]. Ippolito [Law91a]. IR [Wil90a, Pal96b]. Iraq [MM10]. Ireland [Moo92, Old97, Can83, Her83, Her95]. Irish [Law91c, Moo92]. Irwin [Pic98]. Isaac [Bra92a, Hal91a, Hen94b, Ost93a, Bra92a, Fei90, Hal92b, Man93a, NS84, New89, Sno99]. Isabella [Non93b]. Isabelle [Smi93c]. ISBN [Aga93, Aga94a, Aga94b, Aga95a, Aga95b, Aga96a, Aga97b, Aga97c, Aga98, Ait90, Alb92, Alt91, Aly90, Ams92, Ans96, Ari90, Bai91, Bai93, Bal91b, Bar92, Bar90a, Bar93b, Ben90a, Ben93, Ben94b, BV91, Bev90, Bis98, Blo93a, Blo98, Blo93b, Blo93c, Blo94b, Blu90, Boh93, Boh95, Bos92, Bow92a, Bow94a, Bow91a, Bow91b, Bow92c, Bow92b, Bow93d, Bow93a, Bow93b, Bow93c, Bow94b, Bow95, Bra90, Bra91, Bra92a, Bra92b, Bra95, Bra98, BD90, Bro90a, Bro93b, Bro94a, Bro95a, Bro98, Bro96a, Bro91a, Bro92b, Bro93c, Bro93d, Bro94b, Bro90c, Bro92c, Bro96b, Bro99c, Bro90e, Bro91b, Bro91c, Bro95d, Bru91, Bru93, Bru96, Bry91, Bry92a, Buc90, Buc92, Bug96, Bur93, Bur98, But92a, But94, Byn90, CK91, Cam90, Cam92, Can90a, Can90b, Can91a, Can96, Car95, Cas93, Cas94]. ISBN [Cha90c, Chu91, Cla92c, Col93, Coo90, Coo96, Cor93, Cra92a, Cra92c, Cre91b, Cri90, Cro94, Cro95, Cro96b, Cro96c, Cro96d, Cuo95, Dan92, Dav94, Daw98a, Dea92, Dea93, Di 92, Di 93a, Di 93b, Di 93c, Di 93f, Di 93d, Di 93e, Di 95a, Di 98, Div90, Div98, Dol94, Dol98b, Dol99, Dol90a, Dol90b, Eck92, Edg90a, Edg90b, Edg91a, Edg95, Elm91, Elm97, Far94, Far99, Fau96b, Fie91, Fie92, Fie93, Fie94, Fin93a, Fin93b, Fin93c, Fis98b, For90a, For91, Fra90, Fre93, Fuc99, Ful98, Fyf99, Gau99, Ger91, Ger95, Get96, Gol91b, Gol93, Gol94, Gol95b, Goo90a, Goo96, Goo97, Goo91b, Gou91, Gou92, Gou95a, Gou98a, Gou98b, GG90, GG91b, GG91c, GG96, Gra91, Gre91, Gre95, Gre92c, Gre93a, Gre93b, Gre96, Gro93, Hac90d, Hac90e, Hac92a, Hac92b]. ISBN [Hac92c, Hac92d, Hac93a, Hac93b, Hac93c, Hac94, Hac95, Hak95b, Hal91a, Hal91b, Hal92c, Hal92a, Hal92d, Hal94a, Hal94c, Hal95, Har92a, Har98a, Har98b, Har98c, Har99b, Har99a, Har90a, Har93a, Har96a, Har98d, Har90b, Har92b, Har92c, Har92d, Har93b, Har93c, Har95b, Har96c, Har96d, Har97, Har98g, Har94b, Har98h, Hay98a, Hay98b, Hef96, Hel96a, Hel96b, Hem95, Hen90a, Hen90b, Hen91b, Hen92b, Hen93a, Hen93b, Hen94b, Hen95a, Hen95b, Hen96, Hen98a, Hen98b, Hen99, Hes95a, Hes95b, Het95, Hig94, Hil91b, Hil91a, Hil91c, Hil93, Hil98a, Hin91, Hir95, Hoc90a, Hoc91, Hoc92b, Hoc92a, Hoc92c, Hoc92d, Hoc93a, Hoc93b, Hod90a, Hod90b, Hod91, Hod92a, Hod93a, Hod93b, Hod96, HS90, HS91, Hor92, Hor95, Hor96a, Hor96b, Hor97a, Hor97b, Hor98, Hor99a, Hor99b, Hud92, Hug92c, Hug92a]. ISBN [Hug92b, Hug96, Hug98, Hun92, HD97, Hun98, Hut92, Ili95b, Ili97, Ili98, Ink90, Jac96, Jac90, Jac93, Jam90, Jam91, Jam92a, Jam92b, Jam92c, Jam93a, Jam93b, Jam94a, Jam94b, Jam94c, Jam94d, Joh93, Joh96, Joh94, Jon91, Jon92, Jon93, Jon94, Jor91, Jor94, Jor96, Kat90, Kel95, Kla95, Kni90a, Kni90b, Kni90c, Kni92a, Kni92b, Kni95, Kni98a, Kni98b, Koe91, Kra90a, Kra90b, Kra91a, Kra91b, Kra92, Kra93a, Kra93b, Kra94, Kra98, Kre96, Kri91, Kus94, Kus95b, Kus95c, Kus95d, Kus95e, Kus96a, Kus96b, Kus97, Kus98, La 97, Lan90, Lar90, Lar96b, Lar96a, Law90b, Law90c, Law91b, Law91a, Law91c, Law92a, Law92b, Law92c, Law92d, Law92e, Law93g, Law93b, Law93c, Law93d, Law93e, Law93f, Law96b, Law96a, Law96c, Law98b, Law98a, Law98c, Le 91, Lee90, Lev94, Ley90, Lor90, Luc92]. ISBN [Luc93, Luc96b, Luc98, Lyn99, Mac90b, Mag96a, Mal93b, Mal93a, Mal94, Mal95a, Mal96, Mal98, Man93b, Man97, Mar92a, Mar93a, Mar93b, Mar93c, Mar94a, Mar96b, Mar96a, MK98, Mar94b, Mar96c, Mau95, Maz96, Maz98a, Maz98b, McC98, McC95, McE93, McE95, McK96, McK97b, McN92, Met93, Met96a, Met96b, Met99, Mic98, Mil95, Mil94, Mol92, Mol94a, Mol94b, Mol95, Mol96, Mol98a, Mol98b, Mon93, Moo92, Mor90a, Mor92a, Mor90d, Mor91c, Mor92c, Mor92f, Mor93, Mor95c, Mor96a, Mor98, Mor99, Mul91, Mur92a, Mur92b, Mur95b, Mur96, Nea98, Neu92, New92, Nic93a, Nic94a, Nic94b, Nic96a, Nig99b, Nig99c, Non91b, Non91a, Non93a, Non94, Non93b, Nor94a, Nut93, Nye92a, Nye92b, Nye93b, O’B92, O’B96a, O’B96b, Odd91, Olb91a, Olb91b, Old97, Oli92, Ore90, Osb90a, Osb90b, Osl97]. ISBN [Osl98, Ost93b, Ost93c, Ost93d, Ost93a, Ost94, Ost98, Ost99, Out94, Pal90, Pal91a, Pal91b, Pal96a, Pal97, Pal98, Pan91a, Pau90, Per93, Pic92b, Pic93a, Pic98, Poo99, Por90b, Por90a, Por91a, Por91b, Por91c, Por95a, Por97, Pos90, Pri91, Pri93, Pum90, Pum91b, Pum91a, Pum92, Pum94, Pum95a, Pyl91, Pyl92, Pyl93a, Pyl93b, Rag95, Rap92, Ras96a, Ras96b, Rav99, Ree90, Rob93, Roh99b, Roh99a, Rou97, Rud92a, Rus91, Rus94, Sch91a, Sch93c, Sch95b, Sch91b, Sec96, Sec98, Sha97, Sha92, Sha93a, Sha96, She99, She92, Sho93a, Sho93b, Sho93c, Sho96a, Sho99, Shr92, Sim92, Sim94, Slo97, Smi90a, Smi91b, Smi91c, Smi91a, Smi92c, Smi92b, Smi92d, Smi92a, Smi93b, Smi93a, Smi93c, Smi93d, Smi94d, Smi94a, Smi95b, Smi96, Smi98, Smi99a, Smi99b, Soy96, Spa91, Spa95]. ISBN [Spa98a, Spa98b, Sta90b, Ste91a, Ste95, Stu91, Stu98, Sut97, Swa92, Tei93, Tha91, Tim98, Tur93a, Tur93b, Tur91, Tur93c, Twe90, Twe91a, Twe91c, Twe92a, Twe92b, Twe92c, Twe93a, Twe93b, Twe93c, Uim95, Ung92, Ver90, Ver94, Ver95, Ver91, War91a, War90a, War91b, War96a, War96b, Wat91, Wea91, Wei90, Wei94, Whi97, Whi90a, Whi90c, Whi90d, Wil90a, Wil91c, Wil91b, Wil94a, Wil95a, Wil97, Wis96, Wit95a, Woo98, Wri94a, Wri96b, Yav92, Yeo92, Yeo96a, Yeo96b, Yos99, vH92]. ISIS [Ano82c, Ano82b, NW82b, NW82a]. Islam [Hen95a, Hil91a, Huf93, Huf95, Sta90c]. Islamic [HH86, HH92, aHH86, aHH88, aHH92]. Island [Har96d, Gro95b]. Isles [Byn90, All86, Coo89b]. ISN [Twe91b]. Isobel [Bro99c]. isomorphicism [Sch84]. Isomorphismus [Sch84]. Israel [GG91b, Mar93c, Mar96a, Nye93b]. ISSN [Bro93a, Cro96d, Fie92, Hal90d, Hen94a, Law90a, War95b, Woo95]. issue [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Bro93a, Kni92b]. Issues [Luc96b]. Istituto [Bal91b, Di 98]. Italian [Di 98, Hac92b, Ara91, Des89b, Ern91, GGC90, Lan91, Maz95a, Olm92, Sir87, Yat83]. Italianate [Ost94]. Italy [Di 93c, Lar96b, Sha97, Ber95, Fin91, Pan91b, Sir87]. Ithaca [Bro91b, Can91a, Fie94, Fis98b, Har93a, Kra90b, Luc93, Yav92]. iv [Hes95b]. Ivo [Met96a]. IX [Law90a, Aga98, Blo93a, Blo93c, Bow93a, Bra91, Bro95a, Bur93, Car95, Cla92c, Di 93f, Hac92c, Hal92c, Hal94a, Hel96a, Kni90b, Kus95c, Mar93b, Mol95, Ost93b, Out94, Rou97, She92, Smi92b, Smi94d, Soy96, Whi90a]. Ixiv [BD90].

J [Aga95b, Ben90a, BV91, Boh93, Boh95, Bow92c, Bro92b, Bro99c, Bro91c, But94, Dah97, DF97, Di 93e, Fau96b, Fie91, Fie92, Fre93, Goo90a, Gou95a, Gou98b, Gre91, Hac93b, Hal92a, Har93a, Hen94a, Hod90b, Hor96b, Hun98, Jam93a, Kim95, Kra93a, Lan90, Law93g, Law93b, Law93f, Mal95a, Mar92a, Mar96a, Mau95, McC98, Mol95, Mor92f, Mur92b, Neu92, Nic93b, Nic94b, Osl97, Out94, Pyl92, Rag95, Sch91b, Slo97, Smi91c, Str90, Tha91, Twe92c, Twe93a, War95a, War96a, Wei94, Whi90c]. J. [Ack87, Cam90, Dah97, DF97, Hac90d, Hen98b, Hug92c, Kim95, Kra93a, Mag96a, Mal95a, Maz96, MW95, Mol98b, Moo92, Osl97, Out94, Rav99, Sho93c, Tur91]. J.-C. [Out94]. Jaan [Smi93d]. Jaap [Tur93c]. jaar [Bur99]. Jacalyn [Fox92]. Jack [Kra93b]. Jacob [Di 93f]. Jacobean [Daw98b]. Jacques [HC91]. Jahre [AH93, Hes95a]. Jahrhundert [Bro95a, CM90, Dau98, Fuc99, Sch93c, Kle94]. Jahrhunderts [Di 92, Koe91, Jun89, Sch89]. Jaki [Mar96c]. James [Bro93c, Buc92, Goo90a, Har90a, Hen90b, HD97, Law91c, Mal95a, Mar92a, MK98, Mor92f, Nea98, Rob89, Sch91a, Sch91b, Wal88, Yeo96b, Yos99, Bir86, But92a, Car89, Dol98a, Gol83, Hac94, Har98e, Hen86a, MH02, Old94, Ost84, Tay96, Wri87]. Jan [Gre93a, Blo93a, Law93c]. Jane [Di 95a, Bow92b]. Janet [Ben94b, La 97]. Janice [Fie92]. Janus [Man93b, Dob91]. Japan [Bro90c, Hir95, Hoc92c, Luc98, Por95a, BR91, Gou95b, MR89, Nak91, YK94]. Japanese [Bro92c]. Jaramillo [Gle82]. Jarcho [Law91a]. Jardine [Slo97, Ste91a]. Jaroslav [Mur92a]. Jauregg [Por95a, Whi93]. Javier [Oli92]. Jean [Bow93d, Fin93c, GG96, Cro94, Hen93a, Hey83, Pic87, Pum95a, Yeo96a]. Jean-Claude [GG96]. Jean-Francois [Fin93c]. Jean-Pierre [Cro94]. Jean-Robert [Hey83]. Jed [Jam94c, Can90b, Goo97]. Jefferson [Hol86]. Jeffrey [Hor99a]. Jenkin [Mor94]. Jenner [Bru93, Fis91b]. Jens [Mol96]. Jensen [Sho93c]. Jeremy [Hil98a]. Jerrold [Kra93b, Gol92a]. Jersey [Fox92, Law93e, Mor92d, Mor93]. Jesseph [Mil94]. Jesuit [Kra91c]. Jesuiten [Hel96a, Kra91c]. Jevons [GG91c, Sch90a]. Jewish [Gre96, Kat90, Fre92, Pat94, Rud88]. Jim [Hil98a]. Jing [GG90]. Jo [Kni95, Man93b]. Joan [Hor92, Cri90, Pag82, Pal98, Por91b]. Joanne [Hir95]. Jodrell [Aga94c]. Joel [Nic96a]. Johan [Hil91b]. Johann [Twe91c, Lin90]. Johannes [Fin92, Gre92c, Met96a, Met93, Mor91b, Sch93b]. John [Bow92a, Di 93f, Dun92a, Dun92b, Edg90b, Elm91, Fie93, Har99a, Hod91, HD97, Law93e, Mor90a, Ost99, Sch91a, Sho93a, Twe92a, Twe93c, Ver90, ATKL08, Asp90, BB82, CK91, Cla97, Clu88, Cra92c, Day99, DF98, Dol98a, Far94, FC82, FFLW97, Gra91, Jan98, Kni98a, Kno90, Law93d, Le 91, Mor92d, Mun98, Mus98, Pen89, Por90a, Pou93, Pro93, RWD90, She99, Sho99, Sim94, Sta90b, Sul82a, Twe92b, Wil95a, Wri96b, Wri96a]. Johns [Aga98, Bai93, Bow92b, Daw98a, Har98c, Har99b, Har95b, Har98g, Hod90b, Hor96b, Hor99a, Hug92a, Mar93c, Mar96a, Nye93b, Roh99a, Rou97, Sho93b, Smi95b, Spa98b, Wri94a]. Johnson [Twe92a]. Jonas [Joh94, Wil93]. Jonathan [La 97, Luc93, Sho96a]. Jones [Hen94b]. Joppien [BD90]. Jordanus [Mol92, BJ91]. Jørgensen [Kra97]. Jose [Mal98, Goo91b, Kus98]. Joseph [Bro91a, Bro93c, Goo96, AL87, Dav97, Har95a, NR86, Tei93]. Josephson [Woo98]. Jost [Gre93a]. Joule [Har90a, Car89]. Journal [Ano97c, Hen94a, Kni92b]. Journals [Lar96a, Por90a, MD88]. jours [Ano87b, CLR87]. Joy [Pum90]. Joyal [Fin93c]. Jr [Lar96a]. Judgement [Hem95]. judgment [Pet94]. Judith [Law93e]. Judy [Aga98, Wil94a]. Juergen [War95b, War96a]. Juglers [Hil98b]. juin [BG^{+}92]. Julian [Ver95, Elm91, Fra90, Sha93a]. Julien [Mic98]. Julius [Por95a, De 96, Whi93]. July [McK96]. June [Ano97d, BG^{+}92, Di 93f, LHH91, Mar93c, McK96, Out94, OS92b]. Jung [Kus95e, Smi99b, Nol94, Sha98, YED97]. Jungian [Cla99b]. Juni [LHH91]. Junker [Di 92]. Jurgen [HD97, Met93, Wei90]. Just [Bro84]. Justus [Nig99c, Tur93b, ZH92, Bro97a, Mun98, ZH92, vLvHH^{+}88].

K. [Mor99, Tha91]. Kabbalah [Bro96b, Kat90, Cou95, Rud88]. Kaijser [Luc96b]. Kaleckian [Cou96]. Kallinen [Bro96a]. Kalthoff [Lar96a]. Kamerlingh [GG91a, Hac93c]. Kameshwar [Kra92]. Kamminga [Hil91b, Spa98a]. Kant [Smi93a, Cro86, Dic82, Ful91, Hat90, Wal78]. Kapitza [Hug92c, Bad85, KBRS90]. Karen [Whi97]. Kargon [Bro91a]. Karl [Di 95a, Eck92, Hug92a, Mag96b, Mei91, SP98, vBDGO93]. Karrieren [Dei92, Har94b]. Katherine [Smi98]. Kathryn [Yav92]. Kauffman [Hir95]. Kaufmann [Fie94]. Keeler [Ost84]. keeping [Tom84]. Kegan [Hoc92c]. Keith [Hen90a, Mar93c, Mar96a]. Kelly [Aga98, Hir95, Wil91c, YK94]. Kelvin [Mar93b, Smi92b, Tun92, Wil87d]. Ken’ichi [Mol94b]. Kenneth [Twe91a, Ver95, Hor97a, Smi96]. Kensington [FG96]. Kent [Jam90, Ost93b, Twe91d]. Kepler [Bra95, Met93, Dav95, Don94, Fie88, JK84, Ros86, Ste87, Ste94]. Kevin [Wil90a]. Key [Jor96, Kil94]. Keynes [Bro90e, Stu91]. Khwarizmi [vdW85]. Kidd [Cro84]. Kidwell [Twe93b]. Kilmister [Jor96]. Kimé [Bow93d]. King [Jac93, Mar96a]. King-Hele [Jac93]. Kingdom [But94, Por91b, AB89, HAB92]. Kingston [For90a, Rud92a]. Kipnis [Jam94d]. Kirkbride [Tom84]. Kirwan [Cro96c, GGSGB94]. Kisin [Aga94a]. Kitcher [Blo93c]. Kitson [Bra91]. klassieke [Bur99]. klassischen [Hin91]. Klaus [Hac93b, Sch93c, Eck92, Pos90, Wil87b]. Klein [Bro95a, War95a, War96a]. Klincksieck [Osb90b]. Klostermann [Kus95d]. Kluwer [Ams92, Bow94b, Bro96b, Cas94, Div90, Dol90b, Hac93c, Hel96b, Hen98b, Hod93a, Jam91, Jam92c, Mar96c, Mil94, Mul91, Non91a, Non93a, Nor94a, Rus91, Sch95b, Smi94d, Ste95, Tur91]. knew [Tor95]. Knight [Jam94b, Rob93]. Knopf [Bro94b]. Knorr [Whi90a]. Know [Gol94, CP93]. knowing [Pic93b]. Knowledge [Baa90, Bai91, Bai93, Bur93, Dol94, Ful98, Hac93c, O’B92, Ree90, The94, Twe91b, Uim95, Arm83, Bea91, Dol96, G^{+}09, GG91a, Gre92a, Hue88, Lig87, Lun94, Max84, Old86, Olm92, PR88, Rud85, Sir90, Vuc84, War86, WT82, Yeo93]. Known [Coo90, And82, SLT88, vPvP90]. Koenigsberg [Yav92]. Koffka [HK83]. Kolloquium [CM90]. Kolmogorov [Mar94a]. Kommentierte [Tur93b, ZH92]. Komposition [Gro89, Hin91]. Kongelige [Hal94c]. Königsberg [Ole91]. konservativer [Eck92, Mei91]. Kontroverse [New92, Fri91]. Korrespondenz [Eck92, HRP90]. Kostas [Hac93c]. Kovalevsky [Ken83]. Kox [Kra93a, War95a, War96a]. Kragh [Hen91b]. Krayer [Hel96a]. Kremer [Law92e, Van91]. Kremlin [Bad85]. Kriegsgeologen [Mac95]. Krige [Wri96b]. Krisciunas [Wil90a]. Kristallographie [Koe91, Sch89]. Kritik [Di 92, Jun89]. Kroll [Joh93, Ost93d, Wei90]. Kruger [Hod91]. Kuhn [Jor94]. Kuklick [Bow93a]. Kultur [Dau98, Fuc99]. Kumar [Har92c, Har97, Hil93]. Kurt [Jam94c, Wei90, Gre93b, HK83, Kus95c]. kurz [Alt91, Ger89]. Kusukawa [Met96b]. Kyushu [Mol94b].

L [Aga95a, Aga98, Bal91b, BV91, Bro94b, Bro91b, Car95, Cas93, Cle92, Cre91b, Cri90, Di 93a, Fin93b, Goo90a, Gou98b, Hac92b, Har90a, Hoc92a, Kni90a, Lan90, Lar96a, Law92e, Mal95a, Mar92a, Mar96c, Mol92, Mol94a, Mol94b, Mol98b, Mor92f, Mül93, Mur95b, Old97, Pal96a, Pal97, Pan91a, Per93, Rap92, Sch91a, Shr92, Smi96, Van91]. L. [Nic93b]. L.30 [Fin92]. L.75 [Sch91a]. Lab [Pic92a]. L’abbé [FM00]. laboratories [Goo90b, THBH92a]. Laboratory [Col93, De 96, Goo90a, Mar93c, Nye93b, Pal91b, And84, Cit90, Goo91a, Kuz87, Pal96b, HE92]. Lacan [Smi93c, CS92]. Lactic [Rus91, Ben90b]. L’Afrique [G^{+}09]. Lake [BB82]. Lamarck [BM82, Cor88a, Jor84, Lau87b]. Lambourne [Shr92]. lamp [BM84]. Lanaro [Pan91a]. Lancastrian [Rus86a]. Lance [Hor98]. land [FS84, Moy86]. Landmarks [GEMM83]. Landsat [Hug92b, Mac90a]. Langdon [Har98d]. Language [Cro96b, Sho93a, And84, Sut94, Tau93]. languages [Sla82]. Lanham [Ger95, Ost99]. Laplace [Cre91b, Des89b, Des89b]. Large [Hoc93b, GH92]. Large-Scale [Hoc93b, GH92]. Largs [Smi92b, STW90]. Larry [Gol93, Jam93a]. last [Buc85]. Late [Elm97, For90a, Whi90c, Ali86, Ern91, Gol91a, Gou97, Hun95b, May97a, Moy83, O’M89, SP88, Sho86]. later [Mil99]. Latin [Coo90, Lor90, Mol94b, Mon93, BEFH10, BM89, Cop92, EiT92, ET92, Tak92]. Latour [Alb92, Kus95b, Ver90, Ste91c]. Laue [ST94]. launching [Bru87a]. Laura [O’B96a]. Laurent [Cro94, Poi93]. Laurie [Kra91b]. Lavoisier [Cro94, Cro95, Cro96b, Cro96d, Gol95b, McE95, Smi93c, BV90, BV93, Ber93a, B^{+}95, CS92, Don93, Hol85, LF97, Poi93]. Law [Lar90, Ver90, Bur99, Har89, Sha83, Don94]. Lawn [Mol95]. Lawrence [Blu90, Gol91b, Bru96]. leading [Gun84, Pug84]. Learned [Smi99a, Gra98]. Learning [Hal94c, Hil91a, Ost93c, Sha96, BM84, Gol95a, Hal91c, Ped92b]. Leary [Kus97]. Lebensborn [Lil85]. l’École [FD87]. Lecture [Wes95, Hol86]. Lectures [Di 93f, Mor90e, BC98, KT87, Kra91c, Mag96b, NS84, OS92b, SO85, Sch01]. lecturing [Por95b]. Ledger [GL88]. Leeuwenhoeck [vLP82]. Lefanu [Jon91]. Legacy [Har92b, Hug98, BT96, Ste85, TP90]. Lehig [Kni92a]. Lehigh [Rus94]. Lehrstuhl [AH93, Hes95a]. Leibniz [Bro96b, Hal94a, HD97, Ait85, Ber93b, Cou95]. Leicester [Ben94b]. Leiden [Fie92, Hun98, McK96, Mol95, Osl97, Rag95]. Leigh [Law91b]. Leneaux [Mar96c]. Lenore [Jam92a]. L’Enseignement [CLR87, Ano87b]. Leo [Bal91b, Edg95, Gou98a, Mor96a, Maz98b]. Leon [Law92a, Smi93c]. Lesley [HD97]. L’essai [SG80, SG80]. Let [F^{+}88]. Lettere [Di 98, Bal91b, GGC90]. Letters [BKH90, Buc92, Hal94c, Hen93b, Hen94b, Hug92c, Hug96, Law91a, Ped92b, Sha96, Van91, vLvHH^{+}88, AJ^{+}89, CMNH92, FMF85, Ger89, GGC90, HD97, KBRS90, LM23, LHB04, LS76, LHH91, MH02, PvM85, Roc88, SGS00, SL87, WR92, FDLJ89, Gol95a]. Lettres [FDLJ89, Fin93c, HS91, Non94, Non93b, Roc88, SGS00]. l’Europe [vPvP90]. Levels [Smi94d, Bli92]. Levere [Hig94]. Leviathan [SSH85]. Levin [Jor94]. Levine [Rou97]. Levy [Rud92a]. Lewis [Lar96a, Sho99, Ash85, Har95b]. Leyere [Non91a]. Lhote [Fin93c]. liberal [GJ86]. liberalism [Fre90b]. Liberalismo [Goo91b, Fre90b]. Libois [Rus92]. Libraire [Pyl92]. Librairie [Out94]. Library [Bra92a, Elm91, Hac92a, Hen93a, Jam94c, Mil94, Woo95, And84, RWD90, Got92]. Libri [Pan91a]. licences [Lan90, WW88a]. Licht [Sho99]. Lick [Pan97, Wri87]. Liebig [Nig99c, Tur93b, ZH92, Bro97a, Mun98, ZH92]. Liège [Rus92]. Lies [Blo93a, Sap90]. Life [Aga96a, Edg90a, Hal91a, Har99b, Hod91, Hug92c, Hug98, Kra90a, Kra91a, Kra93b, Law92e, Olb91b, Pyl93a, Rou97, Sch95b, Spa91, Tur96, Ver94, War91b, Wis96, Ack87, Ald85, Ash85, A^{+}87, BT96, Bäu84, BB82, BC89, BH90, Bro89, Bud93, Can88, Can93b, Cha91, CB^{+}90, Cro86, Dic82, FB97, Fei90, G^{+}89, Gir86, Gol92a, Goo91a, GJ85, GL88, Gun84, Hey81, Hol85, Jak84, Jam92d, KBRS90, Kre90, Len82, Lev85, Lil85, Llo83, Mac85, Moo89, Pai82, Por95c, Ros90, SSH85, She95, Sou93, Ste85, Ste90a, Sut99, Tei91, dG01, vLP82]. lifting [Ole84]. Light [Alb92, Can90b, Can96, Jam94d, Mal94, Mar93a, Por91c, Buc89, Can83, FIF86, G^{+}09, Hak95a, Hal93, Kip91, Mir89, Roc88, Sch88, Sie91, Ste91c]. Lightman [Nig99a]. Like [Rus86b, Fis98a]. Lilian [Daw98a]. Liliane [Rus92]. Lillian [Kra91b]. Limbrick [Coo90]. Limited [Law91c, Law93d, Twe93a]. Limits [Twe90, Cer89, Lig87]. Linda [Por91b, Nic93c]. Lindberg [Nut93, Pum94]. Lindemann [Har98c]. Lindgren [Twe91c]. Lindqvist [Hes95b]. Linear [Gra85]. Linguistics [Sho93a]. link [Nic96b]. Linnaeus [Spa95, FL^{+}94]. Linné [Spa94]. Linnean [Wal88, GS88]. L’inventario [Olm92]. lire [Gou98a, La 94, Di 98, Mor96a]. Lisa [Fis98b]. Lissa [Bro93a]. list [WW75]. Literary [Sha92, Sho96a, Dea91, Smi94b, WM85]. Literate [Joh93, Kro91]. Literature [Di 98, HW81, Hoc91, Pet90, Por90b, Rou97, Fox87, Hay94, Maz95a, Sha83]. lithium [Joh84]. littéraire [Spa94]. Little [Ken83, vPvP90]. little-known [vPvP90]. Littlefield [Ger95, Hil91a]. Liverpool [Bra92b, Hal91b, Row90b]. Lives [Law98c, Man97, New94, SY96]. Livingstone [Cor93]. Llanymynech [Rob96b]. Lloyd [Bar92]. Loads [FS84]. Locke [Por90b, Ale85, Fox88, Sch01]. Locust [Kni98b]. Lodge [Hal91b, Row90b]. L’œuvre [dGB^{+}89]. logazione [Fin93b]. Logic [Mar94a, Non93a, KI92, PB82, PB83, Wal92, Mun98]. logical [Des89b]. logiche [Cre91b, Des89b]. London [Aga95a, Aga95b, Ans96, Ari90, Bai93, Ben94b, Bev90, Blo98, Blo93c, Blu90, Bow92a, Bow94a, Bow91b, Bow92c, Bow92b, Bra91, BD90, Bro93c, Bro99c, Bro91c, Bru91, Bru93, Bry91, Bur98, But94, Byn90, CK91, Cam90, Can90a, Can90b, Can91a, Col93, Cra92a, Di 93b, Di 93f, Di 93d, Edg90a, Edg95, Elm91, Far94, Far99, Fie94, For91, Fyf99, Get96, Goo90a, Goo96, Gou92, GG91b, Gra91, Gre95, Gro93, Hac90e, Hac92a, Hac92d, Hac93a, Hac94, Hak95b, Hal92c, Hal92a, Har93a, Har96a, Har93b, Har95b, Hay98a, Hen90b, Hen92b, Hen93a, Hen95b, Hir95, Hoc90a, Hoc92c, Hod90a, Hod90b, Hod92a, Hor92, Hor96b, Hor98, Hud92, Hug92b, Jam92a, Jam92b, Jam93a, Jam93b, Jam94a, Joh93, Jon92, Kat90, Kni92a, Kni95, Kra90b, Kra93b, Kus94, Kus96a, La 97, Law90c, Law91c, Law92e]. London [Law93g, Law93b, Law93d, Law96b, Luc92, Mar92a, Mar93b, Mar93c, Mar96a, McC95, McE93, McK97b, Mor90a, Mor92a, Mor92c, Mor92f, Mor99, Mur92b, Nic93a, Nic93b, Nic94b, Nut93, Nye92b, O’B92, O’B96b, Odd91, Osl98, Ost94, Pal91a, Pal97, Pau90, Por90b, Por95a, Pri91, Rob93, Rus91, Rus94, Sha96, Sho93b, Sho93c, Smi91b, Smi92d, Smi92a, Smi93a, Smi94a, Smi99b, Soy96, Twe91a, Twe91b, Twe91c, Twe92a, Twe92b, Twe92c, Twe93a, Twe93b, Ver90, Whi90d, Wil95a, Wri94a, Yav92, Des82b, Des84, Des89a, GS88, Gle91, Ili95c, Kno90, Mor90e, Mor95a, Por95b, Ric97, Rob96a, Sme97, Ste99, Wal88]. long [TBDA10]. Look [Cra92c, Roc90, Sch93a]. López [Goo91b, Mal98]. Lord [Fie93, Mar93b, Pai82, TC90, Tun92]. L’ordine [Des89b]. Loren [Smi99a]. Lorenz [Hod91]. L’organisation [HJF89]. Lorraine [Hod91]. Lost [Di 93b, Gre92b]. Louis [Ait90, Cro96c, Pal96a, Spa98b, Cla89, Gei95, GGSGB94, Nye92b, PR95, SG80]. Louis-Bernard [Cro96c, GGSGB94]. Louise [Hal92c]. Louviers [Cha84]. Lovers [Hal94c, Ped92b]. Low [Hac92b]. Lowell [Cro86]. lower [Bec95]. Löwy [Ams92]. Ltd [Bry91, Hac93a, Hor92, Jam92a, Law93e, Mar93b, Tur93a, Twe93a]. Lubbock [Bro93a, Cla97]. Lucas [Kni98b]. lucida [HA87]. Lucille [Smi92d]. Luckin [Nye92a]. Ludwik [Ams92]. Luigi [Bal91b, GGC90]. Luis [Goo91b, AT87]. Lull [Pri91, Per89]. lumières [Duc82, G^{+}09, Roc88]. lunacy [Sho86]. Lunbeck [Uim95]. luoghi [Fin93b, Olm92]. luyt [BD90]. lvi [Smi92b]. lviii [Bow92c]. lxi [Hod93b]. lxix [Law91a, Sch91a]. lxvi [HD97]. lxx [HD97]. lxxxiii [Mal96, Mon93]. Lyell [Bow92c, MK98, Tau93]. Lyle [Twe92a]. Lyme [Tor95]. Lynette [Lyn99]. Lynn [Kni90a, Pum90, Ras96a]. Lynwood [Nye92b].

M [Ait90, Blu90, Bro93b, Bry91, Buc92, Di 95a, DD95, Goo91b, Hal92a, Har99b, Jac96, Jon92, Kra91b, Kra97, Mal93a, Mar93a, Mil94, Nye99, Poo99, Sho86, Smi94a, Sta90b, Twe91b, War95a, War96b, Wis96, Yav92, Yos99]. M. [Blo82, Hor96a, Hor97b, Ili95b, Pic92b, Ras96b]. M.D [Wal88]. M.D. [Sta84b]. MA [Aga95b, Aga98, Bro98, Dea92, Edg95, Hes95b, Hir95, Joh96, Luc96b, Man97, McE95, Pal98, Pau90, Rou97, Twe90, Wit95a]. Maatschappijwetenschappen [Alb92]. Mach [Wol87, Wol87]. Machian [Fra90, Bar89a]. Machine [Aga98, Hor99a, dLT96, Mic98, Nye93b, Pyl93a, Bak96, CKA96, Cha91, Her88]. Machinery [Kus94, Ric92a]. Machines [Twe91b, Twe93b, MKW92, Mor95a]. Mack [Hug92b]. Mackenzie [Sta90b]. Macleod [Ink90]. Maclure [MD88]. Macmillan [Bow91b, Goo90a, Jam92b, Law90b, Law92c, Mor92f, Sho93c]. Maddin [Buc90]. Made [Kra94, TFC93]. Madison [Gre92c, Jon94, Mau95, Rou97, Sho96a]. Madness [Lar90, Har89]. Madrid [Goo91b, Mal96, Oli92, War90a]. Madure [Por90a]. Maffeo [Ben93]. Maffioli [Hac92b, Ili97]. Magazine [Fyf99, Dav99]. Magda [Por95a]. Magdeburg [Hel96b, vGA94]. Magi [Gol91a]. Magic [Hun90, Hut92, Kat90, Pyl93b, Adl87, Bry88, Gol91a, Hal90c, Rud88, She91, Tre97]. Magie [Gol91a, Gre93b].

Magie-Verständnis [Gol91a, Gre93b]. Magier [Gol91a, Gre93b]. Magierin [Gol91a, Gre93b]. magische [Adl87]. Magnànim [Mal98]. Magnetic [Hak95b, Cra93, Far95, Jor90, Zan86]. magnetism [Nye99]. magnetostatics [Jor90]. magnified [HS99]. Magnus [Zam92]. mai [Ano93q, A^{+}87]. Maienschein [Bow92b]. Maija [Bro96a]. Maila [Hoc92a]. Maimonides [Kel91]. Main [Kus95d, Wei90]. Mainz [Hel96a, Kra91c, Kra91c]. Mainzer [Pos90]. maison [Hil98a]. Maker [Ree90, PR88]. Makers [McC95, Rus94, Tur93a, McC92, Tra93, Wil94b]. Making [Aga94c, ABR93, Bro90a, FB97, Fis98a, Fyf99, Gou95a, Har99b, Hor96b, Jam91, Law96b, Mor92d, Rus91, Ben90b, Can94, Cla97, Dav99, Fri83, Goo90c, MH88, Mis95, Rei85, Wei82, Wil87c, Wil94c, Wri96a]. Malabaricus [Hen86b]. Malebranche [CMDM82]. Malet [Maz98a]. Malicki [Fie92]. Malina [Mur92a]. Malthus [Di 93a, Sch91a]. Mammoths [Mur96, Van93]. Man [Bro92b, Jam94a, Kra94, Law96b, McK96, Mic98, Spa95, App93, Bow90, FL^{+}94, Gar98, Hei86, dLT96, Mur94, PS84a, PS84b, TFC93, Wei82, Wil94c]. Man-Midwifery [Law96b, Wil94c]. Management [Mil95]. managers [H^{+}82]. Manchester [Bar90a, Har90a, Nye92a, Pal90, Sta90b]. Manelli [Fin92]. Manfrass [Sch93c]. Mann [Ver91]. Manning [Hac90e]. Mansell [Mar92a]. Manual [Twe93a, Bry92b]. manufactures [Ber85]. Manufacturing [Mor96b]. manuscript [TW87]. Manuscripts [Bra92a, Hal94a, Hen94b, Law93e, Ber93b, CF86, New89, NJ91, OS92a]. many [Her83]. mappa [La 94]. mapping [Her83]. Mara [War95b]. Marc [Kra90b]. Marcello [Dea92, Mor93]. March [Mar93c]. Marci [Fin92]. Marcia [Pal97, Bra90]. Marco [GG96, Lar96b, McE95]. Marcus [Fin92]. Marek [Mur92a]. Marelene [Hor99b]. Margaret [Bow91b, Hel96b, Hen90b, Can91a, Fre93, GR88, GG91c, Kni98b]. Margin [Nea98, Tay96]. marginados [Goo91b, Rey90]. Marginalia [Hod93b, DDG90]. Marí [Mal96, McC95]. Maria [Joh96, Law93g, Mal98]. Marie [Cas93, Gir86, Ost93c]. Marietti [Fin93a]. Marika [Luc96b]. Marilyn [Hud92]. Marina [Law92d]. Mariner [Aga96a, Chr95]. Mario [Hod93b]. Marjorie [Hen93b]. Mark [Aga94c, Lar96a, Jon94, Law96c]. market [Pri83]. marketing [Far95]. Markets [Kla95]. Marlene [Bro91c]. mars [Ano82a]. Marsh [Lar96a, CMN95]. Marshall [Bai91, Kra98]. Martha [Mar96a]. Martin [Bow92c, Sha93a, Twe93b, Aga98, Law93f, War96a, Wil91c]. marvels [OH85, PP82]. Marx [Edg95]. Mary [But94, Di 93d, Ger91, Hal92c, Har98c, Kni95, Pat83a, Tor95]. Maryland [Hen94b, Hil91a]. Masao [Bro92c]. masked [Gar98]. Maskelyne [Cha90c, How89]. Mass [Bev90, Blo93c, Bow92a, Bra92b, Buc90, CK91, Di 93b, For91, GG91b, Gra91, Hal92d, Hoc90a, Hud92, Hug92b, Kat90, Kni90c, Kra93b, Law91a, Law93g, Luc92, Nye92b, Olb91b, Pyl93b, Smi93a, Twe91a, Twe91b, Twe91c, Twe92a, Twe92b, Twe92c, Twe93b, Whi90d, Row90a]. Mass-Energy [Bra92b, Row90a]. Massachusetts [Hod90a, Hor95]. Masson [Bow95]. master [Sch93b, Sch92b]. Mastery [Fie94]. Matematika [KI92]. Matematiques [Maz98a]. Material [Hen93b, Ili95c, Joh93, Mol98a, CK90, Fre95, Kro91, SW93]. Mathematica [Met96a]. mathematicae [Viè83]. Mathematical [Cri90, Cuo95, GG91c, Hil98a, Mar94a, Maz96, Maz98b, Ric88, Ben82, Bla93, Cor96, DR95, Had94, HF87, Hil98b, KI92, LS76, Sch90a, LHB04, LHH91, vL76].

Mathematician [Smi94a, Can93b, Kno90]. Mathematicians [Fau96b, WW93, Gin88, Ros86, WW75]. Mathematics [Joh94, Mar94a, Mol96, Rav99, Whi90c, Fie97, How82, Høy94, Loa98, O’M89, Ric97, Sch89, WW86, Wil93, KI92, Kra91c]. Mathematik [Hel96a, Koe91, Sch89, Kra91c]. mathématique [Bla93, Hal95, Hil98a]. Mathematischer [HD97, LS76, LHB04, LHH91, vL76]. mathematized [Coh91]. matrix [MR82b]. Matter [Aga97a, Bro99b, Fau96a, Kni95, Kni98a, Kus95a, McE95, McK97a, Mor92e, Mur95a, Smi90b, Sta90a, Ste97b, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ber93a, Bro85, Bro95b, Har82a, Nye93a, Pai86]. Matthew [Dem83, Pic93a]. Matthews [Mag96a]. Matthias [Mor96a]. Maulitz [Fox92]. Maurice [KW87, Mor95c]. Maurizio [Ste95]. Mawe [Dol98a]. Max [Kra91b, Hei86]. máximos [GB95a, Mal96]. maximum [GB95a]. Maxwell [Buc92, Mar93a, Buc85, Gol83, Har96a, Har98e, Hen86a, MH02, MGBE95, Sie91, Yos99]. Maxwellians [Har93a, Hun91, OP87]. May [Ano92p, Ano93q, Di 93f, PCA91, Pic99, A^{+}87, OS92b]. May-December [Ano92p]. Mayer [Smi96, Can93a]. Mazhar [Di 98]. Mazumdar [Jac96, Jon92]. McConnell [Tur93a]. McCready [Lar96a]. McCurdy [Wri94a]. McDowell [Mol94a]. McGill [Bra90, For90a, Rud92a]. McGill-Queen [Bra90, For90a, Rud92a]. McGilvray [GG91b]. McGuire [Hen98b]. McKay [Twe91c]. McNeil [Hor98, Smi91b]. MD [Ger95, Ost99, Smi90a]. meal [Aga94c]. mean [HF87]. Meaning [Boh93, Hal90a, Jam91, Soy96, Coo84, Fri91, Goo90c, Ric92b, Smi92e]. Meanings [Luc92, Nye90]. means [Mun98]. Measure [Mol96, Høy94, Wil87a]. Measurement [Hac92a, Hac93c, Bar93a, Cha90d, Cha94, GG91a, Goo90b, Hun96, Mar92b, Sti86]. Measurements [Mar93b, Tun92]. measures [Con87]. Measuring [Mil95, Van85, Smi94e]. Mechanical [Hil91c, Ole84, Tre97, Whi90b]. Mechanics [CK90, Har96a, Hil91b, Hen84b, Len82, MGBE95, MR82a, MR82c, MR82b, MR82e, MR87a, MR87b, MR00]. Mechanistic [Fis98b, SG96]. medecine [Pyl91]. Medica [Spa98a]. Medical [Bru96, Har99b, Lie87, Mag96a, Olb91a, Pic93a, Por91b, Por95b, Pri93, Arm83, BP02, Cle93, C^{+}95, Deb91, DPP88, FB97, Gos87, Hue88, Jac95a, Mat95, Nic88, Pes84, Ram88, Sir87, War86, WT82, AB89]. medicinal [Goo91b, ML89]. Medicine [Ams92, Blu90, Cra92a, Gre92c, Har87, Har90b, Har93b, Hen93a, Hen95b, Hen99, Hod92a, Jam94c, Lar90, Law90a, Law90b, Law91c, Law92c, Law93g, Law93d, Lyn99, Mol95, O’B92, Pal97, Pau90, Pic93a, Spa98a, Tim98, Aly90, AL87, Bal91a, BP93b, Coo88, Des89a, FL88, Got92, Gou96, Grm90, Har89, Har92e, HH97, Law93a, Mor91a, Pag82, Pic93b, Pro88, Ram88, Ric91, Sir90, Sir97, Sta91, VoC89, WT82, Grm90, AL87]. Médico [War90a, Hue88, Rey90]. médico-sociales [Rey90]. médicosociales [Goo91b]. Medics [Lan90, WW88a]. Medieval [Cam90, Gou92, Hen96, Hil93, Mol92, Mol94b, O’B92, Osl98, Sir90, Bea91, BJ91, Cro90a, Cro96a, EiT92, ET92, Gra94, HW87, Hil84, Kno90, Kre82, Tak92, Wal81]. médiévaux [Bea91, Ung92].

Mediterranean [Cam90, HW87]. Megill [Kus96a]. Meiji [Bro90c, MR89]. Meinel [Eck92]. Melanchthon [Met96b, Kus95g]. melanogaster [Pal96b]. Melbourne [Aga94b, Bro95d, LLDM94]. Melhado [Bro93b]. Meli [Hal94a]. Mellon [Har98d]. Melodie [Mar93c]. Members [McC98, Ano97d, Bro97d, Stu95]. Memoir [Sec96, MS95]. Mémoires [Gol95b, Woo95, BV93]. memoirs [BV93]. Memorial [Bec99]. memorialize [Gay98]. memoriam [HC91]. Memorie [Di 98]. Memories [Pug84]. memory [Mac95]. Men [Jam92c, Mur96, Van93, Mac95, Ros91]. Menachem [McN92, Yeo92]. Mendel [Ore84]. Mendelian [SP98]. Mendenhall [Bro90c, MR89]. Méndez [Fre90b, Goo91b]. Mengal [Bow95]. Menge [Mol94a]. Menlo [Mar96a, EJR94, EJR95]. Menschen [Ver91]. Menso [Mol94b]. Mental [Kus94, Ric92a, Jac95a, Lar91]. mentalities [Vic84]. Mentally [Bis98, Nol95]. mente [Fin92]. Merchants [Cuo95, Had94]. Mercury [Ros82]. Méridiens [Osb90b]. Merle [Mol94a]. Merrill [Kni90a]. Merrit [Edg95]. merry [Söd86]. Merton [Hen92b, Ost94, CDS90]. Meryon [Pal97, EE95]. Mesmer [Hak95b, Cra93]. meson [BR91]. message [SB09]. messager [GP92, Non93b]. Messenger [Ari90, Lar96b, Ber95, GP92, GV89]. Messina [Smi93b]. Metals [Buc90]. Metaphor [Law92a, TC91, You85]. Metaphors [Kus97, Lea90, Lea94]. Metaphysical [Gro93, BGH^{+}87, Fle98, Gar92]. Metaphysics [Har82b, Hen98b, Fur88, McG95, vLMB88]. Metascience [Kni92b]. Metchnikoff [Law92a, TC91]. meteorites [LLDM94]. Meteorology [Aga97b, Kra90b, Fri89, Neb95]. Method [Cla92c, Mal94, Gro91, Kol92, SY86, Sha93c]. methodology [Old86]. Methods [Bar92, Div98, Llo91, RR90, Col87, Fox95, Max84, Tau93]. Metres [Cal90, Hac93a]. Metropolis [Bur93, Baa90]. metropolitan [LLDM94]. Mettrie [Mic98]. Metuchen [Byn90, Law93e]. Metzger [Hal90d]. Micale [Law96c]. Michael [Bro93c, Bru96, Gau99, Goo96, Hil91a, Jam92b, Jam93b, Kni92b, Kni98b, Mur92b, Ost98, Ost99, Shr92, Twe93b, War91b, Wil91b, Ans96, Bro99c, Can91b, Dea93, Elm97, FTG91, Fie91, GJ85, Hef96, Hen94b, Hoc92b, Hun98, Jam92d, Kus95d, Law90a, Law98c, Lee90, Mur95b, Nic94a, Nig99b, Pum91b, Pum91a, Sec96, Sec98, Twe91c]. Michel [Gol95b, Jor91, Gut89, Hal95, Sho98]. Michela [Pri91]. Michelle [Cro96c]. Microchip [Hod90a, Que88]. Microfiche [Hug96, WR92]. microfiches [Hug96]. Microfilm [Hen94b]. microfilms [Hen94b]. microphysics [Buc85]. Microscope [Coo96, Hac92c, Goo91a, Wil95b]. Microscopical [Hac92c]. Microwaves [Hac93a, Cal90]. Mid [Hor99b, And92, Mor95a, RCRC98]. mid-eighteenth-century [Mor95a]. mid-nineteenth [And92]. Mid-Twentieth [Hor99b, RCRC98]. Middle [Ano82c, Bec95, Gol91a, MM10, Mol95, NW82b, Sor83]. Midst [Bis98, Nol95]. Midwifery [Law96b, Wil94c]. Migration [Har98h, AS96]. Mikulás [Mar94b, Har92b]. Milan [Pan91a, Per93, Rap92]. Military [Edg91a, Mac95, Smy89, NPYGDL86]. Milky [Het95, Pau93]. Mill [Mun98]. Millar [Hen90b]. Miller [Dol98b, Dol99, Sec96, Sec98, MH88, MS95, Sho96b]. Mills [Bro91b]. Milton [Stu91, Slo85, Swa92]. Mind [Blo93c, Har90b, Kre96, Law91b, Por97, Soy96, Fig99, Har87, Kit92, Smi92e, Sta89b, Tur94]. Minded [Bis98, Nol95]. miner [AHY90, Win93]. Mineralogie [Sch84]. mineralogy [Lau87a, Sch84]. Minerva [Por97, Ili95a]. mines [AHY90, HS91]. Mineur [HS91, AHY90]. Minister [Por91a, Bru88a]. Ministère [Dav94]. Ministerio [Goo91b]. Minkowski [SO85]. mirabilium [OH85]. Miracle [Swa92, Wai90]. Mirko [Fox92, Pyl91]. Mirowski [Ger95, Kla95, Por91c]. Mirror [Hen99, Sir97]. Misa [Hor96b]. misfortune [dCC09]. Mishneh [Kel91]. misogyny [Out91]. Mission [Har95b, DD94, Pye93]. Misunderstanding [IW96, Pic98]. Mitchell [Har98h, Kus96b]. mitigation [MJ89]. Mitman [Nic93a]. Mitscherlich [Bro94a, Sch92b, Sch92b]. Mixed [Hor95, Seg94]. mobility [BC90]. Mocchi [Fin92]. model [Jor90]. modèles [Duc82, Neu92, Sin91]. Models [Zan86, Sin91]. Moderado [Goo91b, Fre90b]. moderate [Fre90b]. Modern [BBBM90, Bow92a, Coo96, Cor96, CW93, Cuo95, Far99, Fox92, Gou92, Hal92a, Hen92b, Hen95a, Hor96b, Hor97a, Jam94a, Kla95, Kra90b, Kus95b, Luc98, Mal93b, Mal95a, Maz98b, Mil94, Osl98, Pri93, Pum92, Smi90a, Smi93b, Uim95, AH85, App93, Asp90, BA91, BR95, BR96, Bru87a, Bru88b, CDS90, Cro90a, Cro96a, DeV82, Deb91, Eis83, Eis93, FJ93, Fin91, Fri89, Gou95b, Had94, Hea87, Huf93, Huf95, KT87, Kle94, LeG88, Liu92, Lun94, Mac95, Mis95, O^{+}90, Old94, Olm92, Poi04, Poo98, Rus86b, Ste99, Wei83, Wil95b]. moderna [Fin93b, Olm92]. modernes [Poi04]. Modernist [Ros94a, Smi95b]. modernity [Pur99]. modernization [Pri83]. modes [BB84]. modifications [MR82b]. Moewus [Blo93a, Sap90]. Molecular [Blo93a, Bug96, Mar93a, Wri94b, Sap90, Sie91]. molecules [MGBE95, Har96a]. Molland [Lor90]. Moller [Jam94c]. moment [Man93a]. Moments [Kra91a, Ros90]. Monardes [Goo91b]. monde [Bla93, Hal95]. mondo [Fin93b, Olm92]. money [Poi04]. Monographien [Hin91]. Monographs [Nic94a]. monopoly [THBvdB92]. monsters [PP82]. montagne [Ano93q, Dav94]. Montbard [BG^{+}92, Out94]. Montesquieu [Ano07]. Montgolfier [Gil83]. Montreal [For90a, Rud92a, Chu91]. Montréal-Paris [Chu91]. monument [Wri87]. Moon [Fie92, Hen91a, Kol91]. moon-tests [Kol91]. Moorcroft [Ald85]. Moore [Bro93c, Joh94, Kra90a, Wil93]. moral [Liv99]. Morali [Di 98]. Moralism [Hun90]. Moralizing [May97b]. Moran [Gre92c]. Morano [Mül93]. Mordechai [Hal91a]. Morelli [Di 93c]. Morelon [Bur98]. Morgan [Ger91, Ric97, Ros94b]. Moritz [Gre92c, Mor91a]. Morphological [Boh93, Ric92b]. Morphology [Hod92a, Ras96a, Des89a, Hun82, Nyh95]. Morris [Gre91, Hac92d]. Morrison [Edg91a]. mort [BG^{+}92, Out94]. Morveau [Cro96c]. Moscow [Hug92c, KBRS90]. Mosquito [HS99]. Moss [Pum95a]. most [Cha97, Cha83]. Mothers [Lyn99, HH97]. Motion [Fra90, Pyl93b, Bar89a, Dim93, Mai90, Sch95a, She91]. motive [CF86]. Mott [Bai93]. motu [Mil94, BJB92]. mountain [Ano93q, Bry88]. move [Dra81]. Movement [Kus95e, McG84, Nol94]. movers [Mar92b]. moving [Blo94a]. MRC [Por91b, AB89]. Mulino [Cre91b, Fin93b]. Müller [Twe91c, Lin90]. Mulligan [Har95a]. Multiple [Can90a, Cau89]. multiplicatione [BL83]. München [CM90, Fuc99]. mundo [GB95a, Mal96]. Munich [Bro94a, CM90, Met93, Sch93c]. Munksgaard [Hal94c]. Murchison [Smi91a, Sta89a]. Murders [Har89, Lar90]. Murdin [HD97]. Murray [Kni90b]. Muscular [Pal97, EE95]. muse [FD85]. Museum [Ben94b, Bow93c, Bro94a, Bro91c, Bro95d, CM90, Di 93f, Di 93d, Gou91, Hac94, Hor92, Jam92a, Jam93b, Kni92b, Rai91, Bud97, Des82c, Gun84, LLDM94, McC81, Win91].

Museums [Ben94b, Ben97, For90a, Whi97, And92, But92b, SP88, Won93]. Music [Bra95, Gou92, Gou96, Coh84, Cro90a, Ste94]. Mysterious [Jam90, Twy88]. Mysterium [KA81]. Myth [Hil98a, Hod90b, Bow88, FD85, Kol92]. mythes [Hil98a]. mythical [BM82].

N [Aga98, Bro91a, Cor93, Dol90a, Hal92a, Hod93b, Jon93, Mar94a, Mar96c, Pau90, Yeo96a]. N. [Slo97]. N.J [Byn90]. Nachträge [vLvHH^{+}88]. Nachum [Poo99]. Nahum [Jam94d]. Nakayama [Hoc92c]. Nancy [Hen99, O’B92]. Nantucket [Cro96b, Spa95]. Napier [Twe92b, Twe93a, Bry92b]. Naples [Mül93]. Narcotism [Law93d, SE91]. narrative [Söd86]. narratives [Hes99]. NASA [Bro91a, Wri94a, McC93, Smi89]. Nathaniel [Liv87]. Nation [Hor96b, Mis95, Söd86]. National [Dav94, Gro95a, Hir95, Jam94c, Mar94b, Stu98, vH92, Ano82a, Ano93q, May97b, PT92, YK94, Lil85, Pum91c, Web88, Web98]. Nationale [Dav94]. nationalised [H^{+}82]. Nationalism [Cra92b, Hoc93a]. nationalsozialistischer [Lil85]. Nations [Hil93, Ros84b]. Natur [Gol91a, Gre93b, Hei93, Kus95d, Mai88, Pos90, Ver91]. Natur- [Mai88]. Natur-und [Pos90]. natura [Di 93a, Ern91, Fin93b, La 94, Mül93, Olm92, Per93, SSH85]. Natural [Ait90, AH93, Alb92, Bai93, Bow93c, Bro96a, Bro91c, Bro95d, Di 93b, Di 93f, Dol94, For90a, Get96, Gol93, Gre92a, Ili98, Kla95, Kni90a, Kni98b, Mai88, Mal95a, Met96b, Met99, Mor90a, Mor90e, Nig99b, Rai91, Sha93a, Slo97, Smi93a, Spa98b, Whi97, Yeo96b, Yos99, Bru87b, Cav96, CS94, Cla89, Clu88, Dem83, Dol98a, FJ93, Fin93d, Fre94, FC96, Gay92, Gol91a, Gou96, Gow83, Gre92b, Gro89, Har82b, Har98e, Har98f, Hat90, JSS96, Kal95a, Kus95g, Mar92c, Mer89, Mil99, Mor95a, Nor89, PR95, Pow96, Sec94, Sha83, SP88, Ste91c, Ste92, SW95, Ste99, Won93, Yeo93, AH93, Fyf97]. natural-philosophical [Gro89]. naturale [Mül93]. Naturales [Gro89, Hin91]. naturalism [Cla97]. naturalist [Dem83, Jon88]. naturalists [Reh83, Söd86]. Nature [Bro93d, Can96, Cit90, Cuo95, Di 93b, Di 93d, Dol98b, Dol99, Fie94, Get96, Goo91a, Gre92b, Har92a, Hef96, Hen98b, Jon93, LS90, Ley90, Mai88, Nic93a, Nic94a, Non91a, Osb90a, Osb94, Osl97, Osl98, Ost98, Pal91b, Por91c, She99, Sta90b, Ste91b, BL83, BDH96, Bro85, BC98, Coh91, Cro96a, Deg91, Ern91, Fre94, Fur89, Gol91a, Had94, Hak95a, Hei93, Mac88a, McG95, MR96, Mir89, Mit92, Mor96b, Olm92, Ped92a, PS84a, PS84b, Sho93d, Sho93e, Sta84a, Win91, Wyl89, You85, vR95, OH85]. naturelle [Atr86, Bow93d, Gay92, Woo95]. Naturhistoriska [Lar96b, Ber95]. Naturmagie [Gol91a, Gre93b]. naturnal [Atr86]. naturphilosophische [Gro89, Hin91]. Naturwissenschaft [AH93, Gol91a, Gre93b, Hes95a, AH93]. Naturwissenschaften [Eck92, Hes95a, Hug96, Tur93b]. naturwissenschaftliche [Dau98, Fuc99]. naturwissenschaftlicher [HD97, LHB04, LS76, LHH91, vL76]. Naumann [Sho93a]. Naval [Hac90e, Man88]. Navarro [Mal98]. navegar [Lan86]. Navigation [Hac90d, Ben87]. Navy [Hac84]. Nazi [Wei94, Geu92]. Nazionale [Bal91b, Cle92]. Nazis [Pau90, Pro88]. NC [Kus96a]. NE3 [Fau96b]. near [CHG91, Sho93c]. nearly [Adl87]. Nebeker [Aga97b]. Nebulae [Aga96a, Chr95]. needed [Lan87]. Needham [Dav97, NR86]. neglect [Fra86, Rus93]. Negotiating [Cro96b, Hil98b]. Nehemiah [Jon91, Le 90]. Neil [Har96c]. nella [Fin93b, Olm92]. Nelson [Lar96a, NN95]. Nemesis [Chr93]. Nemore [Mol92, BJ91]. Neptune [Fri91].

Neptunismus [Fri91, New92]. Netherlands [McK96]. networks [Pri83, Rus99]. Neue [Bro94a]. Neumann [Bow92a, Twe93c, Asp90, Pou93]. neurasthenia [Gos87]. Neurath [Rei94]. neuroscientific [CJ87]. neutron [Six87]. neuzeitlichen [Bro95a, Kle94]. Neve [Bru96]. Never [Kus95b, Sch95a]. Nevil [Cha90c, How89]. Newberry [Hen93a, Got92]. Newborn [Hor99a, Bak96]. Newbury [Har93b]. Newcastle [Fau96b, Lan90]. Newcastle-upon-Tyne [Lan90]. Newman [Man97]. Newton [Bra90, Bra92a, Bra92b, Bra98, Cas93, Cas94, Hal91a, Hal94a, Hen94b, Hen98b, Mal93a, Mal94, Man93b, Ost93a, Wil91b, Ano92n, Bec91a, Ber93b, Bur83, Can83, De 95, Dob91, F^{+}88, Fei90, Gje86, Gue81, Hal92b, Hal93, Kol91, Kol92, Man93a, MT83, McG95, Mur85, NS84, New89, NJ91, Row90a, Sep88, Sha93c, Sno99, Vic87, WHS92, dG01]. Newtonian [Gol93, Bec82, Bla83, For85, Ste92, Wil87a]. newtonienne [Bla83]. Nicholas [Ben94b, Ste91a, Mor90a]. Nicodemite [Sno99]. Nicolaas [Bow91b]. Nicolas [Fin93c, Goo91b]. Nicolas-Claude [Fin93c]. Nicole [GG90, OH85]. Nicolson [Hen93b]. Niels [Hen92a, Kri91, Aas90]. Nier [Mar93c, Mar96a]. Night [Whi90d, Sch88]. Nihil [Coo90, SLT88, Sut94]. Niigata-Shi [Por95a]. Nikolaos [Div90]. Nikolaus [Hin91]. NIMBY [Wel93]. Nine [Ben93, Maf91, Viè83]. Nineteenth [Bro95d, Can90b, Di 93b, Di 98, For90a, Har90b, Hud92, Jam92b, Jam92c, Law92b, Sho96a, Wil97, And92, Cre91a, CC94, Fis94, Ful91, Gou97, Har92e, Luc96a, Ogi86, Pic90, Sec94, Wri96a, Ali86, Buc85, Buc89, Can91b, Cha82, CJ87, Coo84, Dem83, Dol98a, Eng83, Fin93d, Gol87, Gre82, Gre92b, Hal84, Ham90, Har82a, Har87, Kni86, Kni88, Len82, Maz95a, Moy83, Pri83, Reh83, Roc84, Ros91, Sch88, SP88, Sho86, Smi94b, Tur83].

Nineteenth-Century

[Bro95d, Di 98, Har90b, Jam92c, Sho96a, Fis94, Ful91, Gou97, Har92e, Pic90, Sec94, Wri96a, CJ87, Coo84, Dol98a, Eng83, Fin93d, Har82a, Har87, Maz95a, Pri83, Reh83, Ros91, Sho86, Smi94b, Tur83]. Nishimura [Por95a]. NJ [Dea93, Nic96a]. No [Aga98, Bro94a, Bro96a, Bro92c, Bur93, Cas93, Gay98, Goo91b, Gou95a, Hac93b, Har92d, Hen94b, Kat90, Kni92b, Kra98, Lar96b, Law90a, Mal96, Mal98, McC82, Met99, Mol94b, Non94, Por90b, Rou97, Sha92, Sho93b, Sho93c, Spa94, Twe90, War90a, Wat91, Whi97, Bro91c, Hal90d, Hen94a, Hir95, Tha91, YK94]. Nobel [Cra92b, Hoc93a, Cra84, CHU87]. Noel [Hor97a]. Noether [vdW85]. Nola [Dol90b]. Noll [Bis98, Kus95e]. nombres [Bea91, Ung92]. nominators [CHU87]. nominees [CHU87]. Non [Hod90b, Bow88]. Non-Darwinian [Hod90b, Bow88]. Nonostante [Di 93a]. Nonstandard [Maz96, DR95]. Noor [Di 98]. Noordwijk [Wri96b]. Norberg [Aga98]. Norbert [Aga95b]. Nordic [Luc96b]. Normative [Smi93a, Hat90]. Norst [Bro91c]. North [Bis98, Her95, Old97, Wil95a, Col87]. Northeastern [Hoc91]. Norton [Ras96b]. nos [Ano87b, CLR87]. note [Ano91q, Ano94n, Dim93, Top97]. notebook [MT83]. notebooks [BGH^{+}87]. Notes [Bro90c, For91, Jam93b, Mon93, Wil90b, Ari90, BL83, CLR87, Coo90, Cop92, FTG91, MR89, Met93, Non93b]. Nothing [Coo90, SLT88]. Notion [Osb90b, Ost98, BBC88, BDH96]. Notre [Bro93d, Non91b, Sha97]. Noulens [Lit91]. nouvelle [BN82, Non94, dGB^{+}89]. novâ [Viè83]. novel [Tyt82]. Novelties [Mos93, Pum95a]. Novelty [Smi94d, Bli92]. Novikov [Aga94a]. nuanced [Lee96]. Nuclear [Hoc90a, Hor96a, Hor97b, Kri91, Wea88, Aas90, Wel93, Wyn82, Zan86]. nuestras [Goo91b, ML89]. Nullius [Sut94].

Number [GG91c, Mar94a, Mol92, Mol96, BJ91, Høy94, KI92, Sch90a]. Numbers [Bro94b, Lar96a, Pyl93b, Wat91, Wis96, Bea91, Cru90, Por95c, SO85, She91]. Nuncius [Ari90, Mor96a, GP92, Non93b, SB09, GV89]. Nursery [Hor99a, Bak96]. Nutrition [Spa98a, KC95]. Nutton [Bru96]. Nye [Kni95, Luc92]. Nyhart [Ras96a].

Oakley [Twe91a]. Obervatory [Maf91]. obituaries [WM85]. Obituary [Cha93, Dav97, Day99, Dun92b, Gra92, GG93b, Hal94b, Pic99]. Objectivity [Kus96a, Kus96b, Wis96, Ash95, Het88, Meg94, Por95c]. Objects [Jam93b, FTG91]. Observation [AH85, Jam91, Goo90c]. Observatory [Ben93, Bro93d, Moo92, Sha97, Wil90a, Pan97, Way87, Wri87]. observed [Sto83]. Observers [Sto83]. occasions [MJ89]. Occidentales [Goo91b, ML89]. Occult [Ait90, Cla89, Gre92c, Vic84, Mor91a, Noa99]. occulte [Bal91b, GGC90]. ocean [Men86]. oceanographic [McC82]. Oceanography [Bro91b, McC81, Mil89]. October [CM90]. Odense [Hac95]. Odyssey [Maz96, DR95]. oeuvre [Non94]. Œuvres [VB80, LF97]. Öffentlichkeit [Dau98, Fuc99]. Office [Cro96c, Gol91b, Hac90e, Man88]. Officers [Bro97d]. official [Smy89]. officiels [CLR87]. Offray [Mic98]. Ogilvie [Hud92]. O’Hara [HD97]. Oktober [CM90]. Olaf [Bro93d, Hal94c]. Olaus [Wit95a, Eri94]. Olby [Hal92a]. Oldenbourg [Bro94a, Fuc99]. Olesko [Yav92]. Oliver [Hal91b, Osb90b, Row90b]. Olmi [Fin93b]. Olschki [Bal91b, Gou98a, Mor96a]. Olson [Pum92]. O’Meara [Whi90c]. One [Div90, Hea86, Hea90, War90c]. O’Neill [Aga98]. Onnes [Hac93c, GG91a]. Ontario [Hen98b, BB82]. Ontogenèse [Bow95, CM93, MC^{+}93]. ontogenesis [MC^{+}93]. Open [Bee96, Edg90a, Law98b, Ost93b, Stu91, PB82]. opened [HJF89]. Oppenheimer [Di 95a, Edg91a, Yor89]. Optical [Can90b, Buc89, NS84]. Opticks [Mal94, Hal93]. Optics [AS89, Can83, Can96, Gou92, Hak95a, Smi92a, Cro90a]. Opus [Viè83]. Oral [May99]. Orbits [Jac93, KH92]. Orbs [Hen96, Gra94]. Order [Alb92, Bai91, PS84a, PS84b, And82, Des89b, Ste91c]. ordinary [Law90c, Tro89]. ordine [Cre91b]. Ordoñez [Oli92]. Oresme [OH85]. Orfila [War90a, Hue88]. Organic [Pyl93a, THBB92a, Bur87, Cha91, TT86]. organisation [GG90]. Organization [Hal92d, Coo84, HJF89]. Organizational [Wri94a, McC93]. organized [CM90]. origenes [Lan89]. Origin [Fox92, Hay98a, Mor94, BP95, BR96, Bro84, PB99, Pro93, Wus84]. Original [Bra92a, Daw98a, Jam94c, NR86]. origines [Goo91b]. Origins [Blo93a, Bow92a, CW93, Duh91, Har96d, Hor99a, Jac90, Kus94, Kus95c, Kus95e, Law92a, Mar96c, Nic96a, Rus94, Asp90, Bak96, Ben86, Cat88, Cha88, CJ87, Dan90, Dic82, Eng84, Gro95b, Hag92, Hul99, Lig87, Lan89, Mag96b, Nol94, Pic87, Ric92a, Sap90, Sar87, TC91, Tra93, Wei82, Wis85, Atr86]. Ormerod [Cla92b]. oroscopo [Bal91b, GGC90]. Orthodoxy [Elm97, Hey83, Hun95b]. Osborn [Bow93c, Rai91]. Osborne [Hef96]. Oscar [Di 93e, Wei91]. oscillography [Phi87]. Osiris [Aga95a]. Osler [Fre93]. Ostwald [Mor92d]. Other [Law91a, Mic98, Smi98, Ste99, CS94, GH84, GH96, Gun84, Hol86, dLT96, Mun98, Noa99, Opp85, Yeo96b]. Otté [Roh99a]. Otto [Bro92c, Cam92, Hel96a, Hel96b, Kra91c, Hel96b, Kra91c, Rei94, vGA94]. ottocento [La 94]. our [Ano87b, Ben90a, Bis98, MJ89, ML89, Nol95, Whi88]. outline [Fis94]. Outram [Law96a]. outsider [De 92]. ouverte [GG90, HJF89]. Ovary [Har98b, Cor97]. Overmier [Law93e].

Overseas [Har95b, Pye93]. Owen [Di 93f, Twe91a, GT92]. Oxford [Ait90, Ben90a, BV91, Bro92b, Cas93, Cla92c, Cor93, Cro95, Gou91, Hac90d, Hal94a, Har93c, Har96c, Har97, Hen92a, Hen93b, Hoc92d, Hut92, Jam94b, Jon93, Kni90a, Kni90b, Kus98, Lar90, Law90a, Law90b, Law92a, Law92d, Law93g, Law98b, Law98a, Lee90, Ley90, Mac90b, Mal94, McN92, Mol98a, Mor91c, Nic94a, Ost93a, Pic92b, Por91b, Pyl93a, Rav99, Ree90, Sec98, Smi91c, Smi92c, Smi93d, Str90, Stu98, Swa92, Twe93c, Whi90c, Wil91c, Yeo92, Ben97, Eng83, Gou96, PB^{+}89]. Oxfordshire [Gou91, Har99a, Bee89]. Oxon [War91a].

P [Bro96b, Di 93f, Di 93d, Fau96b, Hor95, Hor99a, Hug92c, Mar94a, Mon93, War96b, Wil94a]. P. [Buc92, Col93, DD95, Hac92d, Lan90, Mal93a, Nye99, Yos99]. P.L.S [Wal88]. Pa [Por91a]. Paget [Law91c, Rob89]. Pais [Hen92a, Mal98, LN95]. palaeontological [Cle83]. Palaeontology [Des82b, Buf87, Des84]. Paléontologie [Lau87b]. Paleontology [Bow93c, Rai91, Lau87b]. Paley [Fyf97]. Palingenesia [Hin91]. Palm [Hac92b]. Palmer [Twe92a]. Pamela [Gre96, Hug92b]. Pamphlets [Lar96a, RT95]. Pan [Jam94a]. Pancaldi [Di 93c]. Pandemic [Fox92]. Pantin [Non93b]. Panza [GG96]. Paola [Cre91b, Nor94a]. Paper [NhTWL87, Twe90, Edg93]. Paperback [Edg90a, Jam94a, Osb90b, Aga95a, Ari90, Bis98, Blo93a, Blo98, Boh95, Bow91a, Bow92c, Bro90c, Bro99c, Bru96, Bug96, Bur93, Can90b, Cor93, Cra92a, Dan92, Daw98a, Di 93f, Edg91a, Edg95, Gol91b, Goo97, Goo91b, Gre92c, Gro93, Har92b, Har96c, Hay98a, Hem95, Hen92b, Hen98a, Hil91c, Hoc92b, Hoc92d, HS90, Hor95, Hug92c, Jon94, Jor91, Jor94, Kus95c, Kus96a, Kus97, Law96a, Lee90, Lor90, Lyn99, Mar94b, Mau95, Met93, Mic98, Moo92, Mor92a, Mor98, Mül93, Mur92a, Mur96, Nic93a, Nig99a, Nig99b, Nut93, O’B92, Ore90, Ost98, Pal91a, Poo99, Pum95a, Ras96a, Rou97, Rud92a, Sho96a, Sho99, Slo97, Smi95b, Smi98, Smi99a, Smi99b, Spa91, Ste91a, Stu91, Stu98, Twe93c, Ver94, Wei90, Whi90d]. Paperbacks [Bra91]. Papers [ATKL08, Bar92, Buc92, Cro96c, Fyf99, Hac93c, Hen94b, Mar93c, Mar96a, Nic93b, War96a, AL87, BH87, BKH90, Bry91, Can99, Dav99, EJ^{+}07, ER93, EJR95, GG91a, MH02, NS84, NJ91, SCS^{+}87, War95a]. Pappus [Jon86]. Paracelsians [Pri93, Deb91]. Paracelsus [Daw98a, Gre93b, Gol91a, PTS^{+}96]. Paradigm [Por97]. Paradigms [Hem95, Le 91, Pet94, Ste90b]. Paradise [Bro95d, Hil91c, Whi90b, Fin93d]. Parakashan [Har92d]. Paranormal [Mau95, Ing84]. parascience [Ing84]. Parchments [Rud92a, Lev90]. Paris [Ano82a, BG^{+}92, Bow93d, Bow93b, Bow95, Bry91, Chu91, CM90, Cro94, Cro96d, Dav94, Gol95b, GG90, GG96, Hal90d, Hal95, Hil98a, HS91, Kel95, Neu92, Non94, Non93b, Osb90b, Out94, Pyl91, Pyl92, vH92, BG^{+}92, Cla92b, CM90]. Parisian [Hun92, Str87]. Park [Har93b, Mar96a, Por91a, Smi90a, EJR94, EJR95]. Parkes [Aga94b, Rob92]. Parot [Bry91]. Paroxysms [Mal94, Sha93c]. Part [Bro93a, Fau96b, Hil91c, Kus94, MR87a, MR87b, WW93, Wil91b, Wil97, Ric92a, War95b, NPYGDL86, WW86]. partial [Eng84]. Particle [Kra91b, Pic84, Ste97a]. partout [Sho98]. Parts [Cro96d]. Pascal [Ost99, Edw87, Loe87, Spa94]. passe [Sho98]. passe-partout [Sho98]. Passions [Mal94, Sha93c]. Past [Smi91c, Bow89, MJ89, May97b, Wil87a]. Pasteur [Pal96a, Gei95]. Pasteurization [Ver90, Lat88]. Patai [Gre96]. Patent [Ink90, Mac88b]. patenting [Luc96a].

Paterson [Hor92, PCA91]. paths [GG92]. Paton [Hal94b]. Patrice [Cro96c]. Patricia [Sch91a, Blo93c]. Patrick [Wil90a, Bow93b, Dem83]. Patron [Mac90b, Alt87]. Patronage [Aga97c, Daw98b, Hun92, Doe96]. Paul [Aga98, Bro91a, Har98d, Het95, Jon91, Lar96a, Mar93c, Mar96a, Nye93b, Sch91b, Bäu84, Bow95, Di 93e, HS99, Jor94, KW87, Mar93b, Por97, Twe90, Woo98]. Paulette [Osb90b]. Pauli [Hen84b, HP85, PHvMW93]. Pauline [Jac96, Jon92]. Pauling [Mor92d, Par99]. Payne [Smi98, GH84, GH96]. Payne-Gaposchkin [Smi98, GH84, GH96]. Payot [Pyl91]. pays [Ano90k, vH92]. Peace [Aga93, GG92, Hod92b, Fis98a]. Peacock [Bar93b, Cro91a]. Pearson [Mag96b, SP98]. Pearsonian [Mag96b]. Pedagogy [Dol98a]. Pedersen [Bro93d, Hal94c, Jam94c]. Peggy [Twe93b]. Peirce [Des89b, Des89b]. Peiresc [Fin93c]. Pencil [Odd91, Pet89]. Penelope [Kni90b, Luc98]. Penguin [MK98]. Penicillin [Bud98, Swa92, Wai90, Wil84]. Pennsylvania [Bro92c, Cam92, Gre91, Por91a, Por97, Sha92]. Penrose [Edg90b]. pensée [Bec91b, Kel95]. Pentagon [Aga98, NOF96]. penury [Goo88]. people [Mor84a, Roc88]. Peppas [Div90]. Pera [Dea92, Mor93]. Perception [Smi93a, Hat90]. Percy [Hoc92a, Wal90]. père [CMDM82]. Peregrinus [Hac92d, Hac93a, Hac94, Hor92, Jam92a, Jam93b, Mar93b]. Pereira [Pri91]. Perez [Ost93d, Ree90]. Pérez-Ramos [Ree90]. Perfectly [Jam92a, BS91]. Performance [Gou96, Mar96b, FG93]. perihelion [Ros82]. Period [Hil91a, Ano92p, Sta90c]. periodical [LP04]. periods [Ano82c, Ano82b, NW82b, NW82a]. Periphery [Hes95b, Lin93]. permutations [Bur87]. perpetual [Sch95a]. Perplexed [Kel91]. Personal [Har96a, Maz96, Sha93b, DR95, Kai94, MGBE95, Men86].

Perspectivas [Oli92, OE^{+}90]. Perspective [Ben90a, Har93b, Bal91a, Cla99b, Kah99, War86, Whi88]. Perspectives [Kla95, Le 91, Luc96b, Por91b, Str90, AB89, KT87, Mal05, OE^{+}90, Ste90b]. Persuading [Dea92, PS91, SW93]. Persuasion [Uim95, Lun94]. Persuasive [PLP92]. Peter [Dol98b, Dol99, Hac92d, Hac93a, Hac94, Hen94b, Hor92, Jam92a, Jam93b, Mar93b, Aga94b, Boh95, Bra92b, Bro92b, Gre91, Hal91b, Hoc93b, Hod90b, Ili98, Mac90b, Mal93b, Mor92a, Mur92b, Nic94b, Sha92, Smi91c, Wei90]. Peterfreund [Hoc91]. Petroleum [FB82]. Petroski [Hem95, Odd91]. Petzold [Eck92]. Petzold-Reichenbach [Eck92]. peu [vPvP90]. Phaenomena [Mol98b, EBT96]. Phaidon [Hac90d]. Phaidon-Christie [Hac90d]. Pharmaceutical [Pic92b, Lie87, Swa88, Wea90]. Pharmacy [Gre92c, Hei88, Mor91b]. Pharmazie [Hei88]. phenomena [Hac91]. phénomènes [Bla83]. phenomenon [Bla83]. PHIBB [Lan90]. Philadelphia [Bai91, Bro92c, Cam92, Di 95a, Gre91, Hil91c, Hun92, HD97, Kra98, Nye93b, Por90a, Por97, Sha92, Sho93a, Tei93]. Philanthropy [Kri91, Aas90]. Philip [Dol98b, Dol99, Edg91a, Met96b, Ger95, Goo88, Kla95, Kus95g, Por91c]. Philological [Sho99]. Philosopher [Rag95, Yeo92, Fis91a, Fre92, Gow83]. Philosopher-Scientist [Rag95, Fre92]. Philosophers [Ley90, Mal95a, FJ93, KRS83, Pum95b, Wyl89]. Philosophical [Bai91, Di 95a, Fyf99, Hil91c, Hun92, Kra98, Mor98, Nut93, Por90a, SB92, SW95, Tei93, Gro89, Hei93, KT87, Lin92, MT83, Mor82, Ner87, Reh83, Ver85, vHC95, Dav99]. Philosophie [Eck92, HRP90, Ano92n, Cas93, Pos90]. Philosophique [Out94, Pyl92]. philosophische [Hei93, Kus95d]. Philosophy [Ait90, Ams92, Bow94b, Bro96a, Bry92a, Cas94, Coo96, Gau99, Gol93, Gre92c, Hac93c, Hen92a, Hen94a, Hen98b, Jor94, Kni92b, Kni95, Kni98a, Kni98b, Mai88, Mal93b, Mar96c, Met96b, Met99, Mic98, Mol98b, Mor90a, Mor90e, Nor94a, Osb90a, Ost93d, Ost98, Pow82, Rus88, Sha93a, She92, Ste95, Whi90c, Yos99, Ano92n, BL83, BA91, Bro95b, Bru87b, Cav96, Cla82, Cla89, Clu88, FC96, Fur89, GA98, Gou96, Gra87, Har82b, Har98e, HRP90, Hom83, JK84, Kal95a, Kus95g, Mar92c, Mil99, Mor91a, Mor95a, Nye93a, O’M89, Old86, PW89, SSK88, Ste92, SW95, Ste99, Wil95b, ZH06]. Photographing [Blu90, FL88]. Photography [Jam93a, Rot93b, Sch92a]. Phrenology [Byn90, Coo89b, Coo84]. phylogenèse [Bow95, CM93, MC^{+}93]. phylogeny [MC^{+}93]. Physical [Duh91, Hal91b, Mar96c, Mor92d, Roh99a, Row90b, AH85, Pai86, Ste87, vH97]. Physician [Kat90, BaJANJ84, Jon88, Rud88, Sta84b]. Physicians [Law91a, Fri81, Daw98b]. physicien [BaJANJ84]. physiciens [Pes84]. Physicist [Hug92c, Smi94a, Can93b, Can94, KBRS90, KW87]. Physicists [Cra92c, Roc90, Gin88, Wil87c]. Physics [Bec91a, Bos92, Bro99c, Cas94, Cra92c, Gro93, Hen92a, Hes95b, Hoc92b, Hug96, HD97, Kel83, Kni92b, Kra91b, Kri91, Mal94, Ole91, Por91c, Smi90a, Yav92, Aas90, Bru88c, CHU87, Gar92, Goo90b, Gou97, Har82a, Har82b, HW81, Hen84a, KT87, Kel91, Lin93, Liu92, Mir89, Moy83, Pic84, Pop82, Pow82, Roc90, Sha93c, WR92, Wil87d, Pes84, War90b]. physicus [SSH85]. Physiognomy [Tyt82]. physiologic [Pyl91]. physiologie [Duc82, Grm90]. Physiology [Law92e, Pic90, Por97, Ful91, Grm90, Kre90, NC81, O’C88]. Physique [Woo95, Pes84]. Physischen [Ver91]. Phytochrome [Nic93c]. Piaget [Cha88]. Picard [Kel95, Pic87]. Pickering [Ans96].

Pictorial [Law93f, Rud92b]. picture [CW93, PB^{+}89]. piece [HS99]. Pierce [Cre91b]. Pierre [O’B96a, Bru87b, Cro94, Gab90, Jak84, Mar96c, Smo94]. Pietsch [Spa98b]. Pigment [Nic93c]. Piltdown [Mur94]. Pinch [Gol94]. Piñero [Goo91b, Mal98]. Pinsky [She92]. Pinto [Har98b]. Pinto-Correia [Har98b]. Pioneer [Edg90b, Fri83, Ald85, Cla92b, Ost84, Pen89]. Pions [Kra91b]. Pitfalls [Olb91a, DPP88]. Pittenger [Jon94]. Place [Goo90a, Hom83, You85]. places [Olm92]. Planck [Hei86, MR82f, MR82g]. Planetary [Don94, TW95, Wil91b, Wil97]. Planets [Gra94, Hen96, Lev90, Rud92a]. Plank [Sho93a]. Planning [Rei94]. Plant [BW90, Pal91b, Sch91b, Cit90, Nic96b, THBH92b]. Plants [Alb92, Ste91c]. plastic [Fri83]. plates [Cam90, Fyf99]. Platonismo [Fin92]. Pleasure [Har98d, CL96]. Plenty [Hor97a, BR95]. ploughshares [Mus98]. Plurality [Dic82, Cro86]. Pneuma [Fre93, Mol98a, Fre95, Osl91]. Poder [War90a, Hue88]. Poe [Tre97]. poem [Ros94b]. Poems [Hüb90]. poet [Mol90]. Poggi [Ste95]. Poincaré [Gra85]. Point [Fra90, Bar89a]. Poirier [Cro94]. Polar [Lar96b, Lev94, Ber95]. Polemic [Ano04]. polemics [Sep88]. policies [Mar10]. Policy [Bug96, CK91, Pal90, Coo89a, Gas99, Hen89, Lil85, Wri94b]. Polish [Ams92, Fie92]. Polite [Law98a, Sut95]. Political [Arm83, Bow94a, Sch91a, Sha97, Stu98, Kuz87, Web98, Wei92]. politicians [H^{+}82]. Politics [Aga98, Bar90b, Bow91b, Bro91a, Bug96, Hac90e, Hoc92d, Hod92a, Jac90, Kus98, Law92c, Mun98, Pal90, Por91a, She92, Tur93c, Bru88a, Bru95, Cat88, Coo89a, Des89a, Edw96, FG96, GR88, Man88, MF86, PW89, Ric91, SY86, Smi89, Wil91a, Wri94b, vG92]. Polity [Hen92a]. Polly [Smi99b]. Polytechnic [Wri96a]. Polytechnical [Lan87]. polytechnique [Lan87, FD87]. Pont [GG96]. Pontificia [Ben93]. Poole [Far99]. Poonam [Har93b]. poor [May97a]. Popes [Ben93, Maf91]. popolazioni [Di 93a]. Popular [Mor98, Pic93a, Coo84, Ram88, Top92, vHC95]. popularization [Dau98]. Population [Hoc93a, Mor92a, Sch91a, CHU87, Cra92b, MJ89, Cro91b]. Porfirevich [FS88]. Porphyry [Gou98a, But96]. Porter [Bar90a, Bru96, Far94, Har92b, Hay98a, Hen95b, Hod91, Kus95b, Mar94b, Wis96]. Portrait [Cra92a, McN92, Ada89, Cro84, FS91, HK83, Ken83, Kno90, O’D83]. positivism [Lan91]. positivismo [Lan91, Pan91a]. possession [SB92]. Post [Ost94, Gay92, Out84]. post-Darwin [Gay92]. Post-Italianate [Ost94]. post-revolutionary [Out84]. Posterior [Non93a, Wal92]. posters [PLP92]. Postface [Ost94]. Postscript [Ost94, PB82, Mer93, PB83]. Postwar [Hir95, Hoc92c, Nak91, YK94]. potent [Tre97]. Potions [Rud92a, Lev90]. Poundstone [Twe93c]. pour [Atr86, BG^{+}92, Out94]. Powell [Cor88b]. Power [Blu90, Car95, Goo88, Hil94, Hor96a, Hor97b, Jam94b, Nig99b, Nye92a, Nye92b, Uim95, CF86, Daw98b, FL88, Hue88, Hun85, Kni92c, Lin84, Luc90, Lun94, Mac83, Mor95a, Sha84, The94]. powerful [Tra92]. Powers [Ait90, Cla89]. Pozzo [Fin93c]. Pp [Aga93, Aga94a, Aga94b, Aga95a, Aga95b, Aga96a, Aga97b, Aga97c, Aga98, Ait90, Alb92, Alt91, Aly90, Ams92, Ans96, Ari90, Bai91, Bai93, Bal91b, Bar92, Bar90a, Bar93b, Ben90a, Ben93, Ben94b, BV91, Bev90, Bis98, Blo93a, Blo98, Blo93b, Blo93c, Blo94b, Blu90, Boh93, Boh95, Bos92, Bow92a, Bow94a, Bow91a, Bow91b, Bow92c, Bow92b, Bow93d, Bow93a, Bow93b, Bow93c, Bow94b, Bow95, Bra90, Bra91, Bra92a, Bra92b, Bra95, Bra98, BD90, Bro90a, Bro93b, Bro93a, Bro94a, Bro95a, Bro98, Bro96a, Bro91a, Bro92b, Bro93c, Bro93d, Bro94b, Bro90c, Bro92c, Bro96b, Bro99c, Bro90e, Bro91b, Bro91c, Bro95d, Bru91, Bru93, Bru96, Bry91, Bry92a, Buc90, Buc92, Bug96, Bur93, Bur98, But92a, But94, Byn90, CK91, Cam90, Cam92, Can90a, Can90b, Can91a, Can96, Car95, Cas93]. Pp [Cas94, Cha90c, Chu91, Cla92c, Cle92, Col93, Coo90, Coo96, Cor93, Cra92a, Cra92c, Cre91b, Cri90, Cro94, Cro95, Cro96b, Cro96c, Cro96d, Cuo95, Dan92, Dav94, Dea92, Dea93, Di 92, Di 93a, Di 93c, Di 93f, Di 93e, Di 95a, Di 98, Div90, Div98, Dol94, Dol98b, Dol99, Dol90a, Dol90b, Eck92, Edg90a, Edg90b, Edg91a, Edg95, Elm91, Elm97, Far94, Far99, Fau96b, Fie91, Fie92, Fie93, Fie94, Fin92, Fin93a, Fin93b, Fin93c, Fis98b, For90a, For91, Fox92, Fra90, Fre93, Fuc99, Ful98, Fyf99, Gau99, Ger91, Ger95, Get96, Gol91b, Gol93, Gol94, Gol95b, Goo90a, Goo96, Goo97, Goo91b, Gou91, Gou92, Gou95a, Gou98a, Gou98b, GG90, GG91b, GG91c, GG96, Gra91, Gre91, Gre95, Gre92c, Gre93a, Gre93b, Gre96, Gro93, Hac90d, Hac90e, Hac92a]. Pp [Hac92b, Hac92c, Hac92d, Hac93a, Hac93b, Hac93c, Hac94, Hac95, Hak95b, Hal90d, Hal90a, Hal91a, Hal92c, Hal92a, Hal92d, Hal94a, Hal94c, Hal95, Har92a, Har98a, Har98b, Har98c, Har99b, Har99a, Har90a, Har93a, Har95a, Har96a, Har98d, Har90b, Har92b, Har92c, Har92d, Har93b, Har93c, Har95b, Har96c, Har96d, Har97, Har98g, Har94b, Har98h, Hay98a, Hay98b, Hef96, Hel96a, Hel96b, Hem95, Hen90a, Hen90b, Hen92a, Hen92b, Hen93a, Hen93b, Hen94a, Hen94b, Hen95a, Hen95b, Hen96, Hen98a, Hen98b, Hen99, Hes95a, Hes95b, Het95, Hig94, Hil91b, Hil91a, Hil91c, Hil93, Hil98a, Hin91, Hir95, Hoc90a, Hoc91, Hoc92b, Hoc92a, Hoc92c, Hoc92d, Hoc93a, Hoc93b, Hod90a, Hod90b, Hod91, Hod92a, Hod93a, Hod93b, Hod96, HS91, Hor92, Hor95, Hor96a, Hor96b, Hor97a, Hor97b, Hor98, Hor99a, Hor99b, Hud92]. Pp [Hug92c, Hug92a, Hug92b, Hug96, Hug98, Hun92, HD97, Hun98, Hut92, Ili95b, Ili97, Ili98, Ink90, Jac96, Jac90, Jac93, Jam90, Jam91, Jam92a, Jam92b, Jam92c, Jam93a, Jam93b, Jam94a, Jam94b, Jam94c, Jam94d, Joh93, Joh96, Joh94, Jon91, Jon92, Jon93, Jon94, Jor91, Jor94, Jor96, Kat90, Kel95, Kla95, Kni90a, Kni90b, Kni90c, Kni92a, Kni92b, Kni95, Kni98a, Kni98b, Koe91, Kra90a, Kra90b, Kra91a, Kra91b, Kra92, Kra93a, Kra93b, Kra94, Kra98, Kre96, Kri91, Kus94, Kus95b, Kus95c, Kus95d, Kus95e, Kus96a, Kus96b, Kus97, Kus98, La 94, La 97, Lan90, Lar90, Lar96b, Lar96a, Law90a, Law90b, Law90c, Law91b, Law91a, Law91c, Law92a, Law92b, Law92c, Law92d, Law92e, Law93g, Law93b, Law93c, Law93d, Law93e, Law93f, Law96b, Law96a, Law96c, Law98b, Law98a, Law98c, Le 91]. Pp [Lee90, Lev94, Ley90, Lor90, Luc92, Luc93, Luc96b, Luc98, Lyn99, Mag96a, Mal93b, Mal93a, Mal94, Mal95a, Mal96, Mal98, Man93b, Man97, Mar92a, Mar93a, Mar93b, Mar93c, Mar94a, Mar96b, Mar96a, MK98, Mar94b, Mar96c, Mau95, Maz96, Maz98a, Maz98b, McC98, McC95, McE93, McE95, McK96, McK97b, McN92, Met93, Met96a, Met99, Mic98, Mil95, Mil94, Mol92, Mol94a, Mol94b, Mol95, Mol96, Mol98a, Mol98b, Mon93, Moo92, Mor90a, Mor92a, Mor90d, Mor91c, Mor92c, Mor92d, Mor92f, Mor93, Mor95c, Mor96a, Mor98, Mor99, Mül93, Mul91, Mur92a, Mur92b, Mur95b, Mur96, Nea98, Neu92, New92, Nic93a, Nic93c, Nic93b, Nic94a, Nic94b, Nic96a, Nig99a, Nig99b, Nig99c, Non91b, Non91a, Non93a, Non94, Non93b, Nor94a, Nut93, Nye92a, Nye92b, Nye93b, O’B92, O’B96a, O’B96b, Odd91]. Pp [Olb91a, Olb91b, Old97, Oli92, Ore90, Osb90a, Osb90b, Osl97, Osl98, Ost93b, Ost93c, Ost93d, Ost93a, Ost94, Ost98, Ost99, Out94, Pal90, Pal91a, Pal91b, Pal96a, Pal97, Pal98, Pan91a, Pau90, Per93, Pic92b, Pic93a, Pic98, Poo99, Por90b, Por90a, Por91a, Por91b, Por91c, Por95a, Por97, Pos90, Pri91, Pri93, Pum90, Pum91b, Pum91a, Pum92, Pum94, Pum95a, Pyl91, Pyl92, Pyl93a, Pyl93b, Rag95, Rap92, Ras96a, Ras96b, Rav99, Ree90, Rob93, Roh99b, Roh99a, Rou97, Rud92a, Rus91, Rus92, Rus94, Sch91a, Sch93c, Sch95b, Sch91b, Sec96, Sec98, Sha93a, Sha96, She99, She92, Sho93a, Sho93b, Sho93c, Sho96a, Sho99, Shr92, Sil90, Sim92, Sim94, Slo97, Smi90a, Smi91b, Smi91c, Smi91a, Smi92c, Smi92b, Smi92a, Smi93b, Smi93a, Smi93c, Smi93d, Smi94d, Smi94a, Smi95b]. Pp [Smi96, Smi98, Smi99a, Smi99b, Soy96, Spa91, Spa94, Spa95, Spa98a, Spa98b, Sta90b, Ste91a, Ste95, Str90, Stu91, Stu98, Sut97, Swa92, Tei93, Tha91, Tim98, Tur93a, Tur93b, Tur91, Tur93c, Twe90, Twe91a, Twe91b, Twe91c, Twe92a, Twe92b, Twe92c, Twe93a, Twe93b, Twe93c, Uim95, Ung92, Ver90, Ver94, Ver95, Ver91, War91a, War90a, War91b, War95b, War96a, War96b, Wat91, Wea91, Wei90, Wei94, Whi97, Whi90a, Whi90c, Whi90d, Wil90a, Wil91c, Wil91b, Wil94a, Wil95a, Wil97, Wis96, Wit95a, Woo95, Woo98, Wri94a, Wri96b, Yav92, Yeo96a, Yeo96b, Yos99, vH92, Sha97, Smi92d]. Practical [Joh94, Mac95, Mil99, Wil93]. Practice [Hoc91, Ili95b, O’B92, Pic93a, Yav92, Ano82d, Bia93, Gou96, Ner87, Ole91, Pet90, Ram88, Sho86, Sir90, Vos97, War86, B^{+}82]. Practices [Bro93a, Mur85]. practitioner [Hil98b]. Prakashan [Har92c]. Prato [Lar96b]. Pre [Hil91b, CK90, Wil87a]. Pre-Industrial [Hil91b, CK90]. pre-Newtonian [Wil87a]. preceding [Can99]. Precision [Goo90b, McC95, Ras96b, Pic87, Wil94b, Pic87]. Preclassical [Bai93, Gre92a]. precursor [BM82]. Predecessor [Por91b, AB89]. Preface [Gol95b, Ost94, Daw98a, Di 92, FC82]. Preformation [Har98b, Cor97]. Prehistoric [Cam90, Law93f, HW87, Rud92b].

Prehistory [McK96, Mur96, NW82b, Ano82c, Har85, Van93]. préliminaire [LSR95]. Preliminary [Bra92a, LSR95, Yeo96a, EO90, New89]. Prelude [Wal81]. première [Grm90, Pyl91]. premiers [Poi04]. Preparation [Sut99]. Present [Jam94a, Sch91b, App93, BW90, MJ89]. presented [ABR93, Bry91, Gou95a]. presenting [Goo94]. Preserved [But94, HAB92]. president [Wal88]. Presidential [Bro99a, Can97, Kni96]. presocratic [KRS83]. Press [Aga93, Aga94a, Aga94b, Aga95a, Aga95b, Aga96a, Aga97b, Aga97c, Aga98, Ait90, Aly90, Ari90, Bai93, Bar92, Bar90a, Bar93b, Ben90a, Ben94b, BV91, Bev90, Bis98, Blo93a, Blo98, Blo93b, Blo93c, Blo94b, Blu90, Boh93, Bos92, Bow92a, Bow94a, Bow91a, Bow92c, Bow92b, Bow93a, Bow93c, Bra90, Bra92a, Bra92b, Bra95, Bra98, Bro90a, Bro93b, Bro93a, Bro91a, Bro93d, Bro90c, Bro92c, Bro99c, Bro90e, Bro91b, Bru96, Bry92a, Buc90, Buc92, Bug96, But92a, Byn90, CK91, Cam90, Cam92, Can90b, Can91a, Can96, Car95, Cha90c, Cla92c, Coo90, Coo96, Cri90, Dan92, Daw98a, Dea93, Di 93b, Di 93c, Di 93d, Dol94, Dol98b, Dol99, Dol90a, Edg90b, Edg91a, Edg95, Elm97, Far99, Fie93, Fie94, Fis98b, For90a, For91, Fox92, Fra90, Fre93, Ful98, Gau99, Ger91, Gol93, Gol94]. Press [Goo90a, Goo97, Gou92, Gou98b, GG91b, GG91c, Gra91, Gre91, Gre96, Gro93, Hac90e, Hac95, Hak95b, Hal91a, Hal91b, Hal94a, Har92a, Har98a, Har98b, Har98c, Har99b, Har90a, Har93a, Har98d, Har90b, Har92b, Har93c, Har95b, Har96d, Har97, Har98g, Har98h, Hay98a, Hay98b, Hef96, Hem95, Hen90a, Hen90b, Hen91b, Hen92a, Hen92b, Hen93b, Hen94a, Hen95a, Hen96, Hen98a, Hen99, Het95, Hig94, Hil91b, Hil93, Hir95, Hoc90a, Hoc91, Hoc92a, Hoc93a, Hoc93b, Hod90a, Hod90b, Hod91, Hod92a, Hod96, HS90, Hor95, Hor96a, Hor96b, Hor97a, Hor97b, Hor99a, Hud92, Hug92a, Hug92b, Hug98, Hun92, Ili95b, Ili98, Ink90, Jac96, Jac90, Jac93, Jam93a, Joh93, Joh96, Joh94, Jon93, Jon94, Jor91, Jor94, Jor96, Kat90, Kla95, Kni90a, Kni92a, Kni95, Kni98b, Kra90a, Kra90b, Kra91a, Kra91b]. Press [Kra92, Kra93b, Kra94, Kre96, Kri91, Kus94, Kus95c, Kus95e, Kus96a, Kus96b, Kus97, Lar90, Law90b, Law91b, Law92a, Law93g, Law93b, Law93c, Law93e, Law93f, Law96b, Law96a, Law96c, Law98b, Law98a, Law98c, Le 91, Lev94, Ley90, Luc92, Luc93, Mag96a, Mal93b, Mal93a, Mal94, Mal95a, Man93b, Man97, Mar93a, Mar93c, Mar96b, Mar96a, Mar94b, Mau95, Maz96, McC98, McN92, Met96b, Met99, Mic98, Mil95, Mol94a, Mol94b, Mol96, Mol98a, Mon93, Mor92a, Mor90d, Mor91c, Mor92d, Mor93, Mor95c, Mor98, Mur92a, Mur96, Nea98, Nic93a, Nic93c, Nic93b, Nic94b, Nic96a, Nig99a, Nig99b, Nig99c, Non91b, Nut93, Nye92a, Nye92b, Nye93b, O’B92, O’B96a, Olb91a, Olb91b, Ore90, Osb90a, Osl98, Ost93c, Ost93d, Ost94, Ost98, Ost99, Pal90, Pal91a, Pal91b, Pal96a, Pal97, Pal98]. Press [Pau90, Pic92b, Pic93a, Pic98, Poo99, Por90b, Por91a, Por91b, Por91c, Por97, Pri93, Pum90, Pum91b, Pum91a, Pum92, Pum94, Pum95a, Pyl93a, Ras96a, Ras96b, Rav99, Ree90, Rob93, Roh99b, Roh99a, Rou97, Rud92a, Rus94, Sch91a, Sec96, Sec98, Sha97, Sha92, Sha93a, Sha96, She92, Sho93b, Sho93c, Sho96a, Sho99, Sim92, Sim94, Slo97, Smi90a, Smi91a, Smi92c, Smi92b, Smi92d, Smi93b, Smi93a, Smi93c, Smi95b, Smi96, Smi98, Smi99a, Smi99b, Spa91, Spa98b, Sta90b, Ste91a, Str90, Stu91, Stu98, Tim98, Tur93c, Twe90, Twe91a, Twe91b, Twe91c, Twe92a, Twe92b, Twe92c, Twe93b, Twe93c, Uim95, Ver90, Ver94, Ver95, War91b, War95b, War96a, War96b, Wat91, Wea91, Wei94, Whi90c, Whi90d, Wil90a, Wil91c, Wil91b, Wil94a, Wil97, Wis96, Woo98, Wri94a, Yav92, Yeo92]. Press [Yeo96a, Yeo96b, Yos99, LP04]. Press/London [Rus94]. Press/Los [Twe92b]. Press/Royal [Har93c]. Presses [Bow93b, Cro96d, Har96a, Kni92a, Rus94]. presuppositions [Ver85]. preussischen [Tur93b, ZH92]. Prévost [FM00, FM00]. Price [Lar96a, Aga98, Bro94a, Bro96a, Bro92c, Bro91c, Bur93, Cas93, Gou95a, Hac93b, Har92d, Hen94b, Kat90, Kra98, Lar96b, Mal96, Mal98, Met99, Mol94b, Non94, Por90b, Rou97, Sha92, Sho93b, Sho93c, Spa94, Twe90, War90a, Whi97, PN95]. prices [Goo91b]. pride [LLDM94]. Priestley [AL87]. prima [Fin93b, Olm92]. prime [Mar92b]. Prince [Gou98a, But96]. Princeton [Bra95, Bra98, Coo96, Fie94, Fox92, Gou98b, GG91c, Gre96, Har90b, Hen99, Kre96, Kus95e, Law96c, Mag96a, Maz96, Mor92d, Mor93, Nea98, O’B96a, Pal96a, Ras96b, Roh99b, Smi96, Uim95, War91b, War96a, Wis96, Woo98]. Principe [Pyl92, BH90]. Principia [Bra92a, Bra98, Hal94a, Ber93b, BN82, De 95, New89]. Principle [Jam94d, Sch91a, BH90, Kip91, MJ89]. Principles [Bow92c, LR91, LS97, MK98, Sch91a]. print [Dol98a]. Printed [Hen93a, Got92]. printing [Eis83, Eis93, NhTWL87]. Priority [Hal94a, Ber93b]. Prisoner [Pou93, Twe93c]. Private [Pal96a, Gei95]. prizes [Cra84, CHU87]. pro [JK84]. probabilità [Cre91b, Des89b]. Probability [Hod91, Mar94a, Sha83, Das88, Des89b, G^{+}89, KI92]. Probing [Hes99]. Problem [Cla92c, Rou97, Smi93c, CS92, Dol98a, Gro91, Har82b, WT82]. problematic [Fis94]. Probleme [Ver91]. Problems [Bar92, Osb90a, RR90, Web88, Whi90a, Fur89, Llo91]. Proceedings [McK96, Ver95]. process [Hul88, Ner87, THBH92b]. Processing [Aga98, NOF96]. Proctor [Pau90]. prodigy [O’D83]. Production [Hor97a, BR95]. profession [Eng83, Gol87, Gre86, Nic88, Pat83b, Pic90]. Professional [Pic93a, Ram88]. Professionalization [Wei94, Geu92]. professor [Ash85, Pic99]. professors [Pro93]. Program [Gre91, Wri94a, McC93, Mor89]. programmatic [Fis94]. Progress [Jac90, Olb91a, Smi91c, Bow89, Cat88, DPP88, Har94a, The94]. progressive [Mus98]. Project [Fau96b, SG80, Sep88, May99]. Projects [Har98a, Cun97]. projet [SG80]. Promethean [HS99]. Promoting [Hal91c, Ost93c]. Proof [Non93a, Wal92]. Prophecy [O’B96a, Smo94]. Prophet [Ash85]. proposed [CMDM82]. proposées [CMDM82]. prospective [Her91]. prospects [MJ89, Mar10]. Prosper [Mor90d, Sto88]. Protestant [Par87]. Protestantism [Ili98, Har98f, Jac83]. proton [Bro85]. protyle [Bro85]. Prout [Bro85]. provenance [JK84]. provided [Hen94b]. Prussian [Por91a, Bru88a, ZH92]. Pseudosciences [Hen93a, Got92]. psyche [Fur88]. Psychiatric [Hay98a, Uim95, BP95, Gol87, Lun94, PB99]. Psychiatry [Hay98a, Ver85, BP95, PB99, Sho86]. psychical [Opp85]. Psychoanalysis [Blo93b, Fig99, Kus98, TR92, Bru95, MF86]. Psychoanalytic [Smi93c, CS92]. Psychological [Hak95b, Kus94, Kus95c, Cra93, Dan90, Ric92a]. Psychology [AW87, Fox87, Hay98b, Hen90a, Kus95c, Kus96b, Kus97, Por90b, Smi99b, Tur93c, Wei94, Ash95, Geu92, Hea87, Lea90, Lea94, Sha98, Shu96, Shu04, Sim88, vG92]. Pt [KN88]. Ptolemaic [GB95a]. ptolemaico [GB95a, Mal96]. Ptolemy [Fie92, Hen91a, PT84]. Public [Bow91b, Dol94, Gol93, Kni92b, Law93c, Pic98, Wis96, Dau98, Gol92b, GR88, Hul99, IW96, Lan87, Lan89, OE^{+}90, Por95c, Pum95b, Ste92, SW95, Sut94, Yeo93]. Publica [Goo91b, Goo91b, Lan89]. publication [Jam94c]. Publications [Bro98, Bro93d, Cro96b, Dea92, Di 93f, Fie91, Hal90a, Hal92d, Har93b, Hen94b, Hes95b, Kni90c, Law90a, Law91a, Luc96b, McE95, Pyl93b, Sha97, Spa95, Wit95a, Wri96b]. Publications/University [Bro93d]. público [Oli92, OE^{+}90]. publics [Lan87]. Published [Har98d, Hen94b, Jam94a, Man93b]. Publishers [Bow94b, Cas94, Div90, Div98, Dol90b, Hel96b, Hod93a, Jam92c, Mar96c, Mil94, Nor94a, Nye93b, Sch95b, Smi94d, Ste95, Twe92b, Twe93b]. Publishing [Bro99c, Hal92c, Har95a, Hen93a, Hod93b, HD97, Ili97, Lar96a, Law92e, Lyn99, Mol98b, Mur95b, BM84]. Publishing/Bhativedanta [Mur95b]. Pufendorf [HD97]. Pugh [Twe92a]. Pullen [Sch91a]. pump [SSH85]. Punishment [Bru91, DH90]. Puritanism [Hen92b, CDS90]. purity [Cla92b]. Purposes [Edg91a, Smy89]. Pursuit [Cri90, Jam93b, Wis96, Mac83, Por95c, Ric88, FTG91]. Puzzle [Twe93c, Pou93]. Pyenson [For90a, Har95b]. Pygmalion [Cro94]. Pygmalion/Gérard [Cro94]. Pythagoras [Whi90c, O’M89].

Quackery [Bar90a, Por89]. quaestio [Mol95, Law93a]. Quaestiones [Hin91, Gro89]. Qualitative [Smi94d, Bli92]. qualities [Ale85]. Quality [LaF82]. Quantification [Mag96a, Mat95]. Quantifying [BV91, Coh84, FHR90]. Quantum [Pop82, Hen84b, MR82a, MR82c, MR82d, MR82e, MR82f, MR82g, MR87b, MR00]. Quarks [Kra91b, Pic84]. quarter [Buc85]. que [Goo91b, ML89]. Queen [Bra90, For90a, Moo92, Rud92a]. Queisser [Hod90a]. Quellen [Hin91, Gro89]. Quest [Kus96b, Law91b, Mag96a, Pal98, Smi93d, Ash95, Fuj96, Mat95, Sta89b, vdVV91]. Questions [Luc90, Nye92a, MT83]. quinine [GL88]. Quod [Coo90, SLT88].

R [Bar92, Bec99, Bro94a, But94, Cla92c, Dea92, Di 93f, Fau96b, Fie93, Fuc99, Gol92a, GG93a, GG93b, Hac92d, Hal90a, Hal92a, Hoc90a, Hug96, Kra93b, Lan90, Mag96b, Man97, Mol98b, Non91a, Pyl93b, Sch91b, Sim94, Smi93b, Twe92a, Twe92b, Twe93b, War95a, Wea91, Whi90a, Wis85, Woo98]. R. [Bro90a, Bro90c, Cuo95, Dol90a, Dol90b, Fau96b, Ful98, Gou95a, Hal92a, Kre96, War96b, Wil91b, Wil94a]. Rabdology [Nap90, Twe92b]. race [CT93, Mur95b, WKB88]. Racial [Pau90, Pro88, Lil85, WKB88]. Rad [Pic92a]. Radar [Hac93a, Cal90]. Radcliffe [Jan98]. Radiation [Can90a, Cau89, OP87]. Radical [Hod92a, Des89a, Jac83, Sch93b]. Radicals [Bec95]. Radio [Aga94b, Aga94c, Poc88, Rob92, Sul82b]. radioactive [De 92]. Radiological [De 92]. Rafael [War90a]. Raff [Spa91]. Raffaella [Bal91b]. ragione [Ern91, Per93]. railroads [Gil83]. Rainbow [Rus94, Tra93]. Rainger [Bow93c]. raison [Bea91, Ung92]. Raisons [Hal95, Bla93]. Raj [Har97, Kum95]. Ramos [Ree90]. Ramsay [Pic93a]. Ramsey [Sho99]. Raphael [Gre96]. Rarest [Spa98b, PR95]. Rashed [Bur98]. Rasse [Wei90, WKB88]. Rassenhygiene [Wei90, WKB88]. Rassenkunde [Ver91]. Rassenpolitik [Lil85]. ratio [HF87]. Rational [PP87, Por97]. Rationalist [Sut97, CD94]. Rationality [Olb91b, Wyn82, Can88]. ray [Hir83, Har93c]. Raymond [Pri91, McK96, Mor90d, Per89]. Rayner [Hor99b]. Rayner-Canham [Hor99b]. Rays [Bro99c, Dah97]. Rd [Fau96b]. Re [Goo94, Roh99b]. Re-enchanted [Roh99b]. Re-presenting [Goo94]. reaction [Gra84]. read [Kla95]. Reading [Di 93d, Win91]. readings [Sch01]. Real [Nic93b, Sin91, Wes95]. Realism [Dol90b, PB83, Rou97, Nol88]. realities [Rou97]. Reality [Smi94d, Bli92]. Reappraisal [AZW87, Bow94b, Sar92]. Reappraisals [Aga95a, LW90, Pum94]. Reardon [But92a]. Reardon-Anderson [But92a]. Reason [Jor91, Nig99b, Por97, Smi93c, CS92, Ern91, Gut89, Pow96]. Reasonable [Hun98, Hey95]. Reasoning [PP87]. reasons [Bla93]. reassessment [Coh96, Mor94]. recapitulation [MC^{+}93, Bow95, CM93, MC^{+}93]. Received [Ano90g, Ano90h, Ano90i, Ano93j, Ano93k, Ano94j, Ano95i, Ano95k, Ano96j, Ano96l, Ost98, Ano91i, Ano91j, Ano91k, Ano91l, Ano92k, Ano92j, Ano92l, Ano93i, Ano93l, Ano94i, Ano94k, Ano94l, Ano95j, Ano96i, Ano96k, Ano97a, Ano97b, Ano98b, Ano98c, Ano98d, Ano98e, Ano99a, Ano99b, Ano99c, Ano99d, BDH96]. Reception [Gli87, Jun89, Fyf97, MR82c, MR82e, MR87b, dG01, dG01]. Rechenmaschinen [Twe93b, MKW92]. Rechenmeister [Met96a, Sch93b]. Recherche [Bry91, Ano82a]. Rechnen [Alt91, Ger89]. Recollections [Smi98, GH84, GH96]. recombinant [Kri82]. reconsideration [Kra97]. Reconsidered [Mur92b, U^{+}91]. Reconstructing [BC98, Jor94, She99]. Reconstruction [Boh93, Hir95, Pic98, IW96, Ric92b, YK94]. record [DD94]. Recreating [Vos97]. recruitment [Don91]. red [Tra92]. Redaction [Mol94b, BEoCF92, ECL92]. Redirecting [Aas90, Kri91]. rediscovered [GJ85]. Reduction [Cla92c, Gro91]. Reede [Hen86b]. reeling [KN88]. réelle [Neu92, Sin91]. Reenchanted [Har96b]. Reflection [Mal94, Sha93c]. reflections [Grm90]. Reflexions [CF86, Pyl91, Grm90]. Reform [Far99, Hal92c, Hod92a, Sha93a, Des89a, Fri81, Gle91, Mar92c, Poo98, Yat83]. reformation [Par87]. reformer [Pag82]. reformers [Who82]. refrigerator [HS99].

Reginald [Cha93]. Regine [Tur93b]. Regiomontanus [Alt91, Ger89, Ger89]. Regiomontanus-Bianchini [Alt91, Ger89]. Regions [Law91b, Sta89b]. Régis [Bur98]. register [Pan97]. Regulatory [Bug96, Wri94b]. Reichenbach [Eck92, HRP90]. Reihe [HD97, LS76]. Reijer [Ost99]. Reill [Dol98b, Dol99]. Reingold [Poo99]. Reinhold [Fie92, Hen91a]. reinterpretation [Don88, Eme84]. Reinterpreting [Hod90b, Bow88]. related [vLMB88]. Relation [Rou97]. Relational [Gra91, Ear89]. Relations [Lar96b, BM82, Ber95, McG84, MM10]. relationship [Sch89]. relationships [Sha83]. Relative [Fra90, Bar89a, Wil87a]. Relativism [Dol90b, Nol88]. relativistic [Liu92]. Relativitätstheorie [Wol87]. Relativity [Kra93a, Pye85, Gli87, Ray87]. Relevance [Hal90a]. Religion [Ern91, Gol91a, Gre93b, Hut92, Jam92b, Mor90a, Ost93d, She99, Bow98, BC98, Can91b, CMDM82, Clu88, Cor88b, Hal90c, Lan91, Lee96, Sha83]. Religione [Ern91, Per93, Lan91, Pan91a]. Religious [Nut93, Dur85, Lin92, MW95]. Reluctant [Mac90b, Alt87]. remember [Tod83]. Reminiscences [Aga93, KW87, Hod92b]. removal [MJ89]. Renacimiento [Lan86]. Renaissance [Bru91, Fie94, Gol91a, Gre93b, Har98a, Hen99, Mal95a, O’B92, Rav99, Wil91b, Wil97, Cun97, DH90, Ern91, FJ93, Fie97, Gol91a, Lan86, Mol95, Sch83, SSK88, Sir87, Sir90, Sir97, TW95, Vic84, Yat83]. Renard [Spa98b, PR95]. René [Pyl93b, Wil97, She91, Smi93d]. Renn [War95a, War95b, War96a]. reorganized [McC85]. replications [BD95, BD95, Mor96a]. Report [Ano91s, Ano92p, Ano93s, Ano96q, Ano99e, Ben94a, Cri94, Cri96, Hug96, May99, Smy89, WR92, Ano98g]. Reppe [THBM92]. Representation [Rou97, BB84, Kel98, Sal85]. Representations [Dol98b, Dol99, Law93f, Rou97, G^{+}09, Hay94, Jac95a, MR96, Rud92b, G^{+}09]. Reprint [Twe93b]. Reprints [Smi90a, Bru88c]. Republic [Jac90, Sha96, War91b, Bro89, Cat88, HD97, Gol95a, Lan87]. républicains [Roc88]. Republicans [Roc88]. République [Lan87]. Research [Aga95a, AB89, CM90, Gre91, Hoc92b, Hoc93b, Kni92b, Kus95c, Law91b, Lev94, Nic93c, Por91b, Tei93, Bec82, BC90, Cre91a, CC94, Dan90, Dei92, ES90, Ern91, Fru90, GH92, Jan98, Law90a, Mor89, Opp85, Rei85, Sta89b, Swa88, THBH92a, Wis85, Hul99]. Resistance [Hor96a, Hor97b]. resisted [Jor90]. resistible [Pal96b]. resolution [EC87]. resonance [Par99, Zan86]. Resource [Mil95, Gay96, Law90a]. Resources [RR90, Pum91c]. respectability [Ver93]. respecting [MJ89]. Restaging [BD95, Mor96a, BD95]. restitutae [Viè83]. Restoration [Joh93, Joh94, Wil93, Kro91]. Restored [Hil91c, Whi90b]. Resurrection [Har98a, Cun97]. Retarded [Bis98, Nol95]. Rethinking [Kus96a, Meg94]. retirees [GG92]. retreat [Smi88]. Retrospect [BBBM90, Smi93b]. Rev [Sho93c]. Rev. [CHG91, Gun84]. revaluation [Don91]. Revealed [Bro95d, Fin93d]. reverence [SB92]. Review [Aga93, Aga94a, Aga94b, Aga95a, Aga95b, Aga97b, Aga97c, Ait90, Alb92, Alt91, Aly90, Ams92, Ans96, Ari90, Bai91, Bai93, Bal91b, Bar92, Bar90a, Bar93b, Ben90a, Ben93, BV91, Bev90, Bis98, Blo93a, Blo98, Blo93b, Blo93c, Blo94b, Blu90, Boh93, Boh95, Bos92, Bow92a, Bow94a, Bow91a, Bow91b, Bow92b, Bow93a, Bow93b, Bow93c, Bow94b, Bow95, Bra90, Bra91, Bra92a, Bra92b, Bra95, Bra98, BD90, Bro90a, Bro93b, Bro93a, Bro94a, Bro95a, Bro98, Bro90b, Bro96a, Bro91a, Bro92b, Bro93c, Bro93d, Bro94b, Bro90c, Bro90d, Bro92c, Bro96b, Bro90e, Bro91b, Bro91c, Bro95d, Bro95e, Bru91, Bru93, Bru96, Bry91, Bry92a, Buc90, Buc92, Bug96, Bur93, Bur98, But92a, But94, Byn90, CK91, Cam90, Cam92, Can90a, Can90b, Can91a, Can96, Car95, Cas93, Cas94, Cha90c, Che90]. Review [Chu91, Cla92c, Cle92, Col93, Col95, Coo90, Coo96, Cor93, Cra92a, Cra92c, Cre91b, Cri90, Cro94, Cro95, Cro96b, Cro96c, Cro96d, Cuo95, Dan92, Dav94, Daw98a, Dea92, Dea93, Di 92, Di 93a, Di 93b, Di 93c, Di 93d, Di 93e, Di 95a, Di 95b, Di 98, Div90, Div98, Dol94, Dol95a, Dol95b, Dol98b, Dol99, Dol90a, Dol90b, Eck92, Edg90a, Edg90b, Edg95, Elm91, Elm97, Far94, Far99, Fau96b, Fie91, Fie92, Fie93, Fin92, Fin93a, Fin93b, Fin93c, Fis98b, For90a, For90b, For91, Fox92, Fra90, Fre90a, Fre93, Fuc99, Ful98, Fyf99, Gau99, Ger91, Ger95, Get96, Gol91b, Gol93, Gol94, Gol95b, Goo90a, Goo96, Goo97, Gou91, Gou92, Gou95a, Gou98a, Gou98b, GG90, GG91b, GG91c, GG96, Gra91, Gre91, Gre95, Gre93a, Gre93b, Gro93, Hac90d, Hac90e]. Review [Hac90b, Hac90c, Hac92a, Hac92b, Hac92c, Hac92d, Hac93a, Hac93b, Hac93c, Hac94, Hac95, Hak95b, Hal90d, Hal90a, Hal90b, Hal91a, Hal91b, Hal92c, Hal92a, Hal92d, Hal94a, Hal94c, Hal95, Har92a, Har98a, Har98b, Har98c, Har99b, Har99a, Har90a, Har93a, Har95a, Har96a, Har98d, Har90b, Har92b, Har92c, Har92d, Har93b, Har93c, Har95b, Har96c, Har96d, Har97, Har98g, Har90c, Har94b, Har98h, Hay98a, Hay98b, Hef96, Hel96a, Hel96b, Hem95, Hen90a, Hen91b, Hen92a, Hen92b, Hen93a, Hen93b, Hen94a, Hen95a, Hen95b, Hen95c, Hen96, Hen98a, Hen98b, Hen99, Hes95a, Hes95b, Het95, Hig94, Hil91b, Hil91a, Hil91c, Hil93, Hil98a, Hin91, Hir95, Hoc90a, Hoc90b, Hoc91, Hoc92b, Hoc92a, Hoc92c, Hoc92d, Hoc93a, Hoc93b, Hod90a, Hod90b, Hod91, Hod92a, Hod93a, Hod93b, Hod96, HS90, HS91, Hor92, Hor95]. Review [Hor96a, Hor96b, Hor97a, Hor97b, Hor98, Hor99a, Hor99b, Hud92, Hug92c, Hug92a, Hug92b, Hug96, Hug98, Hun92, Hun95a, HD97, Hun98, Hut92, Ili95b, Ili97, Ili98, Ink90, Jac95b, Jac96, Jac90, Jac93, Jam90, Jam91, Jam92a, Jam92b, Jam92c, Jam93a, Jam93b, Jam94a, Jam94b, Jam94c, Jam94d, Joh93, Joh96, Joh94, Jon91, Jon92, Jon93, Jon94, Jor94, Jor96, Kat90, Kea90, Kel95, Kla95, Kni90a, Kni90b, Kni90c, Kni90d, Kni90e, Kni92a, Kni92b, Kni95, Kni98a, Kni98b, Koe91, Kra90a, Kra90b, Kra91a, Kra91b, Kra92, Kra93a, Kra93b, Kra94, Kra98, Kre96, Kri90, Kri91, Kus94, Kus95b, Kus95c, Kus95d, Kus95e, Kus95f, Kus96a, Kus96b, Kus97, Kus98, La 94, La 97, Lan90, Lar90, Lar96b, Law90a, Law90b, Law90c, Law90d, Law91b, Law91a, Law92a, Law92b, Law92c, Law92d, Law92e, Law93b]. Review [Law93c, Law93d, Law93e, Law93f, Law96b, Law96a, Law96c, Law98b, Law98a, Law98c, Le 91, Lee90, Lev94, Ley90, Lon90, Lor90, Luc92, Luc93, Luc96b, Luc98, Lyn99, Mac90b, Mag96a, Mal93b, Mal93a, Mal95a, Mal96, Mal98, Man90, Man93b, Man95, Man97, Mar92a, Mar93a, Mar93b, Mar93c, Mar94a, Mar96b, Mar96a, MK98, Mar90, Mar94b, Mar96c, Mau90, Mau95, Maz96, Maz98a, Maz98b, McC98, McC95, McE93, McE95, McK96, McK97b, McN92, McN95, Met93, Met96a, Met96b, Met99, Mic98, Mil95, Mil94, Mol92, Mol94a, Mol94b, Mol95, Mol96, Mol98a, Mol98b, Mon93, Mor92a, Mor90d, Mor91c, Mor92c, Mor92d, Mor92f, Mor93, Mor95c, Mor96a, Mor98, Mor99, Mül93, Mul91, Mur92a, Mur92b, Mur95b, Mur95c, Mur96, Nas90, Nea98, Neu92, New92, Nic93a, Nic93c, Nic93b, Nic94a]. Review [Nic94b, Nic95, Nic96a, Nig99a, Nig99b, Nig99c, Non91b, Non91a, Non93a, Non94, Non93b, Nor94a, Nut90, Nut93, Nye92a, Nye92b, O’B92, O’B96a, O’B96b, Odd91, Olb91a, Olb91b, Old97, Oli92, Osb90a, Osb90b, Osl97, Osl98, Ost93b, Ost93c, Ost93d, Ost93a, Ost94, Ost98, Ost99, Out94, Pal90, Pal91a, Pal91b, Pal96a, Pal97, Pal98, Pan91a, Pau90, Per93, Pic92b, Pic93a, Pic95, Pic98, Poo99, Por90a, Por91a, Por91b, Por91c, Por95a, Por97, Pos90, Pri91, Pri93, Pum90, Pum91b, Pum91a, Pum92, Pum94, Pum95a, Pyl91, Pyl92, Pyl93a, Pyl93b, Rag95, Rap92, Ras96a, Ras96b, Rav99, Ree90, Rob93, Rog90, Roh99b, Roh99a, Rud92a, Rus91, Rus92, Rus94, SG80, Sch90b, Sch93c, Sch95b, Sch91b, Sec96, Sec98, Sha92, Sha93a, Sha96, She99, She92, Sho93a, Sho93b]. Review [Sho95, Sho96a, Shr92, Sil90, Sim92, Sim94, Slo97, Smi90c, Smi90d, Smi91b, Smi91c, Smi91a, Smi92c, Smi92b, Smi92d, Smi92a, Smi93b, Smi93a, Smi93c, Smi93d, Smi94d, Smi94a, Smi95b, Smi95a, Smi96, Smi98, Smi99a, Soy96, Spa91, Spa94, Spa95, Spa98a, Spa98b, Sta90b, Sta95, Ste91a, Ste95, Str90, Stu98, Sut97, Swa92, Tei93, Tha91, Tim98, Tur93a, Tur93b, Tur91, Tur93c, Twe91a, Twe91b, Twe91c, Twe92a, Twe92b, Twe92c, Twe93a, Twe93b, Twe93c, Uim95, Ung92, Van91, Ver90, Ver94, Ver95, Ver91, War91a, War90a, War91b, War95a, War95b, War96a, War96b, Wat91, Wea91, Wei90, Wei94, Whi97, Whi90a, Whi90c, Whi90d, Wil91c, Wil91b, Wil94a, Wil95a, Wil97, Wis96, Wit95a, Wit95b, Woo95, Woo98, Wri94a, Wri95, Wri96b, Yav92, Yeo92, Yeo96a, Yeo96b, Yod90, Yos99, vH92]. review [Pum91c]. Reviewed [Ano96m, Ano91m, Ano92m, Ano93m, Ano94m, Ano98f]. Reviews [Aga96a, Aga98, Ben94b, Bow92c, Bow93d, Bro99c, Di 93f, Edg91a, Fie94, Goo91b, Gre92c, Gre96, Hen90b, Hen94b, Jor91, Lar96a, Law91c, Law93g, Mal94, Moo92, Mor90a, Nye93b, Ore90, Por90b, Rou97, Sch91a, Sha97, Sho93c, Sho99, Smi90a, Smi99b, Stu91, Twe90, Wil90a, WM85]. Revised [Hen93b, Sha97]. revisions [But94]. Revisited [Hal92d, Woo98, Jos97].

Revival [Jon93, Deg91]. Revived [Whi90c, O’M89]. Revolution [BA91, Bec91a, BV90, Bro93a, Can91a, Cas94, Coh85, Cro95, Cro96c, Fie91, Fie94, Hen98a, Hod90b, Hor97a, Ink90, Le 91, Mal93b, Mal95a, Man97, Mar94b, PT86, Pum94, Smi90a, Ano87b, Ath90, Bec95, Ben86, BR95, Bow88, B^{+}95, Bru88b, Coh84, Coh94, Cro90b, Don88, Don93, Edg91b, Eis83, Eis93, FJ93, Fri81, Gle82, Grm90, GGSGB94, Hal83a, HL86, LeG88, LW90, LP04, Mas98, Max84, New94, PLP92, PT92, Ros84a, Ste90b, BV93, B^{+}95, Cro96d, Gol95b, Grm90, Pyl91, Mac88b, NS02, Ano87b, CLR87, NS02]. revolutionary [Lan87, Out84]. révolutionnaires [Lan87]. Revolutions [Hal90a, Jor94, BV93]. Revue [BG80, SG80, Cro96d]. reweaving [Sta85]. rewriting [ZH06]. Rey [Goo91b]. Rezeption [Di 92, Jun89]. Rhetoric [Dea92, Gol93, Pum95a, Lar91, Mos93, PS91, SY86, Slo85, Ste92, Wil87a]. rhetorical [NS02]. rhétorique [NS02]. Rice [Ver95]. Ricerche [Per93, Ern91]. Richard [Cro96c, Di 93f, Ili97, Law93d, Law98c, Mol94a, Mur95b, Ost93d, Wei94, Wri87, Bro90e, Bru91, Bru93, Car95, Dol94, GT92, GGSGB94, Har96d, Joh93, Kus95e, Law92e, Non91b, Ost93d, Pum92, Sho86, Van91]. Richards [Boh93, Cri90, Kus94, Law92c]. Richardson [Twe92b, Ash85]. Riddle [Whi90d]. Rider [BV91, Hug96, Twe92b]. Riemann [Gra85]. Riksmuseet [Lar96b, Ber95]. Riley [Lar96a, RT95]. Rimmer [Lar96a]. Rinascimento [Bal91b, Per93, Ern91]. Riper [Mur96]. Rise [Can90b, Gol93, GG91c, Hen92b, Hen95a, Ili98, Kri91, Law91c, Maz98b, Mol95, Ost93b, Wil91b, Wil97, Aas90, Buc89, CDS90, Cor96, Eng83, GR91, Har98f, Huf93, Huf95, Law93a, MR82f, MR82g, MR87a, Pal96b, Por86, Rob89, Sch90a, Ste92, TW95, THBvdB92, Wes88]. Ritchie [Blo93b]. Ritter [Hil98a]. ritual [Wyn82].

Ritvo [Smi92d]. Rivalry [Hac90e, Man88]. River [BB82]. Rivka [Sha97]. road [Gle82]. roads [FS84]. Robert [Ans96, Ben94b, Bro91a, Cas93, Gou98b, Hen94b, Het95, HS91, Kni92a, Law91c, Mol94b, Mor91c, Mul91, Ost99, Pum91a, Pum94, Sch91a, Smi96, War95b, War96a, Boh93, Bos92, BKH90, BHW94, Bro91a, Buc90, BEoCF92, Can93a, Cla99b, Div98, ECL92, Far99, Fig99, Hey83, Hoo97, Hor92, Hun90, Hun99b, Kah99, Kra90b, Mar93c, Mar96b, Mar96a, Ost94, Ost98, Pau90, SG90, Sha93b, Shr92, Smi91a, Tay89, Ver94, Wil90c, Yeo96b]. Roberts [Bro93a, Elm91, Law91c, Law93g]. Robertson [Aga94b, Blo93b, Hor97a]. Robin [BV91, Hug96, Sim92]. Robinson [Maz96, Mor91c, DR95, Wil90c]. Roche [Cra92c]. Rochester [Sim94]. rocks [Her83]. Roderick [Smi91a, Sta89a]. Rodopi [Hac92b, Spa98a]. Roe [Edg95]. Roger [HC91, Mal93b, Out94, BL83, Blo98, Byn90, Get96, Gow83, Law90b, Met99, Soy96]. Rogers [Bry91, Hen94a, Str90]. Roland [Rus92, Hen90b, Kel95]. Role [Man93b, Por91b, AB89, Dob91, GB95b, Hil98b, Sin97, Sut94, Top92]. Roles [Hor99b, RCRC98]. Roman [Gre96, Osb90a, Smi94c, Fur89, Ole84]. Romance [Kni90a, Mer89]. romano [Wal84]. Romantic [Bro93b, Pum92, MF92, Vig82]. Romanticism [PB94, Ste91a, Ste95, CJ90]. Rome [Ben93, Sha97]. Ronald [Lar96a, Aga97c, Bow93c, Bro94b, Lar96a]. Ronceverte [Gou92]. Rond [Yeo96a]. ronde [Ano82a]. room [Cha84]. Roots [Hak95b, Cra93, THBB92b]. Rosalind [For91]. Rosen [Daw98a]. Rosenberg [Mar93c, Mar96a]. Roshdi [Bur98]. Roskilde [Hac95]. Roslynn [Rou97]. Ross [Jam92c, Smi95b]. Rosser [Mag96a]. Rossi [Kra91a]. Rothmann [GB95b]. Rotterdam [Ili97]. Rousseau [Sch01]. Routledge [Bur98, Edg90a, Far94, Get96, Hal92a, Hen95b, Hor98, Jon92, Law90c, McC95, Mor90a, O’B96b, Smi91b, Smi99b]. Rowan [O’D83]. Rowlands [Bra92b, Hal91b]. Rowman [Ger95, Hil91a]. Roy [Bow94a, Bru96, Far94, Har92b, Hay98a, Hen95b, Bar90a, Mar94b]. Royal [Cra92a, Hac92c, Hal92c, Hal94c, Har93c, Hen94b, Hun92, HD97, Kni90c, Law91c, Law93d, Lyn99, Mor92c, Ost93c, Pal97, Pum91b, Sch91a, Wil90a, FFLW97, AL87, Bar90b, BKH90, DTS^{+}87, Frä89, Gas99, Gle91, Gun84, Hac84, Hal84, Hal91c, Hun82, Hun89, HH97, Mil99, Ped92b, Rus99, STT87, Ste99, Str87, Tay87, TST87, WH91]. royale [Str87]. Rs [Har92c]. Rubber [Gre91, Mor89]. Rubbettino [Fin92]. Rubinger [Bro90c]. Rubinin [Hug92c]. Rudbeck [Wit95a, Eri94]. Ruderman [Kat90]. Rüdiger [BD90]. Rudolf [Di 92, Har95a, Spa91]. Rudwick [Bow92c, Law93f]. Rue [Rot93b]. Ruey [Di 93c]. Ruled [GG91c, Sch90a]. Ruler [Osl97]. runaway [Har94a]. Rupees [Har97]. Rupert [Hut92, Ost93a, Mal94]. Rupke [Bow91b]. rural [Pri83]. Russel [Bro84, Smi92c]. Russell [Fox92, Ost93b]. Russett [Bev90]. Russia [Luc93, Coo92]. Russian [Hug92c, Smi99a, Fri81, Gra98, KBRS90]. Russo [Wri96b]. Rutgers [Dea93, Hen92b, Nic96a]. Ruth [Lar90]. Rutherford [Bad85]. Ryan [Gou95a, Str90, Jam93b].

S [Aga94a, Bal91b, Bow92c, Coo90, Ger91, Gou98a, Hac92b, Hal92a, Har90a, Hen93a, Ili97, Jor94, Kra92, Kra93b, Law93f, Law96c, MW95, Mol98b, Mor96a, Non91b, Nye99, Out94, Por90a, Pum94, Rus94, Smi93b, Tei93]. S. [Fie94, Kra97, Mor90e, Wal91]. S29.90 [Can91a]. sa [Bec91b, Kel95]. Saber [War90a, Hue88]. Sabino [Ben93]. Sabra [Smi92a]. Sachiko [Met96b]. Sachs [De 96]. saga [Wri87]. Sage [Har93b, Nic93c]. Saggi [Mel98]. Sahotra [Bow94b]. Sa’Id [Hil93]. Sale [Bar90a, Por89]. Salem [Hil93]. Sally [Hay98b]. Salud [Goo91b, Lan89]. Sämtliche [HD97, LM23, LS76]. Samuel [HD97]. San [Aga97b, Mur95b, Nic93c, Twe92b, Twe93b]. Sanches [Coo90]. Sandemanian [Jam92b, Can91b]. Sandwich [Jam90, Twy88]. Sangwan [Har92d]. Sanidad [Goo91b]. Sanitarias [Goo91b, Fre90b]. sanitary [Fre90b]. Sansoni [La 94]. sapere [Fin93b, Olm92]. Sapp [Blo93a]. Sarah [Hen93b, Lyn99, Mul91]. Sarasohn [Fis98b]. Sarkar [Bow94b]. Satellite [Hug92b, Mac90a]. Satisfyed [Jam92a, BS91]. Satpal [Har92d]. Saul [Law91a]. Saussure [Woo95]. Savage [Hil91a, Kuk91, Kuk93, Moy86]. Savantes [Dav94, vH92, Ano93q]. Savants [GG96, Lan87, PP95]. Savilian [Pro93]. saviours [Söd86]. savoirs [Bea91, G^{+}09, Ung92]. scala [La 94]. Scale [Hoc93b, GH92]. Scaling [Mil95, Smi94e]. Scarecrow [Byn90, Law93e]. Scenes [Law93f, Rud92b]. Schaaf [Jam93a]. Schabas [GG91c]. Schaffer [McN92, Pum91a, SSH85]. Scheutz [Twe91c, Lin90]. schism [Pop82]. Schleicher [Tau93]. Schloss [Gre93a, Wey92]. Schmidt [HD97]. Schmidt-Biggemann [HD97]. Schmitz [Di 92]. Schneider [Met96a]. Schofield [Kni92a, SG90]. Scholar [Hoc92b, ES90]. scholarly [Vic91]. Scholars [Har98h, Mal95a, Sho99, AS96, FJ93, ZH92]. Scholastic [Mol95, Gra84, Law93a]. schollers [Hil98b]. Scholz [Koe91]. School [Ams92, Lan87, Kim95, Lan87]. Schools [Aga95a]. Schriften [HD97, LM23, LS76]. Schrödinger [Kra90a, MR87a, Moo89]. Schubert [Hoc92b]. Schulmann [War95b, War96a, War95a]. Schuster [Rou97]. Schütt [Bro94a]. Schwab [Yeo96a].

Science [Aga93, Aga97c, Alb92, AL87, Ano82d, Ano87c, Ano97d, Ano99e, Bai91, Bar92, Ben91, Ben94b, BC89, Bev90, Blo93c, Bow91b, Bow94b, Bro90a, Bro93b, Bro98, Bro91a, Bro92c, Bur93, Bur98, But92b, But94, Cas94, Cha90a, Cha90b, CB^{+}90, CM90, Cor88b, Cro90a, Cro96a, Cro92, Cro95, Cro96b, Cro96c, CW93, Dau98, Dav99, Dea92, Di 93b, Di 93c, Dol94, Dol90a, Dol90b, Edg90a, EO90, Elm97, Fie94, For90a, Frä89, Ful98, Fyf99, Gol91b, Gol92b, Gol93, Gol94, Gou91, Gou92, GR88, Gra87, GG91b, Gre96, Hac90e, Hac92a, Hac93c, Hac94, Hal90a, Hal92a, Hal92d, Han85, Har92c, Har92d, Har97, Har98g, Hef96, Hen90a, Hen92b, Hen93a, Hen95a, Hen98b, Hes95b, Het88, Hig94, Hil93, Hir95, Hoc91, Hoc92a, Hoc92c, Hoc92d, Hoc93a, Hoc93b, Hod90a, Hod91, Hod93a, HAB92]. Science [Hom83, Hor92, Hud92, Hug92c, Hug92a, Hul88, Hun95b, Ili95b, Ili97, Ili98, Ing84, Jam92a, Jam92b, Jam92c, Jam93b, Jam94a, Jam94b, Jam94c, Joh96, Joh94, Jor94, Kat90, Kni90c, Kni92a, Kni92b, Kra98, Kri91, KN88, Kum91, Kum95, Lar96b, Law91a, Law92b, Law92d, Law93c, Law98b, Law98a, Law98c, Lee90, LS90, Lev93, Lev94, Llo83, Luc96b, Lyn99, Mac90b, Mal93b, Mar96c, Mau95, McC98, McC85, McE95, Met99, Mil95, Mol95, Mol96, Mol98b, Mor90a, Mor92d, Mor95c, Mor98, Mur96, Nak91, NLH^{+}84, NhTWL87, Nic94a, Nic94b, Nig99a, Nig99b, Non91a, Non94, Nor94a, Nut93, Ols90, OE^{+}90, Osl98, Ost93d, Ost99, Pal96a, Pal98, Pic98, PP87, Pum90, Pum91b, Pum92, Pum95a, Pyl93b, Ree90, RR81, Rob93, Roh99b, Rou97, Rus83a, Rus83b, San91, Sch91b, Sec98, She99, Sho93b]. Science [Sho93c, Sho93d, Sho93e, Sho96a, Shr92, Smi90a, Smi99a, Spa95, Spa98a, Ste95, Stu91, Stu95, Sut95, Top92, Tra92, Tur91, Ver95, Wal90, War95b, Wis96, Wit95a, Woo98, Yeo92, YK94, vHC95, Aas90, AH85, Aga96b, AH93, Ali86, Alt87, Ano82c, Ano82b, Ano93q, App93, Ara91, Atr86, A^{+}87, Baa90, BA91, BM82, Bee96, Ben82, Ben97, Bia93, Bow98, BW90, BM84, BC98, Bro99a, Bru87a, Bru88b, Buc98, Bud97, Bur83, Bur87, Can91b, Cla99a, Cla92a, Cla95a, Cla82, Clu88, Coh84, Coh85, CDS90, CP93, Coo84, Cra84, Cra92b, De 96, Doe96, Dol96, Don93, Dor91, Dra81, Edg91b, EC87, Eri94, Fin91, Fis91a, Fog92, For98, FC96, Fuj96, GH92, Gas99, Gei95, Gie88, G^{+}89, Gil83, Gje84, Gle82, Gol91a, Goo91a, Goo88, Got92, Gra98]. science [Gre92b, GGC90, Hal83a, Ham90, HA87, Har96b, Har98f, Hei86, Hey83, Hod92b, Hol86, Hoo97, Huf91, Huf93, Huf95, HJF89, Hul99, Hun89, HH97, II90, IW96, JK84, Jen91, Jor93, Jos97, Joy87, KC95, Kel91, Kit92, Kni86, Kni88, Kni89, Kni92c, Kni92d, Kni96, Kra87, Kum96, LaF82, LSS90, Lan91, Lau87a, Law93a, Lee96, Lie87, Lin92, Liv87, Llo91, Lan87, LN95, Luc96a, MW94, Man88, Mar92b, Max84, May97b, May99, McG84, Mei91, MF92, MH88, Mor84b, Mor96b, Mos93, NR86, Ner87, NW82b, NW82a, Nol88, Ogi86, O^{+}90, Old86, Osb94, Out84, Out91, Pag82, Pai82, Pan91b, Pat83a, PS91, PR88, Pet90, Pic93b, PB94, PB83, Por95c, Pow96, Pum91c, RM96, Red87, Rei85, Rei94, Ron83, Ros91, Rot93a, Rud88, Rus89, SG90]. science [Sep88, Sha83, SP88, SW89, Sho96b, SY96, Sht96, Sim88, Sin97, Sme97, Smi89, Smi94c, Smi94b, Smi94e, Sta84a, Sta94, Ste92, Sut94, The94, THBM92, Van93, Vic87, Wal84, WM85, Wil91a, Wil93, Wil91d, Wri96a, Yeo93, Zam92, ZH06, dGB^{+}89, Ano97c, Ano98g, B^{+}82, Bro97d, Bud97, Can97, McC81, Rei94, Top97, Ste91c]. Sciences [AZW87, Blo98, Bry91, Bur93, Dav94, Get96, Hac95, Hal94c, Har95b, Hun92, Kni90c, LS90, Mal93a, McC98, ME81, Mor95c, Mur92a, Nic94a, Nic94b, Non91a, Olb91b, Ped92b, Sho93a, Smi92d, Smi95b, Soy96, Ste91a, Tei93, Vuc84, Ano93q, A^{+}87, Bow92d, Bow95, Can88, Car88, CM93, CM90, Cre91a, CC94, Cro96d, CJ90, Fru82, Fru90, GG90, Hil98a, HJF89, Lan87, Llo83, MMVZ90, MC^{+}93, Por83, Pye93, Rit90, Ros94a, Sho93e, Smi92e, Smi97, WHS92, Baa90, Cro92, Frä89, Str87, Stu95]. Scientiarum [Ben93]. Scientific [Ano93q, Bec91a, Bro94b, Bry91, Bry92a, Buc92, But94, Cas94, CW87, Cro96c, Dea92, EMW82, EC87, Fie94, Fyf99, Goo97, Gou95a, Har98h, Hen90a, Hen91b, Hen98a, Jam91, Jor91, Jor94, Kal95b, Lar96b, Lar96a, Law91b, Law92d, Law98c, Mai88, Man97, Mar94b, Nut93, Ost93b, PP87, Pri93, Pum94, Pyl93b, Rud92a, Sha92, Sho93c, Sim94, Sim88, Smi90a, Smi91a, Smi93c, Tei93, Woo95, ABR93, Ano90k, AS96, Bea91, BC90, Ben91, Ber95, Bru88b, Buc94, CF86, CS92, Cla97, Coh84, Coh94, CHG91, Dau98, Dav99, Dea91, Deb91, Edg91b, Eme84, Fig99, Fru90, Gje84, Goo90c, GR91, Gut89, GGSGB94, Hei93, HRP90, HP85, Hoc94, HAB92, Hol85, Hui92, Hul99, Hun82, Kni86, Kni88, Kni89, Kra90c, LeG88, LHB04, LS76, LHH91, LW90, Lin92, MH02, McC85]. scientific [Mor84c, Ner87, New94, PHvMW93, PS91, PB82, PB83, PT92, yQ85, Red87, Ros84a, Rud85, SY86, She91, SY96, Sla82, SP98, Sta89a, Sta89b, Tur83, Vic84, War90c, Zan86, vL76, PvM85]. scientific-philosophical [Hei93]. scientifica [Lev90]. scientifique [Ano82a]. Scientifiques [Dav94, GG90, Ano90k, Ano93q, Bea91, Ung92, vH92]. Scientist [Bro90c, Can91b, Hor92, Jam92b, Kra91a, Kra93b, Rag95, Rou97, Smi91a, Sta89a, Bäu84, Fre92, Gol92a, Hay94, Kah99, MR89, PCA91, Ros90, vLP82]. Scientists [Hac92b, Har98h, Hen90b, Hun96, Hun92, McG84, Ano93q, AS96, Hal84, Kuz87, M^{+}89, Mil90, Moy86, Swa88, PP95]. Scienza [Fin93a, Ara91, Fin92, Lan91, Pan91a]. Scienze [Di 98, Bal91b, GGC90]. Scitur [Coo90, SLT88]. Scolar [Har98a, Met99].

Scotland [FS84, Sha84]. Screen [Rou97, Tur96]. Scriblerians [Por90b, Fox88]. Scrolls [Rud92a, Lev90]. scrutiny [Hom83]. Sculptor [Gou98a, But96]. scurvy [Car86]. se [Goo91b, ML89]. Sea [Lev90, Rud92a, McC82]. Seacraft [Lan86]. Seaman [Cha90c, How89]. Search [Cra92a, Jam90, Jon93, Jor96, McK97b, Pic92b, Sim92, Ada89, Cre93, Deg91, Her90, Kil94, RR91, THBS92, Twy88, Ver93, Wea90]. Seas [Spa98b, PR95]. secolo [Ara91, Fin93a]. Second [HJF89, Sha97, Smi90a, Bru88b, GG90, Lan90, SG80, Smi98, Don94, HJF89]. seconde [SG80]. Secord [MK98, Slo97, Yeo96b]. secret [KA81]. Secretary [Gun84]. Section [Ano93q, Dav94]. Sections [Kel98]. Sedgwick [Spe82]. Seek [Hac84, Met93]. Seeker [Can90a]. seeking [Ver93]. Segal [Hor95]. Segre [Dea93]. sei [La 94]. seiner [Adl87]. Seite [Hei93, Kus95d]. Selected [Bar92, Hac93c, Lar96a, PN95, Smi90a, Sut97, AT87, Bru88c, CD94, DeV82, GG91a]. Selection [Bar93b, Cro91a, Dem83, Gay92, Har94a, KRS83, Bow93d, Gay92]. selective [Rot93a]. self [HK83]. self-portrait [HK83]. selling [Fri83]. Selskab [Hal94c]. Selwyn [Law91c]. Sema’an [Hil93]. Semiconductor [Hac92d, Mor90c]. Seminar [Yav92, Ole91]. Senecas [Hin91, Gro89]. Senior [Law93e]. sense [Cla97]. senses [Cla97]. Sensibility [Law92d, Law93b, BB92, Ben91]. September [Bry91, PCA91, Ver95]. Series [Bry92a, Hac92d, Hac93a, Hac94, Hen98b, Kni92b, Mol96, Mor99, Nic94b, Spa98a, Tha91, Twe92b, Twe93b, Aga95a, Fie92, HD97, LS76]. seris [Lor90]. Serres [Gol95b, Sho98]. Service [Ben93, Stu98, Maf91, Web88, Web98]. Services [Cra92a, Law91c, Law93d]. Servos [Mor92d]. Sessions [Tur93a]. Set [Bar90b, Bur98, Gau99, Gou98a, Lar96a, Spa98b]. Seth [Pro93]. Setting [Bro93b, MF92]. seu [Viè83]. Sevenoaks [Ost93b]. Seventeenth [Elm97, Gau99, Gre92c, Hun92, Hun98, Sha83, Cle93, GA98, Gra84, Hey95, Hun95b, Mor91b, Mur85, Sla82]. Seventeenth-Century [Elm97, Gau99, Hun92, Sha83, Cle93, GA98, Hun95b]. Sewall [Pro86]. Sex [Bru91, DH90, Law93b, BB92]. Sexual [Bar93b, Bev90, Rus89, Cro91a, Har94a]. Sexuality [Bru91, DH90]. SFr [Bro95a, Jam94d, Koe91, Kra93a, Mar94a, Met96a, Mol94b]. shadow [Mur85]. Shadows [Jam93a, Mel98, Sch92a]. Shaler [Liv87]. Shallis [Shr92]. Shamdasani [Smi99b]. Shandean [Ost94, Mer93]. Shanghai [Wri96a]. Shape [Di 93d, Elm97, Hun95b, Win91]. shaped [Pug84]. Shaping [Nye93b, BC92, Hea87, Rud85]. Shapiro [Mal93a, Mal94]. Sharov [Aga94a]. Shea [Dea92, Hal90a, Non91a, Pyl93b]. Sheehan [Aga96a]. Sheets [For90a, Her83]. Sheets-Pyenson [For90a]. Shelwing [Law93e]. Sheridan [Ver90]. Sherry [Rou97]. Shi [Por95a]. Shifting [Ore90, LeG88]. Shigeru [Hoc92c]. Shirley [Law91c]. Shoenberg [Hug92c]. Shop [Nye93b, CW87]. Short [Mar92a, Buf87, WD82]. Shorter [Smi92c, NR86, WS91b, WS91a]. Shortland [Kni92b, Law90a, Law98c, Lee90, Nic94a, Sec96, Sec98, Shr92]. shortly [Ger89]. Should [Gol94, CP93, Pan97, Yox83]. Shoulders [Cuo95, Ost94, Had94, Mer93]. show [Sch95a]. Shteir [Har98g]. Shuttleworth [Hay98b]. Siberian [Woo98, Jos97]. sícek [Mur92a]. side [Hei93]. Sidereal [Ari90, GV89, SB09]. Sidereus [Ari90, GP92, Non93b, GP92, GV89, SB09]. Siècle [Lar90, Bec91b, BaJANJ84, FM00, G^{+}09, GG96, Grm90, Har92b, Har94c, Kel95, PP95, Pic87, Pyl91, Roc88, Sch01, SGS00, TP90, Har89]. siècles [Poi04]. Siegel [Mar93a]. Sigerist [Daw98a, Har99b, FB97]. siglo [Goo91b, Rey90]. Sigma [Eck92]. Sigmaringen [Gre93a]. Sigmund [FMF85]. Significance [Pum92, Ols90, Wel93, JK84]. Silicon [Hod90a]. Silliman [War91b, Bro89]. Silurian [Sec86]. Silverman [Gou98b]. Simcock [Gou91]. Simms [Tur93a, McC92]. Simon [McN92, Pum91a, Rou97, Bur93, SSH85]. Simonton [Hen90a]. Simple [SL87]. simplicity [RR91]. Simpson [Kol92, SL87]. Sinaceur [Neu92]. Since [McK96, Blu90, Bru91, DH90, FL88, Fru82, Hoc92d, Huf91, Hug92a, Lie88, Mac83, Wil91a]. Singapore [Hac92a]. Singer [Can97]. single [Kni92b, War95b]. singular [Kno90]. Sir [Hen94b, Hor92, Joh94, Law90c, Law91c, Smi91a, Smi92b, Hal94b, NJ91, PCA91, Rob89, Rus86a, RR90, Sta89a, STW90, Tro89, Twe91d, Wal88, Wil93]. Siraisi [Hen99, O’B92]. sistemas [GB95a, Mal96]. Sisters [Lyn99, HH97]. sixteen [vLP82]. sixteenth [GJ86, Rud88, Wal81]. sixteenth-century [GJ86, Rud88, Wal81]. Sixteenthcentury [Kat90]. SJ [Ben93, Sha97]. Sketch [Roh99a, vH97]. sketches [O’C88]. Skies [Aga94b, Rob92]. Sleep [Hak95b, Cra93]. Slinn [Wil94a]. Sloan [Di 93f]. Small [Bro97b]. Smith [Blo98, BD90, Bro91a, Edg95, Gre96, Har96c, Hod96, Mil95, Por95a, Sho96a, Soy96, Tur91, Wal88, Ros94b]. Smith-Gordon [Por95a]. Smith-Gordon/Niigata-Shi [Por95a]. Smithsonian [Edg90b]. Smoller [O’B96a]. Smyth [Edg91a]. Snow [Law93d]. So-Called [Hel96b, BEoCF92, ECL92, Mol94b, vGA94]. Sober [Hun98, Hey95]. sobre [GB95a, Goo91b, Lan89, Mal96, Rey90]. Social [BC90, Bow93a, Di 93e, Hac95, Hug92b, Jon93, Kni92b, Luc92, McC98, Nic93a, Pic93a, Por91c, Twe91b, Wat91, Bry88, Cla97, Cro84, Deg91, Hul88, Kri82, Kuk91, Kuk93, Mac90a, McG84, MC^{+}93, Mir89, Mit92, Nye90, Ram88, Rus83a, Rus83b, Sto84, Stu95, The94, Wei91, Wei82, WT82]. sociales [Rey90]. Socialist [Lil85]. Socialists [Jon94, Pit93]. Società [Fin93b]. Societat [Maz98a]. Société [Woo95, Bow93b, Tor92, THBvdB92]. Sociétés [Dav94, vH92, Ano93q]. societies [McC85, Ano93q]. Society [All86, Ano97d, Ano98g, Ano99e, Bai91, BD90, Bro96a, Bro97d, Can97, Cra92a, Di 95a, Elm91, For91, GS88, Gle91, Hac92c, Hal91b, Hal92c, Hen94b, Hil91c, Hoc92c, Hor99b, Hun92, Jon92, Kra98, Lar90, Law91c, Law93b, Law93d, Law98a, Lee90, Lyn99, Mol95, Mol96, Mor92c, Nic93b, Nic94a, Ost93c, Pal97, Por90a, Pum91b, Sch91a, Tei93, Top97, Wal88, WH91, Wil90a, AW87, BB92, Ben86, Har89, Mac83, MS87, McG84, Nak91, Sut95, Tho87, Tor92, WM85, Wil90b, AL87, Bar90b, BKH90, DTS^{+}87, Gas99, Gun84, Hal84, Hal91c, Hun82, Hun89, HH97, Maz92, Mil99, Row90b, Rus99, STT87, SB92, Ste99, Tay87, TST87]. Socio [Har93b, Pal98, She92, Bal91a, PW89]. Socio-Historical [Har93b, Bal91a]. Socio-history [Pal98]. Socio-Politics [She92, PW89]. sociohistory [Fuj96]. sociological [Pic84]. sociologically [Söd86]. sociology [Pic93b]. software [Mor84a]. Sokolik [Mol94a]. Solar [Aga97c, Doe96, Hug92a, Huf91]. Söllner [Har98h]. Solomon [Hal92d]. Somerville [Pat83a]. Sommerfeld [MR82f, MR82g]. sonar [Hac84]. Sonu [Smi99b]. Sophia [Ken83]. Sort [Kra93b, Gol92a]. Soszyński [Fie92]. Soul [Jam94a, Mol98a, App93, Fre95]. Source [Bai91, Cla99a, Cla92a, Cla95a, Jam92d, Pat94]. Sourcebook [Gre96]. Sources [Duh91, How96, Hug96, Jon92, Mar96c, Mol98b, Wal81, BR96, Gro89, Gun84, Maz92, Wal84, WR92]. South [Bis98, FG96, Nol95]. Southern [Aga94b, Rob92]. Southgate [Liv87, Sch95b]. Soveria [Fin92]. Soviet [Gra87]. Sowerby [Dol98a]. Space [Bro91a, Gra91, Hac92a, Wri94a, Wri96b, Bar93a, Ear89, KRS94, McC93, Smi89, Wri96a]. Space-Time [Gra91, Ear89]. Spain [EO90, Goo88, Lan86, Lan89]. Spanish [Fre90b, GB95a, Hue88, Lan86, Lan87, Lan89, LN95, Mal95b, ML89, OE^{+}90]. sparrow [Ken83]. Spary [Slo97]. Spatial [Smi93a, Hat90]. Spätmittelalter [Gol91a, Gre93b]. Speaking [Har98h, AS96]. Special [Bro93a, Mol95, Law93a]. specialists [Rud85]. speciation [Mag96b]. specie [Fin92]. specierum [BL83]. Species [Jac96, Maz95b, A^{+}87, BGH^{+}87, Tau93, Mor94]. specific [Hes99]. Specificity [Jac96, Maz95b]. Specola [Ben93, Non91b]. spectroscopy [Hea86, Hea90]. speculation [Ano82d, B^{+}82, RR91]. speculativa [BM89]. speculative [Lor90]. speculis [BL83, Mol94b, EiT92, ET92, Tak92]. Speculum [Nor94a, Zam92]. Spencer [Gay98, Hoc90a]. Sperm [Har98b, Cor97]. spheres [GB95b]. Spherical [Mol98b, EBT96]. spinal [NC81]. Spinning [KN88]. Spirit [BV91, FHR90]. spiritualism [Opp85]. Spitalfields [SW95]. spokesman [Hei86]. Springford [Bro99c]. spy [Wil87b]. Square [Mol94a]. St [BB82, Hen90b, Jon91, Law90b, Sch91b]. Stability [Bro96a, Kal95a]. stable [Aga94c]. Staden [Aly90]. Stafford [Joh96, Law93g, Smi91a]. stage [Coh84, Kai94]. stagnation [yQ85]. Stanford [Edg91a, Hoc92a, Hoc93b, Law91b, Smi93c, Smi99a]. Stanley [GG91c, Mar96c, Sch90a]. Stanton [Hil91a]. Star [Lar96b, Law91b, Ber95]. Starkey [Man97, New94]. starlight [Hea86, Hea90]. Starry [Lar96b, Ber95]. Stars [Hen96, O’B96a, Pan97, Gra94, Smo94]. start [Lan87]. State [Cuo95, Mac90b, Mol96, Moo92, Nic93a, Por91a, Sha93a, Smi90a, Alt87, Gas99, Mar92c, McG84, Mit92]. statements [Gro89]. States [Nye92b, Smy89, De 92, Hun85, Rot93a, TT86]. Statesman [Ver95]. Statics [Duh91, Mar96c]. statique [BG80, SG80]. Statistical [Har96a, Het95, MGBE95, Pau93, Por86]. Statistics [Ger95, EM94, Mag96b, Sti86, SG80]. Status [McC98, Kel91, Stu95, Ver93]. stay [Dra81]. Stayer [Bra90]. steam [Lin84]. Steel [Hor96b, Mis95]. Steelmaking [Bar84]. Stefano [Ste95]. Steinberg [Har92a]. Steiner [Alt91, Gre92c, Gre93b, Hel96a, Hin91, Lor90, Mol92, New92, Van91]. Stella [Ben94b]. Stemerding [Alb92]. Stengers [Smi93c]. stepchild [Dij90]. Stephen [Har96a, Smi90a, BC90, Law90c, Smi90a]. Stephenson [Bra95]. Sternberg [Spa91]. Steve [Aga95b, Twe92c]. Steven [Kre96, Bis98]. Stevenage [Hac94]. Stewart [Gol93, Le 91, Sil90, TB97]. Stiefkind [Hac93b, Dij90]. Stillman [Non91a, LS90]. Stockholm [Lar96b, Ber95]. Stoic [Fre93, Osl91]. Stokes [Mor90d, Smi92b, STW90, Wil87d]. Stonemason [Sec96, MS95]. Stoner [Can94]. Storia [Cle92, Ara91, Fin93a]. Story [Swa92, Tom84, Gol83, Lee96, Wai90]. storytellers [Sta85]. Stoughton [Ost93b]. Stoughton/The [Ost93b]. strafrecht [Bur99]. Strand [Jam90]. strange [Liu92]. Strangelove [Rou97, Hay94]. Strategic [Twe91a, OO89]. strategies [Nye99, Par99, Sno99]. strategy [Len82]. Straus [Aga96a]. Street [Jam90]. streitbaren [Tur93b, ZH92]. Strickland [Hen92b]. strike [Hac84, Win93]. Stringfellow [Edg90b, Pen89]. Stroud [Lyn99]. Stroup [Hun92]. structural [Sch89]. Structure [Bec91a, Cas94, Fra90, Sha92, Bar89a, BN82, Dea91, Sal85]. Structures [Maz98b, Cha82, Cor96, Pri83].

Stuart [Hen92b, Mun98, Hoc91]. Stubbe [Jac83]. students [AT87, Don91]. Studi [Bal91b, Mor96a, Pan91a, Sha97]. Studia [Fie92]. Studien [Gre93a, Wey92]. Studienplan [Hel96a, Kra91c]. Studies [Blo93b, Boh95, Bow94b, Bro98, Bry91, Bry92a, Cas94, Coo96, Coo88, Cra92c, Di 93f, Fre92, Hac93c, Hac95, Hal90d, Hal91b, Har98d, Har96d, Hoc93a, Kni92b, Kni98a, Kra93a, Kra91c, Kus95c, Kus96b, Kus97, Lar96b, Law90b, Lev94, Mar96c, Mil95, Mol96, Mol98b, Moo92, Mul91, Nor94a, Osl97, Ost99, Por90b, Pum91a, Rag95, Sha92, Sho99, Ste95, Tur93c, Wat91, Wil90a, Wit95a, Ano82d, Ano07, B^{+}82, Bro95b, Bur94, Cra92b, Dea91, EC87, GT92, Høy94, HC90, Kel98, Llo83, Lan89, ME81, Raf88, Roc90, SY86, Viè83, Wey92, vLP82]. Study [Bos92, Bro91a, But92a, Fra90, Har90b, Hoc92b, Jam92b, Jon91, Mol98b, Ost99, Pal90, Pyl93a, Bar89a, BM82, Büh81, Can91b, Cha91, Cle93, Coo89a, ES90, EBT96, Ful91, Gje84, Har87, Hoo97, Le 90, Liu92, Mei91, Mor82, OP87, Old86, OH85, RA91, Sch84, Sha83, Smi89, Sul82a, War90b, Wil87d]. Sturdy [McC98]. Stuttgart [Alt91, Di 92, Di 93e, Gre92c, Gre93b, Hel96a, Hes95a, Hin91, Hug96, Lor90, Mol92, New92, Van91, Ver91, Whi90a, AH93]. Style [Ano04, Tei93, Fru90]. Subject [Kus95c, Dan90, Fle98]. submarine [Hac84]. Subscription [Cro96d, Hen94a, Kni92b]. subscriptions [Lan90, WW88a]. Substance [Fur88, Mol98a, Fre95, Har82b, Sta84a]. Subtle [Pai82]. Sudhoff [Gre92c]. Suez [MM10]. Suffolk [Elm97, Joh94, McC98, Pum91b, Pum91a]. Suggestions [Jam93b, FTG91]. Suhrkamp [Wei90]. suitable [Hun99b]. sul [Bal91b]. sulla [Cre91b, Des89b, Pan91a]. Sumida [Pum90]. Sun [Fie92, Hug92a, Hen91a, Huf91, Rot93b]. SUNY [Mol96]. Sunyer [Maz98a, Mal95b]. Superbomb [Edg91a, Yor89]. Superior [Oli92, War90a]. Supplement [Bow94a, Hod96, BSK^{+}94, Wei92, Sar87]. supply [H^{+}82]. support [Gay96]. sur/Studies [Hal90d]. surgeon [Ald85]. Surgery [Law91c, Rob89]. Survey [Hac90e, Her95, Old97, EO90, Her88, Law90a, Söd86, Man88]. Surveying [Hac90d, Ben87]. surviving [CF86]. Susan [Bug96, For90a, Law91b]. Sutton [Law98a, Lyn99]. Suzanne [Gou98a]. Svante [Hes95b]. Sweden [Kni90c, Lar96b, Ber95, Frä89, Lin84, Söd86]. Swedish [Frä89, Hes95b, Kni90c, Lin93]. Sweet [Bra90]. Swijtink [Hod91]. Swiss [War95a, War96a]. Swords [Mus98]. Sydney [Hac92a, Hod96, Kni92b, Jam92c]. Symmetrian [Pos90]. Symmetrie [Koe91, Sch89]. Symmetrien [Mai88]. symmetry [RR91, Mai88, Sch89]. Symons [Jam92a]. Symposia [Tha91]. Symposium [Bry91, McK96]. syndrome [Wel93]. Synthese [Mil94]. Synthesis [Smi93d, vdVV91]. Synthetic [Gre91, Rus94, Mor89, Tra93]. System [Aga97c, Hug92b, Ink90, Doe96, Mac88b, Mac90a, Pug84]. System/360 [Pug84]. Systematics [Ver93]. Systems [Hac93a, Luc96b, Twe92a, Cal90, GB95a, PJP91, Sal85].

T [Bai93, Bra92a, Cam92, Fie94, Fis98b, Gre91, Gre92c, Jor94, Lan90, Mal93a, Mor90e, She99, Sho93c, Sho99, WHS92, Wil94a]. T. [CHG91, FG96]. table [Ano82a]. Tacitus [Hüb90, Met93, Hüb90]. Tacitus-Übersetzung [Hüb90]. Takahashi [Mol94b]. Takes [Ras96a, Nyh95]. Talbot [Jam93a, Sch92a]. Tale [Smi92d, GG92, Rit90]. Tales [Gin88, Jam92d]. Taming [Dan92, Hac90a]. Tamsyn [O’B96b]. Tanner [GG93b, GG93a]. Tapestry [Jac93, KH92]. tardo [Ern91, Per93]. Tatarewicz [Bro91a].

Taton [Wil91b, Wil97]. Tauber [Law92a]. taxonomist [Ver93]. taxonomy [Sla82]. Taylor [Bro99c, Cas93, Fyf99, Hug98, Law91c, Nea98, BT96, BM84]. Teachers [Smi90a]. Teaching [Lee90, Mol95, SW89, Ano87b, Goo90b, HJF89, Law93a, Par99, Sir87]. Tech [Bro93a]. Technical [CK89, Wil91c, Ano90k, Bec91b, LHB04, LS76, LHH91, vL76, Ano87b]. technician [Sor99a]. Technik [AH93, Eck92, Hes95a, Hug96, Tur93b]. technique [Ano87b, Bec91b, CLR87, Kel95]. Techniques [Bry91, Dav94, Ano90k, Ano93q, Fin93c, vH92]. technischer [HD97, LHB04, LS76, LHH91, vL76]. Technological [Div98, Edg95, Fox95, Ben86, SM94, Wei82, Win93]. Technologie [CM90, Sch93c]. Technology [Aga98, Ben94b, Bro91a, Car95, Cro96c, Div98, Edg90a, Edg95, For91, Gol91b, Gol93, Gre95, Hac92a, Hac92d, Hac93a, Hac94, Hac95, Har92d, Hen93a, Hil91b, Hoc92c, HS90, Hor95, Hor96a, Hor97b, Hor98, Hor99a, Kni92b, Lin84, Luc92, Mar92a, Mar96b, Mol96, Mor99, Nye93b, Pic98, Pyl93a, Smi91b, Smi99a, Sta85, THBB92a, THBS92, Wri94a, AH93, And92, Ano93q, Bak96, Bas88, BC89, BC92, But92b, C^{+}92, Car94, Cha91, CB^{+}90, CM90, CK90, DM96, EC87, FG93, Fox95, Goo88, Got92, Gra98, HSHSJ93, HH86, HH92, Hil94, IW96, KN88, Lie88, Mac83, McC93, Mor96b, Nak91, NLH^{+}84, NPYGDL86, NhTWL87, NOF96, Nye90, Ole84, Pic93b, Rot93a, San91, Seg94, Smi89, SM94, Ste92, Tho87, THBM92, TW87, WD82, Wil90b, YK94, aHH86, aHH88]. technology [aHH92, AH93]. technology-science [THBM92]. Tecknik [Eck92]. Tecnica [Maz98a]. tectonics [Men86]. Teeter [Man93b]. Teich [Har92b, Mar94b]. teksten [Bur99]. Telegraphy [Noa99, Nye93b, Hun96]. teleology [Len82]. Telescope [Aga94b, Bro91a, Aga94c, Rob92, Smi89]. Teller [Edg91a, Yor89]. Telling [Law98c, Lee96, SY96]. Temkin [Daw98a]. Temporal [Ben90a, Whi88]. temps [Poi04]. Temptations [Nye99]. Ten [Lar96a, NBR^{+}95]. Ten-Volume [Lar96a]. ten-volumes [NBR^{+}95]. teologia [Mül93]. Tercentenary [Mul91, HC90]. Terence [Smi99b]. ternary [How96]. Test [SG80]. Testi [Bal91b, Mor96a]. tests [Kol91]. Texas [Bro93a, Hil93]. Text [Lor90, Mol94a, BM89, Coo90, NR86]. Textbooks [BV90]. Texte [Hin91, Van91, Non93b]. textes [CLR87]. Textile [KN88]. Textos [Goo91b]. Texts [Mic98, Ost98, Bur99, Jac95a, KRS83]. Thackray [Di 93f]. Their [Hal90a, Hor99b, Twe93b, BR96, CMNH92, Gro89, Hen93b, MKW92, Par99, Pum91c, RCRC98, Sal85, Wol87, vLvHH^{+}88]. them [Fis98a]. Themes [Div98, Fre93, Fox95, Osl91, Reh83, Spe82]. themselves [Pan97]. Theodor [Bro92c, Hei93, Kus95d]. Theodore [Har99b, Hod91, Har92a, Spa98b, Wis96]. Theologica [Met93, Hüb90]. Theological [Di 93b, Gre92b]. theology [Zam92, Fyf97]. Theophrastus [Daw98a, PTS^{+}96]. Theoretical [Kni95, Can94, Jun89, KT87, Nye93a, Sch89]. theoretische [Di 92, Jun89]. theoretischer [Koe91, Sch89]. Theoricae [Mel98]. théorie [Chu91, Duc87]. Theories [Car88, Cro90b, Fie91, Fra90, Gra91, Mal94, Ore90, Smi93a, BaJANJ84, Bar89a, BaJANJ84, BR96, Bow83, Can83, Cle93, Cor88a, Ear89, Hat90, LeG88, Sha93c, Sor83, Duc82]. Theory [Boh93, Bow94a, Bro93a, Bro98, Can90b, Duh91, Hoc91, Jor96, Law92a, Mar93a, Mar94a, Mar96c, Mol92, Mol98a, Twe93c, Ano82d, B^{+}82, BR91, Buc85, Buc89, BJ91, CD81, Cha82, Cra96, Duc87, Fre95, Gra85, Hen86a, Kil94, KI92, McI85, MR82d, MR82e, MR82f, MR82g, Nye99, Par99, Pet90, Pop82, Pou93, RR91, Ric92b, SO85, Sie91, TC91, Wei92, Wus84, Bro84]. therapeutic [War86]. therapy [Joh84]. Thermodynamics [Law92e, HS99, Kre90, Liu92]. Theroux [HS99]. Thesis [Hen92b, CDS90]. Theunissen [McK96]. Things [Alb92, Mal93a, Ara91, WHS92, Ste91c]. Thinkers [Hen90b, Tur87]. Thinking [Bro95b, Kni98a, Por86]. third [LS76]. Thirteenth [Hil91c, Mar92a, Whi90b]. thirties [Hen84a]. thirty [vLP82]. thirty-two [vLP82]. Thoemmes [Hen94a]. Thomas [Bro90c, Gun84, Hoc92b, Jam94c, Jor94, Lor90, Mar93c, Mar96a, Mol98b, Sch91a, Sch95b, Sho93c, Str90, BM89, Di 92, EJ^{+}07, ER93, EJR95, Gou98b, Hac90e, Hor96b, Jen91, Kol92, MR89, Pro93, Sch91a, SSH85, Shi83, Sou93, Sta84b, Tim98, Tom84, Wat91]. Thompson [Mur95b]. Thomson [Bro99c, Coo90, Mic98, Smi92b, STW90, Dah97, DF97, Kim95]. Thorbecke [Gre93a]. Thoren [Fie93, Gra92]. Thought [Bow91a, Fre93, Gou92, Har90b, Hun90, Jon93, Jon94, Kla95, Kra90a, Man93b, Nic93a, Osl98, Ost93a, Por97, Wal81, AW87, Bec91b, Cha88, Cro90a, Cro96a, Deg91, Dob91, Hal92b, Har87, Kre82, Mit92, Moo89, Osl91, Pit93, Tri89]. threads [SW95]. Three [Di 93f, Ros86, GT92]. thrusting [Mor92b]. tides [BR96]. tigers [Edg93]. Tilting [Edg93]. Tim [Mil95]. Time [Ben90a, Far99, Gra91, Kra93b, Lan82, Law93f, Poo98, Sor83, Whi88, Adl87, Ano87b, Cha83, Ear89, Gol92a, Poi04, Ros84b, Rud92b, Tod83, Wil87a]. Times [Hal91a, Har99b, Hen92a, Hor99b, FB97, Fei90, Hil84, Olm92, RCRC98, Wil87a]. Timms [Blo93b]. Timothy [Ben94b]. Tinkler [Tha91]. TISC [Hac95]. TISC-Conference [Hac95]. tisseurs [Cha84]. title [Gre86]. titre [Gre86]. Toby [Hen95a]. Today [Bar93b, Mur92a, Cro91a, MMVZ90, Rus88]. together [AL87]. Tokyo [Hac92a]. Toland [Sul82a]. Tom [Hoc92d]. Tomash [Twe92b, Twe93b]. Tome [CLR87]. tomes [Bai91]. Tomlinson [Law93g]. Tommaso [Per93, Ern91]. too [McC82]. Tools [Gou98a, But96]. Tooth [Kla95]. Torah [Kel91]. Tore [Bro93b, Hal92d, Kni90c, Spa95, BV91]. Toronto [Hac92a, Kni92a, Rus94, Sho93c]. Torricelli [Non94, dGB^{+}89]. torsion [BD95, Mor96a]. Tort [Bow93b]. Torti [Law91a]. tot [Hen86b]. Total [Old95]. tournant [FM00]. toutes [SGS00]. tra [Lan91, Pan91a]. Tractatus [WW88b]. trade [THBR92]. Tradition [Bru96, Cor93, Hen98b, McG95, Nut93, Pri93, Ree90, Whi90a, Ano07, C^{+}95, Deb91, Lin92, Liv93, PR88, Vic91]. traditional [Bur99, CK90, yQ85]. Traditions [Bow92b, Di 93b, Law93g, Mol94b, She92, Wea91, Dun90, EiT92, ET92, Gre92b, Mai91, PW89, RT92, Tak92]. traduction [BN82, Non93b, Sch01]. traen [Goo91b, ML89]. Tranquility [Fre93]. tranquillity [Osl91]. Trans [Ver90]. Transactions [Di 95a, Hil91c, Hun92]. transfers [TBDA10]. Transformation [Met96b, Tim98, Hey81, Kus95g]. Transforming [Aga98, Bow92b, Mai91, NOF96]. Transistors [Hoc92b, ES90]. transition [Moy83]. Translated [Fox92, Hod90a, Mol94a, Aga94a, Ams92, Ari90, Ben93, Bra98, Bro92c, Cam92, Dav95, Daw98a, Di 93c, GG91b, Hel96b, Hil93, Hoc92b, Jam94c, Jor94, Kus95b, Mal96, Mar96c, Mic98, Mil94, Mur92a, Olb91b, Roh99a, Sha97, Sut97, Twe91c, Twe92b, Twe93b, Wei94, Yeo96a]. Translation [Aly90, Lor90, Mol94b, Mol98b, Mon93, Smi92a, BL83, BM89, Coo90, Cop92, Dra81, EiT92, EBT96, ET92, HS91, Hüb90, Mun98, OH85, Ost99, Sch01, SSH85, Tak92, VoC89, War96a]. translators [Par99]. Transmission [Nye92b, Cle93]. transmutation [BGH^{+}87]. transport [FS84]. Transports [CL96, Har98d]. trapped [Ros86]. trash [Raf88]. Travaux [Dav94, GG90, Lan87]. Travel [Har98d, CL96, Sta84a]. Traveling [GG92]. Travels [Jam92a, BB82, BS91]. Travis [Rus94]. Treasure [Woo95, Far95]. Treatise [Mol92, Mol98b, BJ91, EBT96]. Treatises [Daw98a, Non93a, PTS^{+}96, Wal92, Top92]. treatment [Hun99b]. Treves [Law90c, Tro89]. Trevor [Gol94, Hig94, Mor91c, Non91a]. Trial [Non91b, Lin84, Wes89]. trials [Hüb90]. triangle [Edw87]. tricentenaire [Pic87]. tricentennial [Pic87]. Trigger [Bow91a]. Trigonometric [Ger89]. Trigonometric-astronomical [Ger89]. Trigonometrisch [Alt91, Ger89]. Trigonometrisch-astronomisches [Alt91, Ger89]. trigonometry [Viè83]. Trinity [MT83]. Triumph [Gou98a, But96]. Trollinger [Lar96a]. Trombley [Law90c]. Tropical [Har92e, Har96d, Liv99, Gro95b]. Tropp [Kra94]. troughs [Jan98]. Troughton [Tur93a, McC92]. Trust [Por95c, Wis96, Dea95]. Trustee [Gun84]. Truth [Blo93a, Cra92a, Sch95b, Ada89, Men86, Sap90, Sou93]. tuberculosis [Bry88, Smi88]. Tudor [Jon88]. Tumultuous [McK97b, Cre93]. Tunbridge [Mar93b]. tunnelling [Wes88]. Turin [Di 93a]. Turing [Her88]. Turkle [Rou97]. Turku [Bro96a, Kal95a]. turn [FM00, Kle94]. Turner [Bry91, Bry92a, Gou95a, Hac92c, Hen90b, Kre96, ABR93, Jon88]. Tuscaloosa [Bow93c, Hac90e]. tussles [Ros94b]. Twayne [Boh95]. twelve [Gje84, Spe82]. Tweney [Jam93b]. Twente [Alb92]. Twentieth [Hen90b, Hor99b, Hug96, Lar96a, Pal91a, Cre91a, Jac95a, NBR^{+}95, Arm83, AW87, Bra89, Bry88, Bur99, RCRC98, RT95, Swa88, Tur87, WR92, WD82]. Twentieth-Century [Lar96a, Jac95a, NBR^{+}95, AW87, Bry88, RT95, Swa88, WR92, WD82]. twintigste [Bur99]. Two [Smi92d, Bai91, For98, GB95a, GG92, Gre86, Kol91, Rit90, Smi92a, vLP82]. Twyman [Jam90]. Tycho [Fie93, JK84, Wil91b, Ros86, TC90]. Tychone [JK84]. Tyne [Fau96b, Lan90]. Tynemouth [Kno90]. Tytus [Ams92].

u [PHvMW93, PvM85]. u.a [HP85]. U.S [Hac90e, Wri94a]. U.S. [Man88, McC93, Wis85]. Übersetzung [Hüb90, Met93]. Ucko [Mur92b]. UCL [Far99, Law96b]. Uitgeverij [Hac93b]. Ulfried [Wei94]. Umberto [Ost94]. Umbruchs [Met96a, Sch93b]. unbelief [Lig87]. Uncertain [Dol96, Ful98]. uncertainty [Sti86]. uncharted [GG92]. Underground [For91, Wil90b]. Understanding [App93, Jam94a, Kni92b, Smi93d, vdVV91, Gol91a, Ner87, Sut94]. Uneasy [Jak84]. Unified [Rei94]. Union [Bry91, Mol94a, Sil90, Gra87]. Unit [Law90a]. United [But94, Nye92b, Por91b, AB89, De 92, HAB92, Hun85, Rot93a, Smy89, TT86]. Units [Mar93b, Tun92]. univers [Bla93, Hal95]. Universal [Sla82, Her88]. Universe [Aga94a, Fis98b, Jor96, Kat90, Kra94, Roh99a, Whi90d, Bla93, Car88, Hir83, KA81, Kil94, PB82, Rud88, SG96, SN93, Smi82, TFC93, Van85, vH97]. Universitaires [Bow93b, Cro96d]. Universität [AH93, Hel96a, Hes95a, Kra91c]. Universitäten [Tur93b, ZH92]. Universitatis [Whi97]. Université [Rus92]. Universiteit [Alb92]. Universities [Ras96a, Don91, Nyh95, Sir87, Wei83, ZH92]. University [Aga93, Aga94a, Aga94b, Aga95a, Aga96a, Aga97c, Aga98, Aly90, Ari90, Bai93, Bar92, Bar90a, Bar93b, Ben90a, Ben94b, BV91, Bev90, Bis98, Blo93a, Blo93b, Blo94b, Boh93, Bos92, Bow94a, Bow91a, Bow92c, Bow92b, Bow93a, Bow93c, Bra90, Bra92a, Bra92b, Bra95, Bra98, Bro90a, Bro93b, Bro93a, Bro91a, Bro93d, Bro90c, Bro92c, Bro99c, Bro90e, Bro91b, Bru96, Buc92, Bug96, But92a, Cam90, Cam92, Can90b, Can91a, Can96, Car95, Cha90c, Coo90, Coo96, Cro96c, Cuo95, Dan92, Daw98a, Dea93, Di 93b, Di 93c, Di 93d, Dol94, Dol98b, Dol99, Dol90a, Edg90a, Edg91a, Fie93, Fie94, Fis98b, For90a, Fox92, Fra90, Fre93, Ful98, Gau99, Ger91, Gol91b, Gol93, Gol94, Goo97, Gou98b, GG91c, Gre91, Gre92c, Gre96, Gro93, Hac90e, Hac95, Hak95b, Hal91a, Hal91b, Har92a, Har98b]. University [Har98c, Har99b, Har90a, Har93a, Har96a, Har98d, Har90b, Har92b, Har93c, Har95b, Har96d, Har97, Har98g, Har98h, Hay98b, Hef96, Hem95, Hen90a, Hen91b, Hen92a, Hen92b, Hen94b, Hen95a, Hen96, Hen98a, Hen98b, Hen99, Het95, Hig94, Hil91b, Hil93, Hoc90a, Hoc91, Hoc92a, Hoc93a, Hoc93b, Hod90a, Hod90b, Hod91, Hod92a, Hod96, HS90, Hor95, Hor96a, Hor96b, Hor97a, Hor97b, Hor99a, Hug92a, Hug98, Hun92, Ili95b, Ili98, Ink90, Jac96, Jac90, Jac93, Jam93a, Joh93, Jon93, Jon94, Jor91, Jor94, Jor96, Kat90, Kla95, Kni90a, Kni92a, Kni95, Kra90a, Kra90b, Kra91a, Kra91b, Kra92, Kra94, Kre96, Kri91, Kus95c, Kus95e, Kus96a, Kus96b, Kus97, Law91b, Law92a, Law93b, Law93c, Law93f, Law96a, Law96c, Law98b, Law98c, Le 91, Lev94, Luc93, Mag96a, Mal93b, Mal93a, Mal94, Mal95a]. University [Man93b, Man97, Mar93a, Mar93c, Mar96b, Mar96a, Mar94b, Mau95, Maz96, McK96, Met96b, Mic98, Mil95, Mol94b, Mol96, Mon93, Moo92, Mor92a, Mor90d, Mor91c, Mor92d, Mor93, Mor95c, Mor98, Mur92a, Mur96, Nea98, Nic93a, Nic93b, Nic96a, Nig99a, Nig99b, Nig99c, Nut93, Nye92a, Nye93b, O’B92, O’B96a, Olb91a, Ore90, Osb90a, Ost93b, Ost93c, Ost93d, Ost94, Ost98, Ost99, Pal90, Pal91a, Pal91b, Pal96a, Pal98, Pau90, Pic92b, Pic93a, Pic98, Poo99, Por90b, Por91a, Por91b, Por91c, Por97, Pri93, Pum90, Pum92, Pum94, Pum95a, Pyl93a, Ras96a, Ras96b, Rav99, Ree90, Roh99b, Roh99a, Rou97, Rud92a, Rus94, Sch91a, Sec96, Sec98, Sha97, Sha92, Sha93a, Sha96, She92, Sho93b, Sho93c, Sho96a, Sim92, Sim94, Slo97, Smi90a, Smi91a, Smi92c, Smi92b, Smi92d, Smi93b]. University [Smi93c, Smi95b, Smi96, Smi98, Smi99a, Smi99b, Spa91, Spa98b, Sta90b, Ste91a, Stu91, Stu98, Tim98, Tur93c, Twe93c, Uim95, Ver90, Ver94, Ver95, War91b, War95b, War96a, War96b, Wat91, Wea91, Wei94, Whi90d, Wil90a, Wil91b, Wil94a, Wil97, Wis96, Woo98, Wri94a, Yav92, Yeo96a, Yeo96b, Yos99, Gou97, Ken86, Mor91b, AH93, Fyf97, Kra91c, Ric97, Rob96a, Sme97]. university-industrial [Ken86]. unknown [vPvP90]. Unpublished [Hal94a, Ber93b]. Unreason [Por97]. Unseen [Law93g, Sta91]. untitled [Bro90b, Bro90d, Bro95e, Che90, Col95, Di 95b, Dol95a, Dol95b, For90b, Fre90a, Hac90b, Hac90c, Hal90b, Har90c, Hen95c, Hoc90b, Hun95a, Jac95b, Kea90, Kni90d, Kni90e, Kri90, Kus95f, Law90d, Lon90, Man90, Man95, Mar90, Mau90, McN95, Mur95c, Nas90, Nic95, Nut90, Pic95, Rog90, Sch90b, Sho95, Smi90c, Smi90d, Smi95a, Sta95, Wit95b, Wri95, Yod90]. Unwin [Bra91, Mur92b, Tha91]. unwitting [HK83]. updated [Di 93c]. upon [Fau96b, Lan90]. Uppsala [Bro98, Lar96b, Whi97, Wit95a]. upright [Hei86]. Upsaliensis [Whi97]. Urania [Bra91, Kit89]. Uraniborg [Fie93, TC90]. urban [May97a]. Ursula [Bro95a]. Ursum [JK84]. Ursus [JK84, Ros86]. USA [Hal92d, Har93b]. Usages [Mor96a, BD95]. Use [Buc90, Mol95, Non93a, Law93a, Loa98, SW93, Wal92]. useful [THBS92]. usefulness [Mil99]. Uses [BD95, Ver94, Bud93, Bur83, May97b, Mor95a]. USSR [Vuc84]. Ute [Har94b]. utilitarianism [dCC09]. utilitarisme [dCC09]. utility [Mil99].

V [Ben93, Fau96b, Gou91, Kni98b, Kra94, Lan90, Law98a, Mal95a, Mur92a, Rav99, Sha97, Woo95, Di 95a, Mal93b, Sec96, War90a]. Vagliente [Mar96c]. Valencià [LN95, Mal98, Mal98, LN95]. Valerie [She92]. Vallisneri [Rap92]. Valsiner [Smi93d]. Values [Ras96b]. Vance [Lar96a]. Vapours [Law93d, SE91]. Variorum [Bry92a, Gou95a, Kni98a, Luc98, Ung92, Sch91a]. Varley [Noa99]. Vatican [Ben93, Bro93d, Non91b, Sha97, Maf91]. Vaticana [Ben93, Non91b]. Vaticana/Pontificia [Ben93]. vector [BR91]. Veer [Smi93d]. veka [KI92]. Veneto [Di 98]. Venice [Di 98]. Veragsbuchhandlung [Sch93c]. veranstaltet [CM90]. verba [Sut94]. Verbindung [Bro95a, CM90, Kle94]. verbis [Sut94]. Vergata [Di 93a, Mül93]. verisimilitude [Tre97]. Verlag [Bro94a, Di 92, Fuc99, Gre92c, Gre93a, Gre93b, Har94b, Hel96a, Hes95a, Hug96, HD97, Jam94d, Mol92, Mol94b, New92, Tur93b, Van91, Wei90]. Verlag/Deutsches [Bro94a]. Vernon [Sho93c]. Verrier [Ros82]. vers [GG96, PP95]. Version [Mol94b, BEoCF92, ECL92]. Verständnis [Gol91a, Gre93b]. versus [Ber93b, De 96, Ear89, Gra91, Hal94a, Jan98]. Vertebrate [Bow93c, Buf87, Rai91]. Vertreibung [Dei92, Har94b]. vertues [Far95]. very [May99, Mol90]. Vestiges [CS94, Yeo96b]. veterinary [Ald85, Pat83b]. vi [Bow94a, Cro96c, Hil91c, Hoc91, Kni92b, Moo92, Ore90, Sch91a, Sil90]. vibration [CD81]. Vico [Ros84b]. Victor [Mar96c, Mal98, Fie93, Gra92]. Victoria [Bro95d]. Victorian [Bev90, Cri90, Dol94, Edg90b, Hay98b, Kni90a, Mur96, Nig99a, Sec98, Smi91a, Sta89a, Des82b, Des84, Goo90b, Goo91a, Hey81, Lig87, Liv99, Mer89, Mor96b, Pen89, Ric88, Rus89, Sec86, Sho96b, Sho86, Shu96, Shu04, Van93, Wil87d, Yeo93, You85]. Victorians [Smi91c, Bow89]. Videnskabernes [Hal94c]. vie [A^{+}87, BH90, Pyl92, dG01]. Vienna [MR87a]. View [BV90, Fra90, Bar89a, Hei93, Kni86, Kni88, Kni89, MJ89, Mus98]. Viewing [Hug92b, Mac90a]. Views [BD90, McK96, DJS97, De 92, Gre82, Kel98]. Vignoles [Vig82]. VII [Bos92, Aga97b, Bal91b, BV91, Bro93d, Can96, Cro96b, Div90, Div98, Ger95, Gou91, GG96, Gre92c, Hen90b, Hod96, Pal90, Pic98, Rob93, Spa98a, Twe92a, War90b]. viii [Aga95a, Ams92, Cor93, Elm91, Har98d, Har93c, Hin91, Kni90c, Kni92b, Law90a, Lee90, Luc98, McC95, Mol94b, Mor91c, Mor92c, New92, Nic94a, Nig99a, Nye92a, Nye93b, Pau90, Ras96b, Sha97, Spa98b, War91a, Whi90c, Wit95a, Wri96b]. Vilhelm [Kra90b, Fri89]. Villard [Kel95, Bec91b]. Vincent [Cro96b, Gol95b]. Virginia [Sel10]. Virtual [Rou97]. Viscountess [CMNH92, Hen93b]. visible [Sor99a]. Vision [Can91a, Kre96, Smi92a, Wit95a, AS89, Ath90, Eri94, Par87, Tur94]. Visions [Cri90, Dol98b, Dol99, MR96, Ric88]. visited [Cla97]. Visual [Joh96, Jac95a, Kel98, Sta94]. Vita [Met96a]. Vital [Pyl93a, Sut97, CD94, Cha91]. Vitaly [Aga94a]. Vitamin [Law92c, Ric91, Car86]. Vittorio [Kus95d]. Vivian [Bru96]. vocation [Out84, Sec91]. Voetius [Osl97, vR95]. voie [GG90, HJF89]. Vol [Cam90, Wil91b, Ano82c, Ano82b, Bow94a, Bow92c, Bur93, Fie92, Hen94a, Hil91c, Kni92b, Roh99a, War95a, War95b, BH87, EJR95, KN88, MR82e, NLH^{+}84, STT87, SCS^{+}87]. Volcanism [Fri91]. vols [Bur98, Gau99, Gou98a, Hen95b, Mol92, Mol94b, Roh99a, Sch91a, Smi92a, Spa98b]. Volta [Mor93, Per92]. Voltaire [Cas93, VB80]. Volume [Bai91, Bro93a, Bro99c, Buc92, Cro96d, Fra90, Fyf99, Hod93b, HD97, Kra98, La 97, Lar96a, LM23, Lor90, Mar92a, Mar93c, Mar96a, Nye92b, Por97, Pum92, Rap92, Sil90, Smi92b, Twe93b, War96a, Wil97, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, HP85, Spr97, Tay87, Wei92, Cla95a, NS84]. volumes [Ano97c, NBR^{+}95]. volunteers [Fis98a]. vom [Bro95a, CM90, Gol91a, Gre93b, Kle94]. vor [Alt91, Ger89]. Vorlesung [Hel96a, Kra91c]. Vortices [Mar93a, Sie91]. Voyage [BD90, Sta84a]. Voyages [BD90, Dol98b, Dol99, DJS97, MR96]. Voyaging [Bro95f, La 97]. Vries [The94]. VRIN [Out94, Chu91, Neu92, Pyl92]. vs [Hac90e, Man88]. Vulcan [Gou98a, But96]. Vulgarly [Ost98, BDH96]. Vulkanismus [New92, Fri91]. Vulkanismus-Neptunismus-Kontroverse [New92, Fri91]. Vulkanismusstreit [New92, Fri91]. Vygotsky [Smi93d, vdVV91].

W [AT87, Bow98, Bro90a, Bro91a, But94, Cas93, CMN95, Col93, Cuo95, Di 93f, Gou95a, Har96a, Hod93b, Hug92c, Joh93, Jor96, Lar96a, Law93g, Maz96, McC95, Mor92d, Spa98b, Tra92, Twe92a, Twe92b]. W. [Hen95b, Mag96b]. Wagener [Mar96c]. Wagner [Por95a, Whi93]. Wagner-Jauregg [Por95a, Whi93]. Wahlverwandtschaften [Adl87]. Wainwright [Swa92]. Wake [Dea93, Seg91]. Wall [Kra92]. Wallace [Bec99, Non93a, Smi92c, Bro84, WS91a]. wallcharts [Buc98]. Wallingford [War91a]. Wallis [Fau96b, Lan90, Pro93]. Walter [Hoc92a, Pos90, Kra90a, THBM92]. Walters [Cas93]. Wanted [Gre94]. War [Aga93, Aga98, Edw96, Hor92, Hul99, Lar96a, Num95a, Nye92a, Hod92b, Luc90, PLP92, PCA91]. Warburg [Can90a, Pri91, Smi92a]. Ward [Pro93]. Wardy [Bos92]. warfare [Hac84]. warming [Mar10]. Warren [Rot93b]. wars [Gre86]. Warwick [Lee90]. Was [Mal94, Gar98, Hal93]. Washington [Edg90b, Hor99b, Mal93b, Sil90]. waste [De 92]. Watanabe [Bro92c]. Watelet [Cro94]. Water [Law92b, Sha84, Dra81, Ham90, Ole84]. water-lifting [Ole84]. Waters [Har92a, Ili97, Ver95, MW94, Ste91b]. Watson [Elm91]. Wave [Can90b, Buc89, MR87a, OP87, Ste97a]. Waveforms [Phi87]. Waves [Goo97, Buc94, Jan98]. Way [Het95, HJF89, Pau93]. Wayman [Wil90a]. Ways [Pic93b]. weapons [De 92]. Wear [Bru96]. Weart [Hoc90a]. Weather [Aga97b, Kra90b, Fri89, Neb95]. Weatherall [Pic92b]. weavers [Cha84, SW95]. Web [Aga96b]. Webster [Stu98]. Weekes [Twy88]. Weeks [Jam90]. Weighed [Col93, HE92]. Weight [Mol96, Høy94]. weights [Con87]. Weikersheim [Gre93a, Wey92, Wey92]. Weindling [Di 93e]. Weingart [Wei90]. Weintraub [Bow94a]. Weldon [Mag96b]. Wellcome [Law90a, Spa98a, AL87]. Wellsprings [Gou95b, Luc98]. Welt [Met96a, Sch93b]. Weltauffassung [Kus95d, Hei93]. Wende [Bro95a, Kle94]. were [Cla97]. Werke [HD97, Met93, KvDC^{+}37, vPSBD^{+}04]. Werner [Bro94a, BEFH10, Kra97]. West [Hen95a, Kni98b, Huf93, Huf95]. Western [Bro92c, Bru96, Hen98b, Nut93, Por97, Pum92, Rou97, Rus94, Hay94, Tra93, C^{+}95, Lin92, ML89, Ols90]. Westfall [Ili97, Non91b]. Westman [Pum94]. Westminster [Gay98]. Westport [Blu90]. Westview [Law98a]. Weyer [Gre93a]. Wheaton [Hug96]. Wheatsheaf [Kus95b]. wheels [Wes95]. Wheland [Par99]. Where [Blo93a, Sap90, Ste99]. Whewell [Dol94, McN92, Yeo92, FS91, Fis91a, Yeo93]. which [Fis94, MJ89]. Whiffen [Mor92c]. Whigs [Bec95]. Whiston [For85]. White [Raf88, Sch95b, Sch91b, Sou93]. Whitehouse [Hun96]. Whiteside [Bra92a, Mal93a, WHS92]. Whitney [Hil91c, Wis85]. Whitrow [Ben90a, Por95a]. Whittet [Lan90]. Who [Gar98, Pum95b, Cre91a, CC94, Kra94, TFC93, Yox83]. Whose [Hoc94]. Wide [Aga96b]. Wiener [Aga95b]. wife [Van91]. Wijsbegeerte [Alb92]. Wilbur [Whi90a]. Wilderness [Whi97, Won93]. Wildman [Hun96]. Wilhelm [HD97, Roh99b, FMF85, Har96b]. Wilkie [Hoc92d]. William [Aga98, Dea92, Dol94, GG91c, Hal94b, Lar96a, McN92, Nea98, Non91a, Por90a, Smi92b, Tur93a, Twe91d, Yeo92, Aga96a, Ald85, Bir86, Bow92a, Bro85, BP02, CL87, FS91, Fis91a, For85, Fyf97, Gay96, Hal90a, Jan98, Jon88, Jon91, MD88, Man97, McE93, Nig99c, Non93a, O’D83, Par87, Pyl93b, RT95, Sch90a, Tay96, Twe93c, Yeo93]. Williams [Bry91, For91, Hoc92a, McC95, Mor91c, Twe93b, Wal90]. Willis [Wis85]. Willmoth [HD97, Joh94]. Wilson [Bra98, Coo96, Hod93a, Law96b, Smi92b, Wil91b, Wil97]. Winchester [Jon91, Sch91b]. Wind [Car95, Hil94].

Windmill [Car95, Hil94]. Windscale [Wyn82]. Winsor [Di 93d]. Wire [Nye93b, BC92]. Wisconsin [Jon94, Mau95, Rou97, Sho96a]. wisdom [Max84]. Wise [Ras96b]. Wissenschaften [Van91]. wissenschaftlich [Hei93, Kus95d]. wissenschaftlich-philosophische [Hei93, Kus95d]. wissenschaftlichen [Eck92, HRP90]. wissenschaftlicher [HP85, PHvMW93, PvM85]. Wissenschafts [Pos90]. Wissenschafts-philosophie [Pos90]. Wissenschaftskultur [Eck92, Mei91]. Wissenschaftsphilosophie [Mai88]. Wissenschaftspopularisierung [Fuc99, Dau98]. witch [Hüb90]. Within [Aga96a, Kuk91, Kuk93, She95, Tre97]. Wolfgang [Gre93a, Wey92, HP85, PHvMW93]. Womanhood [Bev90, Rus89]. Women [Gou97, Har98g, Hor99b, Hud92, HH97, Lyn99, Ogi86, RCRC98, SGS00, Ali86, Ano07, Cre91a, CC94, Out91, Sht96, Pur99]. wonderful [Cha97]. Wonders [Whi97]. Wood [Hac94]. Woodbridge [Elm97, Joh94, McC98, Pum91b, Pum91a]. Woodward [Cam90]. Woolf [Sel10]. Word [Joh93, Gil83, Gre94, Kro91]. words [Hul99]. Work [Aga96a, Sch95b, Sho93c, Spa95, Ack87, Ash85, Ber85, Can93b, CHG91, DFG87, FL^{+}94, GJ85, Jak84, Lev85, Mac85, Pum95b, She95, Sou93, dGB^{+}89, vLP82]. working [FG96]. Works [Cas93, Lar96a, War96b, AT87, BCKR89, B^{+}95, BN95, DD95, Gje84, GF^{+}03, KvDC^{+}37, Lan87, LM23, PN95, Tau93, VB80, vPSBD^{+}04, LF97]. Workshop [Mar93c]. World [Aga98, Cle83, Coo96, Ear89, Far94, Fie91, Ful98, GG91c, Gra91, Gre92c, Hac92d, Hil93, Lar96a, Law93f, Pic93a, Tur93a, Wil90a, Ale85, Bla93, BP93a, BP02, Cha97, Cro90b, Dol96, Edw96, GB95a, Gre82, Hei93, Kni86, Kni88, Kni89, Kri88, McC92, Mor91a, Mor90c, Noa99, Olm92, Opp85, Pai86, Pal96b, Pic92a, Ram88, Ron83, Rud92b, Sch90a, Sch93b, Tor95, Wes95, Wil95b, Aga96b, Num95a, Bro97b]. world-view [Kni86, Kni88, Kni89]. worlds [Cro86, Dic82]. Woronoff [Out94]. Worse [DPP88, Olb91a]. Wren [Ben82]. Wright [Bug96, Pro86]. writer [Dem83]. Writing [Rus93, Man93a]. Writings [Jon91, Lar96a, Mic98, Smi92c, Sut97, War95a, Yeo96b, CD94, CS94, CMN95, dLT96, Le 90, LS76, NN95, WS91b, WS91a, vLMB88]. Wyatt [Ben94b]. Wylie [Ley90, She92]. Wynne [Pic98]. Wynter [Twe93a].

X [Aga97c, Aga98, Bar93b, Bev90, Blo93b, Bow93c, Buc90, Cha90c, Cra92a, Cro96b, Di 93f, Edg90b, Fin93a, For91, Gol95b, GG90, Hac92a, Har99a, Har92b, Har93c, Hef96, Hen90b, Hud92, Ili97, Kra90a, Kra93b, Lar96a, Mal93b, MK98, McC98, Met93, Met96a, Mic98, Mil95, Moo92, Mur92b, Nic94a, Non94, Ost93b, Ost99, Out94, Por90a, Por91b, Pri93, Rap92, Smi91a, Smi94a, Smi96, Ste95, Twe91b, Twe93a, Twe93c, Ver95, Wil91b, Wil94a, Wit95a, Aga96a, Ait90, Bos92, Coo90, Coo96, Di 92, Dol90b, Fie94, Hak95b, Har90a, Hen90a, Hil98a, Hoc92a, Hoc92d, Hod90a, Hod90b, Hod92a, Hor99a, Hug96, Ink90, Jac93, Jon92, Jon94, Kra92, Kra94, Kus95b, Kus96a, Lar90, Law90c, Law91c, Law92d, Law98b, Ley90, Mag96a, Mar93a, Met99, Pal98, Pum91a, Pyl93b, Sec98, Smi92c]. x [Smi99b, Sta90b, Str90, Tim98, Wat91, Wil90a, Woo95, Bar90b, Hir83, Law92a, Olb91b, Osb90a]. X-Club [Bar90b]. X-ray [Hir83]. xi [Aga93, Aga94b, Blo93b, Bow92b, Bra95, Dea92, Fie93, For91, Ful98, Hac93a, Hal90a, Har99a, Har98g, Hen94b, Hoc93b, Hun98, Ili98, Jac90, Jam90, Jam92b, Jam92c, Joh94, Jon93, Kni90a, Kra90a, Kus94, Law93c, Luc96b, Mar94b, Mau95, Mil94, Mol98b, Mul91, Non91b, Odd91, Olb91b, Old97, Pal91b, Por91b, Pum92, Pyl93a, Sch95b, Sim92, Smi91c, Smi93d, Swa92, Tei93, Twe90, Twe93c, Wea91]. XII [Met93, Ung92, Ari90, Bar90a, Boh95, Bow92c, Bro92b, Bro99c, Bry92a, Can90a, Can91a, Cra92a, Cra92c, Cro94, Elm97, Fie91, Fis98b, Fox92, Fre93, Fuc99, Gol94, Gou92, GG91c, Gre91, Gre96, Hal91a, Har99b, Har96c, Hem95, Hoc93a, Hor96a, Hor97b, Hug92b, Hut92, Ili95b, Jam93a, Jor96, Kat90, Kni98a, Kra93a, Kus96b, Law91c, Law92e, Law96b, Law96c, Luc93, Man93b, Mar94a, Mil95, Mol98a, Mor99, Nic96a, O’B96a, Pal91a, Por91c, Por97, Ree90, Rud92a, She99, Sho93c, Smi91a, Smi92d, Smi93a, Smi95b, Smi99a, Stu91, Tur93a, Tur93c, Twe92c, Ver95, Wil91c]. xiii [Ben94b, Bis98, Bow93c, Bro98, Cri90, Dan92, Dol94, GG91b, Hal92d, Har98c, Har93a, Har90b, Hen92b, Hil91a, Hor98, Hud92, Jac96, Jam94b, Jor91, Kla95, Kri91, Law92b, Law93f, Law98a, Mar92a, Mar96b, Mor92f, Nea98, Ost93c, Pum90, Ras96a, Roh99b, Rou97, Sha93a, Shr92, Twe91b, Wil90a, Wil97]. XIIIe [Kel95, Bec91b]. XIV [Ait90, Aga96a, Aga98, Bar92, Bar93b, Bow91b, Bra98, Bro93b, Bro92c, Bru96, Buc90, Bur98, Cha90c, Edg91a, Fau96b, Fie92, Goo97, Gou98b, Gra91, Gro93, Har98a, Har92c, Har96d, Hay98b, Hen95a, Hen99, Het95, Hig94, Hod93a, Hor99b, Kra98, Kre96, Kus97, Kus98, Lar96a, Law92c, Law98c, Le 91, McK97b, McN92, Met96a, Mor90a, Mor90d, Mur92a, Mur92b, Nic93a, Nic93b, Nig99c, O’B92, Osb90a, Osl97, Pal96a, Pic92b, Por97, Pum91b, Pum95a, Spa95, Stu98, Uim95, Wil91b, Wil94a, Wis96, Wri94a, Yos99, Cla89]. XIX [Bal91b, Goo91b, KI92, Bai93, But92a, Dea93, Dol98b, Dol99, Far99, Fie94, Fyf99, Gou95a, Jam94a, Maz96, Mor95c, Sha97, Rey90]. XIXe [GG96, PP95]. xl [Lar96a, Sch93c]. xliii [Aly90]. xlix [Hen94b, HD97]. xlvii [Mar96a, MK98]. xlviii [Mar93c, Yeo96b]. xv [Aga94a, Bai91, Boh93, Bow91a, Bow92c, Bro95d, Bru91, Cro95, Di 93c, Edg95, Fra90, Ger91, Goo90a, Gre96, Har97, Hil91b, Hug92a, Hug98, Hun92, Kra91a, Law96a, Luc92, Mur96, Nic94b, Non91a, Oli92, Ost93d, Ost93a, Sha96, Sim94, Smi90a, Smi91b, Ste95, Twe91a]. xvi [Ben93, Col93, Di 93f, Edg91a, Har92a, Hef96, Hen98b, Hoc92c, Hug92c, Hun92, Kus95e, Mal93a, Mal95a, Man97, McC98, McE95, Nor94a, Osl98, Pri93, Sch91a, Sho93a, War91b]. xvii [Blo94b, Bow92a, Bro94b, Bro90c, Bro96b, Bro99c, Bro91b, Gau99, Goo96, Hen92a, Hir95, Hor97a, Jam92a, Jam93b, Jon91, Kni95, La 97, Lar96a, Law93e, Mal94, Nig99b, Pic93a, Pyl91, Twe93b, Ver94, War96a]. XVIIe [Poi04, Grm90, Pic87]. XVIII [Lor90, Blo98, Bro91a, Byn90, CK91, Cas94, Cuo95, Gre95, Har92d, Har98h, Hen98a, Hod91, Hor95, Jam91, Kra93b, La 94, Lar96a, Law91b, Mal94, Mol96, Mor98, Nut93, Pal97, Yav92]. XVIIIe [Poi04, SGS00, Roc88]. XVIIIème [BaJANJ84]. xvix [Hac94]. xx [Aga98, Hay98a, Hoc90a, Ili97, Kni98b, Kra90b, Lan90, Law90b, Law92a, Law93g, Lyn99, Mor92a, Nic93c, Nye93b, Olb91a, Sha97, Sho93b, Sho99, Smi93b]. xxi [Aga97c, Bra92a, Bro93c, Dol90a, Far94, Har95a, Har95b, Law93g, Lev94, Mol94a, Poo99, Rag95, Rus91, Slo97, Wei94]. xxii [Cam90, Cas93, Get96, Hoc92b, Joh93, Lar96a, Smi98, Spa98b, Ste91a, Woo98]. XXIII [Van91, Bug96, Har98b, Hen93a, Hen96, McE93, Non93a, Smi96, Tur91, Mor92d]. xxiv [Aga95b, Cam92, Can90b, Hel96b, Lar96a, Ost99, Por95a, War96b]. xxix [Bro90e, Hen93b, HD97, Joh96]. xxv [Nye92b, O’B96b, Ost94]. xxvi [Bow92c, Hal92a, Hen95b, Hil93, Hor92, Hor96b, Lar96a, Mor93, Smi93c, Smi94a]. xxvi-1081 [Hal92a]. xxvii [Bro90a, Bro99c, Pum94, Wil95a]. xxviii [Buc92]. XXX [Fie92, HS91, Law93d, Mic98, Spa91]. xxxii [Kra91b]. xxxiv [Gol93, Lar96a, Law93b, Smi99b]. xxxix [Sch91b]. xxxvi [Mar96c, Ost98]. xxxvii [Mur95b, Twe92b].

Ya [Kra94]. Yale [Hak95b, Har98d, Jam93a, Pal91a, Sha96, Smi92d]. Yans [HS91]. Year [Ano96o, Ano91s, Ano93s, Ano96q, Ano98g, Ano99e, Ben94a, Cri94, Cri96, Lan87]. Years [AH93, Ben93, Div90, Hal92c, Law91c, Mar96a, Moo92, Mor92c, Old97, War96a, All86, BH87, Bur99, Can97, EJR94, EJR95, FS84, Gle91, Gle82, Hac92c, H^{+}82, Hea86, Hea90, Her95, Maf91, Rus86a, SCS^{+}87, War95a, WH91]. Yellow [Col87]. Yeo [Dol94, Law98c]. Yesterday [Mur92a, MMVZ90]. Yorgos [Hac93c]. York [Aga94b, Aga96a, Aga98, Baa90, Bar90a, Ben94b, Blu90, Boh95, Bro94b, Bur93, Cra92c, Cri90, Cuo95, Edg91a, Fie91, Fie92, Get96, Hac92a, Hac92c, Hal92c, Hal92a, Har95a, Hen93a, Hen95b, Het95, Hoc92b, Hoc92c, Hod93b, HS90, Ink90, Jon92, Kni90a, Lar96a, Law92a, Law92e, Lee90, Mar92a, McK97b, Mol95, Mol96, Mol98b, Mor90a, Nye92a, O’B96b, Osl97, Pal90, Por90b, Pyl93a, Shr92, Sim94, Smi91b, Smi99b, Sta90b, Sut97, Tur93a, Ver91]. Yoshikawa [Hir95]. Young [Ley90, Smi99b, War91b, Wyl89, Bro89, Pye85]. Young-Eisendrath [Smi99b]. Yushkevich [Mar94a]. Yves [Sch93c].

Z [Can90b, Goo97, Jam94c]. Zacharias [Kra93b, Gol92a]. Zagorin [Ost93d]. Zambelli [Nor94a]. Zdenek [Mur92a]. Zeit [Adl87]. Zeno [Hod91]. Zev [Cas94]. Zilboorg [Daw98a]. Zivilisationskritik [Eck92, Mei91]. Zöllner [Eck92, Mei91, Mei91]. Zone [Sut97]. zoologiques [Rus92]. Zoology [Di 93d, Win91]. Zott [Tur93b]. zum [Bro95a, Gol91a, Gre93b, Kle94]. Zur [Eck92, Koe91, Gol91a, Gre93b, Hel96a, HRP90, Hin91, Mai88, Mei91, Pos90, Sch89, Sch84, Tur93b, Van91]. Zurich [MR87a]. Zvelebil [Mur92a]. zwischen [Koe91, Sch89]. zymotechnic [THBB92b].

References

Atran:1987:HDC

[A^{+}87] Scott Atran et al., editors. Histoire du concept d’espèce dans les sciences de la vie: colloque international, mai 1985. (French) [History of the concept of species in science and life: international conference, May 1985]. Editions de la Fondation Singer-Polignac, Paris, France, 1987. ISBN 2-900927-19-6. xiii + 324 pp. LCCN QH83 .H57 1987.

Aaserud:1990:RSN

[Aas90]  Finn Aaserud.  Redirecting

science: Niels Bohr, philanthropy, and the rise of nuclear physics. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-35366-1. xiii + 356 pp. LCCN QC789.2.D4 A27

                                                    1990. URL http://www.loc.
                                                    gov/catdir/description/
                                                    cam024/89048317.html;
                                                    http://www.loc.gov/
                                                    catdir/toc/cam028/
                                                    89048317.html.

Austoker:1989:HPR

[AB89] Joan Austoker and Linda Bryder, editors. Historical perspectives on the role of the MRC: essays in the history of the Medical Research Council of the United Kingdom and its predecessor, the Medical Research Committee, 1913–1953. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1989. ISBN 0-19-261651-X. xi + 259 pp. LCCN R772.A6 H57 1989. US$45.00. URL

http://www.loc.gov/
catdir/enhancements/
fy0637/89003122-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0637/89003122-t.html.

Anderson:1993:MIC

[ABR93]  R.  G.  W.

(Robert Geoffrey William) Anderson, J. A. (James A.) Bennett, and W. F. (William Francis) Ryan, editors. Making instruments count: essays on historical scientific instruments presented to Gerard L’Estrange Turner. Variorum, Aldershot, Hampshire, UK, 1993. ISBN 0-86078-394-4. xix + 492 pp. LCCN Q185 .M255 1993.

Ackerman:1987:JGF

[Ack87]  Robert Ackerman.  J. G.

Frazer: his life and work. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34093-4. x + 348 + 10 pp. LCCN GN21.F65 A27

1987. URL http://www.loc.
gov/catdir/description/
cam023/87009343.html;
http://www.loc.gov/
catdir/toc/cam023/
87009343.html.

Adams:1989:STP

[Ada89] E. B. (Edward Barry) Adams.

In search of truth: a portrait of Don Craib. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1989. ISBN 1-85315-119-X, (paperback), 1-85315-118-1 (paperback). xi + 123 pp. LCCN R654.C73 A35

                                        1989. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/20798932.
                                        html.

Adler:1987:FMA

[Adl87] Jeremy Adler. “Eine fast magische Anziehungskraft”: Goethes “Wahlverwandtschaften” und die Chemie seiner Zeit. (German) [A nearly magic attractive force: Goethe’s “Wahlverwandtschaften” [feeling of affinity] and the chemistry of its time]. Beck, München, Germany, 1987. ISBN 3-406-31559-3. 260 pp. LCCN ???? URL

                                        http://www.gbv.de/dms/bs/
                                        toc/024186546.pdf.

Aftalion:1991:HIC

[Aft91] Fred Aftalion. A history of the international chemical industry. The Chemical sciences in society series. University of Pennsylvania Press, Philadelphia, PA, USA, 1991. ISBN 0-8122-8207-8 (hardcover), 0-8122-1297-5 (paperback). xix + 411 pp. LCCN HD9650.5 .A3813 1991.

Agar:1993:BRA

[Aga93] Jon Agar. Book review: Alan Hodgkin. Chance and Design:

Reminiscences of Science in Peace and War. Cambridge: Cambridge University Press, 1992. Pp. xi + 412. ISBN 0-521-40099-6. £40. British Journal for the History of Science, 26 (2):255–256, June 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027119.

Agar:1994:BRA

[Aga94a]  Jon Agar.  Book review:

Alexander S. Sharov and Igor D. Novikov. Edwin Hubble, the Discoverer of the Big Bang Universe, translated by Vitaly Kisin. Cambridge: Cambridge University Press, 1993. Pp. xv + 187. ISBN 0-521-41617-5. £19.95, $34.95. British Journal for the History of Science, 27(4):479–480, December 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027632.

Agar:1994:BRP

[Aga94b] Jon Agar. Book review: Peter Robertson, Beyond Southern Skies: Radio Astronomy and the Parkes Telescope. Cambridge, New York and Melbourne: Cambridge University Press, 1992. Pp. xi + 357. ISBN 0-521-41408-3. £40.00. British Journal for the History of Science, 27(1):124– 125, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027593.

Agar:1994:MMB

[Aga94c] Jon Agar. Making a meal of the big dish: the construction of the Jodrell Bank Mark 1 radio telescope as a stable edifice, 1946–1957. British Journal for the History of Science, 27(1):3–21, March 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027578.

Agar:1995:BRG

[Aga95a] Jon Agar. Book review: Gerald L. Geison and Frederic L. Homes (eds.), Research Schools: Historical Reappraisals. Osiris 8 (new series). Chicago and London: University of Chicago Press, 1993. Pp. viii + 248. ISBN 0-226-28545-6. $39.00 (hardcover); 0-226-28546-4, $25.00 (paperback). British Journal for the History of Science, 28 (2):245–246, June 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027690.

Agar:1995:BRN

[Aga95b] Jon Agar. Book review: Norbert Wiener, Invention: The Care and Feeding of Ideas, with an Introduction by Steve J. Heims. Cambridge, MA and London: MIT Press, 1993. Pp. xxiv + 159. ISBN 0-262-23167-0. £17.95. British Journal for the History of Science, 28(1): 123–124, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027503.

Agar:1996:BRW

[Aga96a]  Jon Agar.  Book reviews:

William Sheehan, The Immortal Fire Within. The Life and Work of Edward Emerson Barnard. Cambridge: Cambridge University Press, 1995. Pp. xiv + 429, illus. ISBN 0-521-44489-6. £40.00, $49.95. Gale E. Christianson, Edwin Hubble. Mariner of the Nebulae. New York: Farrar, Straus and Giroux, 1995. Pp. x + 420, illus. ISBN 0-374-14660-8. $27.50. British Journal for the History of Science, 29(4): 486–488, December 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027706.

Agar:1996:HSW

[Aga96b] Jon Agar. History of science on the World Wide Web. British Journal for the History of Science, 29(2):223–227, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027836.

Agar:1997:BM

[Aga97a] Jon Agar. Back matter. British Journal for the History of Science, 30(2):254, June 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027732.

Agar:1997:BRF

[Aga97b] Jon Agar. Book review: Frederik Nebeker, Calculating the Weather: Meteorology in the 20th Century. San Diego: Academic Press, 1995. Pp. vii + 255, illus. ISBN 0-12-515175-6. $64.95. British Journal for the History of Science, 30(1): 101–121, March 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027905.

Agar:1997:BRR

[Aga97c] Jon Agar. Book review: Ronald E. Doel, Solar System Astronomy in America: Communities, Patronage, and Interdisciplinary Science, 1920– 1960. Cambridge: Cambridge University Press, 1996. Pp. xxi + 280, illus. ISBN 0-521-41573-X. £40.00, $59.95. British Journal for the History of Science, 30(2):233–249, June 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027720.

Agar:1998:BRM

[Aga98] Jon Agar. Book reviews: Martin Campbell-Kelly and William Aspray, Computer: A History of the Information Machine. New York: Basic Books, 1996. Pp. ix + 340. ISBN 0-465-02989-2. No price given. Paul N. Edwards, The Closed World: Computers and the Politics of Discourse in Cold War America. Cambridge, MA: MIT Press, 1996. Pp. xx + 440. ISBN 0-262-05051-X. £33.95. Arthur L. Norberg and Judy E. O’Neill, Transforming Computer Technology: Information Processing for the Pentagon, 1962– 1986. Baltimore: Johns Hopkins University Press, 1996. Pp. xiv + 360. ISBN 0-8018-5152-1. £41.50. British Journal for the History of Science, 31(3): 361–375, September 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027818.

Achinstein:1985:OEH

[AH85] Peter Achinstein and Owen Hannaway, editors. Observation, experiment, and hypothesis in modern physical science. Studies from the Johns Hopkins Center for the History and Philosophy of Science. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-01083-6. x + 379 pp. LCCN QC6 .O26 1985.

Albrecht:1993:NTG

[AH93] Helmuth Albrecht and Armin Hermann, editors. Naturwissenschaft und Technik in der Geschichte: 25 Jahre Lehrstuhl für Geschichte der Naturwissenschaft und Technik am Historischen Institut der Universität Stuttgart. (German) [Natural science and technology in history: 25 Years of the Chair of History in Natural Science and Technology at the Historical Institut of the University of Stuttgart ]. Verlag für Geschichte der Naturwissenschaften und der Technik, Stuttgart, Germany, 1993. ISBN 3-928186-15-9. 401 pp. LCCN Q183.4.G293 U556 1993.

al-Hassan:1986:ITI

[aHH86] Ahmad Y. al Hassan and Donald R. Hill. Islamic technology: an illustrated history. Cambridge University Press, Cambridge, UK, 1986. ISBN 92-3-102294-6 (UNESCO), 0-521-26333-6 (Cambridge). xiv + 304 pp. LCCN ????

al-Hassan:1988:ITI

[aHH88] Ahmad Y. al Hassan and Donald R. (Donald Routledge) Hill, editors. Islamic technology: an illustrated history. Cambridge University Press, Cambridge, UK, 1988. ISBN 92-3-102294-6 (UNESCO), 0-521-26333-6 (Cambridge). xv + 304 pp. LCCN ????

al-Hassan:1992:ITI

[aHH92] Ahmad Yusuf al Hassan and Donald Routledge Hill. Islamic technology: an illustrated history. Cambridge University Press, Cambridge, UK, 1992. ISBN 92-3-102294-6 (UNESCO, hardback), 92-3-102733-6 (UNESCO, paperback), 0-521-26333-6 (Cambridge, hardback), 0-521-42239-6 (Cambridge, paperback). xv + 304 pp. LCCN ????

Agricola:1990:BMD

[AHY90] Georg Agricola, Robert Halleux, and Albert Yans. Bermannus = Le mineur: un dialogue sur les mines. (French) [Bermannus: the miner: a dialogue on mines]. Science et humanisme, 0761-2885. Belles Lettres, Paris, France, 1990. ISBN 2-251-34504-3. xxx + 185 pp. LCCN TN144 .A3414 1990. 250F.

Aiton:1985:LB

[Ait85] E. J. Aiton. Leibniz: a biography. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-470-6. xiv + 370 pp. LCCN B2597 .A67

1985. URL http://www.loc.
gov/catdir/enhancements/
fy0745/85238184-d.html.

Aiton:1990:BRD

[Ait90] E. J. Aiton. Book review:

Desmond M. Clarke. Occult Powers and Hypotheses: Cartesian Natural Philosophy under Louis XIV. Oxford: Clarendon Press, 1989, Pp. x + 265. ISBN 0-19-824812-1. £27.50. British Journal for the History of Science, 23(1):115– 116, March 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026820.

Albertini:1989:CCL

[AJ^{+}89] Ippolito Francesco Albertini, Saul Jarcho, et al., editors. Clinical consultations and letters. Francis A. Countway Library of Medicine, Boston, MA, USA, 1989. ISBN 0-88135-089-3. lxix + 356 pp. LCCN RC66 .C54 1989.

Anderson:1987:SMD

[AL87] R. G. W. (Robert Geoffrey William) Anderson and Christopher Lawrence, editors. Science, medicine, and dissent: Joseph Priestley (1733–1804): papers celebrating the 250th anniversary of the birth of Joseph Priestley together with a catalogue of an exhibition held at the Royal Society and the Wellcome Institute for the History of Medicine. Wellcome Trust/Science Museum, London, UK, 1987. ISBN 0-901805-28-9. xi + 105 pp. LCCN QD22.P8 S35x 1987.

Albury:1992:BRD

[Alb92] W. R. Albury. Book review: Dirk Stemerding. Plants, Animals and Formulae: Natural History in the Light of Latour’s Science in Action and Foucault’s The Order of Things. Enschede: Faculteit Wijsbegeerte en Maatschappijwetenschappen, Universiteit Twente, 1991. Pp. 202. ISBN 90-365-0379-5. Dfl 40. British Journal for the History of Science, 25(4):468–470, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027059.

Alder:1985:BBL

[Ald85] Garry Alder. Beyond Bokhara:

the life of William Moorcroft, Asian explorer and pioneer veterinary surgeon, 1767– 1825. Century Pub., London, UK, 1985. ISBN 0-7126-0722-6. xiii + 417 + 8 pp. LCCN DS327.7.M67 A43 1985. US$16.95.

Alexander:1985:IQC

[Ale85] Peter Alexander. Ideas, qualities, and corpuscles: Locke and Boyle on the external world. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26707-2. 336 pp. LCCN B1297 .A28

1985. URL http://www.loc.
gov/catdir/description/
cam031/84017457.html;
http://www.loc.gov/
catdir/enhancements/
fy0903/84017457-t.html.

Alic:1986:HHH

[Ali86] Margaret Alic. Hypatia’s heritage: a history of women in science from antiquity to the late Nineteenth Century. Women’s Press, London, UK, 1986. ISBN 0-7043-3954-4 (paperback). ix + 230 + 8 pp. LCCN Q130 .A48 1986. US$4.95.

Allen:1986:BHB

[All86] David Elliston Allen. The botanists: a history of the Botanical Society of the British Isles through a hundred and fifty years. St Paul’s Bibliographies, Winchester, UK, 1986. ISBN 0-906795-36-2. xv + 232 + 8 pp. LCCN QK1 .A3647 1986. US$15.00.

Alter:1987:RPS

[Alt87] Peter Alter. The reluctant patron: science and the state in Britain, 1850–1920. Berg, Oxford, UK, 1987. ISBN 0-907582-67-2. 292 pp. LCCN Q127.G4 A6413 1987. US$45.00.

Alton:1991:BRA

[Alt91] E. J. Alton. Book review: Armin Gerl. Trigonometrischastronomisches Rechnen kurz vor Copernicus: Der Briefwechsel Regiomontanus-Bianchini. Stuttgart: Franz Steiner, 1989. Pp. 357. ISBN 3-515-05373-5. DM 70. British Journal for the History of Science, 24(2): 247–248, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027173.

Aly:1990:BRH

[Aly90] Amal Abou Aly. Book review: Heinrich Von Staden. Herophilus:

The art of medicine in Alexandria: Edition, Translation and Essays. Cambridge: Cambridge University Press, 1989. Pp. xliii + 666. ISBN 0-521-23646-0. £75.00, $140 (hardcover). British Journal for the History of Science, 23(3):340–341, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026765.

Amsterdamska:1992:BRI

[Ams92] Olga Amsterdamska. Book review: Ilana Löwy (ed.). The Polish School of Philosophy of Medicine: From Tytus Chalubinski (1820–1889) to Ludwik Fleck (1896–1961). Compiled, translated and with introductions by the editor. Dordrecht: Kluwer, 1990. Pp. viii + 287. ISBN 0-7923-0958-8. £44.00, $78.00, Dfl. 130.00. British Journal for the History of Science, 25(2):281–282, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027320.

Anderson:1982:CBO

[And82] Lorin Anderson. Charles Bonnet and the order of the known, volume 11 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1982. ISBN 90-277-1389-8. xv + 159 pp. LCCN QH31.B72 A8

Anderson:1984:BLL

[And84] Wilda C. Anderson. Between the library and the laboratory: the language of chemistry in Eighteenth-Century France. The Johns Hopkins University Press, Baltimore, MD, USA, 1984. ISBN 0-8018-3229-2. 190 + 1 pp. LCCN PQ618 .A5 1984. US$22.50.

Anderson:1992:WTE

[And92] R. G. W. Anderson. ‘what is technology?’: education through museums in the mid-nineteenth century. British Journal for the History of Science, 25(2):169–184, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027297.

Anonymous:1982:HFT

[Ano82a] Anonymous. Huygens et la France: table ronde du Centre national de la recherche scientifique, Paris, 27–29 mars 1979. L’Histoire des sciences. Textes et études. J. Vrin, Paris, France, 1982. ISBN 2-7116-2018-2. ix + 268 pp. LCCN Q143.H96 H88 1982. 210F.

Anonymous:1982:ICBb

[Ano82b] Anonymous. ISIS Cumulative Bibliography: a bibliography of the history of science formed from ISIS Critical Bibliographies 1-90, 1913– 1965: vol. 5. Civilizations and periods 15th to 19th Centuries.

Mansell Pub. with the History of Science Society, London, UK, 1982. ISBN 0-7201-1643-0. xi + 593 pp. LCCN Z7405.H2 I2 1982 vol. 5.

Anonymous:1982:ICBa

[Ano82c] Anonymous. ISIS Cumulative Bibliography: a bibliography of the history of science formed from ISIS Critical Bibliographies 1-90, 1913-65: vol. 4. Civilizations and periods: prehistory to Middle Ages. Mansell Pub. with the History of Science Society, London, UK, 1982. ISBN 0-7201-1642-2, 0-7201-0549-8 (set). xviii + 457 pp. LCCN Z7405.H6 I2 1982 vol. 4.

Anonymous:1982:SSS

[Ano82d] Anonymous. Science and speculation: studies in Hellenistic theory and practice. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24689-X, 2-901725-47-3. 351 pp. LCCN ????

Anonymous:1987:CDE

[Ano87a] Anonymous, editor. Conference on documentary editing. Cambridge University Press, Cambridge, UK, 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). 1–128 pp. Papers from the international conference held in Tarrytown, N.Y., April 11–14, 1985, British J. Hist. Sci. 20 (1987), no. 1(64).

Anonymous:1987:LTR

[Ano87b] Anonymous. L’enseignement technique de la Révolution à nos jours. (French) [Technical teaching in the revolution of our time]. Economica, Paris, France, 1987. ISBN 2-7342-0146-1, 2-7178-1323-3. 783 pp. LCCN ????

Anonymous:1987:SCC

[Ano87c] Anonymous. Science and civilisation in China. Cambridge University Press, Cambridge, UK, third edition, 1987. ISBN 0-521-08690-6. xxv + 485 pp. LCCN ????

Anonymous:1990:BVIa

[Ano90a] Anonymous. BJH volume 23 issue 1 cover and back matter. British Journal for the History of Science, 23(1):b1– b3, March 1990. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1990:BVIb

[Ano90b] Anonymous. BJH volume 23 issue 1 cover and front matter. British Journal for the History of Science, 23(1):f1–f2, March 1990. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1990:BVIc

[Ano90c] Anonymous. BJH volume 23 issue 2 cover and back matter. British Journal for the History of Science, 23(2):b1– b6, June 1990. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1990:BVId

[Ano90d] Anonymous. BJH volume 23 issue 2 cover and front matter. British Journal for the History of Science, 23(2):f1–f6, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1990:BVIe

[Ano90e] Anonymous. BJH volume 23 issue 3 cover and back matter. British Journal for the History of Science, 23(3):b1– b8, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1990:BVIf

[Ano90f] Anonymous. BJH volume 23 issue 3 cover and front matter. British Journal for the History of Science, 23(3):f1–f4, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1990:BRa

[Ano90g] Anonymous. Books received.

British Journal for the History of Science, 23(1):125–127, March 1990. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1990:BRb

[Ano90h] Anonymous. Books received.

British Journal for the History of Science, 23(2):253– 255, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026753.

Anonymous:1990:BRc

[Ano90i] Anonymous. Books received. British Journal for the History of Science, 23(3):381–383, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026797.

Anonymous:1990:C

[Ano90j] Anonymous.

Correction. British Journal for the History of Science, 23(1): 99, March 1990. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1990:EIS

[Ano90k] Anonymous.

Échanges d’influences scientifiques et techniques entre pays européens de 1780 à 1830. (French) [Exchanges of scientific and technical influence between the European countries from 1780 to 1830], volume 5 of Actes du 114e Congrès national des sociétés savantes, Paris, 3-9 avril 1989. Section histoire des sciences et des techniques; Colloques du CTHS, 0993-8192. Editions du CTHS, Paris, France, 1990. ISBN 2-7355-0208-2. 170 pp. LCCN Q127.E8 C663 1989. 200F.

Anonymous:1990:E

[Ano90l] Anonymous. Editorial. British Journal for the History of Science, 23(1):1–2, March 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1990:EA

[Ano90m]  Anonymous.  Editorial an-

nouncements. British Journal for the History of Science, 23(2):129, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:BVIa

[Ano91a] Anonymous. BJH volume 24 issue 1 cover and back matter. British Journal for the History of Science, 24(1):b1– b4, March 1991. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:BVIb

[Ano91b] Anonymous. BJH volume 24 issue 1 cover and front matter. British Journal for the History of Science, 24(1):f1–f2, March 1991. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:BVIc

[Ano91c] Anonymous. BJH volume 24 issue 2 cover and back matter. British Journal for the History of Science, 24(2):b1– b6, June 1991. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:BVId

[Ano91d] Anonymous. BJH volume 24 issue 2 cover and front matter. British Journal for the History of Science, 24(2):f1–f2, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:BVIe

[Ano91e] Anonymous. BJH volume 24 issue 3 cover and back matter. British Journal for the History of Science, 24(3):b1– b2, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:BVIf

[Ano91f] Anonymous. BJH volume 24 issue 3 cover and front matter. British Journal for the History of Science, 24(3):f1–f2, September 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:BVIg

[Ano91g] Anonymous. BJH volume 24 issue 4 cover and back matter. British Journal for the History of Science, 24(4):b1–b2, December 1991. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:BVIh

[Ano91h] Anonymous. BJH volume 24 issue 4 cover and front matter. British Journal for the History of Science, 24(4):f1–f2, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:BRa

[Ano91i] Anonymous. Books received.

British Journal for the History of Science, 24(1):129– 130, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027044.

Anonymous:1991:BRb

[Ano91j] Anonymous. Books received.

British Journal for the History of Science, 24(2):271– 273, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027194.

Anonymous:1991:BRc

[Ano91k] Anonymous. Books received.

British Journal for the History of Science, 24(3):397–400, September 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027253.

Anonymous:1991:BRd

[Ano91l] Anonymous. Books received.

British Journal for the History of Science, 24(4):502–504, December 1991. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:BRe

[Ano91m] Anonymous. Books reviewed. British Journal for the History of Science, 24(4):505–508, December 1991. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:C

[Ano91n] Anonymous.

Correction. British Journal for the History of Science, 24(2): 270, June 1991. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:E

[Ano91o] Anonymous. Editorial. British Journal for the History of Science, 24(2):131, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:EAC

[Ano91p] Anonymous. Editorial: Archive collections. British Journal for the History of Science, 24(4): 401–403, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:EN

[Ano91q] Anonymous. Editorial note. British Journal for the History of Science, 24(1):1, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:IA

[Ano91r] Anonymous. Index of authors.

British Journal for the History of Science, 24(4):509, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1991:RCY

[Ano91s] Anonymous. Report of Council for the year 1990–1991. British Journal for the History of Science, 24(4):511– 523, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027161.

Anonymous:1992:AC

[Ano92a] Anonymous. Archive collections. British Journal for the History of Science, 25(4):453– 456, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:BVIa

[Ano92b] Anonymous. BJH volume 25 issue 1 cover and back matter. British Journal for the History of Science, 25(1):b1– b7, March 1992. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:BVIb

[Ano92c] Anonymous. BJH volume 25 issue 1 cover and front matter. British Journal for the History of Science, 25(1):f1–f3, March 1992. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:BVIc

[Ano92d] Anonymous. BJH volume 25 issue 2 cover and back matter. British Journal for the History of Science, 25(2):b1– b4, June 1992. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:BVId

[Ano92e] Anonymous. BJH volume 25 issue 2 cover and front matter. British Journal for the History of Science, 25(2):f1–f2, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:BVIe

[Ano92f] Anonymous. BJH volume 25 issue 3 cover and back matter. British Journal for the History of Science, 25(3):b1– b10, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:BVIf

[Ano92g] Anonymous. BJH volume 25 issue 3 cover and front matter. British Journal for the History of Science, 25(3):f1–f2, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:BVIg

[Ano92h] Anonymous. BJH volume 25 issue 4 cover and back matter. British Journal for the History of Science, 25(4):b1–b6, December 1992. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:BVIh

[Ano92i] Anonymous. BJH volume 25 issue 4 cover and front matter. British Journal for the History of Science, 25(4):f1–f2, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:BRb

[Ano92j] Anonymous. Book received.

British Journal for the History of Science, 25(3):393–396, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027295.

Anonymous:1992:BRa

[Ano92k] Anonymous. Books received.

British Journal for the History of Science, 25(2):295– 298, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027334.

Anonymous:1992:BRc

[Ano92l] Anonymous. Books received.

British Journal for the History of Science, 25(4):501–504, December 1992. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:BRd

[Ano92m] Anonymous. Books reviewed. British Journal for the History of Science, 25(4):505–508, December 1992. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:EPN

[Ano92n] Anonymous.

Elements de la philosophie de Newton. (French) [Elements of the philosophy of Newton], volume 15 of The complete works of Voltaire. The Voltaire Foundation Taylor Institution, Oxford, UK, 1992. ISBN 0-7294-0374-2. xxii + 805 pp. LCCN ????

Anonymous:1992:IA

[Ano92o] Anonymous. Index of authors. British Journal for the History of Science, 25(4):509, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1992:RCP

[Ano92p] Anonymous. Report of Council for the period may-december 1991. British Journal for the History of Science, 25(4):511– 519, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027093.

Anonymous:1993:BVIa

[Ano93a] Anonymous. BJH volume 26 issue 1 cover and back matter. British Journal for the History of Science, 26(1):b1– b7, March 1993. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:BVIb

[Ano93b] Anonymous. BJH volume 26 issue 1 cover and front matter. British Journal for the History of Science, 26(1):f1–f3, March 1993. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:BVIc

[Ano93c] Anonymous. BJH volume 26 issue 2 cover and back matter. British Journal for the History of Science, 26(2):b1– b6, June 1993. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:BVId

[Ano93d] Anonymous. BJH volume 26 issue 2 cover and front matter. British Journal for the History of Science, 26(2):f1–f2, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:BVIe

[Ano93e] Anonymous. BJH volume 26 issue 3 cover and back matter. British Journal for the History of Science, 26(3):b1– b8, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:BVIf

[Ano93f] Anonymous. BJH volume 26 issue 3 cover and front matter. British Journal for the History of Science, 26(3):f1–f2, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:BVIg

[Ano93g] Anonymous. BJH volume 26 issue 4 cover and back matter. British Journal for the History of Science, 26(4):b1–b6, December 1993. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:BVIh

[Ano93h] Anonymous. BJH volume 26 issue 4 cover and front matter. British Journal for the History of Science, 26(4):f1–f3, December 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:BRa

[Ano93i] Anonymous. Books received. British Journal for the History of Science, 26(1):133– 136, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027395.

Anonymous:1993:BRb

[Ano93j] Anonymous. Books received.

British Journal for the History of Science, 26(2):265– 267, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027124.

Anonymous:1993:BRc

[Ano93k] Anonymous. Books received.

British Journal for the History of Science, 26(3):381–383, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:BRd

[Ano93l] Anonymous. Books received.

British Journal for the History of Science, 26(4):497–498, December 1993. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:BRe

[Ano93m] Anonymous. Books reviewed.

British Journal for the History of Science, 26(4):499–503, December 1993. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:C

[Ano93n]  Anonymous.  Correction.

British Journal for the History of Science, 26(3):383, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:E

[Ano93o] Anonymous. Editorial. British Journal for the History of Science, 26(3):269, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:IA

[Ano93p] Anonymous. Index of authors. British Journal for the History of Science, 26(4):505, December 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1993:SMA

[Ano93q] Anonymous. Les scientifiques et la montagne: actes du 116e Congrès national des sociétés savantes, Chambéry, 29 avril-4 mai 1991, Section d’histoire des sciences et des techniques. (French) [The scientists and the mountain: actes of the 116th National Congress of Scientific Societies, 29 April–4 May, 1991, section for the history of science and technology]. Actes du 116e Congrès national des sociétés savantes. Section d’histoire des sciences et des techniques, 0998-5654. Editions du CTHS, Paris, France, 1993. ISBN 2-7355-0272-4. 293 pp. LCCN Q127.E8 C663 1991.

Anonymous:1993:NEB

[Ano93r] Anonymous. New Editor for BJHS. British Journal for the History of Science, 26(3): 385, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027430.

Anonymous:1993:RCY

[Ano93s] Anonymous. Report of Council for the year 1992. British Journal for the History of Science, 26(4):507–516, December 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027479.

Anonymous:1994:BVIa

[Ano94a] Anonymous. BJH volume 27 issue 1 cover and back matter. British Journal for the History of Science, 27(1):b1– b6, March 1994. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:BVIb

[Ano94b] Anonymous. BJH volume 27 issue 1 cover and front matter. British Journal for the History of Science, 27(1):f1–f2, March 1994. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:BVIc

[Ano94c] Anonymous. BJH volume 27 issue 2 cover and back matter. British Journal for the History of Science, 27(2):b1– b8, June 1994. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:BVId

[Ano94d] Anonymous. BJH volume 27 issue 2 cover and front matter. British Journal for the History of Science, 27(2):f1–f2, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:BVIe

[Ano94e] Anonymous. BJH volume 27 issue 3 cover and back matter. British Journal for the History of Science, 27(3):b1– b10, September 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:BVIf

[Ano94f] Anonymous. BJH volume 27 issue 3 cover and front matter. British Journal for the History of Science, 27(3):f1–f2, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:BVIg

[Ano94g] Anonymous. BJH volume 27 issue 4 cover and back matter. British Journal for the History of Science, 27(4):b1–b8, December 1994. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:BVIh

[Ano94h] Anonymous. BJH volume 27 issue 4 cover and front matter. British Journal for the History of Science, 27(4):f1–f2, December 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:BRa

[Ano94i] Anonymous. Books received.

British Journal for the History of Science, 27(1):127– 128, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027596.

Anonymous:1994:BRb

[Ano94j] Anonymous. Books received.

British Journal for the History of Science, 27(2):245– 246, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027459.

Anonymous:1994:BRc

[Ano94k] Anonymous. Books received.

British Journal for the History of Science, 27(3):381–382, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027620.

Anonymous:1994:BRd

[Ano94l] Anonymous. Books received.

British Journal for the History of Science, 27(4):491–493, December 1994. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:BRe

[Ano94m] Anonymous. Books reviewed. British Journal for the History of Science, 27(4):495–497, December 1994. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:EN

[Ano94n] Anonymous. Editorial note. British Journal for the History of Science, 27(1):1, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1994:IA

[Ano94o] Anonymous. Index of authors. British Journal for the History of Science, 27(4):499, December 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BVIa

[Ano95a] Anonymous. BJH volume 28 issue 1 cover and back matter. British Journal for the History of Science, 28(1):b1– b9, March 1995. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BVIb

[Ano95b] Anonymous. BJH volume 28 issue 1 cover and front matter. British Journal for the History of Science, 28(1):f1–f3, March 1995. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BVIc

[Ano95c] Anonymous. BJH volume 28 issue 2 cover and back matter. British Journal for the History of Science, 28(2):b1– b6, June 1995. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BVId

[Ano95d] Anonymous. BJH volume 28 issue 2 cover and front matter. British Journal for the History of Science, 28(2):f1–f2, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BVIe

[Ano95e] Anonymous. BJH volume 28 issue 3 cover and back matter. British Journal for the History of Science, 28(3):b1– b10, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BVIf

[Ano95f] Anonymous. BJH volume 28 issue 3 cover and front matter. British Journal for the History of Science, 28(3):f1–f2, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BVIg

[Ano95g] Anonymous. BJH volume 28 issue 4 cover and back matter. British Journal for the History of Science, 28(4):b1–b6, December 1995. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BVIh

[Ano95h] Anonymous. BJH volume 28 issue 4 cover and front matter. British Journal for the History of Science, 28(4):f1–f2, December 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BRb

[Ano95i] Anonymous. Book received. British Journal for the History of Science, 28(2):251–253, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BRa

[Ano95j] Anonymous. Books received. British Journal for the History of Science, 28(1):127–129, March 1995. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1995:BRc

[Ano95k] Anonymous. Books received. British Journal for the History of Science, 28(3):383–384, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027673.

Anonymous:1996:BVIa

[Ano96a] Anonymous. BJH volume 29 issue 1 cover and back matter. British Journal for the History of Science, 29(1):b1– b3, March 1996. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:BVIb

[Ano96b] Anonymous. BJH volume 29 issue 1 cover and front matter. British Journal for the History of Science, 29(1):f1–f2, March 1996. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:BVIc

[Ano96c] Anonymous. BJH volume 29 issue 2 cover and back matter. British Journal for the History of Science, 29(2):b1– b6, June 1996. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:BVId

[Ano96d] Anonymous. BJH volume 29 issue 2 cover and front matter. British Journal for the History of Science, 29(2):f1–f2, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:BVIe

[Ano96e] Anonymous. BJH volume 29 issue 3 cover and back matter. British Journal for the History of Science, 29(3):b1– b2, September 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:BVIf

[Ano96f] Anonymous. BJH volume 29 issue 3 cover and front matter. British Journal for the History of Science, 29(3):f1–f2, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:BVIg

[Ano96g] Anonymous. BJH volume 29 issue 4 cover and back matter. British Journal for the History of Science, 29(4):b1–b6, December 1996. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:BVIh

[Ano96h] Anonymous. BJH volume 29 issue 4 cover and front matter. British Journal for the History of Science, 29(4):f1–f2, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:BRa

[Ano96i] Anonymous. Books received. British Journal for the History of Science, 29(1):125–127, March 1996. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:BRb

[Ano96j] Anonymous. Books received. British Journal for the History of Science, 29(2):255– 256, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027855.

Anonymous:1996:BRc

[Ano96k] Anonymous. Books received.

British Journal for the History of Science, 29(3):381–383, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027757.

Anonymous:1996:BRd

[Ano96l] Anonymous. Books received.

British Journal for the History of Science, 29(4):493–494, December 1996. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:BRe

[Ano96m] Anonymous. Books reviewed.

British Journal for the History of Science, 29(4):495–497, December 1996. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:C

[Ano96n]  Anonymous.  Correction.

British Journal for the History of Science, 29(3):380, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:IEA

[Ano96o]  Anonymous.  Income and

expenditure account for the year ended 31 December 1995. British Journal for the History of Science, 29(4):508–512, December 1996. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:IA

[Ano96p] Anonymous. Index of authors. British Journal for the History of Science, 29(4):499, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1996:RCY

[Ano96q] Anonymous. Report of Council for the year 1995. British Journal for the History of Science, 29(4):501–506, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027712.

Anonymous:1997:BRa

[Ano97a] Anonymous. Books received. British Journal for the History of Science, 30(1):123–125, March 1997. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1997:BRb

[Ano97b] Anonymous. Books received. British Journal for the History of Science, 30(2):251–253, June 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1997:BJH

[Ano97c] Anonymous. British Journal of History and Science. Index to volumes 1–30 (1962– 1997). British Journal for the History of Science, 30(4):387– 500, December 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027858.

Anonymous:1997:FMM

[Ano97d] Anonymous. Foundation members and members of the British Society for the History of Science on 1 June 1948. British Journal for the History of Science, 30(1):91–94, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1998:AC

[Ano98a] Anonymous. Archive collections. British Journal for the History of Science, 31(3):355– 360, September 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1998:BRa

[Ano98b] Anonymous. Books received.

British Journal for the History of Science, 31(1):103–106, March 1998. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1998:BRb

[Ano98c] Anonymous. Books received.

British Journal for the History of Science, 31(2):251– 252, June 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027772.

Anonymous:1998:BRc

[Ano98d] Anonymous. Books received. British Journal for the History of Science, 31(3):377–379, September 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027827.

Anonymous:1998:BRd

[Ano98e] Anonymous. Books received. British Journal for the History of Science, 31(4):489– 490, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027890.

Anonymous:1998:BRe

[Ano98f] Anonymous. Books reviewed. British Journal for the History of Science, 31(4):491– 493, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027893.

Anonymous:1998:BSH

[Ano98g] Anonymous. British Society for the History of Science. Report of Council for the year 1996. British Journal for the History of Science, 31(1):107– 123, March 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027811.

Anonymous:1999:BRa

[Ano99a] Anonymous. Books received.

British Journal for the History of Science, 32(1):125– 127, March 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027982.

Anonymous:1999:BRb

[Ano99b] Anonymous. Books received.

British Journal for the History of Science, 32(2):252–253, June 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1999:BRc

[Ano99c] Anonymous. Books received.

British Journal for the History of Science, 32(3):379–380, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4028005.

Anonymous:1999:BRd

[Ano99d] Anonymous. Books received.

British Journal for the History of Science, 32(4):489–490, December 1999. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:1999:BSH

[Ano99e] Anonymous. British Society for the History of Science Report of Council for the year 1998. British Journal for the History of Science, 32(4):495– 508, December 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027951.

Anonymous:1999:C

[Ano99f] Anonymous.

Correction. British Journal for the History of Science, 32 (2):251, June 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4028096.

Anonymous:2004:CIE

[Ano04] Anonymous. Correspondence: Images of the Eighteenth Century: Polemic: Style and aesthetics, volume 2004:07 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2004. ISBN 0-7294-0844-2. v + 293 pp. LCCN DC33.4 .C84 2004.

Anonymous:2005:CDH

[Ano05] Anonymous. Correspondence: Dialogue: History of ideas, volume 2005:07 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2005. ISBN 0-7294-0861-2. v + 342 pp. LCCN DC33.4 .C82 2005.

Anonymous:2007:ETW

[Ano07] Anonymous. Enlightenment and tradition: women’s studies: Montesquieu, volume 2007:06 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2007. ISBN 0-7294-0908-2 (paperback). xi + 316 pp. LCCN PQ261 .E55 2007.

Anstey:1996:BRM

[Ans96] Peter Anstey. Book review:

Michael Hunter (ed.), Robert Boyle: By Himself and His Friends. London: Pickering, 1994. Pp. cvii + 188. ISBN 1-85196-085-6. £49.95. British Journal for the History of Science, 29(1):95–96, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027518.

Appleyard:1993:UPS

[App93] Bryan Appleyard. Understanding the present: science and the soul of modern man. Picador. Pan Books, London, UK, 1993. ISBN 0-330-32013-0. xix + 283 pp. LCCN ????

Arato:1991:SDC

[Ara91] Franco Arato. Il secolo delle cose: scienza e storia in Francesco Algarotti. (Italian) [The century of things: science and history in Francisco Algarotti], volume 25 of Ricerche, studi, strumenti; Letteratura. Marietti, Genova, Italy, 1. edition, 1991. ISBN 88-211-8566-X, 88-211-9566-X. 173 pp. LCCN Q127.I8 A72 1991; PQ4683.A2 Z565 1991.

Ariew:1990:BRG

[Ari90] Roger Ariew. Book review:

Galileo Galilei. Sidereus Nuncius or the Sidereal Messenger, translated with introduction, conclusion, and notes by Albert Van Helden. Chicago and London: University of Chicago Press, 1989. Pp. xii + 127. ISBN 0-226-27903-0. £23.95, $34.50 (hardcover). £6.25, $9.25 (paperback). British Journal for the History of Science, 23(3): 355–356, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026777.

Armstrong:1983:PAB

[Arm83] David Armstrong. Political anatomy of the body: medical knowledge in Britain in the Twentieth Century. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24746-2. xii + 146 pp. LCCN R486 .A75 1983.

Alhazen:1989:OIA

[AS89] Alhazen and A. I. Sabra. The Optics of Ibn al-Haytham. Books I–III, On direct vision, volume 40 of Studies of the Warburg Institute. Warburg Institute, University of London, London, UK, 1989. ISBN 0-85481-072-2. ???? pp. LCCN QC353 .A3313 1989.

Ash:1996:FMS

[AS96] Mitchell G. Ash and Alfons Söllner, editors. Forced migration and scientific change: emigre German-speaking scientists and scholars after 1933. Publications of the German Historical Institute. German Historical Institute, Washington, DC, USA, 1996. ISBN 0-521-49741-8 (hardcover). xviii + 301 pp. LCCN E184.G3 F73

1996. URL http://www.loc.
gov/catdir/description/
cam027/95024894.html;
http://www.loc.gov/
catdir/samples/cam031/
95024894.html;   http://
www.loc.gov/catdir/toc/
cam026/95024894.html.

Ashford:1985:PPL

[Ash85] Oliver M. Ashford. Prophet — or professor?: the life and work of Lewis Fry Richardson. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-774-8. xiv + 304 pp. LCCN QC858.R53 A9 1985. US$17.00. URL http:

//www.loc.gov/catdir/
enhancements/fy0745/
84235055-d.html.

Ashworth:1994:CEB

[Ash94] William J. Ashworth. The calculating eye: Baily, Herschel, Babbage and the business of astronomy. British Journal for the History of Science, 27(4): 409–441, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027624.

Ash:1995:GPG

[Ash95] Mitchell G. Ash. Gestalt psychology in German culture, 1890–1967: holism and the quest for objectivity. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-47540-6 (hardcover). xii + 513 pp. LCCN BF203 .A84

                                        1995. URL http://www.loc.
                                        gov/catdir/description/
                                        cam026/94036273.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam026/
                                        94036273.html.

Aspray:1990:JNO

[Asp90] William Aspray. John von Neumann and the origins of modern computing. History of computing. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-01121-2. xvii + 376 pp. LCCN QA76.5 .A77446 1990.

Alvarez:1987:DAS

[AT87] Luis W. Alvarez and W. Peter Trower. Discovering Alvarez: selected works of Luis W. Alvarez, with commentary by his students and colleagues. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-81304-5. ix + 272 + 16 pp. LCCN QC71 .A462

                                        1987. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0608/87012766-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0608/87012766-t.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0609/87012766-b.html.

Atherton:1990:BRV

[Ath90] Margaret Atherton. Berkeley’s revolution in vision. Cornell University Press, Ithaca, NY, USA, 1990. ISBN 0-8014-2358-9. xii + 249 pp. LCCN B1339 .A72 1990.

Adams:1977:PJA

[ATKL08] John Adams, Robert Joseph Taylor, Mary-Jo Kline, and Gregg L. Lint. Papers of John Adams. The Adams papers: Series III, General correspondence and other papers of the Adams statesmen. Belknap Press of Harvard University Press, Cambridge, MA, USA and London, UK, 1977– 2008. ISBN 0-674-65441-2 (vol. 1–2), 0-674-65442-0 (vol. 3– 4), 0-674-65443-9 (vol. 5–6), 0-674-65444-7 (vol. 7–8), 0-674-65445-5 (vol. 9–10), 0-674-01136-8 (vol. 11), 0-674-01281-X (vol. 12), 0-674-01812-5 (vol. 13), (v., 14,, ), 0-674-02607-1 (vol. 14). ???? pp. LCCN E302 .A26 1977. URL http://www.

loc.gov/catdir/toc/fy052/
77004707.html.

Atran:1986:FHN

[Atr86] Scott Atran. Fondements de l’histoire naturelle: pour une anthropologie de la science. (French) [Origins of naturnal history: towards an anthropology of science]. Collection “Le Genre humain”. Editions Complexe, Bruxelles, Belgium, 1986. ISBN 2-87027-180-8. 244 pp. LCCN QH83 .A87 1986.

Ash:1987:PTC

[AW87]  Mitchell  G.

Ash and William Ray Woodward, editors. Psychology in Twentieth-Century thought and society. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-32523-4. ix + 320 pp. LCCN BF105 .P79 1987.

Amrine:1987:GSR

[AZW87] Frederick Amrine, Francis J. Zucker, and Harvey Wheeler, editors. Goethe and the Sciences: a Reappraisal, volume 97 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1987. ISBN 90-277-2265-X, 94-009-3761-X. ISSN 0068-0346. xvii + 442 pp. LCCN Q174 .B67 vol. 97. US$79.00. URL https://

                                        link.springer.com/book/
                                        10.1007/978-94-009-3761-1.

Barnes:1982:SSS

[B^{+}82] Jonathan Barnes et al., editors. Science and speculation: studies in Hellenistic theory and practice. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24689-X (Cambridge), 2-901725-47-3 (Éd. de la Maison des Sciences de l’Homme). xxvi + 351 pp. LCCN Q175 .S4175

                                        1982. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0632/82004221-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0632/82004221-t.html.

Bret:1995:DCA

[B^{+}95] Patrice Bret et al., editors.

Débats et chantiers actuels autour de Lavoisier et de la révolution chimique. (French) [Debates and works about Lavoisier and the chemical revolution], volume 68, 1–2 of Revue d’histoire des sciences. 48e année, Section d’histoire des sciences. Presses universitaires de France, Paris, France, 1995. ISBN 2-13-046972-8. 230 pp. LCCN ????

Barker:1991:RCE

[BA91] Peter Barker and Roger Ariew, editors. Revolution and continuity: essays in the history and philosophy of early modern science, volume 24 of Studies in philosophy and the history of philosophy. Catholic University of America Press, Washington, DC, USA, 1991. ISBN 0-8132-0738-X. v + 222 pp. LCCN B21 .S78 vol. 24.

Baatz:1990:KCS

[Baa90] Simon Baatz. Knowledge, culture, and science in the metropolis: the New York Academy of Sciences, 1817–1970, volume 584 of Annals of the New York Academy of Sciences, 0077-8923. New York Academy of Sciences, New York, NY, USA, 1990. ISBN 0-89766-545-7, 0-89766-546-5 (paperback). ix + 269 pp. LCCN Q11 .N5 vol.

Badash:1985:KRK

[Bad85] Lawrence Badash. Kapitza, Rutherford, and the Kremlin. Yale University Press, New Haven, CT, USA, 1985. ISBN 0-300-01465-1. xi + 129 pp. LCCN QC16.K25 B3 1985.

Baines:1991:BRM

[Bai91] John Baines. Book review: Marshall Clagett. Ancient Egyptian Science: A Source Book. Volume 1: Knowledge and Order (in two tomes). Philadelphia: American Philosophical Society, 1989. Pp. xv + 863. ISBN 0-87169-184-1. $60.00. British Journal for the History of Science, 24(2):243–245, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027169.

Baines:1993:BRM

[Bai93] John Baines. Book review: Mott T. Greene, Natural Knowledge in Preclassical Antiquity. Baltimore and London: Johns Hopkins University Press, 1992. Pp. xix + 182. ISBN 0-8018-4292-1. £18.00. British Journal for the History of Science, 26 (1):77–79, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027342.

Benguigui:1984:TED

[BaJANJ84] Isaac Benguigui, abbé Jean Antoine Nollet, and Jean Jallabert. Théories électriques du XVIIIème siècle. Correspondance entre l’abbé Nollet (1700–1770) et le physicien genevois Jean Jallabert (1712–1768). (French) [Electrical theories of the 18th Century. Correspondence between the Abbey Nollet (1700–1770) and the Genevan physician Jean Jallabert (1712–1768)]. Ed. Georg, Genève, Switzerland, 1984. ISBN 2-8257-0099-1. 234 + 1 pp. LCCN QC516 .B46 1984. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/
11454840.html.

Baker:1996:MNI

[Bak96] Jeffrey P. Baker. The machine in the nursery: incubator technology and the origins of newborn intensive care. The Johns Hopkins University Press, Baltimore, MD, USA, 1996. ISBN 0-8018-5173-4. x + 247 pp. LCCN RJ253.5 .B34

1996. URL http://www.loc.
gov/catdir/bios/jhu052/
95040320.html;
http://www.loc.gov/
catdir/description/
jhu051/95040320.html.

Bala:1991:IMB

[Bal91a] Poonam Bala. Imperialism and medicine in Bengal: a sociohistorical perspective. Sage Publications, New Delhi, India, 1991. ISBN 81-7036-245-8 (India), 0-8039-9100-2 (US). 174 pp. LCCN R607.B46 B35

                                        1991. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0657/91022537-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0657/91022537-t.html.

Baldwin:1991:BRR

[Bal91b] Martha Baldwin. Book review: Raffaella Castagnola (ed.) I Guicciardini e le scienze occulte: L’oroscopo di Francesco Guicciardini, lettere di alchimia, astrologia e cabala a Luigi Guicciardini. Istituto Nazionale di Studi sul Rinascimento, Studi e Testi XIX. Florence: Leo S. Olschki Editore, 1990. Pp. vii+ 397. ISBN 88-222-3761-7. L. 85000. British Journal for the History of Science, 24(3): 374–375, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027238.

Barraclough:1984:SBB

[Bar84] Kenneth Charles Barraclough, editor. Steelmaking before Bessemer, volume 279 of Book / Metals Society. The Metals Society, London, UK, 1984. ISBN 0-904357-53-8, 0-904357-64-3. ???? pp. LCCN ????

Barbour:1989:ARM

[Bar89a] Julian B. Barbour, editor. Absolute or relative motion?: a study from a Machian point of view of the discovery and the structure of dynamical theories. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32467-X. xv + 746 pp. LCCN QC133 .B37 1989.

Barbour:1989:DD

[Bar89b] Julian B. Barbour. The discovery of dynamics, volume 1 of Absolute or relative motion? Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32467-X. xv + 746 pp. LCCN ????

Barry:1990:BRR

[Bar90a] Jonathan Barry. Book review:

Roy Porter. Health for Sale: Quackery in England 1660– 1850. Manchester and New York: Manchester University Press, 1989. Pp. xii + 280. ISBN 0-7190-1903-6. £19.95. British Journal for the History of Science, 23(3):356–357, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026778.

Barton:1990:ISC

[Bar90b] Ruth Barton. ‘an influential set of chaps: The X-Club and Royal Society politics 1864–1885. British Journal for the History of Science, 23(1):53–81, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026802.

Bartholin:1991:EBI

[Bar91] Erasmus Bartholin. Experiments on birefringent Icelandic crystal, volume 40 of Acta historica scientiarum naturalium et medicinalium, 0065-1311. Danish National Library of Science and Medicine, Copenhagen, Denmark, 1991. ISBN 87-7709-010-1. 62, 60 pp. LCCN IN PROCESS.

Barnes:1992:BRG

[Bar92] Jonathan Barnes. Book review: G. E. R. Lloyd. Methods and Problems in Greek Science: Selected Papers. Cambridge: Cambridge University Press, 1991. Pp. xiv + 457. ISBN 0-521-37419-7. £45.00, $59.50. British Journal for the History of Science, 25(3):357– 358, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027259.

Barry:1993:HME

[Bar93a] Andrew Barry. The history of measurement and the engineers of space. British Journal for the History of Science, 26(4): 459–468, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027466.

Bartley:1993:BRH

[Bar93b] Mary M. Bartley. Book review: Helena Cronin. The Ant and the Peacock: Altruism and Sexual Selection from Darwin to Today. Cambridge: Cambridge University Press, 1991. Pp. xiv + 490. ISBN 0-521-32937-X. £27.50, $39.50. British Journal for the History of Science, 26 (1):111–112, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027375.

Barton:1994:AA

[Bar94] Tamsyn Barton. Ancient astrology. Sciences of antiquity. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1994. ISBN 0-415-08066-5, 0-415-11029-7 (paperback). xxv + 245 pp. LCCN BF1671 .B378

1994. URL http://www.loc.
gov/catdir/enhancements/
fy0648/93032441-d.html.

Basalla:1988:ET

[Bas88] George Basalla. The evolution of technology. Cambridge history of science. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-22855-7, 0-521-29681-1 (paperback). viii + 248 pp. LCCN T15 BAS 1988; 509; T; Hist; 609; 125; T15.BAS.

Baumler:1984:PES

[Bäu84] Ernst Bäumler. Paul Ehrlich, scientist for life. Holmes and Meier, New York, NY, USA, 1984. ISBN 0-8419-0837-0. xvi + 288 + 32 pp. LCCN R512.E4 B3313 1984.

Berkeley:1982:LTJ

[BB82]  Edmund  Berke-

ley and Dorothy Smith Berkeley. The life and travels of John Bartram from Lake Ontario to the River St. John. University Presses of Florida, Tallahassee, FL, USA, 1982. ISBN 0-8130-0700-3. xv + 376 pp. LCCN QK31.B3 B47.

Blatt:1984:CCA

[BB84] Sidney J. (Sidney Jules) Blatt and Ethel S. Blatt. Continuity and change in art: the development of modes of representation. L. Erlbaum, Hillsdale, NJ, USA, 1984. ISBN 0-89859-342-5. xix + 411 pp. LCCN N7430.5 .B55

                                        1984. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0742/84004201-d.html.

Barker-Benfield:1992:CSS

[BB92] G. J. Barker-Benfield. The culture of sensibility: sex and society in Eighteenth-Century Britain. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-03713-4. xxxiv + 520 pp. LCCN HQ1593 .B37

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi052/91047945.html;
                                        http://www.loc.gov/
                                        catdir/toc/uchi051/
                                        91047945.html.

Bertotti:1990:MCR

[BBBM90] B. Bertotti, R. Balbinot, S. Bergia, and A. Messina, editors. Modern Cosmology in Retrospect. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37213-5. xx + 426 pp. LCCN QB981 .M774

1990. URL http://www.loc.
gov/catdir/description/
cam024/90041803.html;
http://www.loc.gov/
catdir/toc/cam028/
90041803.html.

Bloch:1988:EFH

[BBC88] Olivier Bloch, Bernard Balan, and Paulette Carrive, editors. Entre forme et histoire: la formation de la notion de développement à l’âge classique. (French) [Between form and history: the formation of the notion of development in the classical age]. Philosophie. Meridiens Klincksieck, Paris, France, 1988. ISBN 2-86563-212-1. 310 pp. LCCN B818 .E58 1988.

Bessel:1989:STE

[BC89] Richard Bessel and Colin Chant, editors. Science, technology and everyday life, 1870– 1950. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1989. ISBN 0-415-00037-8, 0-415-03557-0. 391 pp. LCCN ????

Ben-Chaim:1990:SMS

[BC90] Michael Ben-Chaim. Social mobility and scientific change: Stephen Gray’s contribution to electrical research. British Journal for the History of Science, 23(1):3–24, March 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026800.

Blake-Coleman:1992:CWE

[BC92] B. C. (Barrie Charles) Blake-Coleman. Copper wire and electrical conductors: the shaping of a technology. Harwood Academic Publishers, Chur, Switzerland and New York, NY, USA, 1992. ISBN 3-7186-5200-5. xx + 284 pp. LCCN TS270 .B58

                                        1991. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0653/91040910-d.html.

Brooke:1998:RNE

[BC98] John Hedley Brooke and Geoffrey N. Cantor. Reconstructing nature: the engagement of science and religion: Glasgow Gifford lectures. Glasgow Gifford Lectures. T and T Clark, Edinburgh, Scotland, 1998. ISBN 0-567-08600-3. xii-367 pp. LCCN ????

Babbage:1989:WCB

[BCKR89] Charles Babbage, Martin Campbell-Kelly, and Peter M. Roget, editors. The works of Charles Babbage. The Pickering masters. W. Pickering, London, UK, 1989. ISBN 1-85196-005-8 (set), 1-85196-501-7 (vol. 1), 1-85196-502-5 (vol. 2), 1-85196-503-3 (vol. 3), 1-85196-504-1 (vol. 4), 1-85196-505-X (vol. 5), 1-85196-506-8 (vol. 6), 1-85196-507-6 (vol. 7), 1-85196-508-4 (vol. 8), 1-85196-509-2 (vol. 9), 1-85196-510-6 (vol. 10), 1-85196-511-4 (vol. 11). LCCN QA37.2. 11 volumes.

Bravo-Dobree:1990:BRA

[BD90] Michael Bravo-Dobrée. Book review: Andrew David, Rüdiger Joppien & Bernard Smith (eds). The Charts and Coastal Views of Captain Cook’s Voyages: The Voyage of the Endeavour (1768–1771). London: The Hak-luyt Society, 1988. Pp. Ixiv + 328. ISBN 0-9041-8023-9. £100. British Journal for the History of Science, 23(3): 341–343, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026766.

Blondel:1995:RCU

[BD95] Christine Blondel and Matthias Dörries, editors. Restaging Coulomb: usages, controverses et réplications autour de la balance de torsion. (French) [Restaging Coulomb: Uses, controversies, and replications of the torsion balance], volume 15 of Biblioteca di Nuncius. Studi e testi. Leo S. Olschki, Firenze, Italy, 1995. ISBN 88-222-4196-7. 167 pp. LCCN ????

Boyle:1996:FEV

[BDH96] Robert Boyle, Edward Bradford Davis, and Michael Cyril William Hunter. A free enquiry into the vulgarly received notion of nature. Cambridge texts in the history of philosophy. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56100-0 (hardcover), 0-521-56796-3 (paperback). xxxvi + 171 pp. LCCN Q155 .B68

                                        1996. URL http://www.loc.
                                        gov/catdir/description/
                                        cam027/96010997.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam024/
                                        96010997.html.

Beaujouan:1991:PRN

[Bea91] Guy Beaujouan. Par raison de nombres: l’art du calcul et les savoirs scientifiques médiévaux. (French) [Because of numbers: the art of calculation and medieval scientific knowledge], volume CS344 of Variorum collected studies series. Variorum, Aldershot, Hampshire, UK, 1991. ISBN 0-86078-281-6. ???? pp. LCCN Q124.97 .B43 1991.

Beauchamp:1997:EE

[Bea97] K. G. Beauchamp. Exhibiting electricity, volume 21 of IEE history of technology series. Institution of Electrical Engineers, London, UK, 1997. ISBN 0-85296-895-7. xii + 337 pp. LCCN TK6.A1 B43 1997.

Bechler:1982:CNR

[Bec82] Z. Bechler, editor. Contemporary Newtonian research, volume 9 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1982. ISBN 90-277-1303-0. 241 pp. LCCN QC16.N7 C66 1982.

Bechler:1991:NPC

[Bec91a] Zev Bechler. Newton’s Physics and the Conceptual Structure of the Scientific Revolution, volume 127 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1054-3, 94-011-3276-3. ISSN 0068-0346. xviii + 588 pp. LCCN Q174 .B67

vol. 127. URL https://link.
springer.com/book/10.
1007/978-94-011-3276-3.

Bechmann:1991:VHP

[Bec91b] Roland Bechmann. Villard de Honnecourt: la pensée technique au XIIIe siècle et sa communication. (French) [Villard de Honnecourt: technical thought in the 18th Century and its communication]. Picard, Paris, France, 1991. ISBN 2-7084-0367-2. 384 pp. LCCN NA1053.V6 B4 1991. 420F.

Becquemont:1992:DDE

[Bec92]  Daniel  Bec-

quemont. Darwin, darwinisme, évolutionnisme. (French) [Darwin, Darwinism, and evolutionism]. Histoire des idées, théorie politique et recherches en sciences sociales. Kimé, Paris, France, 1992. ISBN 2-908212-13-7. 341 pp. LCCN QH365.O8 B435 1992. 200F.

Becher:1995:RWC

[Bec95] Harvey W. Becher. Radicals, Whigs and conservatives: the middle and lower classes in the analytical revolution at Cambridge in the age of aristocracy. British Journal for the History of Science, 28(4):405– 426, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027545.

Beccaloni:1999:RWM

[Bec99] George Beccaloni. A. R. Wallace Memorial Fund. British Journal for the History of Science, 32(4):483–484, December 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Beeson:1989:CO

[Bee89] C. F. C. Beeson. Clockmaking in Oxfordshire, 1400–1850. Museum of the History of Science, Oxford, UK, third edition, 1989. ISBN 0-903364-06-9. 212 pp. LCCN ???? With a new introduction and index by A. V. Simcock.

Beer:1996:OFS

[Bee96] Gillian Beer. Open fields: science in cultural encounter. Clarendon Press, New York, NY, USA, 1996. ISBN 0-19-818369-0. viii + 341 pp. LCCN PR468.S34 B44

                                        1996. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0639/95043234-d.html.

Baer:2010:EDL

[BEFH10] Werner Baer, Hadi Salehi Esfahani, Giovanni Facchini, and Geoffrey Hewings, editors. Economic development in Latin America: essay in honor of Werner Baer. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2010. ISBN 0-230-22381-8. xxii + 309 pp. LCCN HC125 .E369 2010.

Bennett:1982:MSC

[Ben82] J. A. (James A.) Bennett.

The mathematical science of Christopher Wren. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24608-3. 148 pp. LCCN Q143.W68 B46

1982. URL http://www.loc.
gov/catdir/description/
cam022/81018045.html;
http://www.loc.gov/
catdir/samples/cam031/
81018045.html;   http://
www.loc.gov/catdir/toc/
cam031/81018045.html.

Beniger:1986:CRT

[Ben86] James R. (James Ralph) Beniger. The control revolution: technological and economic origins of the information society. Harvard University Press, Cambridge, MA, USA, 1986. ISBN 0-674-16985-9. x + 493 pp. LCCN HM258 .B459 1986.

Bennett:1987:DCH

[Ben87] J. A. (James A.) Bennett.

The divided circle: a history of instruments for astronomy, navigation, and surveying. Phaidon, Christie’s, Oxford, UK, 1987. ISBN 0-7148-8038-8. 224 pp. LCCN QB85.8 .B46 1987.

Bennett:1990:BRG

[Ben90a] J. A. Bennett. Book review: G. J. Whitrow. Time in History: The Evolution of our General Awareness of Time and Temporal Perspective. Oxford: Oxford University Press, 1988. Pp. 217. ISBN 0-19-215361-7. £14.95. British Journal for the History of Science, 23 (2):252, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026751.

Benninga:1990:HLA

[Ben90b] H. (Harm) Benninga. A history of lactic acid making: a chapter in the history of biotechnology. Chemists and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0625-2 (US). xxi + 478 pp. LCCN TP248.L3 B46 1990.

Benjamin:1991:SSG

[Ben91] Marina Benjamin, editor. Science and sensibility: gender and scientific enquiry, 1780– 1945. Blackwell, Oxford, UK, 1991. ISBN 0-631-16649-1. x + 295 pp. LCCN Q130 .S327 1991.

Bennett:1993:BRS

[Ben93]  J. A. Bennett.  Book re-

view: Sabino Maffeo, SJ, In the Service of Nine Popes: 100 Years of the Vatican Observatory, translated by George V. Coyne, SJ. Rome: Specola Vaticana/Pontificia Academia Scientiarum, 1991. Pp. xvi + 241. ISBN 88-7761-046-8. $24.95. British Journal for the History of Science, 26(1): 125–126, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027387.

Bennett:1994:RCY

[Ben94a] G. Bennett. Report of Council for the year 1993. British Journal for the History of Science, 27(4):501–507, December 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027643.

Bennett:1994:BRS

[Ben94b] Jim Bennett. Book reviews:

Stella Butler, Science and Technology Museums. Leicester, London and New York: Leicester University Press, 1992. Pp. xiii + 149. ISBN 0-7185-1357-6. £35.00. Janet Carding, Timothy Boon, Nicholas Wyatt and Robert Bud, Guide to the History of Technology in Europe. London: Science Museum, 1992. Pp. 142. ISBN 0-901805-51-3. £8.00. British Journal for the History of Science, 27(1): 125–126, March 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027594.

Bennett:1997:MEH

[Ben97] J. A. Bennett. Museums and the establishment of the history of science at Oxford and Cambridge. British Journal for the History of Science, 30 (1):29–46, March 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027898.

Busard:1992:RCR

[BEoCF92] H.  L.  L.

(Hubertus Lambertus Ludovicus) Busard, Euklides, Robert of Chester, and Menso Folkerts. Robert of Chester’s (?) redaction of Euclid’s Elements, the so-called Adelard II version, volume 9 of Science networks: historical studies, 99-0922970-6. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 3-7643-2727-8, 3-7643-2728-6 (set). 442–959. pp. LCCN ????

Berg:1985:AMI

[Ber85] Maxine Berg. The age of manufactures: industry, innovation and work in Britain, 1700– 1820. Fontana history of England. Fontana Press, London, UK, 1985. ISBN 0-00-686019-2. 378 pp. LCCN HD9720.5 .B47 1985.

Beretta:1993:EMD

[Ber93a] Marco Beretta. The enlightenment of matter: the definition of chemistry from Agricola to Lavoisier, volume 15 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1993. ISBN 0-88135-152-0. xvi + 396 pp. LCCN QD7 .B52 1993.

BertoloniMeli:1993:EPN

[Ber93b] Domenico Bertoloni Meli. Equivalence and priority: Newton versus Leibniz; including Leibniz’s unpublished manuscripts on the Principia. Oxford science publications. Clarendon Press, New York, NY, USA, 1993. ISBN 0-19-853945-2 (hardcover). ix + 318 pp. LCCN QA803 .B47 1993. US$35.00 (US$52.50

U.S.). URL http://www.loc.
gov/catdir/enhancements/
fy0639/92018967-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0639/92018967-t.html.

Beretta:1995:SMP

[Ber95] Marco Beretta, editor. The starry messenger and the polar star: scientific relations between Italy and Sweden from 1500 to 1800: catalogue of an exhibition held at the Naturhistoriska riksmuseet — Stockholm, volume 22 of Uppsala studies in history of science. Giunti, Prato, Italy, 1995.

ISBN 88-09-20793-9. ISSN 0282-1036. 188 pp. LCCN Q127.I8 B466 1995.

Bevan:1990:BRC

[Bev90]  Michael Bevan.  Book re-

view: Cynthia Eagle Russett. Sexual Science: The Victorian Construction of Womanhood. Cambridge, Mass, and London: Harvard University Press, 1989. Pp. 245. ISBN 0-674-80290-X. £15.95, $25.00. British Journal for the History of Science, 23(3):358–359, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026780.

Berthollet:1980:RES

[BG80] Claude-Louis Berthollet and Michelle Goupil. Revue de l’“Essai de statique chimique”. École polytechnique, Palaiseau, France 91128 cedex, 1980. ISBN 2-7302-0019-3. 204 pp. LCCN ????

Beaune:1992:BAD

[BG^{+}92] Jean-Claude Beaune, Jean Gayon, et al., editors. Buffon 88: actes du Colloque international pour le bicentenaire de la mort de Buffon: (Paris, Montbard, Dijon, 14–22 juin 1988). (French) [Buffon 88: acts of the international conference for the bicentenary of the death of Buffon: (Paris, Montbard, Dijon, 14-22 June 1988)]. Science, histoire, philosophie. J. Vrin, Paris, France, 1992. ISBN 2-7116-9755-X. ix + 771 pp. LCCN QH31.B88 B84 1988.

Barrett:1987:CDN

[BGH^{+}87] Paul  H.  Bar-

rett, Peter Jack Gautrey, Sandra Herbert, David Kohn, and Sydney Smith, editors. Charles Darwin’s notebooks, 1836–1844: geology, transmutation of species, metaphysical enquiries. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-35055-7 (hardcover), 0-521-09975-7 (paperback). viii + 747 pp. LCCN QH31.D2; QH365.Z9 B37 1987. US$65.00.

Beck:1987:CPAa

[BH87] Anna Beck and Peter Havas.

The collected papers of Albert Einstein. Vol. 1. The Early years, 1879–1902. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-08475-0 (paperback), 0-691-08463-7 (microfiche). xxii + 196 pp. LCCN QC16.E5 A2 1987a. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/
catdir/description/
prin031/87160800.html;
http://www.loc.gov/
catdir/toc/prin031/
87160800.html.

Bitbol-Hesperies:1990:PVC

[BH90] Annie Bitbol-Hespériès. Le principe de vie chez Descartes. (French) [Descartes principle of life]. Bibliothèque d’histoire de la philosophie. Nouvelle série. J. Vrin, Paris, France, 1990. ISBN 2-7116-1034-9. 236 pp. LCCN B1875 .B48 1990. 198F.

Boyle:1994:RB

[BHW94]  Robert  Boyle,

Michael Cyril William Hunter, and William Wotton. Robert Boyle. W. Pickering, London, UK, 1994. ISBN 1-85196-085-6. cvii + 188 pp. LCCN Q143.B77 A3 1994.

Biagioli:1993:GCP

[Bia93] Mario Biagioli. Galileo, courtier: the practice of science in the culture of absolutism. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1993. ISBN 0-226-04559-5 (hardcover). xii + 402 pp. LCCN QB36.G2 B54

                                        1993. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi051/92033736.html;
                                        http://www.loc.gov/
                                        catdir/toc/uchi051/
                                        92033736.html.

Bird:1986:WJ

[Bir86] Graham Bird. William James. The Arguments of the philosophers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1986. ISBN 0-7100-9602-X. vi + 221 pp. LCCN B945.J24 B39 1986. US$18.95 (US$45.00 U.S. : est.).

Bishop:1998:BRS

[Bis98] Ian Bishop. Book review:

Steven Noll, Feeble-Minded in our Midst: Institutions for the Mentally Retarded in the South, 1900–1940. Chapel Hill: University of North Carolina Press, 1995. Pp. xiii + 254. ISBN 0-8078-2220-5, $39.95 (hardcover); 0-8078-4531-0, $16.95 (paperback). British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027778.

Busard:1991:JNB

[BJ91] H. L. L. (Hubertus Lambertus Ludovicus) Busard and Nemorarius Jordanus. Jordanus de Nemore, De elementis arithmetice artis: a medieval treatise on number theory, volume 22, 1–2 of Boethius. F. Steiner Verlag, Stuttgart, Germany, 1991. ISBN 3-515-05214-3. ???? pp. LCCN QA241 .B835 1991.

Berkeley:1992:BMB

[BJB92] George Berkeley, Douglas M.

Jesseph, and George Berkeley. De motu and, The analyst, volume 41 of The New synthese historical library. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1520-0. x + 230 pp. LCCN Q175 .B4674 1992.

Boyle:1990:LPR

[BKH90] Robert Boyle, Paul Kesaris, and Michael Cyril William Hunter. Letters and papers of Robert Boyle: from the archives of the Royal Society. Collections from the Royal Society. University Publications of America, Frederick, MD, USA, 1990. ISBN 1-55655-170-3 (microfilm), 1-55655-217-3 (guide). LCCN Q143.B77; Microfilm 95/4348 (Q). US$1,395.00, US$60.00. 16 microfilm reels.

Bacon:1983:RBP

[BL83] Roger Bacon and David C. Lindberg. Roger Bacon’s philosophy of nature: a critical edition, with English translation, introduction, and notes, of De multiplicatione specierum and De speculis comburentibus. Clarendon Press, New York, NY, USA, 1983. ISBN 0-19-858164-5. lxxxi + 420 pp. LCCN QC353 .B313 1983. US$74.00.

Blay:1983:CNP

[Bla83] Michel Blay. La conceptualisation newtonienne des phénomènes de la couleur. (French) [The Newtonian conceptualization of the phenomenon of color]. L’Histoire des sciences. Textes et études. J. Vrin, Paris, France, 1983. ISBN 2-7116-0825-5. 304 pp. LCCN QC494.7 .B57 1983. 99.00F.

Blay:1993:RID

[Bla93]  Michel Blay.  Les

raisons de l’infini: du monde clos à l’univers mathématique. (French) [The reasons for infinity: from a closed world to a mathematical universe]. nrf essais. Gallimard, Paris, France, 1993. ISBN 2-07-072992-3. 258 pp. LCCN BD411 .B55

1993. URL http://catalog.
hathitrust.org/api/
volumes/oclc/29673358.
html.

Bliss:1982:DI

[Bli82] Michael Bliss. The discovery of insulin. Paul Harris publishing, Edinburgh, Scotland, 1982. ISBN 0-86228-056-7. 304 + 16 pp. LCCN ????

Blitz:1992:EEQ

[Bli92] David Blitz. Emergent evolution: qualitative novelty and the levels of reality, volume 19 of Episteme. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1658-4. ix + 239 pp. LCCN B818 .B65

1992. URL http://www.loc.
gov/catdir/enhancements/
fy0823/92004950-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0823/92004950-t.html.

Blondel:1982:MAC

[Blo82]  Christine  Blondel.

A-M. Ampère et la création de l’électrodynamique, 1820–1827, volume 10 of Mémoires de la Section des sciences, 0553-0734. Bibliothèque nationale, Paris, France, 1982. ISBN 2-7177-1643-2. 202 pp. LCCN Q46 .F73216 vol. 10.

Bloor:1993:BRJ

[Blo93a] David Bloor. Book review: Jan Sapp, Where the Truth Lies: Franz Moewus and the Origins of Molecular Biology. Cambridge: Cambridge University Press, 1990. Pp. ix + 340. ISBN 0-521-36550-3, £30, $59.50 (hardcover); 0-521-36751-4, £15, $18.95 (paperback). British Journal for the History of Science, 26(2): 260–263, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027122.

Blowers:1993:BRE

[Blo93b] Geoffrey Blowers. Book review: Edward Timms and Ritchie Robertson (eds.), Psychoanalysis and its Cultural Context. Austrian Studies III. Edinburgh: Edinburgh University Press, 1992. Pp. xi + 209. ISBN 0-7486-0359-X. £30.00. British Journal for the History of Science, 26(2):251–252, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027115.

Blowers:1993:BRP

[Blo93c] Geoffrey Blowers. Book review: Patricia Kitcher, Freud’s Dream: A Complete Interdisciplinary Science of Mind. Cambridge, Mass, and London: MIT Press, 1993. Pp. ix + 245. ISBN 0-262-11172-1. £22.50. British Journal for the History of Science, 26(3): 373–374, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027423.

Bloor:1994:MFH

[Blo94a] David Bloor. The moving forces of history. British Journal for the History of Science, 27(3): 351–355, September 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027603.

Blowers:1994:BRA

[Blo94b] Geoffrey Blowers. Book review: Alan Gauld, A History of Hypnotism. Cambridge: Cambridge University Press, 1992. Pp. xvii + 738. ISBN 0-521-30675-2. £75.00. British Journal for the History of Science, 27(2):234–235, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027448.

Bloor:1998:BRR

[Blo98] David Bloor. Book review:

Roger Smith, The Fontana History of the Human Sciences. London: Fontana Press, 1997. Pp. xviii + 1036. ISBN 0-00-686178-4. £12.99 (paperback edition). British Journal for the History of Science, 31(3): 361–375, September 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027819.

Blume:1990:BRD

[Blu90] Stuart Blume. Book review: Daniel M. Fox & Christopher Lawrence. Photographing Medicine: Images and Power in Britain and America since 1840. New York, Westport, London: Greenwood Press, 1988. Pp. 357. ISBN 0-313-23719-0. £36.50. British Journal for the History of Science, 23(3): 364–365, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026785.

Barthelemy-Madaule:1982:LMP

[BM82]  Madeleine  Barthélemy-

Madaule. Lamarck, the mythical precursor: a study of the relations between science and ideology. MIT Press, Cambridge, MA, USA, 1982. ISBN 0-262-02179-X. xv + 174 pp. LCCN QH31.L2 B3613 1982.

Brock:1984:LLT

[BM84] W. H. (William Hodson) Brock and A. J. (Arthur Jack) Meadows. The lamp of learning: Taylor and Francis and the development of science publishing. Taylor and Francis, Boca Raton, FL, USA, 1984. ISBN 0-85066-253-2. xv + 240 pp. LCCN Z286.S4 B76 1984.

Bradwardine:1989:TBB

[BM89]  Thomas  Bradwardine

and George Molland. Thomas Bradwardine, Geometria speculativa: Latin text and English translation with an introduction and a commentary, volume 18 of Boethius. F. Steiner Verlag, Stuttgart, Germany, 1989. ISBN 3-515-05020-5. 176 pp. LCCN QA444 .B7313 1989.

Biarnais:1982:BNG

[BN82] Marie Françoise Biarnais and Sir Isaac Newton. Les “Principia” de Newton: genèse et structure des chapitres fondamentaux avec traduction nouvelle, volume N.S. 2 of Cahiers d’histoire et de philosophie des sciences. Centre de documentation sciences humaines, Paris, France, 1982. ISBN 2-222-03094-3. 287 pp. LCCN ????

Brown:1995:AWA

[BN95] Arthur I. Brown and Ronald L.

Numbers. The antievolution works of Arthur I. Brown, volume 3 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1804-9. xiii + 209 pp. LCCN BS651 .B7874 1995.

Bohlin:1993:BRR

[Boh93] Ingemar Bohlin. Book review:

Robert J. Richards, The Meaning of Evolution: The Morphological Construction and Ideological Reconstruction of Darwin’s Theory. Chicago: University of Chicago Press, 1992. Pp. xv + 205. ISBN 0-226-71202-8. £15.95, $19.95. British Journal for the History of Science, 26 (1):106–108, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027372.

Bohlin:1995:BRP

[Boh95] Ingemar Bohlin. Book review: Peter J. Bowler, Darwinism. Twayne’s Studies in Intellectual History. New York: Twayne, 1993. Pp. xii + 118. ISBN 0-8057-8613-9, $24.95 (hardcover); 0-8057-8638-4, $14.95 (paperback). British Journal for the History of Science, 28 (1):117–119, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027499.

Bostock:1992:BRR

[Bos92] David Bostock. Book review: Robert Wardy. The Chain of Change: A Study of Aristotle’s Physics VII. Cambridge: Cambridge University Press, 1990. Pp. x + 345. ISBN 0-521-37327-1. £35.00, $59.50. British Journal for the History of Science, 25(2):259–263, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027305.

Bowler:1983:EDA

[Bow83] Peter J. Bowler. The eclipse of Darwinism: anti-Darwinian evolution theories in the decades around 1900. The Johns Hopkins University Press, Baltimore, MD, USA, 1983. ISBN 0-8018-2932-1. xi + 291 pp. LCCN QH361 .B68 1983.

Bowler:1988:NDR

[Bow88] Peter J. Bowler. The non-Darwinian revolution: reinterpreting a historical myth. The Johns Hopkins University Press, Baltimore, MD, USA, 1988. ISBN 0-8018-3678-6. x + 238 pp. LCCN QH361 .B693

1988. URL http://www.loc.
gov/catdir/description/
jhu051/88009738.html.

Bowler:1989:IPV

[Bow89] Peter J. Bowler. The invention of progress: the Victorians and the past. Blackwell, Oxford, UK, 1989. ISBN 0-631-16107-4. xi + 231 pp. LCCN GN360 .B67 1989. US$39.00 (U.S. : est.).

Bowler:1990:CDM

[Bow90] Peter J. Bowler. Charles Darwin: the man and his influence. Blackwell scientific biographies. Blackwell, Oxford, UK, 1990. ISBN 0-631-16818-4. xii + 250 pp. LCCN QH31.D2 B743 1991. US$34.00.

Bowler:1991:BRB

[Bow91a] Peter J. Bowler. Book review: Bruce G. Trigger. A History of Archaeological Thought. Cambridge: Cambridge University Press, 1990. Pp. xv + 500. ISBN 0-521-33818-2. £14.95 (paperback). British Journal for the History of Science, 24 (1):120, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027035.

Bowler:1991:BRN

[Bow91b] Peter J. Bowler. Book review: Nicolaas A. Rupke (ed.). Science, Politics and the Public Good: Essays in Honour of Margaret Gowing. London: Macmillan, 1988. Pp. xiv + 260. ISBN 0-333-44159-1. £33.00. British Journal for the History of Science, 24(4): 493–494, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027152.

Bowker:1992:BRW

[Bow92a]  Geof Bowker.  Book re-

view: William Aspray. John von Neumann and the Origins of Modern Computing. Cambridge, Mass, and London: MIT Press, 1991. Pp. xvii + 376. ISBN 0-262-01121-2. £31.50. British Journal for the History of Science, 25(3): 386–387, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027288.

Bowler:1992:BRJ

[Bow92b] Peter J. Bowler. Book review: Jane Maienschein. Transforming Traditions in American Biology, 1880–1915. Baltimore and London: Johns Hopkins University Press, 1991. Pp. xi + 366. ISBN 0-8018-4126-7. £34.50. British Journal for the History of Science, 25(4): 487–488, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027075.

Bowler:1992:BRC

[Bow92c] Peter J. Bowler. Book reviews:

Charles Lyell. Principles of Geology, first edition, vol. I, with a new Introduction by Martin J. S. Rudwick. Chicago and London: University of Chicago Press, 1990. Pp. [lviii] + xv + 511. ISBN 0-226-49794-1. £14.25, $20.75 (paperback, also available in hardback). Vol. II. Chicago and London: University of Chicago Press, 1991. Pp. xii + 330. ISBN 0-226-49797-6. £12.75, $18.25. Vol. III, with a new Bibliography compiled by Martin J. S. Rudwick. Chicago and London: University of Chicago Press, 1991. Pp. xxvi + 398 + 160. ISBN 0-226-49798-4. £31.95 (hardcover); 0-226-49799-2. £14.25 (paperback). British Journal for the History of Science, 25 (3):369, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027268.

Bowler:1992:FHE

[Bow92d] Peter J. (Peter John) Bowler. The Fontana history of the environmental sciences. Fontana history of science. Fontana Press, London, UK, 1992. ISBN 0-00-686184-9. xv + 634 pp. LCCN GE50 .B68 1992.

Bowler:1993:BRH

[Bow93a] Peter J. Bowler. Book review: Henrika Kuklick, The Social History of British Anthropology, 1885–1945. Cambridge: Cambridge University Press, 1992. Pp. ix + 325. ISBN 0-521-41109-2. £30.00, $44.95. British Journal for the History of Science, 26(1): 100–101, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027366.

Bowler:1993:BRP

[Bow93b] Peter J. Bowler. Book review: Patrick Tort (ed.), Darwinisme et société. Paris: Presses Universitaires de France, 1992. Pp.

690. ISBN 2-13-044774-0. FF.

480. British Journal for the History of Science, 26(4):494– 495, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027474.

Bowler:1993:BRR

[Bow93c] Peter J. Bowler. Book review:

Ronald Rainger, An Agenda for Antiquity: Henry Fairfield Osborn and Vertebrate Paleontology at the American Museum of Natural History, 1890– 1935. Tuscaloosa: University of Alabama Press, 1991. Pp. xiii + 360. ISBN 0-8173-0536-X. $37.95. British Journal for the History of Science, 26 (1):116, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027378.

Bowler:1993:BRD

[Bow93d] Peter J. Bowler. Book reviews:

Daniel Becquemont, Darwin, Darwinisme, Evolutionnisme. Paris: Editions Kimé, 1992. Pp. 341. ISBN 2-908212-13-7. 200FF. Jean Gayon, Darwin et l’après Darwin: Une histoire de l’hypothèse de sélection naturelle. Paris: Editions Kimé, 1992. Pp. 453. ISBN 2-908212-14-5. 250FF. British Journal for the History of Science, 26 (1):103–104, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027369.

Bowler:1993:D

[Bow93e] Peter J. Bowler. Darwinism, volume 6 of Twayne’s studies in intellectual and cultural history. Twayne Publishers, New York, NY, USA, 1993. ISBN 0-8057-8613-9, 0-8057-8638-4 (paperback). xii + 118 pp. LCCN QH366.2 .B675 1993. US$24.95, US$14.95.

Bowker:1994:BRR

[Bow94a] Geof Bowker. Book review: E. Roy Weintraub (ed.), Toward a History of Game Theory, Annual Supplement to History of Political Economy, vol.

24. Durham and London: Duke University Press, 1992. Pp. vi + 306. ISBN 0-8223-1253-0. £29.95. British Journal for the History of Science, 27(2): 239–240, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027453.

Bowler:1994:BRS

[Bow94b] Peter J. Bowler. Book review: Sahotra Sarkar (ed.), The Founders of Evolutionary Genetics: A Centenary Reappraisal. Boston Studies in the Philosophy of Science, 142. Dordrecht: Kluwer Academic Publishers, 1992. Pp. 300. ISBN 0-7923-1777-7. £62.00, $115.00, Dfl. 180.00. British Journal for the History of Science, 27(1):122, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027591.

Bowler:1995:BRP

[Bow95] Peter J. Bowler. Book review:

Paul Mengal (ed.), Histoire du concept de récapitulation: Ontogenèse et phylogenèse en biologie et sciences humaines. Paris: Masson, 1993. Pp.

132. ISBN 2-225-84057-1. 180 FF. British Journal for the History of Science, 28(1):116– 117, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027498.

Bowler:1998:EEB

[Bow98] Peter J. Bowler. Evolution and the eucharist: Bishop E. W. Barnes on science and religion in the 1920s and 1930s. British Journal for the History of Science, 31(4):453– 467, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027878.

Brewer:1993:CWG

[BP93a] John Brewer and Roy Porter, editors. Consumption and the world of goods. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1993. ISBN 0-415-03712-3. xix + 564 pp. LCCN HC79.C6 C673 1993.

Bynum:1993:CEH

[BP93b] W. F. (William F.) Bynum and Roy Porter, editors. Companion encyclopedia of the history of medicine. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1993. ISBN 0-415-04771-4 (set), 0-415-09242-6 (vol. 1), 0-415-09243-4 (vol. 2). ???? pp. LCCN R133 .E5

1993. URL http://www.loc.
gov/catdir/enhancements/
fy0648/92026013-d.html.

Berrios:1995:HCP

[BP95] German E. Berrios and Roy Porter, editors. A history of clinical psychiatry: the origin and history of psychiatric disorders. Athlone Press, London, UK, 1995. ISBN 0-485-24011-4 (hardcover), 0-485-24211-7 (paperback). 684 pp. LCCN RC455.2.C4 H57 1999.

Bynum:2002:WHE

[BP02] W. F. (William F.) Bynum and Roy Porter, editors. William Hunter and the Eighteenth-Century medical world. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-52517-9, 0-521-26806-0. xi + 424 pp. LCCN R484 .W55 2002x.

Brown:1991:DVM

[BR91] Laurie M. Brown and Helmut Rechenberg. The development of the vector meson theory in Britain and Japan (1937– 1938). British Journal for the History of Science, 24(4):405– 433, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027128.

Blaxter:1995:DPM

[BR95] Sir K. L. (Kenneth Lyon) Blaxter and Noel F. Robertson. From dearth to plenty: the modern revolution in food production. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-40322-7 (hardcover). xvii + 296 pp. LCCN S494.5.I5 B55

1995. URL http://www.loc.
gov/catdir/description/
cam026/94025259.html;
http://www.loc.gov/
catdir/toc/cam026/
94025259.html.

Bonelli:1996:OMA

[BR96] Federico Bonelli and Lucio Russo. The origin of modern astronomical theories of tides: Chrisogono, de Dominis and their sources. British Journal for the History of Science, 29(4):385–401, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027697.

Bramwell:1989:ETC

[Bra89] Anna Bramwell. Ecology in the Twentieth Century: a history. Yale University Press, New Haven, CT, USA, 1989. ISBN 0-300-04343-0. xiii + 292 pp. LCCN QH541 .B7 1989.

Brackenridge:1990:BRM

[Bra90] J. Bruce Brackenridge. Book review: Marcia Sweet Stayer (ed.). Newton’s Dream. Canada: McGill-Queen’s University Press, 1989. Pp. 135. ISBN 0-7735-0689-6. £15.95. British Journal for the History of Science, 23(1): 107–108, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026812.

Brackenridge:1991:BRA

[Bra91] J. Bruce Brackenridge. Book review: Annabella Kitson (ed.). History and Astrology: Clio and Urania Confer. London: Unwin Paperbacks, 1989. Pp. ix + 272. ISBN 0-04-440522-7. £8.99. British Journal for the History of Science, 24(1):123, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027038.

Brackenridge:1992:BRI

[Bra92a] J. Bruce Brackenridge. Book review: Isaac Newton. The Preliminary Manuscripts for Isaac Newton’s 1687 Principia 1684–1686. Facsimilies of the Original Autographs, now in Cambridge University Library, with an Introduction by D. T. Whiteside. Cambridge: Cambridge University Press, 1989. Pp. xxi + 246. ISBN 0-521-33499-3. £65.00. $115.00. British Journal for the History of Science, 25(3):364–365, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027265.

Brackenridge:1992:BRP

[Bra92b] J. Bruce Brackenridge. Book review: Peter Rowlands. Newton and the Concept of Mass-Energy. Liverpool Historical Essays, 4. Liverpool: Liverpool University Press, 1990. Pp. ii + 34. ISBN 0-85323-187-7. £5.00. British Journal for the History of Science, 25 (2):270–271, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027312.

Brackenridge:1995:BRB

[Bra95] J. Bruce Brackenridge. Book review: Bruce Stephenson, The Music of the Heavens: Kepler’s Harmonic Astronomy. Princeton: Princeton University Press, 1994. Pp. xi + 260. ISBN 0-691-03439-7. £30.00, $39.50. British Journal for the History of Science, 28(4): 464–465, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027551;
 http://www.jstor.org/
stable/4027552.

Brackenridge:1998:BRF

[Bra98] J. Bruce Brackenridge. Book review: François de Gandt, Force and Geometry in Newton’s Principia, translated by Curtis Wilson. Princeton: Princeton University Press, 1995. Pp. xiv + 296. ISBN 0-691-03367-6. $49.50. British Journal for the History of Science, 31 (2):241–250, June 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027765.

Brooks:1984:JBO

[Bro84] John Langdon Brooks. Just before the origin: Alfred Russel Wallace’s Theory of Evolution. Columbia University Press, New York, NY, USA, 1984. ISBN 0-231-05676-1. xiii + 284 + 1 pp. LCCN QH31.W2 B76 1984. US$25.00.

Brock:1985:PPW

[Bro85] W. H. (William Hodson) Brock. From protyle to proton: William Prout and the nature of matter, 1785–1985. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-801-9. xii + 252 pp. LCCN QD22.P85 B76 1985.

Brown:1989:BSL

[Bro89] Chandos Michael Brown. Benjamin Silliman: a life in the young republic. Princeton University Press, Princeton, NJ, USA, 1989. ISBN 0-691-08533-1. xvi + 377 pp. LCCN Q143.S56 B76 1989. US$29.50.

                                        URL http://www.loc.gov/
                                        catdir/description/
                                        prin031/88032086.html.

Brock:1990:BRR

[Bro90a] W. H. Brock. Book review:

R. W. Home (ed.). Australian Science in the Making. Cambridge: Cambridge University Press in association with the Australian Academy of Science, 1988 (1989). Pp. xxvii + 413. ISBN 0-521-35556-7. £35.00, $65.00. British Journal for the History of Science, 23(2):219, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026728.

Brockliss:1990:RU

[Bro90b] L. W. B. Brockliss. Review:

[untitled]. British Journal for the History of Science, 23(4): 482–483, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027216.

Brown:1990:BRR

[Bro90c] Laurie M. Brown. Book review: R. Rubinger. An American Scientist in Early Meiji Japan: The Autobiographical Notes of Thomas C. Mendenhall. Honolulu: University of Hawaii Press, 1989. Pp. xvii + 89. ISBN 0-8248-1177-1. £8.80 (paperback). British Journal for the History of Science, 23 (3):350–351, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026773.

Brown:1990:RU

[Bro90d] Laurie M. Brown. Review: [untitled]. British Journal for the History of Science, 23(4): 478–480, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027213.

Browne:1990:BRR

[Bro90e]  Janet Browne.  Book re-

view: Richard Darwin Keynes. Charles Darwin’s Beagle Diary. Cambridge: Cambridge University Press, 1988. Pp. xxix + 464. ISBN 0-521-23503-0. £35.00, $59.50. British Journal for the History of Science, 23 (3):337–338, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026762.

Bromberg:1991:BRR

[Bro91a] Joan Lisa Bromberg. Book review: Robert W. Smith (with contributions by Paul A. Hanle, Robert H. Kargon and Joseph N. Tatarewicz). The Space Telescope: A Study of NASA, Science, Technology, and Politics. Cambridge: Cambridge University Press, 1989. Pp. xviii + 478. ISBN 0-521-26634-3. £40, $39.50. British Journal for the History of Science, 24 (1):118–119, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027033.

Browne:1991:BRE

[Bro91b]  Janet Browne.  Book re-

view: Eric L. Mills. Biological Oceanography: An Early History, 1870–1960. Ithaca: Cornell University Press, 1989. Pp. xvii + 378. ISBN 0-8014-2340-6. $47.50. British Journal for the History of Science, 24(4): 469–471, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027132.

Browne:1991:BRM

[Bro91c] Janet Browne. Book review:

Marlene J. Norst. Ferdinand Bauer: The Australian Natural History Drawings. Art in Natural History no. 1. London: British Museum of Natural History, 1989. Pp. 120. ISBN 0-565-01048-4. No price given. British Journal for the History of Science, 24(1): 103–104, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027022.

Brock:1992:FHC

[Bro92a] William H. (William Hodson)

Brock, editor. The Fontana history of chemistry. Fontana history of science. Fontana Press, London, UK, 1992. ISBN 0-00-686173-3. xxiii + 744 pp. LCCN QD11 .B75 1992.

Brooke:1992:BRP

[Bro92b] J. H. Brooke. Book review: Peter J. Bowler, Charles Darwin: The Man and His Influence. Oxford: Basil Blackwell, 1990. Pp. xii + 250, illus. ISBN 0-631-16818-4. £19.95. British Journal for the History of Science, 25(4):480– 482, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027069.

Brown:1992:BRM

[Bro92c] Laurie M. Brown. Book review: Masao Watanabe. The Japanese and Western Science, translated by Otto Theodor Benfey. Philadelphia: University of Pennsylvania Press, 1990. Pp. xiv + 141. ISBN 0-8211-8252-3. No price given. British Journal for the History of Science, 25(3):381–382, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027282.

Brock:1993:BRL

[Bro93a]  W. H. Brock.  Book re-

view: Lissa Roberts (ed.), The Chemical Revolution: Context and Practices. Special issue of The Eighteenth Century: Theory and Interpretation, Volume 3, Part 3. Lubbock: Texas Tech University Press, 1992. Pp. 195-286. ISSN 0193-5380. $19.00 (individuals), $32.00 (institutions). British Journal for the History of Science, 26(4):493– 494, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027473.

Brock:1993:BRE

[Bro93b] William H. Brock. Book review: Evan M. Melhado and Tore Frängsmyr (eds.), Enlightenment Science in the Romantic Era: The Chemistry of Berzelius and its Cultural Setting. Cambridge: Cambridge University Press, 1992. Pp. xiv + 246. ISBN 0-521-41775-9. £30.00, $49.95. British Journal for the History of Science, 26(2):245–246, June 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027109.

Brooke:1993:BRA

[Bro93c] John Hedley Brooke. Book review: Adrian Desmond and James Moore, Darwin. London: Michael Joseph, 1991. Pp. xxi + 808. ISBN 0-7181-3430-3. £20.00. British Journal for the History of Science, 26(1): 102–103, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027368.

Brooke:1993:BRO

[Bro93d] John Hedley Brooke. Book review: Olaf Pedersen, The Book of Nature, Vatican City and Notre Dame, Indiana: Vatican Observatory Publications/University of Notre Dame Press, 1992. Pp. vii + 92. ISBN 0-268-00690-3. $9.95. British Journal for the History of Science, 26(4): 486–487, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027469.

Brock:1994:BRH

[Bro94a] W. H. Brock. Book review: Hans Werner Schütt, Eilhard Mitscherlich: Baumeister am Fundament der Chemie. Abhandlungen und Berichte, Neue Folge 8. Munich: R. Oldenbourg Verlag/Deutsches Museum, 1992. Pp. 192. ISBN 3-486-26273-4. No price given. British Journal for the History of Science, 27(2):235– 236, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027449.

Brooke:1994:BRR

[Bro94b] John Hedley Brooke. Book review: Ronald L. Numbers, The Creationists: The Evolution of Scientific Creationism. New York: Alfred A. Knopf, Inc., 1992. Pp. xvii + 458. ISBN 0-679-40104-0. $27.50. British Journal for the History of Science, 27(2):238–239, June 1994.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027452.

Brock:1995:BRU

[Bro95a] W. H. Brock. Book review:

Ursula Klein, Verbindung und Affinität. Die Grundlegung der neuzeitlichen Chemie an der Wende vom 17. zum 18. Jahrhundert. Basel, Boston, Berlin: Birkhäuser, 1994. Pp. ix + 270. ISBN 3-7643-5003-2. £48.00, DM 128.00, SFr 108 00. British Journal for the History of Science, 28(2): 238–239, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027684.

Brooke:1995:TAM

[Bro95b] John Hedley Brooke. Thinking about matter: studies in the history of chemical philosophy, volume CS502 of Collected studies series. Variorum, Aldershot, Hampshire, UK, 1995. ISBN 0-86078-464-9. ???? pp. LCCN QD6 .B76 1995.

Browne:1995:CD

[Bro95c] E. J. (E. Janet) Browne.

Charles Darwin. Alfred A. Knopf, New York, NY, USA, 1995. ISBN 0-224-04202-5 (vol. 1), 0-224-04212-2 (vol. 2). xiii + 605 + 32 pp. LCCN ????

Browne:1995:BRC

[Bro95d] Janet Browne. Book review: Colin Finnet, Paradise Revealed: Natural History in Nineteenth-Century Australia. Melbourne: Museum of Victoria, 1993. Pp. xv + 186. ISBN 0-7306-2494-3. A$ 34.95. British Journal for the History of Science, 28(1):115– 116, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027497.

Browne:1995:RU

[Bro95e] Janet Browne. Review: [untitled]. British Journal for the History of Science, 28(4): 474–475, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027560.

Browne:1995:VCJ

[Bro95f] Janet Browne. Voyaging, volume 1 of Charles Darwin. Cape, London, UK, 1995. ISBN 0-224-04202-5. 605 pp. LCCN ????

Brockliss:1996:BRM

[Bro96a] L. W. B. Brockliss. Book review: Maija Kallinen, Change and Stability: Natural Philosophy at the Academy of Turku, 1640–1713. Helsinki: Finnish Historical Society, 1995. Pp.

439. ISBN 951-710-001-6. No price given. British Journal for the History of Science, 29 (2):233–234, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027840.

Brown:1996:BRA

[Bro96b] Stuart Brown. Book review:

Allison P. Coudert, Leibniz and the Kabbalah. International Archives of the History of Ideas. Dordrecht: Kluwer Academic, 1995. Pp. xvii + 218. ISBN 0-7923-3114-1. £68.00, $109.00, Dfl. 165.00. British Journal for the History of Science, 29(2):234–236, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027841.

Brock:1997:JLC

[Bro97a] W. H. (William Hodson) Brock. Justus von Liebig: the chemical gatekeeper. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56224-4 (hardcover). xiv + 374 pp. LCCN QD22.L7 B76

1997. URL http://www.loc.
gov/catdir/description/
cam027/96025468.html;
http://www.loc.gov/
catdir/samples/cam031/
96025468.html;   http://
www.loc.gov/catdir/toc/
cam031/96025468.html.

Brooke:1997:ISB

[Bro97b] John Brooke. Introduction ‘Small beginnings in a disturbed world’. British Journal for the History of Science, 30(1): 1–4, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027895.

Browne:1997:I

[Bro97c]  Janet Browne.  Introduc-

tion. British Journal for the History of Science, 30(3):256, September 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027860.

Browne:1997:OCM

[Bro97d] Janet Browne. Officers and council members of the British Society for the History of Science, 1947–1997. British Journal for the History of Science, 30(1):77–89, March 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027903.

Brock:1998:BRE

[Bro98]  W. H. Brock.  Book re-

view: Elisabeth Crawford, Arrhenius. From Ionic Theory to the Greenhouse Effect. Uppsala Studies in History of Science, 23. Canton, MA: Science History Publications, 1996. Pp. xiii + 320. ISBN 0-88135-166-0. $49.95. British Journal for the History of Science, 31(4): 469–487, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027876.

Brooke:1999:PAD

[Bro99a] John Hedley Brooke. Presidential address: Does the history of science have a future? British Journal for the History of Science, 32(1): 1–20, March 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027967.

Brown:1999:BM

[Bro99b]  Neil  Brown.

Back matter. British Journal for the History of Science, 32(2):254, June 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4028096.

Brown:1999:BRM

[Bro99c] Neil Brown. Book reviews:

Michael Springford (ed.), Electron: A Centenary Volume. Cambridge: Cambridge University Press, 1997. Pp. xii + 330. ISBN 0-521-56130-2. £37.50, $49.95. Per F. Dahl, Flash of the Cathode Rays: A History of J. J. Thomson’s Electron. Bristol: Institute of Physics Publishing, 1997. Pp. xvii + 526. ISBN 0-7503-0453-7. £29.50, $49.50. Edward A. Davis and Isobel J. Falconer, J. J. Thomson and the Discovery of the Electron. London: Taylor & Francis, 1997. Pp. xxvii + 243. ISBN 0-7484-0696-4. £45.00, $135.00 (hardcover); 0-7484-0720-0. £16.95, $45.00 (paperback). British Journal for the History of Science, 32 (2):237–251, June 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4028086.

Bruce:1987:LMA

[Bru87a] Robert V. Bruce. The launching of modern American science, 1846–1876. The Impact of the Civil War. Knopf, New York, NY, USA, 1987. ISBN 0-394-55394-2. x + 446 + 4 pp. LCCN Q127.U6 B78 1987.

Brundell:1987:PGA

[Bru87b]  Barry  Brun-

dell. Pierre Gassendi: from Aristotelianism to a new natural philosophy, volume 30 of Synthese historical library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1987. ISBN 90-277-2428-8. x + 251 pp. LCCN B1887 .B84 1987.

Brundage:1988:EPM

[Bru88a] Anthony Brundage. England’s “Prussian minister”: Edwin Chadwick and the politics of government growth, 1832–1854. Pennsylvania State University Press, University Park, PA, USA, 1988. ISBN 0-271-00629-3. 208 pp. LCCN DA565.C38 B78 1988. US$24.95.

Brush:1988:HMS

[Bru88b] Stephen G. Brush. The history of modern science: a guide to the second scientific revolution, 1800–1950. The Iowa State University Press series in the history of technology and science. Iowa State University Press, Ames, IA, USA, 1988. ISBN 0-8138-0883-9. xv + 544 pp. LCCN Q125 .B88 1988.

Brush:1988:HPS

[Bru88c] Stephen G. Brush, editor.

History of physics: selected reprints. American Association of Physics Teachers, College Park, MD, USA, 1988. ISBN 0-917853-29-6. 235 pp. LCCN QC7.5 .H57 1988.

Brunton:1991:BRR

[Bru91] Deborah Brunton. Book review: Richard Davenport-Hines. Sex, Death and Punishment: Attitudes to Sex and Sexuality in Britain since the Renaissance. London: Collins, 1990. Pp. xv + 439. ISBN 0-00-217956-3. £20. British Journal for the History of Science, 24(4):496–497, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027154.

Brunton:1993:BRR

[Bru93] Deborah Brunton. Book review: Richard B. Fisher, Edward Jenner 1749–1823. London: André Deutsch, 1991.

Pp. 361. ISBN 0-233-98681-2. £20.00. British Journal for the History of Science, 26(1): 96, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027360.

Brunner:1995:FPP

[Bru95] José Brunner. Freud: the politics of psychoanalysis. Blackwell, Oxford, UK, 1. publ edition, 1995. ISBN 0-631-16404-9. xiv + 238 pp. LCCN BF109.F74 B78 1995.

Brunton:1996:BRL

[Bru96] Deborah Brunton. Book review: Lawrence I. Conrad, Michael Neve, Vivian Nutton, Roy Porter and Andrew Wear, The Western Medical Tradition 800 BC to AD 1800. Cambridge: Cambridge University Press, 1995. Pp. xiv + 556. ISBN, 0-521-38135-5, £60.00, $89.95 (hardcover); 0-521-47564-3, £24.95, $34.95 (paperback). British Journal for the History of Science, 29 (2):253–254, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027853.

Bryder:1988:BMM

[Bry88] Linda Bryder. Below the magic mountain: a social history of tuberculosis in Twentieth-Century Britain. Oxford historical monographs. Clarendon Press, New York, NY, USA, 1988. ISBN 0-19-822947-X. xiv + 298 pp. LCCN RC316.G7 B78 1988. US$27.00 (US$59.00

U.S.). URL http://www.loc.
gov/catdir/enhancements/
fy0606/87031495-d.html.

Bryden:1991:BRC

[Bry91] D. J. Bryden. Book review: C.

Blondel, F. Parot, A. Turner and M. Williams. Studies in the History of Scientific Instruments, papers presented at the 7th Symposium of the Scientific Instruments Commission of the Union Internationale d’Histoire des Sciences, Paris 15-19 September 1987. London: Rogers Turner Books Ltd; for the Centre de Recherche en Histoire des Sciences et des Techniques de la Cité des Sciences et de l’Industrie, 1989. Pp. 290. ISBN 0-9502557-8-5. £35.00. British Journal for the History of Science, 24(4): 490–491, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027150.

Bryden:1992:BRG

[Bry92a] D. J. Bryden. Book review:

Gerard L’Estrange Turner. Scientific Instruments and Experimental Philosophy, 1550–1850. Aldershot, Variorum Press Collected Studies Series. CS 331, 1990. Pp. xii + 329. ISBN 0-86078-280-8. £49.50. British Journal for the History of Science, 25 (3):383, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027284.

Bryden:1992:NBH

[Bry92b] D. J. Bryden, editor. Napier’s bones: a history and instruction manual. Harriet Wynter, London, UK, 1992. ISBN 0-9507258-2-X. 23 pp. LCCN QA71 .B79 1992.

Bowers:1991:CPS

[BS91] Brian Bowers and Lenore Symons, editors. Curiosity perfectly satisfyed: Faraday’s travels in Europe, 1813–1815, volume 16 of IEE history of technology series. P. Peregrinus in association with the Science Museum, London, UK, 1991. ISBN 0-86341-234-3. xvi + 168 pp. LCCN QC16.F2 C87 1991.

Burkhardt:1985:CCD

[BS10] Frederick Burkhardt and Sydney Smith, editors. The correspondence of Charles Darwin. Cambridge University Press, Cambridge, UK, 1985– 2010. ISBN 0-521-25587-2 (vol. 1), 0-521-25588-0 (vol. 2), 0-521-25589-9 (vol. 3), 0-521-25590-2 (vol. 4), 0-521-25591-0 (vol. 5), 0-521-25586-4 (vol. 6), 0-521-84459-2 (vol. 14), 0-521-85931-X (vol. 15 : hardback), 0-521-88195-1 (vol. 16, pt. 1 : hardback), 0-521-88196-X (vol. 16, pt. 2 : hardback), 0-521-51836-9 (vol. 16, 2 vol. hardback set), 0-521-19030-4 (vol. 17), 0-521-76889-6 (vol. 18), 0-521-44241-9. ???? pp. LCCN QH31.D2 A4 1985.

Burkhardt:1985:CCC

[BSK^{+}85] Frederick Burkhardt, Sydney Smith, David Kohn, William Montgomery, and Charles Darwin, editors. A calendar of the correspondence of Charles Darwin, 1821–1882, volume 369 of Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1985. ISBN 0-8240-9224-4. 690 + 1 pp. LCCN QH31.D2 C15 1985.

Burkhardt:1994:CCC

[BSK^{+}94] Frederick Burkhardt, Sydney Smith, David Kohn, William Montgomery, Charles Darwin, et al., editors. A calendar of the correspondence of Charles Darwin, 1821–1882, with supplement. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-43423-8 (hardcover). viii 690, 46 pp. LCCN QH31.D2 C15

1994. URL http://www.loc.
gov/catdir/description/
cam026/94191109.html;
http://www.loc.gov/
catdir/toc/cam029/
94191109.html.

Batchelor:1996:LLG

[BT96] G. K. (George Keith) Batchelor and Sir Geoffrey Ingram Taylor. The life and legacy of G. I. Taylor. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-46121-9. xv + 285 pp. LCCN Q143.T29

B38 1996. URL http://www.
loc.gov/catdir/
description/cam027/
95047579.html;   http://
www.loc.gov/catdir/toc/
cam027/95047579.html.

Buchwald:1985:MMA

[Buc85]  Jed Z. Buchwald.  From

Maxwell to microphysics: aspects of electromagnetic theory in the last quarter of the Nineteenth Century. University of Chicago Press, Chicago, IL, USA, 1985. ISBN 0-226-07882-5 (hardcover), 0-226-07883-3 (paperback). xiii + 339 pp. LCCN QC670 .B83 1985. US$55.00.

Buchwald:1989:RWT

[Buc89] Jed Z. Buchwald. The rise of the wave theory of light: optical theory and experiment in the early Nineteenth Century. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-07886-8 (paperback), 0-226-07884-1 (hardcover). xxiv + 474 pp. LCCN QC403 .B83

1989. URL http://www.loc.
gov/catdir/description/
uchi052/88018647.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/88018647-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/88018647-b.html.

Buchanan:1990:BRR

[Buc90] R. A. Buchanan. Book review:

Robert Maddin (ed.). The Beginning of the Use of Metals and Alloys. Cambridge, Mass.: The MIT Press, 1988. Pp. xiv + 393. ISBN 0-262-13232-X. £40.50. British Journal for the History of Science, 23(3): 343–344, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026767.

Buchwald:1991:EE

[Buc91] Jed Z. Buchwald. Energy and empire. British Journal for the History of Science, 24(1): 85–94, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027018.

Buchwald:1992:BRP

[Buc92] J. Z. Buchwald. Book review: P. M. Harman (ed.). The Scientific Letters and Papers of James Clerk Maxwell. Volume I: 1846–1862. Cambridge: Cambridge University Press, 1990. Pp. xxviii + 748. ISBN 0-521-25625-9. £135.00, $195. British Journal for the History of Science, 25(3):369– 371, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027269.

Buchwald:1994:CSE

[Buc94] Jed Z. Buchwald. The creation of scientific effects: Heinrich Hertz and electric waves. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-07887-6, 0-226-07888-4 (paperback). xiv + 482 pp. LCCN QC661 .B85 1994. URL http://www.

                                        loc.gov/catdir/
                                        description/uchi052/
                                        93041783.html;   http://
                                        www.loc.gov/catdir/toc/
                                        uchi051/93041783.html.

Bucchi:1998:ISC

[Buc98] Massimiano Bucchi. Images of science in the classroom: wallcharts and science education 1850–1920. British Journal for the History of Science, 31(2):161–184, June 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027761.

Bud:1993:ULH

[Bud93] Robert Bud. The uses of life: a history of biotechnology. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-38240-8. xvii + 299 pp. LCCN TP248.18 .B83

                                        1993. URL http://www.loc.
                                        gov/catdir/description/
                                        cam032/92019513.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam032/
                                        92019513.html.

Bud:1997:HSS

[Bud97] Robert Bud. History of science and the Science Museum. British Journal for the History of Science, 30(1):47– 50, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027899.

Bud:1998:PNE

[Bud98] Robert Bud. Penicillin and the new Elizabethans. British Journal for the History of Science, 31(3):305–333, September 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027815.

Buffetaut:1987:SHV

[Buf87] Eric Buffetaut. A short history of vertebrate palaeontology. Croom Helm, London, UK, 1987. ISBN 0-7099-3962-0. 233 + 16 pp. LCCN QE841 .B84 1987. US$52.50.

Bugos:1996:BRS

[Bug96] Glenn E. Bugos. Book review: Susan Wright, Molecular Politics: Developing American and British Regulatory Policy for Genetic Engineering, 1972–1982. Chicago: University of Chicago Press, 1994. Pp. xxiii + 591. ISBN 0-226-91065-2, £59.95. $75.00 (hardcover); 0-226-91066-0, £23.95, $29.95 (paperback). British Journal for the History of Science, 29 (1):118–119, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027536.

Buhler:1981:GBS

[Büh81] W. K. (Walter Kaufmann) Bühler. Gauss: a biographical study. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 3-540-10662-6, 0-387-10662-6. viii + 208 pp. LCCN QA29.G3

B83. URL http://catalog.
hathitrust.org/api/
volumes/oclc/7197827.html.

Burke:1983:USA

[Bur83] John G. Burke, editor. The uses of science in the age of Newton, volume 8 of Publications from the Clark Library professorship, UCLA. University of California Press, Berkeley, CA, USA, 1983. ISBN 0-520-04970-5. 204 pp. LCCN Q127.G4 U83 1983.

Burwick:1987:AOF

[Bur87] Frederick Burwick, editor. Approaches to organic form: permutations in science and culture, volume 105 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1987. ISBN 90-277-2541-1, 94-009-3917-5. ISSN 0068-0346. xix + 286 pp. LCCN Q174 .B67 vol. 105.

US$64.00(U.S.). URL https:

//link.springer.com/book/
10.1007/978-94-009-3917-2.

Burchell:1993:BRS

[Bur93] R. A. Burchell. Book review: Simon Baatz, Knowledge, Culture and Science in the Metropolis: The New York Academy of Sciences, 1817– 1970. Annals of the New York Academy of Sciences, vol. 584. New York: New York Academy of Sciences, 1990. Pp. ix + 269. ISBN 0-89766-545-7 (hardcover); 0-89766-546-5 (paperback). No price given. British Journal for the History of Science, 26 (3):372, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027421.

Burbridge:1994:GEF

[Bur94] David Burbridge. Galton’s 100: An exploration of Francis Galton’s imagery studies. British Journal for the History of Science, 27(4):443– 463, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027625.

Burnett:1998:BRR

[Bur98] Charles Burnett. Book review:

Roshdi Rashed (ed.), Encyclopedia of the History of Arabic Science, in collaboration with Régis Morelon, 3 vols., London: Routledge, 1996. Pp. xiv + 1105. ISBN 0-415-02063-8. £160 (set). British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027779.

Buruma:1999:JSK

[Bur99] Y. Buruma, editor. 100 jaar strafrecht: klassieke teksten van de twintigste eeuw. (Dutch) [100 years of criminal law: traditional texts of the Twentieth Century ]. Amsterdam University Press, Amsterdam, The Netherlands, 1999. ISBN 90-5356-362-8, 90-5356-361-X (paperback). 440 pp. LCCN KKM3800 .A17 1999.

Burkhardt:2000:CCD

[Bur00] Frederick Burkhardt. The correspondence of Charles Darwin, 1837–1843. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-25588-0. xxxiii + 603 pp. LCCN ????

Buttner:1983:HCC

[Büt83] J. (Johannes) Büttner, editor. History of clinical chemistry. W. de Gruyter, Berlin, Germany, 1983. ISBN 3-11-008912-2. 91 pp. LCCN RB40 .H56 1983. DM98.00 (est.) (US$49.00 U.S. : est.).

Butler:1992:BRJ

[But92a] A. R. Butler. Book review: James Reardon-Anderson. The Study of Change: Chemistry in China. 1840–1949. Cambridge: Cambridge University Press, 1991. Pp. xix + 444. ISBN 0-521-39150-4. £45.00, $59.50. British Journal for the History of Science, 25(3): 376, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027276.

Butler:1992:STM

[But92b] Stella V. F. Butler. Science and technology museums. Leicester museum studies series. Leicester University Press, Leicester, UK, 1992. ISBN 0-7185-1357-6. xiii + 149 pp. LCCN T179 .B87 1992.

Butler:1994:BRM

[But94] Stella Butler. Book review:

Mary Holbrook, Science Preserved: A Directory of Scientific Instruments in Collections in the United Kingdom and Eire, with additions and revisions by R. G. W. Anderson and D. J. Bryden. London: HMSO, 1992. Pp. 271. ISBN 0-11-290060-7. £35.00. British Journal for the History of Science, 27(2):242–243, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027456.

Butters:1996:TVS

[But96] Suzanne B. Butters. The triumph of Vulcan: sculptors’ tools, porphyry, and the prince in ducal Florence, volume 14 of Villa i Tatti. L. S. Olschki, Firenze, Italy, 1996. ISBN 88-222-4411-7. 724 pp. LCCN NB615 .B84 1996.

Bensaude-Vincent:1990:VCR

[BV90] Bernadette Bensaude-Vincent. A view of the chemical revolution through contemporary textbooks: Lavoisier, Fourcroy and Chaptal. British Journal for the History of Science, 23(4):435–460, December 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027200.

Bensaude-Vincent:1991:BRT

[BV91] Bernadette Bensaude-Vincent. Book review: Töre Frängsmyr, J. L. Heilbron and Robin E. Rider (eds.). The Quantifying Spirit in the Eighteenth Century. Berkeley, Los Angeles, Oxford: University of California Press, 1990. Pp. vii + 411. ISBN 0-520-07022-4. $35.00. British Journal for the History of Science, 24(2): 259–260, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027182.

Bensaude-Vincent:1993:LMR

[BV93] Bernadette Bensaude-Vincent. Lavoisier: mémoires d’une révolution. (French) [Lavoisier: memoirs of a revolutions]. Figures de la science. Flammarion, Paris, France, 1993. ISBN 2-08-211208-X. 469 pp. LCCN QD22.L4 B28 1993. 160F.

Bridson:1990:PAA

[BW90] Gavin D. R. Bridson and James J. (James Joseph) White. Plant, animal and anatomical illustration in art and science: a bibliographical guide from the 16th century to the present day. St Paul’s Bibliographies in association with Hunt Institute for Botanical Documentation; Detroit : Omnigraphics, Winchester, UK, 1990. ISBN 0-906795-81-8. xxxvii + 450 pp. LCCN QH46.5 .P43

1990. URL http://catalog.
hathitrust.org/api/
volumes/oclc/24505682.
html.

Bynum:1990:BRR

[Byn90] W. F. Bynum. Book review: Roger Cooter. Phrenology in the British Isles: An Annotated, Historical Bibliography and Index. Metuchen, N.J. and London: The Scarecrow Press Inc., 1989. Pp. xviii + 431. ISBN 0-8108-2165-6. £47.25. British Journal for the History of Science, 23(3): 346–347, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026769.

Carding:1992:GHT

[C^{+}92] Janet Carding et al., editors.

Guide to the history of technology in Europe. Science Museum, London, UK, 1992. ISBN 0-901805-51-3. 142 pp. LCCN Q180.E9 G85 1992.

Conrad:1995:WMT

[C^{+}95] Lawrence I. Conrad et al., editors. The Western medical tradition: 800 B.C.–1800 A.D. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-38135-5. xiv + 556 pp. LCCN R131 .W47

1995. URL http://www.loc.
gov/catdir/description/
cam026/94034823.html;
http://www.loc.gov/
catdir/toc/cam021/
94034823.html.

Callick:1990:MMB

[Cal90] E. B. Callick. Metres to microwaves: British development of active components for radar systems, 1937–1944, volume 11 of IEE history of technology series. P. Peregrinus Ltd. on behalf of the IEE, London, UK, 1990. ISBN 0-86341-212-2. xi + 240 pp. LCCN TK6574.4.G7 C35 1990.

Campbell:1990:BRJ

[Cam90] Eila M. J. Campbell. Book review: J. B. Harley & David Woodward (eds). The History of Cartography, Vol. I: Cartography in Prehistoric, Ancient and Medieval Europe and the Mediterranean. Chicago and London: University of Chicago Press, 1987. Pp. xxii + 599, 40 colour plates. ISBN 0-226-31633-5. $100.00. British Journal for the History of Science, 23(1):120–122, March 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026825.

Campbell:1992:BRF

[Cam92] Alex Campbell. Book review:

Fred Aftalion. A History of the International Chemical Industry, translated by Otto T. Benfey. Philadelphia: University of Pennsylvania Press, 1991. Pp. xxiv + 411. ISBN 0-8122-8207-8, £37.95; 0-8122-1297-5, £14.20. British Journal for the History of Science, 25(3): 374–375, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027274.

Cantor:1983:OAN

[Can83] G. N. Cantor. Optics after Newton: theories of light in Britain and Ireland, 1704– 1840. Manchester University Press, Manchester, UK, 1983. ISBN 0-7190-0938-3. ix + 257 pp. LCCN QC352 .C36 1983.

Canguilhem:1988:IRH

[Can88] Georges Canguilhem. Ideology and rationality in the history of the life sciences. MIT Press, Cambridge, MA, USA, 1988. ISBN 0-262-03137-X. xi + 160 pp. LCCN QH305 .C2613 1988.

Cantor:1990:BRC

[Can90a] David Cantor. Book review: Catherine Caufield. Multiple Exposures, Chronicles of the Radiation Age. London: Seeker and Warburg, 1989. Pp. xii + 304. ISBN 0-436-09478-9. £12.95. British Journal for the History of Science, 23(2): 217–218, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026727.

Cantor:1990:BRJ

[Can90b] Geoffrey Cantor. Book review: Jed Z. Buchwald. The Rise of the Wave Theory of Light: Optical Theory and Experiment in the Early Nineteenth Century. Chicago and London: University of Chicago Press, 1989. Pp. xxiv + 474. ISBN 0-226-07884-1 (hardcover) and 0-226-07886-8 (paperback). £59.95, $86.25 (hardcover) £19.95, $28.75 (paperback). British Journal for the History of Science, 23(3): 365–367, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026786.

Cantor:1991:BRM

[Can91a] Geoffrey Cantor. Book review: Margaret Atherton. Berkeley’s Revolution in Vision. Ithaca and London: Cornell University Press, 1990. Pp. xii + 249. ISBN 0-8014-2358-9. S29.90. British Journal for the History of Science, 24(2): 257–258, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027180.

Cantor:1991:MFS

[Can91b] Geoffrey N. Cantor, editor.

Michael Faraday: Sandemanian and Scientist: a study of science and religion in the Nineteenth Century. MacMillan Publishing Company, New York, NY, USA, 1991. ISBN 0-333-55077-3 (hardcover), 0-333-58802-9 (paperback). xi + 359 pp. LCCN QC16.F2.

Caneva:1993:RMC

[Can93a] Kenneth L. Caneva. Robert Mayer and the conservation of energy. Princeton University Press, Princeton, NJ, USA, 1993. ISBN 0-691-08758-X. xxiii + 439 pp. LCCN QC73.8.C6 C36 1993. US$49.50. URL http://www.

loc.gov/catdir/
description/prin031/
92025400.html;   http://
www.loc.gov/catdir/toc/
prin031/92025400.html.

Cannell:1993:GGM

[Can93b] D. M. (Doris Mary) Cannell. George Green: mathematician and physicist, 1793– 1841: the background to his life and work. Athlone Press, London, UK, 1993. ISBN 0-485-11433-X. xxvi + 265 + 8 pp. LCCN QC16.G64 C36 1993. US$70.00.

Cantor:1994:MBT

[Can94] Geoffrey Cantor. The making of a British theoretical physicist — E. C. Stoner’s early career. British Journal for the History of Science, 27(3):277– 290, September 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027599.

Cantor:1996:BRC

[Can96] Geoffrey Cantor. Book review: Casper Hakfoort, Optics in the Age of Euler: Conceptions of the Nature of Light, 1700–1795. Cambridge: Cambridge University Press, 1995, Pp. vii + 243. ISBN 0-521-40471-1. £35.00, $64.95. British Journal for the History of Science, 29(2):236–238, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027842.

Cantor:1997:PAC

[Can97] Geoffrey Cantor. Presidential address: Charles Singer and the early years of the British Society for the History of Science. British Journal for the History of Science, 30 (1):5–23, March 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027896.

Cantor:1999:BCP

[Can99] Geoffrey Cantor. Boyling over:

a commentary on the preceding papers. British Journal for the History of Science, 32(3): 315–324, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027989.

Carpenter:1986:HSV

[Car86] Kenneth J. (Kenneth John)

Carpenter. The history of scurvy and vitamin C. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-32029-1. viii + 288 pp. LCCN RC627.S36 C36

1986. URL http://www.loc.
gov/catdir/description/
cam032/85025464.html;
http://www.loc.gov/
catdir/toc/cam031/
85025464.html.

Carey:1988:TEU

[Car88]  S.  Warren

(Samuel Warren) Carey. Theories of the Earth and universe: a history of dogma in the earth sciences. Stanford University Press, Stanford, CA, USA, 1988. ISBN 0-8047-1364-2. xviii + 413 pp. LCCN QE6 .C37 1988. US$45.00.

Cardwell:1989:JJB

[Car89] D. S. L. (Donald Stephen Lowell) Cardwell. James Joule: a biography. Manchester University Press, Manchester, UK, 1989. ISBN 0-7190-3025-0, 0-7190-3479-5 (paperback). ix + 333 pp. LCCN QC16.J68 C368 1989.

Cardwell:1994:FHT

[Car94] D. S. L. (Donald Stephen Lowell) Cardwell. The Fontana history of technology. Fontana history of science. Fontana Press, London, UK, 1994. ISBN 0-00-686176-8 (paperback). xviii + 565 + 8 pp. LCCN T15 .C37 1994.

Cardwell:1995:BRR

[Car95] Donald Cardwell. Book review: Richard L. Hills, Power from Wind: A History of Windmill Technology, Cambridge: Cambridge University Press, 1994. Pp. ix + 324, illus. ISBN 0-521-41398-2. £45.00, $59.95. British Journal for the History of Science, 28(3):372–373, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027665.

Casini:1993:BRF

[Cas93]  Paolo Casini.  Book

review: François-Marie Arouet de Voltaire, Éléments de la philosophie de Newton, critical edition by Robert L. Walters and W. H. Barber. The Complete Works of Voltaire, 15. Oxford: Voltaire Foundation, Taylor Institution, 1992. Pp. xxii + 850. ISBN 0-7294-0374-2. No price given. British Journal for the History of Science, 26(3):

360–361, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027411.

Casini:1994:BRZ

[Cas94] Paolo Casini. Book review:

Zev Bechler, Newton’s Physics and the Conceptual Structure of the Scientific Revolution. Boston Studies in the Philosophy of Science 127. Dordrecht: Kluwer Academic Publishers, 1991. Pp. xviii + 588. ISBN 0-7923-1054-3. £103.00, $189.00, Dfl. 300.00. British Journal for the History of Science, 27 (2):229–230, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027443.

Caton:1988:PPO

[Cat88] Hiram Caton. The politics of progress: the origins and development of the commercial republic, 1600–1835. University of Florida Press, Gainesville, FL, USA, 1988. ISBN 0-8130-0847-6. xii + 627 pp. LCCN D217 .C33 1988. URL http:

//www.loc.gov/catdir/toc/
fy0714/86030886.html.

Caufield:1989:MEC

[Cau89] Catherine Caufield. Multiple exposures: chronicles of the radiation age. Secker and Warburg, London, UK, 1989. ISBN 0-436-09478-9. xii + 304 + 8 pp. LCCN RA569 .C38 1989b. US$12.95.

Cavendish:1996:GNP

[Cav96] Margaret Cavendish, Duchess of Newcastle. Grounds of natural philosophy, volume 2 of Women in the sciences. Locust Hill Press, West Cornwall, CT, 1996. ISBN 0-933951-66-3. xx + 311 pp. LCCN QC19 .N57 1996.

Chant:1990:STE

[CB^{+}90] Colin Chant, Richard Bessel, et al., editors. Science, technology and everyday life, 1870– 1950. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1990. ISBN 0-415-00037-8 (hardcover), 0-415-03557-0 (paperback). 391 pp. LCCN ????

Creese:1994:BWW

[CC94] Mary R. S. Creese and Thomas M. Creese. British women who contributed to research in the geological sciences in the nineteenth century. British Journal for the History of Science, 27(1):23–54, March 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027579.

Cannon:1981:EDV

[CD81] John T. Cannon and Sigalia Dostrovsky. The evolution of dynamics: vibration theory from 1687 to 1742: with 10 illustrations, volume 6 of Studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 0-387-90626-6. ix + 184 pp. LCCN QA865 .C36.

Canguilhem:1994:VRS

[CD94]  Georges  Canguilhem

and Franøcois Delaporte. A vital rationalist: selected writings from Georges Canguilhem. Zone Books, New York, NY, USA, 1994. ISBN 0-942299-72-8. 481 pp. LCCN Q125 .C34 1994. US$26.95.

Cohen:1990:PRM

[CDS90] I. Bernard Cohen, K. E. Duffin, and Stuart Strickland, editors. Puritanism and the rise of modern science: the Merton thesis. Rutgers University Press, New Brunswick, NJ, USA, 1990. ISBN 0-8135-1529-7, 0-8135-1530-0 (paperback). xiii + 402 pp. LCCN Q127.G4 M4636 1990. US$38.00, US$18.00.

Ceruzzi:1989:BLF

[Cer89]  Paul E. Ceruzzi.  Beyond

the limits: flight enters the computer age. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-03143-4, 0-262-53082-1 (paperback). xi + 270 pp. LCCN TL553 .C38 1989.

Carnot:1986:RMP

[CF86] Sadi Carnot and Robert Fox.

Reflexions on the motive power of fire: a critical edition with the surviving scientific manuscripts. Manchester University Press, Manchester, UK, 1986. ISBN 0-7190-1741-6, 0-936508-16-7. viii + 230 pp. LCCN ???? URL http://

catalog.hathitrust.org/
api/volumes/oclc/
12976087.html.

Cantor:1991:F

[CGJ91] G. N. Cantor, David Gooding, and Frank A. J. L. James. Faraday. MacMillan Publishing Company, New York, NY, USA, 1991. ISBN 0-333-54291-6 (paperback). xiii + 109 pp. LCCN QC16.F2 C36 1991. US$8.99.

Charlton:1982:HTS

[Cha82] T. M. (Thomas Malcolm) Charlton. A history of theory of structures in the Nineteenth Century. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-23419-0. viii + 194 pp. LCCN TA646 .C45

1982. URL http://www.loc.
gov/catdir/description/
cam022/81015515.html;
http://www.loc.gov/
catdir/samples/cam031/
81015515.html;   http://
www.loc.gov/catdir/toc/
cam031/81015515.html.

Chandrasekhar:1983:EMD

[Cha83]  S.  (Subrahmanyan)

Chandrasekhar. Eddington, the most distinguished astrophysicist of his time. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25746-8. 64 pp. LCCN QB36.E33 C48

1983. URL http://www.loc.
gov/catdir/enhancements/
fy0914/84123568-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0914/84123568-t.html.

Chaplain:1984:CTL

[Cha84]  Jean-Michel Chaplain.  La

chambre des tisseurs: Louviers, cité drapière, 1680–1840. (French) [The weavers room: Louviers, drapery city, 1680–1840]. Collection Milieux, 0291-7157. Champ Vallon, Seyssel, France, 1984. ISBN 2-903528-40-3. 302 pp. LCCN HD9862.8.L68 C46 1984. 178F.

Chapman:1988:CEO

[Cha88] Michael Chapman. Constructive evolution: origins and development of Piaget’s thought. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-33163-3, 0-521-36712-3 (paperback). xii + 459 pp. LCCN BF723.C5 C484

1988. URL http://www.loc.
gov/catdir/description/
cam023/87021816.html;
http://www.loc.gov/
catdir/toc/cam029/
87021816.html.

Chalmers:1990:SF

[Cha90a] A. F. (Alan Francis) Chalmers.

Science and its fabrication. Open University, Milton Keynes, UK, 1990. ISBN 0-335-09318-3 (hardcover), 0-335-09317-5 (paperback). xii + 142 pp. LCCN Q175 .C4457

                                        1990. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/20295911.
                                        html.

Chalmers:1990:SFC

[Cha90b] Alan Chalmers. Science and its fabrication. Open Univ. Press, Buckingham, UK, 1990. ISBN 0-335-09318-3, 0-335-09317-5. xii + 142 pp. LCCN ????

Chapman:1990:BRD

[Cha90c] Allan Chapman. Book review: Derek Howse. Nevil Maskelyne, The Seaman’s Astronomer. Cambridge: Cambridge University Press, 1989. Pp. xiv + 280. ISBN 0-521-36261-X. £40.00, $59.50. British Journal for the History of Science, 23(3):377–379, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026793.

Chapman:1990:DCD

[Cha90d] Allan Chapman. Dividing the circle: development and critical measurement of celestial angles (1500–1850). Ellis Horwood library of space science and space technology. Series in astronomy. Ellis Horwood Ltd., New York, NY, USA, 1990. ISBN 0-13-217555-X. 209 pp. LCCN QB147 .C48 1990. US$15.95.

Channell:1991:VMS

[Cha91] David F. Channell. The vital machine: a study of technology and organic life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-506040-7. xi + 192 pp. LCCN T14 .C464 1991. US$22.95. URL http:

//www.loc.gov/catdir/
enhancements/fy0639/
90020273-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0725/90020273-b.html.

Chaldecott:1993:OHR

[Cha93] John A. Chaldecott. Obituary: Henry Reginald Calvert (1904–1992). British Journal for the History of Science, 26 (3):347–349, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027402.

Chapman:1994:DCD

[Cha94] Allan Chapman. Dividing the circle: the development of critical angular measurement in astronomy 1500–1850. Wiley in association with Praxis, Chichester, UK, 1994. ISBN 0-13-217555-X. 209 pp. LCCN ????

Chaloner:1997:MWE

[Cha97] Clinton Chaloner. The most wonderful experiment in the world: a history of the cloud chamber. British Journal for the History of Science, 30(3):

357–374, September 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027867.

Chemla:1990:RU

[Che90] Karine Chemla. Review: [untitled]. British Journal for the History of Science, 23(4): 493–495, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027225.

Collie:1991:HWS

[CHG91] Michael Collie, Thomas Henry Huxley, and George Gordon. Huxley at work: with the scientific correspondence of T. H. Huxley and the Rev. Dr. George Gordon of Birnie, near Elgin. MacMillan Publishing Company, New York, NY, USA, 1991. ISBN 0-333-51059-3. xii + 158 + 6 pp. LCCN QE22.H94 A4 1991.

Christie:1993:ANC

[Chr93] J. R. R. Christie. Aurora, Nemesis and Clio. British Journal for the History of Science, 26(4):391–405, December 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027463.

Christianson:1995:EHM

[Chr95] Gale E. Christianson. Edwin Hubble: mariner of the nebulae. Farrar, Straus, Giroux, New York, NY, USA, 1995. ISBN 0-374-14660-8. x + 420 pp. LCCN QB36.H83 C48 1995.

Crawford:1987:NPC

[CHU87] Elisabeth T. Crawford, J. L.

Heilbron, and Rebecca Ullrich. The Nobel population 1901– 1937: a census of the nominators and nominees for the prizes in physics and chemistry, volume 11; 4 of Berkeley papers in history of science; Uppsala studies in history of science. Office for History of Science and Technology, University of California at Berkeley, Berkeley, CA, USA, 1987. ISBN 0-918102-15-4 (paperback). 337 pp. LCCN QC28 .C74 1987.

Churchill:1991:BRF

[Chu91] Frederick B. Churchill. Book review: François Duchesneau. Genèse de la théorie cellulaire. Montréal-Paris: Bellarmin-Vrin, 1987. Pp. 387 + 22. ISBN 2-89007-646-6 (Bellarmin); 2-7116-9427-5 (Vrin). Can $24.00. British Journal for the History of Science, 24 (1):105–107, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027024.

Cittadino:1990:NLD

[Cit90]  Eugene Cittadino.  Nature

as the laboratory: Darwinian plant ecology in the German Empire, 1880–1900. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-34045-4. xi + 199 pp. LCCN QK900.73.G3 C58

1990. URL http://www.loc.
gov/catdir/description/
cam024/90001422.html;
http://www.loc.gov/
catdir/samples/cam031/
90001422.html;   http://
www.loc.gov/catdir/toc/
cam029/90001422.html.

Clarke:1987:NCO

[CJ87] Edwin Clarke and L. S. Jacyna. Nineteenth-Century origins of neuroscientific concepts. University of California Press, Berkeley, CA, USA, 1987. ISBN 0-520-05694-9. 593 pp. LCCN QP353 .C57 1987.

Cunningham:1990:RS

[CJ90] Andrew Cunningham and Nicholas Jardine, editors. Romanticism and the sciences. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-35602-4, 0-521-35685-7 (paperback). xxii + 345 pp. LCCN Q125 .R735

1990. URL http://www.loc.
gov/catdir/description/
cam023/89017310.html;
http://www.loc.gov/
catdir/toc/cam023/
89017310.html.

Campbell-Kelly:1989:IBT

[CK89] Martin Campbell-Kelly. ICL: a Business and Technical History. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-853918-5. xii + 409 + 36 pp. LCCN HD9696.C64 I463 1989.

US$37.25. URL http://www.

loc.gov/catdir/
enhancements/fy0639/
89016318-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0639/89016318-t.html.

Cotterell:1990:MPI

[CK90] Brian Cotterell and Johan Kamminga. Mechanics of preindustrial technology: an introduction to the mechanics of ancient and traditional material culture. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-34194-9. xv + 325 pp. LCCN QC125.2 .C68

1989. URL http://www.loc.
gov/catdir/description/
cam023/88029238.html;
http://www.loc.gov/
catdir/toc/cam028/
88029238.html.

Campbell-Kelly:1991:BRJ

[CK91] Martin Campbell-Kelly. Book review: John Hendry. Innovating for Failure: Government Policy and the Early British Computer Industry. Cambridge, Mass, and London: MIT Press, 1990. Pp. xviii + 240. ISBN 0-262-08187-3. £31.50. British Journal for the History of Science, 24(4): 479–480, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027140.

Campbell-Kelly:1996:CHI

[CKA96] Martin Campbell-Kelly and William Aspray. Computer: a history of the information machine. The Sloan technology series. Basic Books, New York, NY, USA, 1996. ISBN 0-465-02989-2. ix + 342 pp. LCCN QA76.17 .C36

                                        1996. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0831/96002098-d.html.

Cooper:1987:WCB

[CL87] William Cooper and Stanton J. Linden. William Cooper’s A catalogue of chymicall books, 1673–88, volume 670 of Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, verified. edition, 1987. ISBN 0-8240-8557-4. liv + 159 pp. LCCN Z5522 .C77 1987.

Chard:1996:TTP

[CL96] Chloe Chard and Helen Langdon, editors. Transports: travel, pleasure, and imaginative geography, 1600–1830, volume 3 of Studies in British art. Published for the Paul Mellon Centre for Studies in British Art [and] the Yale Center for British Art [by] Yale University Press, New Haven, 1996. ISBN 0-300-06382-2 (hardcover). viii + 341 pp. LCCN PN56.T7 T67 1996.

Clarke:1982:DPS

[Cla82] Desmond M. Clarke. Descartes’ philosophy of science. Studies in intellectual history. Manchester University Press, Manchester, UK, 1982. ISBN 0-7190-0868-9. xii + 249 pp. LCCN Q175 .C58 1982. US$19.00.

Clarke:1989:OPH

[Cla89] Desmond M. Clarke. Occult powers and hypotheses: Cartesian natural philosophy under Louis XIV. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-824812-1. 265 pp. LCCN Q174.8 .C57 1989. US$30.00. URL http://www.

loc.gov/catdir/
enhancements/fy0637/
88031260-d.html.

Clagett:1992:AES

[Cla92a] Marshall Clagett, editor. Ancient Egyptian science: a source book, volume 184 of Memoirs of the American Philosophical Society, 0065-9738. American Philosophical Society, Philadelphia, PA, USA, 1992. ISBN 0-87169-184-1 (1), 0-87169-214-7 (2), 0-87169-232-5 (3). bd. pp. LCCN ????

Clark:1992:EOE

[Cla92b]  J.  F.  McDi-

armid Clark. Eleanor Ormerod (1828–1901) as an economic entomologist: ‘pioneer of purity even more than of Paris Green’. British Journal for the History of Science, 25(4): 431–452, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027048.

Clarke:1992:BRE

[Cla92c] Desmond M. Clarke. Book review: Emily R. Grosholz. Cartesian Method and the Problem of Reduction. Oxford: Clarendon Press, 1991. Pp. ix + 161. ISBN 0-19-824250-6. £22.50. British Journal for the History of Science, 25(2): 266–267, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027309.

Clagett:1995:AES

[Cla95a]  Marshall Clagett.  Ancient

Egyptian science. Volume II. Calendars, clocks, and astronomy: a source book, volume 214 of Memoirs of the American Philosophical Society, 0065-9738. American Philosophical Society, Philadelphia, PA, USA, 1995. ISBN 0-87169-214-7. xiv + 575 + 144 pp. LCCN Q11 .P612 vol. 184; QB20 .C62 1995.

Clagett:1995:CCA

[Cla95b] Marshall Clagett, editor. Calendars, clocks, and astronomy, volume Vol. 2; 214 of Ancient Egyptian science: a source book / by Marshall Clagett; Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1995. ISBN 0-87169-214-7. xiv + 575 + 70 pp. LCCN ????

Clark:1997:AWD

[Cla97] J. F. M. Clark. ‘the ants were duly visited’: making sense of John Lubbock, scientific naturalism and the senses of social insects. British Journal for the History of Science, 30 (2):151–176, June 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027715.

Clagett:1989:AES

[Cla99a]  Marshall Clagett.  Ancient

Egyptian science: a source book, volume 184, 214, 232 of Memoirs of the American philosophical society. American Philosophical Society, Philadelphia, PA, USA, 1989–1999. ISBN 0-87169-184-1 (vol. 1), 0-87169-214-7 (vol. 2), 0-87169-232-5 (vol. 3). ISSN 0065-9738. various pp. LCCN Q11 .P612 vol. 184, etc.

Clay:1999:RBJ

[Cla99b] John Clay. Robert Boyle: a Jungian perspective. British Journal for the History of Science, 32(3):285–298, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027987.

Cleevely:1983:WPC

[Cle83] R. J. Cleevely. World palaeontological collections. British Museum (Natural History), London, UK, 1983. ISBN 0-565-00850-1, 0-7201-1655-4 (Mansell). 365 pp. LCCN QE716.G72 L633 1983.

Clericuzio:1992:BRF

[Cle92] Antonio Clericuzio. Book review: F. Abbri and F. Crispini (eds.). Atti del IIIe Convegno Nazionale di Storia e Fondamenti della Chimica. Cosenza: Brenner editore, 1991. Pp. 402. L. 35,000. British Journal for the History of Science, 25(3): 372–373, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027272.

Clericuzio:1993:VHB

[Cle93] Antonio Clericuzio. From van Helmont to Boyle. A study of the transmission of Helmontian chemical and medical theories in Seventeenth-Century England. British Journal for the History of Science, 26(3):303– 334, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027400.

Charmasson:1987:LTR

[CLR87] Thérèse Charmasson, Anne-Marie Lelorrain, and Yannick Ripa. L’Enseignement technique de la Révolution à nos jours: textes officiels avec introduction, notes et annexes. Tome I, De la Révolution à 1926. Institut national de recherche pédagogique et Economica, Paris, France, 1987. ISBN 2-7342-0146-1 (INRP: vol. 1), 2-7178-1323-3 (Economica: vol. 1). 783 pp. LCCN LB1028.3. ff 350ff.350.

Clulee:1988:JDN

[Clu88] Nicholas H. Clulee. John Dee’s natural philosophy: between science and religion. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1988. ISBN 0-415-00625-2, 0-415-03122-2 (paperback). xiv + 347 + 24 pp. LCCN BF1598.D5 C58 1988.

Cohen:1990:FDF

[CM90] Yves Cohen and Klaus Manfrass, editors. Frankreich und Deutschland: Forschung, Technologie und industrielle Entwicklung im 19. und 20. Jahrhundert: internationales Kolloquium, veranstaltet vom Deutschen Historischen Institut Paris in Verbindung mit dem Deutschen Museum München und der Cité des sciences et de l’industrie, Paris: München, 12.–15. Oktober 1987. (German) [ France and Germany: Research, technology and industrial development in the 19th and 20th Century: international conference organized by the German Historical Institute, Paris and the City of Science and Industry, Paris, Munich, 12–15 October 1987]. In Kommission bei C.H. Beck, München, Germany, 1990. ISBN 3-406-35092-5. xl + 491 pp. LCCN HC280.T4 F73 1990.

Ceard:1993:HDC

[CM93] Jean Céard and Paul Mengal, editors. Histoire du concept de récapitulation: ontogenèse et phylogenèse en biologie et sciences humaines. Histoire des sciences. Masson Editeur, Masson, France, 1993. ISBN 2-225-84057-1. 132 pp. LCCN ????

Challes:1982:DRP

[CMDM82] Robert Challes, Nicolas Malebranche, Frédéric Deloffre, and Melâhat Menemencioglu, editors. Difficultés sur la religion proposées au père Malebranche. (French) [Difficulties in the religion proposed to Father Malebranche], volume 209 of Studies on Voltaire and the eighteenth century, 0435-2866. The Voltaire Foundation at the Taylor Institution, Oxford, UK, 1982. ISBN 0-7294-0289-4. 576 pp. LCCN PQ2105.A2 S8 vol. 209.

Clark:1995:EWH

[CMN95] Harold W. (Harold Willard) Clark, Frank Lewis Marsh, and Ronald L. Numbers. The early writings of Harold W. Clark and Frank Lewis Marsh, volume 8 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1809-X. xxiii + 531 pp. LCCN BS651 .C54 1995.

Conway:1992:CLC

[CMNH92] Anne Conway, Henry More, Marjorie Hope Nicolson, and Sarah Hutton. The Conway letters: the correspondence of Anne, Viscountess Conway, Henry More, and their friends, 1642–1684. Clarendon Press, New York, NY, USA, revised edition, 1992. ISBN 0-19-824876-8. xxix + 592 pp. LCCN DA378.C6 A4 1992. US$50.00. URL http://www.

loc.gov/catdir/
enhancements/fy0605/
91038838-d.html.

Cohen:1984:QMS

[Coh84] H. F. Cohen. Quantifying music: the science of music at the first stage of the scientific revolution, 1580–1650, volume 23 of The University of Western Ontario series in philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1637-4. xvii + 308 pp. LCCN ML3807 .C63 1984.

Cohen:1985:RS

[Coh85] I. Bernard Cohen. Revolution in science. Belknap Press of Harvard University Press, Cambridge, MA, USA and London, UK, 1985. ISBN 0-674-76777-2. xx + 711 + 1 pp. LCCN Q125 .C542 1985.

Cohen:1991:HCH

[Coh91] H. Floris Cohen. How Christiaan Huygens mathematized nature. British Journal for the History of Science, 24(1): 79–84, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027017.

Cohen:1994:SRH

[Coh94] H. F. Cohen. The scientific revolution: a historiographical inquiry. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-11279-9, 0-226-11280-2 (paperback). xviii + 662 pp. LCCN Q125 .C538 1994. URL http://www.

loc.gov/catdir/
description/uchi052/
93041784.html;   http://
www.loc.gov/catdir/toc/
uchi051/93041784.html.

Cohen:1996:EHC

[Coh96] Clive Cohen. The early history of chemical engineering: a reassessment. British Journal for the History of Science, 29 (2):171–194, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027832.

Coleman:1987:YFN

[Col87]  William Coleman.  Yellow

fever in the North: the methods of early epidemiology, volume 6 of Wisconsin publications in the history of science and medicine. University of Wisconsin Press, Madison, WI, USA, 1987. ISBN 0-299-11110-5, 0-299-11114-8 (paperback). xv + 202 pp. LCCN RA649 .C64 1987. US$45.00, US$22.95.

Coley:1993:BRP

[Col93] Noel Coley. Book review: P. W.

Hammond and Harold Egan, Weighed in the Balance: A History of the Laboratory of the Government Chemist. London: HMSO, 1992. Pp. xvi + 372, illus. ISBN 0-11-515302-0. £9.95. British Journal for the History of Science, 26(2): 248–249, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027112.

Coley:1995:RU

[Col95] Noel G. Coley. Review: [untitled]. British Journal for the History of Science, 28(4): 488–489, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027569.

Connor:1987:WME

[Con87] R. D. Connor. The weights and measures of England. Her Majesty’s Stationary Office, London, England, 1987. ISBN 0-11-290435-1. xxvi + 422 + 41 pp. LCCN QC89.G82 E543 1987. US$30.00.

Cooter:1984:CMP

[Coo84] Roger Cooter. The cultural meaning of popular science: phrenology and the organization of consent in Nineteenth-Century Britain. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-22743-7. xiii + 418 pp. LCCN Q127.G4

C66 1984. URL http://www.
loc.gov/catdir/
description/cam031/
84012078.html;   http://
www.loc.gov/catdir/toc/
cam031/84012078.html.

Cooter:1988:SHA

[Coo88] Roger Cooter, editor. Studies in the history of alternative medicine. St. Antony’s/Macmillan series. MacMillan Publishing Company, New York, NY, USA, 1988. ISBN 0-333-46213-0. xx + 180 pp. LCCN R733 .S87 1988b.

Cooper:1989:BAP

[Coo89a] Andrew Fenton Cooper. British agricultural policy, 1912–36: a study in Conservative politics. Manchester University Press, Manchester, UK, 1989. ISBN 0-7190-2886-8. viii + 240 pp. LCCN HD1927 .C66 1989.

Cooter:1989:PBI

[Coo89b] Roger Cooter. Phrenology in the British Isles: an annotated historical biobibliography and index. Scarecrow Press, Metuchen, NJ, USA, 1989. ISBN 0-8108-2165-6. xviii + 431 pp. LCCN Z7204.P47 C66 1989.

Cook:1990:BRF

[Coo90] Harold J. Cook. Book review:

Francisco Sanches. That Nothing is Known (Quod Nihil Scitur), with Introduction, notes and bibliography by Elaine Limbrick, Latin text, annotations, and translation by Douglas F. S. Thomson. Cambridge: Cambridge University Press, 1988. Pp. x + 310. ISBN 0-521-35077-8. £35.00, $49.50. British Journal for the History of Science, 23(3):360– 361, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026782.

Coopersmith:1992:ER

[Coo92] Jonathan Coopersmith. The electrification of Russia, 1880– 1926. Cornell University Press, Ithaca, NY, USA, 1992. ISBN 0-8014-2723-1. xii + 274 pp. LCCN HD9685.S762 C66 1992.

Cook:1996:BRC

[Coo96] Harold J. Cook. Book review: Catherine Wilson, The Invisible World: Early Modern Philosophy and the Invention of the Microscope. Studies in Intellectual History and the History of Philosophy. Princeton: Princeton University Press, 1995. Pp. x + 280. ISBN 0-691-03418-4. £32.00, $39.50. British Journal for the History of Science, 29(4): 480–481, December 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027702.

Copenhaver:1992:HGB

[Cop92] Brian P. Copenhaver, editor. Hermetica: the Greek Corpus Hermeticum and the Latin Asclepius in a new English translation, with notes and introduction. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-36144-3 (hardcover). lxxxiii + 320 pp. LCCN BF1600 .H475

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/91025703.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        91025703.html.

Corsi:1988:ALE

[Cor88a] Pietro Corsi. The age of Lamarck: evolutionary theories in France, 1790–1830. University of California Press, Berkeley, CA, USA, rev. and updat edition, 1988. ISBN 0-520-05830-5. xiii + 360 pp. LCCN QH15 .C6713 1988.

Corsi:1988:SRB

[Cor88b] Pietro Corsi. Science and religion: Baden Powell and the Anglican debate, 1800– 1860. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-24245-2. ix + 346 pp. LCCN BL245 .C67

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87013845.html;

http://www.loc.gov/
catdir/samples/cam031/
87013845.html;   http://
www.loc.gov/catdir/toc/
cam021/87013845.html.

Cormack:1993:BRD

[Cor93] Lesley B. Cormack. Book review: David N. Livingstone, The Geographical Tradition. Oxford: Blackwell, 1992. Pp. viii + 434. ISBN 0-631-18535-6, £45.00 (hardcover); 0-631-18536-0, £13.95 (paperback). British Journal for the History of Science, 26(4):485– 486, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027468.

Corry:1996:MAR

[Cor96]  Leo Corry.  Modern alge-

bra and the rise of mathematical structures, volume 17 of Science networks. Historical studies. Birkhäuser Verlag, Basel, Switzerland, 1996. ISBN 3-7643-5311-2 (Basel), 0-8176-5311-2 (Boston). 460 pp. LCCN QA151 .C67 1996.

Correia:1997:OEE

[Cor97] Clara Pinto Correia. The ovary of Eve: egg and sperm and preformation. University of Chicago Press, Chicago, IL, USA, 1997. ISBN 0-226-66952-1. xxiii + 396 pp. LCCN Q127.E8 C67 1997. URL http:

//www.loc.gov/catdir/
description/uchi052/

                                        97014011.html;   http://
                                        www.loc.gov/catdir/toc/
                                        uchi051/97014011.html.

Coudert:1995:LK

[Cou95] Allison Coudert. Leibniz and the Kabbalah, volume 142 of Archives internationales d’histoire des idées = International archives of the history of ideas. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3114-1. xvii + 218 pp. LCCN B2599.C33 C68

                                        1995. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0821/94032967-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0821/94032967-t.html.

Courvisanos:1996:ICC

[Cou96] Jerry Courvisanos. Investment cycles in capitalist economies: a Kaleckian behavioural contribution. New directions in modern economics. Edward Elgar Pub., Cheltenham, UK, 1996. ISBN 1-85898-410-6. xi + 272 pp. LCCN HB3720 .C68 1996.

Collins:1993:GWE

[CP93] H. M. (Harry M.) Collins and T. J. (Trevor J.) Pinch. The golem: what everyone should know about science. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-35601-6 (hardcover). xii + 164 pp. LCCN Q125 .C552 1993.

Crawford:1984:BNI

[Cra84] Elisabeth T. Crawford. The beginnings of the Nobel institution: the science prizes, 1901– 1915. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26584-3. ix + 281 + 6 pp. LCCN QC28 .C73

1984. URL http://www.loc.
gov/catdir/description/
cam022/84005844.html;
http://www.loc.gov/
catdir/toc/cam026/
84005844.html.

Cranefield:1992:BRB

[Cra92a] Paul F. Cranefield. Book review: E. B. Adams. In Search of Truth: A Portrait of Don Craib. London: Royal Society of Medicine Services, 1989. Pp. xii + 123, illus. ISBN 1-85315-119-X, £12.95 (hardcover); 1-85315-118-1, £7.95 (paperback). British Journal for the History of Science, 25 (2):285–286, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027323.

Crawford:1992:NIS

[Cra92b] Elisabeth Crawford. Nationalism and internationalism in science, 1880–1939: four studies of the Nobel population. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40386-3 (hardcover). xii + 157 pp. LCCN Q126.9 .C73

Crawford:1992:BRJ

[Cra92c] Elspeth Crawford. Book review: John Roche (ed.). Physicists Look Back: Studies in the History of Physics. Bristol and New York: Adam Hilger, 1990. Pp. xii + 393. ISBN 0-85274-001-8. British Journal for the History of Science, 25 (2):290–291, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027328.

Crabtree:1993:MFM

[Cra93] Adam Crabtree. From Mesmer to Freud: magnetic sleep and the roots of psychological healing. Yale University Press, New Haven, CT, USA, 1993. ISBN 0-300-05588-9. x + 413 pp. LCCN RC497 .C73 1993.

Crawford:1996:AIT

[Cra96] Elisabeth T. Crawford. Arrhenius: from ionic theory to the greenhouse effect, volume 23 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1996. ISBN 0-88135-166-0. xiii + 320 + 9 pp. LCCN Q143.A77 C68 1996.

Creese:1991:BWN

[Cre91a] Mary R. S. Creese. British women of the nineteenth and early twentieth centuries who contributed to research in the chemical sciences. British Journal for the History of Science, 24 (3):275–305, September 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027231.

Crepel:1991:BRP

[Cre91b]  Pierre Crepel.  Book re-

view: Paola Dessi. L’ordine e il caso: Discussione epistemologiche e logiche sulla probabilità de Laplace a Pierce. Bologna: II Mulino, 1989. Pp.

239. ISBN 88-15-02088-8. L. 26,000. British Journal for the History of Science, 24(2): 264–265, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027186.

Crevier:1993:ATH

[Cre93] Daniel Crevier. AI: the tumultuous history of the search for artificial intelligence. Basic Books, New York, NY, USA, 1993. ISBN 0-465-02997-3. xiv + 386 pp. LCCN Q335 .C66 1993. US$27.50, CAN$37.95.

Crilly:1990:BRJ

[Cri90] Tony Crilly. Book review:

Joan L. Richards. Mathematical Visions — The Pursuit of Geometry in Victorian England. New York: Academic Press, 1988. Pp. xiii + 266. ISBN 0-12-587445-6. $34.95. British Journal for the History of Science, 23(3):338–340, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026763.

Critchleys:1994:RAY

[Cri94] Critchleys. Report of the auditors for the year 1993. British Journal for the History of Science, 27(4):508–512, December 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Critchleys:1996:RAY

[Cri96] Critchleys. Report of the auditors for the year 1995. British Journal for the History of Science, 29(4):506–507, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Crook:1984:BKP

[Cro84] D. P. (David Paul) Crook. Benjamin Kidd: portrait of a social Darwinist. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25804-9. vii + 460 pp. LCCN HM106.K52 C76

                                        1984. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0903/83019009-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0903/83019009-t.html.

Crowe:1986:ELD

[Cro86] Michael J. Crowe. The extraterrestrial life debate, 1750– 1900: the idea of a plurality of worlds from Kant to Lowell. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-26305-0. xix + 680 pp. LCCN QB54 .C76

1986. URL http://www.loc.
gov/catdir/description/
cam032/85007842.html.

Crombie:1990:SOM

[Cro90a] A. C. (Alistair Cameron)

Crombie. Science, optics, and music in medieval and early modern thought. Hambledon Press, London [England], 1990. ISBN 0-907628-79-6. xxii + 474 pp. LCCN Q125 .C686 1990. US$55.00. URL http://www.

loc.gov/catdir/
description/hol032/
90030933.html.

Crowe:1990:TWA

[Cro90b] Michael J. Crowe. Theories of the world from antiquity to the Copernican revolution. Dover Publications, Inc., New York, NY, USA, 1990. ISBN 0-486-26173-5. ix + 229 pp. LCCN QB15 .C77 1990. US$5.95.

Cronin:1991:APA

[Cro91a] Helena Cronin. The ant and the peacock: altruism and sexual selection from Darwin to today. Press Syndicate of the University of Cambridge, Cambridge, UK, 1991. ISBN 0-521-32937-X (hardcover). xiv + 490 pp. LCCN QL761 .C76

1991. URL http://www.loc.
gov/catdir/description/
cam032/91007887.html;
http://www.loc.gov/

catdir/toc/cam032/
91007887.html.

Crowcroft:1991:EEH

[Cro91b] Peter Crowcroft. Elton’s ecologists: a history of the Bureau of Animal Population. University of Chicago Press, Chicago, IL, USA, 1991. ISBN 0-226-12146-1, 0-226-12148-8 (paperback). xx + 177 pp. LCCN QL69.G72 U553

1991. URL http://www.loc.
gov/catdir/description/
uchi052/90039469.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/90039469-t.html.

Crosland:1992:SUC

[Cro92] Maurice P. Crosland. Science under control: the French Academy of Sciences, 1795– 1914. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41373-7. xvii + 454 pp. LCCN Q46 .C77

1992. URL http://www.loc.
gov/catdir/description/
cam024/91028748.html;
http://www.loc.gov/
catdir/samples/cam031/
91028748.html;   http://
www.loc.gov/catdir/toc/
cam028/91028748.html.

Crosland:1994:BRJ

[Cro94] Maurice Crosland. Book review: Jean-Pierre Poirier, Antoine Laurent de Lavoisier 1743–1794. Paris: Pygmalion/ Gérard Watelet, 1993. Pp. xii + 545. ISBN 2-85704-384-8. 178FF. British Journal for the History of Science, 27(1): 118, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027588.

Crosland:1995:BRA

[Cro95] Maurice Crosland. Book review: Arthur Donovan, Antoine Lavoisier: Science, Administration and Revolution. Blackwell Science Biographies. Oxford: Blackwell, 1993. Pp. xv + 351. ISBN 0-631-17887-2. £35.00, $29.95. British Journal for the History of Science, 28 (1):111–112, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027494.

Crombie:1996:SAN

[Cro96a] A. C. (Alistair Cameron)

Crombie. Science, art, and nature in medieval and modern thought. Hambledon Press, London, UK, 1996. ISBN 1-85285-067-1 (hardcover). xv + 516 pp. LCCN Q127.E8 C74

1996. URL http://www.loc.
gov/catdir/description/
hol032/95049964.html.

Crosland:1996:BRB

[Cro96b] Maurice Crosland. Book review: Bernadette Bensaude-Vincent and Ferdinando Abbri (eds.), Lavoisier in European Context: Negotiating a New Language for Chemistry. Nantucket: Science History Publications, 1995. Pp. vii + 303. ISBN 0-88135-189-X. $45.95. British Journal for the History of Science, 29 (2):238, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027843.

Crosland:1996:BRE

[Cro96c] Maurice Crosland. Book review: Emmanuel Grison, Michelle Goupil and Patrice Bret (eds.), A Scientific Correspondence during the Chemical Revolution: Louis-Bernard Guyton de Morveau and Richard Kirwan, 1782– 1802. Berkeley Papers in History of Science, 17. Berkeley: Office for History of Science and Technology, University of California at Berkeley, 1995. Pp. vi + 257. ISBN 0-918102-21-9. $10.00. British Journal for the History of Science, 29 (1):98–99, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027521.

Crosland:1996:BRR

[Cro96d] Maurice Crosland. Book review: Revue d’histoire des sciences, Volume 48, Parts 1 and 2: Débats et chantiers actuels autour de Lavoisier et la révolution chimique. Paris: Presses Universitaires de France, 1995. Pp. 230. ISBN 2-13-046972-8. ISSN 0151-4105. Subscription: 510 F per annum outside France. British Journal for the History of Science, 29(4):481–482, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027703.

Crump:1990:AN

[Cru90] Thomas Crump. The anthropology of numbers, volume 70 of Cambridge studies in social anthropology. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-38045-6. x + 201 pp. LCCN GR933 .C78

1990. URL http://www.loc.
gov/catdir/description/
cam032/89034123.html;
http://www.loc.gov/
catdir/toc/cam031/
89034123.html.

Chertok:1992:CPR

[CS92] Léon Chertok and Isabelle Stengers. A critique of psychoanalytic reason: hypnosis as a scientific problem from Lavoisier to Lacan. Stanford University Press, Stanford, CA, USA, 1992. ISBN 0-8047-1950-0. xxvi + 319 pp. LCCN RC497 .C5313 1992. US$35.00. URL http://www.

loc.gov/catdir/
description/cam024/
91022388.html;   http://
www.loc.gov/catdir/toc/
cam027/91022388.html.

Chambers:1994:VNH

[CS94]  Robert  Cham-

bers and James A. Secord. Vestiges of the natural history of creation and other evolutionary writings. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-10072-3, 0-226-10073-1 (paperback). xlviii, vi, 390, 254 pp. LCCN QH363 .C4

                                        1994. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi051/93048415.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0608/93048415-t.html.

Cremo:1993:FAH

[CT93]  Michael  A.

Cremo and Richard L. Thompson. Forbidden archeology: the hidden history of the human race. Bhaktivedanta Institute, San Diego, 1993. ISBN 0-9635309-8-4. xxxviii + 914 pp. LCCN GN741 .C74 1993. US$39.95 (US$49.95 Can.).

Cunningham:1997:ARR

[Cun97] Andrew Cunningham. The anatomical renaissance: the resurrection of the anatomical projects of the ancients. Scolar Press, Aldershot, England, 1997. ISBN 1-85928-338-1. xiv + 283 pp. LCCN QM11 .C86 1997.

Cuomo:1995:BRR

[Cuo95] Serafina Cuomo. Book review: R. W. Hadden, On the Shoulders of Merchants: Exchange and the Mathematical Conception of Nature in Early Modern Europe. Albany: State University of New York, 1994. Pp. xviii + 191. ISBN 0-7914-2011-6. $14.95. British Journal for the History of Science, 28(4): 463–464, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027550.

Cattermole:1987:HDS

[CW87] M. J. G. (Michael J. G.) Cattermole and A. F. (Arthur F.) Wolfe. Horace Darwin’s shop: a history of the Cambridge Scientific Instrument Company, 1878 to 1968. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-569-9. xvi + 285 pp. LCCN Q185 .C365 1987. US$29.50. URL http://www.

loc.gov/catdir/
enhancements/fy0745/
87121867-d.html.

Cunningham:1993:CBP

[CW93] Andrew Cunningham and Perry Williams. De-centring the ‘big picture’: The origins of modern science and the modern origins of science. British Journal for the History of Science, 26(4):407–432, December 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027464.

Dahl:1997:FCR

[Dah97] Per F. Dahl. Flash of the cathode rays: a history of J. J. Thomson’s electron. Institute of Physics Publishers, Bristol, UK, 1997. ISBN 0-7503-0453-7 (hardcover), 0-585-20853-0 (ebook). xvii + 526 pp. LCCN 99.E01389.

Danziger:1990:CSH

[Dan90] Kurt Danziger. Constructing the subject: historical origins of psychological research. Cambridge Studies in the History of Psychology. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-36358-6 (hardcover), 0-521-46785-3 (paperback). vii + 254 pp. LCCN BF181 .D27 1990.

Danziger:1992:BRI

[Dan92] Kurt Danziger. Book review: Ian Hacking. The Taming of Chance. Cambridge: Cambridge University Press, 1990. Pp. xiii + 264. ISBN 0-521-38014-6, £27.50, $44.50 (hardcover); 0-521-38884-8, £9.95, $14.95 (paperback). British Journal for the History of Science, 25(3): 371–372, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027270.

Daston:1988:CPE

[Das88] Lorraine Daston. Classical probability in the Enlightenment. Princeton University Press, Princeton, NJ, USA, 1988. ISBN 0-691-08497-1. xviii + 423 pp. LCCN QA273.A4 D37 1988. US$55.00. URL

                                        http://www.loc.gov/
                                        catdir/description/
                                        prin031/88009869.html.

Daum:1998:WIJ

[Dau98]  Andreas  W.

Daum. Wissenschaftspopularisierung im 19. Jahrhundert: bürgerliche Kultur, naturwissenschaftliche Bildung und die deutsche Öffentlichkeit, 1848– 1914. (German) [Science popularization in the 19th Century: civil culture, scientific education and the German public]. R. Oldenbourg, München, Germany, 1998. ISBN 3-486-56337-8. xii + 617 pp. LCCN QH149 .D39 1998.

URL http://www.h-net.org/
review/hrev-a0b0s0-aa.

Davies:1994:BRS

[Dav94] Gordon L. Herries Davies.

Book review: Les Scientifiques et la montagne. Actes du 116e Congrès National des Sociétés Savantes, Section d’Histoire des Sciences et des Techniques. Paris: Ministère de l’Education Nationale et de la Culture; Comité des Travaux Historiques et Scientifiques, 1993. Pp. 296. ISBN 2-7355-0272-4. FF 300.00. British Journal for the History of Science, 27 (4):487, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027638.

Davis:1995:KT

[Dav95] A. E. L. Davis. Kepler translated. British Journal for the History of Science, 28(3):345– 349, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027649.

Davies:1997:OJN

[Dav97]  Mansel Davies.  Obit-

uary: Joseph Needham (1900– 1995). British Journal for the History of Science, 30(1):95– 100, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027904.

Davis:1995:SMS

[Dav99] E. A. (Edward Arthur) Davis, editor. Science in the making: scientific development as chronicled by historic papers in the Philosophical Magazine, with commentaries and illustrations. Taylor and Francis, Boca Raton, FL, USA, 1995–1999. ISBN 0-7484-0219-5 (vol. 1), 0-7484-0642-5 (vol. 2), 0-7484-0766-9 (vol. 3), 0-7484-0767-7 (vol. 4). ???? pp. LCCN Q125 .S39

                                        1995. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0745/95201153-d.html.

Dawbarn:1998:BRF

[Daw98a] Frances Dawbarn. Book review: Four Treatises of Theophrastus Von Hohenheim Called Paracelsus, translated from the original German, with introductory essays by C. Lilian Temkin, George Rosen, Gregory Zilboorg and Henry E. Sigerist. Edited, with a preface by Henry E. Sigerist, Baltimore: Johns Hopkins University Press, 1996. ISBN 0-8018-5523-3. £13.00, $15.95 (paperback edition). British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027780.

Dawbarn:1998:PPC

[Daw98b] Frances Dawbarn. Patronage and power: the College of Physicians and the Jacobean court. British Journal for the History of Science, 31(1): 1–19, March 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027774.

Day:1999:OJA

[Day99]  Lance Day.  Obitu-

ary: John Anthony Chaldecott (1916–1998). British Journal for the History of Science, 32 (3):343–344, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027991.

deChamps:2009:BFF

[dCC09] Emmanuelle de Champs and Jean-Pierre Cléro, editors. Bentham et la France: fortune et infortunes de l’utilitarisme. (French) [Bentham and France: fortune and misfortune of utilitarianism], volume 2009:09 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2009. ISBN 0-7294-0980-5. xv + 311 pp. LCCN JC223.B5 B46 2009.

Dubowsky:1994:FMH

[DD94] Nathan Dubowsky and Scott Michael Dubowsky. The final mission of HMS Beagle: clarifying the historical record. British Journal for the History of Science, 27 (1):105–111, March 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027584.

Dirac:1995:CWP

[DD95] P. A. M. (Paul Adrien Maurice) Dirac and R. H. (Richard Henry) Dalitz. The collected works of P. A. M. Dirac, 1924– 1948. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-36231-8. xxiii + 1310 pp. LCCN QC21.2 .D57

                                        1995. URL http://www.loc.
                                        gov/catdir/description/
                                        cam026/95001010.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy1006/95001010-t.html.

Darwin:1990:CDM

[DDG90] Charles Darwin, Mario A. Di Gregorio, and N. W. (Nick W.) Gill, editors. Charles Darwin’s marginalia, volume 783 of Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1990. ISBN 0-8240-6639-1 (vol.

1). ???? pp. LCCN QH365 .A1 1990.

DeLaBruheze:1992:RWR

[De 92] Adri De La Bruheze. Radiological weapons and radioactive waste in the United States: insiders’ and outsiders’ views, 1941—1955. British Journal for the History of Science, 25 (2):207–227, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027299.

DeGandt:1995:FGN

[De 95] François De Gandt. Force and geometry in Newton’s Principia. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03367-6. xiv + 296 pp. LCCN QA803 .G3613 1995. URL http://

www.loc.gov/catdir/
description/prin031/
94039315.html;   http://
www.loc.gov/catdir/toc/
prin031/94039315.html.

DeChadarevian:1996:LSV

[De 96] Soraya De Chadarevian. Laboratory science versus countryhouse experiments. the controversy between Julius Sachs and Charles Darwin. British Journal for the History of Science, 29(1):17–41, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027510.

Dear:1991:LSS

[Dea91] Peter Dear, editor. The literary structure of scientific argument: historical studies. University of Pennsylvania Press, Philadelphia, PA, USA, 1991. ISBN 0-8122-8185-3. vi + 211 pp. LCCN Q126.8 .L58 1991.

Dear:1992:BRM

[Dea92] Peter Dear. Book review: Marcello Pera and William R. Shea (eds). Persuading Science: The Art of Scientific Rhetoric. Canton, MA: Science History Publications, 1991. Pp. xi + 212. ISBN 0-88135-071-0. $39.95. British Journal for the History of Science, 25(3): 387–388, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027289.

Dear:1993:BRM

[Dea93] Peter Dear. Book review: Michael Segre, In the Wake of Galileo. New Brunswick, NJ: Rutgers University Press, 1991. Pp. xix + 192. ISBN 0-8135-1700-1. $27.95. British Journal for the History of Science, 26 (3):355–356, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027407.

Dear:1995:TB

[Dea95]  Peter Dear.  Trust Boyle.

British Journal for the History of Science, 28(4):451– 454, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027547.

Debus:1991:FPC

[Deb91] Allen G. Debus. The French Paracelsians: the chemical challenge to medical and scientific tradition in early modern France. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-40049-X. xvi + 247 pp. LCCN QD18.F8 D43

1991. URL http://www.loc.
gov/catdir/description/
cam024/90026163.html;
http://www.loc.gov/
catdir/samples/cam031/
90026163.html;   http://
www.loc.gov/catdir/toc/
cam028/90026163.html.

Degler:1991:SHN

[Deg91] Carl N. Degler. In search of human nature: the decline and revival of Darwinism in American social thought. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-506380-5, 0-19-507707-5 (paperback). x + 400 pp. LCCN HM; HM106 DEG 1991; 300; 301; HM106.DEG.

Deichmann:1992:BHV

[Dei92]  Ute Deichmann.  Biologen

unter Hitler: Vertreibung, Karrieren, Forschung. (German) [Biologists under Hitler: expulsion, careers, research]. Campus, Frankfurt, Germany, 1992. ISBN 3-593-34763-6. 370 pp. LCCN QH305.2.G3 D45 1992. DM38.00.

Deichmann:1996:GG

[Dei96] Ute Deichmann. Genetics in Germany. British Journal for the History of Science, 29(1): 83–87, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027513.

Dempster:1983:PMN

[Dem83]  W. J. Dempster.  Patrick

Matthew and natural selection: Nineteenth Century gentlemanfarmer, naturalist, and writer. P. Harris, Edinburgh, Scotland, 1983. ISBN 0-86228-065-6. 156 + 4 pp. LCCN QH31.M34 D45 1983. US$8.95.

Desmond:1982:AA

[Des82a] Adrian Desmond. Archetypes and ancestors. Blond and Briggs, London, UK, 1982. ISBN 0-85634-121-5. 287 pp. LCCN ????

Desmond:1982:AAP

[Des82b] Adrian J. Desmond. Archetypes and ancestors: Palaeontology in Victorian London, 1850– 1875. Blond and Briggs, London, UK, 1982. ISBN 0-85634-121-5. 287 pp. LCCN QE705.G7

                                        D4. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/9283761.html.

Desmond:1982:IM

[Des82c] Ray Desmond. The India Museum, 1801–1879. Her Majesty’s Stationary Office, London, England, 1982. ISBN 0-11-580088-3. xv + 215 pp. LCCN AM101.L715 D47 1982. US$25.00.

Desmond:1984:AAP

[Des84] Adrian J. Desmond. Archetypes and ancestors: palaeontology in Victorian London 1850–1875. Blond and Briggs, London, UK, 1984. ISBN 0-85634-121-5. 287 pp. LCCN ????

Desmond:1989:PEM

[Des89a]  Adrian J. Desmond.  The

politics of evolution: morphology, medicine, and reform in radical London. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-14346-5. x + 503 pp. LCCN QL810 .D47

1989. URL http://www.loc.
gov/catdir/description/
uchi052/89005137.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/89005137-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/89005137-b.html.

Dessi:1989:LCD

[Des89b] Paola Dessì. L’ordine e il caso: discussioni epistemologiche e logiche sulla probabilità da Laplace a Peirce. (Italian) [The order and the case: epistemological and logical arguments on probability from Laplace to Peirce]. Ricerca. Mulino, Bologna, Italy, 1989. ISBN 88-15-02088-8. 229 pp. LCCN BC141 .D44 1991.

Desmond:1992:EDI

[Des92]  Ray Desmond.  The Eu-

ropean discovery of the Indian flora. Clarendon Press, New York, NY, USA, 1992. ISBN 0-19-854684-X. xii + 355 pp. LCCN QK358 .D44 1992. US$60.00 (US$120.00

U.S.). URL http://www.loc.
gov/catdir/enhancements/
fy0638/91044110-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0638/91044110-t.html.

Desmond:1994:HDD

[Des94] Adrian J. Desmond. Huxley: the devil’s disciple. M. Joseph, London, UK, 1994. ISBN 0-7181-3641-1. xvii + 475 pp. LCCN Q143.H956 D47 1994.

DeVorkin:1982:HMA

[DeV82] David H. DeVorkin. The history of modern astronomy and astrophysics: a selected, annotated bibliography, volume 1; vol. 304 of Bibliographies of the history of science and technology; Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1982. ISBN 0-8240-9283-X. xxvii + 434 pp. LCCN Z5154.H57 D48 1982. US$65.00.

Davis:1997:JJT

[DF97]  Edward  Arthur

Davis and Isolde J. Falconer. J. J. Thomson and the discovery of the electron. Taylor and Francis, Boca Raton, FL, USA, 1997. ISBN 0-7484-0696-4 (hardcover), 0-7484-0720-0 (paperback). xxvii + 243 pp. LCCN QC16.T45

D38 1997. URL http://www.
gbv.de/dms/bowker/toc/
9780748407200.pdf.

Dee:1998:DJD

[DF98] John Dee and Edward Fenton.

The diaries of John Dee. Day Books, Charlbury, Oxfordshire, UK, 1998. ISBN 0-9532213-0-X. xi + 369 pp. LCCN MLCM 2004/02973 (B).

Desargues:1987:GWG

[DFG87]  Gérard  De-

sargues, Judith Veronica Field, and Jeremy Gray. The geometrical work of Girard Desargues. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96403-7. x + 237 pp. LCCN QA485 .D47 1987. US$64.00.

deGandt:2001:CVI

[dG01] François de Gandt, editor.

Cirey dans la vie intellectuelle: la réception de Newton en France. (French) [Cirey in intellectual life: the reception of Newton in France], volume 2001:11 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2001. ISBN 0-7294-0771-3. v + 253 pp. LCCN B2177 .C57 2001.

deGandt:1989:LTS

[dGB^{+}89] François de Gandt, Armand Beaulieu, et al., editors. L’œuvre de Torricelli: science galiléenne et nouvelle géométrie. (French) [The work of Torricelli: Gallic science and new geometry], volume 1re sér., no 32 (1987) of Publications de la Faculté des lettres et sciences humaines de Nice. Diffusion, Les Belles Lettres, Paris, France, 1989. ISBN 2-252-62032-X. 247 pp. LCCN QA444 .O37 1989.

Davenport-Hines:1990:SDP

[DH90] R. P. T. (Richard Peter Treadwell) Davenport-Hines. Sex, death, and punishment: attitudes to sex and sexuality in Britain since the Renaissance. Collins, London, UK, 1990. ISBN 0-00-217956-3. xv + 439 pp. LCCN HQ18.G78 D38 1990. US$20.00.

Davenport-Hines:1992:GH

[DHS92] R. P. T. (Richard Peter Treadwell) Davenport-Hines and Judy Slinn. Glaxo: a history to 1962. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41539-X. xiv + 406 + 4 pp. LCCN HD9667.9.G53

D38 1992. URL http://www.
loc.gov/catdir/
description/cam024/
91042204.html;   http://

www.loc.gov/catdir/toc/
cam028/91042204.html.

DiGregorio:1992:BRT

[Di 92] Mario A. Di Gregorio. Book review: Thomas Junker. Darwinismus und Botanik. Rezeption, Kritik und theoretische Alternativen im Deutschland des 19. Jahrhunderts, with a preface by Rudolf Schmitz. Stuttgart: Deutscher Apotheker Verlag, 1989. Pp. x + 367. ISBN 3-7692-1225-8. British Journal for the History of Science, 25 (4):483–485, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027071.

DiGregorio:1993:BRA

[Di 93a] Mario A. Di Gregorio. Book review: Antonello La Vergata, Nonostante Malthus, Fecondità popolazioni e armonia della natura, 1700–1900. Turin: Bollati Boringhieri 1991. Pp.

312. ISBN 88-339-0563-3. L. 40,000. British Journal for the History of Science, 26(1): 98–99, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027364.

DiGregorio:1993:BRF

[Di 93b] Mario A. Di Gregorio. Book review: Frederick Gregory, Nature Lost? Natural Science and the German Theological Traditions of the Nineteenth Century. Cambridge, Mass. and London, Harvard University Press, 1992. ISBN 0-674-60483-0. British Journal for the History of Science, 26(1):101– 102, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027367.

DiGregorio:1993:BRG

[Di 93c] Mario A. Di Gregorio. Book review: Giuliano Pancaldi, Darwin in Italy: Science across Cultural Frontiers, updated and expanded edition, translated by Ruey Brodine Morelli, Bloomington and Indianapolis, Indiana University Press, 1991. Pp. xv + 222. ISBN 0-253-34287-2. £22.50. British Journal for the History of Science, 26(1):108–110, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027373.

DiGregorio:1993:BRM

[Di 93d] Mario A. Di Gregorio. Book review: Mary P. Winsor, Reading the Shape of Nature: Comparative Zoology at the Agassiz Museum. Chicago and London: University of Chicago Press, 1991. ISBN 0-226-90215-3. British Journal for the History of Science, 26 (1):115, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027377.

DiGregorio:1993:BRP

[Di 93e] Mario A. Di Gregorio. Book review: Paul J. Weindling, Darwinism and Social Darwinism in Imperial Germany: The Contribution of the Cell Biologist Oscar Hertwig (1849– 1922). Stuttgart, Gustav Fischer, 1991. Pp. 355. ISBN 3-437-11305-4. DM 148. British Journal for the History of Science, 26 (1):110–111, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027374.

DiGregorio:1993:BRJ

[Di 93f] Mario A. Di Gregorio. Book reviews: Jacob W. Gruber and John C. Thackray Richard Owen Commemoration: Three Studies. London: Natural History Museum Publications, 1992. Pp. ix + 181. ISBN 0-565-01109-X. £29.95. Richard Owen The Hunterian Lectures in Comparative Anatomy, May and June 1837, edited with introduction and commentary by P. R. Sloan. London: Natural History Publications, 1992. Pp. xvi + 340. ISBN 0-565-01106-5, £37.50 (hardcover), 0-565-011448, £15.95 (paperback). British Journal for the History of Science, 26(3):365– 366, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027416.

DiGregorio:1995:BRC

[Di 95a]  Mario  Di

Gregorio. Book review: Christiane Groeben (ed.), Karl Ernst von Baer (1792–1876), Anton Dohrn (1840–1909): Correspondence. With an Introduction by Jane M. Oppenheimer. Transactions of the American Philosophical Society, 83. Philadelphia: The American Philosophical Society, 1993. Pp. v + 156. ISBN 0-87169-833-1. $15.00. British Journal for the History of Science, 28(3): 353–354, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027652.

DiGregorio:1995:RU

[Di 95b] Mario A. Di Gregorio. Review: [untitled]. British Journal for the History of Science, 28(4): 475–477, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027561.

DiGrigorio:1998:BRN

[Di 98] Mario Di Grigorio. Book review: Noor Giovanni Mazhar, Catholic Attitudes to Evolution in Nineteenth-Century Italian Literature. Memorie di Scienze Morali, Lettere ed Arti, 60. Venice: Istituto Veneto di Scienze, Lettere ed Arti, 1995. Pp. 284. ISBN 88-86166-24-9. 35,000 lire. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027784.

Dick:1982:PWO

[Dic82] Steven J. Dick. Plurality of worlds: the origins of the extraterrestrial life debate from Democritus to Kant. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24308-4. x + 246 pp. LCCN QB54

.D5. URL http://www.loc.
gov/catdir/enhancements/
fy0803/81010165-d.html.

Dijksterhuis:1990:CSD

[Dij90] E. J. (Eduard Jan) Dijksterhuis. Clio’s stiefkind. (Dutch) [Clio’s stepchild]. B. Bakker, Amsterdam, The Netherlands, 1990. ISBN 90-351-0786-1. 321 pp. LCCN Q125 .D47 1990. fl 45.00.

Dimand:1993:CBM

[Dim93]  Robert W. Dimand.  The

case of Brownian motion: a note on Bachelier’s contribution. British Journal for the History of Science, 26(2):233– 234, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027100.

Divall:1990:BRN

[Div90] Colin Divall. Book review:

Nikolaos A. Peppas (ed.). One Hundred Years of Chemical Engineering. Dordrecht: Kluwer Academic Publishers, 1989. Pp. vii + 414. ISBN 0-7923-0145-5. £59.00, $99.00. British Journal for the History of Science, 23 (3):357–358, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026779.

Divall:1998:BRR

[Div98] Colin Divall. Book review: Robert Fox (ed.), Technological Change: Methods and Themes in the History of Technology. Amsterdam: Harwood Academic Publishers, 1996. Pp. vii + 271. ISBN 3-7186-5792-9. £36.00, $54.00. British Journal for the History of Science, 31 (2):241–250, June 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027766.

David:1988:CCV

[DJS97] Andrew David, Rüdiger Joppien, and Bernard Smith, editors. The charts and coastal views of Captain Cook’s voyages, volume extra ser., no. 43–46 of Hakluyt Society, 0951-6158. Hakluyt Society in association with the Australian Academy of the Humanities, London, UK, 1988–1997. ISBN 0-904180-23-9 (vol. 1), 0-904180-31-X (vol. 2), 0-904180-55-7 (vol. 3). ???? pp. LCCN G1036 .C57 1988.

Darwin:1988:CDB

[DK88] Charles Darwin and R. D.

Keynes. Charles Darwin’s Beagle diary. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-23503-0. xxix + 464 pp. LCCN QH365 .A1

1988. URL http://www.loc.
gov/catdir/description/
cam023/87031169.html;
http://www.loc.gov/
catdir/enhancements/
fy0732/87031169-b.html;
http://www.loc.gov/
catdir/toc/cam025/
87031169.html.

LaMettrie:1996:MMO

[dLT96] Julien Offray de La Mettrie and Ann Thomson. Machine man and other writings. Cambridge texts in the history of philosophy. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-47258-X (hardcover), 0-521-47849-9 (paperback). xxx + 179 pp. LCCN B2062 .E5

1996. URL http://www.loc.
gov/catdir/description/
cam027/95014609.html;
http://www.loc.gov/
catdir/samples/cam034/
95014609.html;   http://
www.loc.gov/catdir/toc/
cam024/95014609.html.

Desmond:1991:D

[DM91] Adrian J. Desmond and James R. (James Richard) Moore. Darwin. Michael Joseph, London, UK, 1991. ISBN 0-7181-3430-3. xxi + 807 pp. LCCN QH31.D2 D47 1991.

Day:1996:BDH

[DM96] Lance Day and Ian McNeil, editors. Biographical dictionary of the history of technology. Routledge reference. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1996. ISBN 0-415-06042-7. xiii + 844 pp. LCCN T39 .B49

                                        1996. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0648/95026250-d.html.

Dobbs:1991:JFG

[Dob91] Betty Jo Teeter Dobbs. The Janus faces of genius: the role of alchemy in Newton’s thought. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-38084-7. xii + 359 pp. LCCN QC16.N7 D64

                                        1991. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/91008695.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        91008695.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam026/91008695.html.

Doel:1996:SSA

[Doe96] Ronald Edmund Doel. Solar system astronomy in America: communities, patronage, and interdisciplinary science, 1920–1960. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-41573-X (hardcover). xiv + 280 pp. LCCN QB501 .D64

1996. URL http://www.loc.
gov/catdir/description/
cam027/95023112.html;
http://www.loc.gov/
catdir/toc/cam023/
95023112.html.

Dolby:1990:BRRa

[Dol90a] R. G. A. Dolby. Book review: R. N. Giere. Explaining Science: A Cognitive Approach. Chicago: Chicago University Press, 1988. Pp. xxi + 321. ISBN 0-226-29205-2. £27.95. British Journal for the History of Science, 23(3): 336–337, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026760.

Dolby:1990:BRRb

[Dol90b] R. G. A. Dolby. Book review: R. Nola (ed.). Relativism and Realism in Science. Dordrecht: Kluwer Academic Publishers, 1988. Pp. x + 299. ISBN 90-277-2647-7. £48.00. British Journal for the History of Science, 23(3): 337, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026761.

Dolan:1994:BRR

[Dol94] Brian P. Dolan. Book review:

Richard Yeo, Defining Science: William Whewell, Natural Knowledge and Public Debate in Early Victorian Britain.

Cambridge: Cambridge University Press, 1993. Pp. xiii + 280. ISBN 0-521-43182-4. £35.00, $54.95. British Journal for the History of Science, 27 (3):374–375, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027613.

Dolan:1995:RUa

[Dol95a] Brian P. Dolan. Review: [untitled]. British Journal for the History of Science, 28(4): 471–472, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027557.

Dolan:1995:RUb

[Dol95b] Brian P. Dolan. Review: [untitled]. British Journal for the History of Science, 28(4): 472–473, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027558.

Dolby:1996:UKI

[Dol96] R. G. A. (Riki G. A.) Dolby. Uncertain knowledge: an image of science for a changing world. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56004-7. xi + 365 pp. LCCN Q175 .D688

                                        1996. URL http://www.loc.
                                        gov/catdir/description/
                                        cam027/96022588.html;
                                        http://www.loc.gov/

catdir/samples/cam031/
96022588.html;   http://
www.loc.gov/catdir/toc/
cam022/96022588.html.

Dolan:1998:PTP

[Dol98a]  Brian Dolan.  Pedagogy

through print: James Sowerby, John Mawe and the problem of colour in early Nineteenth-Century natural history illustration. British Journal for the History of Science, 31(3): 275–304, September 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027814.

Dolan:1998:BRD

[Dol98b] Brian P. Dolan. Book review:

David Philip Miller and Peter Hanns Reill (eds.), Visions of Empire: Voyages, Botany, and Representations of Nature. Cambridge: Cambridge University Press, 1996. Pp. xix + 370. ISBN 0-521-48303-4. £40.00, $59.95. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027781.

Dolan:1999:BRD

[Dol99] Brian P. Dolan. Book review:

David Philip Miller and Peter Hanns Reill (eds.), Visions of Empire: Voyages, Botany, and Representations of Nature. Cambridge: Cambridge University Press, 1996. Pp. xix + 370. ISBN 0-521-48303-4. £45.00. British Journal for the History of Science, 32(3): 363–378, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027994.

Donovan:1988:CRE

[Don88] Arthur L. Donovan, editor. The chemical revolution: essays in reinterpretation, volume 2nd ser., vol. 4 (1988) of Osiris, 0369-7827. History of Science Society, Philadelphia, PA, 1988. ISBN 0-934235-10-4, 0-934235-11-2 (paperback). 236 pp. LCCN Q125 .O76 2nd ser., vol. 4, 1988.

Donnelly:1991:IRC

[Don91] James Donnelly. Industrial recruitment of chemistry students from English universities: a revaluation of its early importance. British Journal for the History of Science, 24 (1):3–20, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027013.

Donovan:1993:ALS

[Don93]  Arthur Donovan.  Antoine

Lavoisier: science, administration, and revolution. Blackwell science biographies. Blackwell, Oxford, UK, 1993. ISBN 0-631-17887-2. xv + 351 pp. LCCN QD22.L4 D58 1993.

Donahue:1994:KIS

[Don94] William H. Donahue. Kepler’s invention of the Second Planetary Law. British Journal for the History of Science, 27 (1):89–102, March 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027582.

Dorn:1991:GS

[Dor91] Harold Dorn. The geography of science. The Johns Hopkins University Press, Baltimore, MD, USA, 1991. ISBN 0-8018-4151-8. xx + 219 pp. LCCN Q125 .D65 1991.

Dutton:1988:WTD

[DPP88]  Diana  Barbara  Dut-

ton, Thomas A. Preston, and Nancy E. Pfund. Worse than the disease: pitfalls of medical progress. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-34023-3. xvi + 528 pp. LCCN RA418.5.M4 D88

1988. URL http://www.loc.
gov/catdir/description/
cam023/87034203.html;
http://www.loc.gov/
catdir/toc/cam023/
87034203.html.

Dauben:1995:ARC

[DR95]  Joseph  Warren

Dauben and Abraham Robinson. Abraham Robinson: the creation of nonstandard analysis: a personal and mathematical odyssey. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03745-0. xix + 559 pp. LCCN QA29.R57

D38 1995. URL http://www.
loc.gov/catdir/
description/prin031/
94032715.html;   http://
www.loc.gov/catdir/toc/
prin031/94032715.html.

Dershowitz:1997:CC

[DR97] Nachum Dershowitz and Edward M. Reingold. Calendrical calculations. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56413-1 (hardcover), 0-521-56474-3 (paperback). xxi + 307 pp. LCCN CE12 .D47 1997. URL http:

//www.loc.gov/catdir/
description/cam028/
96045964.html;   http://
www.loc.gov/catdir/toc/
cam023/96045964.html.

Drake:1981:CES

[Dra81] Stillman Drake. Cause, experiment, and science: a Galilean dialogue, incorporating a new English translation of Galileo’s Bodies that stay atop water, or move in it. University of Chicago Press, Chicago, IL, USA, 1981. ISBN 0-226-16228-1. xxix + 237 pp. LCCN QC147 .D7. US$20.00.

Dreyer:1923:HRA

[DTS^{+}87] J. L. E. (John Louis Emil) Dreyer, H. H. (Herbert Hall) Turner, R. A. (Ralph Allen) Sampson, R. J. (Roger John) Tayler, et al., editors. History of the Royal Astronomical Society. Royal Astronomical Society, London, UK, 1923–1987. ISBN 0-632-02173-X (vol. 1), 0-632-02175-6 (vol 1), 0-632-01791-0 (vol. 2.), 0-632-01792-9 (vol. , 2.). ???? pp. LCCN QB1.R75 D73 1923.

Duchesneau:1982:PLE

[Duc82]  François Duchesneau.  La

physiologie des lumières: empirisme, modèles et théories, volume 95 of Archives internationales d’histoire des idées = International archives of the history of ideas. Martinus Nijhoff, Dordrecht and The Hague, The Netherlands, 1982. ISBN 90-247-2500-3. xxi + 611 pp. LCCN QH331 .D79 1982. fl 180.00 (US$90.00 U.S.).

Duchesneau:1987:GTC

[Duc87] François Duchesneau. Genèse de la théorie cellulaire. (French) [Genesis of cellular theory], volume 1 of Collection Analytiques. Bellarmin, Montréal, PQ, Canada, 1987. ISBN 2-89007-646-6. 388 pp. LCCN QH577 .D83 1987.

Duhem:1991:OSS

[Duh91] Pierre Maurice Marie Duhem.

The Origins of Statics: The Sources of Physical Theory, volume 123 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-5658-7, 94-011-3730-7 (e-book).

ISSN 0068-0346. xxxv + 593 pp. LCCN Q174 .B67 vol.

                                        123. URL http://d-nb.info/
                                        101431545X/34;   https://
                                        link.springer.com/book/
                                        10.1007/978-94-011-3730-0.

Dulac:1988:ED

[Dul88] Georges Dulac, editor. Éditer Diderot, volume 254 of Studies on Voltaire and the eighteenth century, 0435-2866. Voltaire Foundation at the Taylor Institution, Oxford, UK, 1988. ISBN 0-7294-0365-3. 555 + 30 pp. LCCN PQ2105.A2 S8 vol. 254.

Dunstan:1990:HEA

[Dun90] Gordon R. (Gordon Reginald) Dunstan, editor. The human embryo: Aristotle and the Arabic and European traditions. University of Exeter Press, Exeter, Devon, UK, 1990. ISBN 0-85989-340-5. xi + 235 pp. LCCN QM601 .H8 1990.

Duncan:1992:EJA

[Dun92a] A. M. Duncan. Eric John Aiton (1923–1991). British Journal for the History of Science, 25 (2(85)):253–254, 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Duncan:1992:OEJ

[Dun92b] A. M. Duncan. Obituary: Eric John Aiton (1923–1991). British Journal for the History of Science, 25(2):253– 254, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027303.

Durant:1985:DDE

[Dur85] John Durant, editor. Darwinism and divinity: essays on evolution and religious belief. Blackwell, Oxford, UK, 1985. ISBN 0-631-14188-X. ix + 210 pp. LCCN BL263 .D27 1985. US$29.95.

Earman:1989:WES

[Ear89] John Earman. World enough and space-time: absolute versus relational theories of space and time. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-05040-4. xiv + 233 pp. LCCN QC173.59.S65 E17 1989.

Euclid:1996:EBT

[EBT96] Euclid, J. L. Berggren, and Robert S. D. Thomas. Euclid’s Phaenomena: a translation and study of a Hellenistic treatise in spherical astronomy, volume 4 of Sources and studies in the history and philosophy of classical science. Garland Publishing, Inc., New York, NY, USA, 1996. ISBN 0-8153-0493-5. x + 132 pp. LCCN QB144 .E9313 1996.

Engelhardt:1987:SCC

[EC87] H. Tristram (Hugo Tristram)

Engelhardt and Arthur L. Caplan, editors. Scientific controversies: case studies in the resolution and closure of disputes in science and technology. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-25565-1, 0-521-27560-1 (paperback). x + 639 pp. LCCN Q181 .S3826

1987. URL http://www.loc.
gov/catdir/description/
cam022/84028480.html;
http://www.loc.gov/
catdir/toc/cam021/
84028480.html.

Eckert:1992:BRK

[Eck92] Michael Eckert. Book review: Klaus Hentschel. Die Korrespondenz Petzold-Reichenbach: Zur Entwicklung der ‘wissenschaftlichen Philosophie’ in Berlin. Berliner Beiträge zur Geschichte der Naturwissenschaften und der Tecknik,

12. Berlin: Sigma, 1990. Pp.

91. ISBN 3-928068-03-2. DM 18, 90. Christoph Meinel. Karl Friedrich Zöllner und die Wissenschaftskultur der Gründerzeit: Eine Fallstudie zur Genese konservativer Zivilisationskritik. Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik,

13. Berlin: Sigma, 1991. Pp.

91. ISBN 3-928068-04-0. DM 16, 50. British Journal for the History of Science, 25(4): 491–492, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027079.

Euclide:1992:RCR

[ECL92] Euclide, Robertus Castrensis, and Hubertus Lambertus Ludovicus. Robert of Chester’s (?) redaction of Euclid’s Elements, the so-called Adelard II version, volume 8–9 of Science networks historical studies. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 0-8176-2728-6 (Boston), 3-7643-2728-6 (Basel), 0-8176-2658-1 (vol. 1, Boston), 3-7643-2658-1 (vol. 1, Basel), 0-8176-2727-8 (vol. 2, Boston), 3-7643-2727-8 (vol. 2, Basel). 959 pp. LCCN ????

Edgerton:1990:BRC

[Edg90a] David Edgerton. Book review:

Colin Chant (ed.). Science and Technology in Everyday Life. London: Routledge in association with the Open University, 1989. Pp. 391. (hardcover). ISBN 0-415-00037-8. £40.00. (Paperback) 0-415-03557-0. £14.95. British Journal for the History of Science, 23(1): 112–113, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026818.

Edgerton:1990:BRH

[Edg90b] David Edgerton. Book review: Harald Penrose. An Ancient Air: A Biography of John Stringfellow of Chard, the Victorian Aeronautical Pioneer. Washington: Smithsonian Institution Press, 1988. Pp. 183. ISBN 0-87474-752-X. $22.50. British Journal for the History of Science, 23(3): 340, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026764.

Edgerton:1991:BRA

[Edg91a] David Edgerton. Book reviews: Atomic Energy for Military Purposes, with a new foreword by Philip Morrison and an essay by Henry De-Wolf Smyth. Stanford: Stanford University Press, 1990. Pp. xvi + 324. ISBN 0-8047-1721-4, $39.50 (hardcover); 0-8047-1722-2, $12.95 (paperback). Herbert F. York. The Advisors: Oppenheimer, Teller and the Superbomb, with a historical essay by Hans A. Bethe. Stanford: Stanford University Press, 1989. Pp. xiv + 201. ISBN 0-8047-1713-3, $32.50 (hardcover); 0-8047-1714-1, $8.95 (paperback). British Journal for the History of Science, 24 (4):476–477, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027137.

Edgerton:1991:HGG

[Edg91b] Samuel Y. Edgerton. The heritage of Giotto’s geometry: art and science on the eve of the scientific revolution. Cornell University Press, Ithaca, NY, USA, 1991. ISBN 0-8014-2573-5 (hardcover). x + 319 pp. LCCN N7430.5 .E34 1991.

Edgerton:1993:TPT

[Edg93] David Edgerton. Tilting at paper tigers. British Journal for the History of Science, 26 (1):67–75, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027341.

Edgerton:1995:BRM

[Edg95] David Edgerton. Book review: Merrit Roe Smith and Leo Marx (eds.), Does Technology Drive History? The Dilemma of Technological Determinism. Cambridge, MA and London: MIT Press, 1994. Pp. xv + 280. ISBN 0-262-19347-7, £31.50 (hardcover); 0-262-69167-1, £14.95 (paperback). British Journal for the History of Science, 28(3):370– 372, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027664.

Edwards:1987:PAT

[Edw87]  A.  W.

F. (Anthony William Fairbank) Edwards. Pascal’s arithmetical triangle. C. Griffin, London, UK, 1987. ISBN 0-85264-283-0, 0-19-520546-4 (Oxford). xii + 174 pp. LCCN QA241 .E37 1987. US$35.00. URL http:

//www.loc.gov/catdir/
enhancements/fy0635/
86019264-d.html;
http://www.loc.gov/

                                        catdir/enhancements/
                                        fy0635/86019264-t.html.

Edwards:1996:CWC

[Edw96] Paul N. Edwards. The closed world: computers and the politics of discourse in Cold War America. Inside technology. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-05051-X. xx + 440 pp. LCCN QA76.17 .E34 1996.

                                        URL http://www.h-net.org/
                                        review/hrev-a0a5x0-aa.

Emery:1995:HGD

[EE95] Alan E. H. Emery and Marcia L. H. Emery. The history of a genetic disease: Duchenne muscular dystrophy or Meryon’s disease. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1995. ISBN 1-85315-249-8, 1-85315-249-3. xvi + 248 pp. LCCN RC935.M7 E46 1995.

Eisenstein:1983:PRE

[Eis83] Elizabeth L. Eisenstein. The printing revolution in early modern Europe. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25858-8, 0-521-27735-3 (paperback). xiv + 297 pp. LCCN Z124 .E374 1983.

Eisenstein:1993:PRE

[Eis93] Elizabeth L. Eisenstein. The printing revolution in early modern Europe. Cambridge University Press, Cambridge, UK, canto edition, 1993. ISBN 0-521-25858-8 (hardcover), 0-521-27735-3 (paperback), 0-521-44770-4 (Canto paperback). xiv + 300 pp. LCCN Z124 .E374 1993.

US$7.95. URL http://www.
loc.gov/catdir/
description/cam026/
94153513.html;   http://
www.loc.gov/catdir/toc/
cam029/94153513.html.

Euclid:1992:MLT

[EiT92] Euclid and Ken ichi Takahashi. The medieval Latin traditions of Euclid’s Catoptrica: a critical edition of De speculis with an introduction, English translation, and commentary. Kyushu University Press, Higashi-ku, Fukuoka-shi, Japan, 1992. ISBN 4-87378-299-6. viii + 373 pp. LCCN QC353 .E813 1992. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/
26940678.html.

Edison:1989:PTEa

[EJ^{+}07] Thomas A. (Thomas Alva) Edison, Reese Jenkins, et al., editors. The papers of Thomas A. Edison. The Johns Hopkins University Press, Baltimore, MD, USA, 1989–2007. ISBN 0-8018-3100-8 (vol. 1), 0-8018-3101-6 (vol. 2), 0-8018-3102-4 (vol. 3), 0-8018-5819-4 (vol. 4), 0-8018-3104-0 (vol. 5), 0-8018-8640-6 (vol. 6). ???? pp. LCCN TK140.E3 A2

1989. URL http://www.loc.
gov/catdir/bios/jhu051/
88009017.html;

                                        http://www.loc.gov/
                                        catdir/description/
                                        jhu051/88009017.html.

Edison:1994:MPE

[EJR94] Thomas Alva Edison, Reese V. Jenkins, and Robert A. Rosenberg. Menlo Park: the early years: April 1876–December 1877, volume Vol. 3 of The papers of Thomas A. Edison / Reese V. Jenkins, dir. and ed. The Johns Hopkins University Press, Baltimore, MD, USA, 1994. ISBN 0-8018-3102-4. xlvii + 727 pp. LCCN ????

Edison:1995:PTE

[EJR95] Thomas Alva Edison, Reese V. Jenkins, and Robert A. Rosenberg. The papers of Thomas A. Edison. Vol. 3, Menlo Park: the early years, April 1876– December 1877. The Johns Hopkins University Press, Baltimore, MD, USA, 1995. ISBN 0-8018-3102-4. xlvii + 727 pp. LCCN ????

Elmer:1991:BRJ

[Elm91] Peter Elmer. Book review: Julian Roberts and Andrew G. Watson (eds.). John Dee’s Library Catalogue. London: The Bibliographical Society, 1990. Pp. viii + 253. ISBN 0-19-721795-8. £60.00. British Journal for the History of Science, 24(2):248–249, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027174.

Elmer:1997:BRM

[Elm97] Peter Elmer. Book review:

Michael Hunter, Science and the Shape of Orthodoxy: Intellectual Change in Late Seventeenth-Century Britain. Woodbridge, Suffolk: Boydell Press, 1995. Pp. xii + 345. ISBN 0-85115-594-4. £55.00, $89.00. British Journal for the History of Science, 30(2):233–249, June 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027721.

Edgeworth:1994:ECE

[EM94] Francis Ysidro Edgeworth and Philip Mirowski. Edgeworth on chance, economic hazard, and statistics. The Worldly philosophy. Rowman and Littlefield, Lanham, MD, USA, 1994. ISBN 0-8476-7751-6 (hardcover). vii + 462 pp. LCCN HB615 .E43 1994.

Emerton:1984:SRF

[Eme84] Norma E. Emerton. The scientific reinterpretation of form. Cornell history of science series. Cornell University Press, Ithaca, NY, USA, 1984. ISBN 0-8014-1583-7. 318 pp. LCCN QC173.3 .E47 1984.

Euclid:1993:DE

[EMMS93a] Euclid, Heinrich Menge, George L. McDowell, and Merle A. Sokolik. The data of Euclid. Union Square Press, Baltimore, MD, USA, 1993. ISBN 0-9635924-1-6, 0-9635924-0-8. xxi + 207 pp. LCCN QA444 .E9313 1993.

Euclides:1993:DE

[EMMS93b] Euclides, Heinrich Menge, George L. McDowell, and Merle A. Sokolik, editors. The data of Euclid. Union Square Press, Baltimore, MD, USA, 1993. ISBN 0-9635924-1-6, 0-9635924-0-8. xxi + 207 pp. LCCN ????

Elias:1982:SEH

[EMW82] Norbert Elias, Herminio Martins, and Richard Whitley, editors. Scientific establishments and hierarchies, volume 6 of Sociology of the sciences: a yearbook. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1982. ISBN 90-277-1322-7, 90-277-1323-5 (paperback). xi + 368 pp. LCCN Q180.A1 S345 1982.

Engel:1983:CDR

[Eng83] A. J. (Arthur Jason) Engel. From clergyman to don: the rise of the academic profession in Nineteenth-Century Oxford. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1983. ISBN 0-19-822606-3. xi + 302 pp. LCCN LF519 .E53 1983. US$45.00.

Engelsman:1984:FCO

[Eng84] Steven B. (Steven Boudewijn) Engelsman. Families of curves and the origins of partial differentiation, volume 93 of North-Holland mathematics studies. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984. ISBN 0-444-86897-6 (paperback). ix + 238 pp. LCCN QA565 .E54 1984.

Elena:1990:HSS

[EO90] Alberto Elena and Javier Ordóñez. History of science in Spain: A preliminary survey. British Journal for the History of Science, 23(2):187–196, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026724.

Edison:1993:PTE

[ER93]  Thomas  A.  Edi-

son and Robert A. Rosenberg. The papers of Thomas A. Edison. The Johns Hopkins University Press, Baltimore, MD, USA, 1993. ISBN 0-8018-3102-4. xlvii + 727 pp. LCCN ????

Eriksson:1994:AVO

[Eri94] Gunnar Eriksson. The Atlantic vision: Olaus Rudbeck and Baroque science, volume 19 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1994. ISBN 0-88135-158-X. viii + 196 pp. LCCN Q125.2 .E75 1994. US$27.95.

Ernst:1991:RRN

[Ern91] Germana Ernst. Religione, ragione e natura: ricerche su Tommaso Campanella e il tardo Rinascimento. (Italian) [Religion, reason and nature: research on Tommaso Campanella and the late Renaissance], volume 41 of Collana di filosofia. F. Angeli, Milano, Italy, 1991. ISBN 88-204-6663-5. 288 pp. LCCN B785.C24 E76 1991. L36000.

Eckert:1990:CET

[ES90] Michael Eckert and Helmut Schubert. Crystals, electrons, transistors: from scholar’s study to industrial research. American Institute of Physics, Woodbury, NY, USA, 1990. ISBN 0-88318-622-5. xxii + 241 pp. LCCN Q125 .E36313 1990.

Euclides:1992:MLT

[ET92] Euclides and Ken’ichi Takahashi. The medieval Latin traditions of Euclid’s Catoptrica: a critical edition of De speculis with an introduction, English translation and commentary. Kyushu University Press, Fukuoka-shi, Japan, 1992. ISBN 4-87378-299-6. viii + 373 pp. LCCN ????

Evans:1987:A

[Eva87] J. D. G. (John David Gemmill) Evans. Aristotle. Philosophers in context. Harvester Press, Brighton, Sussex, UK, 1987. ISBN 0-312-00485-0 (US), 0-7108-1042-3. xii + 208 pp. LCCN B485 .E93

1987. URL http://catalog.
hathitrust.org/api/
volumes/oclc/14691677.
html.

Fauvel:1988:LN

[F^{+}88] John Fauvel et al., editors. Let Newton be! Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-853924-X. 272 pp. LCCN QC16.N7 L47 1988. US$29.95.

URL http://www.loc.gov/
catdir/enhancements/
fy0639/88009911-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0725/88009911-b.html.

Fantoli:1996:GCC

[Fan96] Annibale Fantoli. Galileo, for Copernicanism and for the Church, volume 3 of Studi Galileiani. Vatican Observatory Publications, Vatican City, second edition, 1996. ISBN 0-268-01032-3 (paperback). xx + 567 pp. LCCN QB41.G178 F3613 1996.

Fara:1994:BRJ

[Far94]  Patricia  Fara.

Book review: John Brewer and Roy Porter (eds.), Consumption and the World of Goods. London: Routledge, 1993. Pp. xxi + 564. ISBN 0-415-03712-3. £75.00. British Journal for the History of Science, 27(3): 371–373, September 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027611.

Fara:1995:THV

[Far95] Patricia Fara. ‘A treasure of hidden vertues’: the attraction of magnetic marketing. British Journal for the History of Science, 28(1):5–35, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027482.

Fara:1999:BRR

[Far99] Patricia Fara. Book review: Robert Poole, Time’s Alteration: Calendar Reform in Early Modern England. London: UCL Press, 1998. Pp. xix + 243. ISBN 1-85728-622-7. £45.00, $75.00. British Journal for the History of Science, 32 (3):363–378, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027995.

Fauvel:1996:BM

[Fau96a]  John Fauvel.  Back mat-

ter. British Journal for the History of Science, 29(1): 128, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027541.

Fauvel:1996:BRR

[Fau96b] John Fauvel. Book review: R. V. Wallis and P. J. Wallis, Index of British Mathematicians, Part III: 1701–1800. Newcastle upon Tyne: Project for Historical Bibliography (R. V. Wallis, 43 Briarfield Rd, Newcastle NE3 3UH), 1993, Pp. xiv + 176, illus. ISBN 1-871768-01-2. £22.00. British Journal for the History of Science, 29 (1):121–123, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027539.

Ferrier:1982:HBP

[FB82] R. W. (Ronald W.) Ferrier and J. H. Bamberg. The history of the British Petroleum Company. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24647-4 (vol. 1), 0-521-25950-9 (vol. 2), 0-521-25951-7 (vol. 3: hardback), 0-521-78515-4 (vol. 3: paperback). ???? pp. LCCN HD9571.9.B73 F47 1982; ECON.

Fee:1997:MMH

[FB97] Elizabeth Fee and Theodore M.

Brown, editors. Making medical history: the life and times of Henry E. Sigerist. The Johns Hopkins University Press, Baltimore, MD, USA, 1997. ISBN 0-8018-5355-9. xiii + 387 pp. LCCN R133 .M26

1997. URL http://www.loc.
gov/catdir/bios/jhu051/
96008862.html;
http://www.loc.gov/

                                        catdir/description/
                                        jhu051/96008862.html.

Flamsteed:1982:PJF

[FC82] John Flamsteed and Allan Chapman. The preface to John Flamsteed’s Historia coelestis Britannica, or, British catalogue of the heavens (1725), volume 52–1982 of Maritime monographs and reports, 0307-8590. Trustees of the National Maritime Museum, London, UK, 1982. ISBN 0-905555-60-0 (paperback). vi + 222 pp. LCCN VK553 .F5313 1982.

Fantoli:1994:GCC

[FC94] Annibale Fantoli and George V. Coyne, editors. Galileo: for Copernicanism and for the Church, volume 3 of Studi Galileiani. Vatican Observatory Publications, Vatican City, 1994. ISBN 0-268-01029-3. xix + 540 pp. LCCN QB36.G2 F36713

                                        1994. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/31182040.
                                        html.

French:1996:BSI

[FC96] R. K. (Roger Kenneth) French and Andrew Cunningham, Dr. Before science: the invention of the friars’ natural philosophy. Scolar Press, Aldershot, Hants, 1996. ISBN 1-85928-287-3. 298 pp. LCCN B738.N3 F74 1996.

Friedman:1985:EMM

[FD85] Alan J. Friedman and Carol C. Donley. Einstein as myth and muse. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26720-X. xiv + 224 pp. LCCN QC16.E5 F76

1985. URL http://www.loc.
gov/catdir/description/
cam031/84023839.html;
http://www.loc.gov/
catdir/enhancements/
fy0906/84023839-t.html.

Fourcy:1987:HLP

[FD87] A. (Ambroise) Fourcy and Jean G. Dhombres. Histoire de l’École polytechnique. Librairie du bicentenaire de la Révolution française. Belin, Paris, France, 1987. ISBN 2-7011-0640-0. viii + 516, 198 pp. LCCN T173.P39 F68 1987.

FabridePeiresc:1989:LCD

[FDLJ89]  Nicolas  Claude

Fabri de Peiresc, Cassiano Dal Pozzo, Jean-François Lhote, and Danielle Joyal. Lettres à Cassiano dal Pozzo, 1626– 1637. (French) [Letters to Cassiano dal Pozzo, 1626–1637], volume 3 of Amphion. Adosa, Clermont-Ferrand, France, 1989. ISBN 2-86639-079-2. 280 pp. LCCN DC121.8.P4 A4 1989.

Feingold:1990:BNL

[Fei90] Mordechai Feingold, editor.

Before Newton: the life and times of Isaac Barrow. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-30694-9. xi + 380 pp. LCCN QA29.B37 B44

1990. URL http://www.loc.
gov/catdir/description/
cam024/89038508.html;
http://www.loc.gov/
catdir/enhancements/
fy0632/89038508-t.html.

Flamsteed:1995:CJF

[FFLW97]  John  Flam-

steed, Eric G. Forbes, Murdin Lesley, and Frances Willmoth. The correspondence of John Flamsteed, the first astronomer royal. Institute of Physics Publishing, Bristol, UK, 1995– 1997. ISBN 0-7503-0147-3. ???? pp. LCCN QB36.F6 A4

1995. URL http://www.loc.
gov/catdir/enhancements/
fy0668/95047433-d.html.

Fox:1993:ETI

[FG93]  Robert  Fox

and Anna Guagnini, editors. Education, technology, and industrial performance in Europe, 1850–1939. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-38153-3. xiii + 302 pp. LCCN T105 .E33

1993. URL http://www.loc.
gov/catdir/description/
cam025/92023171.html;
http://www.loc.gov/
catdir/toc/cam028/
92023171.html.

Forgan:1996:CSK

[FG96] Sophie Forgan and Graeme Gooday. Constructing South Kensington: the buildings and politics of T. H. Huxley’s working environments. British Journal for the History of Science, 29(4):435–468, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027699.

Frangsmyr:1990:QSC

[FHR90] Tore Frängsmyr, J. L. Heilbron, and Robin E. Rider, editors. The quantifying spirit in the 18th Century, volume 7 of Uppsala studies in history of science. University of California Press, Berkeley, CA, USA, 1990. ISBN 0-520-07022-4. viii + 411 + 21 pp. LCCN Q125 .Q34 1990.

URL http://ark.cdlib.org/
ark:/13030/ft6d5nb455.

Field:1988:KGC

[Fie88] Judith Veronica Field. Kepler’s geometrical cosmology. University of Chicago Press, Chicago, IL, USA, 1988. ISBN 0-226-24823-2. xx + 243 pp. LCCN QB981 .F47

1988. URL http://www.loc.
gov/catdir/enhancements/
fy0608/87010837-d.html.

Field:1991:BRM

[Fie91] J. V. Field. Book review:

Michael J. Crowe. Theories of the World from Antiquity to the Copernican Revolution. New York: Dover Publications, 1990. Pp. xii + 229. ISBN 0-486-26173-5. £4.45. British Journal for the History of Science, 24(3):376– 377, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027240.

Field:1992:BRJ

[Fie92] J. V. Field. Book review: Janice A. Henderson. On the distance between Sun, Moon and Earth, according to Ptolemy, Copernicus and Reinhold. Leiden, New York, Copenhagen and Cologne. E. J. Brill Studia Copernicana — Brill’s series, vol. I, edited by J. Malicki and J. Soszyński (Polish Academy series, vol. XXX), 1991. Pp. xiv + 220. ISSN 0925-6806. ISBN 90-04-09378-8. British Journal for the History of Science, 25 (4):462, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027053.

Field:1993:BRV

[Fie93] J. V. Field. Book review: Victor E. Thoren, with contributions by John R. Christianson, The Lord of Uraniborg: A Biography of Tycho Brahe. Cambridge: Cambridge University Press, 1990. Pp. xi + 523. ISBN 0-521-35158-8. £40. British Journal for the History of Science, 26(1):83–84, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027347.

Field:1994:BRE

[Fie94] J. V. Field. Book reviews: S.

Y. Edgerton, The Heritage of Giotto’s Geometry: Art and Science on the Eve of the Scientific Revolution. Ithaca and London: Cornell University Press, 1991. Pp. x + 319. ISBN 0-8014-2573-5. $43.95. T. Da C. Kaufmann, The Mastery of Nature: Aspects of Art, Science, and Humanism in the Renaissance. Princeton: Princeton University Press, 1993. Pp. xix + 325, ISBN 0-691-03204-1. $39.95. British Journal for the History of Science, 27 (2):225–226, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027440.

Field:1997:IIM

[Fie97] Judith Veronica Field. The invention of infinity: mathematics and art in the Renaissance. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1997. ISBN 0-19-852394-7 (hardcover). xii + 250 pp. LCCN N7430.5 .F52

1997. URL http://www.loc.
gov/catdir/enhancements/
fy0636/97006805-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0636/97006805-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0724/97006805-b.html.

Friedel:1986:EEL

[FIF86] Robert D. (Robert Douglas)

Friedel,  Paul  Is-

rael, and Bernard S. Finn. Edison’s electric light: biography of an invention. Rutgers University Press, New Brunswick, NJ, USA, 1986. ISBN 0-8135-1118-6. xvi + 263 pp. LCCN TK4351 .F75 1986. US$35.00.

Figlio:1999:PSM

[Fig99] Karl Figlio. Psychoanalysis and the scientific mind: Robert Boyle. British Journal for the History of Science, 32(3):299– 314, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027988.

Findlen:1991:AAC

[Fin91]  Paula Findlen.  From Al-

drovandi to Algarotti: the contours of science in early modern Italy. British Journal for the History of Science, 24(3): 353–360, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027234.

Findlen:1992:BRG

[Fin92] Paula Findlen. Book review: Giuliana Mocchi. Idea, mente, specie. Platonismo e scienza in Johannes Marcus Marci (1595– 1667). Soveria Manelli: Rubbettino Editore, 1990. Pp. 314. L.30,000. British Journal for the History of Science, 25(3): 361–362, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027263.

Findlen:1993:BRF

[Fin93a]  Paula Findlen.  Book re-

view: Franco Arato, II secolo delle cose: Scienza e storia in Francesco Algarotti. Genova: Marietti, 1991. Pp. 173. ISBN 88-211-9566-X. £30.00. British Journal for the History of Science, 26(1):91–92, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027355.

Findlen:1993:BRG

[Fin93b] Paula Findlen. Book review:

Giuseppe Olmi, L’inventario del mondo: Cata-logazione della natura e luoghi del sapere nella prima età moderna. Bologna: Società editrice il Mulino, 1992. Pp. 457. ISBN 88-15-03647-4. L. 50000. British Journal for the History of Science, 26 (2):240–241, June 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027106.

Findlen:1993:BRN

[Fin93c]  Paula Findlen.  Book re-

view: Nicolas-Claude Fabri de Peiresc, Lettres à Cassiano dal Pozzo (1626–1637), éditées et commentées par Jean-François Lhote et Danielle Joyal. Amphion, Études d’histoire des techniques 3. Clermont-Ferrand: Éditions Adosa, 1992. Pp. 280. ISBN 2-86639-079-2. 260 FF. British Journal for the History of Science, 26(3): 356–357, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027408.

Finney:1993:PRN

[Fin93d] C. M. (Colin Michael) Finney. Paradise revealed: natural history in Nineteenth-Century Australia. Museum of Victoria, Melbourne, Victoria, Australia, 1993. ISBN 0-7306-2494-3. xv + 186 pp. LCCN QH21.A8 F56 1993.

Fisch:1991:WWP

[Fis91a] Menachem Fisch. William Whewell, philosopher of science. Clarendon Press, New York, NY, USA, 1991. ISBN 0-19-824240-9. xiv + 220 pp. LCCN Q175 .F5453 1991. US$59.00. URL http://www.

                                        loc.gov/catdir/
                                        enhancements/fy0638/
                                        90046590-d.html.

Fisher:1991:EJ

[Fis91b] Richard B. Fisher. Edward Jenner: 1749–1823. André Deutsch Ltd., London, UK, 1991. ISBN 0-233-98681-2. 361 + 8 pp. LCCN R489.J5 F57 1991x.

Fisch:1994:EWA

[Fis94] Menachem Fisch. “The emergency which has arrived”: the problematic history of nineteenth-century British algebra—a programmatic outline. British Journal for the History of Science, 27(3):247– 276, September 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027598.

Fischer:1998:MTL

[Fis98a] Fritz Fischer. Making them like us: Peace Corps volunteers in the 1960s. Smithsonian Institution Press, Washington, DC, USA, 1998. ISBN 1-56098-889-4. viii + 237 pp. LCCN HC60.5 .F57 1998.

Fisher:1998:BRL

[Fis98b] Saul Fisher. Book review:

Lisa T. Sarasohn, Gassendi’s Ethics: Freedom in a Mechanistic Universe. Ithaca: Cornell University Press, 1996. Pp. xii + 236. ISBN 0-8014-2947-1. £35.50. British Journal for the History of Science, 31(3): 361–375, September 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027820.

Field:1993:RRH

[FJ93] Judith Veronica Field and Frank A. J. L. James, editors. Renaissance and revolution: humanists, scholars, craftsmen, and natural philosophers in early modern Europe. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-43427-0 (hardcover). xv + 291 pp. LCCN Q125.2 .R45

1993. URL http://www.loc.
gov/catdir/description/
cam025/93006624.html;
http://www.loc.gov/
catdir/toc/cam021/
93006624.html.

Field:1998:I

[FJ98] J. V. Field and Frank A. J. L. James. Introduction. British Journal for the History of Science, 31(2):126–128, June 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027759.

Fox:1988:PMI

[FL88] Daniel M. Fox and Christopher Lawrence. Photographing medicine: images and power in Britain and America since 1840, volume 21 of Contributions in medical studies, 0886-8220. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1988. ISBN 0-313-23719-0. 357 pp. LCCN R486 .F69 1988.

Frangsmyr:1994:LMH

[FL^{+}94] Tore Frängsmyr, Sten Lindroth, et al., editors. Linnaeus, the man and his work, volume 18 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, revised edition, 1994. ISBN 0-88135-155-5 (hardcover), 0-88135-154-7 (paperback). xiv + 206 pp. LCCN QH44 .L56 1994. US$19.95, US$19.95.

Fleming:1998:MS

[Fle98] Martha Fleming. A metaphysical subject. British Journal for the History of Science, 31 (4):437, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Francis:2000:LPA

[FM00] R. A. Francis and Jean Mainil, editors. L’abbé Prévost au tournant du siècle. (French) [The Abbey Prévost at the turn of the century], volume 2000:11 of SVEC, 0435-2866. Voltaire foundation, Oxford, UK, 2000. ISBN 0-7294-0733-0. ix + 390 pp. LCCN PQ2021.Z5 A53 2000.

Freud:1985:CLS

[FMF85] Sigmund Freud, J. Moussaieff (Jeffrey Moussaieff) Masson, and Wilhelm Fliess. The complete letters of Sigmund Freud to Wilhelm Fliess, 1887–1904. Belknap Press of Harvard University Press, Cambridge, MA, USA and London, UK, 1985. ISBN 0-674-15420-7. xv + 505 + 20 pp. LCCN BF173.F85 A4 1985.

Fogg:1992:HAS

[Fog92] G. E. (Gordon Elliott) Fogg.

A history of Antarctic science. Studies in polar research. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-36113-3. xxi + 483 pp. LCCN G860 .F64 1992. URL http:

//www.loc.gov/catdir/
description/cam024/
91048063.html;   http://
www.loc.gov/catdir/toc/
cam029/91048063.html.

Force:1985:WWH

[For85] James E. Force. William Whiston, honest Newtonian. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26590-8. xxiii + 208 pp. LCCN BX5199.W52 F67

1985. URL http://www.loc.
gov/catdir/description/
cam022/84004316.html;
http://www.loc.gov/
catdir/samples/cam031/
84004316.html;   http://
www.loc.gov/catdir/toc/
cam029/84004316.html.

Forgan:1990:BRS

[For90a] Sophie Forgan. Book review: Susan Sheets-Pyenson. Cathedrals of Science: The Development of Colonial Natural History Museums during the Late Nineteenth Century. Kingston & Montreal: McGill-Queen’s University Press. 1988. Pp. 144, ill. ISBN 0-7735-0655-1. £19.95. British Journal for the History of Science, 23(1): 117–118, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026822.

Forgan:1990:RU

[For90b] Sophie Forgan. Review: [untitled]. British Journal for the History of Science, 23(4): 483–484, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027217.

Forgan:1991:BRR

[For91] Sophie Forgan. Book review:

Rosalind Williams. Notes on the Underground: An Essay on Technology, Society, and the Imagination. Cambridge, Mass, and London: MIT Press, 1990. Pp. xi + 265. ISBN 0-262-23145-X. £22.50. British Journal for the History of Science, 24(3):391–393, September 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027249.

Forgan:1998:FST

[For98] Sophie Forgan. Festivals of science and the two cultures: science, design and display in the Festival of Britain, 1951. British Journal for the History of Science, 31(2):217– 240, June 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027763.

Fox:1987:PLE

[Fox87] Christopher Fox, editor. Psychology and literature in the Eighteenth Century, volume 8 of AMS studies in the eighteenth century, 0196-6561. AMS Press, New York, NY, USA, 1987. ISBN 0-404-61474-4. xii + 372 + 7 pp. LCCN PR448.P75 P79 1987.

Fox:1988:LSI

[Fox88] Christopher Fox. Locke and the Scriblerians: identity and consciousness in early Eighteenth-Century Britain. University of California Press, Berkeley, CA, USA, 1988. ISBN 0-520-05859-3. x + 174 pp. LCCN BF697 .F68

                                        1988. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0641/88014201-d.html.

Fox:1992:BRM

[Fox92] Daniel M. Fox. Book review: Mirko D. Grmek. History of AIDS: Emergence and Origin of a Modern Pandemic. Translated by Russell C. Maulitz and Jacalyn Duffin. Princeton, New Jersey: Princeton University Press, 1990. Pp. xii + 279. $29.95. British Journal for the History of Science, 25 (4):500, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027090.

Fox:1995:TCM

[Fox95] Robert Fox, editor. Technological change: methods and themes in the history of technology, volume 1 of Studies in the history of science, technology and medicin. Harwood Academic Publishers, Chur, Switzerland and New York, NY, USA, 1995. ISBN 3-7186-5792-9. vi + 271 pp. LCCN ????

Franklin:1986:NE

[Fra86] Allan Franklin. The neglect of experiment. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-32016-X. xii + 290 pp. LCCN Q182.3 .F73

1986. URL http://www.loc.
gov/catdir/description/
cam023/86002604.html;
http://www.loc.gov/
catdir/samples/cam031/
86002604.html;   http://
www.loc.gov/catdir/toc/
cam024/86002604.html.

Frangsmyr:1989:SSR

[Frä89]  Tore  Frängsmyr,

editor. Science in Sweden: the Royal Swedish Academy of Sciences, 1739–1989. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1989. ISBN 0-88135-092-3. viii + 291 + 34 pp. LCCN Q64.S937 S25 1989.

Fraser:1990:BRJ

[Fra90] Craig G. Fraser. Book review:

Julian B. Barbour. Absolute or Relative Motion: A Study from a Machian Point of View of the Discovery and Structure of Dynamical Theories. Volume 1: The Discovery of Dynamics. Cambridge, Cambridge University Press, 1989. Pp. xv + 746. ISBN 0-521-32467X. £60.00, $95.00. British Journal for the History of Science, 23 (3):349–350, September 1990.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026772.

French:1990:RU

[Fre90a] Roger French. Review: [untitled]. British Journal for the History of Science, 23(4): 484–485, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027218.

FresquetFebrer:1990:FMA

[Fre90b] José Luis Fresquet Febrer. Francisco Méndez Álvaro (1806–1883) y las ideas sanitarias del liberalismo moderado. (Spanish) [Francisco Méndez Álvaro (1806–1883) and the sanitary ideas of moderate liberalism ], volume 14 of Textos Clásicos espaänoles de la salud pública. Ministerio de sanidad y consumo, Secretaría general técnica Publicaciones, documentación y biblioteca, Madrid, Spain, 1990. ISBN 84-7670-217-5. 212 pp. LCCN ????

Freudenthal:1992:SGF

[Fre92] Gad Freudenthal, editor. Studies on Gersonides: a Fourteenth-Century Jewish philosopher-scientist, volume 36 of Collection de travaux de l’Acadèmie internationale d’histoire des sciences, 0169-7897. E. J. Brill, Leiden, The Netherlands, 1992. ISBN 90-04-09641-8 (hardcover). xxi + 422 pp. LCCN Q143.L57 S78 1992.

Freudenthal:1993:BRM

[Fre93] Gad Freudenthal. Book review: Margaret J. Osler (ed.), Atoms, Pneuma, and Tranquility. Epicurean and Stoic Themes in European Thought. Cambridge: Cambridge University Press, 1991. Pp. xii + 304. ISBN 0-521-40048-1. £32.50, $49.50. British Journal for the History of Science, 26(1):86–88, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027351.

French:1994:ANH

[Fre94] R. K. (Roger Kenneth) French.

Ancient natural history: histories of nature. Sciences of antiquity. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1994. ISBN 0-415-08880-1 (hardcover), 0-415-11545-0 (paperback). xxii + 357 pp. LCCN 509; 606; QH15 .F74 1994; M96.F00079; 508.093.

Freudenthal:1995:ATM

[Fre95]  Gad Freudenthal.  Aristo-

tle’s theory of material substance: heat and pneuma, form and soul. Clarendon Press, New York, NY, USA, 1995. ISBN 0-19-824093-7. xi + 235 pp. LCCN B491.S8 F74

1995. URL http://www.loc.

                                        gov/catdir/enhancements/
                                        fy0605/94048637-d.html.

Frieden:1981:RPE

[Fri81] Nancy Mandelker Frieden. Russian physicians in an era of reform and revolution, 1856– 1905. Princeton University Press, Princeton, NJ, USA, 1981. ISBN 0-691-05335-9, 0-691-05335-0. xvii + 378 pp. LCCN R532 .F74. US$32.50, US$32.50.

Friedel:1983:PPM

[Fri83] Robert D. (Robert Douglas) Friedel. Pioneer plastic: the making and selling of celluloid. University of Wisconsin Press, Madison, WI, USA, 1983. ISBN 0-299-09170-8. xix + 153 pp. LCCN TP1180.C5 F74 1983. US$19.95.

Friedman:1989:AWV

[Fri89] Robert Marc Friedman. Appropriating the weather: Vilhelm Bjerknes and the construction of a modern meteorology. Cornell University Press, Ithaca, NY, USA, 1989. ISBN 0-8014-2062-8. xx + 251 pp. LCCN QC858.B53 F75 1989.

Fritscher:1991:VGB

[Fri91] Bernhard Fritscher. Vulkanismusstreit und Geochemie: die Bedeutung der Chemie und des Experiments in der Vulkanismus-Neptunismus-Kontroverse. (German) [Volcanism controversy and geochemistry: the meaning of chemistry and experiments in the Volcanism–Neptune controversy], volume 25 of Boethius. F. Steiner Verlag, Stuttgart, Germany, 1991. ISBN 3-515-05865-6. viii + 346 pp. LCCN QE522 .F75 1991.

Fruton:1982:BBH

[Fru82] Joseph S. (Joseph Stewart)

Fruton. A bio-bibliography for the history of the biochemical sciences since 1800. American Philosophical Society, Philadelphia, PA, USA, 1982. ISBN 0-87169-983-4. xiii + 885 pp. LCCN Z5524.B54 F78 1982.

Fruton:1990:CSS

[Fru90] Joseph S. (Joseph Stewart)

Fruton. Contrasts in scientific style: research groups in the chemical and biochemical sciences, volume 191 of Memoirs series / American Philosophical Society, 0065-9738. American Philosophical Society, Philadelphia, PA, USA, 1990. ISBN 0-87169-191-4. ix + 473 pp. LCCN Q11 .P612 vol. 191.

Fenton:1984:LRS

[FS84] Alexander Fenton and Geoffrey Stell, editors. Loads and roads in Scotland and beyond: land transport over 6000 years. A Scottish history and culture paperback. John Donald, Edinburgh, Scotland, 1984. ISBN 0-85976-107-X (paperback). vii + 144 pp. LCCN HE363.S3 L6. US$8.50. URL http://

catalog.hathitrust.org/

                                        api/volumes/oclc/
                                        11225466.html.

Figurovskii:1988:APB

[FS88] Nikolai Aleksandrovich Figurovskii and I. Urii Ivanovich Solovev. Aleksandr Porfirevich Borodin: a chemist’s biography. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 3-540-17888-0, 0-387-17888-0 (US). LCCN ???? Translated from the Russian by Charlene Steinberg and George B. Kauffman; foreword by Martin D. Kamen.

Fisch:1991:WWC

[FS91] Menachem Fisch and Simon Schaffer, editors. William Whewell, a composite portrait. Clarendon Press, New York, NY, USA, 1991. ISBN 0-19-824900-4. xiv + 403 pp. LCCN B1612.W54 W55

                                        1991. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0605/90041928-d.html.

Faraday:1991:MFB

[FTG91] Michael Faraday, Ryan D. Tweney, and David Gooding. Michael Faraday’s ’Chemical notes, hints, suggestions, and objects of pursuit’ of 1822, volume 17 of IEE history of technology series. P. Peregrinus in association with the Science Museum, London, UK, 1991. ISBN 0-86341-255-6. xvii + 152 pp. LCCN QD3 .F22 1991.

Fuchs:1999:BRA

[Fuc99] Eckhardt Fuchs. Book review:

Andreas Daum, Wissenschaftspopularisierung im 19. Jahrhundert. Bürgerliche Kultur, naturwissenschaftliche Bildung und die deutsche Öffentlichkeit, 1848– 1914. München: R. Oldenbourg Verlag, 1998. Pp. xii + 617. ISBN 3-486-56337-8. DM148.00. British Journal for the History of Science, 32(3): 363–378, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027996.

Fujimura:1996:CSS

[Fuj96] Joan H. Fujimura. Crafting science: a sociohistory of the quest for the genetics of cancer. Harvard University Press, Cambridge, MA, USA, 1996. ISBN 0-674-17553-0 (hardcover). x + 322 pp. LCCN RC268.415 .F84 1996.

Fullinwider:1991:DFK

[Ful91] S. P. Fullinwider. Darwin faces Kant: a study in nineteenthcentury physiology. British Journal for the History of Science, 24(1):21–44, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027014.

Fuller:1998:BRR

[Ful98] Steve Fuller. Book review: R.

G. A. Dolby, Uncertain Knowledge: An Image of Science for a Changing World. Cambridge: Cambridge University Press, 1996. Pp. xi + 365. ISBN 0-521-56004-7, £40. British Journal for the History of Science, 31(2):241–250, June 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027767.

Furley:1987:GC

[Fur87] David J. Furley. The Greek cosmologists. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33328-8 (vol. 1). ???? pp. LCCN BD495 .F87

1987. URL http://www.loc.
gov/catdir/description/
cam023/86026384.html;
http://www.loc.gov/
catdir/enhancements/
fy0632/86026384-t.html.

Furth:1988:SFP

[Fur88] Montgomery Furth. Substance, form, and psyche: an Aristotelean metaphysics. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-34143-4. xiv + 300 pp. LCCN B491.M4 F87

1988. URL http://www.loc.
gov/catdir/description/
cam023/87023872.html;
http://www.loc.gov/
catdir/enhancements/
fy0707/87023872-t.html.

Furley:1989:CPE

[Fur89]  David J. Furley.  Cosmic

problems: essays on Greek and Roman philosophy of nature. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33330-X. xiv + 258 pp. LCCN BD581 .F87

1989. URL http://www.loc.
gov/catdir/description/
cam023/88011120.html;
http://www.loc.gov/
catdir/toc/cam023/
88011120.html.

Fyfe:1997:RWP

[Fyf97] Aileen Fyfe. The reception of William Paley’s natural theology in the University of Cambridge. British Journal for the History of Science, 30(3):321– 335, September 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027865.

Fyfe:1999:BRD

[Fyf99] Aileen Fyfe. Book review:

E. A. Davis (ed.), Science in the Making: Scientific Development as Chronicled by Historic Papers in the Philosophical Magazine – with Commentaries and Illustrations, Volume 2: 1850–1900. London: Taylor & Francis, 1997. Pp. xix + 406, 16 plates. ISBN 0-7484-0642-5. £59.95. British Journal for the History of Science, 32(2):237–251, June 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4028087.

Gigerenzer:1989:ECH

[G^{+}89] Gerd Gigerenzer et al., editors. The empire of chance: how probability changed science and everyday life. Ideas in context. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33115-3. xvii + 340 pp. LCCN QA273 .E57

                                        1989. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/88016928.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam022/
                                        88016928.html.

Gallouet:2009:LDS

[G^{+}09] Catherine Gallouët et al., editors. L’Afrique du siècle des lumières: savoirs et représentations. (French) [Africa of the century of light: knowledge and representations], volume 2009:05 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2009. ISBN 0-7294-0959-7. xxix + 307 pp. LCCN PQ265 .A37 2009.

Garber:1998:CHS

[GA98] Daniel Garber and Michael Ayers, editors. The Cambridge history of Seventeenth-Century philosophy. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-58864-2 (set). xvii + 1616 pp. LCCN B801 .C35 1998. URL http:

                                        //www.loc.gov/catdir/
                                        description/cam027/
                                        96025475.html;   http://

www.loc.gov/catdir/toc/
cam025/96025475.html.

Gabbey:1990:PC

[Gab90] Alan Gabbey. Pierre Costabel (1912–1989). British Journal for the History of Science, 23(4):386, 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027197.

Galison:1987:HEE

[Gal87] Peter Galison. How experiments end. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-27915-4 (paperback), 0-226-27914-6. xii + 330 pp. LCCN QC6 .G22 1987. US$12.00 (est.), US$36.00 (est.). URL http:/

/www.loc.gov/catdir/
description/uchi052/
87005928.html;
http://www.loc.gov/
catdir/enhancements/
fy0642/87005928-b.html;
http://www.loc.gov/
catdir/enhancements/
fy0642/87005928-t.html.

Garber:1992:DMP

[Gar92]  Daniel Garber.  Descartes’

metaphysical physics. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-28217-1 (hardcover), 0-226-28219-8 (paperback). xiv + 389 pp. LCCN QC6 .G25 1992. US$60.00, US$23.95.

URL http://www.loc.gov/

                                        catdir/description/
                                        uchi051/91030190.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0608/91030190-t.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0609/91030190-b.html.

Garber:1998:WWM

[Gar98] Daniel Garber. Who was that masked man? British Journal for the History of Science, 31(1):55–62, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027777.

Gascoigne:1999:RSE

[Gas99] John Gascoigne. The Royal Society and the emergence of science as an instrument of state policy. British Journal for the History of Science, 32 (2):171–184, June 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4028082.

Gauld:1992:HH

[Gau92] Alan Gauld. A history of hypnotism. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-30675-2 (hardcover). xvii + 738 pp. LCCN BF1125 .G38

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        cam025/92005106.html;
                                        http://www.loc.gov/

catdir/toc/cam027/
92005106.html.

Gaukroger:1999:BRD

[Gau99]  Stephen Gaukroger.  Book

review: Daniel Garber and Michael Ayers (eds.), The Cambridge History of Seventeenth-Century Philosophy, 2 vols. Cambridge: Cambridge University Press, 1998. Pp. xvii + 1616. ISBN 0-521-58864-2. £90.00 (set). British Journal for the History of Science, 32 (1):111–124, March 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027972.

Gayon:1992:DAD

[Gay92]  Jean Gayon.  Darwin et

l’après Darwin: une histoire de l’hypothèse de sélection naturelle. (French) [Darwin and post-Darwin: a history of the hypothesis of natural selection]. Histoire des idées, théorie politique et recherches en sciences sociales. Editions Kimé, Paris, France, 1992. ISBN 2-908212-14-5. 453 pp. LCCN QH375 .G38 1992. 250F.

Gay:1996:IRW

[Gay96] Hannah Gay. Invisible resource: William Crookes and his circle of support, 1871– 1881. British Journal for the History of Science, 29(3):311– 336, September 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027736.

Gay:1998:NHC

[Gay98]  Hannah Gay.  No ‘Hea-

then’s Corner’ here: the failed campaign to memorialize Herbert Spencer in Westminster Abbey. British Journal for the History of Science, 31(1): 41–54, March 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027776.

Galilei:1995:DSM

[GB95a]  Galileo  Galilei  and

Antonio Beltrán. Diálogo sobre los dos máximos sistemas del mundo ptolemaico y copernicano. (Spanish) [Dialog on the two maximum systems of the world: Ptolemaic and Copernican]. Alianza, Madrid, Spain, 1995. ISBN 84-206-9412-6. lxxxiii + 417 pp. LCCN QB41 .G129418 1995.

Goldstein:1995:RRD

[GB95b] Bernard R. Goldstein and Peter Barker. The role of Rothmann in the dissolution of the celestial spheres. British Journal for the History of Science, 28(4):385–403, December 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027544.

Geison:1995:PSL

[Gei95] Gerald L. Geison. The private science of Louis Pasteur. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03442-7. xiv + 378 + 16 pp. LCCN Q143.P2

G35 1995. URL http://www.
loc.gov/catdir/
description/prin021/
94035338.html;   http://
www.loc.gov/catdir/toc/
prin031/94035338.html.

Greene:1983:LBH

[GEMM83] Edward Lee Greene, Frank N.

Egerton, Robert P. (Robert Patrick) McIntosh, and Rogers Mc-Vaugh. Landmarks of botanical history. Stanford University Press, Stanford, CA, USA, 1983. ISBN 0-8047-1075-9. x + 1139 pp. LCCN QK21.E85 G74 1983.

Gerl:1989:TAR

[Ger89] Armin Gerl. Trigonometrischastronomisches Rechnen kurz vor Copernicus: der Briefwechsel Regiomontanus-Bianchini. (German) [Trigonometricastronomical counting shortly before Copernicus: exchange of letters with Regiomontanus Bianchini], volume 21 of Boethius. F. Steiner Verlag, Stuttgart, Germany, 1989. ISBN 3-515-05373-5. 357 pp. LCCN QB145 .G47 1989.

Gerrard:1991:BRM

[Ger91] Bill Gerrard. Book review:

Mary S. Morgan. The History of Econometric Ideas. Cambridge: Cambridge University Press, 1990. Pp. xv + 296. ISBN 0-521-37398-0. £25.00, $44.50. British Journal for the History of Science, 24(4): 485–486, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027145.

Gerrard:1995:BRP

[Ger95] Bill Gerrard. Book review: Philip Mirowski (ed.), Edgeworth on Chance, Economic Hazard, and Statistics. Lanham, MD: Rowman and Littlefield, 1994. Pp. vii + 462. ISBN 0-8476-7751-6. $45.00. British Journal for the History of Science, 28(3):375–376, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027668.

Getz:1996:BRR

[Get96]  Faye Getz.  Book review:

Roger French, Ancient Natural History: Histories of Nature. Sciences in Antiquity. London and New York: Routledge, 1994. Pp. xxii + 357. ISBN 0-415-11545-0. £15.99. British Journal for the History of Science, 29(3):361–362, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027739.

Geuter:1992:PPN

[Geu92] Ulfried Geuter. The professionalization of psychology in Nazi Germany. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-33297-4 (hardcover). xxi + 335 pp. LCCN BF108.G3 G4813 1992. URL http://

www.loc.gov/catdir/
description/cam024/
91026576.html;   http://
www.loc.gov/catdir/toc/
cam029/91026576.html.

Godel:1986:CW

[GF^{+}03] Kurt Gödel, Solomon Feferman, et al. Collected works. Clarendon Press, New York, NY, USA, 1986–2003. ISBN 0-19-503964-5 (vol. 1), 0-19-503972-6 (vol. 2), 0-19-507255-3 (vol. 3), 0-19-850073-4 (vol. 4), 0-19-850075-0 (vol. 5). ???? pp. LCCN QA9 .G5313

1986. URL http://www.loc.
gov/catdir/enhancements/
fy0603/85015501-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0603/85015501-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0725/85015501-b.html.

Grattan-Guinness:1990:BRN

[GG90] I. Grattan-Guinness. Book review: Nicole Hulin-Jing. L’organisation de l’enseignement des sciences: la voie ouverte par le second Empire. Paris: Comité des Travaux Historiques et Scientifiques, 1989. Pp. 336. ISBN 2-7355-0171-X. 200 F.Fr. British Journal for the History of Science, 23(2):242–245, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026744.

Gavroglu:1991:TMK

[GG91a] Kostas Gavroglu and Yorgos Goudaroulis, editors. Through measurement to knowledge: the selected papers of Heike Kamerlingh Onnes 1853–1926, volume 124 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-7433-X (hardcover), 94-009-2079-2. ISSN 0068-0346. cxv + 570 pp. LCCN Q174 .B67 vol.

124.  URL https://link.
springer.com/book/10.
1007/978-94-009-2079-8.

Grattan-Guinness:1991:BRB

[GG91b] I. Grattan-Guinness. Book review: Bruna Ingrao and Giorgio Israel. The Invisible Hand: Economic Equilibrium in the History of Science, translated by Ian McGilvray. Cambridge, Mass, and London: MIT Press, 1990. Pp. xiii + 491. ISBN 0-262-09028-7. £42.75. British Journal for the History of Science, 24(4): 484–485, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027144.

Grattan-Guinness:1991:BRM

[GG91c] I. Grattan-Guinness. Book review: Margaret Schabas. A World Ruled by Number: William Stanley Jevons and the Rise of Mathematical Economics. Princeton: Princeton University Press, 1990. Pp. xii + 192. ISBN 0-691-08543-9. $29.95. British Journal for the History of Science, 24(4): 486–487, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027146.

Gottfried:1992:TUP

[GG92] Lew Gottfried and Fran Gottfried. Traveling uncharted paths: the tale of two retirees and The Peace Corps. Fairway Press, Lima, OH, USA, 1992. ISBN 1-55673-455-7. 280 pp. LCCN HC60.5 .G68 1992.

Grattan-Guinness:1993:DRC

[GG93a] I. Grattan-Guinness. Dr. R. C.

H. Tanner (1900–1992). British Journal for the History of Science, 26(2(89)):229–231, 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Grattan-Guinness:1993:ODR

[GG93b] I. Grattan-Guinness. Obituary: Dr. R. C. H. Tanner (1900–1992). British Journal for the History of Science, 26 (2):229–231, June 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027099.

Grattan-Guinness:1996:BRM

[GG96] I. Grattan-Guinness. Book review: Marco Panza and Jean-Claude Pont (eds.), Les Savants et l’epistémologie vers la fin du XIXe siècle. Paris: Blanchard, 1995. Pp. vii + 282. ISBN 2-85367-196-8. 150.00 FF. British Journal for the History of Science, 29(3):375–376, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027751.

Guicciardini:1990:GSO

[GGC90] Francesco Guicciardini, Lodovico Guicciardini, and Raffaella Castagnola. I Guicciardini e le scienze occulte: l’oroscopo di Francesco Guicciardini: lettere di alchimia, astrologia e cabala a Luigi Guicciardini. (Italian) [The hidden Guicciardini and science: the horoscope of Francisco Guicciardini: letters on alchemy, astrology and cabala to Luigi Guicciardini ], volume 19 of Studi e testi / Istituto Nazionale di Studi sul Rinascimento, 0394-4409. L. S. Olschki, Firenze, Italy, 1990. ISBN 88-222-3761-7. vii + 397 + 8 pp. LCCN BF1728.G84 G8425 1990.

GuytondeMorveau:1994:SCD

[GGSGB94] Louis-Bernard Guyton de Morveau, Emmanuel Grison, Michelle Sadoun-Goupil, and Patrice Bret. A scientific correspondence during the chemical revolution: Louis-Bernard Guyton de Morveau and Richard Kirwan, 1782– 1802, volume 17 of Berkeley papers in history of science. Office for History of Science and Technology, University of California at Berkeley, Berkeley, CA, USA, 1994. ISBN 0-918102-21-9. vi + 257 pp. LCCN QD22.G94 A4 1994.

Gaposchkin:1984:CPG

[GH84] Cecilia Helena Payne Gaposchkin and Katherine Haramundanis. Cecilia Payne-Gaposchkin: an autobiography and other recollections. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25752-2. vi + 269 pp. LCCN QB36.G37 A33 1984.

Galison:1992:BSG

[GH92]  Peter  Louis  Gal-

ison and Bruce William Hevly, editors. Big science: the growth of large-scale research. Stanford University Press, Stanford, CA, USA, 1992. ISBN 0-8047-2335-4, 0-8047-1879-2. xi + 392 pp. LCCN ????

Gaposchkin:1996:CPG

[GH96] Cecilia Helena Payne Gaposchkin and Katherine Haramundanis. Cecilia Payne-Gaposchkin: an autobiography and other recollections. Cambridge University Press, Cambridge, UK, second edition, 1996. ISBN 0-521-48251-8 (hardcover), 0-521-48390-5 (paperback). xxii + 277 pp. LCCN QB36.G37

                                        A33 1996. URL http://www.
                                        loc.gov/catdir/
                                        description/cam027/
                                        95044001.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam021/95044001.html.

Giere:1988:ESC

[Gie88] Ronald N. Giere. Explaining science: a cognitive approach. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1988. ISBN 0-226-29205-3, 0-226-29206-1 (paperback). xxi + 321 pp. LCCN M88.E00693; 500; 501; Q175 .G4889 1988; J.c.81; 0187.

Gillispie:1983:MBI

[Gil83] Charles Coulston Gillispie. The Montgolfier brothers and the invention of aviation, 1783–1784: with a word on the importance of ballooning for the science of heat and the art of building railroads. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-08321-5. xi + 210 pp. LCCN TL617.M66 G48 1983. US$35.00.

Gindikin:1988:TPM

[Gin88]  S.  G.

(Semen Grigorevich) Gindikin. Tales of physicists and mathematicians. Birkhäuser Verlag, Basel, Switzerland, 1988.

ISBN 0-8176-3317-0. x + 157 pp. LCCN QC7 .G5613 1988. US$30.00.

Giroud:1986:MCL

[Gir86] Françoise Giroud. Marie Curie, a life. Holmes and Meier, New York, NY, USA, 1986. ISBN 0-8419-0977-6, 0-8419-0978-4 (paperback). 291 + 16 pp. LCCN QD22.C8 G4813 1986.

Gooding:1985:FRE

[GJ85] David Gooding and Frank A.

J. L. James, editors. Faraday rediscovered: essays on the life and work of Michael Faraday, 1791–1867. MacMillan Publishing Company, New York, NY, USA, 1985. ISBN 0-333-39320-1, 0-943818-91-5. 258 pp. LCCN ????

Grafton:1986:HHI

[GJ86] Anthony Grafton and Lisa Jardine. From humanism to the humanities: the institutionalizing of the liberal arts in fifteenth-and sixteenth-century Europe. Harvard U.P, Cambridge, Mass, 1986. ISBN 0-674-32460-9, 0-7156-2100-9. xvi + 224 pp. LCCN LA106 .G73 1986.

Gjertsen:1984:CSS

[Gje84] Derek Gjertsen. The classics of science: a study of twelve enduring scientific works. Lilian Barber Press, New York, NY, USA, 1984. ISBN 0-936508-09-4. iii + 374 pp. LCCN Q125 .G54 1984.

Gjertsen:1986:NH

[Gje86] Derek Gjertsen. The Newton handbook. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1986. ISBN 0-7102-0279-2. xiv + 665 pp. LCCN QC16.N7 G54 1986. US$25.00.

Gramiccia:1988:LCL

[GL88] Gabriele Gramiccia and Charles Kaufmann Ledger. The life of Charles Ledger, (1818–1905): alpacas and quinine. MacMillan Publishing Company, New York, NY, USA, 1988. ISBN 0-333-45710-2. xiv + 222 pp. LCCN F2236.L43 G73

                                        1988. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/59101195.
                                        html.

Glen:1982:RJC

[Gle82] William Glen. The road to Jaramillo: critical years of the revolution in earth science. Stanford University Press, Stanford, CA, USA, 1982. ISBN 0-8047-1119-4. xvi + 459 pp. LCCN QE511.4 .G53 1982.

Gleason:1991:RSL

[Gle91] Mary Louise Gleason. The Royal Society of London: years of reform, 1827–1847. Harvard dissertations in the history of science. Garland Publishing, Inc., New York, NY, USA, 1991. ISBN 0-8240-7446-7. ix + 532 pp. LCCN Q41 .G64 1991.

Glick:1987:CRR

[Gli87] Thomas F. Glick, editor. The Comparative Reception of Relativity, volume 103 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1987. ISBN 90-277-2498-9, 94-009-3875-6. ISSN 0068-0346. vii + 412 pp. LCCN Q174 .B67 vol. 103; QC173.5. US$79.00

(US). URL https://link.
springer.com/book/10.
1007/978-94-009-3875-5.

Goldman:1983:DAS

[Gol83] Martin Goldman. The demon in the aether: the story of James Clerk Maxwell. P. Harris, Edinburgh, Scotland, 1983. ISBN 0-86228-026-5. 224 + 12 pp. LCCN QC16.M4 G65 1983. US$15.00.

Goldstein:1987:CCF

[Gol87] Jan Ellen Goldstein. Console and classify: the French psychiatric profession in the Nineteenth Century. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-32279-0. xiii + 414 pp. LCCN RC450.F7 G65 1987.

Goldammer:1991:GMM

[Gol91a] Kurt Goldammer. Der göttliche Magier und die Magierin Natur: Religion, Naturmagie und die Anfänge der Naturwissenschaft vom Spätmittelalter bis zur Renaissance; mit Beiträgen zum Magie-Verständnis des Paracelsus. (German) [ The divine Magi and the Magi nature: Religion, nature magic and the beginnings of the natural science from the late Middle Ages to the Renaissance; with contributions for the magic understanding of Paracelsus], volume 5 of Kosmosophie. F. Steiner Verlag, Stuttgart, Germany, 1991. ISBN 3-515-05584-3. 136 pp. LCCN BF1598.P2 G65 1991.

Golinski:1991:BRF

[Gol91b] Jan Golinski. Book review: Frederic Lawrence Holmes. Eighteenth-Century Chemistry as an Investigative Enterprise. Berkeley: Office for History of Science and Technology, University of California at Berkeley, 1989. Pp. ii + 144. ISBN 0-918102-16-2. $16.00 (paperback). British Journal for the History of Science, 24(1): 102–103, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027021.

Goldstein:1992:DST

[Gol92a] Jack S. Goldstein. A different sort of time: the life of Jerrold R. Zacharias, scientist, engineer, educator. MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-07138-X. xviii + 373 pp. LCCN QC16.Z23 G65 1992.

Golinski:1992:SPC

[Gol92b] Jan Golinski. Science as public culture: chemistry and enlightenment in Britain, 1760–1820. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-39414-7 (hardcover). xi + 342 pp. LCCN QD18.G7 G65

1992. URL http://www.loc.
gov/catdir/description/
cam024/91039024.html;
http://www.loc.gov/
catdir/toc/cam029/
91039024.html.

Golinski:1993:BRL

[Gol93] Jan Golinski. Book review:

Larry Stewart. The Rise of Public Science: Rhetoric, Technology, and Natural Philosophy in Newtonian Britain, 1660–1750. Cambridge: Cambridge University Press, 1992. Pp. xxxiv + 453, illus. ISBN 0-521-41700-7. £45.00, $69.95. British Journal for the History of Science, 26(3):363–364, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027413.

Golinski:1994:BRH

[Gol94]  Jan Golinski.  Book re-

view: Harry Collins and Trevor Pinch, The Golem: What Everyone Should Know about Science. Cambridge: Cambridge University Press, 1993. Pp. xii + 164. ISBN 0-521-35601-6. £10.95, $19.95. British Journal for the History of Science, 27(4):487–488, December 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027639.

Goldgar:1995:ILC

[Gol95a] Anne Goldgar. Impolite learning: conduct and community in the Republic of Letters, 1680– 1750. Yale University Press, New Haven, CT, USA, 1995. ISBN 0-300-05359-2. xiii + 395 pp. LCCN CB203 .G64 1995.

Golinski:1995:BRB

[Gol95b]  Jan Golinski.  Book re-

view: Bernadette Bensaude-Vincent, Lavoisier: Mémoires d’une révolution. Preface by Michel Serres. Paris: Flammarion, 1993. Pp. 469. ISBN 2-08-211208-X. 160.00 FF. British Journal for the History of Science, 28(2):239–240, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027685.

Goodman:1988:PPG

[Goo88] David C. Goodman. Power and penury: government, technology and science in Philip II’s Spain. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-30532-2, 0-521-52477-6. xii + 275 pp. LCCN Q127.S7; M89.E02119.

Gooday:1990:BRF

[Goo90a] Graeme Gooday. Book review: Frank A. J. L. James (ed.). The Development of the Laboratory: Essays on the Place of Experiment in Industrial Civilization. London: Macmillan Press, 1989. Pp. xv + 260. ISBN 0-333-48331-6. £37.50. British Journal for the History of Science, 23(2):226–228, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026734.

Gooday:1990:PMG

[Goo90b] Graeme Gooday. Precision measurement and the genesis of physics teaching laboratories in Victorian Britain. British Journal for the History of Science, 23(1):25–51, March 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026801.

Gooding:1990:EMM

[Goo90c] David Gooding. Experiment and the making of meaning: human agency in scientific observation and experiment, volume 5 of Science and philosophy. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0719-4. xviii + 310 pp. LCCN Q182.3 .G69 1990.

Gooday:1991:NLD

[Goo91a] Graeme Gooday. ‘nature’ in the laboratory: domestication and discipline with the microscope in Victorian life science. British Journal for the History of Science, 24(3):307– 341, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027232.

Goodman:1991:BRJ

[Goo91b] David Goodman. Book reviews: José M. López Piñero. Los origines en España de los Estudios sobre la Salud Publica. Colección Textos Clásicos Españoles de la Salud Pública, No. 1. Madrid: Ministerio de Sanidad y Consumo, 1989. Pp.

245. ISBN 84-7670-164-0 (hardcover); 74-7670-165-9 (paperback). José Luis Fresquet Febrer. Francisco Méndez Alvaro (1806–1883) y las ideas Sanitarias del Liberalismo Moderado. Colección Textos Clásicos Españoles de la Salud Publica, No. 14. Madrid: Ministerio de Sanidad y Consumo, 1990. Pp. 212. ISBN 84-7670-217-5 (hardcover); 84-7670-216-7 (paperback). Antonio M. Rey González. Estudios médicosociales sobre marginados en la España del siglo XIX. Colección Textos Clásicos Españoles de la Salud Pública, No. 17. Madrid: Ministerio de Sanidad y Consumo, 1990. Pp.

237. ISBN 84-7670-204-3 (hardcover); 84-7670-205-1 (paperback). Nicolás Monardes. La historia medicinal de las cosas que se traen de nuestras Indias Occidentales (1565–1574). Madrid: Ministerio de Sanidad y Consumo, 1989. Pp. 291. ISBN 84-7670-191-8. No prices given. British Journal for the History of Science, 24(4):491– 493, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027151.

Gooding:1994:RPF

[Goo94] David Gooding. Re-presenting Faraday. British Journal for the History of Science, 27(3): 361–364, September 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027605.

Gooday:1996:BRA

[Goo96] Graeme Gooday. Book review: Adrian Desmond, Huxley: The Devil’s Disciple. London: Michael Joseph, 1994. Pp. xvii + 475. ISBN 0-7181-3641-1. £20.00. British Journal for the History of Science, 29(1): 106–107, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027528.

Gooday:1997:BRJ

[Goo97] Graeme Gooday. Book review: Jed Z. Buchwald, The Creation of Scientific Effects: Heinrich Hertz and Electric Waves. Chicago: Chicago University Press, 1994. Pp. xiv + 482. ISBN 0-226-07887-6, £59.95, $75.00 (hardcover); 0-226-07888-4, £26.25, $32.95 (paperback). British Journal for the History of Science, 30 (2):233–249, June 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027722.

Gosling:1987:BFN

[Gos87] Francis G. (Francis George) Gosling. Before Freud: neurasthenia and the American medical community, 1870–1910. University of Illinois Press, Urbana, IL, USA, 1987. ISBN 0-252-01406-5. xviii + 192 pp. LCCN RC552.N5 G67 1987.

Gottlieb:1992:CNL

[Got92] Jean S. Gottlieb. A checklist of the Newberry Library’s printed books in science, medicine, technology, and the pseudosciences ca. 1460–1750, volume 1195 of Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1992. ISBN 0-8240-5171-8. xxiii + 312 pp. LCCN Z7405.H6 G67 1992.

Gouk:1991:BRC

[Gou91] Penelope Gouk. Book review: C. F. C. Beeson, Clockmaking in Oxfordshire 1400–1850, 3rd edn, with a new introduction and index by A. V. Simcock. Oxford: Museum of the History of Science, 1989. Pp. vii + 212. ISBN 0-903364-06-9. £22.00. British Journal for the History of Science, 24(4):

489–490, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027149.

Gouk:1992:BRC

[Gou92] Penelope Gouk. Book review:

A. C. Crombie. Science, Optics and Music in Medieval and Early Modern Thought. London and Ronceverte: The Hambledon Press, 1990. Pp. xii + 474, illus. ISBN 0-907628-79-6. £37.50. British Journal for the History of Science, 25(3): 359–360, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027261.

Gouk:1995:BRR

[Gou95a] Penelope Gouk. Book review: R. G. W. Anderson, J. A. Bennett and W. F. Ryan (eds.), Making Instruments Count: Essays on Historical Scientific Instruments presented to Gerard L’Estrange Turner. Aldershot: Variorum, 1993. Pp. xix + 492. ISBN 0-86078-394-4. No price given. British Journal for the History of Science, 28(1):124– 125, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027504.

Gouk:1995:WAC

[Gou95b] Penelope Gouk, editor. Wellsprings of achievement: cultural and economic dynamics in early modern England and Japan. Variorum, Aldershot, Hampshire, UK, 1995. ISBN 0-86078-465-7. viii + 271 pp. LCCN HD8399.E52 W45 1995.

Gouk:1996:PPM

[Gou96] Penelope Gouk. Performance practice: Music, medicine and natural philosophy in interregnum Oxford. British Journal for the History of Science, 29 (3):257–288, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027734.

Gould:1997:WCU

[Gou97] Paula Gould. Women and the culture of university physics in late nineteenth-century Cambridge. British Journal for the History of Science, 30(2): 127–149, June 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027714.

Gouk:1998:BRS

[Gou98a] Penelope Gouk. Book review: Suzanne B. Butters, The Triumph of Vulcan: Sculptors’ Tools, Porphyry, and the Prince in Ducal Florence, 2 vols. Florence: Leo S. Olschki, 1996. Pp. 724, illus. ISBN 88-222-4411-7. 320,000 lire (set). British Journal for the History of Science, 31(1):

63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027782.

Gouk:1998:BRT

[Gou98b] Penelope Gouk. Book review: Thomas L. Hankins and Robert J. Silverman, Instruments and the Imagination. Princeton: Princeton University Press. Pp. xiv + 337. ISBN 0-691-02997-0. £33.50, $39.50. British Journal for the History of Science, 31(1): 63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027783.

Gowing:1983:RCN

[Gow83] Ronald Gowing. Roger Cotes, natural philosopher. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-23741-6. x + 210 pp. LCCN QA804 .G68 1983. US$30.00

(est.). URL http://www.loc.
gov/catdir/description/
cam031/82001154.html;
http://www.loc.gov/
catdir/samples/cam031/
82001154.html;   http://
www.loc.gov/catdir/toc/
cam031/82001154.html.

Galilei:1992:SNM

[GP92] Galileo Galilei and Isabelle Pantin. Sidereus nuncius, le messager céleste. (French) [Sidereus nuncius = The celestial messenger]. Les Belles lettres, Paris, France, 1992. ISBN 2-251-34505-1. civ + 115 pp. LCCN 95.E11231; QB41 .G1714 1992.

Gowing:1988:SPP

[GR88] Margaret Gowing and Nicolaas A. Rupke, editors. Science, politics and the public good: essays in honour of Margaret Gowing. MacMillan Publishing Company, New York, NY, USA, 1988. ISBN 0-333-44159-1. xv + 260 pp. LCCN Q127.G4 S433 1988.

Goodman:1991:RSE

[GR91] David C. (David Charles) Goodman and Colin A. (Colin Archibald) Russell, editors. The rise of scientific Europe, 1500–1800. Hodder and Stoughton, London, UK, 1991. ISBN 0-340-55861-X. ix + 437 pp. LCCN Q127.E8 R574 1991.

Grant:1984:DEC

[Gra84] Edward Grant. In defense of the Earth’s centrality and immobility: scholastic reaction to Copernicanism in the Seventeenth Century, volume 74, part 4 (1984) of Transactions of the American Philosophical Society, 0065-9746. American Philosophical Society, Philadelphia, PA, USA, 1984. ISBN 0-87169-744-0 (paperback). 69 pp. LCCN QB981 .G66 1984.

Gray:1985:LDE

[Gra85]  Jeremy John Gray.  Lin-

ear differential equations and group theory from Riemann to Poincaré. Birkhäuser Verlag, Basel, Switzerland, 1985. ISBN 0-8176-3318-9, 3-7643-3318-9. xxv + 460 pp. LCCN QA372 .G68 1985.

Graham:1987:SPH

[Gra87] Loren R. Graham. Science, philosophy, and human behavior in the Soviet Union. Columbia University Press, New York, NY, USA, 1987. ISBN 0-231-06442-X. xiii + 565 pp. LCCN Q127.S696 G73 1987.

Gray:1991:BRJ

[Gra91] Jeremy Gray. Book review:

John Earman. World Enough and Space-Time: Absolute versus Relational Theories of Space and Time. Cambridge, Mass, and London: MIT Press, 1990. Pp. xiv + 223. ISBN 0-262-05040-4. £22.50. British Journal for the History of Science, 24(4):497, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027155.

Grant:1992:OVT

[Gra92] Edward Grant. Obituary: Victor E. Thoren (1935–1991). British Journal for the History of Science, 25(2):255– 257, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027304.

Grant:1994:PSO

[Gra94] Edward Grant. Planets, stars, and orbs: the medieval cosmos, 1200–1687. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-43344-4 (hardcover). xxiii + 816 pp. LCCN QB981 .G664

1994. URL http://www.loc.
gov/catdir/description/
cam025/93025899.html;
http://www.loc.gov/
catdir/toc/cam025/
93025899.html.

Graham:1998:WWL

[Gra98] Loren R. Graham. What have we learned about science and technology from the Russian experience? Stanford University Press, Stanford, CA, USA, 1998. ISBN 0-8047-2985-9 (hardcover), 0-8047-3276-0 (paperback). xiii + 177 pp. LCCN Q175.5 .G73 1998. URL http:

//www.loc.gov/catdir/
description/cam028/
97041635.html;   http://
www.loc.gov/catdir/
enhancements/fy0710/
97041635-b.html;   http:/
/www.loc.gov/catdir/toc/
cam026/97041635.html.

Greene:1982:GNC

[Gre82] Mott T. Greene. Geology in the Nineteenth Century: changing views of a changing world. Cornell history of science series. Cornell University Press, Ithaca, NY, USA, 1982. ISBN 0-8014-1467-9. 324 pp. LCCN QE11 .G73 1982.

Grelon:1986:ICT

[Gre86] André Grelon, editor. Les Ingénieurs de la crise: titre et profession entre les deux guerres. (French) [The Engineers of the crisis: title and profession between the two wars], volume 21 of Recherches d’histoire et de sciences sociales = Studies in history and the social sciences, 0249-5619. Editions de l’Ecole des hautes études en sciences sociales, Paris, France, 1986. ISBN 2-7132-0875-0. 461 pp. LCCN TA71 .I55 1986. 250F.

Greenaway:1991:BRP

[Gre91] Frank Greenaway. Book review:

Peter J. T. Morris. The American Synthetic Rubber Research Program. Philadelphia: University of Pennsylvania Press, 1989. Pp. xii + 191. ISBN 0-8122-8205-1. British Journal for the History of Science, 24(4):473–474, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027134.

Greene:1992:NKP

[Gre92a]  Mott T. Greene.  Natural

knowledge in preclassical antiquity. The Johns Hopkins University Press, Baltimore, MD, USA, 1992. ISBN 0-8018-4292-1. xix + 182 pp. LCCN BL435 .G74

1992. URL http://www.loc.
gov/catdir/bios/jhu051/
91026952.html;   http://
www.loc.gov/catdir/
description/jhu051/
91026952.html;   http://
www.loc.gov/catdir/toc/
jhu051/91026952.html.

Gregory:1992:NLN

[Gre92b]  Frederick Gregory.  Nature

lost?: natural science and the German theological traditions of the Nineteenth Century. Harvard University Press, Cambridge, MA, USA, 1992. ISBN 0-674-60483-0. viii + 341 pp. LCCN BL245 .G74 1992.

Grell:1992:BRB

[Gre92c] Ole Peter Grell. Book reviews: Bruce T. Moran. The Alchemical World of the German Court: Occult Philosophy and Chemical Medicine in the Circle of Moritz of Hessen (1572–1632). Sudhoff’s Archiv, Beiheft 29. Stuttgart: Franz Steiner Verlag 1991. Pp. 193. ISBN 3-515-05369-7. DM 58. Bruce T. Moran. Chemical Pharmacy Enters the University: Johannes Hartmann and the Didactic Care of Chymiatria in the Early Seventeenth Century. Madison: American Institute for the History of Pharmacy, 1991. Pp. vii + 88. ISBN 0-931292-24-7, $16.50 (hardcover); 0-931292-23-9, $7.50 (paperback). British Journal for the History of Science, 25 (3):360–361, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027262.

Grell:1993:BRJ

[Gre93a] Ole Peter Grell. Book review: Jost Weyer, Graf Wolfgang II von Hohenlohe und die Alchemic: Alchemistische Studien in Schloss Weikersheim, 1587–1610. Sigmaringen: Jan Thorbecke Verlag, 1992. Pp.

516. ISBN 3-7995-7639-8. DM 88.00. British Journal for the History of Science, 26(3):353– 354, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027405.

Grell:1993:BRK

[Gre93b] Ole Peter Grell. Book review:

Kurt Goldhammer, Der göttliche Magier und die Magierin Natur: Religion, Naturmagie und die Anfänge der Naturwissenschaft vom Spätmittelalter bis zur Renaissance; mit Beitragen zum Magie-Verständnis des Paracelsus. Stuttgart: Franz Steiner Verlag, 1991. Pp. 136. ISBN 3-515-05584-3. DM 58.00. British Journal for the History of Science, 26(1):82, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027345.

Greenaway:1994:WNW

[Gre94] Frank Greenaway. Wanted: new word. British Journal for the History of Science, 27(3): 357–360, September 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027604.

Greenaway:1995:BRD

[Gre95] Frank Greenaway. Book review: Donald Cardwell, The Fontana History of Technology. London: Fontana, 1994. Pp. xviii + 565. ISBN 0-00-686176-8. £9.99. British Journal for the History of Science, 28(2):234–235, June 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027680.

Grell:1996:BRP

[Gre96] Ole Peter Grell. Book reviews: Pamela H. Smith, The Business of Alchemy: Science and Culture in the Holy Roman Empire. Princeton: Princeton University Press, 1994. Pp. xii + 308. ISBN 0-691-05691-9. £30.00, $45.00. Raphael Patai, The Jewish Alchemists: A History and Sourcebook. Princeton: Princeton University Press, 1994. Pp. xv + 617. ISBN 0-691-03290-4. £29.95, $35.00. British Journal for the History of Science, 29(1): 93–94, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027517.

Grmek:1990:PRB

[Grm90] Mirko D. (Mirko Dražen)

Grmek. La première révolution biologique: réflexions sur la physiologie et la médicine du XVIIe siècle. (French) [The first biological revolution: reflections on the physiology of medicine in the 17th Century]. Bibliothèque scientifique Payot. Editions Payot, Paris, France, 1990. ISBN 2-228-88277-1. 358 + 8 pp. LCCN QP21 .G76 1990. 220F.

Gross:1989:SNQ

[Gro89] Nikolaus Gross. Senecas Naturales quaestiones: Komposition, naturphilosophische Aussagen und ihre Quellen. (German) [Senecas Naturales quaestiones: Composition, naturalphilosophical statements and their sources], volume 27 of Palingenesia. F. Steiner Verlag, Stuttgart, Germany, 1989. ISBN 3-515-05385-9. 335 pp. LCCN PA6661.N3 G76 1989.

Grove:1990:HHH

[Gro90] David I. Grove. A history of human helminthology. C.A.B. International, Wallingford, UK, 1990. ISBN 0-85198-689-7. viii + 848 + 12 pp. LCCN RC119.7 .G76 1990.

Grosholz:1991:CMP

[Gro91] Emily Grosholz. Cartesian method and the problem of reduction. Clarendon Press, New York, NY, USA, 1991. ISBN 0-19-824250-6. viii + 161 pp. LCCN B1875 .G76

1991. URL http://www.loc.
gov/catdir/enhancements/
fy0605/90035814-d.html.

Grosholz:1993:BRD

[Gro93] Emily Grosholz. Book review: Daniel Garber, Descartes’ Metaphysical Physics. Chicago and London: University of Chicago Press, 1992. Pp. xiv + 389. ISBN 0-226-28217-1, £47.95, $60.00 (hardcover); 0-226-28219-8, £19.25, $23.95 (paperback). British Journal for the History of Science, 26 (1):85–86, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027350.

Gross:1995:BDN

[Gro95a] Joseph Gross. The new dictionary of national biography. British Journal for the History of Science, 28(2):255– 256, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027694.

Grove:1995:GIC

[Gro95b] Richard (Richard H.) Grove. Green imperialism: colonial expansion, tropical island Edens, and the origins of environmentalism, 1600–1860. Studies in environment and history. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-40385-5 (hardcover). xiv + 540 pp. LCCN GE195 .G76

1995. URL http://www.loc.
gov/catdir/description/
cam025/93028654.html;
http://www.loc.gov/
catdir/toc/cam026/
93028654.html.

Gage:1988:BHL

[GS88] Andrew Thomas Gage and William T. (William Thomas) Stearn. A bicentenary history of the Linnean Society of London. Academic Press, New York, USA, 1988. ISBN 0-12-273150-6. ix + 242 pp. LCCN QH1 .G24 1988.

Gruber:1992:ROC

[GT92] Jacob W. Gruber and John C.

Thackray. Richard Owen commemoration: three studies, volume 1 of Historical studies in the life and earth sciences. Natural History Museum Publications, London, UK, 1992. ISBN 0-565-01109-X. ix + 181 pp. LCCN QL31.O9 G78 1992.

Guerlac:1981:NC

[Gue81] Henry Guerlac. Newton on the Continent. Cornell University Press, Ithaca, NY, USA, 1981. ISBN 0-8014-1409-1. 169 pp. LCCN QC16.N7 G82.

Gunther:1984:ILR

[Gun84] Albert Edward (Albert Everard) Gunther. An introduction to the life of the Rev. Thomas Birch D.D., F.R.S. 1705–1766: leading editor of the General Dictionary’ — 1741: Secretary of the Royal Society and Trustee of the British Museum: based on a daily ’Diary’, 1735–1764 and other sources. Halesworth Press, Suffolk, England, 1984. ISBN 0-9507276-1-X. x + 118 pp. LCCN Q143.B49 G85 1984.

Gutting:1989:MFA

[Gut89] Gary Gutting. Michel Foucault’s archaeology of scientific reason. Modern European philosophy. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36619-4, 0-521-36698-4 (paperback). xiii + 306 pp. LCCN B2430.F724 G87

1989. URL http://www.loc.
gov/catdir/description/
cam023/88031881.html;
http://www.loc.gov/
catdir/toc/cam025/
88031881.html.

Galilei:1989:BNB

[GV89] Galileo Galilei and Albert Van Helden. Sidereus nuncius, or, The Sidereal messenger. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-27902-2 (hardcover), 0-226-27903-0 (paperback). xii + 127 pp. LCCN QB41 .G173

1989. URL http://www.loc.
gov/catdir/description/
uchi052/88025179.html;
http://www.loc.gov/
catdir/toc/uchi051/
88025179.html.

Hannah:1982:EMP

[H^{+}82] Leslie Hannah et al. Engineers, managers, and politicians: the first fifteen years of nationalised electricity supply in Britain. MacMillan Publishing Company, New York, NY, USA, 1982. ISBN 0-333-22087-0. xiii + 336 pp. LCCN HD9685.G7 H36

1982b. URL http://catalog.
hathitrust.org/api/
volumes/oclc/8715072.html.

Hartcup:1984:CA

[HA84] Guy Hartcup and T. E. Allibone. Cockcroft and the atom. Adam Hilger Ltd., Bristol, UK, 1984. ISBN 0-85274-759-4. xii + 320 pp. LCCN QC774.C6 H37 1984. US$18.95.

Hammond:1987:CLA

[HA87] John H. Hammond and Jill Austin. The camera lucida in art and science. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-527-3. xii + 200 pp. LCCN NC730 .H26 1987.

Holbrook:1992:SPD

[HAB92] Mary Holbrook, R. G. W.

(Robert Geoffrey William) Anderson, and D. J. Bryden. Science preserved: a directory of scientific instruments in collections in the United Kingdom and Eire. Her Majesty’s Stationary Office, London, England, 1992. ISBN 0-11-290060-7. 271 pp. LCCN Q185.7 .H65 1992. US$35.

Hackmann:1984:SSS

[Hac84] Willem Dirk Hackmann. Seek and strike: sonar, anti-submarine warfare, and the Royal Navy, 1914–54. Her Majesty’s Stationary Office, London, England, 1984. ISBN 0-11-290423-8. xxxv + 487 + 32 pp. LCCN V214 .H33 1984. US$15.95.

Hacking:1990:TC

[Hac90a] Ian Hacking. The taming of chance. Ideas in context. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-38014-6, 0-521-38884-8 (paperback). xiii + 264 pp. LCCN BD595 .H33 1990. URL http:

                                        //www.loc.gov/catdir/
                                        description/cam024/
                                        89036411.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam022/89036411.html.

Hackmann:1990:RUa

[Hac90b] Willem Hackmann. Review: [untitled]. British Journal for the History of Science, 23(4): 485–486, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027219.

Hackmann:1990:RUb

[Hac90c] Willem Hackmann. Review: [untitled]. British Journal for the History of Science, 23(4): 487–488, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027220.

Hackmann:1990:BRJ

[Hac90d] Willem D. Hackmann. Book review: J. A. Bennett. The Divided Circle. A History of Instruments for Astronomy, Navigation and Surveying. Oxford: Phaidon-Christie’s, 1987. Pp. 224. ISBN 0-7148-8038-8. £45. British Journal for the History of Science, 23(2): 223–224, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026732.

Hackmann:1990:BRT

[Hac90e] Willem D. Hackmann. Book review: Thomas G. Manning. U.S. Coast Survey vs. Naval Hydrographic Office. A 19thcentury Rivalry in Science and Politics. Tuscaloosa and London: The University of Alabama Press, 1988. Pp. 202. ISBN 0-8173-0390-1. $21.95. British Journal for the History of Science, 23(3):379–380, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026794.

Hacking:1991:AP

[Hac91] Ian Hacking. Artificial phenomena. British Journal for the History of Science, 24(2): 235–241, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027168.

Hackmann:1992:BRCa

[Hac92a] Willem Hackmann. Book review: A. Chapman. Dividing the Circle. The Development of Critical Angular Measurement in Astronomy 1500–1650. New York, London, Toronto, Sydney, Tokyo, Singapore: Ellis Horwood Library of Space Science and Space Technology, 1990. Pp. 209. ISBN 0-13-217555-X. £15.95. British Journal for the History of Science, 25(2):264–265, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027307.

Hackmann:1992:BRCb

[Hac92b] Willem Hackmann. Book review: C. S. Maffioli and L. C. Palm (eds.). Italian Scientists in the Low Countries in the 17th and 18th Centuries. Amsterdam: Rodopi, 1989. Pp.

334. ISBN 90-5183-121-8. Dfl. 110, $55.00. British Journal for the History of Science, 25 (2):271–272, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027313.

Hackmann:1992:BRG

[Hac92c] Willem Hackmann. Book review: G. L’E. Turner. The Great Age of the Microscope.

The Collection of the Royal Microscopical Society through 150 years. Bristol and New York: Adam Hilger, 1989. Pp. ix + 379. ISBN 0-85274-029-4. £39.50. British Journal for the History of Science, 25 (2):282–284, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027321.

Hackmann:1992:BRP

[Hac92d] Willem Hackmann. Book review: P. R. Morris. A History of the World Semiconductor Industry. IEE History of Technology Series 12. London: Peter Peregrinus on behalf of the Institution of Electrical Engineers, 1990. Pp. 171. ISBN 0-86341-227-0. £32. British Journal for the History of Science, 25(4):495–496, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027084.

Hackmann:1993:BRB

[Hac93a] Willem Hackmann. Book review: E. B. Callick, Metres to Microwaves: British Development of Active Components for Radar Systems 1937 to 1944. IEE History of Technology Series 11. London: Peter Peregrinus Ltd on behalf of the Institution of Electrical Engineers, 1990. Pp. xi + 240. ISBN 0-86341-212-2. £44. British Journal for the History of Science, 26(3):376– 377, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027425.

Hackmann:1993:BRJ

[Hac93b] Willem Hackmann. Book review: E. J. Dijksterhuis, Clio’s Stiefkind, edited with an Introduction and commentary by Klaus van Berkel. Amsterdam: Uitgeverij Bert Bakker, 1990. Pp. 321. ISBN 90-351-0786-1. No price given. British Journal for the History of Science, 26 (3):378–379, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027427.

Hackmann:1993:BRK

[Hac93c] Willem Hackmann. Book review: Kostas Gavroglu and Yorgos Goudaroulis (eds.), Through Measurement to Knowledge: The Selected Papers of Heike Kamerlingh Onnes 1853–1926. Boston Studies in the Philosophy of Science, 24. Dordrecht: Kluwer, 1991. Pp. cxv + 570. ISBN 0-7923-0825-5. £97.00, $149.00, Dfl. 260.00. British Journal for the History of Science, 26(2):249–251, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027114.

Hackmann:1994:BRJ

[Hac94] Willem Hackmann. Book review: James Wood, History of International Broadcasting. IEE History of Technology Series, 19. Stevenage: Peter Peregrinus in association with the Science Museum, London, 1992. Pp. xvix + 258. ISBN 0-86341-281-5. £30. British Journal for the History of Science, 27(1):122–124, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027592.

Hackmann:1995:BRD

[Hac95] Willem Hackmann. Book review: Dan Ch. Christensen (ed.), European Historiography of Technology from the TISC-Conference in Roskilde. Odense University Studies in History and Social Sciences,

156. Odense, Denmark: Odense University Press, 1993. Pp.

242. ISBN 87-7492-910-0. DKK 275.00, $37.00. British Journal for the History of Science, 28 (2):247–249, June 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027692.

Hadden:1994:SME

[Had94] Richard W. Hadden. On the shoulders of merchants: exchange and the mathematical conception of nature in early modern Europe. SUNY series in science, technology, and society. State University of New York Press, Albany, NY, USA, 1994. ISBN 0-7914-2011-6, 0-7914-2012-4 (paperback). xviii + 191 pp. LCCN Q175.52.E85 H33 1994.

Hagen:1992:EBO

[Hag92] Joel Bartholemew Hagen. An entangled bank: the origins of ecosystem ecology. Rutgers University Press, New Brunswick, NJ, USA, 1992. ISBN 0-8135-1823-7 (hardcover), 0-8135-1824-5 (paperback). x + 245 pp. LCCN QH540.8 .H34 1992. US$38.00, US$16.00.

Hakfoort:1995:OAE

[Hak95a] Casper Hakfoort. Optics in the age of Euler: conceptions of the nature of light, 1700– 1795. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-40471-1. vi + 243 pp. LCCN QC403 .H3513

1995. URL http://www.loc.
gov/catdir/description/
cam026/94031656.html;
http://www.loc.gov/
catdir/toc/cam023/
94031656.html.

Hakosalo:1995:BRA

[Hak95b] Heini-Eliisa Hakosalo. Book review: Adam Crabtree, From Mesmer to Freud: Magnetic Sleep and the Roots of Psychological Healing. New Haven and London: Yale University Press, 1994. Pp. x + 413. ISBN 0-300-05588-9. £30. British Journal for the History of Science, 28 (3):355–357, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027654.

Hall:1983:RS

[Hal83a] A. Rupert (Alfred Rupert)

Hall. The revolution in science, 1500–1750. Longman, London, UK, third edition, 1983. ISBN 0-582-49133-9 (paperback). viii + 373 pp. LCCN Q125 .H28 1983.

URL http://www.h-net.org/
review/hrev-a0b9f8-aa.

Hallam:1983:GGC

[Hal83b] A. (Anthony) Hallam. Great geological controversies. Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1983. ISBN 0-19-854431-6, 0-19-854430-8 (paperback). vii + 182 pp. LCCN QE11 .H35 1983. US$34.50, US$18.50.

URL http://www.loc.gov/
catdir/enhancements/
fy0640/83002357-d.html.

Hall:1984:ASN

[Hal84] Marie Boas Hall. All scientists now: the Royal Society in the Nineteenth Century. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26746-3. xii + 261 pp. LCCN Q41.L85 H35

1984. URL http://www.loc.
gov/catdir/description/
cam022/84007705.html;
http://www.loc.gov/

                                        catdir/samples/cam031/
                                        84007705.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam028/84007705.html.

Hall:1990:BRW

[Hal90a] A. Rupert Hall. Book review: William R. Shea (ed.). Revolutions in Science: Their Meaning and Relevance. Canton: Science History Publications, 1988. Pp. xi + 291. British Journal for the History of Science, 23(2):249–250, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026749.

Hall:1990:RU

[Hal90b] A. Rupert Hall. Review: [untitled]. British Journal for the History of Science, 23(4): 461–463, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027201.

Hall:1990:HMM

[Hal90c] A. Rupert (Alfred Rupert) Hall. Henry More — magic, religion, and experiment. Blackwell science biographies. Blackwell, Oxford, UK, 1990. ISBN 0-631-17295-5. xii + 304 pp. LCCN Q143.M77 H35 1990. US$60.00.

Hall:1990:BRG

[Hal90d] Marie Boas Hall. Book review: Gad Freudenthal (ed). Études sur/Studies on Hélène Metzger. Corpus, no. 8/9 Paris, 1988. Pp. 280. ISSN 0296-8916. Fr.F. 70. British Journal for the History of Science, 23(2): 238–239, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026741.

Hall:1991:BRM

[Hal91a] A. Rupert Hall. Book review:

Mordechai Feingold (ed.). Before Newton: The Life and Times of Isaac Barrow. Cambridge: Cambridge University Press, 1990. Pp. xii + 380. ISBN 0-521-30694-9. £35.00, $49.50. British Journal for the History of Science, 24(3):381– 382, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027243.

Hall:1991:BRP

[Hal91b] A. Rupert Hall. Book review: Peter Rowlands. Oliver Lodge and the Liverpool Physical Society, Liverpool Historical Studies, 4. Liverpool: Liverpool University Press, 1990. ISBN 0-85323-027-7. £15.00. British Journal for the History of Science, 24(2):269– 270, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027191.

Hall:1991:PEL

[Hal91c] Marie Boas Hall. Promoting experimental learning: experiment and the Royal Society 1660–1727. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-40503-3. xiii + 207 pp. LCCN Q41 .H33

                                        1991. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/91027181.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        91027181.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam028/91027181.html.

Hall:1992:BRR

[Hal92a] A. Rupert Hall. Book review: R. C. Olby, G. N. Cantor, J. R. R. Christie and M. J. S. Hodge (eds.), Companion to the History of Modern Science. London and New York: Routledge, 1990. Pp. xxvi-1081. ISBN 0-415-01988-5. £65.00. British Journal for the History of Science, 25(4): 457–460, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027050.

Hall:1992:INA

[Hal92b] A. Rupert (Alfred Rupert) Hall. Isaac Newton, adventurer in thought. Blackwell science biographies. Blackwell, Oxford, UK, 1992. ISBN 0-631-17906-2. xv + 468 pp. LCCN QC16.N7 H35 1992.

Hall:1992:BRM

[Hal92c] Marie Boas Hall. Book review: Mary Louise Gleason.

The Royal Society of London: Years of Reform. 1827–1847. New York and London: Garland Publishing, 1991. Pp. ix + 532. ISBN 0-8240-7446-7. £95.00. British Journal for the History of Science, 25(4): 477–478, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027066.

Hall:1992:BRT

[Hal92d] Marie Boas Hall. Book review: Tore Frängsmyr (ed.). Solomon’s House Revisited: The Organization and Institutionalization of Science. Canton, Mass.: Science History Publications (USA), 1990. Pp. xiii + 350. ISBN 0-88135-066-4. $49.50 (USA & Canada), $54.00 (elsewhere). British Journal for the History of Science, 25(2):291–292, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027329.

Hall:1993:AWL

[Hal93] A. Rupert (Alfred Rupert)

Hall. All was light: an introduction to Newton’s Opticks. Clarendon Press, New York, NY, USA, 1993. ISBN 0-19-853985-1. vi + 252 pp. LCCN QC353 .H35 1993. US$52.50.

URL http://www.loc.gov/
catdir/enhancements/
fy0636/92043731-d.html;
http://www.loc.gov/

                                        catdir/enhancements/
                                        fy0636/92043731-t.html.

Hall:1994:BRD

[Hal94a] A. Rupert Hall. Book review: Domenico Bertoloni Meli, Equivalence and Priority: Newton versus Leibniz. Including Leibniz’s Unpublished Manuscripts on the Principia. Oxford: Clarendon Press, 1993. Pp. ix + 318. ISBN 0-19-853945-2. £55.00. British Journal for the History of Science, 27(3):369–371, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027610.

Hall:1994:OSW

[Hal94b] A. Rupert Hall. Obituary: Sir William Paton (1917–1993). British Journal for the History of Science, 27(4):465– 466, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027626.

Hall:1994:BRO

[Hal94c] Marie Boas Hall. Book review: Olaf Pedersen, Lovers of Learning: A History of the Royal Danish Academy of Sciences and Letters, 1742– 1992. Copenhagen: Munksgaard: Det Kongelige Danske Videnskabernes Selskab, 1992. Pp. 348. ISBN 87-7304-236-6. DKK 300.00. British Journal for the History of Science, 27 (2):241–242, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027455.

Hall:1995:BRM

[Hal95] A. Rupert Hall. Book review: Michel Blay. Les Raisons de l’infini: Du monde clos à l’univers mathématique. Paris: Gallimard, 1993. Pp. 258. ISBN 2-07-072992-3. 110 FF. British Journal for the History of Science, 28(2):236–237, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027682.

Hamlin:1990:SIW

[Ham90] Christopher Hamlin. A science of impurity: water analysis in Nineteenth Century Britain. Adam Hilger Ltd., Bristol, UK, 1990. ISBN 0-7503-0042-6. xiii + 342 pp. LCCN TD257 .H35 1990.

Hankins:1985:SE

[Han85] Thomas L. Hankins. Science and the Enlightenment. Cambridge history of science. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-24349-1, 0-521-28619-0 (paperback). viii + 216 pp. LCCN Q125 .H355 1985. URL http:

//www.loc.gov/catdir/
description/cam022/
84016988.html;   http://
www.loc.gov/catdir/toc/
cam029/84016988.html.

Harman:1982:EFM

[Har82a] P. M. (Peter Michael) Harman. Energy, force, and matter: the conceptual development of Nineteenth-Century physics. Cambridge history of science. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24600-8, 0-521-28812-6 (paperback). ix + 182 pp. LCCN QC7

                                        .H257. URL http://www.loc.
                                        gov/catdir/description/
                                        cam032/81017029.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0632/81017029-t.html.

Harman:1982:MNP

[Har82b] Peter M. Harman. Metaphysics and natural philosophy: the problem of substance in classical physics. Harvester Press, Brighton, Sussex, UK, 1982. ISBN 0-389-20305-X, 0-7108-0451-2. xvi + 168 pp. LCCN ????

Hart:1985:PF

[Har85] Clive Hart. The prehistory of flight. University of California Press, Berkeley, CA, USA, 1985. ISBN 0-520-05213-7. xvii + 279 pp. LCCN TL516 .H27 1985.

Harrington:1987:MMD

[Har87] Anne Harrington. Medicine, mind, and the double brain: a study in Nineteenth-Century thought. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-08465-3. xiii + 336 pp. LCCN QP385.5 .H37 1987. US$39.50.

Harris:1989:MMM

[Har89] Ruth Harris. Murders and madness: medicine, law and society in the Fin de Siècle. Oxford historical monographs. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-822991-7, 0-19-820259-8 (paperback). viii + 366 + 4 pp. LCCN HV6133 HAR 1989.

Harman:1990:BRD

[Har90a] P. M. Harman. Book review:

Donald S. L. Cardwell. James Joule: A Biography. Manchester: Manchester University Press, 1989. Pp. x + 333. ISBN 0-7190-3025-0. £35.00. British Journal for the History of Science, 23(2):230, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026737.

Harris:1990:BRA

[Har90b] Ruth Harris. Book review:

Anne Harrington. Medicine, Mind, and the Double Brain: A Study in Nineteenth-Century Thought. Princeton: Princeton University Press, 1987. Pp. xiii + 336. ISBN 0-691-08465-3. £24.70. British Journal for the History of Science, 23(3): 371–373, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026790.

Harwood:1990:RU

[Har90c] Jonathan Harwood. Review: [untitled]. British Journal for the History of Science, 23(4): 480–481, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027214.

Hardy:1992:BRT

[Har92a] Anne Hardy. Book review: Theodore Steinberg, Nature Incorporated: Industrialisation and the Waters of New England. Cambridge: Cambridge University Press, 1991. Pp. xvi + 283. ISBN 0-521-39215-2. £30.00, $34.50. British Journal for the History of Science, 25(4):479–480, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027068.

Harris:1992:BRM

[Har92b] Ruth Harris. Book review: Mikulás Teich and Roy Porter (eds.). Fin de siècle and its Legacy. Cambridge: Cambridge University Press, 1990. Pp. 352. ISBN 0-521-34108-6, £37.50 (hardcover); 0-521-34915-X, £12.95 (paperback). British Journal for the History of Science, 25(4):485– 486, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027073.

Harrison:1992:BRD

[Har92c]  Mark Harrison.  Book re-

view: Deepak Kumar (ed.). Science and Empire: Essays in Indian Context (1700–1947). Delhi: Anamika Prakashan, 1991. Pp. xiv + 205. ISBN 81-85150-19-2. £20.00, $30.00, Rs. 200. British Journal for the History of Science, 25(2): 289–290, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027327.

Harrison:1992:BRS

[Har92d] Mark Harrison. Book review:

Satpal Sangwan. Science, Technology and Colonisation: An Indian Experience, 1757–1857. Delhi: Anamika Parakashan, 1991. Pp. xviii + 196. ISBN 81-85150-11-7. No price given. British Journal for the History of Science, 25(3):382–383, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027283.

Harrison:1992:TMN

[Har92e]  Mark Harrison.  Tropical

medicine in nineteenth-century India. British Journal for the History of Science, 25(3):299– 318, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027255.

Harman:1993:BRB

[Har93a] P. M. Harman. Book review: Bruce J. Hunt, The Maxwellians. Ithaca and London: Cornell University Press, 1991. Pp. xiii + 266. ISBN 0-8014-2641-3. $38.45. British Journal for the History of Science, 26 (1):117–118, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027380.

Harrison:1993:BRP

[Har93b]  Mark Harrison.  Book re-

view: Poonam Bala, Imperialism and Medicine in Bengal: A Socio-Historical Perspective. New Delhi, Newbury Park, London: Sage Publications, 1991. Pp. 1-174. ISBN 81-7036-245-8 (India), 0-8039-9100-2 (USA). British Journal for the History of Science, 26 (1):97–98, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027363.

Harrison:1993:BRR

[Har93c] Mark Harrison. Book review: Ray Desmond, The European Discovery of the Indian Flora. Oxford: Oxford University Press/Royal Botanic Gardens, 1992. Pp. viii + 355. ISBN 0-19-854684-X. £60.00. British Journal for the History of Science, 26(2):237–238, June 1993.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027103.

Hartley:1994:CHP

[Har94a] Mary M. Hartley. Conflicts in human progress: sexual selection and the Fisherian ‘runaway’. British Journal for the History of Science, 27(2): 177–196, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027434.

Harwood:1994:BRU

[Har94b] Jonathan Harwood. Book review: Ute Deichmann, Biologen unter Hitler: Vertreibung, Karrieren, Forschungsförderung. Frankfurt: Campus Verlag, 1992. Pp. 370. ISBN 3-593-34763-6. DM 38.00. British Journal for the History of Science, 27(4): 480–481, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027633.

Harwood:1994:IIF

[Har94c] Jonathan Harwood. Institutional innovation in fin de siècle Germany. British Journal for the History of Science, 27(2): 197–211, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027435.

Harman:1995:BRJ

[Har95a] P. M. Harman. Book review: Joseph F. Mulligan (ed.), Heinrich Rudolf Hertz (1857–1894): A Collection of Articles and Addresses. New York: Garland Publishing, 1994. Pp. xxi + 442. $70.00. British Journal for the History of Science, 28 (3):362, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027659.

Harrison:1995:BRL

[Har95b]  Mark Harrison.  Book re-

view: Lewis Pyenson, Civilizing Mission: Exact Sciences and French Overseas Expansion, 1830–1940. Baltimore and London: Johns Hopkins University Press, 1993. Pp. xxi + 377. ISBN 0-8018-4421-5. £37.00. British Journal for the History of Science, 28(1): 119–120, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027500.

Harman:1996:BRE

[Har96a] P. M. Harman. Book review: Elizabeth Carber, Stephen G. Brush and C. W. F. Everitt (eds.), Maxwell on Heat and Statistical Mechanics: On ‘Avoiding All Personal Enquiries of Molecules’. London: Associated University Presses, 1995, Pp. 550. ISBN 0-934223-34-3. £45.00. British Journal for the History of Science, 29 (1):107–109, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027529.

Harrington:1996:RSH

[Har96b] Anne Harrington. Reenchanted science: holism in German culture from Wilhelm II to Hitler. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-02142-2 (hardcover). xxiv + 309 pp. LCCN QH501 .H37

1996. URL http://www.loc.
gov/catdir/description/
prin021/95048463.html;
http://www.loc.gov/
catdir/toc/prin031/
95048463.html.

Harrison:1996:BRA

[Har96c]  Mark Harrison.  Book re-

view: Anne Godlewska and Neil Smith (eds.), Geography and Empire. Oxford: Blackwell, 1994. Pp. xii + 402. ISBN 0-631-19384-7, £50.00, $64.95 (hardcover); 0-631-19385-5, £15.99, $24.95 (paperback). British Journal for the History of Science, 29 (1):114–116, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027533.

Harrison:1996:BRR

[Har96d]  Mark Harrison.  Book re-

view: Richard H. Grove, Green Imperialism: Colonial Expansion, Tropical Island Edens and the Origins of Environmentalism, 1600–1860. Studies in Environmental History. Cambridge: Cambridge University Press, 1995. Pp. xiv + 540. ISBN 0-521-40385-5. £45.00, $64.95. British Journal for the History of Science, 29 (3):369–372, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027747.

Harrison:1997:BRD

[Har97]  Mark Harrison.  Book re-

view: Deepak Kumar, Science and the Raj, 1857–1905. New Delhi: Oxford University Press, 1995. Pp. xv + 273. ISBN 0-19-563562-0. 425 Rupees. British Journal for the History of Science, 30(1): 101–121, March 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027906.

Harley:1998:BRA

[Har98a] David Harley. Book review: Andrew Cunningham, The Anatomical Renaissance: The Resurrection of the Anatomical Projects of the Ancients. Aldershot: Scolar Press, 1997. Pp. xiv + 283. ISBN 1-85928-338-1. £45.00. British Journal for the History of Science, 31(4): 469–487, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027879.

Harley:1998:BRC

[Har98b] David Harley. Book review: Clara Pinto-Correia, The Ovary of Eve: Egg and Sperm and Preformation. Chicago: University of Chicago Press, 1997. Pp. xxiii + 396. ISBN 0-226-66952-1. $29.95. British Journal for the History of Science, 31(4): 469–487, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027880.

Harley:1998:BRM

[Har98c] David Harley. Book review: Mary Lindemann, Health and Healing in Eighteenth-Century Germany. Baltimore: Johns Hopkins University Press, 1996. Pp. xiii + 506. ISBN 0-8018-5281-1. £41.50. British Journal for the History of Science, 31(4): 469–487, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027881.

Harman:1998:BRC

[Har98d] P. M. Harman. Book review: Chloe Chard and Helen Langdon (eds.), Transports. Travel, Pleasure, and Imaginative Geography, 1600–1830. Studies in British Art, 3. Published for the Paul Mellon Center for Studies in British Art and the Yale Center for British Art. New Haven: Yale University Press, 1996. Pp. viii + 341, illus. ISBN 0-300-06382-2. £35.00. British Journal for the History of Science, 31 (2):241–250, June 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027768.

Harman:1998:NPJ

[Har98e] P. M. (Peter Michael) Harman. The natural philosophy of James Clerk Maxwell. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-56102-7, 0-521-00585-X (paperback). xiv + 232 pp. LCCN QC16.M4 H37

1998. URL http://www.loc.
gov/catdir/description/
cam028/97035703.html;
http://www.loc.gov/
catdir/toc/cam021/
97035703.html;   http://
www.zentralblattmath.org/
zmath/en/search/?an=1005.
01006.

Harrison:1998:BPR

[Har98f] Peter Harrison. The Bible, Protestantism, and the rise of natural science. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-59196-1 (hardcover). xi + 313 pp. LCCN BS650 .H37

1998. URL http://www.loc.
gov/catdir/description/
cam028/97020967.html;

http://www.loc.gov/
catdir/enhancements/
fy0732/97020967-b.html;
http://www.loc.gov/
catdir/toc/cam023/
97020967.html.

Harvey:1998:BRA

[Har98g] Joy Harvey. Book review:

Ann B. Shteir, Cultivating Women, Cultivating Science: Flora’s Daughters and Botany in England 1760 to 1860. Baltimore: Johns Hopkins University Press, 1996. Pp. xi + 301, illus. ISBN 0-8018-5141-6. £25.00. British Journal for the History of Science, 31(1): 63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027785.

Harwood:1998:BRM

[Har98h] Jonathan Harwood. Book review: Mitchell G. Ash and Alfons Söllner (eds.), Forced Migration and Scientific Change: Emigré German-Speaking Scientists and Scholars after 1933. Cambridge: Cambridge University Press, 1996. Pp. xviii + 301. ISBN 0-521-49741-8. £35.00, $59.95. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027786.

Harley:1999:BREb

[Har99a] David Harley. Book review:

Edward Fenton (ed.), The Diaries of John Dee. Charlbury, Oxfordshire: Day Books, 1998. Pp. xi + 369. ISBN 0-9532213-0-X. £18.99. British Journal for the History of Science, 32 (3):363–378, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027997.

Harley:1999:BREa

[Har99b] David Harley. Book review: Elizabeth Fee and Theodore M. Brown (eds.), Making Medical History: The Life and Times of Henry E. Sigerist. Baltimore: Johns Hopkins University Press, 1997. Pp. xii + 387. ISBN 0-8018-5355-9. £33.00. British Journal for the History of Science, 32(1): 111–124, March 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027973.

Hatfield:1990:NNT

[Hat90] Gary C. (Gary Carl) Hatfield. The natural and the normative: theories of spatial perception from Kant to Helmholtz. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-08086-9. xii + 366 pp. LCCN BF469 .H34 1990.

Haynes:1994:FSR

[Hay94]  Roslynn  D.

(Roslynn Doris) Haynes. From Faust to Strangelove: representations of the scientist in western literature. The Johns Hopkins University Press, Baltimore, MD, USA, 1994. ISBN 0-8018-4801-6 (hardcover), 0-8018-4983-7 (paperback). 417 pp. LCCN PN56.5.S35 H39

1994. URL http://www.loc.
gov/catdir/bios/jhu052/
94007023.html;
http://www.loc.gov/
catdir/description/
jhu051/94007023.html.

Hayward:1998:BRG

[Hay98a] Rhodri Hayward. Book review: German Berrios and Roy Porter (eds.), A History of Clinical Psychiatry: The Origin and History of Psychiatric Disorders. London: Athlone Press, 1995. Pp. xx + 684. ISBN 0-485-24011-4, £60.00 (hardcover); 0-485-24211, £19.95 (paperback). British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027787.

Hayward:1998:BRS

[Hay98b] Rhodri Hayward. Book review:

Sally Shuttleworth, Charlotte Brontë and Victorian Psychology. Cambridge: Cambridge University Press, 1996. Pp. xiv + 289. ISBN 0-521-55149-8. £37.50. British Journal for the History of Science, 31(4): 469–487, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027882.

Hutton:1990:HMT

[HC90] Sarah Hutton and Robert Crocker, editors. Henry More (1614– 1687) tercentenary studies, volume 127 = of Archives internationales d’histoire des idées; International archives of the history of ideas; 127. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0095-5. xi + 254 pp. LCCN B1299.M64 H46 1990.

Hall:1991:MJR

[HC91]  A.  Ru-

pert Hall and A. C. Crombie. In memoriam: Jacques Roger (1920–1990). British Journal for the History of Science, 24 (2):231–233, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027167.

Hunter:1997:BRE

[HD97] Michael Hunter and Malcolm De Mowbray. Book review: Essay review The editor in the republic of letters Eric G. Forbes, Lesley Murdin and Francis Willmoth (eds.), The Correspondence of John Flamsteed, First Astronomer Royal. Volume 1: 1666–1682. Bristol and Philadelphia: Institute of Physics Publishing, 1995. Pp. xlix + 955. ISBN 0-7503-0147-3. £140.00, $280.00. Heinz-Jurgen Hess, James G. O’Hara and Herbert Breger (eds.), Gottfried Wilhelm Leibniz: Sämtliche Schriften und Briefe. Dritte Reihe, Mathematischer, naturwissenschaftlicher und technischer Briefwechsel: Volume 3, 1680–1683; Volume 4, 1683– 1690. Berlin: Akademie Verlag, 1991, 1995. Pp. lxx+895; lxvi+747. ISBN 3-05-000766-4, DM 490.00 (Volume 3); 3-05-002602-2, DM 490.00 (Volume 4) (series ISBN: 3-05-000075-9). Wilhelm Schmidt-Biggemann (ed.), Samuel Pufendorf. Gesammelte Werke, Band 1: Briefwechsel (ed. Detlef Döring). Berlin: Akademie Verlag, 1996. Pp. xxix+453. ISBN 3-05-001920-4. DM 298.00. British Journal for the History of Science, 30 (2):221–225, June 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027718.

Hammond:1992:WBH

[HE92] P. W. Hammond and Harold Egan. Weighed in the balance: a history of the Laboratory of the Government Chemist. Her Majesty’s Stationary Office, London, England, 1992. ISBN 0-11-515302-0. 372 pp. LCCN QD51.5.G72 L334 1992. US$9.95.

Hearnshaw:1986:ASO

[Hea86] J. B. Hearnshaw. The analysis of starlight: one hundred and fifty years of astronomical spectroscopy. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25548-1. xv + 531 pp. LCCN QB465 .H43 1986.

Hearnshaw:1987:SMP

[Hea87] L. S. (Leslie Spencer) Hearnshaw. The shaping of modern psychology. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1987. ISBN 0-7102-0576-7. viii + 423 pp. LCCN BF81 .H33 1987. US$16.95 (U.K.), (US$21.95 U.S.).

Hearnshaw:1990:ASO

[Hea90] J. B. Hearnshaw. The analysis of starlight: one hundred and fifty years of astronomical spectroscopy. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-39916-5, 0-521-25548-1. xv + 531 pp. LCCN QB465 .H43

                                        1990. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/91105528.html.

Heffernan:1996:BRM

[Hef96] Michael Heffernan. Book review: Michael A. Osborne, Nature, the Exotic, and the Science of French Colonialism. Bloomington and Indianapolis: Indiana University Press, 1994. Pp. xvi + 216. ISBN 0-253-34266-X. £32.50. British Journal for the History of Science, 29(2):242–244, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027846.

Heilbron:1986:DUM

[Hei86] J. L. Heilbron. The dilemmas of an upright man: Max Planck as spokesman for German science. University of California Press, Berkeley, CA, USA, 1986. ISBN 0-520-05710-4. xiii + 238 + 16 pp. LCCN QC16.P6 H45 1986.

Hein:1988:AHP

[Hei88]  Wolfgang  Ha-

gen Hein. Alexander von Humboldt und die Pharmazie. (German) [Alexander von Humboldt and Pharmacy], volume 56 of Veröffentlichungen der Internationalen Gesellschaft für Geschichte der Pharmazie e. V., 0074-9729. Wissenschaftliche Verlagsgesellschaft, Stuttgart, Germany, 1988. ISBN 3-8047-0984-2. 130 pp. LCCN Q143.H9 H38 1988.

Heims:1991:CG

[Hei91] Steve J. Heims. The cybernetics group. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-08200-4. xii + 334 pp. LCCN H62.5.U5 H45 1991.

Heidelberger:1993:ISN

[Hei93] Michael Heidelberger. Die innere Seite der Natur: Gustav Theodor Fechners wissenschaftlichphilosophische Weltauffassung. (German) [The internal side of nature: Gustav Theodor Fechners scientific-philosophical view of the world], volume 60 of Philosophische Abhandlungen. V. Klostermann, Frankfurt am Main, Germany, 1993. ISBN 3-465-02590-3. 457 + 1 pp. LCCN B3237 .H45 1993.

Hellyer:1996:BRA

[Hel96a] Marcus Hellyer. Book review: Albert Krayer, Mathematik im Studienplan der Jesuiten: Die Vorlesung von Otto Cattenius an der Universität Mainz (1610/11). Beiträge zur Geschichte der Universität Mainz, 15. Stuttgart: Franz Steiner Verlag, 1991. Pp. ix + 434. ISBN 3-515-05874-5. DM 98.00. British Journal for the History of Science, 29 (2):232–233, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027839.

Hellyer:1996:BRO

[Hel96b] Marcus Hellyer. Book review: Otto Von Guericke, The New (So-Called) Magdeburg Experiments of Otto von Guericke, translated by Margaret Glover Foley Ames. Dordrecht: Kluwer Academic Publishers, 1994. Pp. xxiv + 394, illus. ISBN 0-7923-2399-8. £112.00, $170.00, Dfl. 280.00. British Journal for the History of Science, 29(3):365–366, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027743.

Hempstead:1995:BRH

[Hem95] Colin A. Hempstead. Book review: Henry Petroski, Design Paradigms: Case Histories of Error and Judgement in Engineering. Cambridge: Cambridge University Press, 1994. Pp. xii + 209. ISBN 0-521-46108-1, £30.00, $42.95 (hardcover); 0-521-46649-0, £12.95, $17.95 (paperback). British Journal for the History of Science, 28(3):374–375, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027667.

Hendry:1984:CPT

[Hen84a] John Hendry, editor. Cambridge physics in the thirties. Adam Hilger Ltd., Bristol, UK, 1984. ISBN 0-85274-761-6. 209 pp. LCCN QC51.G72 C352 1984.

Hendry:1984:CQM

[Hen84b] John Hendry. The creation of quantum mechanics and the Bohr–Pauli dialogue, volume 14 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1648-X. xi + 177 pp. LCCN QC173.98 .H46 1984.

Hendry:1986:JCM

[Hen86a] John Hendry. James Clerk Maxwell and the theory of the electromagnetic field. Adam Hilger Ltd., Bristol, UK, 1986. ISBN 0-85274-563-X. xix + 305 pp. LCCN QC665.E4 H46

                                        1986. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0745/86196851-d.html.

Heniger:1986:HAV

[Hen86b] J. Heniger. Hendrik Adriaan van Reede tot Drakenstein (1636–1691) and Hortus Malabaricus: a contribution to the history of Dutch colonial botany. A. A. Balkema, Rotterdam, The Netherlands, 1986. ISBN 90-6191-681-X. xvi + 295 pp. LCCN QK358.R43 H46

                                        1986. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0647/87170134-d.html.

Hendry:1989:IFG

[Hen89] John Hendry. Innovating for failure: government policy and the early British computer industry. History of computing. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-08187-3. xvi + 240 pp. LCCN HD9696.C63 G763 1989.

Hendry:1990:BRD

[Hen90a] John Hendry. Book review: Dean Keith Simonton. Scientific Genius: A Psychology of Science. Cambridge: Cambridge University Press, 1989.

Pp. x + 229. ISBN 0-521-35287-8. £22.50. British Journal for the History of Science, 23 (2):246–247, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026746.

Hendry:1990:BRR

[Hen90b] John Hendry. Book reviews:

Roland Turner (ed.). Thinkers of the Twentieth Century. 2nd edition. Chicago and London: St James Press, 1987. Pp. vii and 977. ISBN 0-912289-83-X. £52.50. David, Ian, John & Margaret Millar. Chambers Concise Dictionary of Scientists. Cambridge and Edinburgh: CUP and Chambers, 1989. Pp. 461. ISBN 1-85296-354-9. £14.95. British Journal for the History of Science, 23 (2):228–229, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026735.

Henderson:1991:DBS

[Hen91a] Janice Adrienne Henderson.

On the distances between sun, moon, and Earth1 according to Ptolemy, Copernicus, and Reinhold, volume 30 of Studia Copernicana; Brill’s series, 0925-6806; vol. 1. E. J. Brill, Leiden, The Netherlands, 1991. ISBN 90-04-09378-8. xiv + 220 pp. LCCN QB991.C66 H46

Hendry:1991:BRH

[Hen91b]  John Hendry.  Book re-

view: Helge Kragh. Dirac: A Scientific Biography. Cambridge: Cambridge University Press, 1990. ISBN 0-521-38089-8. £35. British Journal for the History of Science, 24(1): 116, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027031.

Hendry:1992:BRA

[Hen92a] John Hendry. Book review: Abraham Pais. Niels Bohr’s Times, in Physics, Philosophy and Polity. Oxford: Oxford University Press, 1991. Pp. xvii + 565. British Journal for the History of Science, 25(4): 490–491, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027078.

Henry:1992:BRB

[Hen92b] John Henry. Book review: I. Bernard Cohen (ed.). Puritanism and the Rise of Modern Science: The Merton Thesis, edited with the assistance of K. E. Duffin and Stuart Strickland. New Brunswick and London: Rutgers University Press, 1990. Pp. xiii + 402. ISBN 0-8135-1529-7, $45.00 (hardcover); 0-8135-1530-0, $17.00 (paperback). British Journal for the History of Science, 25 (2):269–270, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027311.

Henry:1993:BRJ

[Hen93a] John Henry. Book review:

Jean S. Gottlieb, A Checklist of the Newberry Library’s Printed Books in Science, Medicine, Technology, and the Pseudosciences ca. 1460–1750. New York and London: Garland Publishing, Inc., 1992. Pp. xxiii + 312, illus. ISBN 0-8240-5171-8. $67.00. British Journal for the History of Science, 26(3): 353, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027404.

Henry:1993:BRM

[Hen93b] John Henry. Book review: Marjorie Hope Nicolson (ed.), The Conway Letters: The Correspondence of Anne, Viscountess Conway, Henry More, and their Friends, 1642–1684. Revised Edition with an Introduction and New Material Edited by Sarah Hutton. Oxford: Clarendon Press, 1992. Pp. xxix + 592. ISBN 0-19-824876-8. £55.00. British Journal for the History of Science, 26(3): 357–358, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027409.

Henry:1994:BRB

[Hen94a] John Henry. Book review: British Journal for the History of Philosophy, vol. 1, no. 1, edited by G. A. J. Rogers. Bristol: Thoemmes Press, 1993. Pp. 184. ISSN 0960-8788. Subscription: £28.00 per annum (individual), £36.00 (institutional). British Journal for the History of Science, 27(2): 243–244, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027457.

Henry:1994:BRM

[Hen94b] John Henry. Book reviews: Michael Hunter (ed.), Letters and Papers of Robert Boyle: A Guide to the Manuscripts and Microfilm. Collections from the Royal Society. Bethesda, Maryland: University Publications of America, 1992. Pp. xlix + 90. ISBN 1-55655-217-3. No price given. Peter Jones (ed.), Sir Isaac Newton: A Catalogue of Manuscripts and Papers Collected and Published on Microfilm by Chadwyck-Healey. Cambridge: Chadwyck-Healey, 1991. Pp. xi + 148. ISBN 0-85964-226-7. £50.00 (provided free with the microfilms). British Journal for the History of Science, 27 (1):115–116, March 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027586.

Henry:1995:BRT

[Hen95a] John Henry. Book review:

Toby E. Huff, The Rise of Early Modern Science: Islam, China and the West. Cambridge: Cambridge University Press, 1993. Pp. xiv + 409. ISBN 0-521-43496-3. £35.00, $54.95. British Journal for the History of Science, 28(1): 101–102, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027487.

Henry:1995:BRW

[Hen95b] John Henry. Book review:

W. F. Bynum and Roy Porter (eds.), Companion Encyclopaedia of the History of Medicine, 2 vols. London and New York: Routledge, 1993. Pp. xxvi + 1806. ISBN 0-415-04771-4. £150.00. British Journal for the History of Science, 28(3): 379–381, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027671.

Henry:1995:RU

[Hen95c] John Henry. Review: [untitled]. British Journal for the History of Science, 28(4):466– 468, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027553.

Henry:1996:BRE

[Hen96] John Henry. Book review: Edward Grant, Planets, Stars, and Orbs: The Medieval Cosmos, 1200–1687. Cambridge: Cambridge University Press, 1994. Pp. xxiii + 816. ISBN 0-521-43344-4. £45.00, $69.95. British Journal for the History of Science, 29(3):363–364, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027741.

Henry:1998:BRH

[Hen98a] John Henry. Book review: H. Floris Cohen, The Scientific Revolution: A Historiographical Inquiry. Chicago: University of Chicago Press, 1994. Pp. xviii + 662. ISBN 0-226-11279-9, £59.95, $75.00 (hardcover); 0-226-11280-2, £21.50, $26.95 (paperback). British Journal for the History of Science, 31 (3):361–375, September 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027821.

Henry:1998:BRJ

[Hen98b] John Henry. Book review: J. E. McGuire, Tradition and Innovation: Newton’s Metaphysics of Nature. The University of Western Ontario Series in the Philosophy of Science, 56. Dordrecht: Kluwer Academic, 1995. Pp. xvi + 290. ISBN 0-7923-3617-8. £88.00, $140.00, Dfl. 195.00. British Journal for the History of Science, 31(4): 469–487, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027883.

Henry:1999:BRN

[Hen99]  John Henry.  Book

review: Nancy G. Siraisi, The Clock and the Mirror: Girolamo Cardano and Renaissance Medicine. Princeton: Princeton University Press, 1997. Pp. xiv + 361. ISBN 0-691-01189-3. £37.50, $49.50. British Journal for the History of Science, 32 (1):111–124, March 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027974.

HerriesDavies:1983:SMC

[Her83] G. L. Herries Davies. Sheets of many colours: the mapping of Ireland’s rocks, 1750–1890, volume 4 of Historical studies in Irish science and technology. Royal Dublin Society, Dublin, Ireland, 1983. ISBN 0-86027-014-9. xiv + 242 + 4 pp. LCCN QE265 .H55 1983. US$12.00.

Herken:1988:UTM

[Her88] Rolf Herken, editor. The Universal Turing machine: a halfcentury survey. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-853741-7. xiv + 661 pp. LCCN QA267 .U55

                                        1988. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0636/88213652-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0636/88213652-t.html.

Herman:1990:FSE

[Her90] Robin Herman. Fusion: the search for endless energy. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-38373-0. xi + 267 pp. LCCN QC791.73 .H47

                                        1990. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/90001445.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0827/90001445-b.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam021/
                                        90001445.html.

Herbert:1991:CDP

[Her91] Sandra Herbert. Charles Darwin as a prospective geological author. British Journal for the History of Science, 24 (2):159–192, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027165.

HerriesDavies:1995:NHY

[Her95] G. L. Herries Davies. North from the Hook: 150 years of the Geological Survey of Ireland. Geological Survey of Ireland, Dublin, Ireland, 1995. ISBN 1-899702-00-8. xi + 342 pp. LCCN QE265 .H54 1995.

Hessenbruch:1995:BRH

[Hes95a] Arne Hessenbruch. Book review: Helmuth Albrecht (ed.), Naturwissenschaft und Technik in der Geschichte: 25 Jahre Lehrstuhl für Geschichte der Naturwissenschaft und Technik am Historischen Institut der Universität Stuttgart. Stuttgart: Verlag für Geschichte der Naturwissenschaften und Technik, 1993. Pp. 402. ISBN 3-928186-15-9. DM 40.00. British Journal for the History of Science, 28(3): 362–364, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027660.

Hessenbruch:1995:BRS

[Hes95b] Arne Hessenbruch. Book review: Svante Lindqvist (ed.). Center on the Periphery: Historical Aspects of 20th-century Swedish Physics. Canton, MA: Science History Publications, 1993. Pp. iv + 516. ISBN 0-88135-157-1. $49.95. British Journal for the History of Science, 28(2):246–247, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027691.

Hessenbruch:1999:PHB

[Hes99]  Arne  Hessen-

bruch. Probing historiographical boundaries: analysing contexts employing specific narratives. British Journal for the History of Science, 32(3):345– 358, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027992.

Hetherington:1988:SOE

[Het88] Norriss S. Hetherington. Science and objectivity: episodes in the history of astronomy. Iowa State University Press, Ames, IA, USA, 1988. ISBN 0-8138-1159-7. xii + 168 pp. LCCN QB32 .H48 1988.

Hetherington:1995:BRE

[Het95] Norriss S. Hetherington. Book review: Erich Robert Paul, The Milky Way Galaxy and Statistical Cosmology, 1890– 1924. Cambridge and New York: Cambridge University Press, 1993. Pp. xiv + 262. ISBN 0-521-35363-7. £30.00, $44.95. British Journal for the History of Science, 28(3):361– 362, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027658.

Heyck:1981:TIL

[Hey81] Thomas William Heyck. The transformation of intellectual life in Victorian England. Croom Helm studies in society and history. Croom Helm, London, UK, 1981. ISBN 0-7099-1206-4. 256 pp. LCCN DA533.

Heyd:1983:BOE

[Hey83] Michael Heyd. Between orthodoxy and the Enlightenment: Jean-Robert Chouet and the introduction of Cartesian science in the Academy of Geneva, volume 96 of Archives internationales d’histoire des idées; International archives of the history of ideas; 96. Martinus Nijhoff, Dordrecht and The Hague, The Netherlands, 1983. ISBN 90-247-2508-9. xii + 308 pp. LCCN B4651.C474 H49 1983.

Heyd:1995:SRC

[Hey95] Michael Heyd. Be sober and reasonable: the critique of enthusiasm in the Seventeenth and early Eighteenth Centuries, volume 63 of Brill’s studies in intellectual history, 0920-8607. E. J. Brill, Leiden, The Netherlands, 1995. ISBN 90-04-10118-7 (hardcover). x + 312 pp. LCCN BR112 .H45 1995.

Herz-Fischler:1987:MHD

[HF87] Roger Herz-Fischler. A mathematical history of division in extreme and mean ratio. Wilfrid Laurier University Press, Waterloo, Ont., Canada, 1987. ISBN 0-88920-152-8. xvi + 191 pp. LCCN QA481 .H47 1987. US$65.00.

Hasan:1986:ITI

[HH86] Ahmad Yusuf Hasan and Donald Routledge Hill. Islamic technology: an illustrated history. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-26333-6 (Cambridge), 92-3-102294-6 (UNESCO). xiv + 304 pp. LCCN ????

Hassam:1992:ITI

[HH92] Ahmad Yusuf Hassam and Donald Routledge Hill. Islamic technology: an illustrated history. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-42239-6 92-3-102294-6 (UNESCO). xv + 304 pp. LCCN ????

Hunter:1997:WSM

[HH97] Lynette Hunter and Sarah Hutton, editors. Women, science and medicine 1500–1700: mothers and sisters of the Royal Society. Sutton Pub., Thrupp, Stroud, Gloucestershire, 1997. ISBN 0-7509-1334-7 (hardcover), 0-7509-1343-6 (paperback). xx + 292 pp. LCCN Q41.L85 W66 1997.

Higginson:1994:BRT

[Hig94] Ian Higginson. Book review: Trevor H. Levere, Science and the Canadian Arctic: A Century of Exploration, 1818–1918. Cambridge: Cambridge University Press, 1993. Pp. xiv + 438. ISBN 0-521-41933-6. £40.00, $64.95. British Journal for the History of Science, 27 (3):376–377, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027615.

Hill:1984:HEC

[Hil84] Donald Routledge Hill. A history of engineering in classical and medieval times. Croom Helm, London, UK, 1984. ISBN 0-7099-1209-9. 263 pp. LCCN TA16 .H55x 1984b.

Hill:1991:BRC

[Hil91a]  Donald Hill.  Book

review: Charles Michael Stanton. Higher Learning in Islam: The Classical Period A.D. 700-1300. Savage, Maryland: Rowman & Littlefield, 1990. Pp. xiii + 205. ISBN 0-8476-7637-4. $42.50. British Journal for the History of Science, 24(3): 373–374, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027237.

Hill:1991:BRB

[Hil91b] Donald R. Hill. Book review: Brian Cotterell and Johan Kamminga. Mechanics of Pre-Industrial Technology. Cambridge: Cambridge University Press, 1990. Pp. xv + 325. ISBN 0-521-34194-9. £30.00, $49.50. British Journal for the History of Science, 24(2):245–246, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027170.

Hill:1991:BRE

[Hil91c] Donald R. Hill. Book review: Elspeth Whitney. Paradise Restored: The Mechanical Arts from Antiquity through the Thirteenth Century. Transactions of the American Philosophical Society, vol. 80, Part 1. Philadelphia: The American Philosophical Society, 1990. Pp. vi + 169. ISBN 0-87169-801-3. $20.00 (paperback). British Journal for the History of Science, 24(1): 101–102, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027020.

Hill:1993:BRS

[Hil93] Donald R. Hill. Book review: Sa’Id Ibn Ahmad Al-Andalusi, Science in the Medieval World: Book of the Categories of Nations, translated and edited by Sema’an I. Salem and Alok Kumar. Austin: University of Texas Press, 1992. Pp. xxvi + 118. ISBN 0-292-71339-5. $29.00. British Journal for the History of Science, 26(1):80, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027343.

Hills:1994:PWH

[Hil94] Richard Leslie Hills. Power from wind: a history of windmill technology. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-41398-2 (hardcover). 324 pp. LCCN TJ823 .H55 1994. URL http:/

/www.loc.gov/catdir/
description/cam025/
93008858.html;   http://
www.loc.gov/catdir/toc/
cam021/93008858.html.

Hill:1998:BRC

[Hil98a] Katherine Hill. Book review:

Catherine Goldstein, Jeremy Gray and Jim Ritter (eds.), L’Europe mathématique: Histoires, mythes, identités/ Mathematical Europe: History, Myth, Identity. Paris: Edition de la maison des sciences de l’homme, 1996. Pp. x + 575. ISBN 2-7351-0685-3. 190F. British Journal for the History of Science, 31(4): 469–487, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027884.

Hill:1998:JSN

[Hil98b] Katherine Hill. ‘juglers or schollers?’: Negotiating the role of a mathematical practitioner. British Journal for the History of Science, 31(3): 253–274, September 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027813.

Hine:1991:BRN

[Hin91] Harry M. Hine. Book review:

Nikolaus Gross. Senecas Naturales Quaestiones: Komposition, naturphilosophische Aussagen und ihre Quellen. Palingenesia: Monographien und Texte zur klassischen Altertumswissenschaft,

27. Stuttgart: Franz Steiner, 1989. Pp. viii + 335. ISBN 3-515-05385-9. DM 68. British Journal for the History of Science, 24(2):246–247, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027171.

Hirsh:1983:GIU

[Hir83] Richard F. Hirsh. Glimpsing an invisible universe: the emergence of X-ray astronomy. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25121-4. viii + 186 pp. LCCN QB472 .H57

1983. URL http://www.loc.
gov/catdir/description/
cam031/82023633.html;
http://www.loc.gov/
catdir/enhancements/
fy0906/82023633-t.html.

Hiromasa:1995:BRH

[Hir95] Naohiko Hiromasa. Book review: Hideo Yoshikawa and Joanne Kauffman. Science has no National Borders: Harry C. Kelly and the Reconstruction of Science in Postwar Japan. Cambridge, MA and London: MIT Press, 1994. Pp. xvii + 137. ISBN 0-262-24037-8. £19.95, $29.95. British Journal for the History of Science, 28 (3):368–369, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027662.

Hulin-Jung:1989:LES

[HJF89] Nicole Hulin-Jung and Robert Fox. L’organisation de l’enseignement des sciences: la voie ouverte par le Second Empire. (French) [The organization of teaching of science: the way opened by the Second Empire], volume 6 of Mémoire de la section d’histoire des sciences et techniques. Ed. du Comité des travaux historiques et scientifiques, Paris, France, 1989. ISBN 2-7355-0171-X. ISSN 0769-2218. 335 pp. LCCN Q183.4.F8 H85

1989. URL http://catalog.
hathitrust.org/api/
volumes/oclc/21276320.
html.

Harrower:1983:KKU

[HK83] Molly Harrower and Kurt Koffka. Kurt Koffka, an unwitting self-portrait. University Presses of Florida, Gainesville, FL, USA, 1983. ISBN 0-8130-0760-7. xvi + 334 pp. LCCN BF109.K59 H37 1983.

Herrmann:1984:HAH

[HK84] Dieter B. Herrmann and Kevin Krisciunas. The history of astronomy from Herschel to Hertzsprung. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25733-6. x + 220 pp. LCCN QB28 .H4713

Hindle:1986:ECA

[HL86] Brooke Hindle and Steven D. Lubar. Engines of change: the American industrial revolution, 1790–1860. Smithsonian Institution Press, Washington, DC, USA, 1986. ISBN 0-87474-540-3, 0-87474-539-X (paperback). 309 pp. LCCN HC105 .H66 1986.

Hoch:1990:BRS

[Hoc90a]  Paul K. Hoch.  Book re-

view: Spencer R. Weart. Nuclear Fear: A History of Images. Cambridge, Mass, and London: Harvard University Press, 1988. Pp. xx + 535. ISBN 0-674-62835-7. £23.50. British Journal for the History of Science, 23(3):352–353, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026775.

Hoch:1990:RU

[Hoc90b] Paul K. Hoch. Review: [untitled]. British Journal for the History of Science, 23(4): 470–472, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027207.

Hoch:1991:BRS

[Hoc91] Paul K. Hoch. Book review: Stuart Peterfreund (ed.). Literature and Science: Theory and Practice. Boston: Northeastern University Press, 1990.

Pp. vi + 248. ISBN 1-55553-058-3. £33.65. British Journal for the History of Science, 24 (1):123–124, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027039.

Hoch:1992:BRMb

[Hoc92a] Paul K. Hoch. Book review:

Maila L. Walter. Science and Cultural Crisis: An Intellectual Biography of Percy Williams Bridgman (1882–1961). Stanford: Stanford University Press, 1990. Pp. x + 362. ISBN 0-8047-1796-6. $42.50. British Journal for the History of Science, 25(4): 494–495, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027083.

Hoch:1992:BRMa

[Hoc92b] Paul K. Hoch. Book review:

Michael Eckert and Helmut Schubert. Crystals, Electrons, Transistors: From Scholar’s Study to Industrial Research, translated by Thomas Hughes. New York: American Institute of Physics, 1990. Pp. xxii + 241. ISBN 0-88318-622-5, £45 (hardcover); 0-88318-719-1, £15 (paperback). British Journal for the History of Science, 25(2):288–289, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027326.

Hoch:1992:BRS

[Hoc92c]  Paul K. Hoch.  Book re-

view: Shigeru Nakayama. Science, Technology and Society in Postwar Japan. London and New York: Kegan Paul International, 1991. Pp. xvi + 259. ISBN 0-7103-0428-5. £45.00. British Journal for the History of Science, 25(3): 380–381, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027281.

Hoch:1992:BRT

[Hoc92d] Paul K. Hoch. Book review: Tom Wilkie. British Science and Politics since 1945. Oxford: Basil Blackwell, 1991. Pp. x + 142. ISBN 0-631-16849-4, £24.95 (hardcover); 0-631-16851-6, £8.95 (paperback). British Journal for the History of Science, 25(4):496– 497, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027085.

Hoch:1993:BRE

[Hoc93a]  Paul K. Hoch.  Book re-

view: Elisabeth Crawford, Nationalism and Internationalism in Science, 1880–1939: Four Studies of the Nobel Population. Cambridge: Cambridge University Press, 1992. Pp. xii + 157. ISBN 0-521-40386-3. £27.95, $44.95. British Journal for the History of Science, 26 (1):121–122, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027383.

Hoch:1993:BRP

[Hoc93b]  Paul K. Hoch.  Book re-

view: Peter Galison and Bruce Hevly (eds.), Big Science: The Growth of Large-Scale Research. Stanford: Stanford University Press, 1992. Pp. xi + 392, illus. ISBN 0-8047-1879-2. $45.00. British Journal for the History of Science, 26 (2):259–260, June 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027121.

Hoch:1994:WSI

[Hoc94] Paul K. Hoch. Whose scientific internationalism? British Journal for the History of Science, 27(3):345–349, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027602.

Hoddeson:1990:BRH

[Hod90a] Lillian Hoddeson. Book review:

Hans Queisser. The Conquest of the Microchip: Science and Business in the Silicon Age. Cambridge, Massachusetts, London, England: Harvard University Press, 1988. Translated by Diane Crawford-Burkhardt. Pp. x + 200. ISBN 0-674-16296-X, £19.95. British Journal for the History of Science, 23(1): 104–105, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026810.

Hodge:1990:BRP

[Hod90b] M. J. S. Hodge. Book review: Peter J. Bowler. The Non-Darwinian Revolution: Reinterpreting a Historical Myth. Baltimore and London: The Johns Hopkins University Press, 1988. Pp. x + 238. ISBN 0-8018-3678-6. £17.50. British Journal for the History of Science, 23(3): 331–334, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026758.

Hodge:1991:BRG

[Hod91] M. J. S. Hodge. Book review: Gerd Gigerenzer, Zeno Swijtink, Theodore Porter, Lorraine Daston, John Beatty and Lorenz Kruger. The Empire of Chance. How Probability Changed Science and Everyday Life. Cambridge: Cambridge University Press, 1989. Pp. xviii + 340. ISBN 0-521-33115-3. £32.50. British Journal for the History of Science, 24(1): 124–126, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027040.

Hodge:1992:BRA

[Hod92a] M. J. S. Hodge. Book review:

Adrian Desmond. The Politics of Evolution. Morphology, Medicine, and Reform in Radical London. Chicago and London: The University of Chicago Press, 1989. Pp. x + 503. ISBN 0-226-14346-5. £27.95. British Journal for the History of Science, 25(2):275–278, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027316.

Hodgkin:1992:CDR

[Hod92b] A. L. (Alan Lloyd) Hodgkin.

Chance and design: reminiscences of science in peace and war. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40099-6. xi + 412 pp. LCCN QP355.2 .H63

1992. URL http://www.loc.
gov/catdir/description/
cam024/91027737.html;
http://www.loc.gov/
catdir/toc/cam028/
91027737.html.

Hodge:1993:BRF

[Hod93a] M. J. S. Hodge. Book review:

Fred Wilson, Empiricism and Darwin’s Science. Dordrecht: Kluwer Academic Publishers, 1991. Pp. xiv + 358. ISBN 0-7923-1019-5. £61.00, $99.00, Dfl 175.00. British Journal for the History of Science, 26(1): 104–105, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027370.

Hodge:1993:BRM

[Hod93b] M. J. S. Hodge. Book review: Mario A. Di Gregorio, Charles Darwin’s Marginalia, Volume 1, with the assistance of N. W. Gill, New York: Garland Publishing, 1990. Pp. lxi + 448 + 275. ISBN 0-8240-6639-1. $95.00. British Journal for the History of Science, 26 (1):105–106, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027371.

Hodge:1996:BRF

[Hod96] Jon Hodge. Book review: Frederick Burkhardt and Sydney Smith, A Calendar of the Correspondence of Charles Darwin, 1821–1882, with Supplement. Cambridge: Cambridge University Press, 1994. Pp. vii + 690 + 49. ISBN 0-521-43423-8. £95.00, $150.00. British Journal for the History of Science, 29(3):374–375, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027750.

Holmes:1985:LCL

[Hol85]  Frederic  Lawrence

Holmes. Lavoisier and the chemistry of life: an exploration of scientific creativity, volume 4 of Wisconsin publications in the history of science and medicine. University of Wisconsin Press, Madison, WI, USA, 1985. ISBN 0-299-09980-6, 0-299-09984-9 (paperback). xxiv + 565 pp. LCCN QP511.8.L38 HOL 1985; 920; QP.

Holton:1986:ASB

[Hol86] Gerald James Holton. The advancement of science, and its burdens: the Jefferson Lecture and other essays. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25244-X, 0-521-27243-2 (paperback). xi + 351 pp. LCCN Q173 .H734

1986. URL http://www.loc.
gov/catdir/description/
cam032/86009722.html.

Home:1983:SUS

[Hom83] Roderick Weir Home, editor.

Science under scrutiny: the place of history and philosophy of science, volume 3 of Australasian studies in history and philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1983. ISBN 90-277-1602-1. xvii + 182 pp. LCCN Q175.3 .S323

Hooykaas:1997:RBS

[Hoo97] R. (Reijer) Hooykaas. Robert Boyle: a study in science and Christian belief. Pascal Centre for Advanced Studies in Faith and Science, Redeemer College, Ancaster, Ont., 1997. ISBN 0-7618-0708-X (hardcover). xxiv + 131 pp. LCCN BT1117 .H66 1997.

Horrocks:1992:BRR

[Hor92] Sally M. Horrocks. Book review: Robert Clayton and Joan Algar (eds.). A Scientist’s War: The War Diary of Sir Clifford Paterson, 1939– 45. London: Peter Peregrinus Ltd in association with the Science Museum, 1991. Pp. xxvi + 674. ISBN 0-86341-218-1. £59.00. British Journal for the History of Science, 25(4): 492, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027080.

Horrocks:1995:BRH

[Hor95] Sally M. Horrocks. Book review: Howard P. Segal, Future Imperfect: The Mixed Blessings of Technology in America. Amherst: University of Massachusetts Press, 1994. Pp. xviii + 245. ISBN 0-87023-881-7, £38.00 (hardcover); 0-87023-882-5, £15.20 (paperback). British Journal for the History of Science, 28(3):373– 374, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027666.

Horrocks:1996:BRM

[Hor96a] Sally M. Horrocks. Book review: M. Bauer (ed.), Resistance to New Technology: Nuclear Power, Information Technology and Biotechnology. Cambridge: Cambridge University Press, 1995. Pp. xii + 422. ISBN 0-521-45518-9. £50.00, $74.95. British Journal for the History of Science, 29(4):490–491, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027709.

Horrocks:1996:BRT

[Hor96b] Sally M. Horrocks. Book review: Thomas J. Misa, A Nation of Steel: The Making of Modern America, 1865–1925. Baltimore and London: Johns Hopkins University Press, 1995, Pp. xxvi + 367. ISBN 0-8018-4967-5. $42.50. British Journal for the History of Science, 29(1): 116–117, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027534.

Horrocks:1997:BRK

[Hor97a] Sally M. Horrocks. Book review: Kenneth Blaxter and Noel Robertson, From Dearth to Plenty: The Modern Revolution in Food Production. Cambridge: Cambridge University Press, 1995. Pp. xvii + 296. ISBN 0-521-40322-7. £40.00, $59.95. British Journal for the History of Science, 30(2):233–249, June 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027723.

Horrocks:1997:BRM

[Hor97b] Sally M. Horrocks. Book review: M. Bauer (ed.), Resistance to New Technology: Nuclear Power, Information Technology and Biotechnology. Cambridge: Cambridge University Press, 1995. Pp. xii + 422. ISBN 0-521-45518-9. £50.00, $74.95. British Journal for the History of Science, 30(2):233–249, June 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027724.

Horrocks:1998:BRL

[Hor98] Sally M. Horrocks. Book review: Lance Day and Ian Mc-Neil (eds.), Biographical Dictionary of the History of Technology. London: Routledge, 1996. Pp. xiii + 844. ISBN 0-415-06042-7. £85.00. British Journal for the History of Science, 31(1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027788.

Horrocks:1999:BRJ

[Hor99a] Sally M. Horrocks. Book review: Jeffrey P. Baker, The Machine in the Nursery: Incubator Technology and the Origins of Newborn Intensive Care. Baltimore: Johns Hopkins University Press, 1996. Pp. x + 247. ISBN 0-8018-5173-4. £37.00. British Journal for the History of Science, 32(1): 111–124, March 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027975.

Horrocks:1999:BRM

[Hor99b]  Sally  M.

Horrocks. Book review: Marelene Rayner-Canham and Geoffrey Rayner-Canham, Women in Chemistry: Their Changing Roles From Alchemical Times to the Mid-Twentieth Century. Washington, D.C.: American Chemical Society and Chemical Heritage Foundation, 1998. Pp. xiv + 284. ISBN 0-8412-3522-8. $34.95. British Journal for the History of Science, 32 (2):237–251, June 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4028088.

Howson:1982:HME

[How82] A. G. (Albert Geoffrey) Howson. A history of mathematics education in England. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24206-1. x + 294 pp. LCCN QA14.G7 H68. URL http://

                                        www.loc.gov/catdir/
                                        enhancements/fy0838/
                                        82004175-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0838/82004175-t.html.

Howse:1989:NMS

[How89] Derek Howse. Nevil Maskelyne, the seaman’s astronomer. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36261-X. xiv + 280 pp. LCCN QB36.M378 H69 1989.

Howarth:1996:SHT

[How96] Richard J. Howarth. Sources for a history of the ternary diagram. British Journal for the History of Science, 29(3): 337–356, September 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027737.

Hoyrup:1994:MNW

[Høy94] Jens Høyrup. In measure, number, and weight: studies in mathematics and culture. SUNY series in science, technology, and society. State University of New York Press, Albany, NY, USA, 1994. ISBN 0-7914-1821-9, 0-7914-1822-7 (paperback : acid-free). xviii + 430 pp. LCCN QA10.7 .H69 1994.

Hermann:1985:WPW

[HP85] Armin Hermann and Wolfgang Pauli, editors. Wolfgang Pauli, wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Bd. 2, 1930–1939. (German) [Wolfgang Pauli, scientific correspondence with Bohr, Einstein, Heisenberg and others, volume 2, 1930–1939], volume Vol. 6 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-13609-6 (Berlin), 0-387-13609-6 (New York). xxxix + 783 pp. LCCN ????

Hentschel:1990:KPR

[HRP90] Klaus Hentschel, Hans Reichenbach, and Joseph Petzoldt. Die Korrespondenz Petzoldt–Reichenbach: zur Entwicklung der “wissenschaftlichen Philosophie” in Berlin. (German) [The Petzoldt–Reichenbach correspondence: development “of scientific philosophy” in Berlin], volume 12 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. SIGMA, Berlin, Germany, 1990. ISBN 3-928068-03-2. 91 pp. LCCN Q175.3 .H46 1990.

Hollister-Short:1990:BRG

[HS90] Graham Hollister-Short. Book review: George Basalla. The Evolution of Technology. Cambridge and New York: Cambridge University Press, 1988. ISBN 0-521-29681-1. £8.95 (paperback). British Journal for the History of Science, 23(2): 239–241, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026742.

Hollister-Short:1991:BRG

[HS91] Graham Hollister-Short. Book review: Georg Agricola. Bermannus (Le Mineur): Un dialogue sur les mines, introduction, translation and commentary by Robert Halleux and Albert Yans. Paris: Les Belles Lettres, 1990. Pp. xxx + 185. ISBN 2-251-34504-3. British Journal for the History of Science, 24 (3):375–376, September 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027239.

Hankins:1995:II

[HS95]  Thomas  L.

Hankins and Robert J. Silverman. Instruments and the imagination. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-02997-0. xiv + 337 pp. LCCN Q185 .H25

1995. URL http://www.loc.
gov/catdir/description/
prin021/95004301.html;
http://www.loc.gov/
catdir/toc/prin031/
95004301.html.

Higginson:1999:MPT

[HS99]  Ian  Higginson

and Crosbie Smith. ‘A magnified piece of thermodynamics’: the Promethean iconography of the refrigerator in Paul Theroux’s the mosquito coast. British Journal for the History of Science, 32(3):325– 342, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027990.

Hall:1976:HT

[HSHSJ93] Alfred Rupert Hall, Norman A. F. Smith, Graham Hollister-Short, and Frank A. J. L. James, editors. History of technology. Mansell, London, UK, 1976–1993. ISBN 0-7201-1770-4 (vol. 10), 0-7201-1861-1 (vol. 11), 0-7201-2075-6 (vol. 12), 0-7201-2114-0 (vol. 13), 0-7201-2133-7 (vol. 14), 0-7201-2160-4 (vol. 15). ???? pp. LCCN ????

Hubner:1990:THT

[Hüb90] Jürgen Hübner, editor. Theologica, Hexenprozess, Tacitus-Übersetzung, Gedichte. (German) [Theologica, witch trials, Tacitus translation, Poems], volume 12 of Gesammelte Werke. Beck, München, Germany, 1990. ISBN 3-406-01660-X. 443 pp. LCCN ????

Hudson:1992:BRM

[Hud92] Gillian Hudson. Book review: Marilyn Bailie Ogilvie.

Women in Science: Antiquity through the Nineteenth Century. A Biographical Dictionary with Annotated Bibliography. Cambridge, Mass, and London: The MIT Press, 1988. Pp. xiii + 254. ISBN 0-262-15031-X. £10.95. British Journal for the History of Science, 25(2):292–294, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027330.

HuertasGarcia-Alejo:1988:OSP

[Hue88] Rafael Huertas García-Alejo. Orfila, saber y poder medico. (Spanish) [Orfila, knowledge and medical power], volume 2 of Estudios sobre la ciencia. Consejo Superior de Investigaciones Cientificas, Madrid, Spain, 1988. ISBN 84-00-06745-2. 456 pp. LCCN MLCS 93/04632 (R).

Hufbauer:1982:FGC

[Huf82] Karl Hufbauer. The formation of the German chemical community (1720–1795). University of California Press, Berkeley, CA, USA, 1982. ISBN 0-520-04318-9, 0-520-04415-0. viii + 312 pp. LCCN QD18.G3 H83 1982.

Hufbauer:1991:ESS

[Huf91]  Karl Hufbauer.  Exploring

the sun: solar science since Galileo. New series in NASA history. The Johns Hopkins University Press, Baltimore, MD, USA, 1991. ISBN 0-8018-4098-8. xv + 370 pp. LCCN QB521 .H93 1991.

Huff:1993:REM

[Huf93] Toby E. Huff. The rise of early modern science: Islam, China, and the West. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-43496-3 (hardcover). xiv + 409 pp. LCCN Q127.A5 H84 1993.

Huff:1995:REM

[Huf95] Toby E. Huff. The rise of early modern science: Islam, China, and the West. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-49833-3 (paperback), 0-521-43496-3 (hardcover). xiv + 409 pp. LCCN Q127.A5 H84 1995. URL http:

//www.loc.gov/catdir/
description/cam027/
95237814.html;   http://
www.loc.gov/catdir/toc/
cam025/95237814.html.

Hughes:1992:BRK

[Hug92a] David W. Hughes. Book review: Karl Hufbauer. Exploring The Sun: Solar Science since Galileo. Baltimore: Johns Hopkins University Press, 1991. Pp. xv + 370. ISBN 0-8018-4098-8. £28.50, $46.00. British Journal for the History of Science, 25 (3):383–384, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027285.

Hughes:1992:BRP

[Hug92b] David W. Hughes. Book review: Pamela E. Mack. Viewing the Earth: The Social Construction of the Landsat Satellite System. Cambridge Mass, and London: MIT Press, 1990. Pp. xii + 270. ISBN 0-262-13259-1. £24.75. British Journal for the History of Science, 25(4):499, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027088.

Hughes:1992:BRJ

[Hug92c] Jeff Hughes. Book review: J. W. Boag, P. W. Rubinin and D. Shoenberg (eds.). Kapitza in Cambridge and Moscow: Life and Letters of a Russian Physicist. Amsterdam: Elsevier Science, 1990. Pp. xvi + 429, illus. ISBN 0-444-98753-3, $77.00, Dfl. 150.00 (hardcover); 0-444-98749-5; $38.00, Dfl. 75.00 (paperback). British Journal for the History of Science, 25(2):286–287, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027324.

Hughes:1996:BRB

[Hug96] Jeff Hughes. Book review: Bruce R. Wheaton, Inventory of Sources for History of Twentieth Century Physics: Report and Microfiche Index to 700,000 Letters. With the assistance of Robin E. Rider. Stuttgart: Verlag für Geschichte der Naturwissenschaften und der Technik, 1993. Pp. x + 294, and microfiches. ISBN 3-928186-09-4. $599. British Journal for the History of Science, 29(3): 376–377, September 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027752.

Hughes:1998:BRG

[Hug98] Jeff Hughes. Book review:

G. Batchelor, The Life and Legacy of G. I. Taylor. Cambridge: Cambridge University Press, 1996. Pp. xv + 285. ISBN 0-521-46121-9. £45.00. British Journal for the History of Science, 31(3):361–375, September 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027822.

Huizenga:1992:CFS

[Hui92] John R. (John Robert)

Huizenga. Cold fusion: the scientific fiasco of the century. University of Rochester Press, Rochester, NY, USA, 1992. ISBN 1-878822-07-1. xv + 259 pp. LCCN QC791.775.C64 H85 1992. US$45.00.

Hull:1988:SPE

[Hul88] David L. Hull. Science as a process: an evolutionary account of the social and conceptual development of science. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1988. ISBN 0-226-36050-4. xiii + 586 pp. LCCN Q158.5 .H84 1988. US$39.95.

Hull:1999:WWP

[Hul99] Andrew Hull. War of words: the public science of the British scientific community and the origins of the Department of Scientific and Industrial Research, 1914–1916. British Journal for the History of Science, 32(4): 461–481, December 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027947.

Hunter:1982:RSF

[Hun82] Michael Cyril William Hunter. The Royal Society and its Fellows, 1660–1700: the morphology of an early scientific institution, volume 4 of BSHS monographs. British Society for the History of Science, Chalfont St. Giles, Bucks, England, 1982. ISBN 0-906450-03-9. 270 pp. LCCN Q41.L85 H95 1982.

Hunter:1979:HIP

[Hun85]  Louis C. Hunter.  A his-

tory of industrial power in the United States, 1780–1930. Published for the Eleutherian Mills-Hagley Foundation by the University Press of Virginia, Charlottesville, VA, USA, 1979– 1985. ISBN 0-8139-0782-9 (vol.

1), 0-262-08198-9 (vol. 3). ???? pp. LCCN HD1694.A5 H76.

Hunter:1989:ENS

[Hun89] Michael Cyril William Hunter.

Establishing the new science: the experience of the early Royal Society. Boydell Press, Woodbridge, Suffolk, UK, 1989. ISBN 0-85115-506-5. xiv + 382 pp. LCCN Q41.L85 H93 1989.

Hunter:1990:AMM

[Hun90] Michael Hunter. Alchemy, magic and moralism in the thought of Robert Boyle. British Journal for the History of Science, 23(4):387–410, December 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027198.

Hunt:1991:M

[Hun91]  Bruce J. Hunt.  The

Maxwellians. Cornell history of science series. Cornell University Press, Ithaca, NY, USA, 1991. ISBN 0-8014-2641-3. x + 266 pp. LCCN QC670 .H84 1991.

Hunter:1992:BRA

[Hun92] Michael Hunter. Book review: Alice Stroup. A Company of Scientists: Botany, Patronage, and Community at the Seventeenth-Century Parisian Royal Academy of Sciences. Berkeley: University of California Press, 1990. Pp. xv + 387, illus. $49.95. Alice Stroup. Royal Funding of the Parisian Academic des Sciences during the 1690s. Transactions of the American Philosophical Society, 77. Philadelphia: American Philosophical Society, 1987. Pp. xvi + 167. ISBN 0-87169-774-2. $15.00. British Journal for the History of Science, 25(3):362– 364, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027264.

Hunter:1995:RU

[Hun95a] Michael Hunter. Review: [untitled]. British Journal for the History of Science, 28(4): 469–470, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027555.

Hunter:1995:SSO

[Hun95b] Michael Cyril William Hunter. Science and the shape of orthodoxy: intellectual change in late Seventeenth-Century Britain. Boydell Press, Woodbridge, Suffolk, UK, 1995. ISBN 0-85115-594-4. xii + 345 pp. LCCN Q127.G4 H853 1995.

Hunt:1996:SEW

[Hun96] Bruce J. Hunt. Scientists, engineers and Wildman Whitehouse: measurement and credibility in early cable telegraphy. British Journal for the History of Science, 29(2):155– 169, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027831.

Hunter:1998:BRM

[Hun98] Michael Hunter. Book review: Michael Heyd, ‘Be Sober and Reasonable’: The Critique of Enthusiasm in the Seventeenth and Early Eighteenth Centuries. Leiden: E. J. Brill, 1996. Pp. xi + 312. ISBN 90-04-10118-7. $92.00. Dfl. 142.00. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027789.

Hunter:1999:I

[Hun99a] Michael Hunter. Introduction. British Journal for the History of Science, 32(3):257– 260, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027984.

Hunter:1999:RBS

[Hun99b] Michael Hunter. Robert Boyle (1627–1691): a suitable case for treatment? British Journal for the History of Science, 32(3): 261–275, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027985.

Hutton:1992:BRR

[Hut92] Sarah Hutton. Book review: A. Rupert Hall. Henry More: Magic, Religion and Experiment. Oxford: Blackwell, 1990. Pp. xii + 304. ISBN 0-631-17295-5. £30.00. British Journal for the History of Science, 25(2):267–269, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027310.

Hanle:1982:BAA

[HV82] Paul A. Hanle and Theodore Von Kármán. Bringing aerodynamics to America. MIT Press, Cambridge, MA, USA, 1982. ISBN 0-262-08114-8. xiv + 184 pp. LCCN TL565 .H32 1982.

Heilbron:1981:LHP

[HW81] J. L. Heilbron and Bruce R. Wheaton. Literature on the history of physics in the 20th Century, volume 5 of Berkeley papers in history of science. Office for History of Science and Technology, University of California at Berkeley, Berkeley, CA, USA, 1981. ISBN 0-918102-05-7 (paperback). ISSN 0145-0379. ix + 485 pp. LCCN Z7144.H55 H44; QC7 .H4 1981.

Harley:1987:CPA

[HW87] J. B. (John Brian) Harley and David Woodward, editors. Cartography in prehistoric, ancient, and medieval Europe and the Mediterranean, volume 1 of The history of cartography. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-31633-5. xxi + 599 pp. LCCN GA201 .H53 1987 vol.

1. URL http://www.loc.gov/
catdir/description/
uchi052/98137538.html;
http://www.loc.gov/
catdir/toc/uchi052/
98137538.html.

Harley:1987:HC

[HW03] J. B. (John Brian) Harley and David Woodward, editors. The history of cartography. University of Chicago Press, Chicago, IL, USA, 1987–2003. ISBN 0-226-31633-5 (vol. 1), 0-226-31635-1 (vol. 2, bk. 1), 0-226-31637-8 (vol. 2, bk. 2). ???? pp. LCCN GA201 .H53 1987.

Ingrao:1990:IHE

[II90] Bruna Ingrao and Giorgio Israel. The invisible hand: economic equilibrium in the history of science. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-09028-7. xiii + 491 pp. LCCN HB145 .I5413 1990.

Ilie:1995:AM

[Ili95a] Paul Ilie. The age of Minerva. University of Pennsylvania Press, Philadelphia, PA, USA, 1995. ISBN 0-8122-3307-7 (vol. 1), 0-8122-3308-5 (vol. 2). ???? pp. LCCN B802 .I36

Iliffe:1995:BRM

[Ili95b] Rob Iliffe. Book review: M. Biagioli, Galileo, Courtier: The Practice of Science in the Culture of Absolutism. Chicago: University of Chicago Press, 1993. Pp. xii + 402. ISBN 0-226-04559-5. £23.95, $29.95. British Journal for the History of Science, 28(1):106– 108, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027491.

Iliffe:1995:MDH

[Ili95c] Rob Iliffe. Material doubts: Hooke, artisan culture and the exchange of information in 1670s London. British Journal for the History of Science, 28 (3):285–318, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027646.

Iliffe:1997:BRC

[Ili97] Rob Iliffe. Book review: Cesare S. Maffioli, Out of Galileo: The Science of Waters, 1628– 1718. Foreword by Richard S. Westfall. Rotterdam: Erasmus Publishing, 1994. Pp. xx + 509. ISBN 90-5235-071-X. Dfl. 125.00. British Journal for the History of Science, 30(2): 233–249, June 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027725.

Iliffe:1998:BRP

[Ili98] Robert Iliffe. Book review:

Peter Harrison, The Bible, Protestantism and the Rise of Natural Science. Cambridge: Cambridge University Press, 1998. Pp. xi + 313. ISBN 0-521-59196-1. £40.00, $64.95. British Journal for the History of Science, 31(4):469– 487, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027885.

Inglis:1984:SPH

[Ing84] Brian Inglis. Science and parascience: a history of the paranormal, 1914–1939. Hodder and Stoughton, London, UK, 1984. ISBN 0-340-26325-3. 382 pp. LCCN BF1028 .I54 1984. US$12.95.

Inkster:1990:BRC

[Ink90] Ian Inkster. Book review:

Christine Macleod. Inventing the Industrial Revolution; The English Patent System, 1660– 1800. Cambridge, New York: Cambridge University Press, 1988, Pp. x + 302. ISBN 0-521-30104-1. £25.00, $44.50. British Journal for the History of Science, 23(3):334–336, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026759.

Irwin:1996:MSP

[IW96] Alan Irwin and Brian Wynne, editors. Misunderstanding science?: the public reconstruction of science and technology. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-43268-5 (hardcover). vi + 232 pp. LCCN Q175.5 .M58

1996. URL http://www.loc.
gov/catdir/description/
cam027/95032980.html;
http://www.loc.gov/
catdir/samples/cam034/
95032980.html;   http://
www.loc.gov/catdir/toc/
cam022/95032980.html.

Jacob:1983:HSR

[Jac83] James R. (James Randall) Jacob. Henry Stubbe, radical Protestantism and the early Enlightenment. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24876-0. viii + 222 pp. LCCN CT788.S92 J3

1983. URL http://www.loc.
gov/catdir/description/
cam022/82012919.html;
http://www.loc.gov/
catdir/samples/cam031/
82012919.html;   http://
www.loc.gov/catdir/toc/
cam021/82012919.html.

Jacob:1990:BRH

[Jac90] James R. Jacob. Book review: Hiram Caton. The Politics of Progress. The Origins and Development of the Commerical Republic, 1600–1835. Gainesville, Florida, University of Florida Press: 1988. Pp. xi 627. ISBN 0-8130-0847-6. $49. British Journal for the History of Science, 23(2): 245–246, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026745.

Jacobs:1993:BRD

[Jac93]  J. A. Jacobs.  Book re-

view: Desmond King-Hele, A Tapestry of Orbits. Cambridge: Cambridge University Press, 1992. Pp. x + 244. ISBN 0-521-39323-X. £35.00. British Journal for the History of Science, 26(3):377–378, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027426.

Jackson:1995:IDV

[Jac95a]  Mark Jackson.  Images of

deviance: visual representations of mental defectives in early twentieth-century medical texts. British Journal for the History of Science, 28(3): 319–337, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027647.

Jackson:1995:RU

[Jac95b] Mark Jackson. Review: [untitled]. British Journal for the History of Science, 28(4): 489–490, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027570.

Jackson:1996:BRP

[Jac96]  Mark Jackson.  Book re-

view: Pauline M. H. Mazumdar, Species and Specificity: An Interpretation of the History of Immunology. Cambridge: Cambridge University Press, 1995. Pp. xiii + 457. ISBN 0-521-43172-7. £40.00, $64.95. British Journal for the History of Science, 29(2): 249–250, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027850.

Jaki:1984:UGL

[Jak84] Stanley L. Jaki. Uneasy genius: the life and work of Pierre Duhem, volume 100 of Archives internationales d’histoire des idées = International archives of the history of ideas. Martinus Nijhoff, Dordrecht and The Hague, The Netherlands, 1984. ISBN 90-247-2897-5. xii + 472 pp. LCCN QC16.D925 J34

                                        1984. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1005/83019458-b.html.

James:1990:BRA

[Jam90] Frank A. J. L. James. Book review: Alan Twyman. In Search of the Mysterious Dr Weeks: A Fragment of Sandwich History. Sandwich: The Author (10 Strand Street, Sandwich, Kent, CT13 9DX, England), 1989. Pp. xi + 164. ISBN 0-9513889-0-8. £4.95. British Journal for the History of Science, 23(1): 116–117, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026821.

James:1991:BRD

[Jam91] Frank A. J. L. James. Book review: David Gooding. Experiment and the Making of Meaning: Human Agency in Scientific Observation and Experiment. Dordrecht: Kluwer, 1990. Pp. xviii + 310. ISBN 0-7923-0719-4. £60.00. British Journal for the History of Science, 24(3):386–388, September 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027246.

James:1992:BRB

[Jam92a] Frank A. J. L. James. Book review: Brian Bowers and Lenore Symons (eds.), Curiosity Perfectly Satisfyed: Faraday’s Travels in Europe, 1813– 1815. London: Peter Peregrinus Ltd in association with The Science Museum, 1991. Pp. xvii + 186. ISBN 0-86341-234-3. £19.00. British Journal for the History of Science, 25(4): 473–474, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027063.

James:1992:BRG

[Jam92b] Frank A. J. L. James. Book review: Geoffrey Cantor. Michael Faraday: Sandemanian and Scientist. A Study of Science and Religion in the Nineteenth Century. London: Macmillan, 1991. Pp. xi + 359. ISBN 0-333-55077-3. £40.00. British Journal for the History of Science, 25(4):471–472, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027061.

James:1992:BRS

[Jam92c] Frank A. J. L. James. Book review: Sydney Ross. Nineteenth-Century Attitudes: Men of Science. Dordrecht: Kluwer Academic Publishers, 1991. Pp. xi + 235. ISBN 0-7923-1308-9. £42.00, $69.00, Dfl. 125.00. British Journal for the History of Science, 25 (4):485, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027072.

James:1992:TBA

[Jam92d] Frank A. J. L. James. The tales of Benjamin Abbott: a source for the early life of Michael Faraday. British Journal for the History of Science, 25(2):229–240, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027300.

James:1993:BRL

[Jam93a] Frank A. J. L. James. Book review: Larry J. Schaaf. Out of the Shadows: Herschel Talbot and the Invention of Photography. New Haven and London: Yale University Press, 1992 Pp. xii + 188. ISBN 0-300-05705-9. £45.00 $65.00. British Journal for the History of Science, 26 (2):246–247, June 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027110.

James:1993:BRR

[Jam93b] Frank A. J. L. James. Book review: Ryan D. Tweney and David Gooding (eds.), Michael Faraday’s ‘Chemical Notes, Hints, Suggestions and Objects of Pursuit’ of 1822, London: Peter Peregrinus in association with the Science Museum, 1991. Pp. xvii + 152. ISBN 0-86341-255-6. £29.00. British Journal for the History of Science, 26(1):97, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027362.

James:1994:BRB

[Jam94a] Frank A. J. L. James. Book review: Bryan Appleyard, Understanding the Present: Science and the Soul of Modern Man. Paperback edition, London: Pan Books, 1993 (hardback published 1992). Pp. xix + 283. ISBN 0-330-32013-0. £6.99. British Journal for the History of Science, 27(4):488– 489, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027640.

James:1994:BRD

[Jam94b] Frank A. J. L. James. Book review: David Knight, Humphry Davy: Science and Power. Oxford: Blackwell, 1992. Pp. xiii + 218. ISBN 0-631-16816-8. £30.00. British Journal for the History of Science, 27(2): 230–231, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027445.

James:1994:BRE

[Jam94c] Frank A. J. L. James. Book review: Erasmus Bartholin, Experiments on Birefringent Icelandic Crystal, translated by Thomas Archibald. Introduction by Jed Z. Buchwald and Kurt Moller Pedersen, with a facsimile of the original publication. Copenhagen: Danish National Library of Science and Medicine, 1991. Pp. 63 + 64. ISBN 87-7709-010-1. 160 DKK. British Journal for the History of Science, 27(3): 366–367, September 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027607.

James:1994:BRN

[Jam94d] Frank A. J. L. James. Book review: Nahum Kipnis, History of the Principle of Interference of Light. Basel, Boston and Berlin: Birkhauser Verlag, 1991. Pp. 271. ISBN 3-7643-2316-7. SFr. 118.00. British Journal for the History of Science, 27(3):375–376, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027614.

Jankovic:1998:ICV

[Jan98] Vladimir Jankovic. Ideological crests versus empirical troughs: John Herschel’s and William Radcliffe Birt’s research on atmospheric waves, 1843–1850. British Journal for the History of Science, 31 (1):21–40, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027775.

Jensen:1991:THH

[Jen91] J. Vernon (John Vernon)

Jensen. Thomas Henry Huxley: communicating for science. University of Delaware Press, Newark, DE, USA, 1991. ISBN 0-87413-379-3. 253 pp. LCCN QH31.H9 J46 1991.

Jardine:1984:BHP

[JK84] Nicholas Jardine and Johannes Kepler. The birth of history and philosophy of science: Kepler’s A defence of Tycho against Ursus (Apologia pro Tychone contra Ursum), with essays on its provenance and significance. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25226-1. ix + 301 pp. LCCN Q125.2 .J37

1984. URL http://www.loc.
gov/catdir/enhancements/
fy0907/83014186-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0907/83014186-t.html.

Johnson:1984:HLT

[Joh84] F. Neil (Frederick Neil) Johnson. The history of lithium therapy. MacMillan Publishing Company, New York, NY, USA, 1984. ISBN 0-333-36942-4. xvii + 198 + 15 pp. LCCN RC483.5.L5 J635 1984.

Johns:1993:BRR

[Joh93] Adrian Johns. Book review: Richard W. F. Kroll, The Material Word: Literate Culture in the Restoration and Early Eighteenth Century. Baltimore and London: Johns Hopkins University Press, 1991. Pp. xxii + 420. ISBN 0-8018-4002-3. £30.50. British Journal for the History of Science, 26(2): 243–245, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027108.

Johnston:1994:BRF

[Joh94] Stephen Johnston. Book review: Frances Willmoth, Sir Jonas Moore: Practical Mathematics and Restoration Science. Woodbridge, Suffolk: The Boydell Press, 1993. Pp. xi + 244. ISBN 0-85115-321-6. £35.00, $70.00. British Journal for the History of Science, 27 (3):367–368, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027608.

Johns:1996:BRB

[Joh96] Adrian Johns. Book review:

Barbara Maria Stafford, Artful Science: Enlightenment Entertainment and the Eclipse of Visual Education. Cambridge, MA: MIT Press, 1994. Pp. xxix + 350. ISBN 0-262-19342-6. £24.95. British Journal for the History of Science, 29(3): 368–369, September 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027746.

Jones:1986:BCA

[Jon86] Alexander Jones. Book 7 of the Collection of Alexandria Pappus, volume 8 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-96257-3. x + 748 pp. LCCN QA22 .P3713 1986. US$70.00 (set : est.). Two volumes.

Jones:1988:WTT

[Jon88] Whitney R. D. (Whitney Richard David) Jones. William Turner: Tudor naturalist, physician, and divine. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1988. ISBN 0-415-00359-8. 223 pp. LCCN QH31.T88 J66 1988.

Jones:1991:BRW

[Jon91] Peter Murray Jones. Book review: William Lefanu. Nehemiah Grew: A Study and Bibliography of his Writings. Winchester: St Paul’s Bibliographies, 1990. Pp. xvii + 182. ISBN 0-906795-43-5. £42.00. British Journal for the History of Science, 24(2): 255, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027178.

Jones:1992:BRP

[Jon92] Greta Jones. Book review: Pauline M. H. Mazumdar. Eugenics, Human Failings: The Eugenics Society, its Sources and its Critics in Britain. London and New York: Routledge, 1991. Pp. x + 373. ISBN 0-415-04424-3. £40.00. British Journal for the History of Science, 25(4):486–487, December 1992.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027074.

Jones:1993:BRC

[Jon93] Greta Jones. Book review:

Carl N. Degler, In Search of Human Nature: The Decline and Revival of Darwinism in American Social Thought. Oxford: Oxford University Press, 1991. Pp. xi + 400. ISBN 0-19-506380-5. $13.95. British Journal for the History of Science, 26(4):496, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027476.

Jones:1994:BRM

[Jon94]  Greta Jones.  Book

review: Mark Pittenger, American Socialists and Evolutionary Thought, 1870–1920. Madison: University of Wisconsin Press, 1993. Pp. x + 310. ISBN 0-299-13600-0, £48.00 (hardcover); 0-299-13604-3, £20.00 (paperback). British Journal for the History of Science, 27 (1):120–122, March 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027590.

Jordanova:1984:L

[Jor84] L. J. Jordanova. Lamarck. Past masters. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1984. ISBN 0-19-287588-4, 0-19-287587-6 (paperback). viii + 118 pp. LCCN QH31.L2 J67 1984. US$15.90, US$3.90.

Jordan:1990:MCM

[Jor90] D. W. Jordan. The magnetic circuit model, 1850–1890: the resisted flow image in magnetostatics. British Journal for the History of Science, 23 (2):131–173, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026722.

Jordan:1991:BRG

[Jor91] Tim Jordan. Book reviews: Gary Gutting. Michel Foucault’s Archaeology of Scientific Reason. Cambridge: Cambridge University Press, 1989. Pp. xiii + 306. ISBN 0-521-36619-4, £27.50, $39.50 (hardcover); ISBN 0-521-36698-4, £9.95, $12.95 (paperback). British Journal for the History of Science, 24(1): 126–127, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027041.

Jordanova:1993:GHS

[Jor93] Ludmilla Jordanova. Gender and the historiography of science. British Journal for the History of Science, 26(4):469– 483, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027467.

Jordan:1994:BRP

[Jor94] Tim Jordan. Book review:

Paul Hoyningen-Huene, Reconstructing Scientific Revolutions: Thomas S. Kuhn’s Philosophy of Science, translated by Alexander T. Levin with a foreword by Thomas S. Kuhn; Chicago: University of Chicago Press, 1993. Pp. 330. ISBN 0-226-35551-0, £30.50 (hardcover), 0-226-35551-9, £12.75 (paperback). British Journal for the History of Science, 27 (3):379–380, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027618.

Jordan:1996:BRC

[Jor96] Tim Jordan. Book review:

C. W. Kilmister, Eddington’s Search for a Fundamental Theory: A Key to the Universe. Cambridge: Cambridge University Press, 1994. Pp. xii + 256. ISBN 0-521-37165-1. £35.00, $59.95. British Journal for the History of Science, 29 (3):377–378, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027753.

Josephson:1997:NAR

[Jos97]  Paul R. Josephson.  New

Atlantis revisited: Akademgorodok, the Siberian city of science. Princeton University Press, Princeton, NJ, USA, 1997. ISBN 0-691-04454-6. xxii + 351 pp. LCCN Q127.R9 J67

1997. URL http://www.loc.
gov/catdir/description/
prin021/96045577.html;
http://www.loc.gov/
catdir/toc/prin031/
96045577.html.

Joy:1987:GAA

[Joy87] Lynn Sumida Joy. Gassendi, the atomist: advocate of history in an age of science. Ideas in context. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-30142-4. xiii + 311 pp. LCCN B1887 .J69

1987. URL http://www.loc.
gov/catdir/description/
cam023/87010899.html;
http://www.loc.gov/
catdir/samples/cam034/
87010899.html;   http://
www.loc.gov/catdir/toc/
cam031/87010899.html.

Jardine:1996:CNH

[JSS96] Nicholas Jardine, James A. Secord, and E. C. (Emma C.) Spary, editors. Cultures of natural history. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-45394-1 (hardcover), 0-521-55894-8 (paperback). xxi + 501 pp. LCCN QH15 .C85 1996. URL http:

//www.loc.gov/catdir/
description/cam026/
94039445.html;   http://
www.loc.gov/catdir/toc/
cam021/94039445.html.

Junker:1989:DBR

[Jun89] Thomas Junker. Darwinismus und Botanik: Rezeption, Kritik und theoretische Alternativen im Deutschland des 19. Jahrhunderts. (German) [Darwinism and Botany: Reception, criticism and theoretical alternatives in 19th Century Germany], volume 54 of Quellen und Studien zur Geschichte der Pharmazie. In Kommission, Deutscher Apotheker Verlag, Stuttgart, Germany, 1989. ISBN 3-7692-1225-8. x + 367 pp. LCCN RS61 .Q4 Bd.54.

Kepler:1981:SUM

[KA81] Johannes Kepler and E. J.

Aiton. The secret of the universe = Mysterium cosmographicum, volume 9 of Janus series. Abaris Books, New York, NY, USA, 1981. ISBN 0-913870-64-1. 267 pp. LCCN QB41 .K413 1981.

Kahr:1999:RBF

[Kah99] Brett Kahr. Robert Boyle: a Freudian perspective on an eminent scientist. British Journal for the History of Science, 32 (3):277–284, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027986.

Kaiser:1994:BHA

[Kai94] David Kaiser. Bringing the human actors back on stage: the personal context of the Einstein–Bohr debate. British Journal for the History of Science, 27(2):129–152, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027432.

Kallinen:1995:CSN

[Kal95a] Maija Kallinen. Change and stability: natural philosophy at the Academy of Turku, 1640– 1713, volume 51 of Studia historica, 0081-6493. Finnish Historical Society, Helsinki, Finland, 1995. ISBN 951-710-001-6. 439 pp. LCCN QC9.F5 K35 1995.

Kalthoff:1995:CEE

[Kal95b] Mark A. Kalthoff, editor. Creation and evolution in the early American Scientific Affiliation, volume 10 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1811-1. xxxix + 468 pp. LCCN BS659 .C73 1995.

Katz:1990:BRD

[Kat90] David S. Katz. Book review: David B. Ruderman. Kabbalah, Magic and Science: The Cultural Universe of a Sixteenthcentury Jewish Physician. Cambridge, Mass, and London: Harvard University Press, 1988. Pp. xii + 232. ISBN 0-674-49660-4. No price given. British Journal for the History of Science, 23(3):347– 348, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026770.

Kapitsa:1990:KCM

[KBRS90] P. L. (Petr Leonidovich)

Kapitsa, J. W. Boag, P. E. Rubinin, and David Shoenberg. Kapitza in Cambridge and Moscow: life and letters of a Russian physicist. North-Holland Publishing Co., Amsterdam, The Netherlands, 1990. ISBN 0-444-98749-5 (paperback), 0-444-98753-3 (hardbound). xvi + 429 pp. LCCN QC16.K25 A4 1990.

Kamminga:1995:SCN

[KC95] Harmke Kamminga and Andrew Cunningham, Dr, editors. The science and culture of nutrition, 1840–1940, volume 32 of Clio medica; The Wellcome Institute series in the history of medicine. Rodopi, Amsterdam and Leiden, The Netherlands, 1995. ISBN 90-5183-818-2, 90-5183-819-0 (paperback). vii + 344 pp. LCCN TX353 .S385 1995.

Keas:1990:RU

[Kea90] Michael N. Keas. Review: [untitled]. British Journal for the History of Science, 23(4): 491, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027223.

Keller:1983:IAP

[Kel83]  Alex Keller.  The Infancy

of Atomic Physics: Hercules in His Cradle. Clarendon Press, New York, NY, USA, 1983. ISBN 0-19-853904-5. 230 pp. LCCN QC773 .K4 1983. US$19.50.

Kellner:1991:SAP

[Kel91] Menachem Kellner. On the status of the astronomy and physics in Maimonides’ mishneh torah and guide of the perplexed: A chapter in the history of science. British Journal for the History of Science, 24(4):453–463, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027130.

Keller:1995:BRR

[Kel95] A. G. Keller. Book review: Roland Bechmann, Villard de Honnecourt: La pensée technique au XIIIe siècle et sa communication. Paris: Picard, 1991. Pp. 383, ISBN 2-7084-0367-2. 420 FF. British Journal for the History of Science, 28(4):461–463, December 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027549.

Keller:1998:SVV

[Kel98] Susanne B. Keller. Sections and views: visual representation in eighteenth-century earthquake studies. British Journal for the History of Science, 31(2): 129–159, June 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027760.

Kennedy:1983:LSP

[Ken83] Don H. Kennedy. Little sparrow: a portrait of Sophia Kovalevsky. Ohio University Press, Athens, OH, USA, 1983. ISBN 0-8214-0692-2, 0-8214-0703-1 (paperback). ix + 341 pp. LCCN QA29.K67 K46 1983.

Kenney:1986:BUI

[Ken86] Martin Kenney. Biotechnology: the university-industrial complex. Yale University Press, New Haven, CT, USA, 1986. ISBN 0-300-03392-3. xvi + 306 pp. LCCN TP248.2 .K46 1986.

King-Hele:1992:TO

[KH92] Desmond King-Hele. A tapestry of orbits. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-39323-X. x + 244 pp. LCCN TL1080 .K55

1992. URL http://www.loc.
gov/catdir/description/
cam025/92011674.html;
http://www.loc.gov/
catdir/enhancements/
fy0632/92011674-t.html.

Kolmogorov:1992:MXV

[KI92] A. N. (Andrei Nikolaevich) Kolmogorov and A. P. (Adolf Pavlovich) Iíushkevich, editors. Matematika XIX veka.

Mathematics of the 19th century: mathematical logic, algebra, number theory, probability theory. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 3-7643-2552-6 (Basel), 0-8176-2552-6 (Boston). 308 pp. LCCN QA26 .M27813 1992.

Kilmister:1994:ESF

[Kil94] C. W. (Clive William) Kilmister. Eddington’s search for a fundamental theory: a key to the universe. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-37165-1 (hardcover). xii + 256 pp. LCCN QC6 .K439

1994. URL http://www.loc.
gov/catdir/description/
cam026/94009035.html;
http://www.loc.gov/
catdir/toc/cam021/
94009035.html.

Kim:1995:JJT

[Kim95] Dong-Won Kim. J. J. Thomson and the emergence of the Cavendish School, 1885– 1990. British Journal for the History of Science, 28(2):191– 226, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027677.

Kipnis:1991:HPI

[Kip91]  Naum S. Kipnis.  History

of the principle of interference of light, volume 5 of Science networks historical studies. Birkhäuser Verlag, Basel, Switzerland, 1991. ISBN 3-7643-2316-7 (Basel), 0-8176-2316-7 (US). 271 pp. LCCN QC411 .K53 1990.

Kitson:1989:HAC

[Kit89] Annabella Kitson, editor. History and astrology: Clio and Urania confer. Unwin Paperbacks, London, UK, 1989. ISBN 0-04-440522-7. ix + 272 pp. LCCN BF1671 .H57 1989. US$8.99.

Kitcher:1992:FDC

[Kit92]  Patricia Kitcher.  Freud’s

dream: a complete interdisciplinary science of mind. MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-11172-1. 245 pp. LCCN BF109.F74 K58 1992.

Kors:1987:AEE

[KK87] Alan Charles Kors and Paul J.

Korshin, editors. Anticipations of the Enlightenment in England, France, and Germany. University of Pennsylvania Press, Philadelphia, PA, USA, 1987. ISBN 0-8122-8057-1. viii + 290 pp. LCCN B1302.E65 A57 1987.

Klaes:1995:BRP

[Kla95] Matthias Klaes. Book review:

Philip Mirowski (ed.), Natural Images in Economic Thought: ‘Markets read in Tooth and Claw’. Historical Perspectives on Modern Economics. Cambridge: Cambridge University Press, 1994. Pp. xiii + 618. ISBN 0-521-47884-7. £22.95.

British Journal for the History of Science, 28(3):376–377, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027669.

Klein:1994:VAG

[Kle94]  Ursula Klein.  Verbindung

und Affinität: die Grundlegung der neuzeitlichen Chemie an der Wende vom 17 zum 18 Jahrhundert. (German) [Connection and affinity: the foundation of modern chemistry at the turn of the 17th to 18th Centuries], volume 14 of Science networks, historical studies. Birkhäuser Verlag, Basel, Switzerland, 1994. ISBN 3-7643-5003-2. viii + 270 pp. LCCN QD11 .K526 1994.

Kuhn:1988:SCC

[KN88] Dieter Kuhn and Joseph Needham, editors. Science and civilisation in China. Vol. 5. Chemistry and chemical technology. Pt. 9, Textile technology: Spinning and reeling. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-32021-6. xxxiv + 520 pp. LCCN ????

Knight:1986:ASS

[Kni86] David Knight. The age of science: the scientific worldview in the Nineteenth Century. Blackwell, Oxford, UK, 1986. ISBN 0-631-16176-7. 251 pp. LCCN ????

Knight:1988:ASS

[Kni88] David M. Knight, editor. The age of science: the scientific world-view in the Nineteenth Century. Blackwell, Oxford, UK, 1988. ISBN 0-631-16176-7 (paperback). x + 251 pp. LCCN ????

Knight:1989:ASS

[Kni89] David Knight. The age of science: the scientific world-view in the 19th Century. Blackwell, Oxford, UK, 1989. ISBN 0-631-16176-7, 0-631-15064-1. x + 251 pp. LCCN ????

Knight:1990:BRL

[Kni90a] David Knight. Book review:

Lynn L. Merrill. The Romance of Victorian Natural History. New York and Oxford: Oxford University Press, 1989. Pp. xi + 288. ISBN 0-19-505203X. £24.00. British Journal for the History of Science, 23(3): 359–360, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026781.

Knight:1990:BRP

[Kni90b] David Knight. Book review:

Penelope Murray (ed). Genius: The History of an Idea. Oxford: Basil Blackwell, 1989. Pp. ix + 235. ISBN 0-631-15785-9. £25.00. British Journal for the History of Science, 23(3): 348–349, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026771.

Knight:1990:BRT

[Kni90c] David Knight. Book review: Tore Frängsmyr (ed.). Science in Sweden: the Royal Swedish Academy of Sciences 1739– 1989. Canton, Mass: Science History Publications, 1989. Pp. viii + 291. ISBN 0-88135-092-3. $45.00. British Journal for the History of Science, 23(3): 380, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026795.

Knight:1990:RUa

[Kni90d] David Knight. Review: [untitled]. British Journal for the History of Science, 23(4): 464–465, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027203.

Knight:1990:RUb

[Kni90e] David Knight. Review: [untitled]. British Journal for the History of Science, 23(4): 477–478, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027212.

Knight:1992:BRE

[Kni92a] David Knight. Book review: Elizabeth Garber (ed.). Beyond History of Science: Essays in Honor of Robert E. Schofield. Bethlehem: Lehig University Press, and London and Toronto: Associated University Presses, 1990. Pp. 325. ISBN 0-934223-11-4. British Journal for the History of Science, 25(3):391, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027292.

Knight:1992:BRM

[Kni92b] David Knight. Book review:

Metascience: An International Review Journal for the History, Philosophy and Social Studies of Science, New Series, vol. 1, edited by Michael Shortland. Sydney: Australasian Association for the History, Philosophy and Social Studies of Science, 1991. Pp. viii + 169. Institutional subscription AS50.00, individual subscription A$25.00. Public Understanding of Science: An International Journal of Research in the Public Dimensions of Science and Technology, vol. 1, No. 1. Institute of Physics, in association with the Science Museum, 1992. Pp. vi + 137. ISBN 0963-6625. £23.80 (single issue), £95.00 (annual subscription). British Journal for the History of Science, 25(3): 392, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027293.

Knight:1992:HDS

[Kni92c] David M. Knight. Humphry Davy: science and power. Blackwell science biographies. Blackwell, Oxford, UK, 1992. ISBN 0-631-16816-8. xiii + 218 pp. LCCN QD22.D3 K65 1992.

Knight:1992:ICH

[Kni92d] David M. Knight, editor. Ideas in chemistry: a history of the science. Athlone Press, London, UK, 1992. ISBN 0-485-11390-2. 213 pp. LCCN QD11 .K65 1992. US$35.00.

Knight:1995:BRM

[Kni95] David Knight. Book review: Mary Jo Nye. From Chemical Philosophy to Theoretical Chemistry: Dynamics of Matter and Dynamics of Disciplines, 1800–1950. Berkeley, Los Angeles, London: University of California Press, 1994. Pp. xvii + 328. ISBN 0-520-08210-9. $48.00. British Journal for the History of Science, 28(2):242–243, June 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027688.

Knight:1996:PAG

[Kni96] David Knight. Presidential address: Getting science across. British Journal for the History of Science, 29(2):129– 138, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027829.

Knight:1998:BRJ

[Kni98a] David Knight. Book review:

John Hedley Brooke, Thinking About Matter: Studies in the History of Chemical Philosophy. Aldershot: Variorum, 1995. Pp. xii + 290. ISBN 0-86078-464-9. £47.50. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027790.

Knight:1998:BRM

[Kni98b] David Knight. Book review:

Margaret Lucas Cavendish, Grounds of Natural Philosophy, with an introduction by Colette V. Michael. West Cornwall, CT: Locust Hill Press, 1995. Pp. xx + 311. ISBN 0-933951-66-3. $38.00. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027791.

Knorr:1990:JTA

[Kno90] Wilbur R. Knorr. John of Tynemouth alias John of London: Emerging portrait of a singular medieval mathematician. British Journal for the History of Science, 23(3):293– 330, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026757.

Koetsier:1991:BRE

[Koe91] T. Koetsier. Book review: Erhard Scholz. Symmetrie, Gruppe, Dualität: Zur Beziehung zwischen theoretischer Mathematik und Anwendungen in Kristallographie und Baustatik des 19 Jahrhunderts. Basel, Boston, Berlin: Birkhäuser, 1989. Pp.

406. ISBN 3-7643-1974-7. SFr. 84.00. British Journal for the History of Science, 24(2): 265–266, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027187.

Kollerstrom:1991:NTM

[Kol91] N. Kollerstrom. Newton’s two “moon-tests”. British Journal for the History of Science, 24 (3):369–372, September 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027236.

Kollerstrom:1992:TSN

[Kol92] Nick Kollerstrom. Thomas Simpson and “Newton’s method of approximation”: an enduring myth. British Journal for the History of Science, 25(3):347– 354, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027257.

Kragh:1987:IHS

[Kra87] Helge Kragh. An introduction to the historiography of science. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33360-1. viii + 235 pp. LCCN Q125 .K686

1987. URL http://www.loc.
gov/catdir/enhancements/
fy0642/86028334-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0642/86028334-t.html.

Krach:1990:BRW

[Kra90a]  Helge  Krach.

Book review: Walter Moore. Schrödinger. Life and Thought. Cambridge: Cambridge University Press, 1989. Pp. xi + 513. ISBN 0-521-35434-X. £25.00, $39.50. British Journal for the History of Science, 23(2):231–233, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026738.

Kragh:1990:BRR

[Kra90b] Helge Kragh. Book review:

Robert Marc Friedman. Appropriating the Weather. Vilhelm Bjerknes and the Construction of a Modern Meteorology. Ithaca and London: Cornell University Press, 1989. Pp. xx + 251. ISBN 0-8014-2062-8. $38.45. British Journal for the History of Science, 23 (2):248–249, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026748.

Kragh:1990:DSB

[Kra90c] Helge Kragh. Dirac: a scientific biography. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-38089-8. x + 389 pp. LCCN QC16.D57 K73

                                        1990. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/89017257.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        89017257.html.

Kragh:1991:BRB

[Kra91a] Helge Kragh. Book review: Bruno Rossi. Moments in the Life of a Scientist. Cambridge: Cambridge University Press, 1990. Pp. xv + 181. ISBN 0-521-36439-6. £27.50, $44.50. British Journal for the History of Science, 24(4):478– 479, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027139.

Kragh:1991:BRL

[Kra91b] Helge Kragh. Book review: Laurie M. Brown, Max Dresden and Lillian Hoddeson (eds.). Pions to Quarks: Particle Physics in the 1950s. Cambridge: Cambridge University Press, 1989. Pp. xxxii + 734.

ISBN 0-521-30984-0. £40.00, $59.50. British Journal for the History of Science, 24(4): 477–478, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027138.

Krayer:1991:MIS

[Kra91c] Albert Krayer. Mathematik im Studienplan der Jesuiten: die Vorlesung von Otto Cattenius an der Universität Mainz (1610/11). (German) [Mathematics in Jesuit Studies: the lectures of Otto Cattenius at the University of Mainz (1610/ 11)], volume 15 of Beiträge zur Geschichte der Universität Mainz. Steiner, Stuttgart, Germany, 1991. ISBN 3-515-05874-5. ix + 434 pp. LCCN

????  URL 01; 04; 2;
http://www.gbv.de/dms/
hbz/toc/ht004191413.;
http://www.zentralblatt-
math.org/zmath/en/search/
?an=0790.01006.

Kragh:1992:BRK

[Kra92] Helge Kragh. Book review:

Kameshwar C. Wall. Chandra: A Biography of S. Chandrasekhar. Chicago: University of Chicago Press, 1991. Pp. x + 341, illust. ISBN 0-226-87054-5. $34.50. British Journal for the History of Science, 25(2):287–288, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027325.

Kragh:1993:BRJa

[Kra93a] Helge Kragh. Book review: J. Eisenstaedt and A. J. Kox (eds.), Studies in the History of General Relativity. Boston, Basel, Berlin: Birkhäuser, 1992. Pp. xii + 468. ISBN 0-8176-3479-7. SFr. 198.00. British Journal for the History of Science, 26 (1):123–124, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027385.

Kragh:1993:BRJb

[Kra93b] Helge Kragh. Book review: Jack S. Goldstein, A Different Sort of Time: The Life of Jerrold R. Zacharias, Scientist, Engineer, Educator. Cambridge, Mass, and London: MIT Press, 1992. Pp. xviii + 373. ISBN 0-262-07138-X. $31.50. British Journal for the History of Science, 26(2): 255, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027118.

Kragh:1994:BRT

[Kra94] Helge Kragh. Book review: E. A. Tropp, V. Ya. Frenkel and A. D. Chernin, Alexander A. Friedmann: The Man who Made the Universe Expand. Cambridge: Cambridge University Press, 1993. Pp. x + 267. ISBN 0-521-38470-2. £49.95. British Journal for the History of Science, 27(3): 377–378, September 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027616.

Kragh:1997:MJH

[Kra97] Helge Kragh. S. M. Jørgensen and his controversy with A. Werner: a reconsideration. British Journal for the History of Science, 30(2):203– 219, June 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027717.

Krauss:1998:BRM

[Kra98] Rolf Krauss. Book review: Marshall Clagett, Ancient Egyptian Science, Volume 2: Calendars, Clocks, and Astronomy. Philadelphia: American Philosophical Society, 1995. Pp. xiv + 575, illus. ISBN 0-87169-214-7. No price given. British Journal for the History of Science, 31(1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027792.

Kretzmann:1982:ICA

[Kre82] Norman Kretzmann, editor. Infinity and continuity in ancient and medieval thought. Cornell University Press, Ithaca, NY, USA, 1982. ISBN 0-8014-1444-X. 367 pp. LCCN BD411 .I53.

Kremer:1990:TLE

[Kre90] Richard Lynn Kremer. The thermodynamics of life and experimental physiology, 1770– 1880. Harvard dissertations in the history of science. Garland Publishing, Inc., New York, NY, USA, 1990. ISBN 0-8240-0039-0. xii + 512 pp. LCCN QP517.T48 K74 1990.

Kremer:1996:BRR

[Kre96] Richard L. Kremer. Book review: R. Steven Turner, In the Eye’s Mind: Vision and the Helmholtz-Hering Controversy. Princeton: Princeton University Press, 1994. Pp. xiv + 338. ISBN 0-691-03397-8. £37.50, $49.50. British Journal for the History of Science, 29(4): 482–483, December 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027704.

Krimsky:1982:GAS

[Kri82]  Sheldon Krimsky.  Genetic

alchemy: the social history of the recombinant DNA controversy. MIT Press, Cambridge, MA, USA, 1982. ISBN 0-262-11083-0. xiii + 445 pp. LCCN QH442 .K74 1982.

Krisciunas:1988:ACW

[Kri88] Kevin Krisciunas. Astronomical centers of the world. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-30278-1. x + 320 pp. LCCN QB81 .K75 1988. URL http:

//www.loc.gov/catdir/
description/cam023/
86021542.html;   http://
www.loc.gov/catdir/toc/
cam028/86021542.html.

Krige:1990:RU

[Kri90] John Krige. Review: [untitled]. British Journal for the History of Science, 23(4):472– 473, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027208.

Krige:1991:BRF

[Kri91] John Krige. Book review: Finn Aaserud. Redirecting Science: Niels Bohr, Philanthropy and the Rise of Nuclear Physics. Cambridge: Cambridge University Press, 1990. Pp. xiii + 356. ISBN 0-521-35366-1. £35.00, $47.50. British Journal for the History of Science, 24(4):475–476, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027136.

Kroll:1991:MWL

[Kro91] Richard W. F. Kroll. The material word: literate culture in the Restoration and early Eighteenth Century. The Johns Hopkins University Press, Baltimore, MD, USA, 1991. ISBN 0-8018-4002-3. xxii + 420 pp. LCCN PR437 .K7

1990. URL http://www.loc.

                                        gov/catdir/description/
                                        jhu051/90030586.html.

Kirk:1983:PPC

[KRS83] G. S. (Geoffrey Stephen) Kirk, J. E. (John Earle) Raven, and Malcolm Schofield. The presocratic philosophers: a critical history with a selection of texts. Cambridge University Press, Cambridge, UK, second edition, 1983. ISBN 0-521-25444-2, 0-521-27455-9 (paperback). xiii + 501 pp. LCCN B188 .K5 1983. URL http:/

                                        /www.loc.gov/catdir/
                                        description/cam022/
                                        82023505.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam028/82023505.html.

Krige:1994:ES

[KRS94] John Krige, Arturo Russo, and Lorenza Sebesta. Europe in space, 1960–1973, volume 1172 of ESA SP. European Space Agency, Noordwijk, The Netherlands, 1994. ISBN 92-9092-125-0. viii + 142 pp. LCCN TL789.8.E9 K7497 1994. 70 Dfl.

Kargon:1987:KBL

[KT87] Robert Kargon and Peter Achinstein William Thomson, Baron Kelvin, editors. Kelvin’s Baltimore lectures and modern theoretical physics: historical and philosophical perspectives. Studies from the Johns Hopkins Center for the History and Philosophy of Science. MIT Press, Cambridge, MA, USA, 1987.

ISBN 0-262-11117-9. x + 547 pp. LCCN QC19.6 .K45 1987.

Kuklick:1991:SWS

[Kuk91] Henrika Kuklick. The savage within: the social history of British anthropology, 1885– 1945. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-41109-2. 320 pp. LCCN ????

Kuklick:1993:SWS

[Kuk93] Henrika Kuklick. The savage within: the social history of British anthropology, 1885–1945. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-45739-4, 0-521-41109-2. 325 pp. LCCN ????

Kumar:1991:SEE

[Kum91] Deepak Kumar, editor. Science and Empire: essays in Indian context, 1700–1947. Anamika Prakashan, Delhi, India, 1991. ISBN 81-85150-19-2. xvi + 205 pp. LCCN Q127.I4 S322 1991. Rs200.00 (US$30.00 U.S.).

Kumar:1995:SR

[Kum95] Deepak Kumar. Science and the Raj, 1857–1905. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1995. ISBN 0-19-563562-0. xiv + 273 pp. LCCN Q127.I4 K86 1995.

Kumar:1996:CSC

[Kum96] Deepak Kumar. The ‘culture’ of science and colonial culture, India 1820–1920. British Journal for the History of Science, 29(2):195–209, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027833.

Kusch:1994:BRG

[Kus94]  Martin Kusch.  Book re-

view: Graham Richards, Mental Machinery: The Origins and Consequences of Psychological Ideas, Part 1: 1600– 1850. London: The Athlone Press, 1992. Pp. xi + 490. ISBN 0-485-11394-5. £50.00. British Journal for the History of Science, 27(2):232–233, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027446.

Kusch:1995:BM

[Kus95a] Martin Kusch. Back matter. British Journal for the History of Science, 28(1): 130, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027507.

Kusch:1995:BRB

[Kus95b]  Martin Kusch.  Book re-

view: Bruno Latour, We Have Never Been Modern, translated by Catherine Porter. Hemel Hempstead: Harvester Wheatsheaf, 1993. Pp. x + 157. ISBN 0-7450-0682-X. £13.95. British Journal for the History of Science, 28(1):125– 126, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027505.

Kusch:1995:BRK

[Kus95c] Martin Kusch. Book review:

Kurt Danziger. Constructing the Subject: Historical Origins of Psychological Research. Cambridge Studies in the History of Psychology. Cambridge: Cambridge University Press, 1994. Pp. ix + 254. ISBN 0-521-46785-3. £12.95, $16.95 (paperback edition). British Journal for the History of Science, 28 (2):243–245, June 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027689.

Kusch:1995:BRM

[Kus95d] Martin Kusch. Book review:

Michael Heidelberger, Die innere Seite der Natur: Gustav Theodor Fechners wissenschaftlichphilosophische Weltauffassung. Philosophische Abhandlungen,

60. Frankfurt-on-Main: Vittorio Klostermann, 1993. Pp.

457. ISBN 3-465-02590-3. DM 118. British Journal for the History of Science, 28(1):112– 114, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027495.

Kusch:1995:BRR

[Kus95e] Martin Kusch. Book review: Richard Noll, The Jung Cult: Origins of a Charismatic Movement. Princeton: Princeton University Press, 1994. Pp. xvi + 387. ISBN 0-691-03724-8. £19.95, $24.95. British Journal for the History of Science, 28 (3):357–358, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027655.

Kusch:1995:RU

[Kus95f] Martin Kusch. Review: [untitled]. British Journal for the History of Science, 28(4): 483–484, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027565.

Kusukawa:1995:TNP

[Kus95g] Sachiko Kusukawa. The transformation of natural philosophy: the case of Philip Melanchthon. Ideas in context. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-47347-0 (hardcover). xv + 246 pp. LCCN BR339 .K87

                                        1995. URL http://www.loc.
                                        gov/catdir/description/
                                        cam026/94018291.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam022/
                                        94018291.html.

Kusch:1996:BRA

[Kus96a] Martin Kusch. Book review:

Allan Megill (ed.), Rethinking Objectivity. Durham, NC, and London: Duke University Press, 1994. Pp. x + 342. ISBN 0-8223-1479-7, £43.50 (hardcover); 0-8223-1494-0, £14.95 (paperback). British Journal for the History of Science, 29 (3):378–380, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027754.

Kusch:1996:BRM

[Kus96b] Martin Kusch. Book review:

Mitchell G. Ash, Gestalt Psychology in German Culture, 1890–1967: Holism and the Quest for Objectivity. Cambridge Studies in the History of Psychology. Cambridge: Cambridge University Press, 1995. Pp. xii + 513. ISBN 0-521-47540-6. £35.00, $54.90. British Journal for the History of Science, 29(4):483– 486, December 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027705.

Kusch:1997:BRD

[Kus97]  Martin Kusch.  Book re-

view: David E. Leary (ed.), Metaphors in the History of Psychology. Cambridge Studies in the History of Psychology. Cambridge: Cambridge University Press, 1994. Pp. xiv + 383. ISBN 0-521-42152-7. £12.95, $17.95 (paperback edition). British Journal for the History of Science, 30(1):101– 121, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027907.

Kusch:1998:BRJ

[Kus98] Martin Kusch. Book review: José Brunner, Freud and the Politics of Psychoanalysis. Oxford: Blackwell, 1995. Pp. xiv + 238. ISBN 0-631-16404-9. £40.00, $49.95. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027793.

Kuznick:1987:BLS

[Kuz87] Peter J. Kuznick. Beyond the laboratory: scientists as political activists in 1930s America. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-46583-7 (hardcover), 0-226-46584-5 (paperback). x + 363 pp. LCCN Q175.52.U5 K89

                                        1987. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0608/87005098-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0608/87005098-t.html.

Kepler:1937:GWG

[KvDC^{+}37] Johannes Kepler, Walther von

Dyck,  Max  Caspar,

Franz Hammer, Volker Bialas, Jürgen Hübner, and Helmuth Grössing, editors. Gesammelte Werke. (German) [Collected works]. Beck’sche Verlagsbuchhandlung, München, Germany, 1937. ISBN 3-406-37510-3 (vol. 11,2), 3-406-37511-1 (vol. 11,2), 3-406-01661-8 (vol. 12), 3-406-01660-X (vol. 12), 3-406-31501-1 (vol. 20,1), 3-406-31502-X (vol. 20,1), 3-406-40592-4 (vol. 20,2), 3-406-40593-2 (vol. 20,2). ???? pp. LCCN ????

Kursunoglu:1987:RAG

[KW87] Behram Kurşunoğlu and Eugene Paul Wigner, editors. Reminiscences about a great physicist: Paul Adrien Maurice Dirac. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34013-6. xviii + 297 pp. LCCN QC16.D57 R46

1987. URL http://www.loc.
gov/catdir/description/
cam023/86033409.html;
http://www.loc.gov/
catdir/toc/cam028/
86033409.html.

LopezPinero:1986:ANE

[Lan86] José Mariá López Piñero. El arte de navegar en la Espaãna del Renacimiento. (Spanish) [Seacraft in Renaissance Spain]. Editorial Labor, Barcelona, Spain, 2a edition, 1986. ISBN 84-335-0029-5. 285 pp. LCCN VK87 .L67 1986.

LopezPinero:1987:GCH

[Lan87] José María López Piñero. El grabado en la ciencia hispánica. (Spanish) [Engraving in Spanish science]. Consejo Superior de Investigaciones Científicas, Madrid, Spain, 1987. ISBN 84-00-06699-5. 80 + 60 pp. LCCN Q127.S7 L655 1987.

LopezPinero:1989:OEE

[Lan89] José María López Piñero. Los origenes en España de los estudios sobre la salud publica. (Spanish) [The origins in Spain of studies on public health], volume 1 of Textos clásicos de la salud pública. Ministerio de Sanidad y Consumo, Madrid, Spain, 1989. ISBN 84-7670-164-0, 84-7670-165-9. 245 pp. LCCN RA287 .L66 1989.

LaVergata:1994:BRG

[La 94] Antonello La Vergata. Book review: Giulio Barsanti, La scala, la mappa, l’albero. Immagini e classificazioni della natura fra sei e ottocento. Firenze: Sansoni, 1992. Pp. xviii + 310. 60,000 Lire. British Journal for the History of Science, 27(4): 476–477, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027630.

LaVergata:1997:BRJ

[La 97]  Antonello  La

Vergata. Book review: Janet Browne, Charles Darwin. Voyaging. Volume 1 of a Biography. London: Jonathan Cape, 1995. Pp. xvii + 606, illus. ISBN 0-224-04202-5. £25.00. British Journal for the History of Science, 30(2):233–249, June 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027729.

LaFollette:1982:QS

[LaF82] Marcel C. (Marcel Chotkowski)

LaFollette, editor. Quality in science. MIT Press, Cambridge, MA, USA, 1982. ISBN 0-262-12099-2, 0-262-62040-5 (paperback). xi + 169 pp. LCCN Q172.5.S34 Q34 1982.

Landsberg:1982:ET

[Lan82] Peter Theodore Landsberg, editor. The Enigma of Time. Adam Hilger Ltd., Bristol, UK, 1982. ISBN 0-85274-545-1. xii + 248 pp. LCCN QB209 .E54 1982. US$13.95.

Langins:1987:RAB

[Lan87] Janis Langins. La République avait besoin de savants: les débuts de l’École polytechnique — l’École centrale des travaux publics et les cours révolutionnaires de l’an III. (French) [The Republic needed sciences: the start of the Polytechnical School — the central school for public works and the revolutionary courses of the year III]. Librairie du Bicentenaire de la Révolution française. Belin, Paris, France, 1987. ISBN 2-7011-0641-9. 287 pp. LCCN T173.P39 L36 1987.

Lane:1990:BRP

[Lan90]  Joan Lane.  Book review:

P. J. & R. V. Wallis (with J. G. L. Burnby & T. D. Whittet). Eighteenth Century Medics (subscriptions, licences, apprenticeships). Newcastle-upon-Tyne: PHIBB, second edition, 1988. Pp. xx + 690. ISBN 1-871768-00-4. £80.00. British Journal for the History of Science, 23(3): 363–364, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026784.

Lanaro:1991:PTS

[Lan91] Giorgio Lanaro. Il positivismo tra scienza e religione. (Italian) [The positivism between science and religion], volume 39 of Collana di filosofia. Angeli, Milano, Italy, 1991. ISBN 88-204-6373-3. 173 pp. LCCN IN PROCESS. L24000.

Larson:1990:BRR

[Lar90] Edward J. Larson. Book review: Ruth Harris. Murders and Madness: Medicine, Law, and Society in the Fin de Siècle. Oxford: Clarendon Press, 1989. Pp. x + 366. ISBN 0-19-822991-7. £30.00. British Journal for the History of Science, 23(3): 351–352, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026774.

Larson:1991:REE

[Lar91] Edward J. Larson. The rhetoric of Eugenics: expert authority and the Mental Deficiency Bill. British Journal for the History of Science, 24(1):45– 60, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027015.

Larson:1996:BRR

[Lar96a] Edward J. Larson. Book reviews: Ronald L. Numbers (General Editor), Creationism in Twentieth-Century America: A Ten-Volume Anthology of Documents, 1903–1961. New York: Garland Publishing, 1995. ISBN 0-8153-1801-4. $732.00 set, consisting of: Volume 1: Ronald L. Numbers (ed.), Antievolution Before World War I. Pp. xvii + 403. ISBN 0-8153-1802-2. $65.00. Volume 2: Ronald L. Numbers (ed.), Creation-Evolution Debates. Pp. xiv + 505, illus. ISBN 0-8153-1803-0. $65.00. Volume 3: Ronald L. Numbers (ed.), The Antievolution Works of Arthur I. Brown. Pp. xiv + 209. ISBN 0-8153-1804-9. $65.00. Volume 4: William Vance Trollinger Jr, (ed.), The Antievolution Pamphlets of William Bell Riley. Pp. xxii + 221. ISBN 0-8153-1805-7. $55.00. Volume 5: Paul Nelson (ed.), The Creationist Writings of Byron C. Nelson. Pp. xxvi + 505, illus. ISBN 0-8153-1806-5. $65.00. Volume 6: Edward B. Davis (ed.), The Antievolution Pamphlets of Harry Rimmer. Pp. xxxiv + 482, illus. ISBN 0-8153-1807-3. $84.00. Volume 7: Ronald L. Numbers (ed.), Selected Works of George McCready Price. Pp. xviii + 489. ISBN 0-8153-1808-1. $75.00. Volume 8: Ronald L. Numbers (ed.), The Early Writings of Harold W. Clark and Frank Lewis Marsh. Pp. xxiv + 531, illus. ISBN 0-8153-1809-X. $93.00. Volume 9: Ronald L. Numbers (ed.), Early Creationist Journals. Pp. xiv + 629. ISBN 0-8153-1801-3. $100.00. Volume 10: Mark A. Kalthoff (ed.), Creation and Evolution in the Early American Scientific Affiliation. Pp. xl + 468, illus. ISBN 0-8153-1811-1. $65.00. British Journal for the History of Science, 29 (2):250–252, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027851.

Larson:1996:BRM

[Lar96b] James Larson. Book review: Marco Beretta, The Starry Messenger and the Polar Star: Scientific Relations between Italy and Sweden from 1500 to 1800. Catalogue of an Exhibition held at the Naturhistoriska Riksmuseet, Stockholm. Uppsala Studies in History of Science, 22. Prato: Giunti, 1995. Pp. 188, illus. ISBN 88-09-20793-9. No price given. British Journal for the History of Science, 29(4):479– 480, December 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027701.

Latour:1988:PF

[Lat88] Bruno Latour. The pasteurization of France. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-65760-8. 273 pp. LCCN QR22.F8 L3813 1988.

Laudan:1987:MGF

[Lau87a] Rachel Laudan. From mineralogy to geology: the foundations of a science, 1650–1830. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-46950-6. xii + 278 pp. LCCN QE11 .L35

1987. URL http://www.loc.
gov/catdir/description/
uchi052/86030783.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/86030783-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/86030783-b.html.

Laurent:1987:PEF

[Lau87b] Goulven Laurent. Paléontologie et évolution en France de 1800 à 1860: une histoire des idées de Cuvier et Lamarck à Darwin. (French) [Paleontology and evolution in France from 1800 to 1860: a history of the ideas of Cuvier and Lamarck to Darwin], volume 4 of Mémoires de la Section d’histoire des sciences et des techniques, 0769-2218. Editions du Comité des travaux historiques et scientifiques, Paris, France, 1987. ISBN 2-7355-0126-4. xiv + 553 pp. LCCN QE705.F8 L38 1987. 350F.

Lawrence:1990:BRM

[Law90a] Christopher Lawrence. Book review: Michael Shortland. Medicine and Film: a checklist survey and research resource. Oxford: Research Publications on the Wellcome Unit for the History of Medicine, Oxford, No. IX, 1989. Pp. viii + 54. ISSN 0143 7984. £6.00. British Journal for the History of Science, 23(1):109, March 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026814.

Lawrence:1990:BRR

[Law90b] Christopher Lawrence. Book review: Roger Cooter. (ed.). Studies in the History of Alternative Medicine. Oxford: Macmillan Press in association with St Anthony’s College, Oxford, 1988. Pp. xx + 180. ISBN 0-333-46213-0. £29.50 (hardcover). British Journal for the History of Science, 23(1): 109–110, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026815.

Lawrence:1990:BRS

[Law90c] Christopher Lawrence. Book review: Stephen Trombley. Sir Frederick Treves: the extraordinary Edwardian. London: Routledge, 1989. Pp. x + 218. ISBN 0-415-03423-X. £19.95. British Journal for the History of Science, 23(1):111– 112, March 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026817.

Lawrence:1990:RU

[Law90d] Christopher Lawrence. Review:

[untitled]. British Journal for the History of Science, 23(4): 475, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027210.

Lawrence:1991:BRSb

[Law91a] Christopher Lawrence. Book review: Saul Jarcho (ed.). Clinical Consultations and Letters by Ippolito Francesco Albertini, Francesco Torti, and Other Physicians. Canton, Mass.: Science History Publications, 1989. Pp. lxix + 356. ISBN 0-88135-089-3. $24.95. British Journal for the History of Science, 24 (2):258–259, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027181.

Lawrence:1991:BRSa

[Law91b] Christopher Lawrence. Book review: Susan Leigh Star. Regions of the Mind: Brain Research and the Quest for Scientific Certainty. Stanford: Stanford University Press, 1989. Pp. xviii + 278. ISBN 0-8047-1673-0. British Journal for the History of Science, 24(1): 107–109, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027025.

Lawrence:1991:BRSc

[Law91c] Christopher Lawrence. Book reviews: Shirley Roberts. Sir James Paget: The Rise of Clinical Surgery. London: Royal Society of Medicine Services Limited, 1989. Pp. xii + 223. ISBN 0-905958-91-8. £12.95. Selwyn Taylor. Robert Graves: The Golden Years of Irish Medicine. London: Royal Society of Medicine Services Limited, 1989. Pp. x + 160. ISBN 0-905958-98-5. £12.95. British Journal for the History of Science, 24(2):269, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027190.

Lawrence:1992:BRA

[Law92a] Christopher Lawrence. Book review: Alfred I. Tauber and Leon Chernyak. Metchnikoff and the Origins of Immunology: From Metaphor to Theory. New York and Oxford: Oxford University Press, 1991. Pp. xx + 247. ISBN 0-19-506447-X. British Journal for the History of Science, 25(4): 493, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027081.

Lawrence:1992:BRC

[Law92b] Christopher Lawrence. Book review: Christopher Hamlin. A Science of Impurity: Water Analysis in Nineteenth Century Britain. Bristol: Adam Hilger, 1990. Pp. xiii + 342. ISBN 0-7503-0042-6. £45.00. British Journal for the History of Science, 25(2):279–280, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027318.

Lawrence:1992:BRE

[Law92c] Christopher Lawrence. Book review: Evelleen Richards. Vitamin C and Cancer: Medicine or Politics? Basingstoke: Macmillan, 1991. Pp. xiv + 269. ISBN 0-333-44419-1. £35.00. British Journal for the History of Science, 25(4):499–500, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027089.

Lawrence:1992:BRM

[Law92d] Christopher Lawrence. Book review: Marina Benjamin (ed.). Science and Sensibility: Gender and Scientific Enquiry, 1780– 1945. Oxford: Basil Blackwell, 1991. Pp. x + 295. ISBN 0-631-16649-1. £35.00. British Journal for the History of Science, 25(3):389–391, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027291.

Lawrence:1992:BRR

[Law92e] Christopher Lawrence. Book review: Richard L. Kremer. The Thermodynamics of Life and Experimental Physiology, 1770–1880. New York and London: Garland Publishing, 1990. Pp. xii + 512. ISBN 0-8240-0039-0. $50.00. British Journal for the History of Science, 25 (2):274–275, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027315.

Lawn:1993:RDS

[Law93a] Brian Lawn. The rise and decline of the scholastic Quaestio disputata: with special emphasis on its use in the teaching of medicine and science, volume 2 of Education and society in the Middle Ages and Renaissance, 0926-6070. New York, Leiden, The Netherlands, 1993. ISBN 90-04-09740-6 (hardcover). 176 pp. LCCN AZ321 .L38 1993.

Lawrence:1993:BRG

[Law93b] Christopher Lawrence. Book review: G. J. Barker-Benfield. The Culture of Sensibility: Sex and Society in Eighteenth-Century Britain. Chicago and London: University of Chicago Press, 1992. Pp. xxxiv + 520. ISBN 0-226-03713-4. £39.95, $57.50. British Journal for the History of Science, 26(3):364– 365, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027414.

Lawrence:1993:BRJa

[Law93c] Christopher Lawrence. Book review: Jan Golinski, Science as Public Culture: Chemistry and Enlightenment in Britain, 1760–1820. Cambridge: Cambridge University Press, 1992. Pp. xi + 342. ISBN 0-521-39414-7. £32.50, $54.95. British Journal for the History of Science, 26(1):93, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027357.

Lawrence:1993:BRJb

[Law93d] Christopher Lawrence. Book review: John Snow, On Narcotism by the Inhalation of Vapours, a facsimile edition with an introductory essay by Richard H. Ellis. London: Royal Society of Medicine Services Limited, 1991. Pp. xxx + 112. ISBN 1-85315-158-0. British Journal for the History of Science, 26(1):96– 97, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027361.

Lawrence:1993:BRJc

[Law93e] Christopher Lawrence. Book review: Judith A. Overmier and John Edward Senior, Books and Manuscripts of the Bakken. Metuchen, New Jersey: Scarecrow Press Inc. Folkestone: Shelwing Ltd, 1992. Pp. xvii + 512. ISBN 0-8108-2570-8. £59.65. British Journal for the History of Science, 26(1): 132, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027393.

Lawrence:1993:BRM

[Law93f] Christopher Lawrence. Book review: Martin J. S. Rudwick, Scenes from Deep Time: Early Pictorial Representations of the Prehistoric World. Chicago: University of Chicago Press, 1992. Pp. xiii + 280, illus. ISBN 0-226-73104-9. £35.95, $51.75. British Journal for the History of Science, 26(4): 495–496, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027475.

Lawrence:1993:BRB

[Law93g] Christopher Lawrence. Book reviews: Barbara Maria Stafford, Body Criticism: Imaging the Unseen in Enlightenment Art and Medicine. Cambridge, Mass, and London: MIT Press, 1991. Pp. xxi + 587. ISBN 0-262-19304-3. £49.50. K. B. Roberts and J. D. W. Tomlinson, The Fabric of the Body: European Traditions of Anatomical Illustration. Oxford: Clarendon Press, 1992. Pp. xx + 638, illus. ISBN 0-19-261198-4. £95.00. British Journal for the History of Science, 26 (1):95–96, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027359.

Lawrence:1996:BRD

[Law96a] Christopher Lawrence. Book review: Dorinda Outram, The Enlightenment: New Approaches to European History. Cambridge: Cambridge University Press, 1995. Pp. xv + 141. ISBN 0-521-41522-5, £27.95, $44.95 (hardcover); 0-521-42534-4, £9.95, $14.95 (paperback). British Journal for the History of Science, 29(3): 367–368, September 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027745.

Lawrence:1996:BRA

[Law96b] Christopher J. Lawrence. Book review: Adrian Wilson, The Making of Man-Midwifery: Childbirth in England, 1660–1770. London: UCL Press, 1995 Pp xii + 240. ISBN 1-85728-292-2. £40.00. British Journal for the History of Science, 29(1):96–97, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027519.

Lawrence:1996:BRM

[Law96c] Christopher J. Lawrence. Book review: Mark S. Micale, Approaching Hysteria: Disease and Its Interpretations. Princeton: Princeton University Press, 1995. Pp. xii + 327. ISBN 0-691-03717-5. £24.95, $29.95. British Journal for the History of Science, 29 (1):102–103, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027524.

Lawrence:1998:BRGb

[Law98a] Christopher Lawrence. Book review: Geoffrey V. Sutton, Science for a Polite Society: Gender, Culture, and the Demonstration of Enlightenment. Oxford: Westview Press, 1995. Pp. xiii + 391. ISBN 0-8133-1575-1. £22.50. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027795.

Lawrence:1998:BRGa

[Law98b] Christopher Lawrence. Book review: Gillian Beer, Open Fields, Science in Cultural Encounter. Oxford: Oxford University Press, 1996. Pp. x + 342. ISBN 0-19-818369-0. £25.00. British Journal for the History of Science, 31(1): 63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027794.

Lawrence:1998:BRM

[Law98c] Christopher Lawrence. Book review: Michael Shortland and Richard Yeo (eds.), Telling Lives in Science: Essays on Scientific Biography. Cambridge: Cambridge University Press, 1996. Pp. xiv + 295. ISBN 0-521-43323-1. £50.00. British Journal for the History of Science, 31(1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027796.

LeFanu:1990:NGS

[Le 90] W. R. (William Richard) Le Fanu. Nehemiah Grew: a study and bibliography of his writings. St Paul’s Bibliographies, Winchester, UK, 1990. ISBN 0-906795-43-5. xvii + 182 pp. LCCN Z8369.262 .L4 1990.

LeGrand:1991:BRJ

[Le 91] Homer E. Le Grand. Book review: John A. Stewart. Drifting Continents and Colliding Paradigms: Perspectives on the Geoscience Revolution. Bloomington and Indianapolis: Indiana University Press, 1990. Pp. xiv + 285. ISBN 0-253-35405-6. $35.00. British Journal for the History of Science, 24(3): 394–396, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027251.

Leary:1990:MHP

[Lea90]  David E. Leary,  ed-

itor. Metaphors in the history of psychology. Cambridge studies in the history of psychology, 99-1209266-X. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37166-X, 0-521-42152-7. xiii + 383 pp. LCCN ????

Leary:1994:MHP

[Lea94]  David  E.

Leary, editor. Metaphors in the history of psychology. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-42152-7 (paperback), 0-521-37166-X. xiii + 383 pp. LCCN BF38 .M482 1990.

Lee:1990:BRM

[Lee90]  Peter Lee.  Book review:

Michael Shortland & Andrew Warwick (eds). Teaching the History of Science. Oxford and New York: Basil Blackwell (in conjunction with the British Society for the History of Science), 1989. Pp. viii + 281. ISBN 0-631-16977-6 and 0-631-16978-4 (paperback). £29.50 (hardcover) and £9.95 (paperback). British Journal for the History of Science, 23(2): 233–237, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026739.

Leegwater:1996:TSS

[Lee96] Arie Leegwater. Telling the story of science and religion: a nuanced account. British Journal for the History of Science, 29(3):357–359, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027738.

LeGrand:1988:DCS

[LeG88] H. E. (Homer Eugene)

LeGrand. Drifting continents and shifting theories: the modern revolution in geology and scientific change. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-32210-3, 0-521-31105-5 (paperback). vi + 313 pp. LCCN QE511.5 .L44

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/88030254.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        88030254.html.

Lenoir:1982:SLT

[Len82] Timothy Lenoir. The strategy of life: teleology and mechanics in Nineteenth Century German biology, volume 13 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1982. ISBN 90-277-1363-4. xii + 314 pp. LCCN QH305.2.G3 L46 1982.

Levi:1985:GHL

[Lev85] Hilde Levi. George de Hevesy: life and work: a biography. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-555-9. 147 pp. LCCN QD22.H45 L48 1985.

Levy:1990:PPP

[Lev90] B. Barry Levy. Planets, potions, and parchments: scientifica Hebraica from the Dead Sea Scrolls to the Eighteenth Century. Published for the Jewish Public Library by McGill-Queen’s University Press, Montréal, Québec, Canada, 1990. ISBN 0-7735-0791-4 (paperback), 0-7735-0793-0 (hardcover). xii + 139 pp. LCCN Q128 .L48 1990.

Levere:1993:SCA

[Lev93] Trevor Harvey Levere. Science and the Canadian Arctic: a century of exploration, 1818–1918. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-41933-6. xiv + 438 pp. LCCN Q115 .L554

1993. URL http://www.loc.
gov/catdir/description/
cam025/92012570.html;
http://www.loc.gov/
catdir/samples/cam031/
92012570.html;   http://
www.loc.gov/catdir/toc/
cam029/92012570.html.

Levere:1994:BRG

[Lev94] Trevor H. Levere. Book review: G. E. Fogg, A History of Antarctic Science. Studies in Polar Research. Cambridge: Cambridge University Press, 1992. Pp. xxi + 483. ISBN 0-521-36113-3. £55.00. British Journal for the History of Science, 27(1):118– 120, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027589.

Lewis:1995:H

[Lew95] Mark Edward Lewis. The Huainanzi. British Journal for the History of Science, 28(3): 339–343, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027648.

Leyere:1990:BRI

[Ley90] Trevor H. Leyere. Book review: Ian Wylie. Young Coleridge and the Philosophers of Nature. Oxford: Clarendon Press, 1989. Pp. x + 176. ISBN 0-19-812983-1. £22.50. British Journal for the History of Science, 23(3): 369–370, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026788.

Lavoisier:1955:OLC

[LF97] Antoine Laurent Lavoisier and René Fric. Œuvres de Lavoisier: correspondance. (French) [Works of Lavoisier: correspondence]. A. Michel, Paris, France, 1955– 1997. ISBN 2-7011-1085-8 (vol. 4), 2-85367-184-4 (vol. 5), 2-85367-205-0 (vol. 6). ???? pp. LCCN QD22.L4 A4 1997.

Leibniz:1976:MNT

[LHB04] Gottfried Wilhelm Leibniz, Heinz-Jürgen Hess, and Herbert Breger. Mathematischer naturwissenschaftlicher und technischer Briefwechsel. (German) [Mathematical, scientific and technical letters], volume III of Sämtliche Schriften und Briefe. Akademie-Verlag, Berlin, Germany, 1976–2004. ISBN 3-05-000075-9 (set), 3-05-000235-2 (vol. 2), 3-05-000766-4 (vol. 3), 3-05-002602-2 (vol. 4), 3-05-003471-8 (vol. 5), 3-05-004116-1 (vol. 6). ???? pp. LCCN ????

Leibniz:1991:MNT

[LHH91] Gottfried Wilhelm Leibniz, Heinz-Jürgen Hess, and J. E. Hofmann. Mathematischer, naturwissenschaftlicher und technischer Briefwechsel. 3, 1680–Juni 1683. (German) [Mathematical, scientific and technical letters. 3, 1680–June 1683], volume 3 of Sämtliche Schriften und Briefe / Gottfried Wilhelm Leibniz; Reihe 3. Akademie-Verlag, Berlin, Germany, 1991. ISBN 3-05-000766-4. lix + 895 pp. LCCN ????

Liebenau:1987:MSM

[Lie87] Jonathan Liebenau. Medical science and medical industry: the formation of the American pharmaceutical industry. Studies in business history. MacMillan Press, Basingstoke, Hampshire, 1987. ISBN 0-333-41742-9 (hardcover), 0-8018-3356-6. ix + 207 pp. LCCN HD9666.5 .L54.

Liebenau:1988:CNT

[Lie88] Jonathan Liebenau, editor. The challenge of new technology: innovation in British business since 1850. Business history series. Gower, Aldershot, Hants, England, 1988. ISBN 0-566-05147-8. x + 155 pp. LCCN HC260.T4 C46 1988. US$60.00 (U.S. : est.).

Lightman:1987:OAV

[Lig87] Bernard V. Lightman. The origins of agnosticism: Victorian unbelief and the limits of knowledge. The Johns Hopkins University Press, Baltimore, MD, USA, 1987. ISBN 0-8018-3375-2. x + 249 pp. LCCN BL2747.2 .L54 1987.

Lilienthal:1985:LVI

[Lil85] Georg Lilienthal. Der “Lebensborn e.V.”: ein Instrument nationalsozialistischer Rassenpolitik. (German) [The life font: an instrument of National Socialist racial policy], volume 1 of Forschungen zur neueren Medizin- und Biologiegeschichte. G. Fischer, Stuttgart, Germany, 1985. ISBN 3-437-10939-1. ISSN 0176-330X. vii + 264 pp. LCCN D810.C4 L48 1985.

Lindqvist:1984:TTI

[Lin84] Svante Lindqvist. Technology on trial: the introduction of steam power technology into Sweden, 1715–1736, volume 1 of Uppsala studies in history of science, 0282-1036: 1. Almqvist och Wiksell, Stockholm, Sweden, 1984. ISBN 91-86836-00-5 (paperback), 91-22-00716-4 (hardcover). 392 pp. LCCN ????

Lindgren:1990:GFD

[Lin90] Michael Lindgren. Glory and failure: the difference engines of Johann Müller, Charles Babbage and Georg and Edvard Scheutz. The MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, [2nd] edition, 1990. ISBN 0-262-12146-8. 414 pp. LCCN QA75 .L493 1990.

Lindberg:1992:BWS

[Lin92] David C. Lindberg. The beginnings of Western science: the European scientific tradition in philosophical, religious, and institutional context, 600 B.C. to A.D. 1450. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-48230-8 (hardcover), 0-226-48231-6 (paperback). xviii + 455 pp. LCCN Q124.95 .L55

1992. URL http://www.loc.
gov/catdir/description/
uchi051/91037741.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/91037741-t.html.

Lindqvist:1993:CPH

[Lin93] Svante Lindqvist, editor. Center on the periphery: historical aspects of 20th-Century Swedish physics, volume 17 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1993. ISBN 0-88135-157-1. lv + 516 pp. LCCN QC9.S8 C46 1993. US$50.00.

Lindemann:1996:HHE

[Lin96] Mary Lindemann. Health and healing in Eighteenth-Century Germany, volume 114th series, 4 of The Johns Hopkins University studies in historical and political science. The Johns Hopkins University Press, Baltimore, MD, USA, 1996. ISBN 0-8018-5281-1 (hardcover). xiii + 505 pp. LCCN RA418.3.G3

                                        L54 1996. URL http://www.
                                        loc.gov/catdir/bios/
                                        jhu051/95053137.html;
                                        http://www.loc.gov/
                                        catdir/description/
                                        jhu051/95053137.html.

Litten:1991:ENA

[Lit91]  Frederick S. Litten.  Ein-

stein and the Noulens Affair. British Journal for the History of Science, 24(4):465– 467, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027131.

Liu:1992:ERT

[Liu92] Chuang Liu. Einstein and relativistic thermodynamics in 1952: a historical and critical study of a strange episode in the history of modern physics. British Journal for the History of Science, 25(2):185– 206, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027298.

Livingstone:1987:NSS

[Liv87]  David  N.

Livingstone. Nathaniel Southgate Shaler and the culture of American science. History of American science and technology series. University of Alabama Press, Tuscaloosa, AL, USA, 1987. ISBN 0-8173-0305-7. xiv + 395 pp. LCCN G69.S49 L58 1987.

Livingstone:1993:GTE

[Liv93] David N. Livingstone. The geographical tradition: episodes in the history of a contested enterprise. Blackwell, Oxford, UK, 1993. ISBN 0-631-18535-6, 0-631-18586-0 (paperback). viii + 434 pp. LCCN G80 .L54 1993.

Livingstone:1999:TCM

[Liv99] David N. Livingstone. Tropical climate and moral hygiene: the anatomy of a Victorian debate. British Journal for the History of Science, 32(1): 93–110, March 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027971.

Leddy:2009:EE

[LL09] Neven Leddy and Avi Lifschitz, editors. Epicurus in the Enlightenment, volume 2009, 12 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2009. ISBN 0-7294-0987-2 (paperback). vii + 258 pp. LCCN B512 .E68 2009.

Lucas:1994:CPM

[LLDM94] A. M. Lucas, P. J. Lucas, T. A. Darragh, and S. Maroske. Colonial pride and metropolitan expectations: the British Museum and Melbourne’s meteorites. British Journal for the History of Science, 27(1): 65–87, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027581.

Lloyd:1983:SFI

[Llo83]  G.  E.  R.  (Ge-

offrey Ernest Richard) Lloyd. Science, folklore, and ideology: studies in the life sciences in ancient Greece. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25314-4, 0-521-27307-2 (paperback). xi + 260 pp. LCCN R723 .L58

1983. URL http://www.loc.
gov/catdir/description/
cam022/82019808.html.

Lloyd:1991:MPG

[Llo91]  G.  E.  R.  (Ge-

offrey Ernest Richard) Lloyd. Methods and problems in Greek science. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-37419-7. xiv + 457 pp. LCCN Q127.G7 L593

1991. URL http://www.loc.
gov/catdir/description/
cam024/90001539.html;
http://www.loc.gov/
catdir/toc/cam023/
90001539.html.

Leibniz:1923:SSB

[LM23] Gottfried Wilhelm Leibniz and Uwe Bearb Mayer. Sämtliche Schriften und Briefe. Bd. 5. 1674–1676. (German) [Collected works and letters. Volume 5, 1674–1676]. Akad.-Verl., Berlin, Germany, 1923. ISBN 3-05-004578-7, 3-05-000075-9. xxxviii + 664 pp. LCCN ????

LopezPinero:1995:HCA

[LN95] José María López Piñero and Víctor Navarro Brotóns. Història de la ciència al País Valencià. (Spanish) [History of science in Valencia County]. Edicions Alfons el Magnànim, València, Spain, 1995. ISBN 84-7822-154-9. 661 + 38 pp. LCCN Q127.S7 L67 1995.

Loach:1998:CCU

[Loa98] Judi Loach. Le Corbusier and the creative use of mathematics. British Journal for the History of Science, 31(2): 185–215, June 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027762.

Loeffel:1987:BP

[Loe87] Hans Loeffel. Blaise Pascal, 1623–1662, volume 2 of Vita mathematica. Birkhäuser Verlag, Basel, Switzerland, 1987. ISBN 3-7643-1840-6. 175 pp. LCCN QA29.P2 L64 1987.

Longrigg:1990:RU

[Lon90] James Longrigg. Review: [untitled]. British Journal for the History of Science, 23(4): 467–468, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027205.

Lorch:1990:BRG

[Lor90] R. P. Lorch. Book review: George Molland. Thomas Bradwardine, Geometrica speculative. Latin Text and English Translation with an Introduction and Commentary. Volume XVIII in the seris Boethius. Stuttgart: Steiner, 1989. Pp.

176. ISBN 3-515-05020-5. DM48 (paperback). British Journal for the History of Science, 23(2):237–238, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026740.

Lusebrink:2004:ERP

[LP04] Hans-Jürgen Lüsebrink and Jeremy D. Popkin, editors. Enlightenment, revolution, and the periodical press, volume 2004:06 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2004. ISBN 0-7294-0841-8. v + 226 pp. LCCN PN4808 .E55 2004.

Lyell:1990:PG

[LR91] Sir Charles Lyell and M. J. S. Rudwick. Principles of geology. University of Chicago Press, Chicago, IL, USA, 1990– 1991. ISBN 0-226-49793-3 (hardcover : vol. 1), 0-226-49794-1 (paperback : vol. 1), 0-226-49796-8 (hardcover : vol. 2), 0-226-49797-6 (paperback : vol. 2), 0-226-49798-4 (hardcover : vol. 3), 0-226-49799-2 (paperback : vol. 3). ???? pp. LCCN QE26 .L956

1990. URL http://www.loc.
gov/catdir/enhancements/
fy0608/90011008-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0634/90011008-t.html.

Leibniz:1976:SSB

[LS76] Gottfried Wilhelm Leibniz and Walter Salvator. Sämtliche Schriften und Briefe: dritte Reihe: mathematischer, naturwissenschaftlicher und technischer Briefwechsel. (German) [Collected writings and correspondence: third series: mathematical, scientific, and technical letters]. Akademie-Verlag, Berlin, Germany, 1976. ISBN 3-05-000075-9 (set), 3-05-000235-2 (Vol. 2), 3-05-000766-4 (Vol. 3), 3-05-002602-2 (Vol. 4), 3-05-003471-8 (Vol. 5), 3-05-004116-1 (Vol. 6). ???? pp. LCCN ????

Levere:1990:NES

[LS90] Trevor Harvey Levere and William R. Shea, editors. Nature, Experiment, and the Sciences: Essays on Galileo and the History of Science in Honour of Stillman Drake, volume 120 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0420-9, 94-009-1878-X. ISSN 0068-0346. xiii + 349 pp. LCCN Q174 .B67

vol. 120. URL https://link.

                                        springer.com/book/10.
                                        1007/978-94-009-1878-8.

Lyell:1997:PG

[LS97] Sir Charles Lyell and James A. Secord. Principles of geology. Penguin classics. Penguin Books, Harmondsworth, UK, 1997. ISBN 0-14-043528-X. xlvii + 471 pp. LCCN QE26.2 .L94 1997.

LeRonddAlembert:1995:EDP

[LSR95]  Jean  Le

Rond d’Alembert, Richard N. Schwab, and Walter E. Rex, editors. Encyclopédie. Discours préliminaire. Preliminary discourse to the Encyclopedia of Diderot. University of Chicago Press, Chicago, IL, USA, 1995. ISBN 0-226-13476-8 (paperback). liv + 170 pp. LCCN B1933.E5 S38

                                        1995. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi052/95011562.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0608/95011562-t.html.

Lambourne:1990:CES

[LSS90] Robert J. Lambourne, Michael Shallis, and Michael Shortland. Close encounters?: science and science fiction. Adam Hilger Ltd., Bristol, UK, 1990. ISBN 0-85274-141-3. xiii + 184 pp. LCCN PN3433.6 .L36

                                        1990. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0707/90033934-d.html.

Luckin:1990:QPE

[Luc90]  Bill Luckin.  Questions of

power: electricity and environment in inter-war Britain. Manchester University Press, Manchester, UK, 1990. ISBN 0-7190-3302-0 (hardcover). viii + 200 pp. LCCN HD9685.G7 L83 1990.

Luckin:1992:BRD

[Luc92] Bill Luckin. Book review: David E. Nye. Electrifying America: Social Meanings of a New Technology. Cambridge, Mass, and London: MIT Press, 1990. Pp. xv + 479. ISBN 0-262-14048-9. £26.95. British Journal for the History of Science, 25(3): 379–380, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027280.

Luckin:1993:BRJ

[Luc93] Bill Luckin. Book review:

Jonathan Coopersmith. The Electrification of Russia, 1880– 1926. Ithaca: Cornell University Press, 1992. Pp. xii + 274. ISBN 0-8014-2723-1. $43.95. British Journal for the History of Science, 26(2): 247–248, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027111.

Lucier:1996:CCP

[Luc96a] Paul Lucier. Court and controversy: patenting science in the nineteenth century. British Journal for the History of Science, 29(2):139–154, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027830.

Luckin:1996:BRA

[Luc96b] Bill Luckin. Book review: Arne Kaijser and Marika Hedin (eds.), Nordic Energy Systems: Historical Perspectives and Current Issues. Canton, MA: Science History Publications, 1995. Pp. xi + 246, illus. ISBN 0-88135-164-4. $39.95. British Journal for the History of Science, 29(1):117– 118, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027535.

Luckin:1998:BRP

[Luc98] Bill Luckin. Book review: Penelope Gouk (ed.), Wellsprings of Achievement: Cultural and Economic Dynamics in Early Modern England and Japan. Aldershot: Variorum, 1996. Pp. viii + 271. ISBN 0-86078-465-7. £42.50. British Journal for the History of Science, 31(1): 63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027797.

Lunbeck:1994:PPK

[Lun94] Elizabeth Lunbeck. The psychiatric persuasion: knowledge, gender, and power in modern America. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-04804-5. xiv + 431 pp. LCCN RC437.5 .L89 1994. US$29.95. URL http:

//www.loc.gov/catdir/
description/prin031/
93043818.html;   http://
www.loc.gov/catdir/toc/
prin031/93043818.html.

Lindberg:1990:RSR

[LW90]  David  C.

Lindberg and Robert S. Westman, editors. Reappraisals of the scientific revolution. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-34262-7, 0-521-34804-8 (paperback). xxvii + 551 pp. LCCN Q125 .R34 1990. URL http:

//www.loc.gov/catdir/
description/cam041/
89070859.html;   http://
www.loc.gov/catdir/toc/
cam041/89070859.html.

Lynn:1999:BRL

[Lyn99] Michael R. Lynn. Book review: Lynette Hunter and Sarah Hutton (eds.), Women, Science and Medicine 1500– 1700: Mothers and Sisters of the Royal Society. Stroud: Sutton Publishing, 1997. Pp. xx + 292. ISBN 0-7509-1334-7. £40.00, $72.00 (hardcover); 0-7509-1343-6. £14.99, $22.95 (paperback). British Journal for the History of Science, 32 (2):237–251, June 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4028089.

Millar:1989:CCD

[M^{+}89] David Millar et al., editors. Chambers concise dictionary of scientists. Chambers, Cambridge, UK, 1989. ISBN 1-85296-354-9, 0-550-16032-9. 461 pp. LCCN Q141 .M54

                                        1989. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/20820593.
                                        html.

MacNeill:1983:PPT

[Mac83] William Hardy MacNeill. The pursuit of power: technology, armed force, and society since A.D. 1000. Blackwell, Oxford, UK, 1983. ISBN 0-631-13134-5. x + 405 pp. LCCN ????

MacHale:1985:GBH

[Mac85] Des MacHale. George Boole: his life and work, volume 2 of Profiles of genius series, 0332-3218. Boole Press, Dublin, Ireland, 1985. ISBN 0-906783-05-4, 0-906783-14-3 (paperback). xiii + 304 pp. LCCN QA29.B685 M33 1985. US$16.00.

MacKenzie:1988:ENH

[Mac88a] John M. MacKenzie. The empire of nature: hunting, conservation, and British imperialism. Studies in imperialism. Manchester University Press, Manchester, UK, 1988. ISBN 0-7190-2227-4. x + 340 pp. LCCN SK21 .M33 1988. US$35.00 (U.S. : est.).

MacLeod:1988:IIR

[Mac88b] Christine MacLeod. Inventing the Industrial Revolution: the English patent system, 1660– 1800. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-30104-1. xii + 302 pp. LCCN T257.P2 M33

1988. URL http://www.loc.
gov/catdir/description/
cam023/87037406.html;
http://www.loc.gov/
catdir/samples/cam031/
87037406.html;   http://
www.loc.gov/catdir/toc/
cam021/87037406.html.

Mack:1990:VES

[Mac90a] Pamela Etter Mack. Viewing the Earth: the social construction of the Landsat satellite system. Inside technology. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-13259-1. xii + 270 + 1 pp. LCCN TL796.5.U6 L375 1990.

MacLeod:1990:BRP

[Mac90b] Roy M. MacLeod. Book review: Peter Alter, The Reluctant Patron: Science and the State in Britain, 1850–1920. Oxford: Berg, 1987. ISBN 0-907582-67-2 £30. British Journal for the History of Science, 23(2):224–226, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026733.

Macleod:1995:KPM

[Mac95] Roy Macleod. ‘Kriegsgeologen and practical men’: military geology and modern memory, 1914–1918. British Journal for the History of Science, 28(4): 427–450, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027546.

Maffeo:1991:SNP

[Maf91] Sabino Maffeo. In the service of nine popes: 100 years of the Vatican Obervatory. Pontificia Academia Scientiarum, Vatican City, 1991. ISBN 88-7761-046-8. 241 pp. LCCN QB82.V32 V38513 1991.

Magnello:1996:BRJ

[Mag96a] M. Eileen Magnello. Book review: J. Rosser Matthews, Quantification and the Quest for Medical Certainty. Princeton: Princeton University Press, 1995. Pp. x + 195. ISBN 0-691-03794-9. £32.00, $39.50. British Journal for the History of Science, 29(2):244–246, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027847.

Magnello:1996:KPG

[Mag96b] M. Eileen Magnello. Karl Pearson’s Gresham lectures: W.

F. R. Weldon, speciation and the origins of Pearsonian statistics. British Journal for the History of Science, 29(1):43– 63, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027511.

Mainzer:1988:SNH

[Mai88] Klaus Mainzer. Symmetrien der Natur: ein Handbuch zur Natur- und Wissenschaftsphilosophie. (German) [Symmetry of Nature: a Handbook for Natural and Scientific Philosophy]. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 1988. ISBN 3-11-011507-7. xi + 739 pp. LCCN Q172.5.S95 M3

1988. URL http://catalog.
hathitrust.org/api/
volumes/oclc/22415725.
html.

Maiocchi:1990:CBM

[Mai90] Roberto Maiocchi. The case of Brownian motion. British Journal for the History of Science, 23(3):257–283, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026755.

Maienschein:1991:TTA

[Mai91] Jane Maienschein. Transforming traditions in American biology, 1880–1915. The Johns Hopkins University Press, Baltimore, MD, USA, 1991. ISBN 0-8018-4126-7. xi + 366 pp. LCCN QH305.2.U6 M35 1991.

Malet:1993:BRPb

[Mal93a] Antoni Malet. Book review: P. M. Harman and Alan E. Shapiro (eds.), The Investigation of Difficult Things. Essays on Newton and the History of the Exact Sciences in Honour of D. T. Whiteside. Cambridge: Cambridge University Press, 1992. Pp. xvi + 531. ISBN 0-521-37435-9. £90.00. British Journal for the History of Science, 26(3): 361–363, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027412.

Malet:1993:BRPa

[Mal93b] Antoni Malet. Book review: Peter Barker and Roger Ariew (eds.), Revolution and Continuity: Essays in the History and Philosophy of Early Modern Science. Washington, DC: The Catholic University of America Press, 1991. Pp. v + 222. ISBN 0-8132-0738-X. $42.95. British Journal for the History of Science, 26 (1):88–89, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027352.

Malet:1994:BRR

[Mal94] Antoni Malet. Book reviews: Rupert A. Hall, All Was Light:

An Introduction to Newton’s Opticks. Oxford: Clarendon Press, 1993. Pp. xviii + 252. ISBN 0-19-853985-1. £35.00. Alan E. Shapiro, Fits, Passions, and Paroxysms: Physics, Method, and Chemistry and Newton’s Theories of Colored Bodies and Fits of Easy Reflection. Cambridge: Cambridge University Press, 1993. Pp. xvii + 400. ISBN 0-521-40507-6. £45.00, $69.95. British Journal for the History of Science, 27(4):474–476, December 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027629.

Malet:1995:BRJ

[Mal95a] Antoni Malet. Book review:

J. V. Field and Frank A. J. L. James (eds.), Renaissance and Revolution: Humanists, Scholars, Craftsmen and Natural Philosophers in Early Modern Europe. Cambridge: Cambridge University Press, 1994. Pp. xvi + 291. ISBN 0-521-43427-0. £37.50, $49.95. British Journal for the History of Science, 28(2):235–236, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027681.

Malet:1995:FSB

[Mal95b] Antoni Malet. Ferran Sunyer i Balaguer, 1912–1967. (Spanish) [Ferran Sunyer in Balaguer, 1912–1967]. Societat Catalana de Matemátiques, Barcelona, Spain, 1995. ISBN 84-7283-286-4. 306 pp. LCCN QA29.S793 M35 1995.

Malet:1996:BRG

[Mal96] Antoni Malet. Book review: Galileo Galilei, Diálogo sobre los dos máximos sistemas del mundo ptolemaico y copernicano, edited and translated by Antonio Beltrán Marí. Madrid: Alianza Editorial, 1994. Pp. lxxxiii + 417. ISBN 84-206-9412-6. No price given. British Journal for the History of Science, 29(1):92–93, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027516.

Malet:1998:BRJ

[Mal98]  Antoni Malet.  Book re-

view: Jose María López Piñero and Víctor Navarro Brotons, Història de la ciència al País Valencià. València: Edicions Alfons el Magnànim, 1995. Pp. 661, illus. ISBN 84-7822-154-9. No price given. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027798.

Mallinson:2005:ECN

[Mal05]  G. J. Mallinson,  ed-

itor. The Eighteenth Century now: boundaries and perspectives, volume 2005:10 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2005. ISBN 0-7294-0866-3 (paperback). ix + 348 pp. LCCN PQ261 .E44 2005.

Manning:1988:UCS

[Man88] Thomas G. Manning. U.S.

Coast Survey vs. Naval Hydrographic Office: a 19th-Century rivalry in science and politics. History of American science and technology series. University of Alabama Press, Tuscaloosa, AL, USA, 1988. ISBN 0-8173-0390-1. xii + 202 pp. LCCN VK597.U6 M36 1988.

Mancosu:1990:RU

[Man90] Paolo Mancosu. Review: [untitled]. British Journal for the History of Science, 23(4): 475–477, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027211.

Mandelbrote:1993:DGM

[Man93a] Scott Mandelbrote. ‘A duty of the greatest moment’: Isaac Newton and the writing of biblical criticism. British Journal for the History of Science, 26 (3):281–302, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027399.

Mandelbrote:1993:BRB

[Man93b] Scott Mandelbrote. Book review: Betty Jo Teeter Dobbs.

The Janus Faces of Genius: The Role of Alchemy in Newton’s Thought. Cambridge: Cambridge University Press, 1991 [published 1992]. Pp. xii + 359, illus. ISBN 0-521-38084-7. £30.00, $47.95. British Journal for the History of Science, 26(4):491–493, December 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027472.

Mandelbrote:1995:RU

[Man95] Scott Mandelbrote. Review: [untitled]. British Journal for the History of Science, 28(4): 470–471, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027556.

Mandelbrote:1997:BRW

[Man97] Scott Mandelbrote. Book review: William R. Newman, Gehennical Fire: The Lives of George Starkey, an American Alchemist in the Scientific Revolution. Cambridge, MA: Harvard University Press, 1994. Pp. xvi + 348, illus. ISBN 0-674-34171-6. £39.95. British Journal for the History of Science, 30(1):101– 121, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027908.

Martin:1990:RU

[Mar90] Julian Martin. Review: [untitled]. British Journal for the History of Science, 23(4): 469–470, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027206.

Marsden:1992:BRG

[Mar92a] Ben Marsden. Book review:

Graham Hollister-Short and Frank A. J. L. James (eds.), History of Technology, Thirteenth Annual Volume. London and New York: Mansell, 1991. Pp. xiii + 242. ISBN 0-7201-2114-0. £40.00. British Journal for the History of Science, 25(4):475–476, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027064.

Marsden:1992:ESG

[Mar92b] Ben Marsden. Engineering science in Glasgow: economy, efficiency and measurement as prime movers in the differentiation of an academic discipline. British Journal for the History of Science, 25(3):319– 346, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027256.

Martin:1992:FBS

[Mar92c] Julian Martin. Francis Bacon, the state and the reform of natural philosophy. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-38249-1 (hardcover). xiii + 236 pp. LCCN B1198 .M36

1992. URL http://www.loc.
gov/catdir/description/
cam024/90024242.html;
http://www.loc.gov/
catdir/toc/cam028/
90024242.html.

Marsden:1993:BRD

[Mar93a]  Ben Marsden.  Book re-

view: Daniel M. Siegel, Innovation in Maxwell’s Electromagnetic Theory: Molecular Vortices, Displacement Current, and Light. Cambridge: Cambridge University Press, 1991. Pp. x + 225. ISBN 0-521-35365-3. £30.00, $49.50. British Journal for the History of Science, 26(1):116–117, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027379.

Marsden:1993:BRP

[Mar93b]  Ben Marsden.  Book re-

view: Paul Tunbridge, Lord Kelvin: His Influence on Electrical Measurements and Units. London: Peter Peregrinus Ltd, on Behalf of the Institution of Electrical Engineers, 1992. Pp. ix + 107. ISBN 0-86341-237-8. £19.00. British Journal for the History of Science, 26(3): 371–372, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027420.

Marsden:1993:BRR

[Mar93c] Ben Marsden. Book review:

Robert A. Rosenberg, Paul B. Israel, Keith A. Nier and Melodie Andrews (eds.), The Papers of Thomas A. Edison. Volume 2: From Workshop to Laboratory, June 1873 — March 1876. Baltimore and London: Johns Hopkins University Press, 1991. Pp. xlviii + 842. ISBN 0-8018-3101-6. £54.00. British Journal for the History of Science, 26(1): 118–119, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027381.

Marsden:1994:BRK

[Mar94a] Ben Marsden. Book review:

A. N. Kolmogorov and A. P. Yushkevich (eds.), Mathematics of the 19th Century: Mathematical Logic, Algebra, Number Theory, Probability Theory. Basel, Boston, Berlin: Birkhäuser, 1992. Pp. xii + 308. ISBN 3-7643-2552-6. SFr. 198.00. British Journal for the History of Science, 27(2): 236–237, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027450.

Martin:1994:BRR

[Mar94b] Julian Martin. Book review:

Roy Porter and Mikulás Teich (eds.), The Scientific Revolution in National Context. Cambridge: Cambridge University Press, 1992. Pp. xi + 305. ISBN 0-521-39510-0, £35.00 (hardcover); 0-521-39699-9, £12.99 (paperback). British Journal for the History of Science, 27 (2):228–229, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027442.

Marsden:1996:BRRb

[Mar96a] Ben Marsden. Book review: Robert A. Rosenberg, Paul B. Israel, Keith A. Nier and Martha J. King (eds.), The Papers of Thomas A. Edison. Volume 3: Menlo Park: The Early Years, April 1876 — December 1877. Baltimore and London: Johns Hopkins University Press, 1994. Pp. xlvii + 727. ISBN 0-8018-3102-4. £54.00. British Journal for the History of Science, 29(2): 247–249, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027849.

Marsden:1996:BRRa

[Mar96b] Ben Marsden. Book review: Robert Fox and Anna Guagnini (eds.), Education, Technology and Industrial Performance in Europe, 1850–1939. Cambridge: Cambridge University Press, 1993. Pp. xiii + 302. ISBN 0-521-38153-3. £35.00, $59.95. British Journal for the History of Science, 29(1): 110–112, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027531.

Martin:1996:BRP

[Mar96c] R. N. D. Martin. Book review:

Pierre Duhem, The Origins of Statics: The Sources of Physical Theory, translated from the French by Grant F. Leneaux, Victor N. Vagliente and Guy H. Wagener, with a Foreword by Stanley L. Jaki. Boston Studies in the Philosophy of Science, 123. Dordrecht: Kluwer Academic Publishers, 1991. Pp. xxxvi + 593. ISBN 0-7923-0898-0. £99.00. British Journal for the History of Science, 29 (3):362–363, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027740.

MarquinaBarrio:2010:GWC

[Mar10] Antonio Marquina Barrio, editor. Global warming and climate change: prospects and policies in Asia and Europe. Energy, climate and the environment series. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2010. ISBN 0-230-23771-1 (hardcover). xvii + 510 pp. LCCN QC903.2.E85 G56 2010.

Mason:1998:DDB

[Mas98] Haydn Trevor Mason, editor. The Darnton debate: books and revolution in the Eighteenth Century, volume 359 of Studies on Voltaire and the eighteenth century, 0435-2866. Voltaire Foundation, Oxford, UK, 1998. ISBN 0-7294-0566-4. ix + 305 pp. LCCN PQ2105.A2 S8 vol. 359.

Matthews:1995:QQM

[Mat95] J. Rosser Matthews. Quantification and the quest for medical certainty. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03794-9 (hardcover). x + 195 pp. LCCN R853.C55 M38

                                        1995. URL http://www.loc.
                                        gov/catdir/description/
                                        prin031/94024091.html;
                                        http://www.loc.gov/
                                        catdir/toc/prin031/
                                        94024091.html.

Maulitz:1990:RU

[Mau90] Russell C. Maulitz. Review: [untitled]. British Journal for the History of Science, 23(4): 463–464, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027202.

Mauskopf:1995:BRD

[Mau95] Seymour Mauskopf. Book review: David J. Hess, Science in the New Age: The Paranormal, Its Defenders and Debunkers, and American Culture. Madison: University of Wisconsin Press, 1993. Pp. xi + 243. ISBN 0-229-13820-8, £38.50 (hardcover); 0-229-13824-0, £16.50 (paperback). British Journal for the History of Science, 28 (3):369–370, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027663.

Maxwell:1984:KWR

[Max84] Nicholas Maxwell. From knowledge to wisdom: a revolution in the aims and methods of science. Blackwell, Oxford, UK, 1984. ISBN 0-631-15641-0 (paperback), 0-631-13602-9. 300 pp. LCCN BD220. US$7.95.

May:1997:IUP

[May97a] Maisie May. Inoculating the urban poor in the late eighteenth century. British Journal for the History of Science, 30(3): 291–305, September 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027863.

Mayer:1997:MSU

[May97b] Anna-Katherina Mayer. Moralizing science: the uses of science’s past in national education in the 1920s. British Journal for the History of Science, 30(1):51–70, March 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027900.

Mayer:1999:BVF

[May99] Anna-K. Mayer. ‘I have been very fortunate… ’. brief report on the BSHS Oral History Project: ‘the history of science in Britain, 1945–1965’. British Journal for the History of Science, 32(2):223–235, June 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4028085.

Mazumdar:1992:EHG

[Maz92] Pauline M. H. (Pauline Margaret Hodgson) Mazumdar. Eugenics, human genetics, and human failings: the Eugenics Society, its sources and its critics in Britain. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1992. ISBN 0-415-04424-3. x + 373 pp. LCCN HQ755.5.G7 M39 1991. US$35.00. URL http:

                                        //www.loc.gov/catdir/
                                        enhancements/fy0648/
                                        90028804-d.html.

Mazhar:1995:CAE

[Maz95a] Noor Giovanni Mazhar. Catholic attitudes to evolution in Nineteenth-Century Italian literature, volume 60 of Memorie, 0393-845X. Istituto veneto di scienze, lettere ed arti, Venezia, Italiy, 1995. ISBN 88-86166-24-9. 284 pp. LCCN BX1795.E85 M39 1995.

Mazumdar:1995:SSI

[Maz95b]  Pauline  M.

H. (Pauline Margaret Hodgson) Mazumdar. Species and specificity: an interpretation of the history of immunology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-43172-7 (hardcover). xiii + 457 pp. LCCN QR182 .M39

1995. URL http://www.loc.
gov/catdir/description/
cam025/93031219.html;
http://www.loc.gov/
catdir/samples/cam034/
93031219.html;   http://
www.loc.gov/catdir/toc/
cam023/93031219.html.

Mazzotti:1996:BRJ

[Maz96] Massimo Mazzotti. Book review: J. W. Dauben, Abraham Robinson: The Creation of Nonstandard Analysis, a Personal and Mathematical Odyssey. Princeton: Princeton University Press, 1995. Pp. xix + 559. ISBN 0-691-03745-0. £40.00, $49.50. British Journal for the History of Science, 29(4):489–490, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027708.

Mazzotti:1998:BRA

[Maz98a] Massimo Mazzotti. Book review: Antoni Malet, Ferran Sunyer i Balaguer. Barcelona: Societat Catalana de Matematiques and Societat Catalana d’Historia de la Ciència i de la Tecnica, 1995. Pp. 306. ISBN 84-7283-286-4. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027799.

Mazzotti:1998:BRL

[Maz98b] Massimo Mazzotti. Book review: Leo Corry, Modern Algebra and the Rise of Mathematical Structures. Basel: Birkhäuser, 1996. Pp. 460. ISBN 3-7643-5311-2. DM 178.00. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027800.

Mengal:1993:HDC

[MC^{+}93] Paul Mengal, Jean Céard, et al., editors. Histoire du concept de récapitulation: ontogenèse et phylogenèse en biologie et sciences humaines. (French) [History of the concept of recapitulation: ontogenesis and phylogeny in biology and social sciences]. Histoire des sciences, 1243-5155. Masson Editeur, Masson, France, 1993. ISBN 2-225-84057-1. 132 pp. LCCN QH491 .H578 1993.

McConnell:1981:HIO

[McC81] Anita McConnell. Historical instruments in oceanography: background to the oceanography collection at the Science Museum. Her Majesty’s Stationary Office, London, England, 1981. ISBN 0-11-290324-X (paperback). iv + 51 pp. LCCN GC41 .M4 1981. US$3.95.

McConnell:1982:NST

[McC82] Anita McConnell. No sea too deep: the history of oceanographic instruments. Adam Hilger Ltd., Bristol, UK, 1982. ISBN 0-85274-416-1. 162 pp. LCCN GC41 .M425 1982. US$19.50.

McCosh:1984:BCA

[McC84]  F.  W.  J.  (Fred-

erick William James) McCosh. Boussingault, chemist and agriculturist. Chemists and chemistry. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1682-X. xv + 280 + 2 pp. LCCN S584.52.B68 M33 1984.

McClellan:1985:SRS

[McC85] James E. (James Edward) Mc-Clellan. Science reorganized: scientific societies in the eighteenth century. Columbia University Press, New York, NY, USA, 1985. ISBN 0-231-05996-5, 0-231-05997-3 (paperback). xxix + 413 pp. LCCN Q10 .M38 1985.

McConnell:1992:IMW

[McC92] Anita McConnell. Instrument makers to the world: a history of Cooke, Troughton and Simms. University of York, York, UK, 1992. ISBN 1-85072-096-7. xii + 116 pp. LCCN HD9706.6 .M12 1992.

URL http://books.google.
com/books?id=w8i1AAAAIAAJ.

McCurdy:1993:INH

[McC93] Howard E. McCurdy. Inside NASA: high technology and organizational change in the U.S. space program. New series in NASA history. The Johns Hopkins University Press, Baltimore, MD, USA, 1993. ISBN 0-8018-4452-5. xiv + 215 pp. LCCN TL521.312 .M33

1993. URL http://www.loc.
gov/catdir/bios/jhu051/
92018753.html;
http://www.loc.gov/
catdir/description/
jhu051/92018753.html.

McConnell:1995:BRM

[McC95] Anita McConnell. Book review: Mari E. W. Williams, The Precision Makers: A History of the Instruments Industry in Britain and France, 1870–1939. London: Routledge, 1993. Pp. viii + 216. ISBN 0-415-03732-8. £40.00. British Journal for the History of Science, 28(1): 120–121, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027501.

McClaughlin:1998:BRD

[McC98] Trevor McClaughlin. Book review: David J. Sturdy, Science and Social Status: The Members of the Académie des Sciences, 1666–1750. Woodbridge, Suffolk: Boydell Press, 1995. Pp. xvi + 461. ISBN 0-85115-395-X. £65.00, $117.00. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027801.

McEvoy:1993:BRW

[McE93] John G. McEvoy. Book review: William H. Brock, The Fontana History of Chemistry. London: Fontana, 1992, Pp. xxiii + 744. ISBN 0-00-686173-3. £8.99. British Journal for the History of Science, 26(3): 351–353, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027403.

McEvoy:1995:BRM

[McE95] John G. McEvoy. Book review:

Marco Beretta, The Enlightenment of Matter. The Definition of Chemistry from Agricola to Lavoisier. Canton, MA: Science History Publications, 1993. Pp. xvi + 396, illus. ISBN 0-88135-152-0. $49.95. British Journal for the History of Science, 28 (1):109–111, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027493.

McGucken:1984:SSS

[McG84] William McGucken. Scientists, society, and state: the social relations of science movement in Great Britain, 1931– 1947. Ohio State University Press, Columbus, OH, USA, 1984. ISBN 0-8142-0351-5. 381 pp. LCCN Q175.52.G7 M38 1984. US$22.50.

McGuire:1995:TIN

[McG95] J. E. McGuire. Tradition and innovation: Newton’s metaphysics of nature, volume 56 of The University of Western Ontario series in philosophy of science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3617-8. xvi + 290 pp. LCCN B1299.N34 M37

                                        1995. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0821/95030104-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0821/95030104-t.html.

McIntosh:1985:BEC

[McI85] Robert P. (Robert Patrick) McIntosh. The background of ecology: concept and theory. Cambridge studies in ecology. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-24935-X. xiii + 383 pp. LCCN QH541 .M386

                                        1985. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/84027490.html;
                                        http://www.loc.gov/

catdir/toc/cam031/
84027490.html.

McKenzie:1996:BRR

[McK96] Mark McKenzie. Book review: Raymond Corbey and Bert Theunissen (eds.), Ape, Man, Apeman: Changing Views Since 1600. Evaluative Proceedings of the Symposium Ape, Man, Apeman: Changing Views Since 1600, Leiden, the Netherlands, 28 June — 1 July, 1993. Leiden: Department of Prehistory, Leiden University, 1995. Pp. 408, illus. ISBN 90-73368-05-7. Dfl. 45.00. British Journal for the History of Science, 29 (1):119–120, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027537.

McKevitt:1997:BM

[McK97a] Paul McKevitt. Back matter. British Journal for the History of Science, 30(1): 126, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027917.

McKevitt:1997:BRD

[McK97b] Paul McKevitt. Book review:

Daniel Crevier, AI: The Tumultuous History of the Search for Artificial Intelligence. London and New York: Basic Books, 1993. Pp. xiv + 386. ISBN 0-465-02997-3. £17.99, $27.50. British Journal for the History of Science, 30(1): 101–121, March 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027909.

McNally:1992:BRM

[McN92] Don McNally. Book review: Menachem Fisch and Simon Schaffer (eds.). William Whewell: A Composite Portrait. Oxford: Clarendon Press, 1991. Pp. xiv + 403. ISBN 0-19-824900-4. £47.50. British Journal for the History of Science, 25(3): 365–368, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027266.

McNaught:1995:RU

[McN95] Sarah McNaught. Review: [untitled]. British Journal for the History of Science, 28(4): 468–469, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027554.

Maclure:1988:EJW

[MD88] William Maclure and John S. Doskey. The European journals of William Maclure, volume 171 of Memoirs of the American Philosophical Society, 0065-9762. American Philosophical Society, Philadelphia, PA, USA, 1988. ISBN 0-87169-171-X. xlix + 815 + 1 pp. LCCN Q11 .P612 vol. 171.

Mendelsohn:1981:SCA

[ME81] Everett Mendelsohn and Yehuda Elkana, editors. Sciences and cultures: anthropological and historical studies of the sciences, volume 5 of Sociology of the sciences. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1981. ISBN 90-277-1234-4, 90-277-1235-2 (paperback). xvii + 270 pp. LCCN Q175.55 .S3.

Megill:1994:RO

[Meg94] Allan Megill, editor. Rethinking objectivity. Post-contemporary interventions. Duke University Press, Durham, NC, USA, 1994. ISBN 0-8223-1479-7, 0-8223-1494-0 (paperback). ix + 342 pp. LCCN BD220 .R48 1994.

Meinel:1991:KFZ

[Mei91]  Christoph  Meinel.

Karl Friedrich Zöllner und die Wissenschaftskultur der Gründerzeit: eine Fallstudie zur Genese konservativer Zivilisationskritik. (German) [Karl Friedrich Zöllner and the explosive growth of the culture of science: a case study on the genesis of a more conservative critique of civilization], volume 13 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. Sigma, Berlin, Germany, 1991.

ISBN 3-928068-04-0. 59 pp. LCCN Q127.G35 M45 1991.

Meli:1998:SDB

[Mel98] Domenico Bertoloni Meli. Shadows and deception: from Borelli’s Theoricae to the saggi of the cimento. British Journal for the History of Science, 31(4):383–402, December 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027873.

Menard:1986:OTP

[Men86] Henry W. (Henry William) Menard. The ocean of truth: a personal history of global tectonics. Princeton series in geology and paleontology. Princeton University Press, Princeton, NJ, USA, 1986. ISBN 0-691-08414-9. xiv + 353 + 24 pp. LCCN QE501.3 .M45 1986. US$29.50.

Mendoza:1990:DBE

[Men90] Eric Mendoza. Delaroche and Bérard and experimental error. British Journal for the History of Science, 23(3):285– 292, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026756.

Merrill:1989:RVN

[Mer89] Lynn L. Merrill. The romance of Victorian natural history. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1989. ISBN 0-19-505203-X. x + 288 + 12 pp. LCCN PR756.N38 M47 1989; QH138.A1. URL http://www.

loc.gov/catdir/
enhancements/fy0638/
88009956-d.html.

Merton:1993:SGS

[Mer93] Robert King Merton. On the shoulders of giants: a Shandean postscript. University of Chicago Press, Chicago, IL, USA, Post-Italianate, University of Chicago Press edition, 1993. ISBN 0-226-52086-2 (paperback). xxv + 319 pp. LCCN QC16.N7 M38

1993. URL http://www.loc.
gov/catdir/description/
uchi052/92043871.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/92043871-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/92043871-b.html.

Methuen:1993:BRJ

[Met93] Charlotte Methuen. Book review: Johannes Kepler, Gesammelte Werke, Bd XII: Theologica, Hexenprozess, Tacitus Übersetzung, Gedichte. Edited with notes by Jürgen Hübner, Helmuth Grössing, Friederike Boockmann, Friedrich Seek. Munich: CH Beck, 1990. Pp. 444. ISBN 3-406-01660-X, DM 148.00 (hardcover); 3-406-01661-8, DM 118.00 (paperback). British Journal for the History of Science, 26(2): 238–239, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027104.

Methuen:1996:BRI

[Met96a] Charlotte Methuen. Book review: Ivo Schneider, Johannes Faulhaber, 1580–1635: Rechenmeister in einer Welt des Umbruchs. Vita Mathematica, 7. Basel, Boston, Berlin: Birkhäuser, 1993. Pp. xiv + 271. ISBN 3-7643-2919-X. SFr. 88.00, DM 98.00. British Journal for the History of Science, 29(3):364– 365, September 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027742.

Methuen:1996:BRS

[Met96b]  Charlotte Methuen.  Book

review: Sachiko Kusukawa, The Transformation of Natural Philosophy: The Case of Philip Melanchthon. Cambridge: Cambridge University Press, 1995. ISBN 0-521-47347-0. £35.00. British Journal for the History of Science, 29 (1):91–92, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027515.

Methuen:1999:BRR

[Met99] Charlotte Methuen. Book review: Roger French and Andrew Cunningham, Before Science: The Invention of the Friars’ Natural Philosophy. Aldershot: Scolar Press, 1996. Pp. x + 298. ISBN 1-85928-287-3. No price given. British Journal for the History of Science, 32(2):237–251, June 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4028090.

McGrath:1986:FDP

[MF86] William J. McGrath and Sigmund Freud. Freud’s discovery of psychoanalysis: the politics of hysteria. Cornell University Press, Ithaca, NY, USA, 1986. ISBN 0-8014-1770-8. 336 pp. LCCN BF173.F85 M226 1986.

Melhado:1992:ESR

[MF92] Evan Marc Melhado and Tore Frängsmyr, editors. Enlightenment science in the romantic era: the chemistry of Berzelius and its cultural setting, volume 10 of Uppsala studies in history of science. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41775-9. xiv + 246 pp. LCCN QD22.B5

E55 1992. URL http://www.
loc.gov/catdir/
description/cam024/
91045939.html;   http://
www.loc.gov/catdir/toc/
cam029/91045939.html.

Maxwell:1995:MHS

[MGBE95] James Clerk Maxwell, Elizabeth Garber, Stephen G. Brush, and C. W. F. (C. W. Francis) Everitt. Maxwell on heat and statistical mechanics: on “avoiding all personal enquiries” of molecules. Lehigh University Press, Bethlehem, PA, USA, 1995. ISBN 0-934223-34-3. 550 pp. LCCN QC310.2 .M39 1995.

Miller:1988:MCS

[MH88] George A. (George Armitage) Miller and William Hirst, editors. The making of cognitive science: essays in honor of George A. Miller. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-34255-4. x + 284 pp. LCCN BF311 .M196

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87012492.html.

Maxwell:1990:SLP

[MH02] James Clerk Maxwell and P. M. (Peter Michael) Harman. The scientific letters and papers of James Clerk Maxwell. Cambridge University Press, Cambridge, UK, 1990–2002. ISBN 0-521-25625-9 (vol. 1), 0-521-25626-7 (vol. 2), 0-521-25627-5 (vol. 3). ???? pp. LCCN QC670 .M385 1990. URL http://www.

                                        loc.gov/catdir/
                                        description/cam023/
                                        89000452.html;   http://
                                        www.loc.gov/catdir/
                                        enhancements/fy0732/
                                        89000452-b.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        cam021/89000452.html.

Micale:1995:AHD

[Mic95] Mark S. Micale. Approaching hysteria: disease and its interpretations. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03717-5. xii + 327 pp. LCCN RC532 .M53 1995. US$29.95. URL http:

//www.loc.gov/catdir/
description/prin021/
94016596.html;   http://
www.loc.gov/catdir/toc/
prin031/94016596.html.

Michael:1998:BRJ

[Mic98]  Emily Michael.  Book re-

view: Julien Offray de la Mettrie, Machine Man and Other Writings, translated and edited by Ann Thomson, Cambridge Texts in the History of Philosophy. Cambridge: Cambridge University Press, 1996. Pp. xxx + 179. ISBN 0-521-47258-X, £35.00, $54.95 (hardcover), 0-521-47849-9, £12.95, $18.95 (paperback). British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027802.

Mills:1989:BOE

[Mil89]  Eric L. Mills.  Biological

oceanography: an early history, 1870–1960. Cornell University Press, Ithaca, NY, USA, 1989. ISBN 0-8014-2340-6. xvii + 378 pp. LCCN QH91.25 .M55 1989.

Millar:1990:CCD

[Mil90] David Mitarb Millar, editor.

Chambers concise dictionary of scientists. Chambers, Cambridge, UK, 1990. ISBN 1-85296-354-9, 0-550-16032-9. 461 pp. LCCN ????

Milton:1994:BRG

[Mil94]  J. R. Milton.  Book re-

view: George Berkeley, De motu and The Analyst. A Modern Critical Edition with Introductions and Commentary, edited and translated by Douglas M. Jesseph. The New Synthese Historical Library, 41. Dordrecht: Kluwer Academic Publishers, 1992. Pp. xi + 230. ISBN 0-7923-1520-0. £51.00, $86.00, Dfl. 150.00. British Journal for the History of Science, 27(1):116– 117, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027587.

Mills:1995:BRT

[Mil95] Eric L. Mills. Book review: Tim D. Smith, Scaling Fisheries: The Science of Measuring the Effects of Fishing, 1855–1955. Cambridge Studies in Applied Biology and Resource Management. Cambridge: Cambridge University Press, 1994. Pp. xii + 392. ISBN 0-521-39032-X. £50.00, $74.95. British Journal for the History of Science, 28(3):354– 355, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027653.

Miller:1999:UNP

[Mil99]  David Philip Miller.  The

usefulness of natural philosophy: the Royal Society and the culture of practical utility in the later eighteenth century. British Journal for the History of Science, 32(2):185– 201, June 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4028083.

Mirowski:1989:MHT

[Mir89] Philip Mirowski. More heat than light: economics as social physics, physics as nature’s economics. Historical perspectives on modern economics. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-35042-5. xii + 450 pp. LCCN HB74.P49 M57

1989. URL http://www.loc.
gov/catdir/description/
cam032/89009728.html;
http://www.loc.gov/
catdir/toc/cam031/
89009728.html.

Misa:1995:NSM

[Mis95] Thomas J. Misa. A nation of steel: the making of modern America, 1865–1925, volume new ser., no. 17 of Studies in the history of technology. The Johns Hopkins University Press, Baltimore, MD, USA, 1995. ISBN 0-8018-4967-5. xxvi + 367 pp. LCCN HD9515 .M576 1995. URL http://www.

loc.gov/catdir/bios/

jhu052/94038681.html;
http://www.loc.gov/
catdir/description/
jhu051/94038681.html.

Mitman:1992:SNE

[Mit92] Gregg Mitman. The state of nature: ecology, community, and American social thought, 1900–1950. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-53236-4 (hardcover), 0-226-53237-2 (paperback). xiv + 290 pp. LCCN QH541.26 .M58

1992. URL http://www.loc.
gov/catdir/description/
uchi051/91045638.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/91045638-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/91045638-b.html.

Malthus:1989:EPP

[MJ89] T. R. (Thomas Robert) Malthus and Patricia James. An essay on the principle of population: or a view of its past and present effects on human happiness, with an inquiry into our prospects respecting the future removal or mitigation of the evils which it occasions. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32361-4 (set), 0-521-32360-6 (vol. 1), 0-521-32363-0 (vol. 2), 0-521-24775-6. 2 v. pp. LCCN HB861 .E7

1989. URL http://www.loc.

gov/catdir/description/
cam041/87009372.html.

Marten:1998:BRC

[MK98] Jan Marten and Ivo Klaver.

Book review: Charles Lyell, Principles of Geology, edited by James A. Secord. Harmondsworth: Penguin Books, 1997. Pp. xlvii + 472. ISBN 0-14-043528-X. £9.99. British Journal for the History of Science, 31(4):469– 487, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027886.

Martin:1992:CMR

[MKW92] Ernst Martin, Peggy Aldrich Kidwell, and Michael R. Williams. The calculating machines (Die Rechenmaschinen): their history and development, volume 16 of The Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-13278-8. xvii + 367 pp. LCCN QA75 .M3513 1992.

Monardes:1989:HMC

[ML89]  Nicolás  Monardes  and

José María López Piänero. La historia medicinal de las cosas que se traen de nuestras Indias Occidentales (1565–1574). (Spanish) [The medicinal history of our Western Indians]. Ministerio de Sanidad y Consumo, Madrid, Spain, 1989.

ISBN 84-7670-191-8. 291 pp. LCCN RS169 .M65 1989.

Mockli:2010:EAR

[MM10] Daniel Möckli and Victor Mauer, editors. European–American relations and the Middle East: from Suez to Iraq. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2010. ISBN 0-415-47664-X (hardcover), 0-203-85144-7 (ebook). ???? pp. LCCN DS63.2.U5 E88 2010.

Malina:1990:AYT

[MMVZ90] Jaroslav Malina, Jaroslav Malina, Zdeněk Vašíček, and Marek Zvelebil. Archaeology yesterday and today: the development of archaeology in the sciences and humanities. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-26621-1 (hardcover), 0-521-31977-3 (paperback). xiv + 320 pp. LCCN CC100 .V3313 1990.

Molland:1990:VAP

[Mol90] A. G. Molland. A very astronomical poet. British Journal for the History of Science, 23 (1):95–98, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026804.

Molland:1992:BRH

[Mol92] George Molland. Book review: H. L. L. Busard. Jordanus de Nemore, De elementis arithmetice artis: A Medieval Treatise on Number Theory, 2 vols. Stuttgart: Franz Steiner Verlag, 1991. Pp. 372,

188. ISBN 3-515-05214-3. DM 138,-. British Journal for the History of Science, 25(4): 460, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027051.

Molland:1994:BRG

[Mol94a] George Molland. Book review: George L. McDowell and Merle A. Sokolik, The Data of Euclid: Translated from the Text of Menge, with Introduction by Richard Delahide Ferrier. Baltimore: Union Square Press, 1993. Pp. xxi + 207. ISBN 0-9635924-0-8. $40.00. British Journal for the History of Science, 27(4):473– 474, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027628.

Molland:1994:BRH

[Mol94b] George Molland. Book review:

Hubert L. L. Busard and Menso Folkerts, Robert of Chester’s (?) Redaction of Euclid’s Elements, the so-called Adelard II Version, 2 vols. Basel, Boston, Berlin: Birkhäuser Verlag, 1992. Pp. 959. ISBN 3-7643-2728-6. SFr. 348.00. Ken’ichi Takahashi, The Medieval Latin Traditions of Euclid’s Catoptrica: A Critical Edition of De speculis with an Introduction, English Translation and Commentary. Fukuolca: Kyushu University Press, 1992. Pp. viii + 373. ISBN 4-87378-299-6. No price given. British Journal for the History of Science, 27(2):223, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027438.

Molland:1995:BRB

[Mol95] George Molland. Book review: Brian Lawn, The Rise and Decline of the Scholastic ‘quaestio disputata’, With Special Emphasis on its Use in the Teaching of Medicine and Science. Education and Society in the Middle Ages and Renaissance,

2. Leiden, New York, Cologne: E. J. Brill, 1993. Pp. ix + 176. ISBN 90-04-09740-6. $51.50, Gld. 90.00. British Journal for the History of Science, 28 (1):102–103, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027488.

Molland:1996:BRJ

[Mol96] George Molland. Book review: Jens Høyrup, In Measure, Number, and Weight: Studies in Mathematics and Culture. SUNY Series in Science, Technology and Society. Albany: State University of New York Press, 1994. Pp. xviii + 430. ISBN 0-7914-1821-9. $16.95. British Journal for the History of Science, 29 (2):229–230, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027837.

Molland:1998:BRG

[Mol98a] George Molland. Book review:

Gad Freudenthal, Aristotle’s Theory of Material Substance: Heat and Pneuma, Form and Soul. Oxford: Clarendon Press, 1995. Pp. xii + 235. ISBN 0-19-824093-7. £30.00. British Journal for the History of Science, 31(3):361–375, September 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027823.

Molland:1998:BRJ

[Mol98b] George Molland. Book review:

J. L. Bergren and R. S. D. Thomas. Euclid’s Phaenomena: A Translation and Study of a Hellenistic Treatise in Spherical Astronomy. Sources and Studies in the History and Philosophy of Classical Science, 4. New York: Garland Publishing Inc., 1996. Pp. xi + 132. ISBN 0-8153-0493-5. $36.00. British Journal for the History of Science, 31(2):241–250, June 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027769.

Monfasani:1993:BRB

[Mon93] John Monfasani. Book review: Brian P. Copenhaver, Hermetica: The Greek Corpus Hermeticum and the Latin Asdepius in a New English Translation with Notes and Introduction. Cambridge: Cambridge University Press, 1992. Pp. lxxxiii + 320. ISBN 0-521-36144-3. £45.00, $69.95. British Journal for the History of Science, 26(4):487– 489, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027470.

Moore:1989:SLT

[Moo89]  Walter  John

Moore. Schrödinger, life and thought. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-35434-X. xi + 513 pp. LCCN QC16.S265 M66

                                        1989. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/88025807.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam021/
                                        88025807.html.

Moore:1992:BRJ

[Moo92]  Patrick Moore.  Book re-

views: J. A. Bennett. Church, State and Astronomy in Ireland: 200 Years of Armagh Observatory. Armagh and Belfast: Armagh Observatory in association with The Institute of Irish Studies, The Queen’s University of Belfast, 1990. Pp. vi + 277. ISBN (hardcover) 0-9504478-1-1 (Armagh Observatory); 0-85389-350-0 (Queen’s University). ISBN (paperback) 0-9504478-2-X (Armagh Observatory); 0-85389-351-9 (Queen’s University). £17.95 (hardcover). British Journal for the History of Science, 25 (2):284, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027322.

Morsink:1982:AGA

[Mor82] Johannes Morsink. Aristotle: On the generation of animals: a philosophical study. University Press of America, Washington, DC, USA, 1982. ISBN 0-8191-2606-3, 0-8191-2607-1 (paperback). viii + 184 pp. LCCN QL41.A717 M67 1982. US$19.50, US$9.50.

Moreau:1984:CCA

[Mor84a] René Moreau. The computer comes of age: the people, the hardware, and the software. MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-13194-3. x + 227 pp. LCCN QA76.17 .M6713 1984.

Morrell:1984:GSE

[Mor84b] Jack Morrell, editor. Gentlemen of science: early correspondence of the British Association for the Advancement of Science, volume 30 of Camden fourth series. Offices of the Royal Historical Society, University College London, London, UK, 1984. ISBN 0-86193-103-3. 382 pp. LCCN DA20 .R913 vol. 30. US$10.00.

Morselli:1984:AAS

[Mor84c] Mario Morselli. Amedeo Avogadro, a scientific biography. Chemists and chemistry. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1624-2. xi + 375 pp. LCCN QD22.A8 M67 1984.

Morris:1989:ASR

[Mor89] Peter John Turnbull Morris. The American synthetic rubber research program. The chemical sciences in society series. University of Pennsylvania Press, Philadelphia, PA, USA, 1989. ISBN 0-8122-8205-1. xii + 191 pp. LCCN TS1925 .M58 1989.

Moran:1990:BRN

[Mor90a] Bruce T. Moran. Book reviews: Nicholas H. Clulee. John Dee’s Natural Philosophy: Between Science and Religion. London and New York: Routledge, 1988. Pp. xiv + 347. ISBN 0-415-00625-2. £40.00. ISBN 0-415-03122-2. £14.95. British Journal for the History of Science, 23(3):361–363, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026783.

Morgan:1990:HEI

[Mor90b] Mary S. Morgan. The history of econometric ideas. Historical perspectives on modern economics. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37398-0, 0-521-42465-8. xv + 296 pp. LCCN HB139 .M66 1990.

Morris:1990:HWS

[Mor90c] P. R. (Peter Robin) Morris. A history of the world semiconductor industry, volume 12 of IEE history of technology series. P. Peregrinus on behalf of the Institution of Electrical Engineers, London, UK, 1990. ISBN 0-86341-227-0. ix + 171 pp. LCCN HD9696.S42 M67 1990.

Morris:1990:BRR

[Mor90d] Peter Morris. Book review:

Raymond G. Stokes. Divide and Prosper: The Heirs of I. G. Farben under Allied Authority. Berkeley: University of California Press, 1988. Pp. xiv + 290. ISBN 0-520-06248-5. £24.55. British Journal for the History of Science, 23(3): 370–371, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026789.

Morton:1990:LNP

[Mor90e] A. Q. Morton. Lectures on natural philosophy in London, 1750–1765: S. C. T. Demainbray (1710–1782) and the ’inattention’ of his countrymen. British Journal for the History of Science, 23(4):411– 434, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027199.

Moran:1991:AWG

[Mor91a] Bruce T. Moran. The alchemical world of the German court: occult philosophy and chemical medicine in the circle of Moritz of Hessen, 1572–1632, volume Heft 29 of Sudhoffs Archiv. Beihefte, 0341-0773. F. Steiner Verlag, Stuttgart, Germany, 1991. ISBN 3-515-05369-7. 193 pp. LCCN BF1434.G5 M67 1991. DM58.00.

Moran:1991:CPE

[Mor91b] Bruce T. Moran. Chemical pharmacy enters the university: Johannes Hartmann and the didactic care of chymiatria in the early Seventeenth Century, volume new ser., no. 14 of Publication / American Institute of the History of Pharmacy. American Institute of the History of Pharmacy, Madison, WI, USA, 1991. ISBN 0-931292-23-9 (paperback), 0-931292-24-7. vii + 88 pp. LCCN RS65 .M67 1991.

Morris:1991:BRT

[Mor91c] Peter J. T. Morris. Book review: Trevor I. Williams. Robert Robinson: Chemist Extraordinary. Oxford: Oxford University Press, 1990. Pp. viii + 201. ISBN 0-19-858180-7. British Journal for the History of Science, 24(1):111– 112, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027028.

Morrell:1992:BRP

[Mor92a] Jack Morrell. Book review: Peter Crowcroft. Elton’s Ecologists: A History of the Bureau of Animal Population. Chicago and London: University of Chicago Press, 1991. Pp. xx + 177, illus. ISBN 0-226-12146-1, £27.95, $35.00 (hardcover); 0-226-12148-8, £12.75, $14.95 (paperback). British Journal for the History of Science, 25 (4):488–489, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027076.

Morrell:1992:GT

[Mor92b]  Jack Morrell.  Geological

thrusting. British Journal for the History of Science, 25(2): 241–246, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027301.

Morris:1992:BRD

[Mor92c] Peter Morris. Book review:

David H. Whiffen. The Royal Society of Chemistry: The First 150 Years. London: Royal Society of Chemistry, 1991. Pp. viii + 270. ISBN 0-85186-294-2. £14.95. British Journal for the History of Science, 25(3): 375–376, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027275.

Morris:1992:BRJ

[Mor92d]  Peter Morris.  Book re-

view: John W. Servos. Physical Chemistry from Ostwald to Pauling. The Making of a Science in America. Princeton, New Jersey: Princeton University Press, 1990. Pp. 402 + xxiii. British Journal for the History of Science, 25(3):373– 374, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027273.

Morris:1992:BM

[Mor92e] Peter J. T. Morris. Back matter. British Journal for the History of Science, 25(1): i–v, March 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027010.

Morus:1992:BRG

[Mor92f] Iwan Morus. Book review: Geoffrey Cantor, David Gooding and Frank A. J. L. James. Faraday. London: Macmillan, 1991. Pp. xiii + 109. ISBN 0-333-54291-6. £9.99. British Journal for the History of Science, 25(4):472–473, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027062.

Morus:1993:BRM

[Mor93] Iwan Rhys Morus. Book review: Marcello Pera, The Ambiguous Frog: The Galvani-Volta Controversy on Animal Electricity. Princeton, New Jersey: Princeton University Press, 1992. Pp. xxvi + 203. ISBN 0-691-08512-9. British Journal for the History of Science, 26(1):92–93, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027356.

Morris:1994:FJB

[Mor94] Susan W. Morris. Fleeming Jenkin and the origin of species: a reassessment. British Journal for the History of Science, 27 (3):313–343, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027601.

Morton:1995:CPN

[Mor95a] Alan Q. Morton. Concepts of power: natural philosophy and the uses of machines in mid-eighteenth-century London. British Journal for the History of Science, 28(1):63– 78, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027484.

Morton:1995:I

[Mor95b] Alan Q. Morton. Introduction. British Journal for the History of Science, 28(1):1– 3, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027481.

Morus:1995:BRM

[Mor95c] Iwan Rhys Morus. Book review: Maurice Crosland, Science Under Control: The French Academy of Sciences, 1795– 1914. Cambridge: Cambridge University Press, 1992. Pp. xix + 454. ISBN 0-521-41373-7. £60.00, $120.00. British Journal for the History of Science, 28(2):241–242, June 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027687.

Morus:1996:BRC

[Mor96a]  Iwan  Rhys

Morus. Book review: Christine Blondel and Matthias Dörries (eds.), Restaging Coulomb: Usages, controverses et réplications autour de la balance de torsion. Biblioteca di Nuncius, Studi e Testi, 15. Florence: Leo S. Olschki, 1994. Pp. 167. ISBN 88-222-4196-7. 37,000 lire. British Journal for the History of Science, 29 (1):99–101, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027522.

Morus:1996:MNS

[Mor96b] Iwan Rhys Morus. Manufacturing nature: science, technology and Victorian consumer culture. British Journal for the History of Science, 29(4): 403–434, December 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027698.

Morus:1998:BRH

[Mor98] Iwan Rhys Morus. Book review: Hermann von Helmholtz, Science and Culture: Popular and Philosophical Essays, edited with an introduction by David Cahan. Chicago: University of Chicago Press, 1995. Pp. xviii + 418. ISBN 0-226-32658-6, £41.50, $52.00 (hardcover); 0-226-32659-4, £14.25, $17.95 (paperback). British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027803.

Morus:1999:BRK

[Mor99] Iwan Rhys Morus. Book review: K. G. Beauchamp, Exhibiting Electricity. IEE History of Technology Series, 21. London: Institution of Electrical Engineers, 1997. Pp. xii + 337. ISBN 0-85296-895-7. £45.00. British Journal for the History of Science, 32(1): 111–124, March 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027976.

Moss:1993:NHR

[Mos93]  Jean Dietz Moss.  Novel-

ties in the heavens: rhetoric and science in the Copernican controversy. University of Chicago Press, Chicago, IL, USA, 1993. ISBN 0-226-54234-3, 0-226-54235-1 (paperback). xiv + 353 pp. LCCN QB29 .M67 1993. URL http://www.

                                        loc.gov/catdir/
                                        description/uchi052/
                                        92021608.html;   http://
                                        www.loc.gov/catdir/toc/
                                        uchi051/92021608.html.

Moyer:1983:APT

[Moy83] Albert E. Moyer. American physics in transition: a history of conceptual change in the late Nineteenth Century, volume 3 of The History of modern physics, 1800–1950. Tomash Publishers, Los Angeles, 1983. ISBN 0-938228-06-4. xx + 218 pp. LCCN QC9.U5 M68 1983. US$30.00 (est.).

Moyal:1986:BSL

[Moy86] Ann Mozley Moyal. A bright and savage land: scientists in colonial Australia. Collins, London, UK, 1986. ISBN 0-00-217555-X. 192 pp. LCCN QH26 .M68 1986.

Mehra:1982:DQM

[MR82a] Jagdish Mehra and Helmut Rechenberg. The discovery of quantum mechanics, 1925, volume 2 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90674-6. vi + 355 pp. LCCN QC173.98 .M44 vol. 2. See other volumes in this series [MR82c, MR82b, MR82f, MR82g, MR87a, MR00].

Mehra:1982:FMM

[MR82b] Jagdish Mehra and Helmut Rechenberg. The formulation of matrix mechanics and its modifications, 1925–1926, volume 3 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90675-4, 3-540-90675-4. viii + 334 pp. LCCN QC173.98 .M44 vol.

3. URL http://www.loc.gov/
catdir/enhancements/
fy1005/82003253-d.html;
http://www.loc.gov/
catdir/enhancements/
fy1005/82003253-t.html. See

other volumes in this series [MR82a, MR82c, MR82f, MR82g, MR87a, MR00].

Mehra:1982:FEQ

[MR82c] Jagdish Mehra and Helmut Rechenberg. The fundamental equations of quantum mechanics, 1925–1926. The reception of the new quantum mechanics, 1925–1926, volume 4 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90680-0. viii + 322 pp. LCCN QC173.98 .M44 vol. 4. See other volumes in this series [MR82a, MR82b, MR82f, MR82g, MR87a, MR00].

Mehra:1982:HDQa

[MR82d] Jagdish Mehra and Helmut Rechenberg. The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-90675-4 (Berlin), 0-387-90675-4 (New York). vi + 334 pp. LCCN QC173.98 .M44 vol. 3.

Mehra:1982:HDQb

[MR82e] Jagdish Mehra and Helmut Rechenberg. The historical development of quantum theory. Vol. 4,1, The fundamental equations of quantum mechanics 1925– 1926. Vol. 4,2, The reception of the new quantum mechanics 1925–1926. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90680-0, 3-540-90680-0. viii + 222 pp. LCCN ????

Mehra:1982:QTPa

[MR82f] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 1) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-90642-8 (Berlin), 0-387-90642-8 (New York). 372 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82g, MR87a, MR00].

Mehra:1982:QTPb

[MR82g] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 2) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90667-3 (New York), 3-540-90667-3 (Berlin). 375– 878 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82f, MR87a, MR00].

Mehra:1987:ESR

[MR87a] Jagdish Mehra and Helmut Rechenberg. Erwin Schrödinger and the rise of wave mechanics.

Part 1, Schrödinger in Vienna and Zurich 1887–1925, volume 5, part 1 of The Historical Development of Quantum Theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96284-0 (vol. 1, New York), 3-540-96284-0 (vol. 1, Berlin), 0-387-96377-4 (v. 2, New York), 3-387-96377-4 (v. 2, Berlin). xx + 366 pp. LCCN QC173.98 .M44 vol. 5. US$48.00, US$49.50. See other volumes in this series [MR82a, MR82c, MR82b, MR82f, MR82g, MR00].

Mehra:1987:FEQ

[MR87b] Jagdish Mehra and Helmut Rechenberg. The fundamentals equations of quantum mechanics, 1925–1926. Part 2, The reception of the new quantum mechanics, 1925–1926, volume 4 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-90680-0 (New York), 3-540-90680-0 (Berlin). viii + 322 pp. LCCN ????

Mendenhall:1989:ASE

[MR89] Thomas C. (Thomas Corwin) Mendenhall and Richard Rubinger. An American scientist in early Meiji Japan: the autobiographical notes of Thomas C. Mendenhall, volume 35 of Asian studies at Hawaii. University of Hawaii Press, Honolulu, HI, USA, 1989. ISBN 0-8248-1177-1. xvii + 89 pp. LCCN QC16.M483 .A3 1989; DS21 .A831 no.35. US$11.00. URL http://

catalog.hathitrust.org/
api/volumes/oclc/
18521314.html.

Miller:1996:VEV

[MR96] David Philip Miller and Peter Hanns Reill, editors. Visions of empire: voyages, botany, and representations of nature. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-48303-4 (hardcover). xix + 370 pp. LCCN QH11 .V57

1996. URL http://www.loc.
gov/catdir/description/
cam027/95019085.html;
http://www.loc.gov/
catdir/toc/cam021/
95019085.html.

Mehra:2000:CQM

[MR00] Jagdish Mehra and Helmut Rechenberg. The completion of quantum mechanics, 1926– 1941, volume 6 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-95086-9 (part 1), 0-387-95182-2 (part 2). xxxvi + 1612 pp. LCCN QC173.98 .M44 vol. 6. See other volumes in this series [MR82a, MR82c, MR82b, MR82f, MR82g, MR87a].

MartinezAlier:1987:EEE

[MS87] Juan Martínez Alier and Klaus Schlüpmann. Ecological economics: energy, environment, and society. Blackwell, Oxford, UK, 1987. ISBN 0-631-15739-5. ix + 286 pp. LCCN HD85.S7 M363513 1987. US$39.95.

Miller:1995:HMM

[MS95] Hugh Miller and Michael Shortland. Hugh Miller’s memoir: from stonemason to geologist. Edinburgh University Press, Edinburgh, Scotland, 1995. ISBN 0-7486-0521-5. v + 266 pp. LCCN QE22.M5 A3 1995.

McGuire:1983:CPQ

[MT83] J. E. McGuire and Martin Tamny. Certain philosophical questions: Newton’s Trinity notebook. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-23164-7. xii + 519 pp. LCCN Q175 .M417

                                        1983. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/82022200.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam031/
                                        82022200.html.

Mulligan:1991:BRS

[Mul91] Lotte Mulligan. Book review: Sarah Hutton (ed.). Henry More (1614– 1687): Tercentenary Studies, with a biography and bibliography by Robert Crocker. Dordrecht: Kluwer, 1989. Pp. xi + 254. ISBN 0-7923-0095-5. $95.00, Dfl. 185.00. British Journal for the History of Science, 24(2):250–252, June 1991.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027176.

Muller:1993:BRA

[Mül93] Gerhard H. Müller. Book review: Antonello La Vergata, L’equilibrio e la guerra della natura. Dalla teologia naturale al darwinismo. Naples: Morano, 1990. Pp. 700. L. 60,000 (paperback edition). British Journal for the History of Science, 26 (3):366–368, September 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027417.

Munday:1998:POM

[Mun98]  Pat Munday.  Politics by

other means: Justus von Liebig and the German translation of John Stuart Mill’s logic. British Journal for the History of Science, 31(4):403– 418, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027874.

Murdin:1985:UNS

[Mur85] Lesley Murdin. Under Newton’s shadow: astronomical practices in the Seventeenth Century. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-456-0. viii + 152 pp. LCCN QB33.G7 M87 1985. US$18.00. URL

http://www.loc.gov/

                                        catdir/enhancements/
                                        fy0745/85120279-d.html.

Murray:1989:GHI

[Mur89] Penelope Murray, editor. Genius: the history of an idea. Blackwell, Oxford, UK, 1989. ISBN 0-631-15785-9. ix + 235 pp. LCCN BF412 .G44 1989. US$50.00.

Murray:1992:BRJ

[Mur92a] Tim Murray. Book review: Jaroslav Malina and Zdenek V sícek. Archaeology Yesterday and Today: The Development of Archaeology in the Sciences and the Humanities, translated and edited by Marek Zvelebil. Cambridge: Cambridge University Press, 1990. Pp. xiv + 320. ISBN 0-521-26621-1, £40.00, $65 (hardcover); 0-521-31977-3, £15.00, $19.95 (paperback). British Journal for the History of Science, 25(4): 489–490, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027077.

Murray:1992:BRP

[Mur92b] Tim Murray. Book review: Peter J. Ucko, Michael Hunter, Alan J. Clark and Andrew David, Avebury Reconsidered: From the 1660s to the 1990s. London: Unwin Hyman, 1991. Pp. xiv + 293, illus. ISBN 0-04-445919-X. £60.00. British Journal for the History of Science, 25(4):

463–464, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027055.

Murray:1994:PM

[Mur94] Tim Murray. The Piltdown man? British Journal for the History of Science, 27(1): 103–104, March 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027583.

Murray:1995:BM

[Mur95a]  Tim Murray.  Back mat-

ter. British Journal for the History of Science, 28(4):503– 514, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027575.

Murray:1995:BRM

[Mur95b] Tim Murray. Book review:

Michael A. Cremo and Richard L. Thompson, Forbidden Archeology: The Hidden History of the Human race. San Diego: Govardhan Hill Publishing/Bhativedanta Institute, 1993. Pp. xxxvii + 914. ISBN 0-9635309-8-4. £28.95. British Journal for the History of Science, 28(3):377–379, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027670.

Murray:1995:RU

[Mur95c] Tim Murray. Review: [untitled]. British Journal for the History of Science, 28(4): 491, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027572.

Murray:1996:BRB

[Mur96] Tim Murray. Book review: A. Bowdoin Van Riper, Men Among the Mammoths: Victorian Science and the Discovery of Human Prehistory. Chicago: University of Chicago Press, 1993. Pp. xv + 267. ISBN 0-226-84991-0, $51.75 (hardcover); 0-226-84922-9, $19.50 (paperback). British Journal for the History of Science, 29 (2):240–242, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027845.

Musselman:1998:SPJ

[Mus98] Elizabeth Green Musselman. Swords into ploughshares: John Herschel’s progressive view of astronomical and imperial governance. British Journal for the History of Science, 31(4): 419–435, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027875.

Maffioli:1994:GSW

[MW94]  Cesare  S.

Maffioli and Richard S. Westfall. Out of Galileo: the science of waters: 1628–1718, volume 49 of Nieuwe Nerderlandse bijdragen tot de geschiedenes der geneeskunde en der natuurwetenschappen. Erasmus Publ., Rotterdam, The Netherlands, 1994. ISBN 90-5235-071-X. xx + 509 pp. LCCN GB737 .M34 1994.

McOuat:1995:JBH

[MW95] Gordon McOuat and Mary P.

Winsor. J. B. S. Haldane’s Darwinism in its religious context. British Journal for the History of Science, 28(2):227– 231, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027678.

Nakayama:1991:STS

[Nak91] Shigeru Nakayama. Science, technology, and society in postwar Japan. Japanese studies. Kegan Paul International, London, UK, 1991. ISBN 0-7103-0428-5. xv + 259 pp. LCCN Q127.J3 N27 1991.

Napier:1990:R

[Nap90] John Napier. Rabdology, volume 15 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-14046-2. xxxvi + 135 pp. LCCN QA75 .N2713 1990.

Napoli:1996:HC

[Nap96] Andrea Napoli. Hobbes’ correspondence. British Journal for the History of Science, 29 (2):211–215, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027834.

Nastasi:1990:RU

[Nas90] Pietro Nastasi. Review: [untitled]. British Journal for the History of Science, 23(4): 488–489, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027221.

Numbers:1995:CTC

[NBR^{+}95] Ronald L. Numbers, Arthur I. Brown, William Bell Riley, William Vance Trollinger, George McCready Price, Harold Willard Clark, and Frank Lewis Marsh, editors. Creationism in twentiethcentury America: a ten-volumes anthology of documents, 1903–1961. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1802-2 (1), 0-8153-1803-0 (2), 0-8153-1804-9 (3), 0-8153-1805-7 (4), 0-8153-1806-5 (5), 0-8153-1807-3 (6), 0-8153-1808-1 (7), 0-8153-1809-X (8), 0-8153-1810-3 (9), 0-8153-1811-1 (10). ???? pp. LCCN ????

Neuburger:1981:HDE

[NC81] Max Neuburger and Edwin Clarke. The historical development of experimental brain and spinal cord physiology before Flourens. The Johns Hopkins University Press, Baltimore, MD, USA, 1981. ISBN 0-8018-2380-3. xxv + 391 pp. LCCN QP353 .N4813.

Neary:1998:BRE

[Nea98] Francis Neary. Book review:

Eugene Taylor, William James on Consciousness beyond the Margin. Princeton: Princeton University Press, 1996. Pp. xiii + 215. ISBN 0-691-01136-2. £27.95, $35. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027804.

Nebeker:1995:CWM

[Neb95] Frederik Nebeker. Calculating the weather: meteorology in the 20th Century, volume 60 of International geophysics series. Academic Press, New York, USA, 1995. ISBN 0-12-515175-6. vii + 255 pp. LCCN QC871 .N344 1995. URL http:

//www.loc.gov/catdir/
description/els032/
94037625.html;   http://
www.loc.gov/catdir/toc/
els032/94037625.html.

Nersessian:1987:PSC

[Ner87]  Nancy  J.

Nersessian, editor. The process of science: contemporary philosophical approaches to understanding scientific practice. Science and philosophy. Martinus Nijhoff, Dordrecht and The Hague, The Netherlands, 1987. ISBN 90-247-3425-8. xii + 221 pp. LCCN Q175.3 .P785 1987.

Neumann:1992:BRH

[Neu92] Peter M. Neumann. Book review: Hourya Sinaceur. Corps et modèles. Essai sur l’histoire de l’algèbre réelle. J. Vrin: Paris, 1991. Pp. 496. ISBN 2-7116-1038-1. 198 F. British Journal for the History of Science, 25 (3):372, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027271.

Newton:1989:PMI

[New89] Sir Isaac Newton. The preliminary manuscripts for Isaac Newton’s 1687 Principia, 1684– 1685, volume 2 of Newton manuscripts series. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33499-3. xxi + 246 pp. LCCN QA803 .N413 1989.

Newcomb:1992:BRB

[New92] Sally Newcomb. Book review: Bernhard Fritscher. Vulkanismusstreit und Geochemie: Die Bedeutung der Chemie und des Experiments in der Vulkanismus-Neptunismus-Kontroverse. Stuttgart: Franz Steiner Verlag, 1991. Pp. viii + 346. ISBN 3-515-05865-6. DM 86,-. British Journal for the History of Science, 25(4):467– 468, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027058.

Newman:1994:GFL

[New94]  William  Royall  New-

man. Gehennical fire: the lives of George Starkey, an American alchemist in the scientific revolution. Harvard University Press, Cambridge, MA, USA, 1994. ISBN 0-674-34171-6. xiv + 348 pp. LCCN QD24.S73 N49 1994.

Needham:1987:SCC

[NhTWL87] Joseph Needham, Tsuen hsuin Tsien, Ling Ed Wang, and Gwei Djen Lu. Science and civilisation in China. 05, 01, Chemistry and chemical technology, Paper and printing. Cambridge University Press, Cambridge, UK, third edition, 1987. ISBN 0-521-08690-6. xxv + 485 pp. LCCN ????

Nicholls:1988:HMP

[Nic88] Phillip A. Nicholls. Homoeopathy and the medical profession. Croom Helm, London, UK, 1988. ISBN 0-7099-1836-4. 298 pp. LCCN RX51 .N53 1988.

Nicolson:1993:BRG

[Nic93a] Malcolm Nicolson. Book review: Gregg Mitman, The State of Nature: Ecology, Community and American Social Thought, 1900–1950. Chicago and London: University of Chicago Press, 1992. Pp. xiv + 290. ISBN 0-226-53236-4, £46.50, $66.75 (hardcover); 0-226-53237-2, £18.75, $66.75 (paperback). British Journal for the History of Science, 26 (2):253–255, June 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027117.

Nicolson:1993:BRR

[Nic93b] Malcolm Nicolson. Book review: L. A. Real and J. H. Brown (eds.), Foundations of Ecology: Classic Papers with Commentaries. Chicago and London: University of Chicago Press (with the Ecological Society of America), 1991. Pp. xiv + 905. British Journal for the History of Science, 26 (1):129–130, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027391.

Nicolson:1993:BRL

[Nic93c] Malcolm Nicolson. Book review: Linda Sage, Pigment of the Imagination: A History of Phytochrome Research. San Diego: Academic Press, 1992. Pp. xx + 562. $99.50.

British Journal for the History of Science, 26(1):127– 128, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027389.

Nicolson:1994:BRM

[Nic94a] Malcolm Nicolson. Book review: Michael Shortland (ed.), Science and Nature: Essays in the History of the Environmental Sciences. BSHS Monographs, 8. Oxford: British Society for the History of Science, 1993. Pp. viii + 291. ISBN 0-906450-08-X. £10.00. British Journal for the History of Science, 27(4):484– 487, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027637.

Nicolson:1994:BRP

[Nic94b] Malcolm Nicolson. Book review: Peter J. Bowler, The Fontana History of Environmental Sciences. Fontana History of Science Series. London: Fontana Press, 1992. Pp. xv + 634. ISBN 0-00-686184-9. £8.99. British Journal for the History of Science, 27(2): 221–223, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027437.

Nicolson:1995:RU

[Nic95] Malcolm Nicolson. Review:

[untitled]. British Journal for the History of Science, 28(4): 486–487, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027567.

Nicolson:1996:BRJ

[Nic96a] Malcolm Nicolson. Book review: Joel B. Hagen, An Entangled Bank: The Origins of Ecosystem Ecology. New Brunswick, NJ: Rutgers University Press, 1992. Pp. xii + 245. ISBN 0-8135-1823-7, $38. British Journal for the History of Science, 29(1):112– 114, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027532.

Nicolson:1996:HPG

[Nic96b]  Malcolm Nicolson.  Hum-

boldtian plant geography after Humboldt: the link to ecology. British Journal for the History of Science, 29(3):289– 310, September 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027735.

Nightingale:1999:BRB

[Nig99a] Simon Nightingale. Book review: Bernard Lightman (ed.), Victorian Science in Context. Chicago: University of Chicago Press, 1997. Pp. viii + 489. 0-226-48111-5, £55.95, $70.00 (hardcover); 0-226-48112-3, £17.95, $22.50 (paperback). British Journal for the History of Science, 32 (3):363–378, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027998.

Nightingale:1999:BRM

[Nig99b] Simon Nightingale. Book review: Michael Power (ed.), Accounting and Science: Natural Inquiry and Commercial Reason. Cambridge: Cambridge University Press, 1996. Pp. xvii + 288. ISBN 0-521-55325-3, £40.00, $54.95 (hardcover); 0-521-55699-6, £14.95, $19.95 (paperback). British Journal for the History of Science, 32 (3):363–378, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027999.

Nightingale:1999:BRW

[Nig99c] Simon Nightingale. Book review: William H. Brock, Justus von Liebig: The Chemical Gatekeeper. Cambridge: Cambridge University Press, 1997. Pp. xiv + 374. ISBN 0-521-56224-4. £50.00, $79.95. British Journal for the History of Science, 32(1):111– 124, March 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027977.

Newton:1991:SIN

[NJ91] Sir Isaac Newton and Peter Jones. Sir Isaac Newton: manuscripts and papers. Chadwyck-Healey, Cambridge, England, 1991. ISBN 0-85964-226-7 (guide). 43 microfilm reels pp. LCCN Q158.5; Microfilm 95/4567 (Q).

Needham:1984:SCC

[NLH^{+}84] Joseph Needham, Gwei-Djen Lu, H. T. Huang, Francesca Bray, Christian Daniels, Nicholas K. Menzies, and Nathan Sivin. Science and civilisation in China. Vol. 6, Biology and biological technology. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-08731-7 (pt. 1), 0-521-25076-5 (pt. 2), 0-521-41999-9 (pt. 3), 0-521-65270-7 (pt. 5), 0-521-63262-5 (pt. 6). ???? pp. LCCN HD2098.NEE.

Nelson:1995:CWB

[NN95] Byron C. (Byron Christopher) Nelson and Paul Nelson. The creationist writings of Byron C. Nelson, volume 5 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1806-5. xxvi + 505 pp. LCCN BS651 .N44 1995.

Noakes:1999:TOA

[Noa99] Richard J. Noakes. Telegraphy is an occult art: Cromwell Fleetwood Varley and the diffusion of electricity to the other world. British Journal for the History of Science, 32(4):421– 459, December 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027946.

Norberg:1996:TCT

[NOF96] Arthur L. (Arthur Lawrence)

Norberg, Judy E. (Judy Elizabeth) O’Neill, and Kerry J. Freedman. Transforming computer technology: information processing for the Pentagon, 1962–1986, volume [new ser., no. 18] of Johns Hopkins studies in the history of technology. The Johns Hopkins University Press, Baltimore, MD, USA, 1996. ISBN 0-8018-5152-1. xiv + 360 pp. LCCN QA76.17 .N67

1996. URL http://www.loc.
gov/catdir/bios/jhu051/
95023820.html;
http://www.loc.gov/
catdir/description/
jhu051/95023820.html.

Nola:1988:RRS

[Nol88] Robert Nola, editor. Relativism and realism in science, volume 6 of Australasian studies in history and philosophy of science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2647-7. ix + 299 pp. LCCN Q175 .R385 1988.

Noll:1994:JCO

[Nol94] Richard Noll. The Jung cult:

origins of a charismatic movement. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03724-8. xv + 387 pp. LCCN BF109.J8 N65 1994. US$24.95. URL http://

                                        www.loc.gov/catdir/
                                        description/prin031/
                                        94004831.html;   http://
                                        www.loc.gov/catdir/toc/
                                        prin031/94004831.html.

Noll:1995:FMO

[Nol95] Steven Noll. Feeble-minded in our midst: institutions for the mentally retarded in the South, 1900–1940. University of North Carolina Press, Chapel Hill, NC, USA, 1995. ISBN 0-8078-2220-5, 0-8078-4531-0 (paperback). xiii + 254 pp. LCCN HV3006.S85 N65 1995.

Nonnoi:1991:BRT

[Non91a] Giancarlo Nonnoi. Book review: Trevor H. Leyere and William R. Shea (eds.). Nature, Experiment and the Sciences, Essays on Galileo and the History of Science in Honour of Stillman Drake. Dordrecht: Kluwer, 1990. Pp. xv + 349. ISBN 0-7923-0420-9. Dfl. 180, £60.00. British Journal for the History of Science, 24(3): 379–381, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027242.

Nonnoi:1991:BRR

[Non91b] Glancarlo Nonnoi. Book review: Richard S. Westfall. Essays on the Trial of Galileo.

Vatican City: Specola Vaticana, and Notre Dame Press, 1989. Pp. xi + 103. ISBN 0-268-00923-6. $9.95. British Journal for the History of Science, 24 (2):249–250, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027175.

Nonnoi:1993:BRW

[Non93a] Giancarlo Nonnoi. Book review: William A. Wallace, Galileo’s Logic of Discovery and Proof: The Background, Content and Use of His Appropriated Treatises on Aristotle’s Posterior Analytics. Dordrecht: Kluwer Academic, 1992. Pp. xxiii + 323. ISBN 0-7923-1577-4. £84.00, $139.00, Dfl. 240.00. British Journal for the History of Science, 26(4): 489–491, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027471.

Nonnoia:1993:BRG

[Non93b] Giancarlo Nonnoia. Book review: Galileo Galilei, Sidereus nuncius, le messager celeste, texte, traduction et notes établis par Isabella Pantin. Paris: Les Belles Lettres, 1992. Pp. civ + 115. ISBN 2-251-34505-1. 230 FF. British Journal for the History of Science, 26(3):354– 355, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027406.

Nonnoi:1994:BRF

[Non94] Glancarlo Nonnoi. Book review: F. De Gandt (ed.), L’oeuvre de Torricelli: Science Galiléenne et nouvelle géométrie. Paris: Les Belles Lettres, 1989 Pp. 247. ISBN 2-252-62032-X. No price given. British Journal for the History of Science, 27(3):365–366, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027606.

Norst:1989:FBA

[Nor89]  Marlene J. Norst.  Ferdi-

nand Bauer: the Australian natural history drawings, volume 1 of Art in natural history. British Museum (Natural History), London, UK, 1989. ISBN 0-565-01067-0 (hardcover), 0-565-01048-4 (paperback). 120 pp. LCCN NC994.5.B34 N67 1989. US$16.95, US$8.95.

North:1994:BRP

[Nor94a] J. D. North. Book review: Paola Zambelli, The Speculum astronomiae and its Enigma. Boston Studies in the Philosophy of Science 135. Dordrecht: Kluwer Academic Publishers, 1992. Pp. xvi + 352. ISBN 0-7923-1380-1. £66.00, $112.00, Dfl. 195.00. British Journal for the History of Science, 27 (2):224–225, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027439.

North:1994:FHA

[Nor94b]  John David North.  The

Fontana history of astronomy and cosmology. Fontana history of science. Fontana Press, London, UK, 1994. ISBN 0-00-686177-6 (paperback). xxvii + 697 + 8 pp. LCCN ????

Needham:1986:CCT

[NPYGDL86] Joseph Needham, Ho Ping-Yu, Lu Gwei-Djen, and Wang Ling. Chemistry and chemical technology. Part 7, Military technology: the gunpowder epic, volume 5, pt. 7 of Science and civilisation in China. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30358-3. xxxiii + 703 pp. LCCN DS721 .N39 vol. 5 pt. 7.

Needham:1986:SSCa

[NR86] Joseph Needham and C. A.

(Colin Alistair) Ronan. The shorter science and civilisation in China: an abridgement of Joseph Needham’s original text. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25272-5, 0-521-31560-3 (paperback). ix + 298 pp. LCCN ????

Newton:1984:OPI

[NS84] Sir Isaac Newton and Alan E.

(Alan Elihu) Shapiro. The optical papers of Isaac Newton. Volume I. The optical lectures 1670–1672. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25248-2. xix + 627 pp. LCCN QC353 .N52 1984.

Negrel:2002:ERE

[NS02] Eric Négrel and Jean-Paul Sermain, editors. Une expérience rhétorique: l’éloquence de la Révolution. (French) [A rhetorical experiment: the eloquence of the Revolution], volume 2002:02 of SVEC 0435-2866. Voltaire Foundation, Oxford, UK, 2002. ISBN 0-7294-0788-8. vi + 333 pp. LCCN DC155 .E9 2002.

Numbers:1992:C

[Num92] Ronald L. Numbers. The creationists. A. A. Knopf, New York, NY, USA, 1992. ISBN 0-679-40104-0. xvii + 458 pp. LCCN BS651 .N85 1992. US$27.50.

Numbers:1995:ABW

[Num95a] Ronald L. Numbers, editor. Antievolutionism before World War I, volume 1 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1802-2. xvii + 403 pp. LCCN BT712 .A57 1995.

Numbers:1995:CED

[Num95b]  Ronald  L.  Numbers,

editor. Creation-evolution debates, volume 2 of Creationism in twentieth-century America.

Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1803-0. xiv + 505 pp. LCCN BS651 .C6924 1995.

Nutton:1990:RU

[Nut90] Vivian Nutton. Review: [untitled]. British Journal for the History of Science, 23(4): 481–482, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027215.

Nutton:1993:BRD

[Nut93] Vivian Nutton. Book review:

David C. Lindberg, The Beginnings of Western Science. The European Scientific Tradition in Philosophical, Religious, and Institutional Context, 600 B.C. to A.D. 1450. Chicago and London: University of Chicago Press, 1992. Pp. xviii + 455. ISBN 0-226-48230-8, £45.50, $57.00 (hardcover); 0-226-48231-6, £15.95, $19.95 (paperback). British Journal for the History of Science, 26 (1):80–82, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027344.

Neu:1982:ICBb

[NW82a] John Neu and Magda Whitrow, editors. ISIS cumulative bibliography. 1913–65, Formed from ISIS bibliographies 1–90. 5, Civilizations and periods, 15th to 19th centuries: a bibliography of the history of science.

Mansell, London, UK, 1982. ISBN 0-7201-1643-0. xi + 593 pp. LCCN ????

Neu:1982:ICBa

[NW82b] John Neu and Magda Whitrow, editors. ISIS cumulative bibliography. 1913–65, Formed from ISIS critical bibliographies 1– 90. 4, Civilizations and periods, Prehistory to Middle Ages: a bibliography of the history of science, volume 1913–65 of ISIS cumulative bibliography. Mansell, London, UK, 1982. ISBN 0-7201-1642-2. xviii + 457 pp. LCCN ????

Nye:1990:EAS

[Nye90] David E. Nye. Electrifying America: social meanings of a new technology, 1880–1940. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-14048-9. xv + 479 pp. LCCN HD9685.U5 N94 1990.

Nye:1992:BRB

[Nye92a]  David  E.

Nye. Book review: Bill Luckin. Questions of Power: Electricity and Environment in Inter-War Britain. Manchester and New York: Manchester University Press, 1990. Pp. viii + 200. ISBN 0-7190-3302-0. £29.95. British Journal for the History of Science, 25(3): 378–379, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027279.

Nye:1992:BRL

[Nye92b]  David E. Nye.  Book re-

view: Louis C. Hunter and Lynwood Bryant. A History of Industrial Power in the United States, 1780–1930. Volume 3: The Transmission of Power. Cambridge, Mass, and London: MIT Press, 1991. Pp. xxv + 596 ISBN 0-262-08198-9. British Journal for the History of Science, 25(4):476– 477, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027065.

Nye:1993:CPT

[Nye93a] Mary Jo Nye. From chemical philosophy to theoretical chemistry: dynamics of matter and dynamics of disciplines, 1800– 1950. University of California Press, Berkeley, CA, USA, 1993. ISBN 0-520-08210-9. xvii + 328 pp. LCCN QD452 .N94

1993. URL http://www.loc.
gov/catdir/bios/ucal051/
92043114.html;
http://www.loc.gov/
catdir/description/
ucal041/92043114.html.

Nye:1993:BRB

[Nye93b] Robert A. Nye. Book reviews:

B. C. Blake-Coleman, Copper Wire and Electrical Conductors: The Shaping of a Technology. Philadelphia: Harwood Academic Publishers, 1992. Pp. xx + 283. ISBN 3-7186-5200-5. £24.00, $44.00. Paul Israel.

From Machine Shop to Industrial Laboratory: Telegraphy and the Changing Context of American Invention, 1830– 1920. Baltimore: Johns Hopkins University Press, 1992. Pp. viii + 251. ISBN 0-0818-4379-0. £29.00. British Journal for the History of Science, 26(3): 369–371, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027419.

Nye:1999:TTS

[Nye99] Mary Jo Nye. Temptations of theory, strategies of evidence: P. M. S. Blackett and the earth’s magnetism, 1947– 1952. British Journal for the History of Science, 32(1):69– 92, March 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027970.

Nyhart:1995:BTF

[Nyh95] Lynn K. Nyhart. Biology takes form: animal morphology and the German universities, 1800– 1900. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1995. ISBN 0-226-61086-1 (hardcover), 0-226-61088-8 (paperback). xiii + 414p. pp. LCCN QL799.5 .N94

1995. URL http://www.loc.
gov/catdir/description/
uchi051/95003227.html;
http://www.loc.gov/
catdir/enhancements/

fy0609/95003227-b.html;
http://www.loc.gov/
catdir/toc/uchi051/
95003227.html.

Olby:1990:CHM

[O^{+}90] Robert C. (Robert Cecil) Olby et al., editors. Companion to the history of modern science. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1990. ISBN 0-415-01988-5. xxvi + 1081 pp. LCCN Q125 .C565

1990. URL http://www.loc.
gov/catdir/enhancements/
fy0648/89010483-d.html.

OBoyle:1992:BRN

[O’B92] Cornelius O’Boyle. Book review: Nancy G. Siraisi. Medieval and Early Renaissance Medicine: An Introduction to Knowledge and Practice. Chicago and London: University of Chicago Press, 1990. Pp. xiv + 250. ISBN 0-226-76129-0, £29.95 (hardcover); 0-226-76130-4, £8.75 (paperback). British Journal for the History of Science, 25(2): 263–264, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027306.

OBoyle:1996:BRL

[O’B96a]  Cornelius  O’Boyle.

Book review: Laura Ackerman Smoller, History, Prophecy, and the Stars: The Christian Astrology of Pierre d’Ailly, 1350–1420. Princeton: Princeton University Press, 1994. Pp. xii + 233. ISBN 0-691-08788-1. £26.50, $35.00. British Journal for the History of Science, 29 (2):231–232, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027838.

OBoyle:1996:BRT

[O’B96b] Cornelius O’Boyle. Book review: Tamsyn Barton, Ancient Astrology. London and New York: Routledge, 1994. Pp. xxv + 245. ISBN 0-415-11029-7. £12.99. British Journal for the History of Science, 29(1):89–91, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027514.

OConnor:1988:FBP

[O’C88] Walter John O’Connor, editor. Founders of British physiology: a biographical dictionary, 1820–1885: a collection of biographical sketches. Manchester University Press, Manchester, UK, 1988. ISBN 0-7190-2537-0. ix + 278 pp. LCCN QP25 .O28 1988. US$25.00. URL http:

                                        //catalog.hathitrust.org/
                                        api/volumes/oclc/
                                        17506405.html.

ODonnell:1983:WRH

[O’D83]  Seán O’Donnell.  William

Rowan Hamilton: portrait of a prodigy, volume 1 of Profiles of genius series, 0332-3218. Boole Press, Dublin, Ireland, 1983. ISBN 0-906783-06-2. xvi + 224 pp. LCCN MLCM 91/04283 (Q).

Oddy:1991:BRH

[Odd91]  Nicholas Oddy.  Book re-

view: Henry Petroski. The Pencil — a History. London: Faber and Faber, 1990. Pp. xi + 434. ISBN 0-571-16182-0. £14.99. British Journal for the History of Science, 24(4): 499–501, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027158.

Ordonez:1990:CSP

[OE^{+}90] Javier Ordóñez, Alberto Elena, et al., editors. La ciencia y su público: perspectivas históricas. (Spanish) [Science and its public: historical perspectives], volume 12 of Estudios sobre la ciencia. Consejo Superior de Investigaciones Científicas, Madrid, Spain, 1990. ISBN 84-00-07056-9. xv + 443 pp. LCCN Q175.55 .C55 1990.

Ogilvie:1986:WSA

[Ogi86]  Marilyn  Bailey

Ogilvie. Women in science: antiquity through the nineteenth century: a biographical dictionary with annotated bibliography. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-15031-X. xi + 254 pp. LCCN Q141 .O34 1986.

Oresme:1985:NOB

[OH85]  Nicole  Oresme  and

Bert Hansen. Nicole Oresme and The marvels of nature: a study of his De causis mirabilium with critical edition, translation, and commentary, volume 68 of Studies and texts, 0082-5328. Pontifical Institute of Mediaeval Studies, Toronto, ON, Canada, 1985. ISBN 0-88844-068-5 (paperback). xi + 478 pp. LCCN Q153 .O7413 1985.

Olby:1991:BRD

[Olb91a] Robert Olby. Book review: Diana B. Dutton. Worse than the Disease. Pitfalls of Medical Progress. Cambridge: Cambridge University Press, 1988. Pp. xx + 528. ISBN 0-521-34023-3. £25, $29.95. British Journal for the History of Science, 24 (1):114–116, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027030.

Olby:1991:BRG

[Olb91b] Robert Olby. Book review: Georges Canguilhem. Ideology and Rationality in the History of the Life Sciences, translated by Arthur Goldhammer. Cambridge, Mass.: MIT Press, 1988, Pp. xi + 160. ISBN 0-262-03137-X. British Journal for the History of Science, 24(4): 494–496, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027153.

Olby:1993:I

[Olb93] Robert Olby. Introduction. British Journal for the History of Science, 26(1):1–3, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027336.

Oldroyd:1986:AKI

[Old86] D. R. (David Roger) Oldroyd.

The arch of knowledge: an introductory study of the history of the philosophy and methodology of science. Methuen, New York, NY, USA, 1986. ISBN 0-416-01331-7, 0-416-01341-4 (paperback). 413 pp. LCCN Q174.8 .O43 1986.

Oldroyd:1994:JHF

[Old94] David Oldroyd. James Hutton: the founder of modern geology? British Journal for the History of Science, 27 (2):213–219, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027436.

Oldroyd:1995:TH

[Old95] David Oldroyd. ‘total’ history. British Journal for the History of Science, 28(4):455– 459, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027548.

Oldroyd:1997:BRG

[Old97]  David R. Oldroyd.  Book

review: Gordon L. Herries Davies, North from the Hook: 150 Years of the Geological Survey of Ireland. Dublin: Geological Survey of Ireland, 1995. Pp. xi + 342, illus. ISBN 1-899702-00-8. £34.00, $57.00. British Journal for the History of Science, 30(1):101– 121, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027910.

Oleson:1984:GRM

[Ole84] John Peter Oleson. Greek and Roman mechanical waterlifting devices: the history of a technology, volume 16 of Phoenix: journal of the Classical association of Canada, Supplementary volume. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1693-5. ISSN 0079-1784. xiv + 458 + 149 pp. LCCN TJ840.

Olesko:1991:PCD

[Ole91] Kathryn Mary Olesko. Physics as a calling: discipline and practice in the Königsberg seminar for physics. Cornell history of science series. Cornell University Press, Ithaca, NY, USA, 1991. ISBN 0-8014-2248-5. xviii + 488 pp. LCCN QC47.G3 O43 1991.

Oliveira:1992:BRJ

[Oli92] Carlos Roberto Oliveira. Book review: Javier Ordoñez and Albert Elena (eds.). La ciencia y su público: Perspectivas históricas. Madrid: Consejo Superior de Investigaciones Científicas, 1990. Pp. xv + 443. ISBN 84-00-07056-9. British Journal for the History of Science, 25(4):470– 471, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027060.

Olmi:1992:LDM

[Olm92] Giuseppe Olmi. L’inventario del mondo: catalogazione della natura e luoghi del sapere nella prima et à moderna. (Italian) [The inventory of the world: cataloging the nature and places the knowledge from the first to modern times], volume 17 of Annali dell’Istituto storico italo-germanico. Monografia. Il Mulino, Bologna, Italy, 1992. ISBN 88-15-03647-4. 457 pp. LCCN QH83 .O46 1992. L50000.

Olson:1982:SDS

[Ols90] Richard Olson. Science deified and science defied: the historical significance of science in Western culture. University of California Press, Berkeley, CA, USA, 1982–1990. ISBN 0-520-04621-8 (vol. 1), 0-520-06846-7 (vol. 2). ???? pp. LCCN Q125 .O44 1982.

OMeara:1989:PRM

[O’M89] Dominic J. O’Meara. Pythagoras revived: mathematics and philosophy in late antiquity. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-824485-1. xii + 251 pp. LCCN B243 .O46 1989. US$49.95.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0637/88025246-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0637/88025246-t.html.

Oakley:1989:ABS

[OO89] Brian Oakley and Kenneth Owen. Alvey, Britain’s strategic computing initiative. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-15038-7. xv + 337 pp. LCCN QA76 .O225 1989.

OHara:1987:HMS

[OP87] J. G. (James G.) O’Hara and W. (Willibald) Pricha. Hertz and the Maxwellians: a study and documentation of the discovery of electromagnetic wave radiation, 1873–1894, volume 8 of IEE History of technology series. P. Peregrinus Ltd. in association with the Science Museum, London, UK, 1987. ISBN 0-86341-101-0. xiv + 154 pp. LCCN QC670 .O47 1987.

Oppenheim:1985:OWS

[Opp85] Janet Oppenheim. The other world: spiritualism and psychical research in England, 1850– 1914. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-26505-3. xii + 503 + 16 pp. LCCN BF1242.G7

O66 1985. URL http://www.
loc.gov/catdir/
description/cam022/
84005887.html;   http://
www.loc.gov/catdir/toc/
cam029/84005887.html.

Orel:1984:M

[Ore84] Vítezslav Orel. Mendel. Past masters. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1984. ISBN 0-19-287625-2, 0-19-287624-4 (paperback). vii + 111 pp. LCCN QH31.M45 O6813 1984. US$15.90, US$3.90.

Oreskes:1990:BRH

[Ore90] Naomi Oreskes. Book reviews: H. E. Le Grand. Drifting Continents and Shifting Theories. Cambridge: Cambridge University Press, 1988. Pp. vi.+ 313. ISBN 0-521-32210-3, £30.00 (hardcover). ISBN 0-521-31105-5, £10.95 (paperback). British Journal for the History of Science, 23(1): 113–115, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026819.

Overmier:1992:BMB

[OS92a]  Judith  A.

Overmier and John Edward Senior. Books and manuscripts of the Bakken. Scarecrow Press, Metuchen, NJ, USA, 1992. ISBN 0-8108-2570-8. xvii + 512 + 4 pp. LCCN Z6664.8.E4 O94 1992.

Owen:1992:HLC

[OS92b]  Sir  Richard

Owen and Phillip R. Sloan. The Hunterian lectures in comparative anatomy: May–June, 1837. Natural History Museum publications. Natural History Museum, London, UK, 1992. ISBN 0-565-01106-5 (hardcover), 0-565-01144-8 (paperback). 340 pp. LCCN M93.E02368; M93.E02370; QL808 OWE 1992.

Ogee:2006:DEC

[OS06] Frédéric Ogée and Anthony Strugnell, editors. Diderot and European culture, volume 2006:09 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2006. ISBN 0-7294-0883-3. xii + 246 pp. LCCN PQ1979 .D43 2006.

Osborne:1990:BRD

[Osb90a] Catherine Osborne. Book review: David Furley. Cosmic Problems: Essays on Greek and Roman Philosophy of Nature. Cambridge: Cambridge University Press, 1989. Pp. xiv + 258. ISBN 0-521-33330-X. British Journal for the History of Science, 23(3):367– 368, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026787.

Osborne:1990:BRO

[Osb90b] Michael A. Osborne. Book review: Oliver Bloch, Bernard Balan & Paulette Carrive (eds). Entre Forme et Histoire: La Formation de la Notion de Développment à l’Âge Classique. Paris: Méridiens Klincksieck, 1988. Pp. 310. ISBN 2-86563-212-1. (Paperback) 140 FF. British Journal for the History of Science, 23(2):221– 222, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026731.

Osborne:1994:NES

[Osb94] Michael A. Osborne. Nature, the exotic, and the science of French colonialism. Science, technology, and society. Indiana University Press, Bloomington, IN, USA, 1994. ISBN 0-253-34266-X. xvi + 216 pp. LCCN Q127.F8 O78 1994.

Osler:1991:APT

[Osl91]  Margaret  J.

Osler, editor. Atoms, pneuma, and tranquillity: Epicurean and Stoic themes in European thought. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-40048-1. xii + 304 pp. LCCN B181 .A76

1991. URL http://www.loc.
gov/catdir/description/
cam024/90045068.html;

                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        90045068.html.

Osler:1997:BRJ

[Osl97] Margaret J. Osler. Book review: J. A. van Ruler, The Crisis of Causality: Voetius and Descartes on God, Nature and Change. Brill’s Studies in Intellectual History, 66, Leiden, New York, Cologne: E. J. Brill, 1995. Pp. xiv + 353. ISBN 90-04-10371-6. $85.75. British Journal for the History of Science, 30(2):233–249, June 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027726.

Osler:1998:BRC

[Osl98] Margaret J. Osler. Book review: A. C. Crombie, Science, Art and Nature in Medieval and Modern Thought. London: Hambledon Press, 1996. Pp. xvi + 516. ISBN 1-85285-067-1. £40.00. British Journal for the History of Science, 31(1): 63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027805.

Osterbrock:1984:JKP

[Ost84] Donald E. Osterbrock. James E. Keeler, pioneer American astrophysicist, and the early development of American astrophysics. Cambridge University Press, Cambridge, UK, 1984.

ISBN 0-521-26582-7. xii + 411 pp. LCCN QB36.K37 O88

1984. URL http://www.loc.
gov/catdir/description/
cam022/84007727.html;
http://www.loc.gov/
catdir/samples/cam034/
84007727.html;   http://
www.loc.gov/catdir/toc/
cam021/84007727.html.

Oster:1993:BRRb

[Ost93a] Malcolm Oster. Book review:

A. Rupert Hall, Isaac Newton: Adventurer in Thought. Oxford: Blackwell, 1992. Pp. xv + 468. ISBN 0-631-17906-2. £19.95. British Journal for the History of Science, 26(3): 358–360, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027410.

Oster:1993:BRD

[Ost93b] Malcolm Oster. Book review:

David Goodman and Colin A. Russell (eds.), The Rise of Scientific Europe, 1500–1800. Sevenoaks, Kent: Hodder & Stoughton/The Open University, 1991. Pp. ix + 437. ISBN 0-340-55861-X. £14.99. British Journal for the History of Science, 26(2):235–236, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Oster:1993:BRM

[Ost93c] Malcolm Oster. Book review:

Marie Boas Hall, Promoting Experimental Learning: Experiment and the Royal Society, 1660–1727. Cambridge: Cambridge University Press, 1991. Pp. xiii + 207. ISBN 0-521-40503-3. £35.00, $59.95. British Journal for the History of Science, 26(1):90–91, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027354.

Oster:1993:BRRa

[Ost93d]  Malcolm Oster.  Book re-

view: Richard Kroll, Richard Ashcraft and Perez Zagorin (eds.), Philosophy, Science, and Religion in England 1640–1700. Cambridge: Cambridge University Press, 1991. Pp. xv + 287. ISBN 0-521-41095-9. £37.50, $59.95. British Journal for the History of Science, 26(1):89–90, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027353.

Oster:1994:BRR

[Ost94]  Malcolm Oster.  Book re-

view: Robert K. Merton, On the Shoulders of Giants: A Shandean Postscript, the Post-Italianate Edition, with a foreword by Umberto Eco, an afterword by Denis Donoghue and a Preface and Postface by the Author. Chicago and London: University of Chicago Press, 1993. Pp. xxv + 320. ISBN 0-226-52086-2. $14.95.

British Journal for the History of Science, 27(3):368–369, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027609.

Oster:1998:BRR

[Ost98] Malcolm Oster. Book review:

Robert Boyle, A Free Enquiry into the Vulgarly Received Notion of Nature, edited by Edward B. Davis and Michael Hunter. Cambridge Texts in the History of Philosophy. Cambridge: Cambridge University Press, 1996. Pp. xxxvi + 171. ISBN 0-521-56100-0, £37.50, $54.95 (hardcover); 0-521-56796-3 (paperback), £13.95, $18.95. British Journal for the History of Science, 31(2):241–250, June 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027770.

Oster:1999:BRR

[Ost99] Malcolm Oster. Book review:

Reijer Hooykaas, Robert Boyle: A Study in Science and Christian Belief. English translation by Harry Van Dyke. Foreword by John H. Brooke and Michael Hunter. Lanham, MD and Ancaster, ON: University Press of America and The Pascal Centre for Advanced Studies in Faith and Science, 1997. Pp. xxiv + 156. ISBN 0-7618-0708-X. $36.50. British Journal for the History of Science, 32(3):

363–378, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4028000.

Outram:1984:GCV

[Out84] Dorinda Outram. Georges Cuvier: vocation, science, and authority in post-revolutionary France. Manchester University Press, Manchester, UK, 1984. ISBN 0-7190-1077-2. viii + 299 pp. LCCN QH31.C89 O95 1984.

Outram:1991:FGM

[Out91] Dorinda Outram. Fat, gorillas and misogyny: women’s history in science. British Journal for the History of Science, 24(3): 361–367, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027235.

Outram:1994:BRJ

[Out94] Dorinda Outram. Book review: J.-C. Beaune, S. Benoit, J. Gayon, J. Roger and D. Woronoff (eds.), Buffon 88: Actes du Colloque international pour le bicentenaire de la mort de Buffon (Paris, Montbard, Dijon, 14-22 June 1988). Paris: Librairie Philosophique VRIN, 1992. Pp. ix + 771. ISBN 2-7116-9755-X. 195FF. British Journal for the History of Science, 27 (2):230, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027444.

Outram:1995:E

[Out95] Dorinda Outram. The Enlightenment, volume 7 of New approaches to European history. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-41522-5, 0-521-42534-4 (paperback). xv + 141 pp. LCCN B802 .O98

1995. URL http://www.loc.
gov/catdir/description/
cam026/94035014.html;
http://www.loc.gov/
catdir/toc/cam022/
94035014.html.

Ogborn:2010:GB

[OW10] Miles Ogborn and Charles W. J. Withers, editors. Geographies of the book. Ashgate Pub., Farnham, England, 2010. ISBN 0-7546-7850-4 (hardcover), 0-7546-9675-8 (ebook). xiv + 302 pp. LCCN Z278 .G45 2010.

Pagel:1982:JBV

[Pag82] Walter Pagel. Joan Baptista van Helmont: reformer of science and medicine. Cambridge monographs on the history of medicine. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24807-8. x + 219 pp. LCCN R524.H4 P313

1982. URL http://www.loc.
gov/catdir/description/
cam031/81024193.html;
http://www.loc.gov/
catdir/samples/cam031/
81024193.html;   http://

                                        www.loc.gov/catdir/toc/
                                        cam031/81024193.html.

Pais:1982:SLS

[Pai82] Abraham Pais. “Subtle is the Lord”: the science and the life of Albert Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1982. ISBN 0-19-853907-X. xvi + 552 + 9 pp. LCCN QC16.E5 P26 1982.

Pais:1986:IBM

[Pai86] Abraham Pais. Inward bound: of matter and forces in the physical world. Clarendon Press, New York, NY, USA, 1986. ISBN 0-19-851971-0, 0-19-851997-4 (paperback). xiv + 666 pp. LCCN QC7 .P27 1986. US$24.95.

Palladino:1990:BRA

[Pal90] Paolo Palladino. Book review: Andrew Fenton Cooper. British Agricultural Policy, 1912–1936. A Study in Conservative Politics. Manchester and New York: Manchester University Press, 1989. Pp. vii + 240. ISBN 0-7190-2886-8. £29.95. British Journal for the History of Science, 23(3):375–376, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026792.

Palladino:1991:BRA

[Pal91a] Paolo Palladino. Book review: Anna Bramwell. Ecology in the Twentieth Century: A History. New Haven and London: Yale University Press, 1989. Pp. xii + 292. ISBN 0-300-04343-0, £27.50 (hardcover); 0-300-04521-2 £9.95 (paperback). British Journal for the History of Science, 24(4):471– 473, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027133.

Palladino:1991:BRE

[Pal91b] Paolo Palladino. Book review: Eugene Cittadino. Nature as the Laboratory: Darwinian Plant Ecology in the German Empire, 1880–1990. Cambridge: Cambridge University Press, 1990. Pp. xi + 199. ISBN 0-521-34045-4. £27.50, $44.50. British Journal for the History of Science, 24 (3):389–391, September 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027248.

Palladino:1996:BRG

[Pal96a] Paolo Palladino. Book review: Gerald L. Geison, The Private Science of Louis Pasteur. Princeton: Princeton University Press, 1995. Pp. xiv + 378. ISBN 0-691-03442-7. £24.95, $29.95. British Journal for the History of Science, 29(2):239–240, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027844.

Palladino:1996:BWL

[Pal96b] Paolo Palladino. Bringing the world into the laboratory, or the (ir)resistible rise of Drosophila melanogaster. British Journal for the History of Science, 29 (2):217–221, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027835.

Palladino:1997:BRA

[Pal97] Paolo Palladino. Book review: Alan E. H. Emery and Marcia L. H. Emery, The History of a Genetic Disease: Duchenne Muscular Dystrophy or Meryon’s Disease. London: Royal Society of Medicine Press, 1995. Pp. xviii + 248, illus. ISBN 1-85315-249-8. £20.00, $40.00. British Journal for the History of Science, 30(1):101–121, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027911.

Palladino:1998:BRJ

[Pal98] Paolo Palladino. Book review: Joan H. Fujimura, Crafting Science: A Socio-history of the Quest for the Genetics of Cancer. Cambridge, MA: Harvard University Press, 1996. Pp. x + 322, illus. ISBN 0-674-17553-0. £29.95. British Journal for the History of Science, 31(3): 361–375, September 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027824.

Pancaldi:1991:BRG

[Pan91a] Giuliano Pancaldi. Book review: Grorgio Lanaro. II positivismo tra scienza e religione: Studi sulla fortuna di Comte in Gran Bretagna Milan: Franco Angeli Libri, 1990. Pp. 173. ISBN 88-204-6373-3. L. 24000. British Journal for the History of Science, 24(3): 388–389, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027247.

Pancaldi:1991:DIS

[Pan91b] Giuliano Pancaldi. Darwin in Italy: science across cultural frontiers. Indiana University Press, Bloomington, IN, USA, updated and expanded edition, 1991. ISBN 0-253-34287-2. xv + 222 pp. LCCN QH31.D2 P36313 1991.

Pang:1997:SSH

[Pan97] Alex Soojung-Kim Pang. ‘stars should henceforth register themselves’: astrophotography at the early Lick Observatory. British Journal for the History of Science, 30(2):177–202, June 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027716.

Parry:1987:PVW

[Par87] G. J. R. Parry. A Protestant vision: William Harrison and the reformation of Elizabethan England. Cambridge studies in the history and theory of politics. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-32997-3. ix + 348 pp. LCCN BX5071 .P37

                                        1987. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/86017091.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        86017091.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam031/86017091.html.

Park:1999:CTP

[Par99] Buhm Soon Park. Chemical translators: Pauling, Wheland and their strategies for teaching the theory of resonance. British Journal for the History of Science, 32(1):21–46, March 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027968.

Patterson:1983:MSC

[Pat83a] Elizabeth Chambers Patterson. Mary Somerville and the cultivation of science, 1815–1840, volume 102 of Archives internationales d’histoire des idées = International archives of the history of ideas. Martinus Nijhoff, Dordrecht and The Hague, The Netherlands, 1983. ISBN 90-247-2823-1. xiii + 264 pp. LCCN Q143.S72 P37 1983.

Pattison:1983:BVP

[Pat83b] Iain Pattison. The British veterinary profession, 1791–1948. J. A. Allen, London, UK, 1983. ISBN 0-85131-379-5. x + 207 + 4 pp. LCCN SF657 .P37 1984.

Patai:1994:JAH

[Pat94]  Raphael Patai.  The Jew-

ish alchemists: a history and source book. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03290-4. xiv + 617 pp. LCCN QD24.A2 P38 1994. US$49.50. URL http://

www.loc.gov/catdir/
description/prin031/
93035687.html;   http://
www.loc.gov/catdir/toc/
prin031/93035687.html.

Paul:1990:BRR

[Pau90] Diane B. Paul. Book review:

Robert N. Proctor. Racial Hygiene: Medicine Under the Nazis. Cambridge, MA, and London: Harvard University Press, 1988. Pp. viii + 414. ISBN 0-674-74580-9. £25.25, $34.95. British Journal for the History of Science, 23(1): 118–119, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026823.

Paul:1993:MWG

[Pau93] Erich Robert Paul. The Milky Way galaxy and statistical cosmology, 1890–1924. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-35363-7 (hardcover). xiv + 262 pp. LCCN QB857.7 .P38

1993. URL http://www.loc.
gov/catdir/description/
cam025/92036530.html;
http://www.loc.gov/
catdir/toc/cam029/
92036530.html.

Popper:1982:OUA

[PB82] Sir Karl R. (Karl Raimund) Popper and William Warren Bartley. The open universe: an argument for indeterminism: from the Postscript to the logic of scientific discovery, volume 2 of Postscript to the logic of scientific discovery / Karl R. Popper; Postscript to “The logic of scientific discovery” / Karl R. Popper; Postscript to “The logic of scientific discovery”. Hutchinson, London, UK, 1982. ISBN 0-09-146180-4 (hardcover) 0-09-146181-2 (paperback). xxii + 185 pp. LCCN QC6.2 .P6 1982.

Popper:1983:RAS

[PB83] Sir Karl Raimund Popper and William Warren Bartley. Realism and the aim of science: from the postscript of the logic of scientific discovery, volume 1 of Postscript to the logic of scientific discovery / Karl R. Popper. Hutchinson, London, UK, 1983. ISBN 0-09-151450-9, 0-09-151451-7. xxxix + 420 pp. LCCN

Q175. URL http://catalog.
hathitrust.org/api/
volumes/oclc/9771811.html.

Parnwell:1989:NOP

[PB^{+}89]  E.  C.  Parnwell,

Raymond Burns, et al. The new Oxford picture dictionary. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, english/spanish edition, 1989. ISBN 0-19-434355-3. iii + 144 pp. LCCN PE2835 .P37

1989. URL http://www.loc.
gov/catdir/enhancements/
fy1004/88034532-d.html.

Poggi:1994:RSS

[PB94] Stefano Poggi and Maurizio Bossi, editors. Romanticism in science: science in Europe, 1790–1840, volume 152 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2336-X, 94-017-2921-2. ISSN 0068-0346. xv + 245 pp. LCCN Q174 .B67 vol.

152. URL http://www.loc.
gov/catdir/enhancements/
fy0821/93001728-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0821/93001728-t.html;
https://link.springer.
com/book/10.1007/978-94-
017-2921-5.

Porter:1999:HCP

[PB99] Roy S. Porter and German E. Berrios, editors. A history of clinical psychiatry: the origin and history of psychiatric disorders. New Brunswick, N.J., London, UK, 1. paperback edition, 1999. ISBN 0-485-24211-7 (paperback), 0-485-24011-4 (hardcover). xx + 684 pp. LCCN ????

Paterson:1991:SWW

[PCA91]  Sir  Clifford  Paterson,

Sir Robert Clayton, and Joan Algar. A scientist’s war: the war diary of Sir Clifford Paterson, 1939–45: 1st September 1939–9th May 1945, volume 14 of IEE history of technology series. P. Peregrinus, in association with the Science Museum, London, UK, 1991. ISBN 0-86341-218-1. xxvi + 674 pp. LCCN TK140.P38 A3 1991.

Pedersen:1992:BN

[Ped92a] Olaf Pedersen. The book of nature. Vatican Observatory Publications, Vatican City State, 1992. ISBN 0-268-00690-3. vi + 92 pp. LCCN BL240.2 .P354 1992.

Pedersen:1992:LLH

[Ped92b] Olaf Pedersen. Lovers of learning: a history of the Royal Danish Academy of Sciences and Letters 1742–1992. Munksgaard, Copenhagen, Denmark, 1992. ISBN 87-7304-236-6. 348 pp. LCCN ????

Penrose:1989:AAB

[Pen89]  Harald Penrose.  An an-

cient air: a biography of John Stringfellow of Chard, the Victorian aeronautical pioneer. Smithsonian Institution Press, Washington, DC, USA, 1989. ISBN 0-87474-752-X. 183 pp. LCCN TL540.S837 P46 1989.

Pereira:1989:ACA

[Per89]  Michela Pereira.  The al-

chemical corpis attributed to Raymond Lull, volume 18 of Warburg Institute surveys and texts. Warburg Institute, University of London, London, UK, 1989. ISBN 0-85481-078-1 (paperback). ISSN 0266-1772. 118 pp. LCCN B765.L84; Z8527.7 .P47 1989. URL http://

catalog.hathitrust.org/
api/volumes/oclc/
20485312.html.

Pera:1992:AFG

[Per92] Marcello Pera. The ambiguous frog: the Galvani–Volta controversy on animal electricity. Princeton University Press, Princeton, NJ, USA, 1992. ISBN 0-691-08512-9. xxvi + 203 pp. LCCN QP341 .P4713 1992. US$29.95. URL http:

//www.loc.gov/catdir/
description/prin031/
91003077.html.

Peruzzi:1993:BRG

[Per93] Enrico Peruzzi. Book review:

Germana Ernst, Religione, ragione e natura: Ricerche su Tommaso Campanella e il tardo Rinascimento, Milan: Franco Angeli, 1991. Pp. 288. ISBN 88-204-6663-5. L. 36000. British Journal for the History of Science, 26(2):239–240, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027105.

Pestre:1984:PPF

[Pes84] Dominique Pestre. Physique et physiciens en France, 1918– 1940. (French) [Physics and medical doctors in France, 1918–1940]. Histoire des sciences et des techniques, 0761-1102. Editions des archives contemporaines, Paris, France, 1984. ISBN 2-903928-08-8. xv + 356 pp. LCCN QC9.F7 P47x 1984. 225.00F.

Petroski:1989:PHD

[Pet89] Henry Petroski. The pencil: a history of design and circumstance. Faber and Faber, London, UK, 1989. ISBN 0-571-16182-0. xi + 434 pp. LCCN ????

Peterfreund:1990:LST

[Pet90] Stuart Peterfreund, editor. Literature and science: theory and practice. Northeastern University Press, Boston, MA, USA, 1990. ISBN 1-55553-058-3. vi + 248 pp. LCCN PN55 .L5 1990. US$37.50 (est.).

Petroski:1994:DPC

[Pet94]  Henry  Pet-

roski. Design paradigms: case histories of error and judgment in engineering. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-46108-1, 0-521-46649-0 (paperback). xii + 209 pp. LCCN TA174 .P473

1994. URL http://www.loc.
gov/catdir/description/
cam025/93032560.html;
http://www.loc.gov/
catdir/toc/cam024/
93032560.html.

Phillips:1987:WHE

[Phi87] Vivian J. Phillips. Waveforms:

a history of early oscillography. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-274-6. xii + 259 pp. LCCN TK381 .P47

1987. URL http://www.loc.
gov/catdir/enhancements/
fy0745/87180307-d.html.

Pauli:1979:WPW

[PHvMW93] Wolfgang Pauli, Armin Hermann, K. von Meyenn, and Victor F. (Victor Frederick) Weisskopf. Wolfgang Pauli, wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a.. (German) [Wolfgang Pauli, scientific correspondence with Bohr, Einstein, Heisenberg, and others], volume 2, 6, 11 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979–1993. ISBN 0-387-08962-4, 3-540-08962-4, 0-387-13609-6, 3-540-13609-6, 0-387-54911-0, 3-540-54911-0. ???? pp. LCCN QC16.P37 W64.

Pickering:1984:CQS

[Pic84] Andrew Pickering. Constructing quarks: a sociological history of particle physics. Edinburgh University Press, Edinburgh, Scotland, 1984. ISBN 0-85224-458-4, 0-85224-535-1 (paperback). xi + 468 pp. LCCN QC793.16 .P53 1986. US$20.00. URL http:

                                        //catalog.hathitrust.org/
                                        api/volumes/oclc/
                                        11167232.html.

Picolet:1987:JPD

[Pic87]  Guy  Picolet,  edi-

tor. Jean Picard et les débuts de l’astronomie de précision au XVIIe siècle: actes du colloque du tricentenaire. (French) [Jean Picard and the origins of precision astronomy in the 17th Century: acts of the tricentennial conference]. Editions du Centre national de la recherche scientifique, Paris, France, 1987. ISBN 2-222-04104-X. 382 pp. LCCN QB36.P45 J43 1987. 150F.

Pickstone:1990:PDP

[Pic90] John V. Pickstone. A profession of discovery: Physiology in nineteenth-century history. British Journal for the History of Science, 23(2):207– 216, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026726.

Pickering:1992:RLW

[Pic92a] Andy Pickering. The Rad Lab and the world. British Journal for the History of Science, 25 (2):247–251, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027302.

Pickstone:1992:BRM

[Pic92b] John Pickstone. Book review:

M. Weatherall. In Search of a Cure. A History of Pharmaceutical Discovery. Oxford: Oxford University Press, 1990. Pp. xiv + 298. ISBN 0-19-261747-8. £19.50. British Journal for the History of Science, 25(4): 493–494, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027082.

Pickstone:1993:BRM

[Pic93a] John Pickstone. Book review: Matthew Ramsay, Professional and Popular Medicine in France, 1770–1830. The Social World of Medical Practice. Cambridge: Cambridge University Press, 1988. Pp. xvii + 406. ISBN 0-521-30517-9. £35. British Journal for the History of Science, 26(3): 365, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027415.

Pickstone:1993:WKT

[Pic93b] John V. Pickstone. Ways of knowing: towards a historical sociology of science, technology and medicine. British Journal for the History of Science, 26(4):433–458, December 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027465.

Pickstone:1995:RU

[Pic95] John V. Pickstone. Review: [untitled]. British Journal for the History of Science, 28(4): 484–486, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027566.

Pickstone:1998:BRA

[Pic98] John Pickstone. Book review: Alan Irwin and Brian Wynne (eds.), Misunderstanding Science? The Public Reconstruction of Science and Technology. Cambridge: Cambridge University Press, 1995. Pp. vii + 229. ISBN 0-521-43268-5. £35.00, $59.95. British Journal for the History of Science, 31 (1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027806.

Pickstone:1999:OPD

[Pic99] John Pickstone. Obituary: Professor Donald Cardwell (4 August 1919–8 May 1998). British Journal for the History of Science, 32(4):485– 488, December 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027948.

Pittenger:1993:ASE

[Pit93] Mark Pittenger. American socialists and evolutionary thought, 1870–1920. History of American thought and culture. University of Wisconsin Press, Madison, WI, USA, 1993. ISBN 0-299-13600-0, 0-299-13604-3 (paperback). x + 310 pp. LCCN HX83 .P58 1993. US$60.00, US$24.95.

Pugh:1991:IES

[PJP91] Emerson W. Pugh, Lyle R.

Johnson, and John H. Palmer. IBM’s 360 and early 370 systems. History of computing. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-16123-0. xx + 819 pp. LCCN QA76.8.I12 P84 1991.

Paret:1992:PIP

[PLP92] Peter Paret, Beth Irwin Lewis, and Paul Paret. Persuasive images: posters of war and revolution from the Hoover Institution archives. Princeton University Press, Princeton, NJ, USA, 1992. ISBN 0-691-03204-1. xv + 233 pp. LCCN D522.25 .P37 1992. US$29.95. URL http://www.

loc.gov/catdir/

                                        description/prin031/
                                        92004388.html;   http://
                                        www.loc.gov/catdir/toc/
                                        prin031/92004388.html.

Price:1995:SWG

[PN95] George McCready Price and Ronald L. Numbers. Selected works of George Mc-Cready Price, volume 7 of Creationism in twentieth-century America. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1808-1. xvii + 489 pp. LCCN BS651 .P77 1995.

Pocock:1988:EBR

[Poc88]  Rowland F. Pocock.  The

early British radio industry. Manchester University Press, Manchester, UK, 1988. ISBN 0-7190-2621-0. viii + 184 pp. LCCN HD9696.R363 G747 1988. US$35.00 (est.).

Poirier:1993:ALL

[Poi93]  Jean Pierre Poirier.  An-

toine Laurent de Lavoisier, 1743–1794. Pygmalion, Paris, France, 1993. ISBN 2-85704-384-8. x + 545 pp. LCCN QD22.L4 P65 1993. 178F.

Poirson:2004:AAF

[Poi04] Martial Poirson, editor. Art et argent en France au temps des premiers modernes (XVIIe et XVIIIe siècles). (French) [Art and money in France at the time of the first modern 17th and 18th Centuries], volume 2004:10 of SVEC, 0435-2866.

Voltaire Foundation, Oxford, UK, 2004. ISBN 0-7294-0849-3. xi + 341 pp. LCCN PQ245 .A77 2004.

Poole:1998:TAC

[Poo98]  Robert Poole.  Time’s al-

teration: calendar reform in early modern England. UCL Press, London, UK, 1998. ISBN 1-85728-622-7. xix + 243 pp. LCCN CE73 .P66

1998. URL http://www.loc.
gov/catdir/enhancements/
fy0909/99159352-d.html.

Poole:1999:BRN

[Poo99] Robert Poole. Book review:

Nachum Dershowitz and Edward M. Reingold, Calendrical Calculations. Cambridge: Cambridge University Press, 1997. Pp. xxi + 307. ISBN 0-521-56413-1, £40, $64.95 (hardcover); 0-521-56474-3, £14.95, $22.95 (paperback). British Journal for the History of Science, 32(1): 111–124, March 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027978.

Popper:1982:QTS

[Pop82] Karl R. (Karl Raimund) Popper. Quantum theory and the schism in physics. The Postscript to The logic of scientific discovery. Unwin Hyman, London, UK, 1982. ISBN 0-04-445555-0 (paperback), 0-09-146170-7 (hardcover). 230 pp. LCCN QC174.12 .P67 1982.

Porter:1983:ESA

[Por83] Roy Porter. The earth sciences: an annotated bibliography, volume 3; vol. 315 of Bibliographies of the history of science and technology; Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1983. ISBN 0-8240-9267-8. xix + 192 + 15 pp. LCCN Z6031 .P67 1983. US$30.00.

Porter:1986:RST

[Por86] Theodore M. Porter. The rise of statistical thinking, 1820– 1900. Princeton University Press, Princeton, NJ, USA, 1986. ISBN 0-691-08416-5. xii + 333 pp. LCCN QA276.15 .P67 1986. US$35.00.

Porter:1989:HSQ

[Por89] Roy Porter. Health for sale: quackery in England, 1660– 1850. Manchester University Press, Manchester, UK, 1989. ISBN 0-7190-1903-6. xi + 280 pp. LCCN R730 .P67 1989.

Porter:1990:BRJ

[Por90a] Charlotte M. Porter. Book review: John S. Doskey (ed.). The European Journals of William Madure. Philadelphia: American Philosophical Society, 1988. Pp. 815. ISBN: 0-87169-171-X. $40. British Journal for the History of Science, 23(1):103–104, March 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026809.

Porter:1990:BRC

[Por90b] Roy Porter. Book reviews:

Christopher Fox. Locke and the Scriblerians. Identity and Consciousness in Early Eighteenth Century Britain. Berkeley, Los Angeles and London: University of California Press, 1988. Pp. 174. ISBN 0-520-05859-3. No price given. Christopher Fox. (ed.) Psychology and Literature in the Eighteenth Century. (AMS Studies in the Eighteenth Century no. 8). New York AMS Press, 1987. Pp. 372. ISBN 0-404-61474-4. $42.50. British Journal for the History of Science, 23(1): 110–111, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026816.

Porter:1991:BRA

[Por91a] Dorothy E. Porter. Book review: Anthony Brundage. England’s ‘Prussian Minister’: Edwin Chadwick and the Politics of Government Growth, 1832–1854. University Park, Pa.: Pennsylvania State University Press, 1988. Pp. 208, ISBN 0-271-00629-3 £20.25, $22.50. British Journal for the History of Science, 24(1): 109, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027026.

Porter:1991:BRJ

[Por91b]  Dorothy E. Porter.  Book

review: Joan Austoker and Linda Bryder (eds.). Historical Perspectives on the Role of the MRC. Essays in the History of the Medical Research Council of the United Kingdom and its Predecessor, the Medical Research Committee, 1913–1953. Oxford: Oxford University Press, 1989. Pp xi + 259. ISBN 0-19-261651-X. £30.00. British Journal for the History of Science, 24(1): 112–114, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027029.

Porter:1991:BRP

[Por91c] Theodore M. Porter. Book review: Philip Mirowski, More Heat than Light: Economics as Social Physics, Physics as Nature’s Economics. Cambridge: Cambridge University Press, 1989. Pp. xii + 450. ISBN 0-521-35042-5. £35.00, $59.50. British Journal for the History of Science, 24(1):110– 111, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027027.

Porter:1995:BRM

[Por95a]  Roy  Porter.

Book review: Magda Whitrow, Julius Wagner-Jauregg (1857– 1940). London: Smith-Gordon/ Niigata-Shi, Japan: Nishimura, 1993. Pp. xxiv + 221. ISBN 1-85463-012-1. £20.00. British Journal for the History of Science, 28 (1):114–115, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027496.

Porter:1995:MLG

[Por95b] Roy Porter. Medical lecturing in Georgian London. British Journal for the History of Science, 28(1):91–99, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027486.

Porter:1995:TNP

[Por95c] Theodore M. Porter. Trust in numbers: the pursuit of objectivity in science and public life. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03776-0. xiv + 310 pp. LCCN Q175.5 .P67

1995. URL http://www.loc.
gov/catdir/description/
prin031/94021440.html;
http://www.loc.gov/
catdir/toc/prin031/
94021440.html.

Porter:1997:BRP

[Por97] Roy Porter. Book review: Paul Ilie, The Age of Minerva. Volume 1: Counter-rational Reason in the Eighteenth Century. Goya and the Paradigm of Unreason in Western Europe; Volume 2: Cognitive Discontinuities in Eighteenth Century Thought. From Body to Mind in Physiology and the Arts. Philadelphia: University of Pennsylvania Press, 1995. Pp. xiv + 411; xii + 382. ISBN 0-8122-3307-7, £52.50 (Volume 1); 0-8122-3308-5, £47.50 (Volume 2). British Journal for the History of Science, 30(2): 233–249, June 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027727.

Post:1990:BRK

[Pos90] Heinz Post. Book review: Klaus Mainzer. Symmetrian Der Natur. Ein Handbuch zur Naturund Wissenschafts-philosophie. Walter de Gruyter & Co., 1988. Pp. 739. ISBN 3-11-011507-7. DM 365. British Journal for the History of Science, 23(1): 108–109, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026813.

Poundstone:1993:PDJ

[Pou93] William Poundstone. Prisoner’s dilemma: John von Neumann, game theory, and the puzzle of the bomb. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-286162-X (paperback), 0-385-41580-X. xi + 294 + 8 pp. LCCN QA29.V66 P68 1993. US$12.95 (US$15.95 Can.).

Powers:1982:PNP

[Pow82] Jonathan Powers. Philosophy and the new physics, volume 792 of Ideas; University paperbacks. Methuen, London, UK, 1982. ISBN 0-416-73470-7, 0-416-73480-4 (paperback). xvii + 203 pp. LCCN QC6 .P694 1982.

Power:1996:ASN

[Pow96] Michael Power, editor. Accounting and science: natural inquiry and commercial reason, volume 26 of Cambridge studies in management. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55325-3 (hardcover), 0-521-55699-6 (paperback). xvii + 288 pp. LCCN HF5657 .A2565 1996. URL http://

www.loc.gov/catdir/
description/cam027/
95038403.html;   http://
www.loc.gov/catdir/toc/
cam022/95038403.html.

Pare:1982:MM

[PP82] Ambroise Paré and Janis L.

Pallister. On monsters and marvels. University of Chicago Press, Chicago, IL, USA, 1982. ISBN 0-226-64562-2. xxxii + 224 + 1 pp. LCCN QM691 .P3713 1982. US$16.50.

Pitt:1987:RCS

[PP87] Joseph C. Pitt and Marcello Pera, editors. Rational Changes in Science: Essays on Scientific Reasoning, volume 98 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1987. ISBN 90-277-2417-2, 94-009-3779-2. ISSN 0068-0346. xiii + 224 pp. LCCN Q174 .B67 vol. 98. US$64.00. URL https:

//link.springer.com/book/
10.1007/978-94-009-3779-6.

Panza:1995:SEV

[PP95] Marco Panza and Jean-Claude Pont, editors. Les savants et l’épistémologie vers la fin du XIXe siècle. (French) [Scientists and epistemology at the end of the 19th Century]. Bibliothèque scientifique Albert Blanchard. Libr. scientifique et technique Albert Blanchard, Paris, France, 1995. ISBN 2-85367-196-8. 282 pp. LCCN B2185 .S38 1995.

Perez-Ramos:1988:FBI

[PR88] Antonio Pérez-Ramos. Francis Bacon’s idea of science and the maker’s knowledge tradition. Clarendon Press, New York, NY, USA, 1988. ISBN 0-19-824979-9. xi + 334 pp. LCCN B1198 .P44 1988. US$32.50.

URL http://www.loc.gov/
catdir/enhancements/
fy0637/88009835-d.html.

Perez-Ramos:1990:FBA

[PR90] Antonio Pérez-Ramos. Francis Bacon and astronomical inquiry. British Journal for the History of Science, 23(2): 197–205, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4026725.

Pietsch:1995:FCC

[PR95] Theodore W. Pietsch and Louis Renard, editors. Fishes, crayfishes, and crabs: Louis Renard’s natural history of the rarest curiosities of the seas of the Indies. Foundations of natural history. The Johns Hopkins University Press, Baltimore, MD, USA, 1995. ISBN 0-8018-4790-7 (set : hardback). ???? pp. LCCN QL334.E27

F57 1995. URL http://www.
loc.gov/catdir/bios/
jhu051/93049411.html;
http://www.loc.gov/
catdir/description/
jhu051/93049411.html.

Price:1983:MRF

[Pri83] Roger Price. The modernization of rural France: communications networks and agricultural market structures in Nineteenth-Century France. [Hutchinson university library]. Hutchinson, London, UK, 1983. ISBN 0-09-145180-9. 503 pp. LCCN HD1945 .P737x 1983b.

Principe:1991:BRM

[Pri91] Lawrence M. Principe. Book review: Michela Pereira. The Alchemical Corpus Attributed to Raymond Lull. London: Warburg Institute, 1989. Pp. 114.

ISBN 0-85481-078-1. £8. British Journal for the History of Science, 24(2):247, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027172.

Principe:1993:BRA

[Pri93]  Lawrence  M.

Principe. Book review: Allen G. Debus, The French Paracelsians: The Chemical Challenge to Medical and Scientific Tradition in Early Modern France. Cambridge: Cambridge University Press, 1991. Pp. xvi + 247. ISBN 0-521-40049-X. £40.00, $59.95. British Journal for the History of Science, 26(1):82–83, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027346.

Provine:1986:SWE

[Pro86] William B. Provine. Sewall Wright and evolutionary biology. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1986. ISBN 0-226-68474-1. xvi + 545 pp. LCCN QH361 .P87 1986.

Proctor:1988:RHM

[Pro88] Robert Proctor. Racial hygiene: medicine under the Nazis. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-74580-9. viii + 414 pp. LCCN RA418.3.G3 P76 1988.

Probst:1993:ICO

[Pro93] Siegmund Probst. Infinity and creation: The origin of the controversy between Thomas Hobbes and the Savilian professors Seth Ward and John Wallis. British Journal for the History of Science, 26(3):271– 279, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027398.

Prigogine:1984:OCMa

[PS84a] I. (Ilya) Prigogine and Isabelle Stengers. Order out of chaos: man’s new dialogue with nature. Heinemann, London, UK, 1984. ISBN 0-434-60395-3. xxxi + 349 pp. LCCN Q175 PRI 1984.

Prigogine:1984:OCMb

[PS84b] I. (Ilya) Prigogine and Isabelle Stengers. Order out of chaos: man’s new dialogue with nature. Bantam Books, New York, NY, USA, 1984. ISBN 0-553-34082-4. xxxi + 349 pp. LCCN Q175.P8823; Q175 .P8823 1984.

Pera:1991:PSA

[PS91] Marcello Pera and William R.

Shea, editors. Persuading science: the art of scientific rhetoric. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1991. ISBN 0-88135-071-0. xi + 212 pp. LCCN Q175 .P384 1991. US$40.00.

Ptolemy:1984:PB

[PT84] Ptolemy and G. J. Toomer. Ptolemy’s Almagest. Duckworth classical, medieval, and renaissance editions. Duckworth, London, UK, 1984. ISBN 0-7156-1588-2. ix + 693 pp. LCCN QB41 .P957x 1984b. US$55.

Porter:1986:RH

[PT86] Roy Porter and Mikuláš Teich, editors. Revolution in history. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25978-9, 0-521-27784-1 (paperback). x + 341 pp. LCCN D32 .R44

1986. URL http://www.loc.
gov/catdir/description/
cam023/86006085.html;
http://www.loc.gov/
catdir/toc/cam028/
86006085.html.

Porter:1992:SRN

[PT92] Roy Porter and Mikuláš Teich, editors. The scientific revolution in national context. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-39510-0 (hardcover), 0-521-39699-9 (paperback). x + 305 pp. LCCN Q127.E8 S36

1992. URL http://www.loc.
gov/catdir/description/
cam024/91034348.html;
http://www.loc.gov/
catdir/toc/cam021/
91034348.html.

Paracelsus:1996:FTT

[PTS^{+}96]  Paracelsus,  C.  Lilian

(Clarice Lilian) Temkin, Henry E. (Henry Ernest) Sigerist, et al. Four treatises of Theophrastus von Hohenheim, called Paracelsus. The Johns Hopkins University Press, Baltimore, MD, USA, johns hopkins paperbacks edition, 1996. ISBN 0-8018-5523-3 (paperback). xii + 256 pp. LCCN R128.6 .P2213

1996. URL http://www.loc.
gov/catdir/bios/jhu051/
96030930.html;
http://www.loc.gov/
catdir/description/
jhu052/96030930.html.

Pugh:1984:MSI

[Pug84] Emerson W. Pugh. Memories that shaped an industry: decisions leading to IBM System/ 360. The MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-16094-3. x + 323 pp. LCCN QA76.8.I1015 P83 1984.

Pumfrey:1990:BRL

[Pum90] Stephen Pumfrey. Book review:

Lynn Sumida Joy. Gassendi the Atomist: Advocate of History in an Age of Science. Cambridge: University Press, 1988. Pp. xiii + 311. ISBN 0-521-30142-4. £27.50, $34.50. British Journal for the History of Science, 23(1): 99–102, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026807.

Pumfrey:1991:BRMb

[Pum91a] Stephen Pumfrey. Book review: Michael Hunter and Simon Schaffer (eds.). Robert Hooke. New Studies. Woodbridge, Suffolk: The Boydell Press, 1989. Pp. x + 310. ISBN 0-85115-523-5. £39.50. British Journal for the History of Science, 24(3): 382–384, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027244.

Pumfrey:1991:BRMa

[Pum91b] Stephen Pumfrey. Book review: Michael Hunter. Establishing the New Science: The Experience of the Early Royal Society. Woodbridge, Suffolk: The Boydell Press, 1989. Pp. xiv + 382. ISBN 0-85115-506-5. £45.00. British Journal for the History of Science, 24 (2):252–254, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027177.

Pumfrey:1991:HSN

[Pum91c] Stephen Pumfrey. History of science in the National Science Curriculum: a critical review of resources and their aims. British Journal for the History of Science, 24(1):61– 78, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027016.

Pumfrey:1992:BRR

[Pum92] Stephen Pumfrey. Book review:

Richard Olson. Science Deified and Science Defied: The Historic Significance of Science in Western Culture. Volume 2: From the Early Modern Age through the Early Romantic Era ca. 1640 to ca. 1820. Berkeley: University of California Press, 1991. Pp. xi + 445, illus. ISBN 0-520-06846-7. $45.00. British Journal for the History of Science, 25(4): 464–466, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027056.

Pumfrey:1994:BRD

[Pum94] Stephen Pumfrey. Book review: David C. Lindberg and Robert S. Westman (eds.), Reappraisals of the Scientific Revolution. Cambridge: Cambridge University Press, 1990. Pp. xxvii + 551. ISBN 0-521-34804-8. £14.95. British Journal for the History of Science, 27(2):226–228, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027441.

Pumfrey:1995:BRJ

[Pum95a] Stephen Pumfrey. Book review: Jean Dietz Moss, Novelties in the Heavens: Rhetoric and Science in the Copernican Controversy. Chicago: University of Chicago Press, 1993. Pp. xiv + 353. ISBN 0-226-54234-3, £39.95 (hardcover); 0-226-54235-1, £14.95 (paperback). British Journal for the History of Science, 28(1): 105–106, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027490.

Pumfrey:1995:WDW

[Pum95b] Stephen Pumfrey. Who did the work? Experimental philosophers and public demonstrators in Augustan England. British Journal for the History of Science, 28(2):131–156, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027675.

Pursell:1999:DME

[Pur99] Carroll Pursell. Domesticating modernity: the Electrical Association for Women, 1924– 1986. British Journal for the History of Science, 32(1):47– 67, March 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027969.

Pauli:1985:WBB

[PvM85] Wolfgang Pauli and Karl von Meyenn. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a. 2, 1930–1939. (German) [Scientific letters with Bohr, Einstein, Heisenberg and others, 2, 1930–1939], volume 6 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-13609-6 (Berlin), 0-387-13609-6 (New York). xxxix + 783 pp. LCCN ????

Pinsky:1989:CTC

[PW89]  Valerie  Pinsky  and

Alison Wylie, editors. Critical traditions in contemporary archaeology: essays in the philosophy, history, and socio-politics of archaeology. New directions in archaeology. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32109-3. ix + 160 pp. LCCN CC72 .C75 1989.

Pyenson:1985:YEA

[Pye85] Lewis Pyenson. The young Einstein: the advent of relativity. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-779-9. xiv + 255 + 9 pp. LCCN QC16.E5 P93

1985. URL http://www.loc.
gov/catdir/enhancements/
fy0745/85120280-d.html.

Pyenson:1993:CME

[Pye93] Lewis Pyenson. Civilizing mission: exact sciences and French overseas expansion, 1830–1940.

The Johns Hopkins University Press, Baltimore, MD, USA, 1993. ISBN 0-8018-4421-5. xxi + 377 pp. LCCN Q127.F8 P94

                                        1993. URL http://www.loc.
                                        gov/catdir/description/
                                        jhu051/92022577.html.

Pyle:1991:BRM

[Pyl91] Andrew Pyle. Book review: Mirko D. Grmek. La première révolution biologique: Reflexions sur la physiologic et la medecine du xvii siècle. Paris: Editions Payot, 1990. Pp. 358. ISBN 2-228-88277-1. British Journal for the History of Science, 24(2):255–257, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027179.

Pyle:1992:BRA

[Pyl92] Andrew Pyle. Book review: Annie Bitbol-Hespériès. Le Principe de vie chez Descartes. Paris: Libraire Philosophique J. Vrin, 1990. Pp. 236. ISBN 2-7116-1034-9. British Journal for the History of Science, 25 (2):265–266, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027308.

Pyle:1993:BRD

[Pyl93a] Andrew Pyle. Book review: David F. Channell, The Vital Machine: A Study of Technology and Organic Life. New York and Oxford: Oxford University Press, 1991. Pp. xi + 192. ISBN 0-19-506040-7. $22.95. British Journal for the History of Science, 26(1): 128–129, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027390.

Pyle:1993:BRW

[Pyl93b]  Andrew Pyle.  Book re-

view: William R. Shea, The Magic of Numbers and Motion: The Scientific Career of René Descartes. Canton, Mass.: Science History Publications, 1991. Pp. x + 371. ISBN 0-88135-098-2. $54.95. British Journal for the History of Science, 26(2):242–243, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027107.

Queisser:1988:CM

[Que88] Hans J. Queisser. The conquest of the microchip. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-16296-X. x + 200 + 12 pp. LCCN TK7874 .Q3613 1988.

Russell:1985:RDH

[R^{+}85] Colin Archibald Russell et al., editors. Recent developments in the history of chemistry. Royal Society of Chemistry, London, UK, 1985. ISBN 0-85186-917-3. x + 333 pp. LCCN QD15. US$27.50.

Reardon-Anderson:1991:SCC

[RA91] James Reardon-Anderson. The study of change: chemistry in China, 1840–1949. Studies of the East Asian Institute, Columbia University. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-39150-4 (hardcover). xix + 444 pp. LCCN QD18.C5 R43

1991. URL http://www.loc.
gov/catdir/description/
cam024/90034975.html;
http://www.loc.gov/
catdir/toc/cam029/
90034975.html.

Rafter:1988:WTE

[Raf88] Nicole Hahn Rafter, editor. White trash: the eugenic family studies, 1877–1919. Northeastern University Press, Boston, MA, USA, 1988. ISBN 1-55553-030-3. x + 382 pp. LCCN HQ755.5.U5 W52 1988. US$40.00.

Ragep:1995:BRG

[Rag95] F. Jamil Ragep. Book review: Gad Freudenthal (ed.), Studies on Gersonides, a Fourteenth-Century Philosopher-Scientist. Leiden: E. J. Brill, 1992. Pp. xxi + 422. ISBN 90-04-09641-8. $96.00, Gld. 168.00. British Journal for the History of Science, 28(1):103– 105, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027489.

Rainger:1991:AAH

[Rai91] Ronald Rainger. An agenda for antiquity: Henry Fairfield Osborn and vertebrate paleontology at the American Museum of Natural History, 1890– 1935. History of American science and technology series. University of Alabama Press, Tuscaloosa, AL, USA, 1991. ISBN 0-8173-0536-X. xiii + 360 pp. LCCN QE841 .R24 1991.

Ramsey:1988:PPM

[Ram88]  Matthew Ramsey.  Profes-

sional and popular medicine in France, 1770–1830: the social world of medical practice. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-30517-9. xvii + 406 pp. LCCN R504 .R36

1988. URL http://www.loc.
gov/catdir/description/
cam023/87015882.html;
http://www.loc.gov/
catdir/samples/cam031/
87015882.html;   http://
www.loc.gov/catdir/toc/
cam031/87015882.html.

Rappaport:1992:BRA

[Rap92] Rhoda Rappaport. Book review: Antonio Vallisneri. Epistolario. Volume I, 1679–1710, edited by Dario Generali. Milan: Franco Angeli, 1991. Pp.

639. ISBN 88-204-6733-X. L. 70,000. British Journal for the History of Science, 25(4): 466–467, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027057.

Rasmussen:1996:BRL

[Ras96a] Nicolas Rasmussen. Book review: Lynn K. Nyhart, Biology Takes Form: Animal Morphology and the German Universities. Chicago: University of Chicago Press, 1995. Pp. xiii + 414. ISBN 0-226-61086-1, £59.95, $75.00 (hardcover); 0-226-61088-8, £21.95, $27.50(paperback). British Journal for the History of Science, 29(3):373– 374, September 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027749.

Rasmussen:1996:BRM

[Ras96b] Nicolas Rasmussen. Book review: M. Norton Wise (ed.), The Values of Precision. Princeton: Princeton University Press, 1995. Pp viii + 372. ISBN 0-691-03759-0. £35.00, $49.50. British Journal for the History of Science, 29 (1):120–121, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027538.

Raven:1999:BRJ

[Rav99] Diederick Raven. Book review: J. V. Field, The Invention of Infinity: Mathematics and Art in the Renaissance. Oxford: Oxford University Press, 1997. Pp. 264. ISBN 0-19-852394-7. £29.50, $35.00. British Journal for the History of Science, 32(2):237–251, June 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4028091.

Ray:1987:ER

[Ray87] Christopher Ray. The evolution of Relativity. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-423-4. x + 211 pp. LCCN QC173.52 .R39

1987. URL http://www.loc.
gov/catdir/enhancements/
fy0745/87021137-d.html.

Rayner-Canham:1998:WCT

[RCRC98] Marelene F. Rayner-Canham and Geoffrey Rayner-Canham. Women in chemistry: their changing roles from alchemical times to the mid-Twentieth Century. History of modern chemical sciences, 1069– 2452. American Chemical Society, Washington, DC, 1998. ISBN 0-8412-3522-8. xiv + 284 pp. LCCN QD20 .R39 1998.

Redondi:1983:GE

[Red83] Pietro Redondi. Galileo eretico, volume 7 of Microstorie. G. Einaudi, Torino, Italy, 1983. ISBN 88-06-05632-8. x + 463 + 16 pp. LCCN Q125.2 .R43 1983. L25000.

Redner:1987:ESE

[Red87] Harry Redner. The ends of science: an essay in scientific authority. Westview Press, Boulder, CO, USA, 1987. ISBN 0-8133-0452-0. xiv + 344 pp. LCCN Q158.5 .R43 1987.

Rees:1990:BRA

[Ree90] Graham Rees. Book review: Antonio Pérez-Ramos. Francis Bacon’s Idea of Science and the Maker’s Knowledge Tradition. Oxford: Oxford University Press, 1988. Pp. xii + 334. ISBN 0-19-824979-9. £37.50. British Journal for the History of Science, 23(1): 106–107, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026811.

Rehbock:1983:PNT

[Reh83] Philip F. Rehbock. The philosophical naturalists: themes in early Nineteenth-Century British biology, volume 3 of Wisconsin publications in the history of science and medicine. University of Wisconsin Press, Madison, WI, USA, 1983. ISBN 0-299-09430-8. xv + 281 pp. LCCN QH305.2.G7 R43 1983. US$30.00.

Reich:1985:MAI

[Rei85] Leonard S. Reich. The making of American industrial research: science and business at GE and Bell, 1876–1926. Studies in economic history and policy. Cambridge University Press, Cambridge, UK, 1985.

ISBN 0-521-30529-2. xvi + 309 pp. LCCN HD30.42.U5 R44

1985. URL http://www.loc.
gov/catdir/description/
cam023/85006645.html;
http://www.loc.gov/
catdir/samples/cam031/
85006645.html;   http://
www.loc.gov/catdir/toc/
cam031/85006645.html.

Reisch:1994:PSO

[Rei94] George A. Reisch. Planning science: Otto Neurath and the International Encyclopedia of Unified Science. British Journal for the History of Science, 27(2):153–175, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027433.

ReyGonzalez:1990:EMS

[Rey90] Antonio M. Rey González.

Estudios médico-sociales sobre marginados en la Espaäna del siglo XIX, volume 17 of Textos clásicos espaänoles de la salud pública. Ministerio de Sanidad y Consumo, Madrid, Spain, 1990. ISBN 84-7670-204-3, 84-7670-205-1. 237 pp. LCCN RA418.3.S7 R49

1990. URL http://catalog.
hathitrust.org/api/
volumes/oclc/21796295.
html.

Richardson:1987:DDD

[Ric87]  Ruth Richardson.  Death,

dissection, and the destitute. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1987. ISBN 0-7102-0919-3. xvii + 426 pp. LCCN GT3243 .R53 1987.

Richards:1988:MVP

[Ric88] Joan L. Richards. Mathematical visions: the pursuit of geometry in Victorian England. Academic Press, New York, USA, 1988. ISBN 0-12-587445-6. xiii + 266 pp. LCCN QA443.5 .R53 1988.

Richards:1991:VCC

[Ric91] Evelleen Richards. Vitamin C and cancer: medicine or politics? MacMillan Publishing Company, New York, NY, USA, 1991. ISBN 0-333-44419-1. xii + 269 pp. LCCN RC271.A78 R53 1991b.

Richards:1992:MMO

[Ric92a] Graham Richards, editor. Mental machinery: the origins and consequences of psychological ideas: part 1: 1600–1850. Athlone Press, London, UK, 1992. ISBN 0-485-11394-5. xi + 490 pp. LCCN BF95 .R53 1992. US$50.00.

Richards:1992:MEM

[Ric92b]  Robert  J.

(Robert John) Richards. The meaning of evolution: the morphological construction and ideological reconstruction of Darwin’s theory. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-71202-8. xv + 205 pp. LCCN QH366.2 .R52

1992. URL http://www.loc.
gov/catdir/description/
uchi053/91019017.html;
http://www.loc.gov/
catdir/enhancements/
fy0666/91019017-b.html;
http://www.loc.gov/
catdir/enhancements/
fy0666/91019017-t.html.

Rice:1997:IDA

[Ric97] Adrian Rice. Inspiration or desperation? Augustus de Morgan’s appointment to the chair of mathematics at London University in 1828. British Journal for the History of Science, 30 (3):257–274, September 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027861.

Ritvo:1990:DIF

[Rit90] Lucille B. Ritvo. Darwin’s influence on Freud: a tale of two sciences. Yale University Press, New Haven, CT, USA, 1990. ISBN 0-300-04131-4. xii + 267 pp. LCCN BF173 .R515 1990.

Ramsey:1997:CBC

[RL97] J. T. (John T.) Ramsey and A. Lewis Licht. The comet of 44 B.C. and Caesar’s funeral games, volume 39 of American classical studies. Scholars Press, Atlanta, Ga., 1997. ISBN 0-7885-0273-5 (hardcover), 0-7885-0274-3 (paperback). xx + 236 pp. LCCN GT4852.R6 R36

1997. URL http://www.loc.

                                        gov/catdir/enhancements/
                                        fy0603/96017535-d.html.

Raaashid:1996:EHA

[RM96] Rushdåi Råashid and Régis Morelon, editors. Encyclopedia of the history of Arabic science. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1996. ISBN 0-415-02063-8 (set), 0-415-12410-7 (vol. 1), 0-415-12411-5 (vol. 2), 0-415-12412-3 (vol. 3). xiv + 1105 pp. LCCN Q127.A5 E53

                                        2000. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0648/00362029-d.html.

Roberts:1989:SJP

[Rob89] Shirley Roberts. Sir James Paget: the rise of clinical surgery. Eponymists in medicine. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1989. ISBN 0-905958-99-3 (hardcover), 0-905958-91-8. ISSN 0955-8640. 192 pp. LCCN RA489.R62 R62 1989. US$7.50. URL http://

                                        catalog.hathitrust.org/
                                        api/volumes/oclc/
                                        22547173.html.

Robertson:1992:BSS

[Rob92]  Peter Robertson.  Beyond

southern skies: radio astronomy and the Parkes telescope. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41408-3. xi + 357 pp. LCCN QB479.2 .R63

                                        1992. URL http://www.loc.

gov/catdir/description/
cam025/92003077.html;
http://www.loc.gov/
catdir/toc/cam025/
92003077.html.

Roberts:1993:BRD

[Rob93] Gerrylynn K. Roberts. Book review: David Knight, Ideas in Chemistry: A History of the Science. London: Athlone Press, 1992. Pp. vii + 213. ISBN 0-485-11390-2. £38.00. British Journal for the History of Science, 26(2):236–237, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

//www.jstor.org/stable/
4027101;   http://www.
jstor.org/stable/4027102.

Roberts:1996:CKI

[Rob96a] Gerrylynn K. Roberts. C. K.

Ingold at University College London: educator and department head. British Journal for the History of Science, 29 (1):65–82, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027512.

Roberts:1996:DLE

[Rob96b] Michael B. Roberts. Darwin at Llanymynech: the evolution of a geologist. British Journal for the History of Science, 29(4):469–478, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027700.

Rocke:1984:CAN

[Roc84] Alan J. Rocke. Chemical atomism in the Nineteenth Century: from Dalton to Cannizzaro. Ohio State University Press, Columbus, OH, USA, 1984. ISBN 0-8142-0360-4. xviii + 386 pp. LCCN QD15 .R63 1984. US$27.50.

Roche:1988:RLG

[Roc88] Daniel Roche. Les républicains des lettres: gens de culture et lumières au XVIIIe siècle. (French) [Republicans of letters: people of culture and light in the 18th Century]. Fayard, Paris, France, 1988. ISBN 2-213-02122-8. 393 pp. LCCN Z1003.5.F7 R63 1988. 160F.

Roche:1990:PLB

[Roc90] John J. (John James) Roche, editor. Physicists look back: studies in the history of physics. Adam Hilger Ltd., Bristol, UK, 1990. ISBN 0-85274-001-8. xii + 393 pp. LCCN QC7 .P475

                                        1990. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0705/91148368-d.html.

Rogers:1990:RU

[Rog90] G. A. J. Rogers. Review: [untitled]. British Journal for the History of Science, 23(4): 490–491, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027222.

Rohkramer:1999:BRAb

[Roh99a] Thomas Rohkrämer. Book review: Alexander von Humboldt, Cosmos: A Sketch of the Physical Description of the Universe, translated by Elise C. Otté, 2 vols. Baltimore: Johns Hopkins University Press, 1997. Pp. 375 and 367. ISBN 0-8018-5502-0, £13.00, $15.95 (vol. 1); 0-8018-5503-9, £13.00, $15.95 (vol. 2). British Journal for the History of Science, 32(3):363– 378, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4028001.

Rohkramer:1999:BRAa

[Roh99b] Thomas Rohkrämer. Book review: Anne Harrington, Reenchanted Science. Holism in German Culture from Wilhelm II to Hitler. Princeton: Princeton University Press, 1996. Pp. xiii + 309, illus. ISBN 0-691-02142-2. £13.99, $29.95. British Journal for the History of Science, 32(1):111– 124, March 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027979.

Ronan:1983:CIH

[Ron83] Colin A. Ronan. The Cambridge illustrated history of the world’s science. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-600-38423-3 (Newnes) 0-521-25844-8 (Cambridge). 543 pp. LCCN Q125 R645.

Roseveare:1982:MPV

[Ros82] N. T. Roseveare. Mercury’s perihelion, from Le Verrier to Einstein. Oxford science publications. Clarendon Press, New York, NY, USA, 1982. ISBN 0-19-858174-2. viii + 208 pp. LCCN QB611 .R67 1982. US$20.00.

Rosen:1984:CSR

[Ros84a] Edward Rosen. Copernicus and the scientific revolution. Anvil series. Krieger, Malabar, FL, USA, original edition, 1984. ISBN 0-89874-573-X (paperback). 220 pp. LCCN QB36.C8 R65 1984. US$5.95.

Rossi:1984:DAT

[Ros84b] Paolo Rossi. The dark abyss of time: the history of the earth and the history of nations from Hooke to Vico. University of Chicago Press, Chicago, IL, USA, 1984. ISBN 0-226-72835-8. xvi + 338 pp. LCCN D16.8 .R6813 1984. US$37.50.

Rosen:1986:TIM

[Ros86] Edward Rosen. Three imperial mathematicians: Kepler trapped between Tycho Brahe and Ursus. The New horizon series. Abaris Books, New York, NY, USA, 1986. ISBN 0-89835-242-8. 384 pp. LCCN QB35 .R67 1986.

Rossi:1990:MLS

[Ros90] Bruno Benedetto Rossi. Moments in the life of a scientist. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-36439-6. xiii + 181 + 32 pp. LCCN QC16.R495 R6713 1990. URL http://

www.loc.gov/catdir/
description/cam023/
89022283.html;   http://
www.loc.gov/catdir/toc/
cam028/89022283.html.

Ross:1991:NCA

[Ros91] Sydney Ross. Nineteenth-Century attitudes: men of science, volume 13 of Chemists and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1308-9 (hardcover). xi + 235 pp. LCCN Q174.8 .R67 1991.

Ross:1994:MIH

[Ros94a] Dorothy Ross, editor. Modernist impulses in the human sciences, 1870– 1930. The Johns Hopkins University Press, Baltimore, MD, USA, 1994. ISBN 0-8018-4744-3, 0-8018-4745-1 (paperback). xii + 379 pp. LCCN H53.U5

M63 1994. URL http://www.
loc.gov/catdir/bios/
jhu051/93038354.html;
http://www.loc.gov/
catdir/description/
jhu051/93038354.html.

Ross:1994:MTS

[Ros94b] Sydney Ross. De Morgan tussles with Smith’s harmonics in a comic poem. British Journal for the History of Science, 27(4):467–471, December 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027627.

Rothenberg:1982:HST

[Rot93a] Marc Rothenberg. The history of science and technology in the United States: a critical and selective bibliography, volume 2, 17; vol. 308, 815 of Bibliographies of the history of science and technology; Garland reference library of the humanities. Garland Publishing, Inc., New York, NY, USA, 1982– 1993. ISBN 0-8240-9278-3 (vol. 1), 0-8240-8349-0 (vol. 2). 2 v. pp. LCCN Z7405.H6 R67 1982. US$24.00.

Rothermel:1993:ISW

[Rot93b] Holly Rothermel. Images of the sun: Warren De la Rue, George Biddell Airy and celestial photography. British Journal for the History of Science, 26(2):137–169, June 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027095.

Rousseau:1997:BRV

[Rou97] G. S. Rousseau. Book reviews: Virtual realities Roslynn D. Haynes, From Faust to Strangelove: Representations of the Scientist in Western Literature. Baltimore: Johns Hopkins University Press, 1994. Pp. ix + 417. ISBN 0-8018-4801-6, £16.50. George Levine (ed.), Realism and Representation: Essays on the Problem of Realism in Relation to Science, Literature and Culture. Madison: University of Wisconsin Press, 1993. Pp. xiii + 330. ISBN 0-229-13630-2, £40.00 (hardcover); 0-229-13634-5, £19.00 (paperback). Sherry Turkle, Life on the Screen: Identity in the Age of the Internet. Cambridge, MA: Simon and Schuster, 1995. Pp.

347. ISBN 0-297-81514-8. No price given. British Journal for the History of Science, 30 (2):227–232, June 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027719.

Rowlands:1990:NCM

[Row90a] Peter Rowlands. Newton and the concept of mass-energy, volume 4 of Liverpool historical essays. Published for the Dept. of History, University of Liverpool [by] Liverpool University Press, Liverpool, UK, 1990. ISBN 0-85323-187-7. 34 pp. LCCN QC73 .R69 1990.

Rowlands:1990:OLL

[Row90b] Peter Rowlands. Oliver Lodge and the Liverpool Physical Society, volume 4 of Liverpool historical studies. Published for the Department of History, University of Liverpool [by] Liverpool University Press, Liverpool, UK, 1990. ISBN 0-85323-027-7 (paperback). iv + 310 pp. LCCN QC16.L6 R69 1990. US$12.50.

Reingold:1981:SAD

[RR81] Nathan Reingold and Ida H. Reingold, editors. Science in America, a documentary history, 1900–1939. The Chicago history of science and medicine. University of Chicago Press, Chicago, IL, USA, 1981. ISBN 0-226-70946-9. xii + 490 pp. LCCN Q125 .S43433. US$39.50

                                        (est.). URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0608/81002584-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0608/81002584-t.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0609/81002584-b.html.

Russell:1990:ASE

[RR90] Colin A. Russell and Shirley P. Russell. The archives of Sir Edward Frankland: Resources, problems and methods. British Journal for the History of Science, 23(2):175–185, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026723.

Rhodes:1991:DST

[RR91] Frank H. T. Rhodes and Frank H. T. Rhodes. Darwin’s search for a theory of the Earth: symmetry, simplicity and speculation. British Journal for the History of Science, 24(2): 193–229, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027166.

Roberts:1992:FBE

[RT92] K. B. Roberts and J. D. W.

Tomlinson. The fabric of the body: European traditions of anatomical illustrations. Clarendon Press, New York, NY, USA, 1992. ISBN 0-19-261198-4. xx + 638 pp. LCCN R836 .R63 1990. US$180.00. URL http://www.

loc.gov/catdir/
enhancements/fy0640/
90007650-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0640/90007650-t.html.

Riley:1995:CTC

[RT95]  William  Bell  Riley

and William Vance Trollinger, editors. Creationism in Twentieth-Century America. 4: the antievolution pamphlets of William Bell Riley. Garland Publishing, Inc., New York, NY, USA, 1995. ISBN 0-8153-1801-4, 0-8153-1805-7. xxii + 221 pp. LCCN BS651 .R546 1995.

Rudwick:1985:GDC

[Rud85] M. J. S. Rudwick. The great Devonian controversy: the shaping of scientific knowledge among gentlemanly specialists. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1985. ISBN 0-226-73101-4. xxxiii + 494 pp. LCCN QE665 .R83

1985. URL http://www.loc.
gov/catdir/description/
uchi052/84016199.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/84016199-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/84016199-b.html.

Ruderman:1988:KMS

[Rud88] David B. Ruderman. Kabbalah, magic, and science: the cultural universe of a sixteenth-century Jewish physician. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-49660-4. viii + 232 pp. LCCN BM755.J273 R83 1988.

Ruderman:1992:BRB

[Rud92a] David B. Ruderman. Book review: Barry B. Levy. Planets, Potions and Parchments: Scientific Hebraica from the Dead Sea Scrolls to the Eighteenth Century. Montreal and Kingston: McGill-Queen’s University Press, 1990. Pp. xii + 140, illus. ISBN 0-7735-0793-0, £47.45 (hardcover); 0-7735-0791-4, £28.45 (paperback). British Journal for the History of Science, 25(3):355– 357, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027258.

Rudwick:1992:SDT

[Rud92b] M. J. S. Rudwick. Scenes from deep time: early pictorial representations of the prehistoric world. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-73104-9 (hardcover). xiii + 280 pp. LCCN QE705.A1 R82

1992. URL http://www.loc.
gov/catdir/description/
uchi052/91047677.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/91047677-t.html.

Russell:1983:SS

[Rus83a] Colin A. (Colin Archibald)

Russell. Science and social, 1700–1900. MacMillan Publishing Company, New York, NY, USA, 1983. ISBN 0-333-29272-3. xvii + 307 pp. LCCN ????

Russell:1983:SSC

[Rus83b] Colin Archibald Russell. Science and social change: 1700– 1900. Themes in comparative history. MacMillan Publishing Company, New York, NY, USA, 1983. ISBN 0-333-29272-3 (hardcover), 0-333-29273-1 (paperback). xvii + 307 + 16 pp. LCCN Q175 .5. US$15.00.

Russell:1986:LCE

[Rus86a] Colin Archibald Russell. Lancastrian chemist: the early years of Sir Edward Frankland.

Open University Press, Milton Keynes, England, 1986. ISBN 0-335-15175-2. ix + 187 pp. LCCN QD22.F65 R87 1986.

Russell:1986:LEL

[Rus86b] Nicholas Russell. Like engend’ring like: heredity and animal breeding in early modern England. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30657-4. ix + 271 pp. LCCN SF53 .R87

                                        1986. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/85021350.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0665/85021350-t.html.

Ruse:1988:PBT

[Rus88]  Michael Ruse.  Philosophy

of biology today. SUNY series in philosophy and biology. State University of New York Press, Albany, NY, USA, 1988. ISBN 0-88706-910-X, 0-88706-911-8 (paperback). x + 155 pp. LCCN QH331 .R8793 1988.

Russett:1989:SSV

[Rus89] Cynthia Eagle Russett. Sexual science: the Victorian construction of womanhood. Harvard University Press, Cambridge, MA, USA, 1989. ISBN 0-674-80290-X. 245 pp. LCCN QP81.5 .R87 1989. US$25.00.

Russell:1991:BRH

[Rus91] Colin A. Russell. Book review: H. Benninga. A History of Lactic Acid Making: A Chapter in the History of Biotechnology. Dordrecht, Boston and London: Kluwer, 1990. Pp. xxi + 478. Illus. ISBN 0-7923-0625-2. £53.00, Dfl. 150.00. British Journal for the History of Science, 24(4):474– 475, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027135.

Russell:1992:BRL

[Rus92] Nicholas Russell. Book review: Liliane Bodson and Roland Libois (eds.). Contributions à l’histoire des connaissances zoologiques. Colloques d’histoire des connaissances zoologiques, 2. Liège: Université de Liège, 1991. Pp. 123. 450FB. British Journal for the History of Science, 25(3): 385–386, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027287.

Russell:1993:WEH

[Rus93] S. Russell. Writing energy history: explaining the neglect of CHP/DH in Britain. British Journal for the History of Science, 26(1):33–54, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027339.

Russell:1994:BRA

[Rus94] Colin A. Russell. Book review:

Anthony S. Travis, The Rainbow Makers: The Origins of the Synthetic Dyestuffs Industry in Western Europe. Bethlehem: Lehigh University Press/ London and Toronto: Associated University Presses, 1993. Pp. 335. ISBN 0-934223-18-1. £41.95. British Journal for the History of Science, 27(4): 481–482, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027634.

Rusnock:1999:CNR

[Rus99] Andrea Rusnock. Correspondence networks and the Royal Society, 1700–1750. British Journal for the History of Science, 32(2):155–169, June 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4028081.

Roberts:1990:JDL

[RWD90] R. J. (Richard Julian) Roberts, Andrew G. Watson, and John Dee, editors. John Dee’s library catalogue. Bibliographical Society, London, UK, 1990. ISBN 0-19-721795-8. viii + 253 + 170 pp. LCCN Z997.D298 J64

                                        1990. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0637/90228267-d.html.

Salthe:1985:EHS

[Sal85] Stanley N. Salthe. Evolving hierarchical systems: their structure and representation. Columbia University Press, New York, NY, USA, 1985. ISBN 0-231-06016-5. x + 343 pp. LCCN QH83 .S26 1985.

Sangwan:1991:STC

[San91] Satpal Sangwan. Science, technology and colonisation: an Indian experience, 1757–1857. Anamika Prakashan, Delhi, India, 1991. ISBN 81-85150-11-7. xviii + 196 pp. LCCN MQ.

Sapp:1990:WTL

[Sap90] Jan Sapp. Where the truth lies: Franz Moewus and the origins of molecular biology. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-36550-3, 0-521-36751-4 (paperback). ix + 340 pp. LCCN QH506 .S27 1990. URL http:

//www.loc.gov/catdir/
description/cam023/
89032582.html;   http://
www.loc.gov/catdir/toc/
cam026/89032582.html.

Sarjeant:1987:GHG

[Sar87] William Antony S. Sarjeant.

Geologists and the history of geology: an international bibliography from the origins to 1978. Supplement, 1979–1984 and additions. R.E. Krieger Pub. Co., Malabar, Fla., U.S.A., 1987. ISBN 0-89874-939-5. 2 v. pp. LCCN Z6031 .S28 Suppl.

Sarkar:1992:FEG

[Sar92] Sahotra Sarkar, editor. The founders of evolutionary genetics: a centenary reappraisal, volume 142 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346. 300 pp. LCCN Q174 .B67 vol. 142. URL http:

//www.loc.gov/catdir/
enhancements/fy0823/
92012823-d.html;   http:/
/www.loc.gov/catdir/
enhancements/fy0823/
92012823-t.html;   https:
//link.springer.com/book/
10.1007/978-94-011-2856-8.

Smith:1992:DRA

[SB92] Murphy D. Smith and David Borodin. Due reverence: antiques in the possession of the American Philosophical Society, volume 203 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1992. ISBN 0-87169-184-1. vi + 77 pp. LCCN Q11 .P612 vol. 203.

Shea:2009:GBN

[SB09] William R. Shea and Tiziana Bascelli. Galileo’s Sidereus Nuncius, or, a sidereal message. Science History Publications, Sagamore Beach, MA, USA, 2009. ISBN 0-88135-375-2. viii + 124 pp. LCCN QB41 .G173 2009.

Schmitt:1983:AR

[Sch83] Charles B. Schmitt. Aristotle and the Renaissance, volume 27 of Martin classical lectures. Published for Oberlin College by Harvard University Press, Cambridge, MA, USA, 1983. ISBN 0-674-04525-4. viii + 187 pp. LCCN PA25 .M3 vol. 27.

Schutt:1984:EIF

[Sch84] Hans-Werner Schütt. Die Entdeckung des Isomorphismus: eine Fallstudie zur Geschichte der Mineralogie und der Chemie. (German) [The discovery of isomorphicism: a case study in the history of mineralogy and chemistry], volume 9 of Arbor scientiarum. Reihe A, Abhandlungen. Gerstenberg Verlag, Hildesheim, Germany, 1984. ISBN 3-8067-0902-5. 324 pp. LCCN QD951 .S48 1984.

Schivelbusch:1988:DNI

[Sch88] Wolfgang Schivelbusch. Disenchanted night: the industrialisation of light in the Nineteenth Century. Berg, Oxford, UK, 1988. ISBN 0-85496-191-7. 227 pp. LCCN GT440 .S3613x 1988b.

Scholz:1989:SGD

[Sch89] Erhard Scholz, editor. Symmetrie, Gruppe, Dualität: zur Beziehung zwischen theoretischer Mathematik und Anwendungen in Kristallographie und Baustatik des 19. Jahrhunderts. (German) [Symmetry, group, duality: on the relationship between theoretical mathematics and applications in crystallography and structural design of the 19th Century]. Science networks historical studies. Birkhäuser Verlag, Basel, Switzerland, 1989. ISBN 3-7643-1974-7. ISSN 1421-6329. 406 pp. LCCN QD911 .S324 1989b.

Schabas:1990:WRN

[Sch90a]  Margaret  Sch-

abas. A world ruled by number: William Stanley Jevons and the rise of mathematical economics. Princeton University Press, Princeton, NJ, USA, 1990. ISBN 0-691-08543-9. xii + 192 pp. LCCN HB103.J5 S33 1990. US$29.95. URL http://

                                        www.loc.gov/catdir/
                                        description/prin031/
                                        89070245.html.

Schaffer:1990:RU

[Sch90b] Simon Schaffer. Review: [untitled]. British Journal for the History of Science, 23(4): 465–467, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027204.

Schabas:1991:BRT

[Sch91a]  Margaret  Sch-

abas. Book reviews: Thomas Robert Malthus. An Essay on the Principle of Population, edited by Patricia James, 2 vols. Cambridge: Cambridge University Press for the Royal Economic Society, 1990. Pp. xvi + 446; vi + 397. ISBN 0-521-32361-4. L.75, $120.00.

Thomas Robert Malthus. Principles of Political Economy, variorum edition, edited by John Pullen, 2 vols. Cambridge: Cambridge University Press for the Royal Economic Society, 1989. Pp. lxix + vi + 601; 507. ISBN 0-521-24775-6. L. 75.00, $120.00. British Journal for the History of Science, 24(4): 482–484, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027143.

Schupbach:1991:BRG

[Sch91b] William Schupbach. Book review: Gavin D. R. Bridson and James J. White. Plant, Animal and Anatomical Illustration in Art and Science: A Bibliographical Guide from the 16th Century to the Present Day. Winchester: St Paul’s Bibliographies in association with Hunt Institute for Botanical Documentation, 1990. Pp. xxxix + 450. ISBN 0-906795-81-8. £75.00. British Journal for the History of Science, 24(4):488–489, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027148.

Schaaf:1992:SHT

[Sch92a] Larry J. (Larry John) Schaaf.

Out of the shadows: Herschel, Talbot and the invention of photography. Yale University Press, New Haven, CT, USA, 1992. ISBN 0-300-05705-9 (hardcover). xii + 188 pp. LCCN TR57 .S32 1992. US$65.00.

Schutt:1992:EMB

[Sch92b]  Hans-Werner Schütt.  Eil-

hard Mitscherlich, Baumeister am Fundament der Chemie. (German) [Eilhard Mitscherlich, master builder of the basis of chemistry], volume n.F., Bd. 8 of Abhandlungen und Berichte. In Kommission bei R. Oldenbourg, München, Germany, 1992. ISBN 3-486-26273-4. 192 pp. LCCN AM101 .M9743 Bd. 8.

Schiemenz:1993:HLB

[Sch93a]  Günter P. Schiemenz.  A

heretical look at the Benzolfest. British Journal for the History of Science, 26(2): 195–205, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027097.

Schneider:1993:JFR

[Sch93b] Ivo Schneider. Johannes Faulhaber: 1580–1635; Rechenmeister in einer Welt des Umbruchs. (German) [Johannes Faulhaber: 1580–1635: Computing master in a world of radical change], volume 7 of Vita Mathematica. Birkhäuser Verlag, Basel, Switzerland, 1993. ISBN 3-7643-2919-X. xiv + 271 pp. LCCN QA29.F38 S35

                                        1993. URL http://catalog.
                                        hathitrust.org/api/

volumes/oclc/29955228.
html.

Schubring:1993:BRY

[Sch93c] Gert Schubring. Book review:

Yves Cohen and Klaus Manfrass (eds.), Frankreich und Deutschland: Forschung, Technologie und industrielle Entwicklung im 19. und 20. Jahrhundert. Munich: C. H. Beck’schen Veragsbuchhandlung, 1990. Pp. xl + 491. ISBN 3-406-35092-5. DM. 88. British Journal for the History of Science, 26 (1):119–121, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027382.

Schaffer:1995:SNE

[Sch95a] Simon Schaffer. The show that never ends: perpetual motion in the early eighteenth century. British Journal for the History of Science, 28(2): 157–189, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027676.

Schuhmann:1995:BRB

[Sch95b] Karl Schuhmann. Book review:

Beverley C. Southgate, ‘Covetous of Truth’: The Life and Work of Thomas White, 1593– 1676. International Archives of the History of Ideas, 134. Dordrecht: Kluwer Academic Publishers, 1993. Pp. xi + 189. ISBN 0-7923-1926-5. £60.00.

British Journal for the History of Science, 28(1):108– 109, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027492.

Schosler:2001:DLF

[Sch01] Jörn Schösler, editor. La diffusion de Locke en France: traduction au dix-huitiéme siècle: lectures de Rousseau. (French) [The diffusion of Locke in France: 18th Century translation of the readings of Rousseau], volume 2001:04 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2001. ISBN 0-7294-0748-9. v + 364 pp. LCCN MLCM 2005/02528 (B).

Stachel:1987:CPA

[SCS^{+}87] John J. Stachel, David C. Cassidy, Robert Schulmann, Jürgen Renn, Olga Griminger, Gary Smith, and Robert Summerfield, editors. The collected papers of Albert Einstein. Vol. 1. The early years, 1879–1902. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-08407-6. lxvi + 433 pp. LCCN QC16.E5 A2 1987. US$52.50. URL ftp:/

/uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/
catdir/description/
prin021/86043132.html;
http://www.loc.gov/
catdir/toc/prin032/

86043132.html.  Published

with a translation into English by Anna Beck, ISBN 0-691-08475-0 [BH87], which is not available separately. Consultant for the translation: Peter Havas.

Snow:1991:NIV

[SE91] John Snow and Richard H. Ellis. On narcotism by the inhalation of vapours. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1991. ISBN 1-85315-158-0. xxx + 112 pp. LCCN RD81 .S663 1991.

Secord:1986:CVG

[Sec86] James A. Secord. Controversy in Victorian geology: the Cambrian–Silurian dispute. Princeton University Press, Princeton, NJ, USA, 1986. ISBN 0-691-08417-3. xvii + 363 pp. LCCN QE656 .S4 1986. US$49.50.

Secord:1991:DVD

[Sec91] James A. Secord. The discovery of a vocation: Darwin’s early geology. British Journal for the History of Science, 24 (2):133–157, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027164.

Secord:1993:I

[Sec93] James A. Secord. Introduction. British Journal for the History of Science, 26(4):387– 389, December 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027462.

Secord:1994:CIA

[Sec94] Anne Secord. Corresponding interests: artisans and gentlemen in nineteenth-century natural history. British Journal for the History of Science, 27(4): 383–408, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027623.

Secord:1996:BRM

[Sec96] Anne Secord. Book review: Michael Shortland (ed.), Hugh Miller’s Memoir: From Stonemason to Geologist. Edinburgh: Edinburgh University Press, 1995. Pp. v + 266. ISBN 0-7486-0521-5. £12.95. British Journal for the History of Science, 29(1): 105–106, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027527.

Secord:1998:BRM

[Sec98] Jim Secord. Book review: Michael Shortland (ed.), Hugh Miller and the Controversies of Victorian Science. Oxford: Oxford University Press, 1996, Pp. x + 401. ISBN 0-19-854053-1. £49.50. British Journal for the History of Science, 31(2):241– 250, June 1998. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Segre:1991:WG

[Seg91] Michael Segre. In the wake of Galileo. Rutgers University Press, New Brunswick, NJ, USA, 1991. ISBN 0-8135-1700-1, 0-8135-1701-X (paperback). xix + 192 pp. LCCN Q127.I8 S44 1991.

Segal:1994:FIM

[Seg94] Howard P. Segal. Future imperfect: the mixed blessings of technology in America. University of Massachusetts Press, Amherst, MA, USA, 1994. ISBN 0-87023-881-7 (hardcover), 0-87023-882-5 (paperback). xviii + 245 pp. LCCN T21 .S43 1994. US$40.00, US$15.95.

Sellers:2010:CCV

[Sel10] Susan Sellers, editor. The Cambridge companion to Virginia Woolf. Cambridge companions to literature. Cambridge University Press, Cambridge, UK, second edition, 2010. ISBN 0-521-89694-0 (hardcover), 0-521-72167-9 (paperback). xxi + 272 pp. LCCN PR6045.O72 Z5655 2010. URL http://

assets.cambridge.org/
97805218/96948/cover/
9780521896948.jpg.

Sepper:1988:GCN

[Sep88] Dennis L. Sepper. Goethe contra Newton: polemics and the project for a new science of color. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-34254-6. xvi + 222 pp. LCCN QC495 .S43

1988. URL http://www.loc.
gov/catdir/description/
cam031/87021463.html;
http://www.loc.gov/
catdir/toc/cam031/
87021463.html.

Sadoun-Goupil:1980:RBS

[SG80] Michelle Sadoun-Goupil. Revue de L’essai de statique chimique, Claude-Louis Berthollet: édition critique d’un projet de seconde édition de L’essai de statique chimique. (French) [Review of the Test of chemical statistics, Claude-Louis Berthollet: the critical edition of a project of the second edition of Test of chemical statistics]. L’École Polytechnique, Paris, France, 1980. ISBN 2-7302-0019-3. vii + 204 pp. LCCN QD28

.B533. URL http://catalog.
hathitrust.org/api/
volumes/oclc/9259834.html.

Schofield:1990:BHS

[SG90] Robert E. Schofield and Elizabeth Garber, editors. Beyond history of science: essays in honor of Robert E. Schofield. Lehigh University Press, Bethlehem, PA, USA, 1990. ISBN 0-934223-11-4. 325 pp. LCCN Q126.9 .B49 1990.

Sarasohn:1996:GEF

[SG96] Lisa T. Sarasohn and Pierre Gassendi. Gassendi’s ethics: freedom in a mechanistic universe. Cornell University Press, Ithaca, NY, USA, 1996. ISBN 0-8014-2947-1 (hardcover). xii + 236 pp. LCCN B1887 .S27 1996.

Silver:2000:FTL

[SGS00]  Marie-France  Alberte

Silver and Marie-Laure Girou-Świderski, editors. Femmes en toutes lettres: les épistolières du XVIIIe siècle. (French) [Women of all letters: the epistolieres of the 18th Century], volume 2000:04 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2000. ISBN 0-7294-0741-1. xiii + 277 pp. LCCN PQ711 .F45 2000.

Shapiro:1983:PCS

[Sha83] Barbara J. Shapiro. Probability and certainty in seventeenthcentury England: a study of the relationships between natural science, religion, history, law, and literature. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-05379-0, 0-691-10146-9 (lim. paperback ed.). x + 347 pp. LCCN DA380 .S53 1983. US$32.50, US$12.50.

Shaw:1984:WPS

[Sha84] John Shaw. Water power in Scotland, 1550–1870. John Donald, Edinburgh, Scotland, 1984. ISBN 0-85976-072-3. xi + 606 pp. LCCN TJ859 .S53; TC147. US$25.00.

URL Biblioth\`;   http://

                                        catalogue.bnf.fr/ark:/
                                        12148/cb34808984k.

Shapin:1992:BRP

[Sha92]  Steven Shapin.  Book re-

view: Peter Dear (ed.). The Literary Structure of Scientific Argument: Historical Studies. Philadelphia: University of Pennsylvania Press, 1991. ISBN 0-8122-8185-3. No price given. British Journal for the History of Science, 25(3):388– 389, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027290.

Shapin:1993:BRJ

[Sha93a] Steven Shapin. Book review: Julian Martin, Francis Bacon, the State, and the Reform of Natural Philosophy. Cambridge: Cambridge University Press, 1992. Pp. xiii + 236. ISBN 0-521-38249-1. £35,00, $49.95. British Journal for the History of Science, 26(1): 84–85, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027349.

Shapin:1993:PDI

[Sha93b] Steven Shapin. Personal development and intellectual biography: the case of Robert Boyle. British Journal for the History of Science, 26(3):335– 345, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027401.

Shapiro:1993:FPP

[Sha93c] Alan E. (Alan Elihu) Shapiro.

Fits, passions, and paroxysms: physics, method, and chemistry and Newton’s theories of colored bodies and fits of easy reflection. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-40507-6. xvii + 400 pp. LCCN QC402 .S53

1993. URL http://www.loc.
gov/catdir/description/
cam031/92014762.html;
http://www.loc.gov/
catdir/enhancements/
fy0907/92014762-t.html.

Shapin:1996:BRA

[Sha96] Steven Shapin. Book review:

Anne Goldgar, Impolite Learning: Conduct and Community in the Republic of Letters, 1680–1750. New Haven, CT, and London: Yale University Press, 1995. Pp. xv + 395. ISBN 0-300-05359-2. £25. British Journal for the History of Science, 29(1):97– 98, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027520.

Shank:1997:BRA

[Sha97]  Michael H. Shank.  Book

reviews: Annibale Fantoli, Galileo: For Copernicanism and for the Church, translated by George V. Coyne, SJ. Studi Galileiani, 3. Rome: Vatican Observatory Publications, 1994 (distributed outside Italy by University of Notre Dame Press, Notre Dame, IN). First edition: pp. xix + 540. ISBN 0-268-01029-3. Second edition, revised and corrected, 1996, pp. xx + 567. ISBN 0-268-01032-3. $21.95. Rivka Feldhay, Galileo and the Church: Political Inquisition or Critical Dialogue? Cambridge: Cambridge University Press, 1995. Pp. viii + 303. ISBN 0-521-34468-8. £35.00, $54.95. British Journal for the History of Science, 30 (1):101–121, March 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027912.

Shamdasani:1998:CFC

[Sha98] Sonu Shamdasani. Cult fictions: C. G. Jung and the founding of analytical psychology. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1998. ISBN 0-415-18613-7 (hardcover), 0-415-18614-5 (paperback). x + 121 pp. LCCN BF173 .S486

1998. URL http://www.loc.
gov/catdir/enhancements/
fy0649/97032830-d.html.

Shea:1991:MNM

[She91] William R. Shea. The magic of numbers and motion: the scientific career of René Descartes. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 1991. ISBN 0-88135-098-2. x + 371 pp. LCCN Q125.2 .S54 1990.

Shennan:1992:BRV

[She92] Stephen J. Shennan. Book review: Valerie Pinsky and Alison Wylie (eds.). Critical Traditions in Contemporary Archaeology: Essays in the Philosophy, History and Socio-Politics of Archaeology. Cambridge: Cambridge University Press, 1990. Pp. ix + 160. ISBN 0-521-32109-3. £27.50, $54.50. British Journal for the History of Science, 25(3):384– 385, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027286.

Sheehan:1995:IFW

[She95] William Sheehan. The immortal fire within: the life and work of Edward Emerson Barnard. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-44489-6 (hardcover). xiv + 429 pp. LCCN QB36.B3 S48

1995. URL http://www.loc.
gov/catdir/description/
cam026/94026742.html;
http://www.loc.gov/
catdir/toc/cam021/
94026742.html.

Shea:1999:BRJ

[She99]  William Shea.  Book re-

view: John Brooke and Geoffrey Cantor, Reconstructing Nature. The Engagement of Science and Religion. Edinburgh: T. & T. Clark, 1998. Pp. xii + 367. ISBN 0-567-08600-3. £24.95, $49.95. British Journal for the History of Science, 32(2):237–251, June 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4028092.

Shirley:1983:THB

[Shi83] John William Shirley. Thomas Harriot, a biography. Clarendon Press, New York, NY, USA, 1983. ISBN 0-19-822901-1. xii + 508 pp. LCCN Q143.H36 S46 1983. US$59.00.

Shortt:1986:VLR

[Sho86] S. E. D. (Samuel Edward Dole) Shortt. Victorian lunacy: Richard M. Bucke and the practice of late Nineteenth-Century psychiatry. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30999-9. xvi + 207 pp. LCCN RC448.O53 S56

                                        1986. URL http://www.loc.
                                        gov/catdir/description/
                                        cam032/85026919.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy1009/85026919-t.html.

Shortland:1993:BRB

[Sho93a] Michael Shortland. Book review: Bernd Naumann, Frans Plank and Gottfried Hofbauer (eds.), Language and Earth:

Elective Affinities Between the Emerging Sciences of Linguistics and Geology. Amsterdam and Philadelphia: John Benjamin, 1992. Pp. xvi + 445. ISBN 90-272-4553-3. Dfl. 200.00. British Journal for the History of Science, 26(1):100, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027365.

Shortland:1993:BRH

[Sho93b] Michael Shortland. Book review: Harold Dorn. The Geography of Science. Baltimore and London: Johns Hopkins University Press, 1991. Pp. xx + 219. ISBN 0-8018-4151-8. No price given. British Journal for the History of Science, 26 (1):130–132, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027392.

Shortland:1993:BRJ

[Sho93c]  Michael Shortland.  Book

reviews: J. Vernon Jensen, Thomas Henry Huxley: Communicating for Science. London and Toronto: Associated University Press, 1991. Pp. 253. ISBN 0-87413-379-3. No price given. Michael Collie, Huxley at Work, with the Scientific Correspondence of T. H. Huxley and the Rev. Dr George Gordon of Birnie, near Elgin. London: Macmillan, 1991. Pp. xii +158. ISBN 0-333-51059-3. No price given. British Journal for the History of Science, 26 (1):112–114, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027376.

Shortland:1993:SN

[Sho93d] Michael Shortland, editor. Science and nature. British Society for the History of Science, Chalfont St. Giles, Bucks, England, 1993. ISBN 0-906450-08-X (paperback). viii + 291 pp. LCCN ????

Shortland:1993:SNE

[Sho93e] Michael Shortland, editor. Science and nature: essays in the history of the environmental sciences, volume 8 of BSHS monographs. British Society for the History of Science, Chalfont St. Giles, Bucks, England, 1993. ISBN 0-906450-08-X. viii + 291 pp. LCCN GF13 .S33

                                        1993. URL http://catalog.
                                        hathitrust.org/api/
                                        volumes/oclc/28748717.
                                        html.

Shortland:1995:RU

[Sho95] Michael Shortland. Review: [untitled]. British Journal for the History of Science, 28(4): 477–480, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027562.

Shortland:1996:BRJ

[Sho96a] Michael Shortland. Book review: Jonathan Smith, Fact and Feeling: Baconian Science and the Nineteenth-Century Literary Imagination. Madison: University of Wisconsin Press, 1995. Pp. 277. ISBN 0-299-14350-3, £46.95 (hardcover); 0-299-14354-6, £20.50 (paperback). British Journal for the History of Science, 29(1): 101–102, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027523.

Shortland:1996:HMC

[Sho96b] Michael Shortland, editor. Hugh Miller and the controversies of Victorian science. Clarendon Press, New York, NY, USA, 1996. ISBN 0-19-854053-1. viii + 401 pp. LCCN Q143.M645 A3

1996. URL http://www.loc.
gov/catdir/enhancements/
fy0637/95041168-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0637/95041168-t.html.

Shortland:1998:MSP

[Sho98] Michael Shortland. Michel Serres, passe-partout. British Journal for the History of Science, 31 (3):335–353, September 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027816.

Shotter:1999:BRJ

[Sho99]  David Shotter.  Book re-

views: John T. Ramsey and A. Lewis Licht, The Comet of 44 BC and Caesar’s Funeral Games, American Philological Association: American Classical Studies, 39. Atlanta: The Scholars Press, 1997. Pp. xx + 236. ISBN 0-7885-0273-5. $27.95 (hardcover); ISBN 0-7885-0274-3. $17.95 (paperback). British Journal for the History of Science, 32(2): 237–251, June 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4028093.

Shrum:1992:BRR

[Shr92] Wesley Shrum. Book review: Robert Lambourne, Michael Shallis and Michael Shortland. Close Encounters? Science and Science Fiction. Bristol and New York: Adam Hilger, 1990. Pp. xiii + 184. ISBN 0-85274-141-3. L. 12.95. British Journal for the History of Science, 25(2):294, June 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027331.

Shteir:1996:CWC

[Sht96]  Ann B. Shteir.  Cultivat-

ing women, cultivating science: Flora’s daughters and botany in England, 1760–1860. The Johns Hopkins University Press, Baltimore, MD, USA, 1996. ISBN 0-8018-5141-6 (hardcover). xi + 301 pp. LCCN QK21.G7 S57

1996. URL http://www.loc.
gov/catdir/bios/jhu051/
95012736.html;
http://www.loc.gov/
catdir/description/
jhu051/95012736.html.

Shuttleworth:1996:CBV

[Shu96] Sally Shuttleworth. Charlotte Brontë and Victorian psychology, volume 7 of Cambridge studies in nineteenth-century literature and culture. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55149-8 (hardcover). xiv + 289 pp. LCCN PR4169 .S48

1996. URL http://www.loc.
gov/catdir/description/
cam026/95004681.html;
http://www.loc.gov/
catdir/toc/cam023/
95004681.html.

Shuttleworth:2004:CBV

[Shu04] Sally Shuttleworth. Charlotte Brontë and Victorian psychology, volume 7 of Cambridge studies in nineteenth-century literature and culture. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-61717-0 (paperback), 0-521-55149-8. xiv + 289 pp. LCCN PR4169 .S48

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0632/2005278193-d.html;
 http://www.loc.gov/
catdir/enhancements/
fy0632/2005278193-t.html.

Siegel:1991:IME

[Sie91]  Daniel M. Siegel.  Inno-

vation in Maxwell’s electromagnetic theory: molecular vortices, displacement current, and light. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-35365-3. x + 225 pp. LCCN QC670 .S48

                                        1991. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/90042511.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam029/
                                        90042511.html.

Silverman:1990:BRC

[Sil90] Sam Silverman. Book review: C. Stewart Gillmor (ed.). History of Geophysics. Volume 2. Pp. vi + 191. Washington, D.C.: American Geophysical Union, 1986. British Journal for the History of Science, 23 (2):229–230, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026736.

Simonton:1988:SGP

[Sim88] Dean Keith Simonton. Scientific genius: a psychology of science. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-35287-8. x + 229 pp. LCCN Q172.5.C74 S57

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87027823.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam031/
                                        87027823.html.

Simon:1992:BRR

[Sim92] Bart Simon. Book review:

Robin Herman. Fusion: The Search For Endless Energy. Cambridge: Cambridge University Press, 1990. Pp. xi + 267. ISBN 0-521-38373-0. £14.95. British Journal for the History of Science, 25(3): 377–378, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027278.

Simon:1994:BRJ

[Sim94] Bart Simon. Book review: John R. Huizenga, Cold Fusion: The Scientific Fiasco of the Century. Rochester, New York: University of Rochester Press, 1992. Pp. xv + 259. ISBN 1-878822-07-1. £29.50. British Journal for the History of Science, 27 (3):378–379, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027617.

Sinaceur:1991:CME

[Sin91] Hourya Sinaceur. Corps et modèles: essai sur l’histoire de l’algèbre réelle. (French) [Bodies and models: essay on the history of real algebra]. Mathesis. J. Vrin, Paris, France, 1991. ISBN 2-7116-1038-1. 496 pp. LCCN QA155 .S56 1991. 198F.

Singer:1997:RHS

[Sin97] Charles Singer. The role of the history of science. British Journal for the History of Science, 30(1):71–73, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027901.

Siraisi:1987:ARI

[Sir87] Nancy G. Siraisi. Avicenna in Renaissance Italy: the Canon and medical teaching in Italian universities after 1500. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-05137-2. xii + 410 + 8 pp. LCCN R791.A6 S57 1987. US$65.00.

Siraisi:1990:MER

[Sir90] Nancy G. Siraisi. Medieval and early Renaissance medicine: an introduction to knowledge and practice. University of Chicago Press, Chicago, IL, USA, 1990. ISBN 0-226-76129-0, 0-226-76130-4 (paperback). xiv + 250 pp. LCCN R141 .S546

1990. URL http://www.loc.
gov/catdir/description/
uchi052/89020368.html;
http://www.loc.gov/
catdir/toc/uchi051/
89020368.html.

Siraisi:1997:CMG

[Sir97] Nancy G. Siraisi. The clock and the mirror: Girolamo Cardano and Renaissance medicine. Princeton University Press, Princeton, NJ, USA, 1997. ISBN 0-691-01189-3 (hardcover). xiv + 361 pp. LCCN R520.C32 S57

1997. URL http://www.loc.
gov/catdir/description/
prin021/96046354.html;
http://www.loc.gov/
catdir/toc/prin031/
96046354.html.

Six:1987:DDN

[Six87] Jules Six. La découverte du neutron. (French) [The discovery of the neutron]. Editions du Centre national de la recherche scientifique, Paris, France, 1987. ISBN 2-222-04052-3. vii + 174 pp. LCCN MLCM 90/06183 (Q). 110F.

Simpson:1987:SCL

[SL87] George Gaylord Simpson and Leo F. Laporte. Simple curiosity; letters from George Gaylord Simpson to his family, 1921– 1970. University of California Press, Berkeley, CA, USA, 1987. ISBN 0-520-05792-9. x + 340 pp. LCCN QE707.S55 A4 1987.

Slaughter:1982:ULS

[Sla82] M. M. (Mary M.) Slaughter.

Universal languages and scientific taxonomy in the Seventeenth Century. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24477-3. x + 277 pp. LCCN PM8009 .S77

1982. URL http://www.loc.
gov/catdir/enhancements/
fy1002/81020147-d.html;
http://www.loc.gov/
catdir/enhancements/
fy1002/81020147-t.html.

Sloane:1985:DME

[Slo85] Thomas O. Sloane. Donne, Milton, and the end of humanist rhetoric. University of California Press, Berkeley, CA, USA, 1985. ISBN 0-520-05212-9, 0-520-05213-7. xv + 332 pp. LCCN PR545.H86 S58 1985.

Sloan:1997:BRJ

[Slo97] Phillip R. Sloan. Book review: N. Jardine, J. A. Secord and E. C. Spary (eds.), Cultures of Natural History. Cambridge: Cambridge University Press, 1995. Pp. xxi + 501. ISBN 0-521-45394-1, £65.00, $89.95 (hardcover); 0-521-55894-8, £22.90, $29.95 (paperback). British Journal for the History of Science, 30 (2):233–249, June 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027728.

Sanchez:1988:NKQ

[SLT88] Francisco Sánchez, Elaine Limbrick, and D. F. S. (Douglas Ferguson Scott) Thomson. That nothing is known = (Quod nihil scitur). Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-35077-8. x + 310 pp. LCCN B785.S23 Q6813

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87033395.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        87033395.html;   http://

www.loc.gov/catdir/toc/
cam031/87033395.html.

Smith:1994:DTD

[SM94] Merritt Roe Smith and Leo Marx, editors. Does technology drive history?: the dilemma of technological determinism. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-19347-7, 0-262-69167-1 (paperback). xv + 280 pp. LCCN T14.5 .D64 1994.

Smeaton:1997:HSU

[Sme97] William A. Smeaton. History of science at University College London: 1919–1947. British Journal for the History of Science, 30(1):25–28, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027897.

Smith:1982:EUA

[Smi82] Robert W. (Robert William)

Smith. The expanding universe: astronomy’s “Great Debate,” 1900–1931. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-23212-0 (US). xiv + 220 pp. LCCN QB981 .S69 1982.

Smith:1988:RT

[Smi88] F. B. (Francis Barrymore)

Smith. The retreat of tuberculosis, 1850–1950. Croom Helm, London, UK, 1988. ISBN 0-7099-3383-5. 271 pp. LCCN RC310.5 .S65 1988. US$25.00 (est.).

Smith:1989:STS

[Smi89] Robert W. (Robert William) Smith. The space telescope: a study of NASA, science, technology, and politics. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-26634-3. xviii + 478 pp. LCCN QB500.267 .S55

1989. URL http://www.loc.
gov/catdir/enhancements/
fy0803/89000707-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0803/89000707-t.html.

Smith:1990:BRS

[Smi90a] Crosbie Smith. Book reviews: Stephen G. Brush (ed.). History of Physics: Selected Reprints. College Park, MD: American Association of Physics Teachers, 1988. Pp. 235. ISBN 0-917853-29-6. $16.00. Stephen G. Brush, The History of Modern Science. A Guide to the Second Scientific Revolution, 1800–1950. Ames, Iowa: Iowa State University Press, Pp. xv + 544. ISBN 0-8138-0883-9. $39.95. British Journal for the History of Science, 23(1): 119–120, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026824.

Smith:1990:BM

[Smi90b]  Roger Smith.  Back mat-

ter. British Journal for the History of Science, 23(4):505– 514, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027229.

Smith:1990:RUa

[Smi90c] Roger Smith. Review: [untitled]. British Journal for the History of Science, 23(4): 491–493, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027224.

Smith:1990:RUb

[Smi90d] Roger Smith. Review: [untitled]. British Journal for the History of Science, 23(4): 495–496, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027226.

Smith:1991:BRR

[Smi91a] Crosbie Smith. Book review:

Robert A. Stafford. Scientist of Empire: Sir Roderick Murchison, Scientific Exploration and Victorian Imperialism. Cambridge: Cambridge University Press, 1990. Pp. xii + 293. ISBN 0-521-33537-X. £30.00, $49.50. British Journal for the History of Science, 24(2): 267–269, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027189.

Smith:1991:BRI

[Smi91b] N. A. F. Smith. Book review: Ian McNeil (ed.). An Encyclopaedia of the History of Technology. London and New York: Routledge, 1990. Pp. xv + 1062. ISBN 1-415-01306-2. £65.00. British Journal for the History of Science, 24 (1):119, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027034.

Smith:1991:BRP

[Smi91c] Roger Smith. Book review: Peter J. Bowler. The Invention of Progress: The Victorians and the Past. Oxford: Basil Blackwell, 1989. Pp. xi + 231. ISBN 0-631-16107-4. £25.00. British Journal for the History of Science, 24(2):266–267, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027188.

Smith:1992:BRS

[Smi92a] A. Mark Smith. Book review: A. I. Sabra. The Optics of Ibn al-Haytham. Books I, II, II: On Direct Vision. With Translation, Introduction, Commentary, Glossaries. London: The Warburg Institute, 1989. Pp. 735 (in two vols.). ISBN 0-85481-072-2. British Journal for the History of Science, 25 (3):358–359, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027260.

Smith:1992:BRD

[Smi92b]  Crosbie Smith.  Book re-

view: David B. Wilson (ed.). The Correspondence Between Sir George Gabriel Stokes and Sir William Thomson, Baron Kelvin of Largs. Volume 1: 1846–1869; Volume 2: 1870– 1901. Cambridge: Cambridge University Press, 1990. Pp. lvi + ix + 783. ISBN 0-521-32831-4. £125.00, $195.00. British Journal for the History of Science, 25(2):278–279, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027317.

Smith:1992:BRA

[Smi92c] Roger Smith. Book review: Alfred Russel Wallace. An Anthology of His Shorter Writings, edited by Charles H. Smith. Oxford: Oxford University Press, 1991. Pp. x + 551. ISBN 0-19-857725-7. £40.00. British Journal for the History of Science, 25(4): 482–483, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027070.

Smith:1992:BRL

[Smi92d] Roger Smith. Book review:

Lucille B. Ritvo. Darwin’s Influence on Freud: A Tale of Two Sciences. New Haven and London: Yale University Press, 1990. pp. xii + 267. ISBN 0-300-04131-4. £19.95, $35.00. British Journal for the History of Science, 25(2): 280–281, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027319.

Smith:1992:IHM

[Smi92e] Roger Smith. Inhibition: history and meaning in the sciences of mind and brain. Free Association Books, London, UK, 1992. ISBN 1-85343-181-8. xi + 333 pp. LCCN QP353 .S65 1992b.

Smith:1993:BRG

[Smi93a] A. D. Smith. Book review: Gary Hatfield, The Natural and the Normative: Theories of Spatial Perception from Kant to Helmholtz. Cambridge, Mass, and London: MIT Press, 1991. Pp. xii + 366. ISBN 0-262-08086-9. £31.50. British Journal for the History of Science, 26(1):93–95, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027358.

Smith:1993:BRB

[Smi93b] Robert W. Smith. Book review: B. Bertotti, R. Balbinot, S. Bergia and A. Messina (eds.), Modern Cosmology in Retrospect. Cambridge: Cambridge University Press, 1990. Pp. xx + 426. ISBN 0-521-37213-5.

£45.00, $69.50. British Journal for the History of Science, 26 (1):124–125, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027386.

Smith:1993:BRL

[Smi93c] Roger Smith. Book review:

Léon Chertok and Isabelle Stengers, A Critique of Psychoanalytic Reason: Hypnosis as a Scientific Problem from Lavoisier to Lacan. Stanford: Stanford University Press, 1992. Pp. xxvi + 319. ISBN 0-8047-1950-0. $35.00. British Journal for the History of Science, 26(2): 252–253, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027116.

Smith:1993:BRR

[Smi93d] Roger Smith. Book review:

René van der Veer and Jaan Valsiner, Understanding Vygotsky: A Quest for Synthesis. Oxford: Blackwell, 1991. Pp. xi + 450. ISBN 0-631-16528-2. £35.00. British Journal for the History of Science, 26(1): 122–123, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027384.

Smith:1994:BRDb

[Smi94a] Crosbie Smith. Book review:

D. M. Cannell, George Green, Mathematician and Physicist, 1793–1841. London: Athlone, 1993. Pp. xxvi + 265. ISBN 0-485-11433-X. £35.00. British Journal for the History of Science, 27(4):477–479, December 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027631.

Smith:1994:FFB

[Smi94b] Jonathan Smith. Fact and feeling: Baconian science and the Nineteenth-Century literary imagination. Science and literature. University of Wisconsin Press, Madison, WI, USA, 1994. ISBN 0-299-14350-3, 0-299-14354-6 (paperback). ix + 277 pp. LCCN PR468.S34 S65 1994.

Smith:1994:BAS

[Smi94c] Pamela H. Smith. The business of alchemy: science and culture in the Holy Roman Empire. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-05691-9. xii + 308 pp. LCCN Q125.2 .S58 1994. US$45.00. URL http:

//www.loc.gov/catdir/
description/prin031/
93044856.html;   http://
www.loc.gov/catdir/toc/
prin031/93044856.html.

Smith:1994:BRDa

[Smi94d] Roger Smith. Book review: David Blitz, Emergent Evolution: Qualitative Novelty and the Levels of Reality. Episteme 19. Dordrecht: Kluwer Academic Publishers, 1992. Pp. ix + 239. ISBN 0-7923-1658-4. £52.00, $89.00, Dfl. 150.00. British Journal for the History of Science, 27(2):237– 238, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027451.

Smith:1994:SFS

[Smi94e] Tim D. (Tim Denis) Smith.

Scaling fisheries: the science of measuring the effects of fishing, 1855–1955. Cambridge studies in applied ecology and resource management. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-39032-X. x + 392 pp. LCCN SH329.F56 S58

1994. URL http://www.loc.
gov/catdir/description/
cam026/93038292.html;
http://www.loc.gov/
catdir/toc/cam029/
93038292.html.

Smith:1995:RU

[Smi95a] David Smith. Review: [untitled]. British Journal for the History of Science, 28(4): 487–488, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027568.

Smith:1995:BRD

[Smi95b] Roger Smith. Book review: Dorothy Ross (ed.), Modernist Impulses in the Human Sciences 1870– 1930. Baltimore: Johns Hopkins University Press, 1994. Pp. xii + 379. ISBN 0-8018-4744-3, £48.00 (hardcover); 0-8018-4745-1. £15.50 (paperback). British Journal for the History of Science, 28(3):359– 361, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027657.

Smith:1996:BRK

[Smi96] Crosbie Smith. Book review: Kenneth L. Caneva, Robert Mayer and the Conservation of Energy. Princeton: Princeton University Press, 1993. Pp. xxiii + 439. ISBN 0-691-08758-X. £33.00, $49.50. British Journal for the History of Science, 29(3):372–373, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027748.

Smith:1997:FHH

[Smi97] Roger Smith. The Fontana history of the human sciences. Fontana history of science. Fontana Press, London, UK, 1997. ISBN 0-00-686178-4. xviii + 1036 + 8 pp. LCCN H51 .S64 1997.

                                        URL http://books.google.
                                        com/books?id=4395AAAAIAAJ.

Smith:1998:BRC

[Smi98] Robert W. Smith. Book review: Cecilia Payne-Gaposchkin, An Autobiography and Other Recollections, second edition, edited by Katherine Haramundanis. Cambridge: Cambridge University Press, 1996. Pp. xxii + 277. ISBN 0-521-48251-8, £35.00 (hardcover); 0-521-48930-5, £12.95 (paperback). British Journal for the History of Science, 31(3):361– 375, September 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027825.

Smith:1999:BRL

[Smi99a] Roger Smith. Book review:

Loren Graham, What Have We Learned about Science and Technology from the Russian Experience? Stanford: Stanford University Press, 1998. Pp. xii + 177. ISBN 0-8047-2985-9, £27.95, $39.50 (hardcover); 0-8047-3276-0, £10.95, $14.95 (paperback). British Journal for the History of Science, 32 (1):111–124, March 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027980.

Smith:1999:BRP

[Smi99b] Roger Smith. Book reviews:

Polly Young-Eisendrath and Terence Dawson (eds), The Cambridge Companion to Jung. Cambridge: Cambridge University Press, 1997. Pp. xxxiv + 332. ISBN 0-521-47309-8 (hardcover). £35.00, $49.95; 0-521-47889-8 (paperback). £12.95, $17.95. Sonu Shamdasani, Cult Fictions: C. G. Jung and the Founding of Analytical Psychology. London and New York: Routledge, 1998. Pp. x + 121. ISBN 0-415-18613-7 (hardcover). £35.00; 0-415-18641-5 (paperback). £12.99. British Journal for the History of Science, 32(3):363–378, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4028002.

Smoller:1994:HPS

[Smo94] Laura Ackerman Smoller. History, prophecy, and the stars: the Christian astrology of Pierre d’Ailly, 1350–1420. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-08788-1. xii + 233 pp. LCCN BF1679.8.A453 S56 1994. US$35.00. URL http:

//www.loc.gov/catdir/
description/prin021/
93044907.html;   http://
www.loc.gov/catdir/toc/
prin031/93044907.html.

Smyth:1989:AEM

[Smy89] Henry De Wolf Smyth. Atomic energy for military purposes: the official report on the development of the atomic bomb under the auspices of the United States Government, 1940–1945. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1721-4 (hardcover), 0-8047-1722-2 (paperback). xv + 324 pp. LCCN QC773.A1 S69 1989.

Sharov:1993:EHD

[SN93]  A.  S.

(Aleksandr Sergeevich) Sharov and I. D. (Igor Dmitrievich) Novikov. Edwin Hubble, the discoverer of the big bang universe. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-41617-5. xv + 187 pp. LCCN QB36.H83 S5313

1993. URL http://www.loc.
gov/catdir/description/
cam025/92024683.html;
http://www.loc.gov/
catdir/toc/cam028/
92024683.html.

Snobelen:1999:INH

[Sno99] Stephen D. Snobelen. Isaac Newton, heretic: The strategies of a Nicodemite. British Journal for the History of Science, 32(4):381–419, December 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027945.

Scharlau:1985:FML

[SO85] Winfried Scharlau and Hans Opolka. From Fermat to Minkowski: lectures on the theory of numbers and its historical development. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-90942-7. xi + 184 pp. LCCN QA241 .S2813 1985.

Soderqvist:1986:EMN

[Söd86] Thomas Söderqvist. The ecologists: from merry naturalists to saviours of the nation: a sociologically informed narrative survey of the ecologization of Sweden 1895–1975. Almqvist and Wiksell International, Stockholm, Sweden, 1986. ISBN 91-22-00827-6 (paperback), 91-22-00812-8. vii + 330 pp. LCCN QH541 .S62x 1986.

Sorabji:1983:TCC

[Sor83] Richard Sorabji. Time, creation and the continuum: theories in antiquity and the early Middle Ages. Duckworth, London, UK, 1983. ISBN 0-7156-1693-5. xviii + 473 pp. LCCN BD638 .S67 1983. US$29.50.

Sorell:1986:H

[Sor86] Tom Sorell. Hobbes. The Arguments of the philosophers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1986. ISBN 0-7100-9845-6. xii + 163 pp. LCCN B1247 .S63 1986. US$25.00.

Sorell:1987:D

[Sor87] Tom Sorell. Descartes. Past masters; Oxford paperbacks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987. ISBN 0-19-287635-X (paperback), 0-19-287636-8 (hardcover). 120 pp. LCCN B1875 .S67 1987. US$4.50; US$15.00. URL http://www.

loc.gov/catdir/
enhancements/fy0637/
86023662-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0724/86023662-b.html.

Sorrenson:1999:GGV

[Sor99a]  Richard  Sor-

renson. George Graham, visible technician. British Journal for the History of Science, 32 (2):203–221, June 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4028084.

Sorrenson:1999:I

[Sor99b] Richard Sorrenson. Introduction. British Journal for the History of Science, 32(2):130– 132, June 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4028079.

Southgate:1993:CTL

[Sou93] Beverley C. Southgate. “Covetous of truth”: the life and work of Thomas White, 1593– 1676, volume 134 of Archives internationales d’histoire des idées. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 0-7923-1926-5 (hardcover). xi + 189 pp. LCCN B1299.W454 S68

1993. URL http://www.loc.

                                        gov/catdir/enhancements/
                                        fy0821/92026439-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0821/92026439-t.html.

Soyland:1996:BRR

[Soy96]  A. J. Soyland.  Book re-

view: Roger Smith, Inhibition: History and Meaning in the Sciences of Mind and Brain. London: Free Association Books, 1992. Pp. ix + 333. ISBN 1-85343-181-8. £37.50. British Journal for the History of Science, 29(1): 103–104, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027525.

Sheets-Pyenson:1988:CSD

[SP88] Susan Sheets-Pyenson. Cathedrals of science: the development of colonial natural history museums during the late Nineteenth Century. McGill-Queen’s University Press, Kingston, ON, Canada, 1988. ISBN 0-7735-0655-1. xii + 144 + 20 pp. LCCN QH70.A1 S54 1988. US$24.95.

Spencer:1998:FSC

[SP98] Hamish G. Spencer and Diane B. Paul. The failure of a scientific critique: David Heron, Karl Pearson and Mendelian eugenics. British Journal for the History of Science, 31(4): 441–452, December 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027877.

Spalding:1991:BRC

[Spa91] David A. E. Spalding. Book review: Charles H. Sternberg. The Life of a Fossil Hunter. Introduction by Rudolf A. Raff. Bloomington and Indiana: University of Indiana Press, 1990. Pp. xxx + 286. ISBN 0-253-35549-4, $29.95 (hardcover); 0-253-20571-9, $12.95 (paperback). British Journal for the History of Science, 24(3):393– 394, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027250.

Spary:1994:BRP

[Spa94] Emma Spary. Book review:

Pascal Duris, Linné et la France (1780–1850). Histoire des idées et critique littéraire, 318. Geneva: Droz, 1993. Pp. 218. No price given. British Journal for the History of Science, 27 (3):373–374, September 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027612.

Spary:1995:BRT

[Spa95] E. C. Spary. Book review:

Tore Frängsmyr (ed.), Linnaeus: The Man and His Work. Nantucket: Science History Publications, 1994. Pp. xiv + 206. ISBN 0-88135-154-7. $14.95. British Journal for the History of Science, 28 (2):237–238, June 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027683.

Spary:1998:BRH

[Spa98a] Emma Spary. Book review: Harmke Kamminga and Andrew Cunningham (eds.), The Science and Culture of Nutrition, 1840–1940. Clio Medica, 32/Wellcome Institute Series in the History of Medicine. Amsterdam: Rodopi, 1995. Pp. vii + 344. ISBN 90-5183-818-2. $33.00, Dfl. 50.00. British Journal for the History of Science, 31(1):63–102, March 1998. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027807.

Spary:1998:BRT

[Spa98b] Emma Spary. Book review: Theodore W. Pietsch (ed.), Fishes, Crayfishes, and Crabs: Louis Renard’s Natural History of the Rarest Curiosities of the Seas of the Indies, 2 vols. Baltimore: Johns Hopkins University Press, 1996. Pp. viii + 224, xxii + 214, illus. ISBN 0-8018-4790-7. £78.50 (set). British Journal for the History of Science, 31(1):63– 102, March 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027808.

Speakman:1982:ASG

[Spe82] Colin Speakman. Adam Sedgwick, geologist and dalesman, 1785–1873: a biography in twelve themes. Broad Oak Press, Broad Oak, Heathfield, East Sussex, UK, 1982. ISBN 0-906716-01-2. 145 pp. LCCN QE22.S5 S64 1982. US$5.75.

Springford:1997:ECV

[Spr97] Michael Springford, editor. Electron: a centenary volume. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56130-2 (hardcover). xii + 330 pp. LCCN QC793.5.E62 E39

1997. URL http://www.loc.
gov/catdir/description/
cam028/96044744.html;
http://www.loc.gov/
catdir/toc/cam021/
96044744.html.

Shapin:1985:LAP

[SSH85] Steven Shapin, Simon Schaffer, and Thomas Hobbes. Leviathan and the air-pump: Hobbes, Boyle, and the experimental life: including a translation of Thomas Hobbes, Dialogus physicus de natura aeris by Simon Schaffer. Princeton University Press, Princeton, NJ, USA, 1985. ISBN 0-691-08393-2. xiv + 440 pp. LCCN QC166 .S47 1985. US$59.50.

Schmitt:1988:CHR

[SSK88] Charles B. Schmitt, Quentin Skinner, and Eckhard Kessler, editors. The Cambridge history of Renaissance philosophy. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-25104-4. xiii + 968 pp. LCCN B775 .C25

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87005212.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam023/
                                        87005212.html.

Schroder:1994:ELD

[ST94] W. Schröder and H.-J. Treder. The ‘Einstein–Laue’ discussion. British Journal for the History of Science, 27(1):113– 114, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027585.

Stafford:1984:VSA

[Sta84a] Barbara Maria Stafford. Voyage into substance: art, science, nature, and the illustrated travel account, 1760– 1840. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-19223-3. xxiii + 645 pp. LCCN G468 .S73 1984.

Stansfield:1984:TBM

[Sta84b] Dorothy A. Stansfield. Thomas Beddoes, M.D., 1760–1808: chemist, physician, democrat. Chemists and chemistry. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1984. ISBN 90-277-1686-2. xix + 306 pp. LCCN QD22.B263 S83 1984.

Staudenmaier:1985:TSR

[Sta85] John M. Staudenmaier. Technology’s storytellers: reweaving the human fabric. Society for the History of Technology and the MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-19237-3. xxvii + 282 pp. LCCN T20 .S73 1985.

Stafford:1989:SES

[Sta89a] Robert A. Stafford. Scientist of empire: Sir Roderick Murchison, scientific exploration and Victorian imperialism. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33537-X. xii + 293 pp. LCCN QE22.M8 S82

1989. URL http://www.loc.
gov/catdir/description/
cam023/88034134.html;
http://www.loc.gov/
catdir/samples/cam031/
88034134.html;   http://
www.loc.gov/catdir/toc/
cam027/88034134.html.

Star:1989:RMB

[Sta89b] Susan Leigh Star. Regions of the mind: brain research and the quest for scientific certainty. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1673-0. xiii + 278 pp. LCCN QP385 .S66 1989. US$27.50. URL http:

//www.loc.gov/catdir/

                                        description/cam023/
                                        89004432.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam025/89004432.html.

Stafford:1990:BM

[Sta90a] Robert A. Stafford. Back matter. British Journal for the History of Science, 23(1): 128, March 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026828.

Stafford:1990:BRJ

[Sta90b] Robert A. Stafford. Book review: John M. Mackenzie. The Empire of Nature: Hunting, Conservation and British Imperialism. Manchester and New York: Manchester University Press, 1988. Pp. x + 340. ISBN 0-7190-2227-4. £35.00. British Journal for the History of Science, 23(1): 122–124, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026826.

Stanton:1990:HEI

[Sta90c] Charles M. Stanton. Higher education in Islam: the classical period. Rowman and Littlefield, Lanham, MD, USA, 1990. ISBN 0-8476-7637-4. ???? pp. LCCN BP.

Stafford:1991:BCI

[Sta91] Barbara Maria Stafford. Body criticism: imaging the unseen in Enlightenment art and medicine. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-19304-3. xxi + 587 pp. LCCN BF697.5.B63 S73 1992.

Stafford:1994:ASE

[Sta94] Barbara Maria Stafford. Artful science: enlightenment, entertainment, and the eclipse of visual education. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-19342-6. xxix + 350 pp. LCCN QP495 .S73 1994. US$35.00.

Staley:1995:RU

[Sta95] Richard Staley. Review: [untitled]. British Journal for the History of Science, 28(4): 482–483, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027564.

Stein:1985:ALL

[Ste85] Dorothy Stein. Ada, a life and a legacy. MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-19242-X. xix + 321 + 16 pp. LCCN QA29.L72 S74 1985. US$19.95.

Stephenson:1987:KPA

[Ste87] Bruce Stephenson. Kepler’s physical astronomy, volume 13 of Studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 3-540-96541-6 (Berlin) 0-387-96541-6 (New York). v + 216 pp. LCCN ????

Sternberg:1990:LFH

[Ste90a] Charles H. (Charles Hazelius) Sternberg. The life of a fossil hunter. Indiana University Press, Bloomington, IN, USA, 1990. ISBN 0-253-35549-4, 0-253-20571-9 (paperback). xxvii + 286 + 46 pp. LCCN QE707.S8 A3 1990.

Stewart:1990:DCC

[Ste90b]  John A. Stewart.  Drift-

ing continents and colliding paradigms: perspectives on the geoscience revolution. Science, technology, and society. Indiana University Press, Bloomington, IN, USA, 1990. ISBN 0-253-35405-6. xii + 285 pp. LCCN QE511.5 .S86 1990.

Steigerwald:1991:BRA

[Ste91a] Joan Steigerwald. Book review: Andrew Cunningham and Nicholas Jardine (eds.). Romanticism and the Sciences. Cambridge: Cambridge University Press, 1990. Pp. xxii + 345. ISBN 0-521-35602-4, £40.00, $59.50 (hardcover); 0-521-35685-7, £15.00, $19.95 (paperback). British Journal for the History of Science, 24 (3):384–386, September 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027245.

Steinberg:1991:NII

[Ste91b] Theodore Steinberg. Nature incorporated: industrialization and the waters of New England. Studies in environment and history. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-39215-2. xv + 284 pp. LCCN TS1323.15 .S74

1991. URL http://www.loc.
gov/catdir/description/
cam024/90048485.html;
http://www.loc.gov/
catdir/samples/cam031/
90048485.html;   http://
www.loc.gov/catdir/toc/
cam029/90048485.html.

Stemerding:1991:PAF

[Ste91c] Dirk Stemerding. Plants, animals and formulae: natural history in the light of Latour’s Science in action and Foucault’s The order of Things, volume 4 of WMW-publikatie. Faculteit Wijsbegeerte en Maatschappijwetenschappen, Universiteit Twente, Enschede, The Netherlands, 1991. ISBN 90-365-0379-5. ISSN 0925-0131. 202 pp. LCCN QH83 .S7 1991.

URL http://arno.unimaas.
nl/show.cgi?fid=5638.

Stewart:1992:RPS

[Ste92] Larry R. Stewart. The rise of public science: rhetoric, technology, and natural philosophy in Newtonian Britain, 1660– 1750. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41700-7. xxxiv + 453 pp. LCCN Q127.G4 STE 1992; Sci.315; Q; R/STE; Hist.

Stephenson:1994:MHK

[Ste94] Bruce Stephenson. The music of the heavens: Kepler’s harmonic astronomy. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03439-7. xi + 260 pp. LCCN QB361 .S74 1994. US$39.50. URL http://www.

                                        loc.gov/catdir/
                                        description/prin031/
                                        93044916.html;   http://
                                        www.loc.gov/catdir/toc/
                                        prin031/93044916.html.

Steigerwald:1995:BRS

[Ste95] Joan Steigerwald. Book review: Stefano Poggi and Maurizio Bossi (eds.), Romanticism in Science: Science in Europe, 1790–1840. Boston Studies in the Philosophy of Science, 152. Dordrecht: Kluwer Academic Publishers, 1994. Pp. xv + 245. ISBN 0-7923-2336-X. £70.00, $107.00, Dfl. 175.00. British Journal for the History of Science, 28(2):240–241, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027686.

Stepansky:1997:AAW

[Ste97a] Barbara K. Stepansky. Ambiguity: aspects of the wave– particle duality. British Journal for the History of Science, 30 (3):375–385, September 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027868.

Stepansky:1997:BM

[Ste97b] Barbara K. Stepansky. Back matter. British Journal for the History of Science, 30(3): 386, September 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027869.

Stewart:1999:OCC

[Ste99] Larry Stewart. Other centres of calculation, or, where the Royal Society didn’t count: commerce, coffee-houses and natural philosophy in early modern London. British Journal for the History of Science, 32 (2):133–153, June 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4028080.

Stigler:1986:HSM

[Sti86] Stephen M. Stigler. The history of statistics: the measurement of uncertainty before 1900. Belknap Press of Harvard University Press, Cambridge, MA, USA and London, UK, 1986. ISBN 0-674-40340-1. xvi + 410 pp. LCCN QA276.15 .S75 1986.

Stocking:1983:OOE

[Sto83] George W. Stocking, editor.

Observers observed: essays on ethnographic fieldwork, volume 1 of History of anthropology. University of Wisconsin Press, Madison, WI, USA, 1983. ISBN 0-299-09450-2. vi + 242 pp. LCCN GN346 .O27 1983. US$19.95.

Stocking:1984:FHE

[Sto84] George W. Stocking, editor. Functionalism historicized: essays on British social anthropology, volume 2 of History of anthropology. University of Wisconsin Press, Madison, WI, USA, 1984. ISBN 0-299-09900-8. vi + 244 pp. LCCN GN363 .F86 1984. US$19.95.

Stokes:1988:DPH

[Sto88] Raymond G. Stokes. Divide and prosper: the heirs of I. G. Farben under Allied authority, 1945–1951. University of California Press, Berkeley, CA, USA, 1988. ISBN 0-520-06248-5. xiv + 290 pp. LCCN HD9654.9.I5 S76 1988.

Stroup:1987:RFP

[Str87] Alice Stroup. Royal funding of the Parisian Académie royale des sciences during the 1690s, volume 77, 4 of Transactions of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1987. ISBN 0-87169-774-2. ISSN 0065-9746. 167 pp. LCCN Q46 .S77 1987.

Strong:1990:BRG

[Str90] Tracy B. Strong. Book review: G. A. J. Rogers & Alan Ryan (eds). Perspectives on Thomas Hobbes. Clarendon Press: Oxford, 1988. Pp. x + 209. IBSN 0-19-824998-5. £22.50. British Journal for the History of Science, 23(3): 353–355, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026776.

Sampson:1987:HRA

[STT87] R. A. (Ralph Allen) Sampson, R. J. (Roger John) Tayler, and H. H. (Herbert Hall) Turner, editors. History of the Royal Astronomical Society, Vol. 2, 1920–1980. Published for the Society by Blackwell Scientific Publications, Oxford, UK, 1987. ISBN 0-632-01791-0 (hardback:v. 2), 0-632-01792-9 (paperback:v. 2). vii + 262 + 14 pp. LCCN QB1.R5.

Sturdy:1991:BRA

[Stu91] Stephen Sturdy. Book reviews:

Alan Chalmers. Science and its Fabrication. Milton Keynes: Open University Press, 1990. Pp. xii + 142. ISBN 0-335-09318-3, £22.50 (hardcover); ISBN 0-335-09317-5, £6.99 (paperback). British Journal for the History of Science, 24(1): 128, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027042.

Sturdy:1995:SSS

[Stu95] D. J. (David J.) Sturdy. Science and social status: the members of the Academie des sciences 1666–1750. Boydell Press, Woodbridge, Suffolk, UK, 1995. ISBN 0-85115-395-X. xvi + 461 pp. LCCN Q149.F7 S78 1995.

Sturdy:1998:BRC

[Stu98] Steve Sturdy. Book review: Charles Webster, The National Health Service: A Political History. Oxford: Oxford University Press, 1998. Pp. xiv + 241. ISBN 0-19-289296-7. £9.99 (paperback edition). British Journal for the History of Science, 31(4):469–487, December 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027887.

Stokes:1990:CBS

[STW90] George Gabriel Stokes, William Thomson (Baron Kelvin), and David B. Wilson. The correspondence between Sir George Gabriel Stokes and Sir William Thomson, Baron Kelvin of Largs. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-32831-4 (set). lvi + ix + 783 pp. LCCN QC7 .S79

                                        1990. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/89017463.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        89017463.html.

Sullivan:1982:JTD

[Sul82a]  Robert E. Sullivan.  John

Toland and the Deist controversy: a study in adaptations, volume 101 of Harvard historical studies. Harvard University Press, Cambridge, MA, USA, 1982. ISBN 0-674-48050-3. viii + 355 pp. LCCN BL2790.T6 S94.

Sullivan:1982:CRA

[Sul82b] Woodruff Turner Sullivan, editor. Classics in radio astronomy, volume 10 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1982. ISBN 90-277-1356-1, 90-277-1363-4. xxiv + 348 pp. LCCN QB475 .C55 1982.

Sutton:1994:NVN

[Sut94] Clive Sutton. ‘nullius in verba’ and ‘nihil in verbis’: public understanding of the role of language in science. British Journal for the History of Science, 27(1):55–64, March 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027580.

Sutton:1995:SPS

[Sut95] Geoffrey V. Sutton. Science for a polite society: gender, culture, and the demonstration of enlightenment. Westview Press, Boulder, CO, USA, 1995. ISBN 0-8133-1575-1 (hardcover), 0-8133-1576-X (paperback). xiii + 391 pp. LCCN Q125 .S982

1995. URL http://www.loc.
gov/catdir/enhancements/
fy0832/95020157-b.html;

                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0832/95020157-d.html.

Sutton:1997:BRG

[Sut97] John Sutton. Book review: Georges Canguilhem, A Vital Rationalist: Selected Writings from Georges Canguilhem, edited by François Delaporte and translated by Arthur Goldhammer. New York: Zone Books, 1994. Pp. 481. ISBN 0-942299-72-8. £24.25, $36.25. British Journal for the History of Science, 30(1):101– 121, March 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027913.

Sutton:1999:PAL

[Sut99] Clive Sutton. Preparation for adult life? British Journal for the History of Science, 32(3): 359–361, September 1999. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027993.

Shortland:1989:THS

[SW89] Michael Shortland and Andrew Warwick, editors. Teaching the history of science. British Society for the History of Science, Chalfont St. Giles, Bucks, England, 1989. ISBN 0-631-16977-6, 0-631-16978-4 (paperback). 281 pp. LCCN Q126.9 .T4 1989. US$59.00 (U.S. : est.), US$19.90 (U.S. : est.).

Sprenger:1993:PHU

[SW93] Elizabeth Sprenger and Pauline Webb. Persuading the housewife to use electricity? An interpretation of material in the Electricity Council archives. British Journal for the History of Science, 26(1):55–65, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027340.

Stewart:1995:PTN

[SW95] Larry Stewart and Paul Weindling. Philosophical threads: natural philosophy and public experiment among the weavers of Spitalfields. British Journal for the History of Science, 28 (1):37–62, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027483.

Swann:1988:ASP

[Swa88] John Patrick Swann. Academic scientists and the pharmaceutical industry: cooperative research in Twentieth-Century America. The Johns Hopkins University Press, Baltimore, MD, USA, 1988. ISBN 0-8018-3558-5. xi + 249 pp. LCCN RS122 .S93 1988.

Swann:1992:BRM

[Swa92] John P. Swann. Book review: Milton Wainwright. Miracle Cure: The Story of Penicillin and the Golden Age of Antibiotics. Oxford: Basil Blackwell, 1990. Pp. xi + 196. ISBN 0-631-16492-8. £16.95. British Journal for the History of Science, 25(3):376–377, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027277.

Schuster:1986:PRS

[SY86] John Andrew Schuster and Richard R. Yeo, editors. The politics and rhetoric of scientific method: historical studies, volume 4 of Australasian studies in history and philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1986. ISBN 90-277-2152-1. xxxix + 305 pp. LCCN Q175.3 .P65 1986. US$64.50.

Shortland:1996:TLS

[SY96]  Michael  Shortland

and Richard R. Yeo, editors. Telling lives in science: essays on scientific biography. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-43323-1 (hardcover). xiv + 295 pp. LCCN Q141 .T45

1996. URL http://www.loc.
gov/catdir/description/
cam027/95047399.html;
http://www.loc.gov/
catdir/toc/cam021/
95047399.html.

Thackray:1985:CAH

[T^{+}85] Arnold Thackray et al., editors.

Chemistry in America, 1876– 1976: historical indicators. Chemists and chemistry. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1985. ISBN 90-277-1720-6. xxiii + 564 pp. LCCN QD18.U6 C47 1985.

Takahashi:1992:MLT

[Tak92] Ken’ichi Takahashi. The medieval Latin traditions of Euclid’s Catoptrica: a critical edition of De speculis with an introduction, English translation and commentary. Kyushu University Press, Fukuoka-shi, Japan, 1992. ISBN 4-87378-299-6. viii + 373 pp. LCCN ????

Taub:1993:EIE

[Tau93] Liba Taub. Evolutionary ideas and ‘empirical’ methods: the analogy between language and species in works by Lyell and Schleicher. British Journal for the History of Science, 26 (2):171–193, June 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027096.

Taylor:1987:HRA

[Tay87] R. J. (Roger John) Taylor, editor. History of the Royal Astronomical Society: volume 2, 1920–1980. Reprinted for the Society by Blackwell Scientific Publications, Palo Alto, CA, USA, 1987. ISBN 0-632-01791-0, 0-632-01792-9 (paperback). vii + 262 pp. LCCN QB1.R75 H57 1987.

Taylor:1989:RG

[Tay89] Selwyn Taylor. Robert Graves. Royal Society of Medicine Services, London, UK and New York, NY, USA, 1989. ISBN 0-905958-98-5 (hardcover), 0-905958-90-X (paperback). 160 pp. LCCN ???? US$15.00; US$7.50.

Taylor:1996:WJC

[Tay96] Eugene Taylor. William James on consciousness beyond the margin. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-01136-2. xiii + 215 pp. LCCN BF109.J28

                                        T38 1996. URL http://www.
                                        loc.gov/catdir/
                                        description/prin021/
                                        96000898.html;   http://
                                        www.loc.gov/catdir/toc/
                                        prin031/96000898.html.

Tannoch-Bland:1997:DSI

[TB97]  Jennifer  Tannoch-Bland.

Dugald Stewart on intellectual character. British Journal for the History of Science, 30(3): 307–320, September 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027864.

Thomson:2010:CTF

[TBDA10] Ann Thomson, Simon Burrows, Edmond Dziembowski, and Sophie Audidière, editors. Cultural transfers: France and Britain in the long Eighteenth Century, volume 2010:04 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2010. ISBN 0-7294-0993-7 (paperback). xi + 326 pp. LCCN DA47.1 .C86 2010.

Thoren:1990:LUB

[TC90] Victor E. Thoren and J. R.

(John Robert) Christianson. The Lord of Uraniborg: a biography of Tycho Brahe. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-35158-8 (hardcover). xi + 523 pp. LCCN QB36.B8 T49

1990. URL http://www.loc.
gov/catdir/description/
cam032/90001477.html;
http://www.loc.gov/
catdir/samples/cam034/
90001477.html;   http://
www.loc.gov/catdir/toc/
cam028/90001477.html.

Tauber:1991:MOI

[TC91] Alfred I. Tauber and Leon Chernyak. Metchnikoff and the origins of immunology: from metaphor to theory. Monographs on the history and philosophy of biology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-506447-X. xviii + 247 pp. LCCN QR180.72.M48 T38

1991. URL http://www.loc.
gov/catdir/enhancements/
fy0603/90026245-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0603/90026245-t.html.

Teich:1991:LE

[Tei91] Mikulás Teich. A life with enzymes. British Journal for the History of Science, 24(1): 95–99, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027019.

Teich:1993:BRJ

[Tei93]  Mikulás Teich.  Book re-

view: Joseph S. Fruton, Contrasts in Scientific Style: Research Groups in the Chemical and Biochemical Sciences. Philadelphia: American Philosophical Society, 1990. Pp. xi + 473. ISBN 0-87169-191-4. $40.00. British Journal for the History of Science, 26(2): 256–259, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027120.

Tropp:1993:AFM

[TFC93] E. A. (Eduard Abramovich) Tropp, Viktor Iakovlevich Frenkel, and A. D. (Artur Davidovich) Chernin. Alexander A. Friedmann: the man who made the universe expand. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-38470-2. x + 267 pp. LCCN QC16.F73 T7613

                                        1993. URL http://www.loc.
                                        gov/catdir/description/
                                        cam025/92028315.html;
                                        http://www.loc.gov/

catdir/toc/cam028/
92028315.html.

Thackray:1991:BRK

[Tha91] John Thackray. Book review:

K. J. Tinkler (ed.). History of Geomorphology from Hutton to Hack. The Binghamton Symposia in Geomorphology International Series, no. 19. Boston: Unwin Hyman, 1989. Pp. 344. ISBN 0-040551138-1. £40. British Journal for the History of Science, 24(1): 121–122, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027037.

Travis:1992:OCH

[THBB92a] Anthony S. Travis, Willem J.

Hornix, Robert Bud, and John J. Beer. Organic chemistry and high technology, 1850–1950. British Journal for the History of Science, 25(1):1–4, March 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027001.

Travis:1992:ZRB

[THBB92b] Anthony S. Travis, Willem J.

Hornix, Robert Bud, and Robert Bud. The zymotechnic roots of biotechnology. British Journal for the History of Science, 25(1): 127–144, March 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027008.

Travis:1992:ERL

[THBH92a] Anthony S. Travis, Willem J. Hornix, Robert Bud, and Ernst Homburg. The emergence of research laboratories in the dyestuffs industry, 1870–1900. British Journal for the History of Science, 25 (1):91–111, March 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027006.

Travis:1992:PPI

[THBH92b] Anthony S. Travis, Willem J. Hornix, Robert Bud, and Willem J. Hornix. From process to plant: innovation in the early artificial dye industry. British Journal for the History of Science, 25(1):65–90, March 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027005.

Travis:1992:TSI

[THBM92] Anthony S. Travis, Willem J. Hornix, Robert Bud, and Peter J. T. Morris. The technologyscience interaction: Walter Reppe and cyclooctatetraene chemistry. British Journal for the History of Science, 25(1): 145–167, March 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027009.

Travis:1992:BCI

[THBR92] Anthony S. Travis, Willem J.

Hornix, Robert Bud, and Peter Reed. The British chemical industry and the indigo trade. British Journal for the History of Science, 25(1): 113–125, March 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027007.

Travis:1992:TDS

[THBS92] Anthony S. Travis, Willem J.

Hornix, Robert Bud, and Johan Schot. Technology in decline: a search for useful concepts. British Journal for the History of Science, 25(1): 5–26, March 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027002.

Travis:1992:WMF

[THBvdB92] Anthony S. Travis, Willem J.

Hornix, Robert Bud, and Henk van den Belt. Why monopoly failed: the rise and fall of Société La Fuchsine. British Journal for the History of Science, 25(1):45–63, March 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027004.

Theunissen:1994:KPH

[The94] Bert Theunissen. Knowledge is power: Hugo de Vries on science, heredity and social progress. British Journal for the History of Science, 27(3): 291–311, September 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027600.

Thomas:1987:DTS

[Tho87] Donald E. Thomas. Diesel: technology and society in industrial Germany. University of Alabama Press, Tuscaloosa, AL, USA, 1987. ISBN 0-8173-0295-6. xii + 279 pp. LCCN TJ140.D5 T48 1987.

Timmermann:1998:BRT

[Tim98] Carsten Timmermann. Book review: Thomas H. Broman, The Transformation of German Academic Medicine, 1750– 1820. Cambridge: Cambridge University Press, 1997. Pp. x + 209. ISBN 0-521-55321-1. £35.00, $54.95. British Journal for the History of Science, 31(4):469–487, December 1998. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027888.

Todd:1983:TRA

[Tod83] Alexander R. (Alexander Robertus) Todd. A time to remember: the autobiography of a chemist. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25593-7. viii + 257 pp. LCCN QD22.T74 A35

1983. URL http://www.loc.
gov/catdir/enhancements/
fy1001/83005172-d.html;
http://www.loc.gov/
catdir/enhancements/
fy1001/83005172-t.html.

Tomes:1984:GCT

[Tom84] Nancy Tomes. A generous confidence: Thomas Story Kirkbride and the art of asylumkeeping, 1840–1883. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-24172-3. xv + 387 pp. LCCN RC445.P4 P685 1984.

Topham:1992:SPE

[Top92] Jonathan Topham. Science and popular education in the 1830s: the role of the Bridgewater Treatises. British Journal for the History of Science, 25(4):397–430, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027047.

Topham:1997:NAB

[Top97] Jonathan R. Topham. A note on the archives of the British Society for the History of Science, 1947–1997. British Journal for the History of Science, 30(1):75–76, March 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027902.

Tort:1992:DSF

[Tor92] Patrick Tort, editor. Darwinisme et société. (French) [Darwinism and society]. Presses universitaires de France, Paris, France, 1re édition, 1992. ISBN 2-13-044774-0. 690 pp. LCCN HM106 .D26 1992. 480 FF.

Torrens:1995:MAL

[Tor95] Hugh Torrens. Mary Anning (1799–1847) of Lyme: ‘the greatest fossilist the world ever knew’. British Journal for the History of Science, 28(3):257– 284, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027645.

Teich:1990:FSL

[TP90] Mikuláš Teich and Roy Porter, editors. Fin de siècle and its legacy. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-34108-6 (hardcover), 0-521-34915-X (paperback). xii + 345 pp. LCCN D395 .F55 1990. URL http:

                                        //www.loc.gov/catdir/
                                        description/cam023/
                                        89022347.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam021/89022347.html.

Timms:1992:PCC

[TR92] Edward Timms and Ritchie Robertson, editors. Psychoanalysis in its cultural context, volume 3 of Austrian studies. Edinburgh University Press, Edinburgh, Scotland, 1992. ISBN 0-7486-0359-X. xi + 209 pp. LCCN BF173 .P7756 1992.

Travis:1992:SPC

[Tra92]  Anthony S. Travis.  Sci-

ence’s powerful companion: A. W. Hofmann’s investigation of aniline red and its derivatives. British Journal for the History of Science, 25(1):27– 44, March 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027003.

Travis:1993:RMO

[Tra93] A. S. Travis. The rainbow makers: the origins of the synthetic dyestuffs industry in western Europe. Lehigh University Press, Bethlehem, PA, USA, 1993. ISBN 0-934223-18-1. 335 pp. LCCN HD9660.D843 E858 1993.

Tresch:1997:PMV

[Tre97] John Tresch. ‘the potent magic of verisimilitude’: Edgar Allan Poe within the mechanical age. British Journal for the History of Science, 30(3): 275–290, September 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027862.

Trigger:1989:HAT

[Tri89] Bruce G. Trigger. A history of archaeological thought. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-32878-0, 0-521-33818-2 (paperback). xv + 500 pp. LCCN CC100 .T75 1989. URL http:

//www.loc.gov/catdir/
description/cam023/
88016926.html;   http://
www.loc.gov/catdir/toc/
cam021/88016926.html.

Trombley:1989:SFT

[Tro89] Stephen Trombley. Sir Frederick Treves: the extra-ordinary Edwardian. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1989. ISBN 0-415-03423-X. ix + 218 pp. LCCN RD27.35.T74 T76 1989.

Turner:1987:HRA

[TST87] H. H. (Herbert Hall) Turner, R. A. (Ralph Allen) Sampson, and R. J. (Roger John) Tayler, editors. History of the Royal Astronomical Society. Published for the Society by Blackwell Scientific Pub., Oxford, UK, 1987. ISBN 0-632-02173-X (vol. 1), 0-632-01791-0 (vol. 2), 0-632-02175-6 (paperback: vol. 1), 0-632-01792-9 (paperback: vol. 2). ???? pp. LCCN QB1.R5.

Tarbell:1986:EHO

[TT86]  D.  Stan-

ley Tarbell and Ann Tracy Tarbell. Essays on the history of organic chemistry in the United States, 1875–1955. Folio Publishers, Nashville, Tenn., 1986. ISBN 0-939454-03-3. x + 433 pp. LCCN QD248.5.U6 T37 1986. US$22.00 (est.).

Tunbridge:1992:LKH

[Tun92] Paul Tunbridge. Lord Kelvin, his influence on electrical measurements and units, volume 18 of IEE history of technology series. P. Peregrinus on behalf of the Institution of Electrical Engineers, London, UK, 1992. ISBN 0-86341-237-8. ix + 107 pp. LCCN QC16.K3 T86 1992.

Turner:1983:NCS

[Tur83] Gerard L’Estrange Turner.

Nineteenth-Century scientific instruments. Sotheby Publications, London, UK, 1983. ISBN 0-520-05160-2, 0-85667-170-3. 320 + 20 pp. LCCN Q185 .T86 1983.

Turner:1987:TTC

[Tur87]  Roland Turner,  editor.

Thinkers of the Twentieth Century. St. James Press, Chicago, IL, USA, second edition, 1987. ISBN 0-912289-83-X. vii + 977 pp. LCCN CT120 .T45 1987.

Turtle:1991:BRJ

[Tur91] Alison M. Turtle. Book review:

J. C. Smith (ed.). Historical Foundations of Cognitive Science. Dordrecht: Kluwer, 1990. Pp. xxiii + 303. ISBN 0-7923-045109. British Journal for the History of Science, 24(4): 498–499, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027156.

Turner:1993:BRA

[Tur93a] G. L’E. Turner. Book review: Anita McConnell, Instrument Makers to the World: A History of Cooke, Troughton & Simms. York: William Sessions Ltd, 1992. Pp. xii + 116. ISBN 1-85072-096-7. £16.00. British Journal for the History of Science, 26(2):249, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027113.

Turner:1993:BRR

[Tur93b] Steven Turner. Book review: Regine Zott and Emil Heuser, Die streitbaren Gelehrten. Justus Liebig und die preussischen Universitäten. Kommentierte Edition eines historischen Disputes. Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik,

15. Berlin: ERS Verlag, 1992. Pp. 289. ISBN 3-928577-13-1. DM 51.00. British Journal for the History of Science, 26(3): 368–369, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027418.

Turtle:1993:BRJ

[Tur93c] Alison M. Turtle. Book review: Jaap Van Ginneken, Crowds, Psychology, and Politics, 1871– 1899. Cambridge Studies in the History of Psychology. Cambridge: Cambridge University Press, 1992. Pp. xii + 269.

ISBN 0-521-40418-5. £35.00, $59.95. British Journal for the History of Science, 26(3):372– 373, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027422.

Turner:1994:EMV

[Tur94] R. Steven (Roy Steven) Turner.

In the eye’s mind: vision and the Helmholtz–Hering controversy. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03397-8. xiv + 338 pp. LCCN QP491 .T87 1994. US$49.50. URL http:

//www.loc.gov/catdir/
description/prin031/
93037258.html;   http://
www.loc.gov/catdir/toc/
prin031/93037258.html.

Turkle:1996:LSI

[Tur96] Sherry Turkle. Life on the screen: identity in the age of the Internet. Weidenfeld and Nicolson, London, UK, 1996. ISBN 0-297-81514-8. 347 pp. LCCN QA76.9.C66 T87 1996.

Tselos:1987:GAM

[TW87] George D. Tselos and Colleen Wickey. A guide to archives and manuscript collections in the history of chemistry and chemical technology, volume 7 of Center for History of Chemistry publication. Center for History of Chemistry, Philadelphia, PA, USA, 1987. ISBN 0-941901-05-X. viii + 198 pp. LCCN QD18.U6 T84

                                        1987. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0630/95167507-d.html.

Taton:1989:PAR

[TW95] René Taton and Curtis Wilson, editors. Planetary astronomy from the Renaissance to the rise of astrophysics, volume 2 of The General history of astronomy. Cambridge University Press, Cambridge, UK, 1989–1995. ISBN 0-521-24254-1 (pt. A), 0-521-35168-5 (pt. B). ???? pp. LCCN QB15 .G38 1984 vol. 2. URL http://www.

                                        loc.gov/catdir/
                                        description/cam023/
                                        88025817.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam022/88025817.html.

Tweedale:1990:BRP

[Twe90] Geoffrey Tweedale. Book reviews: Paul E. Ceruzzi. Beyond the Limits: Flight Enters the Computer Age. Cambridge, MA: MIT Press, 1989. Pp. xi + 270. Cloth ISBN 0-262-03143-4. Paper ISBN 0-262-53082-1. No price given. British Journal for the History of Science, 23(2):220–221, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026730.

Tweedale:1991:BRB

[Twe91a] Geoffrey Tweedale. Book review: Brian Oakley and Kenneth Owen. Alvey: Britain’s Strategic Computing Initiative. Cambridge, Mass, and London: The MIT Press, 1989. Pp. xv + 337. ISBN 0-262-15038-7. £24.95. British Journal for the History of Science, 24(1): 116–117, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027032.

Tweedale:1991:BRH

[Twe91b] Geoffrey Tweedale. Book review: Harry M. Collins. Artificial Experts: Social Knowledge and Intelligent Machines. Cambridge, Mass, and London: MIT Press, 1990. Pp. xiii + 266. ISN 0-262-13168-X. £17.95. British Journal for the History of Science, 24(4): 481–482, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027142.

Tweedale:1991:BRM

[Twe91c] Geoffrey Tweedale. Book review: Michael Lindgren. Glory and Failure: The Difference Engines of Johann Müller, Charles Babbage and Georg and Edvard Scheutz, translated by Craig C. McKay. Cambridge, Mass, and London: MIT Press, 1990. Pp. 414. ISBN 0-262-12146-8. £40.50. British Journal for the History of Science, 24(2): 261–263, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4027184.

Tweedale:1991:GIC

[Twe91d] Geoffrey Tweedale. Geology and industrial consultancy: Sir William Boyd Dawkins (1837– 1929) and the Kent Coalfield. British Journal for the History of Science, 24(4):435– 451, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027129.

Tweedale:1992:BRE

[Twe92a] Geoffrey Tweedale. Book review: Emerson W. Pugh, Lyle R. Johnson and John H. Palmer. IBM’s 360 and Early 370 Systems. Cambridge, Mass, and London: MIT Press, 1991. Pp. vii + 819. ISBN 0-262-16123-0. £33.75. British Journal for the History of Science, 25(4):498, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027087.

Tweedale:1992:BRJ

[Twe92b] Geoffrey Tweedale. Book review: John Napier, Rabdology, translated by W. F. Richardson, introduction by R. E. Rider. Charles Babbage Institute Series for the History of Computing, 15. Cambridge, Mass, and London: MIT Press/ Los Angeles and San Francisco:

Tomash Publishers, 1990. Pp. xxxvii + 135. ISBN 0-262-14046-2. £35.95. British Journal for the History of Science, 25(4):462–463, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027054.

Tweedale:1992:BRS

[Twe92c] Geoffrey Tweedale. Book review: Steve J. Heims, The Cybernetics Group. Cambridge, Mass, and London: MIT Press. 1991. Pp. xii + 334. ISBN 0-262-08200-4. £22.50. British Journal for the History of Science, 25(4):497– 498, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027086.

Tweedale:1993:BRD

[Twe93a] Geoffrey Tweedale. Book review: D. J. Bryden, Napier’s Bones: A History and Instruction Manual. London: Harriet Wynter Ltd, 1992. Pp. 24. ISBN 0-9507258-2-X. £12; limited edition, 750 copies. British Journal for the History of Science, 26(1):84, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027348.

Tweedale:1993:BRE

[Twe93b] Geoffrey Tweedale. Book review: Ernst Martin, The Calculating Machines (Die Rechenmaschinen): Their History and Development, translated and edited by Peggy Aldrich Kidwell and Michael R. Williams. Volume 16 in the Charles Babbage Institute Reprint Series for the History of Computing. Cambridge, Mass, and London: MIT Press; Los Angeles and San Francisco: Tomash Publishers, 1992. Pp. xvii + 367, illus. ISBN 0-262-13278-8. £44.95. British Journal for the History of Science, 26(1): 126–127, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027388.

Tweedale:1993:BRW

[Twe93c] Geoffrey Tweedale. Book review: William Poundstone, Prisoner’s Dilemma: John von Neumann, Game Theory, and the Puzzle of the Bomb. Oxford: Oxford University Press, 1992. Pp. xi + 290. ISBN 0-19-286162-X. £7.99 (paperback edition). British Journal for the History of Science, 26(3): 375–376, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027424.

Twyman:1988:SMD

[Twy88] Alan Twyman. In search of the mysterious Doctor Weekes: (a fragment of Sandwich history). Field’s Secretarial Services, Dover, UK, 1988. ISBN 0-9513889-0-8 (paperback). xi + 162 pp. LCCN QC515.W4 T99 1988.

Tytler:1982:PEN

[Tyt82] Graeme Tytler. Physiognomy in the European novel: faces and fortunes. Princeton University Press, Princeton, NJ, USA, 1982. ISBN 0-691-06491-1. xix + 436 pp. LCCN PN3500 .T9. US$25.00.

Ucko:1991:AR

[U^{+}91] Peter J. Ucko et al., editors.

Avebury reconsidered: from the 1660s to the 1990s. Unwin Hyman, London, UK, 1991. ISBN 0-04-445919-X. xiv + 293 pp. LCCN GN776.22.G7 A94 1991. US$100.00.

Uimonen:1995:BRE

[Uim95] Minna Uimonen. Book review:

Elizabeth Lunbeck, The Psychiatric Persuasion: Knowledge, Gender, and Power in Modern America. Princeton: Princeton University Press, 1994. Pp. xiv + 431. ISBN 0-691-04804-5. £25.00, $29.95. British Journal for the History of Science, 28 (3):358–359, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027656.

Unguru:1992:BRG

[Ung92] S. Unguru. Book review: Guy Beaujouan. Par raison de nombres; L’art du calcul et les savoirs scientifiques médiévaux. Aldershot: Variorum, 1991. Pp. XII + 300. ISBN 0-86078-281-6. £47.50. British Journal for the History of Science, 25(4): 460–462, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027052.

VanHelden:1985:MUC

[Van85] Albert Van Helden. Measuring the universe: cosmic dimensions from Aristarchus to Halley. University of Chicago Press, Chicago, IL, USA, 1985. ISBN 0-226-84881-7, 0-226-84882-5 (paperback). viii + 203 pp. LCCN QB15 .V33

1985. URL http://www.loc.
gov/catdir/description/
uchi052/84016397.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/84016397-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/84016397-b.html.

VanHeteren:1991:BRR

[Van91] G. Van Heteren. Book review: Richard L. Kremer (ed.). Letters of Hermann von Helmholtz to his wife 1847–1859 (Boethius-Texte und Abhandlungen zur Geschichte der exakten Wissenschaften, Bd XXIII). Stuttgart: Franz Steiner Verlag, 1990. British Journal for the History of Science, 24 (2):263–264, June 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027185.

VanRiper:1993:MAM

[Van93] A. Bowdoin Van Riper. Men among the mammoths: Victorian science and the discovery of human prehistory. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1993. ISBN 0-226-84991-0 (hardcover), 0-226-84992-9 (paperback). xv + 267 pp. LCCN GN805 .V36

1993. URL http://www.loc.
gov/catdir/description/
uchi052/93017006.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/93017006-t.html.

Voltaire:1980:CWV

[VB80] Voltaire and Theodore Besterman. The complete works of Voltaire = Les Œuvres completes de Voltaire. Voltaire Foundation at the Taylor Institution, Oxford, UK, 1980. ISBN 0-7294-0745-4 (vol. 1A), 0-7294-0774-8 (vol. 1B), 0-7294-0833-7 (vol. 1C), 0-7294-0790-X (vol. 3A), 0-7294-0474-9 (vol. 3B), 0-7294-0507-9 (vol. 4), 0-7294-0567-2 (vol. 5), 0-7294-0359-9 (vol. 8), 0-7294-0646-6 (vol. 9), 0-7294-0325-4 (vol. 10), 0-7294-0360-2 (vol. 14), 0-7294-0374-2 (vol. 15), 0-7294-0794-2 (vol. 16), 0-7294-0421-8 (vol. 17), 0-7294-0893-0 (vol. 18A), 0-7294-0894-9 (vol. 18B), 0-7294-0508-7 (vol. 19), 0-7294-0753-5 (vol. 20A), 0-7294-0754-3 (vol. 20B), 0-7294-0874-4 (vol. 22), 0-7294-0871-X (vol. 28A), 0-7294-0899-X (vol. 28B), 0-7294-0800-0 (vol. 30A), 0-7294-0814-0 (vol. 30B), 0-7294-0828-0 (vol. 30C), 0-7294-0426-9 (vol. 31A), 0-7294-0427-7 (vol. 31B), 0-7294-0842-6 (vol. 32A), 0-7294-0843-4 (vol. 32B), 0-7294-0338-6 (vol. 33), 0-7294-0376-9 (vol. 35), 0-7294-0377-7 (vol. 36), 0-7294-0854-X (vol. 38), 0-7294-0922-8 (vol. 40), 0-7294-0942-2 (vol. 45A), 0-7294-0568-0 (vol. 46– 47), 0-7294-0717-9 (vol. 46), 0-7294-0718-7 (vol. 47), 0-7294-0118-9 (vol. 48), 0-7294-0968-6 (vol. 49B), 0-7294-0334-3 (vol. 50), 0-7294-0729-2 (vol. 56A), 0-7294-0644-X (vol. 56B), 0-7294-0744-6 (vol. 56C), 0-7294-0348-3 (vol. 62), 0-7294-0361-0 (vol. 63A), 0-7294-0896-5 (vol. 63B), 0-7294-0824-8 (vol. 63C), 0-7294-0312-2 (vol. 64), 0-7294-0476-5 (vol. 66), 0-7294-0900-7 (vol. 67), 0-7294-0823-X (vol. 68), 0-7294-0475-7 (vol. 69), 0-7294-0863-9 (vol. 71A), 0-7294-0825-6 (vol. 71B), 0-7294-0813-2 (vol. 73), 0-7294-0719-5 (vol. 74A), 0-7294-0892-2 (vol. 74B), 0-7294-0943-0 (vol. 75A), 0-7294-0982-1 (vol. 75B), 0-7294-0931-7 (vol. 80A), 0-7294-0936-8 (vol. 136), 0-7294-0937-6 (set, vol. 137A-137B), 0-7294-0963-5 (v. 137A), 0-7294-0964-3 (vol. 137B), 0-7294-0938-4 (vol. 138), 0-7294-0777-2 (vol. 141), 0-7294-0887-6 (vol. 142). ???? pp. LCCN PQ2070 1980.

vonBaer:1993:CKE

[vBDGO93] Karl Ernst von Baer, Anton Dohrn, Christiane Groeben, and Jane M. (Jane Marion) Oppenheimer. Correspondence, Karl Ernst von Baer (1792– 1876), Anton Dohrn (1840– 1909), volume 83, part 3 of Transactions of the American Philosophical Society, 0065-9746. American Philosophical Society, Philadelphia, PA, USA, 1993. ISBN 0-87169-833-1. 156 pp. LCCN QH31.B15 A4 1993.

vanderVeer:1991:UVQ

[vdVV91] René van der Veer and Jaan Valsiner. Understanding Vygotsky: a quest for synthesis. Blackwell, Oxford, UK, 1991. ISBN 0-631-16528-2. x + 450 pp. LCCN BF109.V95 V44 1991. US$35.00.

vanderWaerden:1985:HAA

[vdW85] B. L. (Bartel Leendert) van der Waerden. A history of algebra: from al-Khwarizmi to Emmy Noether. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-13610-X (US). xi + 271 pp. LCCN QA151 .W343 1985.

Verwey:1985:PAB

[Ver85] Gerlof Verwey. Psychiatry in an anthropological and biomedical context: philosophical presuppositions and implications of German psychiatry, 1820– 1870, volume 15 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1985. ISBN 90-277-1713-3. xix + 316 pp. LCCN RC450.G3 V4713 1985.

Vernon:1990:BRB

[Ver90] Keith Vernon. Book review: Bruno Latour. The Pasteurization of France. London: Harvard University Press, 1988. Trans Alan Sheridan and John Law. Pp. 273. ISBN 0-674-65760-8. £23.95. British Journal for the History of Science, 23(3): 344–346, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026768.

Verwey:1991:BRG

[Ver91] Gerlof Verwey. Book review: Gunter Mann and Franz Dumont (eds.). Die Natur des Menschen: Probleme der Physischen Anthropologie und Rassenkunde (1750–1850). Stuttgart & New York: Gustav Fischer, 1990. Pp. 457. ISBN 3-437-11271-6. DM 128. British Journal for the History of Science, 24(2):260–261, June 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027183.

Vernon:1993:DSS

[Ver93] Keith Vernon. Desperately seeking status:

Evolutionary Systematics and the taxonomists’ search for respectability 1940–1960. British Journal for the History of Science, 26(2):207–227, June 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027098.

Vernon:1994:BRR

[Ver94] Keith Vernon. Book review:

Robert Bud, The Uses of Life: A History of Biotechnology. Cambridge: Cambridge University Press, 1993. Pp. xvii + 299. ISBN 0-521-38240-8. £12.95, $19.95 (paperback edition). British Journal for the History of Science, 27(4):482– 483, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027635.

Vernon:1995:BRC

[Ver95]  Keith Vernon.  Book re-

view: C. Kenneth Waters and Albert Van Helden (eds.), Julian Huxley. Biologist and Statesman of Science. Proceedings of a Conference held at Rice University 25-27 September 1987. Houston: Rice University Press, 1992. Pp. xii + 344. ISBN 0-89263-314-X. $32.50. British Journal for the History of Science, 28(1): 121–123, March 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027502.

vanGinneken:1992:CPP

[vG92] Jaap van Ginneken. Crowds, psychology, and politics, 1871– 1899. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40418-5 (hardcover). xii + 269 pp. LCCN HM281 .G46

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        cam024/91042797.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam029/
                                        91042797.html.

Vallisnieri:1991:E

[VG98] Antonio Vallisnieri and Dario Generali. Epistolario, volume 9 of Filosofia e scienza nel Cinquecento e nel Seicento; Pubblicazioni del “Centro di studi del pensiero filosofico del Cinquecento e del Seicento in relazione ai problemi della scienza” del Consiglio nazionale delle ricerche. Serie III, Testi. F. Angeli, Milano, Italy, 1991– 1998. ISBN 88-204-6733-X (vol. 1). ???? pp. LCCN QH31.V35 A4 1991. L70000.

vonGuericke:1994:NCM

[vGA94] Otto von Guericke and Margaret Glover Foley Ames. The new (so-called) Magdeburg experiments of Otto von Guericke, volume 137 of Archives internationales d’histoire des idées. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2399-8. xxiv + 394 pp. LCCN Q155 .G9313 1994.

vanHeteren:1992:BRE

[vH92] G. van Heteren. Book review: Échanges d’influences scientifiques et techniques entre pays Européens de 1780 à 1830. Acres du 114ème Congrès National des Sociétés Savantes (Paris 3-9 Avril 1989). Colloques du CTHS, 5. Paris: Editions du CTHS. 1990. Pp. 171. ISBN 2-7355-0208-2. British Journal for the History of Science, 25(2):272–274, June 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027314.

vonHumboldt:1997:CSP

[vH97] Alexander von Humboldt. Cosmos: a sketch of a physical description of the universe. Foundations of natural history. The Johns Hopkins University Press, Baltimore, MD, USA, 1997. ISBN 0-8018-5502-0 (vol. 1 : paperback), 0-8018-5503-9 (vol. 2 : paperback). ???? pp. LCCN QB981 .H8613

1997. URL http://www.loc.
gov/catdir/bios/jhu051/
96036421.html;
http://www.loc.gov/
catdir/description/
jhu052/96036421.html.

vonHelmholtz:1995:SCP

[vHC95] Hermann von Helmholtz and David Cahan. Science and culture: popular and philosophical essays. University of Chicago Press, Chicago, IL, USA, 1995. ISBN 0-226-32658-6, 0-226-32659-4 (paperback). xviii + 418 pp. LCCN Q174.8 .H45 1995. URL http://www.

                                        loc.gov/catdir/
                                        description/uchi051/
                                        95012217.html;   http://
                                        www.loc.gov/catdir/toc/
                                        uchi051/95012217.html.

Vickers:1984:OSM

[Vic84] Brian Vickers, editor. Occult and scientific mentalities in the Renaissance. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25879-0, 0-521-33836-0 (paperback). xiv + 408 pp. LCCN H7.5VIC; BF1429 OCC 1984.

Vickers:1987:ESB

[Vic87] Brian Vickers, editor. English science, Bacon to Newton. Cambridge English prose texts. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-30408-3, 0-521-31683-9 (paperback). xii + 244 pp. LCCN Q127.G4 E54

                                        1987. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0642/86019265-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0642/86019265-t.html.

Vickers:1991:STC

[Vic91] Brian Vickers. A scholarly tradition continued. British Journal for the History of Science, 24 (3):343–351, September 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027233.

Viete:1983:AAN

[Viè83] François Viète. The analytic art: nine studies in algebra, geometry, and trigonometry from the Opus restitutae mathematicae analyseos, seu, Algebrâ novâ. Kent State University Press, Kent, OH, USA, 1983. ISBN 0-87338-282-X. 450 pp. LCCN QA33 .V4813 1983.

Vignoles:1982:CBV

[Vig82] Keith H. Vignoles. Charles Blacker Vignoles, romantic engineer. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-23930-3 (US). xii + 187 pp. LCCN TA140.V6 V59 1982.

vonLeibniz:1976:MNT

[vL76] Freiherr Gottfried Wilhelm von Leibniz. Mathematischer, naturwissenschaftlicher und technischer Briefwechsel. (German) [Mathematical, scientific, and technical correspondence]. Sämtliche Schriften und Briefe / Gottfried Wilhelm Leibniz. Akademie-Verlag, Berlin, Germany, 1976. ISBN 3-05-000235-2 (vol. 2), 3-05-000766-4 (Vol. 3), 3-05-002602-2 (vol. 4), 3-05-003471-8 (vol. 5), 3-05-004116-1 (vol. 6). ???? pp. LCCN ????

vonLeibniz:1988:DMR

[vLMB88] Freiherr Gottfried Wilhelm von Leibniz, R. N. D. (R. Niall D.) Martin, and Stuart C. Brown. Discourse on metaphysics and related writings. Classics of philosophy and science series. Manchester University Press, Manchester, UK, 1988. ISBN 0-7190-2338-6. ix + 170 pp. LCCN B2572.E5 M37 1988. US$25.00 (US$35.00 U.S.).

vanLeeuwenhoek:1982:AVL

[vLP82] Antoni van Leeuwenhoek and L. C. Palm, editors. Antoni van Leeuwenhoeck, sixteen hundred and thirty-two 1632– 1723: studies on the life and work of the Delft scientist commemorating the 350th anniversary of his birthday, volume 8 of Nieuwe Nederlandse bijdragen tot de geschiedenis der geneeskunde en der natuurwetenschappen. Rodopi, Amsterdam and Leiden, The Netherlands, 1982. ISBN 90-6203-824-7. 209 pp. LCCN ????

vonLiebig:1988:JLA

[vLvHH^{+}88] Freiherr Justus von Liebig, August Wilhelm von Hofmann, Emil Heuser, Regine Zott, and Emil Erlenmeyer. Justus von Liebig und August Wilhelm Hofmann in ihren Briefen: Nachträge 1845–1869. Justus von Liebig und Emil Erlenmeyer in ihren Briefen von 1861–1872. (German) [Justus von Liebig and August Wilhelm Hofmann in their Letters: Appendix 1845–1869. Justus von Liebig und Emil Erlenmeyer in their letters of 1861–1872]. Bionomica-Verlag, Mannheim, Germany, 1988. ISBN 3-88208-012-4. 54 + 32 pp. LCCN QD22.L7 A4 1988.

VonStaden:1989:HAM

[VoC89]  Heinrich  Von

Staden and Herophilus of Chalcedon. Herophilus: the art of medicine in early Alexandria: edition, translation and essays. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-23646-0. xliii + 666 pp. LCCN R126.H3/; 0120; N.i.224; C.Gr.H.313.

Voskuhl:1997:RHA

[Vos97]  Adelheid Voskuhl.  Recre-

ating Herschel’s actinometry: An essay in the historiography of experimental practice. British Journal for the History of Science, 30(3):337–355, September 1997. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027866.

vonPufendorf:1996:GWG

[vPSBD^{+}04] Samuel Freiherr von Pufendorf, Wilhelm Schmidt-Biggemann, Detlef Döring, Gerald Hartung, and Frank Böhling. Gesammelte Werke. (German) [Collected works]. Akademie-Verlag, Berlin, Germany, 1996– 2004. ISBN 3-05-002887-4 (set), 3-05-001920-4 (vol. 1), 3-05-002880-7 (vol. 2), 3-05-003441-6 (vol. 3), 3-05-003181-6 (vol. 4), 3-05-003605-2 (vol. 5). ???? pp. LCCN B2603 .A3 1996.

vonProschwitz:1990:BCE

[vPvP90] Gunnar von Proschwitz and Mavis von Proschwitz, editors. Beaumarchais et le Courier de l’Europe: documents inédits ou peu connus. (French) [Beaumarchais and the Courier of Europe: unknown or little-known documents], volume 273–274 of Studies on Voltaire and the eighteenth century, 0435-2866. Voltaire Foundation, Oxford, UK, 1990. ISBN 0-7294-0397-1 (set), 0-7294-0398-X (vol. 1), 0-7294-0399-8 (vol. 2). xxvii + 1289 pp. LCCN PQ2105.A2 S8 vol. 273–274.

vanRuler:1995:CCV

[vR95] J. A. van Ruler. The crisis of causality: Voetius and Descartes on God, nature, and change, volume 66 of Brill’s studies in intellectual history, 0920-8607. E. J. Brill, Leiden, The Netherlands, 1995. ISBN 90-04-10371-6. xii + 353 pp. LCCN B4000.V64 R85 1995.

Vucinich:1984:EKA

[Vuc84] Alexander Vucinich. Empire of knowledge: the Academy of Sciences of the USSR (1917– 1970). University of California Press, Berkeley, CA, USA, 1984. ISBN 0-520-04871-7. x + 484 pp. LCCN AS262.A68 V79 1984.

Wainwright:1990:MCS

[Wai90] Milton Wainwright. Miracle cure: the story of penicillin and the golden age of antibiotics. Blackwell, Oxford, UK, 1990. ISBN 0-631-16492-8. xi + 196 pp. LCCN RM267 .W24 1990. US$30.00 (U.S. : est.).

Walker:1978:K

[Wal78]  Ralph  Charles  Suther-

land Walker. Kant. Arguments of the philosophers. Routledge and K. Paul, London, UK, 1978. ISBN 0-7100-0009-X (paperback), 0-7100-8994-5. xii + 201 pp. LCCN B2798.

Wallace:1981:PGE

[Wal81] William A. Wallace. Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought, volume 62 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1981. ISBN 90-277-1215-8, 90-277-1216-6 (paperback), 94-009-8404-9. ISSN 0068-0346. xvi + 369 pp. LCCN Q174 .B67

vol. 62. URL https://link.
springer.com/book/10.
1007/978-94-009-8404-2.

Wallace:1984:GHS

[Wal84] William A. Wallace. Galileo and his sources: the heritage of the collegio romano in Galileo’s science. Princeton University Press, Princeton, NJ, USA, 1984. ISBN 0-691-08355-X. xiv + 371 pp. LCCN QB36.G2 W35 1984. US$42.50.

Walker:1988:SJE

[Wal88] Margot Walker. Sir James Edward Smith, M.D., F.R.S., P.L.S., 1759–1828: first president of the Linnean Society of London. Linnean Society of London, London, UK, 1988. ISBN 0-9506207-1-8. 60 pp. LCCN QK31.S64 W34 1988.

Walter:1990:SCC

[Wal90] Maila L. Walter. Science and cultural crisis: an intellectual biography of Percy Williams Bridgman (1882–1961). Stanford University Press, Stanford, CA, USA, 1990. ISBN 0-8047-1796-6. x + 362 pp. LCCN QC16.B73 W35 1990. US$39.50

(est.). URL http://www.loc.
gov/catdir/toc/cam023/
90030511.html.

Wali:1991:CBC

[Wal91] K. C. (Kameshwar C.) Wali. Chandra: a biography of S. Chandrasekhar. University of Chicago Press, Chicago, IL, USA, 1991. ISBN 0-226-87054-5, 0-226-87055-3 (paperback). x + 341 + 32 pp. LCCN QB36.C46 W35

1991. URL http://www.loc.
gov/catdir/description/
uchi052/90010845.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/90010845-t.html;

http://www.loc.gov/
catdir/enhancements/
fy0609/90010845-b.html.

Wallace:1992:GLD

[Wal92] William A. Wallace. Galileo’s logic of discovery and proof: the background, content, and use of his appropriated treatises on Aristotle’s Posterior analytics, volume 137 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1577-4, 0-7923-1579-0 (set), 94-015-8040-5. ISSN 0068-0346. xxiii + 323 pp. LCCN Q174 .B67

vol. 137. URL https://link.
springer.com/book/10.
1007/978-94-015-8040-3.

Warner:1986:TPM

[War86] John Harley Warner. The therapeutic perspective: medical practice, knowledge, and identity in America, 1820– 1885. Harvard University Press, Cambridge, MA, USA, 1986. ISBN 0-674-88330-6. x + 367 pp. LCCN RM47.U6 W37 1986.

Ward:1990:BRR

[War90a] Jenny Ward. Book review:

Rafael Huertas. Orfila, Saber y Poder Médico. Madrid: Consejo Superior de Investigaciones Cientificas, 1988. Pp. v + 435. ISBN 84-00-06745-2. No price given. British Journal for the History of Science, 23(2): 220, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026729.

Wardy:1990:CCS

[War90b] Robert Wardy. The chain of change: a study of Aristotle’s Physics VII. Cambridge classical studies. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37327-1. x + 345 pp. LCCN Q151 .W36

                                        1990. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/89009804.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam023/
                                        89009804.html.

Warner:1990:WSI

[War90c] Deborah Jean Warner. What is a scientific instrument, when did it become one, and why? British Journal for the History of Science, 23(1):83–93, March 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

                                        URL http://www.jstor.org/
                                        stable/4026803.

Warboys:1991:BRD

[War91a] Michael Warboys. Book review: David I. Grove. A History of Human Helminthology. Wallingford, Oxon: CAB International, 1990. Pp. viii + 848. ISBN 0-85198-689-7. £55.00, $96.25. British Journal for the History of Science, 24(4): 499, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027157.

Warner:1991:BRC

[War91b] John Harley Warner. Book review: Chandos Michael Brown, Benjamin Silliman: A Life in the Young Republic. Princeton: Princeton University Press, 1989. Pp. xvi + 377. ISBN 0-691-08533-1. $29.95. British Journal for the History of Science, 24(1): 104–105, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027023.

Warwick:1995:BRB

[War95a] Andrew Warwick. Book review:

The collected papers of Albert Einstein, vol 3, The Swiss years: Writings, 1909–1911 — M. J. Klein, A. J. Kox, J. Renn, R. Schulmann. British Journal for the History of Science, 28(4):480–482, December 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027563.

Warwick:1995:BRM

[War95b] Andrew Warwick. Book review:

Mara Beller, Juergen Renn and Robert Schulmann (eds.), Einstein in Context, Science in Context, vol. 6, part 1. Cambridge: Cambridge University Press, 1993. Pp. 368. ISSN 0-691-10250-3. £25.00, $43.00 (single part). British Journal for the History of Science, 28 (3):364–368, September 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027661.

Warwick:1996:BRM

[War96a] Andrew Warwick. Book review: Martin J. Klein, Anne J. Kox, Juergen Renn and Robert Schulmann (eds.), The Collected Papers of Albert Einstein. Volume 5: The Swiss Years: Correspondence, 1902– 1914. English translation by Anna Beck and Don Howard. Princeton: Princeton University Press, 1995. Pp. xvii + 384. ISBN 0-691-00099-9. £29.95, $19.50. British Journal for the History of Science, 29(1): 109–110, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027530.

Warwick:1996:BRR

[War96b] Andrew Warwick. Book review: R. H. Dalitz (ed.), The Collected Works of P. A. M. Dirac, 1924–48. Cambridge: Cambridge University Press, 1995. Pp. xxiv + 1310. ISBN 0-521-36231-8. £175.00, $250.00. British Journal for the History of Science, 29(4): 488–489, December 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027707.

Watson:1991:BRT

[Wat91] Helen Watson. Book review:

Thomas Crump. The Anthropology of Numbers. Cambridge Studies in Social Anthropology, No. 70. Cambridge: Cambridge University Press, 1990. Pp. x + 201. ISBN 0-521-38045-6. £30.00, $49.50. British Journal for the History of Science, 24(1):121, March 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027036.

Wayman:1987:DOB

[Way87] Patrick A. Wayman. Dunsink Observatory, 1785–1985: a bicentennial history, volume 7 of Historical studies in Irish science and technology. Dublin Institute for Advanced Studies, Dublin, Ireland, 1987. ISBN 0-86027-020-3. xiii + 353 pp. LCCN QB82.I732 D869 1987. US$25.00.

Womersley:1997:EGB

[WBP97]  David  Womer-

sley, J. W. (John Wyon) Burrow, and J. G. A. (John Greville Agard) Pocock, editors. Edward Gibbon: bicentenary essays, volume 355 of Studies on Voltaire and the eighteenth century, 0435-2866. Voltaire Foundation, Oxford, UK, 1997. ISBN 0-7294-0552-4. xii + 429 pp. LCCN PQ2105.A2 S8 vol.

Williams:1982:SHT

[WD82]  Trevor  Illtyd  Williams

and T. K. (Thomas Kingston) Derry. A short history of Twentieth-Century technology c. 1900–c. 1950. Clarendon Press, New York, NY, USA, 1982. ISBN 0-19-858159-9. xix + 411 pp. LCCN T20 .W55 1982. US$25.00. URL http:

                                        //www.loc.gov/catdir/
                                        enhancements/fy0604/
                                        82004362-d.html.

Weart:1988:NFH

[Wea88] Spencer R. Weart. Nuclear fear: a history of images. Harvard University Press, Cambridge, MA, USA, 1988. ISBN 0-674-62835-7. 535 pp. LCCN QC773 .W43 1988.

Weatherall:1990:SCH

[Wea90] M. (Miles) Weatherall. In search of a cure: a history of pharmaceutical discovery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-261747-8 (hardcover). xiv + 298 + 12 pp. LCCN RM45 .W43 1990. US$37.50. URL http://www.

                                        loc.gov/catdir/
                                        enhancements/fy0636/
                                        90007426-d.html;
                                        http://www.loc.gov/
                                        catdir/enhancements/
                                        fy0636/90007426-t.html.

Wear:1991:BRG

[Wea91] Andrew Wear. Book review: G. R. Dunstan (ed.). The Human Embryo. Aristotle and the Arabic and European Traditions. Exeter: University of Exeter Press, 1990. Pp. xi + 235. ISBN 0-85989-340-5. £25.00. British Journal for the History of Science, 24(4): 487–488, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027147.

Webster:1988:PHC

[Web88] Charles Webster. Problems of health care: the National Health Service before 1957, volume 1 of The Health services since the war; Peacetime history. Her Majesty’s Stationary Office, London, England, 1988. ISBN 0-11-630942-3. x + 479 pp. LCCN RA412.5.G7 W43 1988.

Webster:1998:NHS

[Web98] Charles Webster. The National Health Service: a political history. Oxford paperbacks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-289296-7. xiii + 241 pp. LCCN RA413.5.G7 W43 1998.

Weiss:1982:MTM

[Wei82] John Hubbel Weiss. The making of technological man: the social origins of French engineering education. MIT Press, Cambridge, MA, USA, 1982. ISBN 0-262-23112-3. xviii + 377 pp. LCCN T121 .W44.

Weisz:1983:EMU

[Wei83] George Weisz. The emergence of modern universities in France, 1863–1914. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-05375-8. xii + 397 pp. LCCN LA698 .W44 1983. US$35.00.

Weindling:1990:BRP

[Wei90] Paul Weindling. Book review: Peter Weingart, Jürgen Kroll & Kurt Bayertz. Rasse, Blut und Gene. Geschichte der Eugenik und Rassenhygiene in Deutschland. Frankfurt am Main: Suhrkamp Verlag, 1988. Pp. 746. ISBN 3-518-57887-1. DM48 (paperback). British Journal for the History of Science, 23(2):250–251, June 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026750.

Weindling:1991:DSD

[Wei91] Paul Weindling. Darwinism and social Darwinism in imperial Germany: the contribution of the cell biologist, Oscar Hertwig, 1849–1922, volume 3 of Forschungen zur neueren Medizin- und Biologiegeschichte, 0176-330X. G. Fischer, Stuttgart, Germany, 1991. ISBN 3-437-11305-4. 355 pp. LCCN QH31.H38 W38 1991. DM148.00.

Weintraub:1992:THGa

[Wei92] E. Roy (Eliot Roy) Weintraub, editor. Towards a history of game theory: annual supplement to volume 24, History of political economy, volume 24 of History of political economy. Duke University Press, Durham, NC, USA, 1992. ISBN 0-8223-1253-0. ISSN 0018-2702. vi + 306 pp. LCCN HB1 .H55 suppl. vol. 24 1992; HB144 .T68 1992.

Weindling:1994:BRU

[Wei94] Paul Weindling. Book review: Ulfried Geuter, The Professionalization of Psychology in Nazi Germany, translated by Richard J. Holmes. Cambridge: Cambridge University Press, 1992. Pp. xxi + 335. ISBN 0-521-33297-4. £35, $54.95. British Journal for the History of Science, 27 (2):233–234, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027447.

Welsh:1993:NSS

[Wel93] Ian Welsh. The NIMBY syndrome: its significance in the history of the nuclear debate in Britain. British Journal for the History of Science, 26 (1):15–32, March 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027338.

West:1988:IRT

[Wes88] Graham West. Innovation and the rise of the tunnelling industry. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-33512-4. xv + 355 pp. LCCN TA803 .W47

                                        1988. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/87008089.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam028/
                                        87008089.html.

Westfall:1989:ETG

[Wes89] Richard S. Westfall. Essays on the trial of Galileo, volume 1, no. 5 of Vatican Observatory publications. Special series. Studi Galileiani. Vatican Observatory, Vatican City State, 1989. ISBN 0-268-00923-6. xi + 103 pp. LCCN QB36.G2 W47 1989.

Wess:1995:LDR

[Wes95] Jane Wess. Lecture demonstrations and the real world: the case of cart-wheels. British Journal for the History of Science, 28(1):79–90, March 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027485.

Weyer:1992:GWI

[Wey92] Jost Weyer. Graf Wolfgang II von Hohenlohe und die Alchemie: alchemistische Studien in Schloss Weikersheim, 1587–1610. (German) [Count Wolfgang II of Hohenlohe and Alchemy: alchemical studies in Weikersheim Castle], volume 39 of Forschungen aus Württembergisch Franken. J.

Thorbecke, Sigmaringen, Germany, 1992. ISBN 3-7995-7639-8 (J. Thorbecke), 3-921429-39-0 (Historischer Verein für Württembergisch Franken). 516 pp. LCCN QD24.H64 W49 1992.

Whiffen:1991:RSC

[WH91] D. H. (David Hardy) Whiffen and D. H. (Donald Holroyde) Hey. The Royal Society of Chemistry: the first 150 years. Royal Society of Chemistry, London, UK, 1991. ISBN 0-85186-294-2. viii + 266 pp. LCCN M91.E02912.

Whitrow:1988:THE

[Whi88] G. J. Whitrow. Time in history: the evolution of our general awareness of time and temporal perspective. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-215361-7. x + 217 pp. LCCN D21.3 .W47 1988. US$22.00. URL http://www.

loc.gov/catdir/
enhancements/fy0640/
88004245-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0726/88004245-b.html.

Whiteside:1990:BRW

[Whi90a] D. T. Whiteside. Book review: Wilbur R. Knorr. The Ancient Tradition of Geometric Problems. Boston, Basel & Stuttgart: Birkhäuser, 1986. Pp. ix + 381 + 30 (indexes). £57.75. ISBN (US) 0-8176-3148-8/Europe 3-7643-3148-8.

$69.00. British Journal for the History of Science, 23(3):373– 375, September 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026791.

Whitney:1990:PRM

[Whi90b]  Elspeth  Whitney.

Paradise restored: the mechanical arts from Antiquity through the Thirteenth Century, volume 80:1 of Transactions of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1990. ISBN 0-87169-801-3. ISSN 0065-9746. 169 pp. LCCN T15 .W57 1990; Q11 .P6 n.s. vol. 80 pt.1.

Whitrow:1990:BRD

[Whi90c] G. J. Whitrow. Book review: Dominic J. O’Meara. Pythagoras Revived: Mathematics and Philosophy in Late Antiquity. Oxford: Clarendon Press, 1989. Pp. viii + 251. ISBN 0-19-824485-1. £27.50. British Journal for the History of Science, 23(2):247–248, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4026747.

Whitrow:1990:BRH

[Whi90d] G. J. Whitrow. Book review: E. Harrison. Darkness at Night, a Riddle of the Universe. Cambridge, Mass, and London, Harvard University Press:

1987. Pp. 293. ISBN 0-674-19274-0. £19.95 cloth, £10.25 paperback. British Journal for the History of Science, 23 (2):241–242, June 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026743.

Whitrow:1993:JWJ

[Whi93] Magda Whitrow. Julius von Wagner-Jauregg: (1857–1940). Smith-Gordon and Co., London, UK, 1993. ISBN 1-85463-012-1. xxiv + 221 + 8 pp. LCCN RC339.52.W34 W45

1993. URL http://catalog.
hathitrust.org/api/
volumes/oclc/29571664.
html.

White:1997:BRK

[Whi97] Paul S. White. Book review:

Karen Wonders, Habitat Dioramas: Illusions of Wilderness in Museums of Natural History. Uppsala: Acta Universitatis Upsaliensis, 1993. Pp.

262. ISBN 91-554-3157-7. No price given. British Journal for the History of Science, 30 (2):233–249, June 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027730.

Whorton:1982:CFH

[Who82] James C. Whorton. Crusaders for fitness: the history of American health reformers. Princeton University Press, Princeton, NJ, USA, 1982. ISBN 0-691-04694-8. 359 + 4 pp. LCCN RA418.3.U6 W5 1982. US$19.50.

Whiteside:1992:IDT

[WHS92] D. T. (Derek Thomas) Whiteside, P. M. (Peter Michael) Harman, and Alan E. (Alan Elihu) Shapiro, editors. The investigation of difficult things: essays on Newton and the history of the exact sciences in honour of D. T. Whiteside. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-37435-9. xvi + 531 pp. LCCN QC16.N7 I58

                                        1992. URL http://www.loc.
                                        gov/catdir/description/
                                        cam025/92248567.html;
                                        http://www.loc.gov/
                                        catdir/samples/cam031/
                                        92248567.html;   http://
                                        www.loc.gov/catdir/toc/
                                        cam028/92248567.html.

Wiener:1993:ICF

[Wie93] Norbert Wiener. Invention: the care and feeding of ideas. MIT Press, Cambridge, MA, USA, 1993. ISBN 0-262-23167-0. xxiv + 159 pp. LCCN T15 .W65 1993.

Williams:1981:HBG

[Wil81] Trevor Illtyd Williams. A history of the British gas industry. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1981. ISBN 0-19-858157-2. xvii + 304 + 16 pp. LCCN HD9581.G72 W54.

Wilson:1982:D

[Wil82]  Margaret  Dauler

Wilson. Descartes. The Arguments of the philosophers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1982. ISBN 0-7100-9208-3 (paperback). xvii + 255 pp. LCCN B1875 .W58 1982.

Williams:1984:HFP

[Wil84] Trevor Illtyd Williams. Howard Florey, penicillin and after. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1984. ISBN 0-19-858173-4 (US). ???? pp. LCCN RB17.F57 W55 1984. US$22.95. URL http://www.

loc.gov/catdir/
enhancements/fy0635/
84007845-t.html.

Wilcox:1987:MTP

[Wil87a] Donald J. Wilcox. The measure of times past: pre-Newtonian chronologies and the rhetoric of relative time. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-89721-4, 0-226-89722-2 (paperback). ix + 292 pp. LCCN D11 .W59 1987. US$38.00, US$14.00.

URL http://www.loc.gov/
catdir/enhancements/
fy0608/87001669-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/87001669-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/87001669-b.html.

Williams:1987:KFA

[Wil87b]  Robert  Chad-

well Williams. Klaus Fuchs, atom spy. Harvard University Press, Cambridge, MA, USA, 1987. ISBN 0-674-50507-7. vi + 267 + 10 pp. LCCN UB271.R9 F838 1987.

Williamson:1987:MP

[Wil87c] Rajkumari Williamson, editor. The making of physicists. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-524-9. viii + 200 pp. LCCN QC47.E85 M35

                                        1987. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0745/87204779-d.html.

Wilson:1987:KSC

[Wil87d] David B. Wilson. Kelvin and Stokes: a comparative study in Victorian physics. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-526-5. xvi + 253 pp. LCCN QC16.K3 W55

                                        1987. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0745/87181742-d.html.

Williams:1990:BRK

[Wil90a] Mari E. W. Williams. Book reviews: Kevin Krisciunas, Astronomical Centres of the World. Cambridge: Cambridge University Press 1988. Pp. x + 320 ISBN 0-521-30278-1 £17.50. Patrick A. Wayman, Dunsink Observatory, 1785– 1985: A Bicentennial History. Dublin: Dublin Institute for Advanced Studies and Royal Dublin Society, 1987. Pp. xiii + 353. ISBN 0-86027-020-3. IR £25. British Journal for the History of Science, 23(1): 102–103, March 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4026808.

Williams:1990:NUE

[Wil90b] Rosalind H. Williams. Notes on the underground: an essay on technology, society, and the imagination. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-23145-X. xi + 265 pp. LCCN TA712 .W55 1990.

Williams:1990:RRC

[Wil90c] Trevor Illtyd Williams. Robert Robinson, chemist extraordinary. Clarendon Press, New York, NY, USA, 1990. ISBN 0-19-858180-7. viii + 201 + 7 pp. LCCN QD22.R63 W54 1990. US$37.00. URL http:

//www.loc.gov/catdir/
enhancements/fy0635/
89022123-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0635/89022123-t.html.

Wilkie:1991:BSP

[Wil91a]  Tom Wilkie.  British sci-

ence and politics since 1945. Making contemporary Britain. Blackwell, Oxford, UK, 1991. ISBN 0-631-16849-4, 0-631-16851-6 (paperback). vii + 142 pp. LCCN Q127.G4 W62 1991. US$39.95, US$18.95.

Williams:1991:BRR

[Wil91b] M. E. W. Williams. Book review: R. Taton and C. Wilson (eds.). Planetary Astronomy from the Renaissance to the Rise of Astrophysics. Part A: Tycho Brahe to Newton. The General History of Astronomy, Vol. 2. General Editor, Michael Hoskin. Cambridge: Cambridge University Press, 1989. Pp. X + 274 + xiv. ISBN 0-521-24254-1. £27.50, $49.50. British Journal for the History of Science, 24(3):377– 379, September 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027241.

Williams:1991:BRM

[Wil91c] Mari E. W. Williams. Book review: Martin Campbell-Kelly. ICL: A Business and Technical History. Oxford: Clarendon Press, 1989. Pp. xii + 409. ISBN 0-19-853918-5. £30.00. British Journal for the History of Science, 24(4):480– 481, December 1991. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027141.

Wilson:1991:EDS

[Wil91d] Fred Wilson. Empiricism and Darwin’s science, volume 47 of The University of Western Ontario series in philosophy of science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1019-5 (hardcover). xiv + 358 pp. LCCN QH331 .W53 1991.

Willmoth:1993:SJM

[Wil93] Frances Willmoth. Sir Jonas Moore: practical mathematics and restoration science. Boydell Press, Woodbridge, Suffolk, UK, 1993. ISBN 0-85115-321-6. xi + 244 pp. LCCN QA29.M65 W55 1993. US$35 (US$70.00).

Williams:1994:BRR

[Wil94a] Mari Williams. Book review:

R. P. T. Davenport-Hines and Judy Slinn, Glaxo: A History to 1962. Cambridge: Cambridge University Press, 1992. Pp. xiv + 405. ISBN 0-521-41539-X. £55.00. British Journal for the History of Science, 27(2):240–241, June 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027454.

Williams:1994:PMH

[Wil94b] Mari E. W. Williams. The precision makers: a history of the instruments industry in Britain and France, 1870–1939. Comparative and international business series. Modern histories. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1994. ISBN 0-415-03732-8. viii + 216 pp. LCCN HD9706.6.G72 W55 1994.

Wilson:1994:MMM

[Wil94c] Adrian Wilson. The making of man-midwifery: childbirth in England, 1660–1770. UCL Press., London, UK, 1994. ISBN 1-85728-292-2. xii + 239 pp. LCCN RG518.G7. URL http://

catalog.hathitrust.org/
api/volumes/oclc/
32910640.html.

Williams:1995:BRJ

[Wil95a] Mari E. W. Williams. Book review: John North. The Fontana History of Astronomy and Cosmology. London: Fontana, 1994. Pp. xxvii + 697. ISBN 0-00-686177-6. £12.99. British Journal for the History of Science, 28(2):233–234, June 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027679.

Wilson:1995:IWE

[Wil95b] Catherine Wilson. The invisible world: early modern philosophy and the invention of the microscope. Studies in intellectual history and the history of philosophy. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-03418-4. x + 280 pp. LCCN B67 .W55 1995.

URL http://www.h-net.org/
review/hrev-a0a0i9-aa;
http://www.loc.gov/
catdir/description/
prin031/94024556.html;
http://www.loc.gov/

catdir/toc/prin031/
94024556.html.

Williams:1997:BRR

[Wil97] Mari Williams. Book review:

Rene Taton and Curtis Wilson (eds.), The General History of Astronomy, Volume 2. Planetary Astronomy from the Renaissance to the Rise of Astrophysics. Part B: The Eighteenth and Nineteenth Centuries. Cambridge: Cambridge University Press, 1995. Pp. xiii + 281. ISBN 0-521-35168-5. £45.00, $69.95. British Journal for the History of Science, 30 (1):101–121, March 1997. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027914.

Winsor:1991:RSN

[Win91] Mary P. Winsor. Reading the shape of nature: comparative zoology at the Agassiz Museum. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1991. ISBN 0-226-90214-5 (hardcover), 0-226-90215-3 (paperback). xviii + 324 pp. LCCN QL71.U62 C358

1991. URL http://www.loc.
gov/catdir/enhancements/
fy0608/91008742-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0608/91008742-t.html;
http://www.loc.gov/
catdir/enhancements/
fy0609/91008742-b.html.

Winterton:1993:MST

[Win93]  Jonathan Winterton.  The

1984–1985 miners’ strike and technological change. British Journal for the History of Science, 26(1):5–14, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027337.

Wise:1985:WRW

[Wis85] George Wise. Willis R. Whitney, General Electric, and the origins of U.S. industrial research. Columbia University Press, New York, NY, USA, 1985. ISBN 0-231-06044-0, 0-231-06016-5. 375 pp. LCCN T40.W45 W57 1985.

Wise:1996:BRT

[Wis96] M. Norton Wise. Book review: Theodore M. Porter, Trust in Numbers: The Pursuit of Objectivity in Science and Public Life. Princeton: Princeton University Press, 1995. Pp. xiv + 310. ISBN 0-691-03776-0. £19.95, $24.95. British Journal for the History of Science, 29(2):246–247, June 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027848.

Withers:1995:BRG

[Wit95a] Charles W. J. Withers. Book review: Gunnar Eriksson. The Atlantic Vision: Olaus Rudbeck and Baroque Science. Uppsala Studies in History of Science, 19. Canton, MA: Science History Publications, 1994. Pp. viii + 196. ISBN 0-88135-158-X. $27.95. British Journal for the History of Science, 28(3): 351–353, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027650.

Withers:1995:RU

[Wit95b] Charles W. J. Withers. Review: [untitled]. British Journal for the History of Science, 28(4):473–474, December 1995. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027559.

Weingart:1988:RBG

[WKB88] Peter Weingart, Jürgen Kroll, and Kurt Bayertz. Rasse, Blut und Gene: Geschichte der Eugenik und Rassenhygiene in Deutschland. (German) [Race, blood and genes: History of eugenics and racial cleanliness in Germany]. Wissenschaftsforschung. Suhrkamp, Frankfurt am Main, Germany, 1988. ISBN 3-518-57887-1, 3-518-57886-3. 746 pp. LCCN

???? URL http://www.gbv.
de/dms/bs/toc/025182471.

Wiener:1985:CSS

[WM85] Norbert Wiener and Pesi Rustom Masani. Cybernetics, science, and society: Ethics, aesthetics, and literary criticism: Book reviews and obituaries, volume 23; vol. 4 of Mathematicians of our time; Collected works with commentaries / Norbert Wiener; Collected works with commentaries. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-23123-9. xx + 1083 pp. LCCN QA3 .W53 vol. 4.

Wolters:1987:MMI

[Wol87]  Gereon Wolters.  Mach I,

Mach II, Einstein und die Relativitätstheorie: eine Fälschung und ihre Folgen. (German) [Mach I, Mach II, Einstein and their consequences]. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 1987. ISBN 3-11-010825-9. xv + 474 s pp. LCCN ???? URL 01;

                                         http://www.zentralblatt-
                                        math.org/zmath/en/search/
                                        ?an=0643.01029.

Wonders:1993:HDI

[Won93] Karen Wonders. Habitat dioramas: illusions of wilderness in museums of natural history, volume nova ser. 25 of Acta Universitatis Upsaliensis. Figura, 0071-481X. Acta Universitatis Upsaliensis, Uppsala, Sweden, 1993. ISBN 91-554-3157-7. 262 pp. LCCN N25 .F5 no. 25.

Wood:1992:HIB

[Woo92] James Wood. History of international broadcasting, volume 19< 23 > of IEE history of technology series. P. Peregrinus Ltd. in association with the Science Museum, London, UK, 1992. ISBN 0-86341-281-5, 0-85296-920-1 (vol. 2). ???? pp. LCCN HE8697.4 .W66

1992. URL http://www.loc.
gov/catdir/enhancements/
fy0631/93171088-d.html;
http://www.loc.gov/
catdir/enhancements/
fy0631/93171088-t.html.

Wood:1995:BRA

[Woo95] Paul Wood. Book review: Albert V. Carozzi and Gerda Bouvier, The Scientific Library of Horace-Bénédict de Saussure (1797): Annotated Catalog of an 18th-Century Bibliographic and Historic Treasure. Mémoires de la Société de Physique et d’Histoire naturelle de Genève, 46. Genève: Société de Physique et d’Histoire naturelle, 1994. Pp. x + 201. ISSN 0252-7960. $40 (US). British Journal for the History of Science, 28 (3):353, September 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027651.

Wood:1998:BRP

[Woo98] Alan Wood. Book review: Paul R. Josephson, New Atlantis Revisited: Akademgorodok, the Siberian City of Science. Princeton: Princeton University Press, 1997. Pp. xxii + 351. ISBN 0-691-04454-6. £27.50, $39.50. British Journal for the History of Science, 31(4):469–487, December 1998.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027889.

Wheaton:1992:ISH

[WR92]  Bruce  R.  Wheaton

and Robin E. Rider, editors. Inventory of sources for history of Twentieth-Century physics: report and microfiche index to 700,000 letters. Verlag für Geschichte der Naturwissenschaften und der Technik, Stuttgart, Germany, 1992. ISBN 3-928186-09-4. 62 microfiches (16,035 frames) pp. LCCN QC15 .I68 1992.

Wright:1987:JLM

[Wri87] Helen Wright. James Lick’s monument: the saga of Captain Richard Floyd and the building of the Lick Observatory. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-32105-0. xvi + 231 pp. LCCN QB82.U62 M689

                                        1987. URL http://www.loc.
                                        gov/catdir/description/
                                        cam031/86004179.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam031/
                                        86004179.html.

Wright:1994:BRH

[Wri94a] Nigel Wright. Book review: Howard E. McCurdy, Inside NASA: High Technology and Organizational Change in the U.S. Space Program. Baltimore and London: Johns Hopkins University Press, 1993. Pp. xiv + 215. ISBN 0-8018-4452-5. £27.50. British Journal for the History of Science, 27(4): 483–484, December 1994. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027636.

Wright:1994:MPD

[Wri94b] Susan Wright. Molecular politics: developing American and British regulatory policy for genetic engineering, 1972–1982. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-91065-2 (hardcover), 0-226-91066-0 (paperback). xxii + 591 pp. LCCN QH442 .W75

1994. URL http://www.loc.
gov/catdir/description/
uchi051/93047054.html;
http://www.loc.gov/
catdir/toc/uchi051/
93047054.html.

Wright:1995:RU

[Wri95] Nigel Wright. Review: [untitled]. British Journal for the History of Science, 28(4): 490–491, December 1995. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027571.

Wright:1996:JFS

[Wri96a] David Wright. John Fryer and the Shanghai Polytechnic: making space for science in nineteenthcentury China. British Journal for the History of Science, 29(1):1–16, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027509.

Wright:1996:BRJ

[Wri96b] Nigel Wright. Book review: John Krige and Arturo Russo, Europe in Space, 1960–1973. Noordwijk: European Space Agency Publications Division, 1994. Pp. viii + 142, illus. ISBN 92-9092-125-0. Dfl. 70.00. British Journal for the History of Science, 29(2): 252–253, June 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027852.

Wallace:1991:ARW

[WS91a] Alfred Russel Wallace and Charles H. (Charles Hyde) Smith. Alfred Russel Wallace: an anthology of his shorter writings. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991. ISBN 0-19-857725-7. x + 551 pp. LCCN Q158.5 .W3525 1991. US$62.50 (U.S.). URL http:

//catalog.hathitrust.org/
api/volumes/oclc/
22422509.html.

Wallace:1991:AHS

[WS91b] Alfred Russel Wallace and Charles H. (Charles Hyde) Smith. An anthology of his shorter writings. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1991.

ISBN 0-19-857725-7. x + 551 pp. LCCN ????

Wright:1982:PMK

[WT82] Peter Wright and Andrew Treacher, editors. The problem of medical knowledge: examining the social construction of medicine. Edinburgh University Press, Edinburgh, Scotland, 1982. ISBN 0-85224-417-7 (paperback). 232 pp. LCCN RA418 .P77 1982.

Wussing:1984:GAG

[Wus84] Hans Wussing. The genesis of the abstract group concept: a contribution to the history of the origin of abstract group theory. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-23109-3. 331 pp. LCCN QA171 .W8713 1984.

Wallis:1975:IBM

[WW75] Peter John Wallis and R. V.

Wallis. An index of British mathematicians: a check-list. Issued by the Project for Historical Biobibliography at the University of Newcastle upon Tyne, Newcastle upon Tyne, UK, 1975. ISBN 0-7017-0000-9 (paperback: vol. 2), 1-871768-01-2 (pt. 3). LCCN QA30 .W35 1976.

Wallis:1986:BBM

[WW86] Ruth V. Wallis and Peter John Wallis. Biobibliography of British mathematics and its applications. Part II, 1701–1760. Epsilon Press [for the] Project for Historical Biobibliography, Letchworth, UK, 1986. ISBN 1-85027-003-1. xxiii + 502 pp. LCCN Z6651.

Wallis:1988:ECM

[WW88a] Peter John Wallis and R. V. Wallis. Eighteenth Century medics: subscriptions, licences, apprenticeships. Project for Historical Biobibliography, Newcastle upon Tyne, UK, second edition, 1988. ISBN 1-871768-00-4. xxvi + 690 + xix pp. LCCN R713.29 .W34 1988.

Wodeham:1988:TIC

[WW88b] Adam Wodeham and Rega Wood. Tractatus de indivisibilibus: a critical edition, volume 31 of Synthese historical library. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2424-5. vii + 333 pp. LCCN B105.I53 W63 1988.

Wallis:1993:IBM

[WW93] Ruth V. Wallis and Peter John Wallis. Index of British Mathematicians Part 3, 1701–1800. PHIBB, Project for Historical Biobibliography, Newcastle upon Tyne,UK, 1993. ISBN 1-871768-01-2. xiv + 175 pp. LCCN ????

Wylie:1989:YCP

[Wyl89] Ian Wylie. Young Coleridge and the philosophers of nature. Oxford English monographs. Clarendon Press, New York, NY, USA, 1989. ISBN 0-19-812983-1. x + 176 pp. LCCN PR4487.P6 W9 1989. US$33.75. URL http://www.

loc.gov/catdir/
enhancements/fy0636/
88021084-d.html.

Wynne:1982:RRW

[Wyn82]  Brian Wynne.  Rationality

and ritual: the Windscale Inquiry and nuclear decisions in Britain, volume 3 of BSHS monographs. British Society for the History of Science, Chalfont St. Giles, Bucks, England, 1982. ISBN 0-906450-02-0 (paperback). x + 222 pp. LCCN HD9698.G72 W96x 1982.

Yates:1983:RRI

[Yat83] Frances Amelia Yates. Renaissance and reform: the Italian contribution, volume 2 of Collected essays. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1983. ISBN 0-7100-9530-9. xii + 273 + 16 pp. LCCN DG445 .Y37 1983.

Yavetz:1992:BRK

[Yav92] Ido Yavetz. Book review:

Kathryn M. Olesko. Physics as a Calling: Discipline and Practice in the Koenigsberg Seminar for Physics. Ithaca and London: Cornell University Press, 1991. Pp. xviii + 488. ISBN 0-8014-2248-5. $39.95. British Journal for the History of Science, 25(4): 478–479, December 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027067.

Young-Eisendrath:1997:CCJ

[YED97] Polly Young-Eisendrath and Terence Dawson, editors. The Cambridge companion to Jung. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-47309-8 (hardcover), 0-521-47889-8 (paperback). xxxiv + 332 pp. LCCN BF109.J8 C36

                                        1996. URL http://www.loc.
                                        gov/catdir/description/
                                        cam027/96029101.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam023/
                                        96029101.html.

Yeo:1992:BRM

[Yeo92]  Richard  Yeo.

Book review: Menachem Fisch. William Whewell, Philosopher of Science. Oxford: Clarendon Press, 1991. ISBN 0-19-824240-9. £30.00. British Journal for the History of Science, 25(3): 368–369, September 1992. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027267.

Yeo:1993:DSW

[Yeo93] Richard R. Yeo. Defining science: William Whewell, natural knowledge, and public debate in early Victorian Britain. Ideas in context. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-43182-4 (hardcover). xiii + 280 pp. LCCN Q127.G4 Y46

1993. URL http://www.loc.
gov/catdir/description/
cam025/92030620.html;
http://www.loc.gov/
catdir/toc/cam028/
92030620.html.

Yeo:1996:BRJ

[Yeo96a] Richard Yeo. Book review:

Jean Le Rond D’alembert, Preliminary Discourse to the Encyclopedia of Diderot, translated by Richard N. Schwab. Chicago: University of Chicago Press, 1995. Pp. 224. ISBN 0-226-13476-8. £8.75. British Journal for the History of Science, 29(3):366–367, September 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4027744.

Yeo:1996:BRR

[Yeo96b] Richard Yeo. Book review:

Robert Chambers, Vestiges of the Natural History of Creation and other Evolutionary Writings. Edited with a new Introduction by James A. Secord. Chicago: University of Chicago Press, 1994. Pp. xlviii + 390 + 254. ISBN 0-226-10073-1. £15.95, $22.95. British Journal for the History of Science, 29 (1):104–105, March 1996. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.
jstor.org/stable/4027526.

Yoshikawa:1994:SNN

[YK94] Hideo Yoshikawa and Joanne Kauffman. Science has no national borders: Harry C. Kelly and the reconstruction of science and technology in postwar Japan. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-24037-8. xvii + 137 + 9 pp. LCCN Q127.J3 Y677 1994.

Yoder:1990:RU

[Yod90] Joella G. Yoder. Review: [untitled]. British Journal for the History of Science, 23(4): 473–474, December 1990. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/4027209.

York:1989:AOT

[Yor89] Herbert F. (Herbert Frank) York. The advisors: Oppenheimer, Teller, and the superbomb. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1713-3, 0-8047-1714-1 (paperback). xiv + 201 pp. LCCN UG1282.A8 Y67

                                        1989. URL http://www.loc.
                                        gov/catdir/description/
                                        cam023/88062671.html;
                                        http://www.loc.gov/
                                        catdir/toc/cam027/
                                        88062671.html.

Yost:1999:BRP

[Yos99] Robinson M. Yost. Book review: P. M. Harman, The Natural Philosophy of James Clerk Maxwell. Cambridge: Cambridge University Press, 1998. Pp. xiv + 232. ISBN 0-521-56102-7. £35.00, $59.95.

British Journal for the History of Science, 32(3):363–378, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/
stable/4028003.

Young:1985:DMN

[You85] Bob (Robert M.) Young. Darwin’s metaphor: nature’s place in Victorian culture. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-30083-5, 0-521-31742-8 (paperback). xvii + 341 pp. LCCN QH361 .Y68

1985. URL http://www.loc.
gov/catdir/description/
cam031/84021377.html.

Yoxen:1983:GBW

[Yox83] Edward Yoxen. The gene business who should control biotechnology? Crucible. Science in society; A Channel Four book. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1983. ISBN 0-330-28112-7 (paperback). 264 pp. LCCN TP248.3.

Qian:1985:GIS

[yQ85] Wen yuan Qian. The great inertia: scientific stagnation in traditional China. Croom Helm, London, UK, 1985. ISBN 0-7099-2104-7. xii + 155 pp. LCCN Q127.C5 Q24 1985. US$27.00.

Zambelli:1992:SAE

[Zam92] Paola Zambelli. The Speculum astronomiae and its enigma: astrology, theology, and science in Albertus Magnus and his contemporaries, volume 135 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1380-1 (hardcover), 94-017-3467-4. ISSN 0068-0346. xvi + 352 pp. LCCN Q174 .B67 vol. 135. URL http://

www.loc.gov/catdir/
enhancements/fy0826/
91024566-d.html;   http:/
/www.loc.gov/catdir/
enhancements/fy0826/
91024566-t.html;   https:
//link.springer.com/book/
10.1007/978-94-017-3467-7.

Zandvoort:1986:MSD

[Zan86] Henk Zandvoort. Models of scientific development and the case of nuclear magnetic resonance, volume 184 of Synthese library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; and Tokyo, Japan, 1986. ISBN 90-277-2351-6. xiii + 305 pp. LCCN Q175 .Z3 1986.

Zott:1992:SGJ

[ZH92] Regine Zott and Emil Heuser, editors. Die streitbaren Gelehrten: Justus Liebig und die preussischen Universitäten: kommentierte Edition eines historisches Disputes. (German) [The contentious scholars: Justus Liebig and the Prussian universities: commentated edition of a historical dispute], volume 15 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. ERS-Verlag, Berlin, Germany, 1992. ISBN 3-928577-13-1. 289 pp. LCCN QD49.G3 S77 1992.

Zinsser:2006:EDC

[ZH06] Judith P. Zinsser and Julie Candler Hayes, editors. Emilie Du Châtelet: rewriting Enlightenment philosophy and science, volume 2006:01 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2006. ISBN 0-7294-0872-8. xi + 325 pp. LCCN PQ1981.D55 Z66